A Complete Bibliography of Publications in the Journal of Symbolic Computation

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

05 March 2024 Version 2.87

Title word cross-reference

#CSP [MM16]. #P [Bac99]. #P-Hard [Bac99].

(1 + i) [Wei00]. (1, 1) [FES11]. (C_{2}, 0) [Gen22]. (i, j) [MP04]. (k, 3) [HJ15]. (N, N) [Sha01]. (n^{2} − 1) [RW90]. (w, σ) [AP10]. (x − α)^{m} [BKSV20]. (x − β)^{n} [BKSV20]. (xy)z = y(zx) [HJM93, HJM94]. ∗ [Con93]. 0/0 [Chi08]. 1 [ES20, FFP98, Mos08]. 1000 [RDU03]. 112 [KLZA12]. 12 [BCE^{+}01]. 15 [KM00b]. 17 [Wil93]. 2 [AKS12, BdlCRLS19, Cre01, DS02, Eic10, GS12, HLM95, JPPSG09, Kos07, Sha01, SW97b, Wil90]. 23 ≤ n ≤ 26 [Ng89]. 2^{6} [Wur93]. 2A [Mat01b]. 3 [AOW23, BNT18, GKR23, JWC^{+}16, LN13, LMM05, Wil90]. 32 [Ano01f]. 3 × 3 [HKS21]. 5 [GSHPBS12, Har13]. 6 [Har13]. 64 [NV07]. 8 [HTZ04, dGPS09]. 81 [NV07]. 9 [Gaá00]. [P]: M^{∞} [Mor99]. ^{2} [Alc16, AOW23]. _{.}0 [CF09b]. _{n} [BP00]. A [FC04, BP00, Chi01, CC07]. A_{n}(k) [GV03]. AC1 [Con93]. arctan [Str12]. B [Hel96, Nak09, NN10]. β [Wir12]. BŜO_{n}⟨16⟩ [Ng89]. Z_{p} [Ree98]. C^{1} [GHC92]. C^{2} [JPNP23]. C^{n} [KKM15]. C^{r} [MP89]. C_{3,5} [OT13]. C_{n} [GLW99]. D [CJUE06]. EL [LW10]. L_{∞} [KS06]. Co_{3} [SW97b]. D [CJUE01, CvHKK18, CKLZ19, JPP19, NOT18, OTW00, OT01, Pan89, TW01, Ngu09]. D_{6}(p^{f}) [LMP19]. DD [JPP19]. ddet [MKF93]. Δ [FHR99]. δ^{+} [Wir12]. Δ_{1} [PS95a]. d ≤ 15 [Mal87]. E [DJ92, Bee01, Naw16]. E_{6}(p^{f}) [LMP19]. E_{8}(C) [GR01]. ϵ [OPP93]. exp [Str12]. F [PW85, HT17, Mur23]. F (x, y) = G(x, y) [Gaá93a, Gaá93b]. F_{2} [GR12]. F_{5} [BFS15, GHMA13, HA10, Kin14]. F_{p} [Vir93]. FP_{3} [CO94, CO96]. G [PW85, BFG07, BF11, LH18]. G/H [Hel96]. G^{1} [Far19]. G_{20} [Mar19b]. G_{21} [Mar19b]. G_{4} [BCCK20]. G_{5} [BCCK20]. G_{6} [BCCK20]. G_{7} [BCCK20]. G_{8} [BCCK20]. G_{a} [van93]. GU(3, q^{2}) [FF17]. GU(4, q^{2}) [FF17]. H [AP08]. h^{∗} [BEZ23]. D̂ [Nak09]. I_{1}(XY) [SST18]. J_{1} [CH96]. J_{3} [Wil90, Wil93]. K [Gün90, BE13, BMQS06, Hel00, JH21]. K(x) [KST93]. K(x, y) [HP91]. K[[x]] [AGSM17]. K[x] [AGSM17, GSSV12]. λ [SP10, Tak89]. ⟨T ⟩ [LMPX11]. ≤ 50 [BL96]. ≤ 7 [FMM07]. lim + [Wir12]. LLL [Van00]. log [Str12]. LU [Car15]. Ly [Gol01]. m [Koe95, RZAG99]. M_{11} [Kön17, MZM87]. C^{n} [Wal05]. CP^{3} [HS09]. F_{4} [HK10]. F_{p} [Sal08]. P^{2} [DJ07]. P^{3} [HMXD07, MHXD09, MPSXD09, BG05]. P^{n} [Dum09]. Q [AHS18]. Q(ζ_{7} + ζ̄_{7}) [HLSW16]. R^{3} [GV16]. R^{4} [FG16]. RHom_{D}(M, N) [Wal05]. Z[x] [JY17, Abb13]. Z_{q}[u]/⟨u^{2}⟩ [TS24]. D [Mon05, Wal05]. μ [CW03, CCL05, CR19, DJ96, Doh09, Eit94, HHK17, Kal11, MP89, ZS01]. N [KP91, BM01, BY23, HR19, BF91, BPH07, CIL07, EL12, FKM10, GHL16, LFD19, Nor95b, Nor95c, YXL99]. n ≤ 10^{6} [Kan91]. O [CLM16]. O(n^{3} + n log n log(log n) log q) [Pol95]. O^{+}(4, q) [FF17]. P [Guo20, hHL21, MS10, AIRR12, BCI13, BR87, Con90a, Dor21, EG21, LMP19, Led00a, Lim93, MF90, O’B90, O’B93, O’B94, SW02, Sla86, Win88]. p^{2} [BvzGZ13]. P^{4} [DS00, GEL05, DMS21]. P_{5} [Par22]. π [WZ23]. ΠΣ^{∗} [ABPS21]. p ≤ 31991 [Ree98]. PSL(2, 41) [GR01]. PSL(2, 49) [GR01]. Q [MZM87, Mau87, BR87, DV00, AU21, AR13, AGRZ99, BLM22, BK99a, BY23, CHM05, CX09, FKTS12, GG92, GHLZ21, Kra95, MZ05, Mu08, PP11, Rie03, SK12, WZ23, Zha03]. Q(\sqrt{2}) [Deu93]. Q(\sqrt{3}) [Deu93]. QD [Cuy97]. QPC_{2} [Tak93]. R [Sht88, dG01, HZ19]. R[[t]][x] [MR17]. R[x] [Sht88]. R^{n} [Ric99]. R_{2} [BELT21]. R_{3} [BELT21]. RΠΣ^{∗} [Sch17b]. r × 3 × 3 [BR06a]. S [BP00, EYZ21]. S(F_{4}) [GG92]. S(F_{4})^{∨} [GG92]. S^{2} [BELT21]. S_{k}(Γ_{1}(N)) [Wan94b]. S_{n} [Ber04, Cla21, Mau87]. S_{n−k} [Cla21]. sat(T) [LMPX11]. SL(2, C) [Chu99]. SL(n, K) [PS89]. SL_{2}(11) [Klü00]. Sp(4, q) [FF17]. T [PS95e, BPD19]. τ [KPT15]. θ [dG11]. U [MKF93, BR22]. Uq(sl(2)) [JJ01]. V [DMY16]. V [X]^{m} [DMY16]. ∥ [JKP98, PS95c]. x [GSSV12, MR17]. x^{p} [ES23]. x_{m} mod p(x) [Pan96]. Y′ = AY + B, Y′ = AY [Ber02]. y^{2} = x^{p} − x + d [LLL08]. Z [BM00, LMZ23, WyW93]. Z[i] [Wei00]. Z^{n} [BL12].

-action [MP89, van93]. -adic

[AIRR12, BR87, Dor21, MF90, Win88].

-adics [Lim93]. -algebras [BMQS06, LH18]. -Algorithm [Cuy97]. -analogue [CHM05, PP11]. -analogues [WZ23]. -ary [Wei00]. -bases [Doh09, HHK17]. -basis [CW03, CCL05, CR19, GSSV12, ZS01, Pan89]. -beam [KP91]. -binomial [Kra95]. -Calculus [Tak89, Kal11]. -canonical [AP10]. -cases [GG92]. -conjecture [KPT15]. -constant [MP89]. -cube [FKM10]. -D [YXL99]. -descent [Cre01]. -difference [AR13]. -Dimensional [Nor95c, AOW23, ES20, JWC^{+}16, Nor95b]. -discriminants [CC07]. -ellipse [JH21]. -Expressions [OPP93]. -extensions [Sch17b]. -field [ABPS21]. -finite [CvHKK18, CKLZ19, JPP19, JPNP23]. -Fold [Koe95]. -function [BPD19]. -functions [Eit94, NN10]. -gonal [BCI13]. -Gosper [BLM22, Mu08]. -graphs [BFG07, BF11]. -Group [O’B90]. -Groups [O’B93, O’B94, AOW23, Con90a, JPPSG09, Led00a, SW02, Sla86, dG11, BM01]. -Hahn [AGRZ99]. -holonomic [BY23, SK12].

-homogeneous [MR17]. -Hypergeometric [BK99a, Rie03, Zha03, Mat01b]. -integer [GHLZ21]. -invariant [Cla21]. -invariants [BR22]. -Laguerre-Hahn [FHR99]. -lattices [JY17]. -log-convexity [HZ19]. -matrices [HKS21, dG01]. -Modular [HLM95, SW97b, Wil90]. -module [PS89]. -modules [CJUE01, CJUE06, Mon05, NOT18, OTW00, OT01, TW01, Wal05]. -norm [KS06]. -orbits [Hel96]. -orthogonal [FKTS12]. -Partitioning [Gün90]. -periodics [GKR23]. -polynomials [BEZ23]. -power [SK12]. -problems [Con93]. -puzzle [RW90]. -radicals [EYZ21]. -Rational [WZ23]. -recursive [hHL21, MS10]. -residuals [Mur23]. -resultant [MKF93, FC04]. -resultants [Chi01]. -ring [EG21]. -rings [EG21]. -Roots [BF91]. -saturation [DMY16]. -schemes [Guo20]. -sequences [CLM16]. -series [AU21]. -set [LFD19]. -simple [BE13]. -solvability [Ngu09]. -step [BdlCRLS19]. -Steps [Wir12]. -subgroups [Hel00, LMP19]. -subnormality [Mur23]. -sums [Naw16]. -systems [AP08]. -th [BY23]. -theorem [PS95e]. -threshold [HT17]. -transformations [WyW93]. -types [EL12]. -unification [DJ92, Con93]. -uniformity [MP04]. -varieties [Hel00]. -view [BNT18]. -WZ [CX09]. -Zeilberger [MZ05].

1 [Kal90]. 16 [HJM94].

2 [Sti03]. 2-nomial [Li10]. 20 [Nor95b]. 2009 [JKP12]. 2011 [ES13]. 2014 [NSW16]. 2017 [EY21, MMO18]. 2019 [BDT22]. 2021 [CL23, DKM23]. 2023 [Ano23l, Ano23k, Ano23j]. 2024 [Ano23h, Ano24f, Ano24e, Ano24g, Ano24h].

3 [CLL17]. 33 [AP04, DHM11]. 37 [AK06, HZ15]. 39 [Hil05b].

41 [Ano06, Fer06a]. 42 [HP08, MHXD09]. 45 [FS13, KMR18, RS11b].

5 [Sag89]. 50 [HdC16]. 5th [CIM17].

60th [GW11]. 6R [HLSS15].

7 [KL90].

87b [Kal90].

9 [KT94]. 91c [KL90]. 91h [KT94]. 94i [HJM94]. 99 [Ano01d, AJ01].

abelian

[AH13, BJT22, Gau09, LR15, AH01, BJSS89, CDO01, GN04, Pie24, Wal02a, Wal02b]. Aberration [NSW85, WK91]. Abortion [SK91]. Absolute [BR87, Hon98a, Kal85, BS10, BCG10, CG06, EGW09, Kal90, KMR18, MR10, PS20, Rup04]. Abstract [BB93b, CP97, FLOR00, GSA^{+}12, Kal97b, MMO94, OPP93, BSC12, QHL^{+}13, SJ12]. AC [AK92, Dom92, Fer96, LM94b, MU04, NR97]. AC-termination [MU04]. AC-Theories [NR97, LM94b]. AC-Unification [AK92]. AC1 [Dom92]. academic [GK12a]. Accelerated [Web96]. Acceleration [BSC12, SJ12]. Accelerator [AB89]. accelero [vdH07b]. accelero-summation [vdH07b]. Acceptors [FB93]. accessibility [LLTPT^{+}11]. accessible [BPT11]. according [FS16]. Accurate [AT96]. Achievements [Gre00b]. aCM [CLM16]. across [GK18]. Acting [Hel00]. Action [Hv95, GES05, HK07, Hub09a, Kud22, MP89, van93]. Actions [BH00, Bay03, BJT22, Cla22, CGP23]. Activity [AGMT98]. Acyclic [Tha93]. adaptation [Sal08]. adaptive [BF20, JKKK20]. Adding [SO89]. Addition [HI94, Fuk04]. Additive [BLL^{+}16, HK10, vzGGZ21]. adhesive [BBC^{+}11]. adic [AIRR12, BR87, Dor21, MF90, Win88]. adics [Lim93]. Adjacency [ACM88, MC02, HDPS11]. Adjoint [Mňu97, SV92, DPS23, SS11, Vil11]. adjoints [CQ12, Wei13]. Advances [Ano00b, GR98, Yan99]. Adventures [LC89]. Advert [Ano03a]. Affine [CM93, HJS13, Wal00, BGMSG07, BJS04, BS21, FFP98, GMKP21, HJ18, NNPZN19, RDU03, SSS23, Sch07, Sch17a, Shp14, Tab11, VVY21, Wib07, van93]. Affinity [BGH93]. AG [MRG13, MRG17]. AI [PdRAEC24]. Aided [Wil95, DJ89, KD90]. aiding [Sil04]. Alexander [KP15]. Algebra [AGM97, Ano99c, AM88b, BC01, BG01, BGH93, BB92, Ber93, BK99a, Bos01, BK99b, CP97, Cap90, CO01, CL00, Cav86, CF91a, DGS96, Dav88, Dav02, Die92, Div91, Ebe01, Eis90, EC87, FS95, GV99, Grä95, Gre00b, Gri88, HSS02, HS95, HKL99, Hub00, Jac97, KS98, Kal97b, KKL92, Koe92, KG03, KR97, Lab95, LS02, MC97, NMM90, OT87, OZ94, Pro00, Rob97, Rd91, RST01, San95, Sch96, She92, Vei97, WS98, Web95, WBM99, YP91, AP90, AG91, BW05, BC06, BBF17, BDPR13, BCP97, BKG21, BKHG21, BR12, CS05a, CLS91, CM09, DHH^{+}04, DPS16, DT06, DW22, ES23, EPY98, FGPGP14, FS98, FK89, GH05b, GSZ85, Ger19, GTLN17, GSSV12, HV22, JMPR04, KL19, KASW05, LL13, Mas16, NSW16, NPP17, OdR03]. algebra [Oll88, Par08, Píl07, SM18, SMB03, SS88, Sti03, TM89, BD15, KK09, Tra07b, Yok17]. Algebraic [AF88, ACM88, Arn88b, AM88a, BGLGM17, BF91, BGK86, BDS17, Boy93c, Bro01a, BK90, BW87, BEM00, CFRS23, CDF92, Cha00, Chi96, CL23, CH91b, CJK02, Com98c, CR88, CS06, Cza89, DGLM^{+}24, Duv94, EY21, Enc95, For87b, GSS05, GPWZ02, GSST98, GKO09, GVGC99, Gre00b, HHK^{+}23, HPRS11, Hes02, HSW97, HKSS17, Hub99, IS10, JKP12, Kal93, Kal01b, KR94, KFF88, KLZ96, LM89, Laz92b, Lee08, MM00, McC88, MC02, Mil87, Mňu97, NT17, Nau98, Poh97, Ren92a, Ric92a, Rie93, Ris88, Roc22, RS90, SF90, SJA01, SGD97, Sed02, SME87, SW97a, Sen02, SS95, SU93b, Smi02, Str97, Str06, Stu91, TM89, Tra98, Ura24, YNT92, Zha90, van94, van97c, vdP99, AV11, AMT09, AS05, ASS07, AS07, Alc08a, Alc08b, Alc12, BGLHR12, Bas06]. algebraic [BR10, Bay03, BKV24, BS22a, BJS04, BP11, BC22, BDE^{+}16, BK15, BS09, BCGY12, CR98, CFMMP10, CDM^{+}13b, CDM^{+}13a, CM16, CJL13, DJO^{+}11, DM09, DGPP10, DM05, DJK05, DMR12, DDM15, DR23b, El 08, EGS23, ES13, EBD21, ES20, FGLH^{+}23, FLE^{+}23, FGVN06, Fit85, FGPT03, FGL04, FGP05, FGT09, FK09, GPS09, Giu88, HHS23, Har13, HS17a, Hil87, HS21b, Hor24, Hub19, IT10, KZ08, KS12b, KBRV24, KÁ21, KKM15, LPT20, LXZZ23, MSY21, MM16, MCMMPR14, MSW15, Mer10, Mor91, NÁS^{+}24, PdRAEC24, Pra13, PES24, Qi06, Rei06, Rio03, Sag88, Sag89, LT22, SS11, Sha13, SvE14, SvE21, Sor22, Ste05, Sti03, Sto17, Str16, SvH19, TN09, VGW18, VL10, Wal05, Wan18, XY02, Zen06, dG09]. algebraically [Ste10]. Algebras [BS90b, BR87, BH02, DR92, Drä01, Dün94, EG00, GK96a, HT95, Let01, Rón90, Ros93, VL93, Vel00, dG01, vdP99, AL88, BdlCRLS19, BMQS06, BDM17, CGGO09, CGS97, CS98, DFdG13, Dra03, EG04, EKP22, Eic10, GIL88, HL18, IK13, IL09, JMV23, KP13, KRW90, Kut19, LPT20, Lab90, Lea06, LZS11, LH18, LW03a, LW03b, Li10, LG21, Mad14, MS09, MM04b, NOT18, OdR09, Orl22, PRR18, QR07, RR05, Roo13, RR08, Sch19, Shi04, Wur93, ZGG23, dG09, dGPS09]. algebroid [MSS21]. Algorithm [AB22, AS01, ACM88, Arn88b, AM88a, Bah01, BO10, Baj86, BP99a, BL98a, BEM97, BL85, BC93, CM96, CGG89, CS90, CD00, Col02, CF94, CKS99, Cuy97, Cza89, DR00, Die92, For02, Gaá93a, GM88, HNVL90, Hem02, Hen98, HHK17, Jeb93, Jeb95, Kal93, KT02, Kem99, KM99, KM01, Kov86, Kri85, LS00a, Lim93, LPS93, Lo98b, LO99, MS95, MM97, Mau00, MW91, Mic88, Mon02b, Mul90, MF96, MO95, Nak09, Nie94a, Nor90, O’B90, OTW00, OPP93, PS95b, Pic98, Ple87, Poh87a, PW94, Rol86, Ros93, Sch90b, Sch85, Sed02, SL92, SS94, SS98b, Sho95, Sim90b, Sit92, Sny93, Ste97, Sto99, Tak92, Tak95, Tho02, Tra00, Tra96, UW96, Van00, VL93, Web96, Wei00, XY02, ZSY93, Zha95, van94, AV11, AGR95, AL88, BV03]. algorithm [BFS15, Bay03, BSSY18, Bel04, BFPT21, BGMSG07, BvdE03, BLV16, BLV18, BMQS06, BLM22, BLS23, BC24b, Buc06a, BK12b, But88, CL17, CHM05, CHM12, CvH04, CK12c, DA05, DHM11, DH17, Dum09, DLLP08a, Ebe19, EP10, ELME23, FDS13, FP09, FG06, FSW10a, FW15, GLLdR19, GHMA13, GS03, GKMO08, GOT05, HBN95, Har12a, HJS16, HM23, HJA17, HTX15, HM21, IMP17, JY17, JV09, KSW13a, KS12a, Kin14, KS86, Lab90, Lec19, LS12, LH18, LR15, MM06, MRG17, MO21, MS11b, MPT20, Min98, Moe05, MD24, Mu08, MH06, NT21, Nag21b, NWW21, OdR03, OdR09, PDS08, PS89, Pop15, Qi23, Ree98, Ren04, RT17, Rue11, RSS13, ST89a, Sch04, ST20, Smo21, TM85, TV18, Tsu09, Ung19, Ura24, Vil11, Wan06, WWWX23, WS09, Wur93, Yes21, YZ21, Zha03]. algorithm [Zhe21, de 98, van93, CCD^{+}09, Rou09]. Algorithmic [BGLHR12, BENW06, CDO01, Kal98, Ley01, Ley04, LH17, Mar02, MMY00, Pas86, PRR18, RS00, SK12, Wal00, Wal05, CS06, DT23, GS07b, HY23b, JLW13, KMY24, LGM21, Rad15, Sch17b]. Algorithmical [FGL04]. algorithmically [BM04]. Algorithmics [JKP12, ES13]. Algorithms [AGS16, AT96, BZ85, BP99b, BKRW17, BTW93, BM01, BK99a, BDS17, Bou93, CM10, CM12, CDO97, DG14, DF08, ER95, ES98, EC95, FES11, FGPT03, GvPS00, GKsL03, GSHPBS12, Grä95, HJS18, HH13, HM05, Hel16, HLS01a, HS98, HI94, Hub00, IPS11, KLM^{+}21, KN11, KL98b, Koe95, Koh08, Kop08, KL89, Leo91, LRW97, Mv90, MO88, MR98, MES19, MC93, MC02, MNJ94, Mon97, MR15, Mro96, MQS99, Naw16, Nie94b, Nip91, Oak13, Ost99, Pan02, Pic00, RS21b, SS92, Sch01, Sha90b, SH17a, Tun09, WBM99, Wor94, ACFP12, AAB^{+}18, AGS18a, Arn03, BS18, BP09a, BS17a, BdlCRLS19, BF20, BF22, BCG10, BDL^{+}13, BCGY12, CS22, CCG06, CCD^{+}09, CL07, DF05, DMW17, DJ15, DJS18, DPS23, EF17, EH16, EMSS16, FM17, Gal87, GH12, GMKP21, Gen07]. algorithms [GSPB17, GSSV12, GOP18, HL04, HNE21, HdC13, HdC16, HG20, HNRS21, JWW23, JB04, KsL03, Kau07, Ker17, KMR18, KL90, KL17b, Lec07, LY05, MO85, Mal21, MR10, MS21, MZ05, MTV21, OO13, OS04a, Rob04, Ros23, RdC13, RS13, Sht88, Vac17, Vac18, VVY21, Ye18, ZL12, vdHS06, vdH07d, vdP05]. Aligning [GK18]. alkane [LMM05]. Almost [Fas10, Wei88, Wei90, BLV18, BK12b, Li10]. along [Gal16, NT21, NNPZN19]. Alternant [BF01]. alternating [Val11, WO06]. Alternative [BH02, Gar95, Mar96a, JML^{+}13, SS03b]. always [BLV18]. ambient [GTLN17]. Amenability [DMW17]. amoebae [BHH23]. among [HR19, Mor13, Ye17]. amortization [Bur16]. Analogical [YX95]. Analogs [Mil96]. Analogue [Wei00, AGS18b, CHM05, PP11]. analogues [WZ23]. analyse [JSC13]. analyses [BLV16]. Analysing [DS96, PS97a]. Analysis [ABP96, CL00, CM96, Cra91, Eis90, GV99, GKLM91, KC01, KG03, Mv90, Mag89, Mro96, Nor90, RT85, SS94, WKB86, APS12, BL06a, BSC12, BR12, CCG06, CCD^{+}09, Cip08, CTY10, DE06, Ede13, FK09, HS06, HJS16, HJS22, HTX15, JT03, LS16a, Lic21, MdCW17, MV13, Oll88, PT14, RH18, Roq13, SJ12, Str19, Wan86, XL13]. Analytic [Eck87, Ful90, GLLdR21, HH09, McC97, OT87, Whi91b, CMV13, GGG06, HH13, LLZ24, Lem03, vdH05]. Analytical [Mer01, VV97, PNM13]. analytical-experimental [PNM13]. Analytically [DH00]. analyze [GES05]. Ancilla [STDD16]. Ancilla-free [STDD16]. ANF [HK21]. Angle [AI90, WW94]. anisotropic [KR23]. Annihilating [TN09, GVHHUE05]. Annihilators [KZ14]. Annotated [Frü96]. Annotations [ACGR01]. answers [KSD16]. Anti [CKKM10]. Anti-patterns [CKKM10]. any [DW18]. apolarity [Sta23]. Apparent [CKLZ19, BJ21]. appendix [Sza08]. Applicability [CHM05]. Application [Ape98, Baj86, BF01, CD87, CD85, Cow92, DR00, DT95, DTGV01, Eis90, ES18, EC87, GV99, HS89, JKP98, KC01, Mer10, Mer01, Miy01, MR02, Pal13, Pan96, PZ96, She92, She97b, Tri86, UYSA89, VGT90, Vor89, WKA94, YP91, ZBH96, AAKM21, AHKY09, AMW12, BGL14, BCR15, CCD^{+}09, Eit94, FK89, GSZ85, HJ15, Kin14, KH23, KS86, LMR94, MBC^{+}10, MSW15, MS03b, MKF93, NOF10, Naw16, PS95c, Sch17a, STW18, WWWX23, Wur93]. Applications [Ano02a, BB00, BF91, Bro01b, CH97a, Cha00, CRAB91, CS90, DR86, Gat03, HSW97, HL97, JKP12, Key01, KM01, Li04, MC97, MR87, NSW85, Sch94, Tra00, Wan94a, AU21, AAB^{+}18, BP24, BCE11, BKW20, BBCM13, BY23, BBN18, BF11, BW03, Bur16, BG05, CFMMP10, CS22, CES23, CM09, DHH^{+}04, EK19, ES13, FP09, FRR06, GGAVRC13, GGMFVT13, HDHX17, HGKV11, KASW05, LH18, LH98, LLL19, LR15, MV10, PB07, PR12, PWZ18, SA89, ST19, Win14, GTLN16, GTLN19, Tra07b]. Applied [Dav88, MQS00, Rol86, AP90, Bar13, Par08]. Applying [BKG21, GV96, SJ12].

Approach

[AK92, Ape95, BT98, CK99, Du 99, For87b, FKM95, Ful90, HY96, Ley01, MM97, Mňu97, OZ94, RS00, Sau01, Sch93, Sod96, Tak92, VV97, Wer98, WG94, Wor94, YNT94, ZS01, BSW21a, BKW20, BPH07, BKSS12, BJT22, Brå24, BSW21b, CR98, Cla22, CS06, DIdW18, DT23, El 03, FGVN06, GS07b, Guo20, HY23b, JMV18, Kho08, KPT15, KZ10, KMY24, KT23, MPH17, PV13, Rad15, Sch03a, SS03b, Win88]. Approaches [MPS02]. approximant [JNV21]. Approximate [EGB12, HKPP09, KMYZ08, KL98a, Nag11, Nag21a, Tun02, vzGMS10, AV11, AS23, CG06, Der13, Lia13a, Lia13b, MRSW07, MSW15, Nag21b, Nak09, Sag14]. approximately [RSS13]. Approximating [For02, Hon04]. Approximation [Far97, FF92, Mil87, BS22a, BC05, CJL13, LOOR^{+}03]. Approximations [BX97, GR10, GHL21a, KLR93, RS21b]. April [Ano24g, Ano23l]. Aquarius [BH95]. arbitrarily [DO06]. Arbitrary [FGT02, Kem96, SS89a, Tra98, Bîl11, Bur04, FS10, FS13, Har17, HJ18, MRH23]. arbitrary-precision [MRH23]. arc [Far19]. Architecture [EW00]. Architectures [CM96]. areas [Tsa16]. Arising [GH02, Deu93, Ye17]. Aristotelian [Moz89]. Arithmetic [CJK02, CW90, von87, Abb12, BPZ06, BGLGM17, CV11, CH17, EPY98, GJT13, Har14, HLO23, LMS09, LMRS11, Nal18, OT13, Sag14, Smo21, ZWM15]. arithmetically [DH16]. arithmeticity [DFdG15]. arithmetics [DS12, HL21a, HL21b]. arithmetization [BBN18]. arrangement [HDPS11]. arrangements [BDPR13, SS24]. Array [Sak88]. Arrays [CM93, MG88, Tor93, Joh15, LH98]. Artificial [FL11]. Artin [AS24, AK00, Bok08, DS12, ME21, Sut13, Sut16]. Artinian [KZ14]. ary [Wei00]. Askey [FKT13]. aspect [HKSS17]. Aspects [CM93, Ris88, MCMMPR14, Mor91, Poz15, Sch17b]. assertions [JML^{+}13]. assess [PNM13]. assessment [GGdR^{+}13]. assignment [AP10, Cla22, JKKK20]. assistant [GK18]. Assistants [BC01]. assisted [FM02]. Associated [Bah01, BH87, DiP16, FHR99, AH13, ASS20, BGLGM17, Jam11, Joh15, MS03a, NN10, PV05]. Association [Miy01, KLZA12]. Associative [BP85, EG00, Fag87, JM93, LC89, Pau92b, CM17a, CdG09, DKM21, GIL88, Ger06, LL13, Mor20, Raj06, Sch19, Wid01]. Associative-Commutative [Fag87, LC89]. Associativity [For87b]. Associator [BH02]. assume [AB05]. astro [DJ89]. astro-geophysics [DJ89]. astroid [DP19]. Astronautics [WKA94]. Asymmetric [Lab92]. asymptotes [LPR17]. Asymptotic [BEM97, Die92, HZ19, KP91, Nor90, PY05, Sal94, San96, SS95, DET09, vdH09]. Asymptotics [SS98a, SS99, MS21]. asynchronous [KZ10]. Atiyah [MS22]. atom [SSSK18]. atom-variables [SSSK18]. Atomic [Pic00]. attacks [LT22]. attributed [Ore11]. augmentation [Mu08]. August [Ano24f, Ano23i]. Author [Ano01b, Ano01f, Ano02c]. auto [Sto17]. Autocovariance [AP22]. Automata [CH91a, DCC95, KFK97, RW94, BCR15, DJK05, GR10, LM94b, RV05]. AUTOMATE [CH91a]. Automated [BBK14, Bib85, BS00, Bou97, BK99b, CL20, CCM95, Col05, FOT00, LW03a, LW03b, LS02, MR87, Zha90, BKR19, CS05a, DP19, GSSST10, IKGT11, KKK^{+}16, KS86, Liu19, Win06]. Automatic [BBB92, Bee01, BB93a, Bie85, BD04, EHR91, HTZ04, HH99, JB04, LJ09, SS98b, Sod96, WyW93, GGL06, HV16, Wan86]. Automating [Ebe19, KNZ91]. automation [CGO88, ZWM15]. Automorphism [CH03, Dic92, FFP98, Wil90].

Automorphisms

[Hul99, AP10, ABMN10, BCI13, Bor22]. Autonomous [Sch85, CFS22, CFRS23, FG06, NW10, NW11]. autotopisms [FMM07]. Avalanches [Sav90]. Average [CR90, AP22]. averaging [HY23b]. avoid [NPD09]. avoiding [BBCF22, BP23, KMY24]. aware [Ran12]. AXE [LMR94]. axes [BFMS87]. AXIOM [BT94]. Axiomatic [Sch93]. Axioms [Lai24].

Baby [BS18]. Baby-step [BS18]. background [SA89]. Bäcklund [FK89, WS09]. Backtrack [BL85, LT89]. Backtracking [PW94, Bec03, JWW23]. Bad [Kal01b, Nau98]. Bailey [BLM22, PP11]. balancing [GMS09]. Ballot [Ges92]. balls [BR10]. Bar [Cer18]. Barcelona [DGPP10]. Barnett [DTGV02]. Baruah [DT23]. Base [AS01, CF94, CGZ00, LMP89, MO95, AHW05, ACMB19, AR06]. Based [AGM97, ABP96, Arn88b, BG01, BB93b, CGG89, DS96, Dün94, HS95, Leo91, LHD96, Pau92b, Soc91, WG94, YI94, You89, Zha94, dB89, ASS07, BF95, BSSY18, BF22, Bro12, Bur16, CM16, CvHKK18, CJ15, CKKM10, Cla22, CLS91, DKLP21, Ede13, EF17, ELME23, EP04, FMR04, HJA17, Hua23, HC12, KZ10, KMY24, MMW11, Mal21, MRG17, MdCW17, Moz89, OB03, PDS08, Qi23, Rue11, Sch07, SWF11, SS03a, Vis05, Ye18]. Bases [AF00, ABL93, AHLM99, Ape95, Arn95, AGSM17, BCE^{+}94, Bec90, Bec93, BGK86, BTBQM00, BF01, CRAB91, CKM97, Cow92, Czi95, FGLM93, GV03, Gar95, GG99, Göb95, Gre00a, GS98, HT95, HSS02, HKL99, JL91, JM95, Kal97a, Kal99, Kal01a, Laz85, Lev00, MN02, Mil96, Mon02b, MQS00, Nor02, PZ96, Ros93, Rut92, STA94, Sei02, Smi02, Sne98, Wei92, dG01, vG90, AFT08, AB92, ÁAF^{+}18, ATY08, AT08, AK86, Arn03, AKR11, BM88, BBF17, BE22, BCR11, BR22, BR06a, BL12, BL17, BDLP22, BV06, BR88, BDM^{+}16, BDM17, BP09b, CJUE06, CMR15, CdG09, CR11, Cip08, DJ05, Dah22, DHM11, DO06, DE06, Doh09, Dön13, DW21, DL06, EP10, EF17, EH21, EPP21, ES20, FMM07, FL11, FES11, FEV16, Fer88, FFP98, FD14, FD18, FK04]. bases [GS24, GH05a, Gat03, GHMA13, Ger06, GTZ88, GKsL03, Göb98, GMP22, GMP13, GSZ13, Hal13, HPS22, HM23, HOS23, HH07, HM09, HP91, HV22, HHK17, IvH17, IL09, JNSV17, JNV21, JGF09, KRW90, Kap86, KSW13b, Kha14, LL09, LL13, Lai24, LLM^{+}13, LS04, LO09, LS11, LS12, LMZ23, Lev07b, LG21, Lia13a, Lia13b, Lia22, LH98, LLL19, Mad14, Mar08, MRW17, MM04b, MRG17, Mau87, MR13, Möl88, MW10, MS03b, NT17, Pau07, PPR20, Pol95, QR07, Raa12, Raj06, Rap06, RS16, Rei06, RR05, Rou08, Sak21, SIS^{+}11, Sch17a, Sch05, SS88, Sta18, Ste13, SS03b, Szi17, Vac18, WO06, Wal03, Wei03, Wei06, Wib07, Win88, ZGG23, ZW08, WRI09]. Basic [Buc87, MQS99, NRS89, Kra95, Naw16]. Basis [FT95, FF92, FD14, GHC92, Hon98b, HS00, Hre94, JL91, KM99, KM01, MR98, MM00, ŌS94, Pan89, Pau92a, Tay02, Tra00, van94, AFdCS15, ACFP12, AH05, BFS15, Bok08, BD09, BM16b, Buc06a, CW03, CCL05, CR19, DS09, Ede13, FMTT13, Gon17, GSW11, GSSV12, JPP23, KRK88, Kho08, LO08, Li10, LOOR^{+}03, Mal21, MAN^{+}10, MD24, SS16, TUÖ05, Tsa16, Val11, ZS01, ZL12]. Baumslag [Sim90b]. Baumslag-Cannonito-Miller [Sim90b]. Baxter [GIM07]. Bayesian [GSS05]. be [KMN88, MS09, vdH02]. beam [KP91]. behavior [Alc08a, Alc08b]. behavioral [NOF10]. Benchmarks [AK92, FOT00]. bending [Loj13, Roq13]. Bendix [EHR91, KMN88, Sim91]. Bergmann [TM85]. Berlekamp [BF20, CK12c, Gen07].

Berlekamp/Massey [CK12c]. Bernoulli [BBV15, FW15, KP91]. Bernstein [Bah01, BO10, JJ01, Ley01, NOT18, UCJ04]. Bertini [Ang18]. best [Nag21b]. Better [KSD16, LXZZ23]. Betti [AC19, Bas06, dCW09, SSS23, TV18, dAM17]. Between [ASJ97, MR02, Soc91, TW01, AH13, ACFP12, Coo09, HKL24, Hir89, KLZA12, KLV10, SPZ10, WW94, aZGS05, Zhe21]. Beyond [Dav02, LY18, ZWM15, HNR24, ST24b]. Bezout [CZG02, DTGV02, AL10, BU99, HHPS21]. Bezout-like [DTGV02]. Bezoutian [Mou05]. BGG [Pos18]. Bi [LA96, AHS21]. bi-homogeneous [AHS21]. Bi-rewrite [LA96]. Bibliography [Ano87, Arn88a, BA85, CH85, CH86]. bicubic [GS07a]. bidegree [FES11]. Bidirectional [KJ96]. bifiltered [Fer06a, Fer06b]. Bifurcation [GKLM91, LP90, Mag89, GH05b]. Bifurcations [EW00, GES05]. Big [Lim93]. bigraded [BBC^{+}17]. bihomogeneous [FES11, SJG13]. bikei [CN19]. bilinear [Hul13]. Billiard [GKR23]. bimodules [RR05]. Binary [Dür89, FKM10, Lip93, SS94, Wei00, ZSY93, ADCZ21, ABF09, Bed07, Bed09, BFPT21, BP10a, BP10b, CF09a, CTY10, DL88, DJS18, Sal08, STDD16]. Binding [BGH93]. Binomial [AP93, CTV16, GHY17, KM99, KM01, MS00b, PS95b, Tak95, BL12, BLS17, CM17b, Kra95, KW24, dAM17]. binomials [CTV16]. biological [BDE^{+}21, HTX15]. biquandles [CN07]. Birational [El 05, BBC^{+}17, SS88, Sta16]. Birkhoff [CIL07, GV96]. Birth [KKK17]. birthday [GW11]. bisection [Bur16, Col17]. bisection-based [Bur16]. Bit [ES18, EGS23, DP24, EMT21, vdHL13]. bit-complexity [vdHL13]. bitangents [GP24, LM20, PSV11]. Bitfield [Ric92b]. Bits [BS90a]. Bivariate [DE02, Ger19, LPR17, Lev21, McC97, Sch99, YNT94, vW95, Ave09, BLPR15, BMT21, CMV13, CWZ23, CK04a, Dah22, DET09, Fer06a, Fer06b, HW24, Khe03, LWXZ23, Sal08, Wei13, ZX20]. Black [BP00, KT90a]. Blackboard [DJ96]. blending [PDS03]. Block [HNRS21, Tho02, HJS22, LT22]. Block-Krylov [HNRS21]. Blockhandler [Ric92b]. Blocks [HLM95, Moe05]. blood [Sad17]. blowup [NNPZN19]. blowups [BS21]. BMM [BCCK20]. BMR [Mar19b]. Board [Ano18g, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano18a, Ano18b]. Board [Ano18c, Ano18d, Ano18e, Ano18f, Ano18h, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano20a, Ano20c, Ano20b, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21a, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano24a, Ano24b, Ano24c, Ano24d]. bodies [SPZ10]. body [CC24, Oll88, PY05]. Boer [Toh10]. Bombieri [Boy93a, Boy93b]. bonds [HLSS15]. Boolean [BJSS89, BD09, BD13, BS87, Eit94, HcL21, Kon95, MN89, SIS^{+}11, VB03, Zha94]. Border [BL17, BM16b, AFT08, LLM^{+}13]. Bordiga [BNT18]. Borel [BLR13].

Borwein [PP11, And95]. Bott [MS22]. Bottom [DS15]. Bottom-up [DS15]. Bound [Yap91, BMS17, BE17, Col15, Col16, Dum09, Eng10, GKMO08, GTLN19, KL21, KS12b, KMR18, MR10]. boundable [DP24]. boundaries [CDM^{+}13b]. Boundary [Cou00, Mil92b, BR12, Ros05, RR08, ST24b]. Boundary-Value [Mil92b]. Bounded [BS86, BP09b, Dur94, Gre16, Ric92a, Yam94, CHSS05, SSS05]. Bounded-degree [Gre16]. boundedness [Mic13]. Bounding [BR10, CR19, Cou18, SH17b, BW22]. Bounds [Abb13, BS10, BTW93, BDM23, Boy93a, Boy93b, DS00, GZ89, HH98, HM02a, Hon98a, Laz92a, Mig00, Wei94, Yam94, BDS17, Bus09, Col04, DJ07, EMT20, GMP22, GS18, GOP18, HS21a, HPS22, HH16, HHT18, KM06, Lia22, MR13, MS10, MZ05, Pap23, PS20, dCW09, WS23, YZ21, vdHS06, vzG13]. Box [BP00, MC02]. Boxes [KT90a]. bracket [LW03a, LW03b, NWW21]. Bradley [PP11]. braid [BV06, Bok08, PY94]. Braids [Bur01, Bur03]. branch [BKW20]. branches [HH07, HH09, HH13]. Branching [Dur94, SS96a]. Brauer [Bri06, Wil93]. breadth [LZ12]. bridge [KP15]. brief [GK12a]. Brieskorn [Sch04]. Bringing [LMRS11]. Broker [ABP96]. browser [KBRV24]. Bruce [NT21]. Bruhat [Dra05, DPS17]. Brun [BLV18]. Bruno [AKR11, Buc06a, HKP^{+}06]. Buchberger [HP08, AL88, AKR11, AT96, BMQS06, Buc06a, CM17a, Cza89, FD14, GM88, Hem02, HKP^{+}06, HP07, KS86, MM88, NP20, Ree98, Tak92, Tra96]. Buckling [RT85]. Budan [Gal13a]. Build [Str01]. Building [BT98, Pel97, Pel03a, Pel03b]. Built [NR97, Pet00, GN04]. Built-in [NR97, Pet00, GN04]. Bulk [CG02]. Bundles [LN13, BBV15, CP10, Lun16, Pos18].

C [RW94, BFK02, KZ08]. C-finite [KZ08]. cactus [BR13b]. CAD [Bro01b, HJX16, HDHX17, HLS01b, MH16, MPP19]. CAL [SA89]. Calabi [BR13c, Hie16]. Calculate [Kem99]. Calculating [Ber02, BGK86, Car01, Con90a, Gri90, IK13, Kem96]. Calculation [HI08, Kem02]. Calculations [BT94, CJMP97, Hre94, AL88, CGL07]. Calculemus [Ano01d, AJ01]. Calculemus-99 [Ano01d, AJ01]. Calculi [BF93, Fuc00a]. Calculus [BM00, CCM95, HSS98, MO98, Pau85, Pet00, ST89b, Tak89, BW05, BKSS12, DG14, DaZZ04, Gau09, GK21, HS17b, Kal11, LH17, Pel03a, SP10, Tak93, TM89]. calls [LCQ^{+}10]. Can [Mon97, Ebe19]. Canal [LSW01, PP97, DZ09, VL16]. Cancellative [Wal02a, Wal02b]. Cancellativity [NÓ89]. Cannonito [Sim90b]. Cannot [BCE^{+}94]. Canonical [Cha00, Dür89, Ste97, Wei03, van97c, AC24, AP10, JNV21, LM92, MM09, Sht88, Sto20]. Cantor [Eid23]. Capabilities [Pel97]. Capable [Sak88]. card [KT23]. Cardinality [LM94a]. career [GK12a]. Carlitz [Top14]. Carmichael [Arn95]. Cartan [DFdG13, DaZZ04, HT91, NPD09]. Cartesian [DDR11]. CAS [Kad13]. Case [BGH93, Bür89, EH16, HM02b, Hre94, Laz85, LZ12, Min02, Nip91, PH87, Sha90b, von90a, von90b, BR12, CIM17, DHKS09, DW22, HT17, KS12b, KS12c, Lev21, Min03, PdRAEC24, PS95c]. Cases [Ott91, GG92]. Castelnuovo [BGM06, HH04]. cat [AOW23]. Catalan [DV21]. Catalan-many [DV21]. Catalecticant [ABC22]. Catalogue [Le 86]. catalytic [Gon17]. Catastrophe [Cow92]. Categorical [Sto99, SLK11]. categorically [BGS11]. Categories [BH00, BBC^{+}11, Der13, DDR11, FGR03]. Cauchy [LR01, PR22, Sch17a]. Cause [SK91]. Cayley [BF11, CKM09, Deh94, LW03a, LW03b, Lun16, Whi91a]. CCC [Hir89]. Cell [Aur87, RS00, BK15, CLL17, Dra05, NÁS^{+}24]. cells [AH23, CRSW22]. Cellular [MS00b]. census [HR20]. Center [GV16, JLR03, MPH17]. center-focus [MPH17]. central [CC24, OdR03, Píl07]. centralisers [BNRW22]. centralizer [Poz24]. Centralizers [BC91, BR13a]. Certain [Köl85, Mau00, Wan99, ZD02, Abo10, AJGVS09, AIRR12, CJP22, Mau87]. certifiable [HS17b]. certificate [MWZ16]. certificates [DP24, DLMM11, DKLP21, GS24, GMW23, MW23]. Certification [DHS22, KLYZ12, Sto03]. Certified [AS23, CVY17, CJL13, CWZ23, HL16, MS04, Mul04, IMP17, MMS23]. Certifying [AHS18, HL17, JGF09]. chain [HP07, HP08, Nor15]. Chains [Rol90, BLPM19, CM16, HdC13, HdC16, SH17b, VK16]. Challenge [BCE^{+}94, DHK^{+}95, CTY10, KK09]. Challenges [Kal00, BB11]. chance [TZ21]. Change [CF94, FGLM93, Fit97, BLM10]. Changing [GGdR^{+}13]. Character [Con90b, Sch90b, Ber04, PS89, Sla86, Ung06]. Characteristic [Can90, Chi96, FGT02, GVYZ09, GH12, Kem02, Mat01a, NG93, Alu03, BLM10, Cox21a, Cox21b, DW21, Eic10, FF17, GLY09, Gol06, Gol08, GKO09, Hel16, HMXD07, HcL21, JLW13, Kos07, Mag17, MHXD09, Roo13, RdC13, Ste05, Sut15, Wal03, Wil93, vdPT15]. characteristics [BV03]. Characterization [AK04, AK06, Dön13, SPD14, FDS13]. Characterizations [Sch17b]. Characterizing [GS22]. Characters [AH86, BP98, Con90a, Gal22, LMP19, SW97b, Ung19, Wil90]. Chardin [Sza08]. Chebyshev [Boy93c, GKsL03, IK21, KPR10, KPRT18]. Check [LT89]. Checking [Ebe19, BGMSG07, BM04, DEG^{+}21, HGKV11, Kal11, MM10]. Chemical [GH02, Gon17, Lic21]. Chemically [Tri86]. Chern [BEP13, DEPS11, Hel16]. Chevalley [LMP19]. Chief [CH97b, Buc92]. Choice [Sch92]. Choosing [Hre94, LXZZ23]. Chordal [MBL21, LXZZ23]. Chow [FKO18, Qi23]. chromatic [Ger19]. Chunky [Roc11]. CIO [GRV17]. cipher [BGP09]. ciphers [LT22]. circuit [DIdW18, Pap23]. circuits [Bih15, MM16]. Circular [BFHS92, Hon97a, BFMS87, Hag89b]. circulations [Tsa23]. circumradius [Tsa16]. circumscription [Wer12]. claim [Col15]. Class [BdS01, BEM97, CP00, CE85, CDO97, Eck87, FHR99, GO90, Kno92, Kno93, Wan91, YNT92, dv96, tW91, AC04, Buc06a, FD18, Hel16, Hul13, Koh08, LWX23, TN09, Tra07a, XLY15]. Classes [MPS02, MMO94, Wid01, Alu03, BEP13, BNRW22, GU21, Har17, HV16, JPPSG09, MS16, SS03a, Vat12]. Classical [Cre01, Laz88, MC97, MMO94, Bro03, MRD11, Ye17]. Classification [BS21, FM02, GZ90, Sch98b, Zha93, BdlCRLS19, BMS20, Cou22, DP09, DLLP08b, FS12, GSZ85, HRdWY22, HH09, HH13, HY23a, LJ09, MS15, MS16, MV15, aZGS05]. classified [WK91]. Classifying [Deh94]. Clausal [Fuc00a, BL06a, Bec03]. Clause [BH95, Bon96, Boy92, PG86, Pel03b, Sid93]. Clause-Diffusion [BH95, Bon96]. Clauses [NR95, Gal87]. Clifford [HT95, LN13, TM89, Vel00]. Closed [Arn88a, Cha14, HLSS15, Lis95, Man93b, Man93a, GGMAMF19, Ste10]. Closure [KR89, Tsa00, Eli04, SSS23, de 98]. Closures [Hal01, Vas00, Sta18]. CLP [Sid93]. Cluster [Arn88b, Wan23]. Cluster-Based [Arn88b]. Clusters [HS97]. CNF [HK21, Pic03]. Co [LBM98, SJ12]. co- [SJ12]. Co-operative [LBM98]. cocyclic [AAFR09, ÁAF^{+}18]. Codatatypes [Hag89a]. Code [PH87, Cer18, MCMMPR14]. Codes [BF01, Key01, ABF09, BB10, BM10, BU14, BK13, BP09b, CCQ18, Cox19, DS09, FMTT13, FKM10, GTLN17, GTLN21, GS23, HK10, IS10, LO08, LO09, MPS16, MBC^{+}10, MRG13, MRG17, Rua09, Sop13, TS24]. codewords [MPS16]. Codimension [DS02, Mat01b, Tay02, AS24, AC24, AKS12, BGM06, RS16]. codimensions [BFK18]. Coding [BU09, CR88, CFMMP10, MS03b, PWZ18]. Coefficient [Bro92, Bro00, JL91, PS95b, Zha95, Col04, FK11, MS21, Wol00b]. Coefficients [Bea92, GK94, HH98, KST93, Kal02, MR02, Pet92, Sin91, Vir93, van97a, van97b, AB09, ABK15, AV11, ARST09, BVW18, BC24a, DT23, EK21, FGG^{+}16, FKT13, JPP23, KLYZ12, Pau07, Rup04, Ye17]. cofactor [HV22]. Cogeneric [MSY00]. Cohen [DH16, MS11a, ST20]. Coherence [Web95]. coherent [CLQ10, NP20]. Cohn [Pic98, Sau93]. Cohomology [Bac99, Car01, EMS00, EK21, Hol85, Mus00, Wal00, Wal03, ÀL06, DS18b, ES11, HM05, Hub09b, Kud22, NT17, Ng89, TN09]. Coisotropic [Koh21]. Collapse [Yel87]. Collapse-free [Yel87]. Collection [Geb02, LGS90, VL90]. collisions [BvzGZ13, Zie16]. column [BELP13]. Combination [FH86, SS89a, TRRK10]. Combinations [MT93, Mid94, Yel87]. Combinator [Sta89]. Combinatorial [BO99, BL85, Cer21, Dub93, ES23, FS95, Giu88, Lab95, NWW21, RS00, Rod15, STW18, CLM16, DLMM11, DJS18]. Combinatorics [PS95a, BC24b, Buc92, PZ92]. Combinators [Gib87]. combined [GS03, QHL^{+}13]. Combining [BS96, Bou93, CH95, CP00, Cza89, KR94, Nip91, Pel03a]. Combustion [GKLM91]. comes [BM04]. coming [CCQ18]. commandments [Sto11]. Comments [Buc06b]. Common [BFHS92, DTGV02, GPGO16, KT90a, LM89, AV11, BFMS87, KLM^{+}21, MBPLRR10]. communicating [SR07]. Communication [BC01, LNP^{+}21]. Commutative [Baa89, BP85, Fag87, HS90, JM93, KM01, LC89, Pau92b, Pri96, RST01, Wra88, Yap91, AB92, BP24, BTBQM00, CS98, DKM21, GTLN16, HKSS17, KRW90, Kin14, LL09, Lab90, LMZ23, Ros23, Sch19, Sti03, ZGG23]. Commutativity [For87b, Zha90, HL21a]. Commutator [HKL99]. commuting [Hre06]. Compact [Nor95a, GK21]. Comparative [AM99]. Comparing [BF93, KKSd96, MPS02]. Comparison [CJUE01, GVHHUE05, Gre95, Lav91, BF20]. compartment [MS14]. compatible [KZ14, LLW03, SS06]. Compatibly [BCS97, KS12a]. compiler [ST89a]. Compiling [PS93]. Complement [Fer96, BGS11]. Complements [Wal00]. Complete [BGM15, BCGY12, CGY09, EW07, KP99, Le 86, Lyn97, SG89, Bed07, Bed09, BGMSG07, BM04, BSW21b, FS12, GS89, HDPS11, LJ09, RS10, RS11b, Sop13, Sta16, Vac17, Win14]. Completely [Ber02, CS99, BJM17, GN04]. Completeness [MT93, Mid94, PP91a, Pau92b, TRRK10]. Completion [BD88, Buc87, BGK96, Car01, Com98a, Com98b, Fri89, Gan91, KMN88, KT08, ML92, Soc91, Sto99, Che18, GR10]. Complex [BKHG21, Gaá95, SME87, Str97, Wal00, Zha96, BSSY18, DNS21, DGW19, GMMM17, Lou08]. Complexes [Aur87, RS00, AKL17, BC05, CVY17, DiP16, DE03, RdC13, DRN24]. Complexities [SW95, GVHHUE05]. Complexity [BS88, BKN87, Bir98, BK90, CR90, Gao01, Gri88, Gri90, GK16, HK95, HS90, Kal99, Meg90, Nie94b, Ren92a, Ren92b, Ren92c, Ris88, RdC13, RS90, Sch03b, Tak95, Van00, Vor92, Vor99, Wei94, Wei88, Wei90, vdHS06, Ahn08, AKS12, BFS15, BP09a, Ber04, BDS17, BKSV20, BS09, BGT20, CH17, DET09, ES18, EGS23, EMT21, FES11, FES13, FEV16, GSSST10, GPMS20, HvdH18, Lec19, MOP15, MT20, Mor91, NPP17, Pic03, Pol95, PS13, Sag14, SvH19, WY20, vdHL13]. Compliant [BGK96]. complicated [DO06]. component [APS12, BR10, CS22, EW00, EGS23, JLR03]. Components [Hub99, BS09, KN11, KL17b]. Composable [Ohl95]. Composed [HM02b, Min02, CKM09, Min03, Min06]. Composing [Ber98a]. Composite [LS94, SK91]. composites [GPGO16, Naw16]. Composition [CH97b, GS98, Hon97b, Hon98b, Nor02, AR06, BCR15, Kan91, LHK^{+}13, LLW03]. Compositions [BvzGZ13]. Comprehensive [NOT18, Wei92, Wei06, KSW13a, KSW13b, MM09, SS03b, Wei03]. Comput [AP04, AK06, AP17, CS09, DHM11, FS13, Fer06a, HZ15, HJM94, Hil05b, HdC16, HP08, Kal90, KT94, MHXD09, RS11b, Sag89]. Comput. [KL90, Nor95b, KMR18]. Computability [Bac94a]. computable [JPP19]. Computation [AS05, AC19, Ano01d, AJ01, AH86, BVW18, BFK02, BS92, BCGR92, Bod04, Bor22, BE02, Bre86, BW03, Buc92, Bur92, CV00, CL00, CGG89, CZG02, Chi96, CL23, CCM95, CR88, Dav94, DT95, DR86, Eck87, Edi85, EKP22, EY21, FGT95, FGLM93, Fit89, FGT15, FFP98, GK00, GH97, GL92, GK94, GS06, HPS22, Hol85, Hol91, Hon96, Hug90, JPPSG09, JKP12, JKP98, Kal00, LPY01, LM89, Lev99, LR01, Lüb02, Mar02, MF90, Mau87, Ng89, Ous91, PSZ91, PC98, PZ92, PS95a, PD07, PdRAEC24, QR07, RT85, RS90, Sal94, Sav90, STA94, She97b, Shi04, Ste97, Szi17, Tes99, Tho02, UYSA89, VV97, Vas00, VGT90, Vor89, Wal00, Wei00, WBM99, Zha96, dB89, ACMB19, AAKM21, AH05, Arr16, ACS13, BHLGO15, BGH^{+}04, Bar13, BE13]. computation [BC24a, Bau15, BBRS24, BCLR13, BFH17, BFSS06, BY23, BBKK15, BSC12, BR88, BS15, CFS07, CH03, CVY17, CLL17, CM17b, CN07, DM09, DK18, DEG^{+}21, DDM15, Dor21, DPS17, EH21, Eid23, Eli04, Els12, ES13, EFG16, ERSG05, EH16, Fox18, GS03, Gue18, HJS16, HKPP09, HMN06, HKP^{+}06, JFMRS12, JNV21, KS06, KS12b, Kin14, KS04, Kut10, LLM^{+}13, LR90, LLL08, LL16, LS12, LFD19, LJ09, Lia13a, LHK^{+}13, LLL13, MM04a, MJK17, MS03a, NSW16, NWW21, Orl22, Par04, PT14, PS95c, PNM13, PH11, Poz24, RS21a, Sha12, Smo21, Sut12, Sut13, Sut16, TM85, UCJ04, Vac18, Vil11, WyW93, WC12, Win88, Yok17, YW87, ZL12, vdH07a, Ano01b, Ano02c]. Computational [BdlCRLS19, BS01, BdS01, Bos01, BK90, Cha00, EL12, GH05b, Jac97, KP13, Kal97b, MCMMPR14, Mic90, Ren92a, Ren92b, Ren92c, Sha13, Stu91, YXL99, Bar13, Ber04, BJT22, BEG09, CR98, DPS16, GKMW21, GKM05, Kem16, Ker17, Mer10, Poz15, TZ21]. Computations [Ape98, Bec90, BDPR13, But85, CDO97, Cuy97, Dab01, DGS10, DSV01, HKK98, Kal02, Lim93, Mee94, Mro96, Pau92a, Whi91b, von87, ABFS15, BD09, CHU19, CC24, DZY22, DK16, Ebe19, Ede13, Els15, EK19, EPY98, GIM07, HPRS11, LR15, MS22, The06]. Compute [BH00, Die92, Mau00, NY99, BV03, BC22, EMSS16, FSW10a, FGPT03, Hel16, JY17, MPT20, OdR03, OdR09, Pop15, van93]. Computer [AP90, AGM97, AP93, Ano99c, BC01, Bel03, BGH93, Ber93, BK99a, CP97, Cap90, CO01, Cav86, CF91a, DGS96, Dav88, Dav02, Die92, Div91, DJ89, Ebe01, FS95, FM02, FK89, GV99, GSZ85, GZ90, Gre00b, HS95, JT03, KS98, KKL92, Koe92, KG03, KR97, KD90, Lab95, LS02, NMM90, OT87, OZ94, Par08, PW90, Pro00, Rob97, Rd91, Ryb90, San95, Sch96, Tra07b, Vei97, Web95, Wil95, WBM99, YP91, Zha93, AG91, BW05, BKG21, BKHG21, CS05a, CLS91, DT06, FGPGP14, FKO18, Ger19, JMPR04, KASW05, MSZ09, NSW16, Oll88, PS89, SMB03, Yok17, KK09]. Computer-Aided [Wil95, DJ89, KD90].

Computer-Algebra [KKL92]. Computer-assisted [FM02]. Computer-Generated [AP93]. Computers [Bos97, She97a]. Computing [ABKR00, AKR05, ABPR21, AK00, AH13, Ald23, Alu03, ÀL06, AHS21, ABMN10, AOW23, AE05, AV00, Bah01, BO04, BR09a, BP24, BP99a, Bas06, BEP13, BS88, BGI11, BMQS06, BLR99, BCR11, BJS04, BR87, BL12, BJS^{+}07, BDLP22, BTBQM00, BLM10, BEM97, BLW03, BC89, BC91, CH97b, CCH97, CCH01, CH04, CHSS05, CGS97, CDF92, CMP87, CH91a, CGK^{+}21, CD00, CWL08, CDM^{+}13b, CVZ21, CS99, Con90b, CJ97, Dab97a, Dab97b, DH07, DLMM11, DFdG13, DdG21, DS18b, DW21, DDL^{+}23, DW22, Drt06, DMY16, Dun99, Dür89, EENMP19, EW86, EW02, Eic02, EG21, ES98, EYZ21, Ell04, ES11, EK19, Enc95, EG07, FLE^{+}23, FS98, FS16, FS23, FGVN06, Gaá95, GS05, GP24, GS02, GHL21a, GHL21b, GHL^{+}00, GS90, Göb95, GTLN17, GTLN19, GDRV21]. Computing [Hal13, Hal01, Har17, Hel96, Hel00, HM09, HMXD07, Hes02, Höf01, HH99, Hul99, Hul13, IvH17, Jam11, JMV23, JNSV17, JS07, Kal93, KT90a, KRK88, KRTZ23, KZ08, Kem22, KP97a, KC18, KPRT18, KBRV24, KS97, KW88, KW24, Kud22, LS98, LGPS91, LS11, LMZ23, LZ12, Lin18, LPRR02, LV14, LM94a, Man93b, Man93a, Mat01a, MHXD09, MPSXD09, Miy01, Mňu97, Mon02a, Mou98, NS90, NP95, Nor01, Oki23, Pan96, PDS03, PP97, Pis04, PS97b, Poh97, Poz19, PES24, Qur17, Ric92a, Rie93, Rin13, RRS06, RR12, Rón90, Roo13, dC10, SM16, Sch90a, SS16, Sek11, Sha90b, SW02, Sim90a, Sla86, Sla01, Smi00, Ste10, Sti03, Str97, Sut15, Tab13, TV18, TW01, Ung06, VL93, VJ07, Vil95, Vor99, Wan00, Wen06, Yah23, YNT89, ZS01]. Computing [ZW08, dG01, dG11, van94, ASS07, AAFR09, Arn03, BR13a, BCE11, BKW20, Bay03, BBF17, BF22, BLS23, CKR04, CS22, CM12, CvH04, CK12b, Col17, DF08, EF17, ELME23, EGS23, FEV16, FW15, GVHHUE05, GMF13, GKsL03, GR11, GMMM17, GMP22, HM23, HDPS11, HHK17, IK21, JLW13, JGF09, KSW13a, KSW13b, KLM^{+}21, KS12a, KN11, KP97b, KMR18, Lia13b, MWZ16, MR10, MAN^{+}10, NT21, Nak09, Naw16, PS13, Ren04, RT17, ST20, SvH19, Toh10, Ung19, YY03, Zhe21, de 98]. concavity [CCG06, hHL21]. Concept [BN01, BS90b, Bel03]. concepts [GK18]. Concerning [AP93]. Conchoid [GP13]. Concurrent [Fis96, LC96, SJG96, LMA11]. Condensation [LW01, Ryb01, LMR94, Ryb90]. Condensed [GKLM91]. Condensed-Phase [GKLM91]. Condition [TCT23, CdG09, CO94, CO96, GGG06, HP07, HP08, HY21]. Conditional [DJ96, Gan91, Mid94, WG94, ABFS15, Kap87, Mor13, PSA23, Wir09]. conditionals [SS06]. Conditions [Vir99, EFRS06, FPT04, Lem03, Li04]. cone [DS18a, GOT05, HQS19, MW23, Rob09]. cones [BFMS87, DR23a, dC10]. conference [NSW16, Yok17, Bos01]. Configurations [Stu91, Brå24, CC24]. confinement [ABC22]. Confluence [Kah95, Kap87, Wir09]. conformal [Kol08]. Congruence [Fit97, KR89, She92, VM14]. congruences [Hem18, RS21a]. Congruent [BFHS92, BFMS87]. Conic [Far97, GO00, LW03b]. conics [FS16, SS24]. Conjecture [And95, BP00, Rei99, AML19, AGS18b, Ang24, ASS13, BCCK20, BKG21, BST16, CK19, CIM17, Col05, GKMW21, GS24, GG92, GR22, HLO22, Kli90, KPT15, Mar19b, Kho08]. conjectures [DT23, vdH06]. Conjugacy [BNRW22, PY94, DMW17, GGM10]. conley [BR09a]. connected [BR10, EGS23, Kem22, aZGS05]. Connection [FKT13, Fuc00a, Pet00, Sch01, BR09a, OB03, Sid93]. connection-based [OB03]. connections [EG07].

Connectivity [PES24, CES23].

Conquering [Ste05]. Consequences

[CR90]. Conservation

[Fit89, WBM99, BC22, PH11, Wol03].

conservative [LW10]. Conserved [GH97]. Considered [KMN88, Pro00]. Consistency [LT89, Vor92, GS18, HGKV11, SWF11]. Consistent [Ott91]. Consolution [BF93]. Constant [Wol00b, GG92, HNR24, MP89, Roc22]. Constant-coefficient [Wol00b]. constants [HPS22]. constituents [Pre06]. Constrained [KFK97, NR95, DIdW18, Nal18]. Constraint [ABP96, AR03, CZ92, Frü96, HLS01b, HLS01a, HJA97, KR94, LC96, Pel97, RSS10, SJG96, AB05, SA89, Com98b]. Constraint-Based [ABP96]. Constraints [Com98a, Com98b, DH00, HJA97, NR97, Rat02, BVE21, EBD21, LM92, Ore11, TM85, XZ10, ZWH11]. constructed [LC16, SMJ19]. Constructibility [Ley01, SM18]. Constructible [CP00]. Constructing [AK86, Arn95, Aur87, BK15, CK04b, CFTY97, Dra03, EP02, Gla88a, GHS01, Har92, HPS97, Hul05, KM99, KM01, Let01, Lin91a, Pau86, PS97a, Pos18, Ros93, RP89, Smi93, dGN02, dG09, vG90, DA05]. Construction [AAKM21, Ber98b, BE99a, BU99, Bro01b, For87a, GK96a, GSPB17, HR19, LG21, PW90, Sho94, Yap91, ZGG23, Els17, Fuk04, GSHPBS12, IKGT11, Lab90, MH16, MPP19, Möl88, NÁS^{+}24, Pol95, Ren17, WY11, HK07]. Constructions [DS00, DV21, Ebe01, BGS11, FS23, FGP05]. Constructive [BP00, Bro03, CH85, CH86, GPWZ02, HJA97, HL21a, HL21b, JL91, MMO94, MM88, MRD11, Tak91, Abr17b, AABdG21, Göb98, HT91, Per04, Tak93, U.05, PP17]. Constructor [MT93, Mid94, SS96a, WG94, You89]. Constructor-Based [WG94, You89]. containing [Piq91]. containment [KK17]. Contents [Ano99a, Ano00a, Ano01a, Ano01b, Ano01f, Ano02b, Ano02c, Ano04a, Ano04b]. Context [GGSST10, KF01, SSS02, ACGL04, HNRS21, LPT20, KLV10]. context-free [LPT20]. Contexts [FGT95]. Contextual [Str01, AR03, AB05]. continuation [DEPS11]. Continued [Col16, CK12b, Sad16]. Continuity [ST89b]. Continuous [Bur16, JWC^{+}16, MCJ21, RH18, SLX^{+}13, CCG06]. continuous-time [RH18]. Continuously [Hem02]. Contour [ABY90]. Contraction [BH95]. contracts [MM10]. contributed [Kap06]. Control [ACGR01, Jir97, LHD96, UYSA89, AHKY09, Pal13, PS18a]. Controlled [Fuc00a, WKB86]. Convection [Mag89]. Conventional [Sit97]. convergence [CFS22, PT14]. Convergent [MSKO93, OKK98]. Conversion [Kal99, Tra00]. conversions [HK21]. Converting [CKM97, Kha14]. Convex [ABY90, AC01, JZ04, DHTY04, Fuk04, GSA^{+}12, TRRK10]. convexity [CCG06, HZ19, Sor22]. Convolution [HC12, JB04]. convolutional [GTLN21]. Convolutions [APZ20, VL10]. Conway [CF09b, HL04, KP15, Mar19a, SW97b]. Cooperation [ASJ97]. Cooperative [DGS96]. coordinate [BvdE03]. Coordinates [Hav91]. Coq [DM05, GPWZ02, PMW93]. corank [BP07, BMS20]. Corner [Chi01, FC04, CK03]. corner-cut [CK03]. Corners [DDD95]. Correct [FLOR00]. correcting [BP09b]. corrector [HL16]. Correlation [PSZ91]. correspondence [Pos18]. Corrigenda [Nor95b]. Corrigendum [AP17, CS09, Der87a, FS13, HZ15, HdC16, HP08, KMR18, MHXD09, RS11b, KT94]. Coset [CG02, HH99, Lin91b, Sim91, BGHW06].

Cosets [Sla01]. cosine [GR98]. cost [GMP22]. could [MS09]. Count [BF91, Bry21]. Countably [Sne98]. counterexample [Ang24]. Counterexamples [vdH06]. Counting [AH01, BS24, BR06b, HK95, HI98, Kal01a, Lab92, VW08, vzGGZ21, BS09, CK12c, DHTY04, GS12, GPMS20, Hub09b, MV10]. Coupled [FM02, GZ90, Zha93, KP91]. Course [Jac97, Sit97]. Courses [AGM97, CP97]. covariants [Bed09]. Cover [HLM95, Str01, BLR13]. covering [Wil90]. Covers [RS00, BKW20, Poz15]. Coxeter [BB92, BDPR13, CLW95, Du 99, dM99]. CP [Alc16]. Cracker [CCT11]. Creative [Zei91, CK12a, CvHKK18, GHLZ22]. Criteria [BD88, Pic00, WW94, BBC^{+}17, HA10, HP07, HP08, LLL19, MMS18]. criterion [AP11b, AP17, GHMA13]. Critical [BD88, Bec93, Brå24, Buc87, NMM90, Sto99, dCR17, ASS07, BNT18, Bod04, CKR04, FLE^{+}23, GP20, Hor24, Kad13]. Critical-Pair [Buc87, Sto99]. Critical-Pair/Completion [Buc87]. Cross [OKK98]. Cross-Sections [OKK98]. Crossed [AOW23, BW03, OUI16]. crossing [Mon05]. crossings [BJM17]. cryptographic [BD04, KKK^{+}16]. Cryptography [BCE^{+}94, FP09]. cryptology [FGPGP14]. cryptosystems [CCT11]. cube [FKM10, KC09, TCT23]. Cubic [Gaá00, Lip93, PSV13, CK12d, DF05, FFP98, GSPB17, KT04, Sei20, SS24, SJS06, ZWH11, VM14]. Cubical [BT94]. cubics [BR13b, Eng10, HMXD07, MHXD09, MPSXD09]. Cumulant [PW06]. CUP [JPS13]. Curried [Kah95, KKSd96]. curvature [Hor24]. Curve [AM88a, BE02, HI94, MM00, Ric92a, Ahn08, ASS13, BO04, BGMSG07, BE17, CGL07, CK12d, FS16, FGT09, GS05, Gau09, GMMM17, HC12, Hub09b, JWC^{+}16, MP89, PSA23, RSV09, Sto17, SS24, Zen06].

Curved [GK94]. Curves [AH01, ACOR00, AF88, CLQ10, Cre01, GR02, GSST98, GV97, HSW97, Hon97a, HI98, Kal01b, Kid02, Mic13, Mňu97, RS97, SF90, Sch92, SGD97, SW91a, SW97a, Sen02, Sha01, ZS01, van97c, AC24, AK04, AK06, AS05, AS07, Alc08a, Alc12, AHM18, AGS16, BGLGM17, BGI18, Ber98b, BCGY12, CV11, CWL08, CVZ21, CJL13, CS16, CS05c, CGP23, DP19, DS18b, DGW19, DR23b, Eid23, El 08, Els15, FGS09a, FGS09b, FS10, FS12, FS13, FDS13, FM24, FG08, FGP05, FL21, GMF13, GP24, Gen22, Har12a, Har13, HJ18, HLO23, HJ15, Hie16, HLSW16, HS98, IMP17, JWG10, KRTZ23, KS12b, KH23, LSY07, LN13, LLL08, Lou21, Lub14, MSS21, MR15, OT13, PD07, Pis04, PSV11, PV13, Qi23, Ren17, Roc22, RSS13, Sad16, SSS23, SS09, SS11, SJG13, Sor22]. curves [Tab11, VL10, Wan04]. Curvilinear [GV99]. Cut [BL00, EW86, BL06a, CK03, Sak21]. Cut-elimination [BL00, BL06a]. Cutting [Chi01, FC04]. cycle [Bor22]. Cycles [LP90, CFS07, HY23b, SH17b]. Cyclic [AF00, BF91, Gaá95, GTLN21, Hil05a, Hil05b, RZAG99, ABF09, BCI13, BM10, GGM10, Lüb23, TS24]. cyclically [BC06, NWW21]. cyclicity [LS16b, Sha12]. Cyclotomic [BCE^{+}01, CM10]. Cylinders [BFHS92]. Cylindrical [ACM88, Arn88b, Bro01a, CH91b, CJK02, EBD21, McC88, MC02, PS00, Str11, Str16, BDE^{+}16, BK15, CM16, KÁ21, LXZZ23, MSW15, NÁS^{+}24, PdRAEC24, Str06].

D [BPH07, LMM05, YXL99]. D’Alembertian [AB09]. Dancing [Hem18]. Data [MMO94, Yan98, DE06, Far19, Hir89, HR17, RT17, SLK11]. data-discriminants [RT17]. Database [GP96b, FGR03, JR06, Moz89]. Databases [AB00b]. Dataflows [YI94]. Davenport [SS90]. Dealing [DM05, LO96]. death [KKK17]. Debugging [DL93]. December [Ano23j]. Decidability [GSSST10, SS96b, SSS05, BHSS89]. Decidable [ARS10, Ott91, SSS02, Sta89]. Deciding [EW00, GRW16, Gri88, LW10, MW12, VGW18, Vor92, dNdR03, DH16, DF09]. decimic [BP10a]. Decision [ARS02, BS96, CCG06, Ren92a, Ren92b, BE10, Bus09, DGLM^{+}24, GX04, STDD16, SS03a, Ye17]. Declarations [MGL00]. declarative [AHH^{+}05]. Decoding [BF01, CP10, ABF09, BB10, BP09b, DS09, LO08, LO09, MRG13, MRG17]. Decomposable [SO89, Sha01, GGMFVT13, HJA17, vzG13]. decompose [BFPT21]. Decomposing [Gri20, Püs02, Wan98, AABdG21, FP09]. Decomposition [AF88, Arn88b, BZ85, BR87, BRM01, BCRS89, Bro01a, CH91b, CJK02, DTGV01, EG00, GR02, GRS02, HLM95, HLS01b, HLS01a, HS00, Hub00, KL89, KLZ96, LSW01, Laz85, MS00b, McC88, Mon02a, PS00, PS93, RZAG99, Rus87, Rut92, Rut93, Sau01, SY96, ZG09, von90a, von90b, AGR95, AF08, BGLHR12, BBCM13, BE10, BDE^{+}16, BK15, BIS16, CM12, CDM^{+}13a, CM16, CHKL22, CFS24, DPS17, Dur09, EG04, EBD21, FGT05, GTZ88, GHLZ21, Gol08, GIJ14, HOP06, JPS13, KMYZ08, KN11, KMM22, KL90, KÁ21, Li04, LXZZ23, MES19, MSW15, MBL21, Nag21a, NY04, OO13, OdR03, OdR09, PS21, PdRAEC24, Rou09, Ste05, Str11, Str16, Str06]. Decompositions [Ae02, ACM88, Bea92, CFM96, Kal94, Klü99, MC02, vW95, AP04, Ang18, ABC22, ADCZ21, ABM^{+}23, Bur04, Cer21, CJ22, DFdG13, GKO09, GSSV12, LPR17, MRSW07, Rob09, Zie16, DRN24, vzGGZ21]. Decoupling [BNN17, Wol02]. Decreasing [BFHT85]. Dedekind [For87a, HR19, del95]. Deduction [Ano01d, AJ01, BH95, CH85, CH86, DS96, She97b, Tak91, GSSST10, TRRK10, Com98a]. Deductive [AB01, CP00, DR93, GP96b, Tra89, Moz89]. Default [SJG96, Sto11]. Defect [HLM95, CC07]. defectivity [Abo10]. defects [Mor11]. Deficiency [GRV17]. Defined [Ma94, MG88, Tor93, AB99, EG21, EFG16, GMF13, Kol08, Lou21, MMO94, Oak13, Ryb03]. defines [Qi06]. defining [Ahn08, BDM23, DG20, LR98]. Definite [Dey21, Köl85, JPP23]. definite-sum [JPP23]. Definition [CG02]. deflation [HMS17, LLZ24]. deflation-one [LLZ24]. Deformation [Gal13b, MP89]. Deformations [AC24, Mar02]. Degeneracies [Yap90]. degeneracy [Mer10]. Degenerate [Roj99, GP20, HNE21]. Degenerations [CM97, CM21]. Degree [AF00, BL96, But93, Dür89, Gaá00, KST93, KM00b, Lia22, LPS93, Mal87, Mau00, Roy87, Ahn08, AKS12, ABB^{+}19, BW22, Bay03, Bed07, Bed09, BvzGZ13, CG23, CK12a, CM17b, Cou18, Cou22, DL88, EGB12, FGS09b, FS10, FS13, FFP98, GKL04, GSHPBS12, Gre16, HPS22, HTZ04, Hor24, HKYY18, IK21, Kan91, Lee08, Lee17, LV14, Nor15, Par22, Píl07, RDU03, SS09, dGPS09, vzGMS10]. degree-complexity [Ahn08]. Degrees [Con90b, SSV23, ZD02, BDM23, CR19, Sla86]. Del [GSHPBS12, dGPS09]. delete [BGS11]. Deletion [AB00b]. delineability [ASS07]. delineability-based [ASS07]. Delivery [Nor95a]. Demjanenko [FZ87]. dendriform [Mad14]. Denesting [Lan92]. d’enfants [HJ15]. Denominators [KT90a]. Dense [AV96, Min03, Lec07, MS04]. Densities [GH97]. Dependencies [Sch91, vdH13]. Dependency [GAO02]. dependent [HS21a, MR13]. Depending [DTGV01, AR13]. Depth [Bec03, BFHT85, von90c, BF20, HS21a, Pop15]. Depth-first [Bec03]. depths [OS24]. derandomization [GSSV12]. Derivation [Bee01, CP93, GHC92, Mau00, Sof94, CO94, CO96, DJ89, FPT04, GMP13, JB04, WS09]. Derivations [FGT02, GL92, You89]. derivatives [Gal13b, Hua23]. Derived [AB00a, OPP93]. Deriving [BB93b, CSS96]. Derksen [Kem16]. Descartes [KM06, Sag14]. Descent [BB92, FG08, BDPR13, Cre01, HKYY18]. describing [CE19]. Description [BN01, CLM16, Göb98, LW10]. Descriptions [NNN98]. Descriptive [Ave86]. Design [CM93, DYA97, GKW98, HNS95, Jir97, Pad96, UYSA89, AHKY09, LS04]. Designs [Key01]. Desingularization [CKS16, Bec09, BE11, U.05, PP17]. dessins [HJ15]. Detecting [AH05, BL98a, GR11, Kal01b, RSV09, Sch91, KL17b]. Detection [HS97, AHM18, JWC^{+}16]. determinant [Vil11]. Determinantal [PV13, Dey21, HESV21, PS21]. determinants [HNE16, HHLQ13, MM04a, Oki23]. Determination [LM90, LW01, Zen06, FGL04, PS89, SK12]. determine [HBN95, SM18, SS88, CC24]. determined [Sza08, Tsa16, Wer12]. Determining [Hen98, LS16b, Mic88, Sch85, WZ12, Yan99, YXL99, ZX20, FK89, GS24, LH17]. Deterministic [Gao01, GKL04, Guo20, HSS18, Kal87, MO98, MS11b, Pol95, ST89a]. Development [AB00a]. devoted [HKP^{+}06]. DeWitt [GK94]. DeWitt-Seeley-Gilkey [GK94]. diagonal [Bri06, LS11, LS12]. Diagonalization [HM97]. diagonals [BDS17]. diagrams [BKV24, CC24, KPRT18, STDD16]. diameter [BGT20]. Diatomic [OT87]. dictionary [STW18]. Dieudonné [Oki23]. Difference [Bro00, FHR99, GV99, Hen98, HS99, Lev00, Wan18, Wol00b, ABvHP11, AR13, ABPS21, BJ21, Cha14, Dun99, FGH08, FSW10a, FSW10b, GLY09, GVYZ09, GHY17, GMF13, HW24, Lev21, LYG15, OS24, LT22, Sch08, Sch16, SvE14, SvE21, YZ21, ZG09, ZW08]. Difference-Differential [Lev00, Lev21, ZW08]. Different [Egl96]. Differential [Ano01e, BP99a, Bar99, BRM01, Bro92, BEM97, CDF92, CV00, Com98c, CS99, CSTU02, Die92, Dun99, GC93, Gri90, GSZ13, Hv95, Hub99, Hub00, Hub09a, Kov86, Lev00, LS01, MC97, Mil87, Mil92b, Mor99, ŌS94, RT89, SV92, Sch99, Sch85, SS95, Sin90, Sin91, SU93a, SU93b, Sit97, Tra06, Tsa00, Ulm94, Van02, VRUW99, Vid99, Wan91, Wan99, WBM99, Zha96, dv96, van97a, van97b, vdP99, ABK15, Abr17b, Ald23, AAB^{+}18, AMW12, Arr16, BGLHR12, BP09a, BCE11, BE13, BELP13, BM19, BSW21a, BC24a, Bîl11, BD12, BLM10, BLL^{+}16, BLPM19, Bro90b, CFS22, CvH04, CQ12, Cou18, DJO^{+}11, DS86, Dra03, DW22, DP09, ES23, FGLH^{+}23, FSW10a, FSW10b, Fre04, FL21, FK89, GH05a, Gao03, GVYZ09, GGG06, GLLdR19, GHL16, GV16, Gol06, Gol08]. differential [GKO09, GOT05, GS18, GOP18, HT91, HI08, Hil87, HY23b, IvH17, JLR03, Lev21, LG21, LGM21, MS03a, Mil93, Nak16, NNvdPT15, Ngu09, PH11, RR08, RS19, RS10, RS11b, Rue11, Ulm03, WK20, ZW08, vdH07a, vdH07c, vdP05, vdPT15]. Differential-difference [Dun99]. differentials [HH07, LN21]. differentiated [Vil11]. Differentiating [AZ90]. Differentiation [Wan94a, HLXL18]. diffgrob2 [MC97]. Diffusion [BH95, Bon96]. Diffusive [Mag89]. Digit [Jeb95, Rou08]. dilated [BVW18]. dilogarithmic [KS19]. dilogarithms [Bad06]. Dimension [Chi96, GHL^{+}00, HS21a, Lev00, MR13, Meg90, Vor99, BMNB^{+}11, CFRS23, DJ07, Eng10, FD18, Gau09, Giu88, HJ18, KW88, Lev07b, Lev21, LV14, Lou21, MRSW07, MV13, PS89, Shi04, ZW08]. Dimension-dependent [MR13].

Dimensional

[ACM88, ARE02, FGLM93, Laz92b, McC88, Nor95c, Sak88, AKR05, ACMB19, AOW23, BRM01, Buc06a, CES23, CGY09, CGG12, CJ15, Dur09, ES20, FMR04, HOP06, HKPP09, HSV08, HKYY18, JWC^{+}16, KMH89, LST03, MWZ16, MM04b, MP04, Mon02a, Mos08, NT17, NY99, Nor95b, PP17, PS13, SS90, ST20, Shp14, XWL23]. Dimensions [AP08, EW86, EPW90, BLV18, EKP22, MS11a, PH11]. Dimer [BGH93]. Diophantine [AP11a, Cip08, CF89, CKS99, FT95, PT98, PV00, PV02, Wei88, Wei90]. Dirac [RS19, TM85]. Direct [ZS01, CJ22]. Dirichlet [OZ94]. Disc [Pan96]. discipline [CLS91]. discontinuities [AP90, FL11]. discontinuous [JT03]. Discovering [APS12, Fox18, SLX^{+}13]. Discovery [Wil95, Sil04]. Discrete [MM00, OZ94, BJT22, CM09, Gau09, KKK17, MS03a, Vac17, vzGM22, VM14]. discretizations [WRI09]. Discriminant [HM02a, Lip93, Sma96, BHM^{+}23, IMP17, MM06, Mor11]. Discriminants [McC99, Nie12, BMT21, CC07, LM09, RT17]. Discriminator [Bur92]. Discussing [Mon02b, DHM11]. Discussion [YYZ12]. Disjoint [BS96, CJ22, SS89a]. disks [KC09]. DISPGB [MM06]. Dissections [Lis95]. dissipative [KP91]. Distance [CFSGL21, LN21, BP09b, BGT20, ÇJM^{+}21, Lee17, Toh10]. Distances [Hav91, GGMAMF19, GMMM17, GTLN21]. Distinct [MM00, GR12, GKL04, XXZZ23]. distinct-degree [GKL04]. Distributed [BH95, Bon96, CM96, DGS96, DS96, LC96, STA94, ZBH96, SMB03]. distribution [BP07]. distributions [BM10, KP91, RS19]. distributive [SS03a]. Distributivity [SS96b, TA87, Con93]. Disunification [CL89, Lug95]. Division [Jeb93, KJ96, Rol90, BF22, Eid23, MP11b, Nak09, VK16, Vil11]. division-free [Vil11]. Divisions [Ape98, Rol86]. Divisor [DTGV02, MG94a, van97c, GGMAMF19, JPPSG09]. divisor-closed [GGMAMF19]. Divisorial [Vas00]. Divisors [KT90a, LM89, Lüb02, IS10, KLM^{+}21, LLL08]. Dixon [Chi01, CZG02, CK03, CK04b, CK04a, CKM09, FC04, Pal13, QZYC22, Sch90b]. Do [She97a]. documentations [SWF11]. Documents [CC01]. does [LMPX11]. Domain [For87a, MPS02, CK90, DMY16, KRK88, QHL^{+}13, Sek11]. Domains [AHLM99, KR94, Pan89, del95, DF08, EPP21, FD18, HL21b, LS12, MS11a, Oak13, VK16]. dominant [Hon04]. Don’t [vdH02]. Dotted [MW91]. Double [AP93, Jeb95, LP03, Lin91b, Mag89, Sla01, BGHW06, DHKS07, DHKS09, FGS09a, FGS09b, KS12c, RS11a, RS20]. Double-Diffusive [Mag89]. Double-Digit [Jeb95]. Double-Sum [AP93]. Doubly [CC91, DH88, LH98]. down [Fre13, MBL21]. DPPEs [RS11b, RS10, Rue11]. Dragging [SMJ19]. drift [KLR93]. Drinfeld [DNS21]. Driven [WKB86, HT23]. drops [ERSG05]. DSC [DHK^{+}95]. Dual [BE02, BDM^{+}16, BU14, CC07, DP24, HK10, KL17a]. Duality [JJ01, Pap23, Jou09, Wan23]. DVR [Car15]. Dynamic [DTL10, Duv94, GMS09, Pas86, BKR19, KRV19, SMJ19, Str19]. Dynamical [CD00, LP02, Ous91, CDSS09, KS06, SLX^{+}13]. Dynamically [HKK98, GPMS20]. Dynamics [GG99, HS89, Gat03, MJST22, Sad17, SMJ19].

E-CCC [Hir89]. E-services [Ran12]. each [EGS23]. Early [KsL03]. Easy [LHK^{+}13]. economy [Pra13]. Ecosystems [MRS96]. edge [EFG16, FC04, HDPS11]. edge-adjacency [HDPS11]. edge-skeleton [EFG16]. Editing [CH95, vdH15]. editor [DGPP10, Bos01, Buc92, Hon96, Kut10, Lev07a, Smo98]. Editor-in-Chief [Buc92]. Editorial [Ano18g, BK90, Cav00, CGT04, DS06, Edi85, Hon00, Lam97, LS02, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17a, Ano17b, Ano17c, Ano17d]. Editorial [Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18h, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano20a, Ano20c, Ano20b, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21a, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano24a, Ano24b, Ano24c, Ano24d]. editors [CJGV09, JMPR04, Ano99c, Ano00b, AJ01, BB93a, CH97a, CJS01, CL00, DGR07, DDHS13, HSW97, HL97, KM98, KW13, KR97, LPPR12, MMY00, MNJ94, PZ92, PS95a, SB99, WS98]. education [GGAVRC13]. effect [DDR11]. Effective [ACMB19, Ano01e, BS22a, BBC^{+}17, But85, DGPP10, DHTY04, Dub93, GPS09, GS18, HS01, MMS18, MS21, MS10, MGRR23, MS22, NP20, VB03, vdH05, AMDW16, BCLR13, CM17a, CDM^{+}13b, CFS24, DDM15, DJ07, GX04, GDRV21, HV16, Kad13, Mor20, RRS19]. Efficiency [BPH07, Gre95, HcL21]. Efficient [BTW93, CO01, CF89, CM93, CE95, DF05, DSV01, EG00, EG04, EH21, Eid23, ES98, EC95, EFG16, Geb02, GHLZ21, GHLZ22, GKW98, HI94, Lim93, Mag17, PH87, SS96a, Sut12, Sut13, Sut16, ZSY93, ZL12, vdH07b, FP09, HDPS11, JGF09, KSW13a, KSW13b, MTV21, PS18b, Sal08, YZ21, FGLM93]. Efficiently [CKR04, DMS21]. Ehrhart [BEZ23, BS15]. eigenfrequency [KP91]. eigenrings [BSW21a]. Eigenspaces [MT01]. Eigenvalue [For02, CGK09, HHLQ13]. Eigenvalues [Qi05, Koz23, Qi06]. Eigenvectors [OO13]. eight [DS18a]. Eilenberg [MGRR23]. Einstein [ACS13, ASS20]. Eisenstein [DF05]. Elastic [Eis90]. Electronic [BC01]. Element [TL96, Wan86]. Elementary [AM88b, Bro90a, Bur04, DS86, Lüb02, Bad06, DHKS07, El 03, PV13, SR07, ZWM15, Zha03]. Elements [BBB92, CH96, Gaá95, KT04, YNT89, Buc06a, CF09b, GTLN17, SS16]. Élie [NPD09]. Eliminating [KL17a]. Elimination [Arn88a, AM88b, CH97a, CH91b, DH88, DS02, DYA97, DKLP21, EW00, EM99, Fer96, GV96, GVGC99, HL97, HLS97, Jir97, KFF88, Laz88, LS95, Mul01, PS00, Ren92b, Ren92c, SK91, Vir99, Wan93, Wei97, BL00, BL06a, BGG13, CM16, CS98, EH16, Fer98, GGL06, GOP18, HSV08, HE12, JPS13, Sch07, SD05, Tra07a, XLY15, Zan94]. Elimination-based [DKLP21]. ELISE [Die92]. Elkies [BW22]. ellipse [JH21]. ellipses [GKR23]. Elliptic [Car99, Cre01, Gar95, HH98, Kid02, MV10, Ye18, BGH^{+}04, CV11, FG08, Gau09, Hub09b, Sad16, GKR23]. embed [DW18]. Embedded [BE11, BCS97, KL17b]. Embedding [AAB^{+}18, BS87, Rd91, Vel00, LL13]. Embeddings [GR01, BELT21]. Empirical [AGMT98]. empty [Fer98]. encoding [Cox19]. encodings [Vat12]. Endomorphism [GHS01, Sch90a]. endomorphisms [DL06, HLSW16]. enforcement [LMA11]. Engel [CdG09]. Engine [WKB86]. Engineering [KC01, Mer01]. enjoying [HL21a]. enough [HL21a]. Enriched [Lab92]. entries [MM04a]. entropic [STVvR24]. Entropy [Gri22]. Enumerating [You89, KMY24]. Enumeration [CP00, CG02, Lin91b, LMM05, Sim91, HV16, Pel03a, TS24, FGS09b]. enumerative [DaZZ04]. Enveloping [dG01, AL88, FS98, IL09]. Environment [DGS96, HL98, BPT11]. environments [SMB03]. epidemic [BENW06]. epimorphisms [Poz19]. equal [LMPX11, Roc11]. equal-spaced [Roc11]. Equality [Dav02, Hsi87, NR95, BPT12]. Equation [BTG02, FT95, FHR99, Hv95, Hub99, Wol00b, Arr16, BL06b, Bro90b, Che23, DZ09, FGG^{+}16, GH12, GIM07, KP91, LL16, Maw88, Mil93]. Equational [AB99, BS96, BS86, BHSS89, CZ92, CL89, DR93, GR10, HK95, HKK98, JM95, Lyn97, McN92, Pau85, QW96, SS89a, SS89b, WG94, AHL03, DKM21, EBD21, LM94b, Pic03]. Equations [AP89, AK00, AHW05, BP99a, Bar99, BGH93, BF91, BGK86, Bro92, Bro00, CDF92, CV00, Chu99, CF89, CM93, Com98c, CS99, CSTU02, CKS99, Cza89, Die92, DG20, FM02, GPP93, GP96a, Gaá00, Gaá02, Gan91, GC92, GH97, HH98, Hen98, HS99, HPT02, KST93, KFF88, Kov86, LS01, MC97, Mar96a, Mil87, Mil92b, Nau98, PV00, PV02, RT89, SV92, SSS02, Sch99, Sch85, SS95, Sin90, Sin91, SU93a, SU93b, Sit97, Sma96, Sny93, SBB^{+}89, Tra98, Tun02, Ulm94, VRUW99, Vid99, Wan99, Wol02, YNT92, Zha96, dv96, vdP99, ABvHP11, AR13, ABPS21, Adl16, AAB^{+}18, AP11a, AC04, AHL03, BV03, BP09a, BC24a, BB10, BPZ06, Bîl11, BS22a, BD12, BR06b, CFS22, Cha14, Cou18, DJO^{+}11, DS86, DP09, Drt06, DJ89, EG15, FGH08, FSW10a]. equations [FSW10b, Fre04, FK89, Gal87, GH05a, Gao03, GGG06, GSZ85, GS18, GR98, HL17, Heu98, HTZ04, Heu06, HW24, IvH17, Izu16, Kut07, KKM15, LE22, LGM21, MP09, MRV21, Nak16, NNvdPT15, Ngu09, PH11, Roc22, RT17, RR08, LT22, ST24a, SvE14, SvE21, Tun09, Ulm03, WK20, ZWH11, vdH07c, vdP05, vdPT15]. equilibria [BENW06, Tsa23]. Equisingular [CGL07]. Equivalence [BL93, HS90, LWXZ23, NNvdPT15, Bîl11, BLPM19, CF09a, DKLP21, LWX23, MS16, MV15, SS88]. equivalences [HJ18]. equivalent [CO96, Nak16]. Equivariant [GG99, Wor94, BK16, Pos18]. Errata [Kal90, KL90]. Erratum [AP04, DHM11, Fer06a, HJM94, Hil05b, Sag89]. Error [Che85, Kno92, Kno93, Mro96, VGT90, BP09b, KS19]. error-correcting [BP09b]. Error-Functions [Kno92, Kno93]. errors [GLLZ23]. especially [Sut15]. Essential [Hub99, FKO18]. essentially [Ber98a]. Estimates [Bea92, Sha90a]. Estimating [KO17, Tsa16]. Estimations [Ded97]. estimator [CS21]. Eta [LLL08, HR19, Smo21]. Euclid [CCD^{+}09, Gir21, Jeb95]. Euclidean [Bus09, BL98a, BBN18, Col02, EPP21, Hav91, Kal93, KRK88, Lee17, Mv90, Nor90, Orl22, Rol86, RS13, VK21]. Euclidean-like [BL98a]. Euler [FW15, Hel16, KP91, RdC13, XW20]. Evaluating [OPP93, Pra13]. Evaluation [Ant05, AB89, DSW09, Dur09, Duv94, Köl85, Mer01, vdH01, BES13, CCQ18, Cox21a, Piq91]. Evaluations [KT90a, KY16]. Even [BCE^{+}94]. every [BR10]. Evolution [BPT12, GH97, Sch94, GSZ85]. evolutionary [Adl16]. Exact [CK03, Dor21, Eit94, HNE21, Jeb93, KLYZ12, KJ96, LZS11, Wol00a, BGH^{+}04, CG06, GPMS20, HDPS11, ME21, MW23, Nal18]. exactly [BS24, Wan06]. exactness [LR23]. Example [Duv94, Vor89, HTZ04, Kos07, MSS21]. Examples [BGK86, Hav91, Laz88, Sta16, BO10, Sch06]. Except [BE99b]. Exceptional [HRT01]. excess [Rod15]. Exchange [GKW98]. exclusion [HLXL18]. Excursion [Gar95].

Executable [Ano96, ET96, Mal21]. Exhibition [EW02]. Existence [CFS22, Far19, Gol01, CCF^{+}15, CDWZ21, EG07, VGW18]. Existential [Ren92a]. Exp [Sha90a]. Exp-log [Sha90a]. expander [HL18]. expansion [vdH11]. expansions [PPR13, WK91]. Experimental [Gre95, PP91b, PNM13]. Experimenting [HJM93, HJM94]. Experiments [ML92]. Explainable [PdRAEC24]. explanation [TRRK10]. Explicit [CJUE01, Cou00, Els15, HH98, Har13, KM00b, Köl85, KS21, Led00a, Vel00, XW20, BDM17, Jou09]. Exploiting [EP02, BH21, LXZZ23]. Exploration [MPS02, TZ21]. Exploring [ABFS15, HL98, Tsa23]. Exponent [Vor92]. Exponential [DH88, GV88, Bas06, CvH04, Fre04, HL17, XLY15]. Exponentiation [GvPS00]. Exponents [Pel97]. expressible [BGH^{+}04]. Expression [CJMP97, NPD09, Sto11]. Expressions [Bac94b, BFHT85, BS87, GKW98, OPP93, RW94, WBM99, Zip85, BBK14, DM05, KR22, MCJ21]. expressive [Hir89]. Ext [CGS97, Kos07]. Ext-quiver [Kos07]. Extended [AAKM21, BB93b, HA10, HR11, MS95, SL92, CHM12, Kem16, WWWX23]. Extending [DJ96, LG21]. Extension [GS23, Ott91, Smi00, YNT89, AL88, AB05, Har12a, JPP19, JPNP23, Kun18, ST19]. Extensions [BH00, CLW95, CcK02, Dab01, Gre95, Led00b, MQS00, SW97a, Vas00, del95, ABPS21, Bur16, BLW03, Fre04, KT04, LW10, Mau87, NP20, Sch17b, Sut12, Sut13, Sut16]. exterior [HT91]. External [SP10]. Extracting [AGT13, Pel03b]. Extraction [Tak91]. extraneous [QZYC22]. extrapolation [vdH09]. extremal [AC19]. Extreme [BS17b]. Extremely [BM16a]. EZ [Tsu09]. EZ-GCD [Tsu09].

F5

[AP17, AP11b, EP10, Vac17, Vac18, VVY21].

F5C [EP10]. face [Tsa16]. facility [AB05]. Factor [Boy93a, Boy93b, Col04, CTY10, GNP12]. Factorial [Pau95, JPP23]. factorial-basis [JPP23]. Factoring [BS90a, Baj86, CE96, CD85, Gao01, Gie98, GHL16, Gri90, LST03, OP05, Pau01, Poh05, VH98, Wei13, vzGP01, CL17, DNS21, DMS21, Gen07, Guo20, RR08]. factorisation [Sal08]. factorise [DEP22]. Factorization [BTW93, Cza89, Hub00, KT90a, LSS19, NSW85, NG93, Pan02, Pau95, Sho95, SW91b, Vir93, YNT94, del95, van97a, van97b, AHKY09, BCG10, BU14, BDM^{+}16, Car15, CG06, EP04, EGW09, GKL04, GTLN19, GNP12, KMYZ08, Lec07, LH18, LWX23, MS11a, MSW15, Min98, MT20, Rob04, Rup04, Sch19, Ste05, VK16, Whi91a]. Factorization-free [Hub00]. factorizing [MKF93]. Factors [CTR99, McC99, Pan96, Abb13, AV11, AGT13, CGK^{+}21, GR12, Gre16, HJS22, Kan91, QZYC22, Sek09]. Failed [BCE^{+}94]. fails [VK21]. Faithful [dGN02]. fall [CG23, HKYY18]. Families [BS21, LPY01, RT89, ASS13, BKW20, CR11, FK04, HTZ04, Lou08, Lub14, Pet10, PT98, Wib07]. Family [GH02, HPT02, ACFP12, EG21, GEL05, Heu98, Heu06, Ren17]. fan [Aue05, Gol06, MR88]. Fano [Yah23]. Fans [CM97, MR17]. Far [MN89]. Fast [BL98a, BFSS06, BY23, BP00, Bre86, Bro12, CC91, CZG02, Col02, Cox19, Cox21a, Cox21b, DS12, DPS17, FT95, Gaá93a, Gal87, JNV21, Kal85, Kal90, LMS09, LR01, MS03a, Mul90, OS04a, Poz24, PWZ18, Ros23, Sal94, Sho94, Sny93, Tra00, vdH01, BLS23, CCD^{+}09, CM04, DPS23, HM21, JB04, LS16a, LMRS11, RS13, Zhe21]. Faster [AGR16, AIRR12, BF91, Har09, Har14, Hre94, HG20, KMR18, MR10, Roj99, CL17, Rio03]. fat [BDS^{+}18, Szp22]. Faugère [EP10]. Fault [Abb17]. Fault-tolerant [Abb17]. Favorite [Kal00]. feasibility [AIRR12]. Feasible [von87, HNE21]. Feature [Bac94b, DR92]. Features [Buc87]. February [Ano24e]. Feedback [DYA97]. Fermat [AML19, HS09, HJ15, Lee17, SS24]. Fermions [Hug90]. Few [KM00a, Bas06]. Fewer [BS90a]. Feynman [BKSS12]. FFT [Cla22, Van02, vdH10]. FFT-based [Cla22]. FFT-like [Van02]. FGLM [BF20, BTBQM00, DH17, FM17, HNRS21]. fiber [DS18b, HQS19, RS16]. fibers [CTV16]. fibrations [RRS19]. Field [Bro92, Bro00, Gre95, HJ15, LM89, McN92, MQS00, Rut93, SW97a, Str97, van94, ABPS21, BFH17, CM21, CK12c, DM05, EPY98, HW24, KR23, PZ12, Sch08, ST19, SvH19, Wen06]. Fields [AF00, AH01, Arn88a, BCS97, Bru01, BW87, CcK02, Dav94, DS97, EG00, Enc95, FGT02, Gaá93a, GPP93, Gaá95, GP96a, Gaá00, GvPS00, Gie98, HM02a, Hes02, HPT02, HI98, Kal87, Kem96, LPY01, Mv90, MQS99, NG93, PW90, Pau01, Poh97, Rol90, Sho94, Smi02, Ste97, Stu00, Wor94, YNT89, Zha95, vG90, von90c, vzGP01, AP11a, AP22, AL88, BES13, Bau15, Bel04, CL17, CCQ18, CM10, Cox21a, Cox21b, DS12, DFO13, DT06, DNS21, DFS11, Eid23, EK19, Fie04, FS23, GKL04, GH12, GS12, Gen07, GNP12, Guo20, GS06, HL04, Heu06, HL18, JT03, JPPSG09, JR06, KO17, KT04, Kli90, KC18, KS21, Kut19, Lüb23, NT21, NY04, OP05, Poh05, Rob04, Roo13, Sad16, Shp14, Ste05, Ste10, Sut15, Top14, Ura24, Vac17, VK21]. fields [VV18, Win14, YY03]. Fifteen [But93]. filiform [CGGO09, EKP22]. filtration [DS18b]. Filtrations [MS02]. final [HY23a]. Finance [BTG02]. find [BvdE03, SJG13]. Finding [AF00, BP98, FT95, Gaá93a, GLW99, Ges95, Jeb95, Kan91, KKM15, Lo98a, LO99, MM00, Sak88, Tak95, Tra98, Vat12, AV11, BN04, Bîl11, Buc06a, FG06, HNE16, JPP23, KH23, Pan02, Raa12]. fine [DH07]. FINGER [Wan86]. Finite [AH01, BB92, BE99a, BM01, BCS97, BL96, BL93, CH91a, Che85, CO94, Dav94, DHS98, EG00, Ebe01, EW02, Gal22, GvPS00, Gie98, Gla88a, Gla88b, Gre95, HS99, Höf01, HI98, Kal87, Kar85, Kem96, LM90, Mv90, NG93, Nie94a, Nor95d, Ous91, PW90, Rón90, Sak88, Sho94, Tha93, TL96, Wei94, Zha92, vG90, von90c, vzGP01, Bad06, BES13, BP24, Bel03, Ber98a, Bor22, Bur04, CH03, CH04, CELG04, CHSS05, CHU19, CGS97, CL17, CvHKK18, CKLZ19, CN19, CLM16, Cla22, CK12c, Cox21a, Cox21b, CN07, CO96, DA05, DS12, DFO13, DW18, DNS21, Drä01, DFS11, EENMP19, Eid23, GKL04, GH12, Gen07, GMP13, Guo20, HL04, HMN06, HY23a, HL18, JMV23, JPP19, JPNP23, KZ08, KS19, KO17, Kin13, Lea06, LS16b, LST03, Lüb23]. finite [Mag17, MM04b, Mur23, Nie03, NY04, Pel03a, Poz19, Shp14, Sil04, SH17a, Top14, Ung06, Vac17, Wan86, Win14]. finite-dimensional [LST03, MM04b]. finite-precision [Vac17]. Finitely [BRM01, CDO01, GK96a, Let01, Lin91a, Lo98a, MO88, MQS00, NÓ89, OS92, OKK98, PS97a, Vat06, dGN02, BMQS06, CE19, CdG09, DMY16, GGMAMF19, Lab90, MO85, Sch17a]. Finitely-generated [dGN02, DMY16]. finitely-presented [CdG09]. Finitely-valued [OS92]. Finiteness [HdC13, HdC16, CC24, CO94, CO96, DF09, Ric91]. finitistic [Shi04]. First [Ano00b, BZ03, Hol85, Hsi87, Lab95, Man93a, OS92, Pau85, PSS12, Ren92a, Ren92b, Ren92c, ST89a, Sch85, Tre92, BE13, Bas06, Bec03, CFS22, Dra03, FG06, HW24, KPR10, NW10, Str11, VGW18]. First-Order [Ren92a, Ren92b, Ano00b, OS92, PSS12, Ren92c, BE13, VGW18]. Fitzpatrick [LOOR^{+}03]. Five [SW95, DT23, Oll88]. five-body [Oll88]. Fixed [Ley01, Pan96]. Fixing [WBM99]. fixpoint [BSC12]. fKenzo [HPRS11]. flag [ACS13, ASS20, PPR20, Qur17]. Flajolet [SSS^{+}11]. Flat [CR11, Kun18, Kut08, BBV15]. Flatness [Ass94]. Flattenings [Mar02]. flex [CK12d]. float [Abb12]. Floating [Cuy97]. Floating-Point [Cuy97]. floats [Lia13a]. Flow [Fit89, Sav90, YP91]. Fluid [CJMP97, NMM90, YP91]. Flynn [LOOR^{+}03]. Focus [Nie94b, MPH17]. Fold [BB93b, Koe95, IT10]. Fold/Unfold [BB93b]. foliations [Alc16, CS06, Cou22]. FOR-loops [KW10]. Forcing [PP91a]. Foreward [Buc92, PZ92]. Foreword [Ano99c, Ano00b, AJ01, AFP09, BK11, BDT22, BB93a, BK12a, Bos01, CH97a, CJS01, CJGV09, CL00, DGR07, DDHS13, DOR17, FR09, FLB00, HSW97, Hon96, HL97, KM98, Kal21, KOS21, KASW05, KR97, Kut10, Lev07a, LPPR12, LL23, MMO18, MMY00, PS95a, PSS12, PEGS11, SB99, SS18, Smo98, WS98, DGPP10, KW13, Par08]. Form [Boy92, CD00, Dür89, Egl96, GPP93, GP96a, Gaá00, Man93a, MF90, PG86, Sma96, Van00, Vil95, Von98, AMW12, Bed07, Bed09, BPZ06, BLS23, BGG13, Cha14, DL88, FKO18, Hul13, JNV21, JY17, KR23, LM92, LS11, LLW24, Rau11, RZ09, SST18, Sch04, Stu17, WWWX23, WY11, DSV01]. Formal [AC01, BP99a, BJM17, BM19, BSW21a, Bec09, Bie85, CO01, Kal02, LMA11, RH18, Sin90, van97b, Abr17b, HI08, IvH17, VB03]. Formalization [BBN18, FLOR00, Mal21]. Formalizing [IGT15]. Formally [HK10, MRH23]. format [BR06a, ST24b]. Formations [Mur23]. Forms [Ae02, CK99, CD87, Drä01, FH94, GS02, Lip93, Lis95, Man93b, Mar96b, San96, Sch98b, SS95, Sne98, Ste97, Wei94, tW91, AP04, AP10, AH13, ABC22, ADCZ21, BCE11, BE13, BLV06, BFPT21, BC06, BR13b, BS17b, BLL^{+}16, BLPR15, BD13, BST16, CGO88, CV11, CL07, Deu93, DFdG13, GLLdR19, GHL21b, HR12, Kop08, KC18, LS12, LWXZ23, MTV21, PRR18, WZ12, YY03]. Formula [Mul97, Wol00b, MS22, PR22].

Formulae [CH95, DS97, GV97, DE03, EK11, EM12, GHS03, LM94b, SS09, BMT21]. Formulas [tW91, Bro12, KBRV24, XLY15, XW20, ZWH11]. formulation [CK03, CK04b, CK04a, HS17b]. FORTRAN [SR86]. Forward [Dur94, SS96a, JMPR04]. Forward-Branching [Dur94, SS96a]. Foulkes [CIM17]. Found [Lab95]. Foundation [Eis90]. Foundations [ES13, Fre13, JKP12]. Four [AM99, AS24, AC24, BDPR13, BR13c, DEP22, GR12, Tsa16, aZGS05]. Fourier [CR98, CM04, Cla21, Cox22, DE06, JSC13, KS16, PT14, RH18]. Fourteen [But93]. Fourth [FHR99]. Fourth-order [FHR99]. FP [YI94]. fractals [HT17]. Fraction [BCL06, LS95, LS12, Mul01, Col16]. Fraction-free [BCL06, LS12, Mul01]. fractional [Gal13b, GKS12, VM14]. fractions [BLL^{+}16, CK90, CK12b, Sad16]. Fraenkel [Win06]. fragments [ARS10, CCG06, dNdR03]. frame [CS16, FDS13]. frames [FS10, FS12, FS13, Olv03]. Framework [AB99, BFK02, BF93, CH95, DJ96, DH00, OS92, Str01, BD09, CMR19, SLK11]. frameworks [KS18]. Franca [LHK^{+}13]. Free [AS24, HKL99, JM93, LS98, LS95, Lev00, OT01, Ros93, Wra88, AFdCS15, BCL06, Bec03, BD13, CE19, CMR15, CS09, DS97, Gal16, Gal22, HV22, Hub00, LL13, LPT20, LS12, LMZ23, LG21, Mul01, Nag21a, QR07, dCW09, Sch19, Sla07, Smi05, STDD16, Tak91, Vil11, Yel87, ZGG23, dGN02]. free-variable [Bec03]. frequencies [Sad17]. Freyd [DDR11]. Freyd-categories [DDR11]. Frobenius [KS12a, KZ14, Kud22, MD24, Rou08]. frontiers [The06]. fuchsian [CvHKK18]. full [ABK15, SS06]. Fully [KÁ21, LW12]. Function [Ape98, Cap90, Che85, GLLZ23, GRS02, Hes02, Mul97, MQS99, PSZ91, Von98, dM99, van94, Ald23, AGR95, Bau15, BPD19, BR09b, CCQ18, HI08, KY16, KT23, KS21, Nak09, Rei06, Ste05, Sto17, Sut15, Wen06, Ye18]. Functional [FH86, Gib87, HCB96, Sal94, Sch96, von90a, von90b, Ant05, Izu16, JGF09, LW12]. functionally [Loj13]. Functions [BBB92, BS92, Bro90a, Che85, Czi95, DTGV01, Gar95, GHC92, Jef97, Kno92, Kno93, Koh92, KLZ96, LS94, MS95, Mer01, Pro00, SS98a, SS99, Ste95, Tra96, Von98, van97a, vdH01, AJGVS09, AF08, Bad06, BGH^{+}04, BBCF22, Bîl11, BM10, BDM^{+}16, CMV13, CCG06, CS05b, CvHKK18, CDWZ21, Cla21, DHH^{+}04, Eit94, GS03, GPGO16, Gri20, Gue18, GP20, HR19, HT17, JML^{+}13, JPP19, KLYZ12, KS19, Kön17, Kut07, LP03, LR90, Lou21, MSY21, MS21, MAN^{+}10, NN10, Oak13, Piq91, Raa12, Ryb03, Sek11, Sha90a, SLX^{+}13, STDD16, SK12, Str12, VW08, WK20, Ye17, ZX20, vdH05, vdH07b]. Functors [BCF19]. Fundamental [RS00, DL88, FGLH^{+}23, MPSXD09]. further [KR22]. Future [Cav86].

Galerkin [AG91]. Galois

[Arr16, AV00, BKW20, BW22, Ber02, BF01, CS99, DR00, DT06, DTL10, DV00, DW22, Els12, EK19, FS23, GK00, Hv95, Hen98, Kli90, KM00b, Klü00, KS21, Mal00, MZM87, MMY00, SU93a, Sut15, Vel00, Yah23, vdH07a, vdP99, vdP05]. Galoisgruppe [MZM87]. galoisian [Val11]. game [AGS18a, FRR06, KT23]. games [GKS12, TZ21]. Gamma [GK96b]. Gamma-Operation [GK96b]. gap [El 05, HT17, BM16a, OdR09]. Garcia [DS09]. Garside [Bok08, DG14, GGM10]. Gathen [GP12]. Gauge [WBM99]. Gauss [MV10, Sch01]. Gauss-Manin [Sch01]. Gaussian [AH23, Col02, FL04, JPS13, KMM22, LS95, MBC^{+}10, Mul01, Rol86]. gcd [BLV16, BLV18, DF05, CCD^{+}09, CGG89, EGB12, Gri90, HM21, Jeb95, Nag11, Nag21b, SS92, SL92, SS94, Tsu09, Web96, Wei00, vzGMS10]. GCDHEU [CGG89]. GCDs [Enc95, KL98a]. GCRD [WWWX23]. Geddes [GW11]. genera [CLM16]. General [ASJ97, BBCM13, BL85, GMW23, JL91, KFF88, NS90, Ore01, Pau92b, PP17, Ren92b, SO89, Str01, Wol00b, BO04, BLV06, CS22, CMR19, DHKS09, DS00, DJ92, FG06, KS12c, LLL08, MRG17, MO21, NW10, NW11, PR22, VGW18]. generalisation [LR15]. Generalised [Can90, BR88]. Generalization [MR98, MRG13, Sti87, Ang15, CR11, Sch10, Yes21]. Generalizations [BZ93, Win14, DS18a]. Generalize [Pue89]. Generalized [ASS97, BB00, BEM00, CS90, JKP98, Kal93, Key01, Kri85, MTV21, Mul01, PZ96, SML91, Tes99, Vel00, Vil95, vdH07c, Alc12, BGLGM17, Bîl11, BS15, DGW19, DPS17, EK11, FES13, Hal13, HJ15, JY17, LM92, Rua09]. generate [FG16]. Generated [AP93, BRM01, CDO01, FH94, Lo98a, MO88, MQS00, BMQS06, CE19, DMY16, FES11, GGMAMF19, Sch17a, dGN02]. Generating [ACOR00, CM04, DHS98, HL18, MSKO93, MP11a, Ous91, RCK07, Sak88, Sny93, Tho02, dM99, vHKN13, BBCF22, CELG04, CF91b, FW14, HI08, HL04, HM09, HP91, HJA17, Hub09a, JMV23, KT90b, KT94, Kin13, KMY24, KT23, Vat06, VW08]. Generation [BBB92, KL98b, O’B90, Sla07, LW03a, LW03b, Wan86]. Generator [Chu99]. Generators [FGT09, HRT01, LM94a, RT98, BO04, BJS04, Bok08, Hul13, IK13, JWG10, Lüb23, PS18b, VV18, Wen06]. Generic [Ass94, BT98, CH95, CS05c, FH94, KM00a, Led00b, Led00a, Ma94, MSY00, AGS18a, CJL13, CJ15, DJ05, DMW17, DLLP08a, FJLT07, Kal11, Mal21, Zhe21]. genericity [HSS18]. genetic [HS06]. genotypes [Sad17]. Genus [Bau15, GS12, HLSW16, PV00, PV02, Sha01, AP11a, Har13, HLSS15, JH21, KH23, Qi23, Sto20]. GeoBench [Sch94]. Geobucket [Yan98]. Geometric [Baj86, CM97, DH00, HLS01b, HLS01a, HJA97, Meg90, Sch93, Yap90, BM88, DJO^{+}11, DIdW18, LW03a, LW03b, Mor91, QSGB19, RZ09, RR12, WZ12]. geometrical [NPD09, TM89]. Geometries [Deh94]. Geometry [AM88b, CL00, Cha00, Ebe01, FG16, GO91, GO00, GVGC99, Gre00b, HHK^{+}23, Hav91, LR98, Ren92a, Ren92b, Ren92c, Ris88, Stu91, Whi91b, AV11, BKR19, BBN18, CK12d, DGPP10, DR23b, FGLH^{+}23, GS07a, GPS09, GSS05, GMS09, HS17a, IS10, Kap86, KRTZ23, KRV19, KS86, Lan10, LW03a, LW03b, MCMMPR14, SM18, SMJ19, Sha13, Sti03, Str19, STVvR24, The06]. geophysics [DJ89]. Gerdt [MSV22]. German [MZM87]. GFUN [HR11]. GI [YW87]. GI/S [YW87]. giant [BS18]. giant-step [BS18]. Gilkey [GK94]. gimbal [KLR93]. GiNaC [BFK02]. Given [Gaá95, KT90a, Mal00, Sak88, AC19, Bay03, BP07, HG20, PS09, Sek11, VL16]. Global [GGEZ12, Smi00, Alc08a, Bau15, HJX16, KLYZ12, OP05, Poh05, Wen06]. Goals [CSS96, FOT00]. Golay [BKHG21]. Goldbach [BP00]. gonal [BCI13, Har13]. Good [Alc08a, Alc08b, Kid02, PR12, Sch05, Sto11]. Gorenstein [AS24, AC24]. Gosper [BP99b, BLM22, CS05b, LPS93, MS95, Mu08, PS95d]. Gosper-Type [BP99b]. grade [BDM23]. Graded [HNVL90, Rob86, She92, VL93, AC19, BC06, CGGO09, EKP22, HM05, Loj13, MM88]. Grail [RW94]. Gram [Vil23]. grammars [LPT20]. Granular [Sav90]. Granularity [LO96, LHD96, MRS96]. Granularity-Based [LHD96]. Graph [Der13, DBG89, KS97, DK16, FKM10, GES05, HDPS11, IT10, Mar19a, MP14, Ore11]. Graphic [CH95]. Graphical [Che92, KM98, YW87]. Graphs [CI07, Rie93, RP89, BELT21, BFG07, BF11, DMW17, GIM07, GRV17, HJA17, HR20, HL18, KLZA12, Lin18, MP11a, MBC^{+}10, MBL21, Ore11, PSV13, Poz15, XXZZ23]. Grassmann [HT95]. Grassmannians [BBRS24, Coo09, Koh21, ST24a, Wan23]. Greatest [DTGV02, LM89, Pau95, KLM^{+}21, KT90a]. Green [LM90, Mer10]. grid [KS18]. grid-like [KS18]. Grids [GV99, Her94]. Griffiths [Mer10]. Gröbner [BTBQM00, FGLM93, Gre00a, Lev00, MQS00, Tra00, AB92, ABL93, AHLM99, ACFP12, ÁAF^{+}18, Ape95, AK86, Arn03, AKR11, Aue05, BFS15, BCE^{+}94, BS90b, BBF17, BE22, BCR11, BGK86, BR06a, BL12, BV06, Bok08, BD09, BD13, BP09b, BF01, CJUE06, CRAB91, CdG09, CR11, Cip08, CKM97, Czi95, Dah22, DHM11, DS09, Dön13, DW21, DL06, EP10, Ede13, EF17, EH21, FMM07, FMTT13, FL11, FES11, FEV16, Fer88, FF92, FFP98, FD14, FD18, FJLT07, GS24, GG99, Ger06, GTZ88, Gol06, GKMO08, Gon17, GMP22, GSW11, GSZ13, GS98, Hal13, HT95, HPS22, HP91, HKL99, HV22, Hon98b, Hre94, IL09, JGF09, Kal97a, Kal99, Kal01a, KRK88, KRW90, Kap86, KSW13a, KSW13b, Kho08, KM99, KM01, LL09, LL13, Lea06, LS04, LO08]. Gröbner [LO09, LS11, LS12, LMZ23, Lev07b, Li10, LG21, Lia13a, Lia13b, Lia22, LOOR^{+}03, LH98, LLL19, MN02, Mad14, MR98, Mal21, MM09, MR17, MM04b, MRG17, MR13, Mil96, Möl88, Mon02b, MW10, MR88, MS03b, MD24, NOT18, ŌS94, Pau92a, PZ96, Pau07, PPR20, Raa12, Raj06, Rei06, RR05, Ros93, Rou08, Rut92, Sak21, SIS^{+}11, STA94, Sch07, Sch17a, Sei02, SS88, Smi02, Sne98, Ste13, SS03b, Szi17, Tay02, Vac18, Val11, WO06, Wal03, Wei92, Wei03, Wei06, Wib07, Win88, ZGG23, ZW08]. Gröbner-based [Sch07]. Gröbner-basis [BD09]. Gröbner-free [BD13]. Groebner [Tsa16]. Ground [Sny93, AHL03, Gal87]. Group [AH86, Ber02, Bos97, BP00, But85, CCH01, Cap90, Car01, CDO97, CH96, DHS98, GK00, Gol01, Hen98, HLM95, HPS97, KKL92, Klü00, Le 86, LGPS91, Leo91, LM94a, MR98, Mal00, Mic88, O’B90, OT13, RZAG99, Ros93, She97a, Sim90a, Tes99, WK91, dM99, AC04, AT08, Arr16, BS18, Bok08, BK16, CH03, CHSS05, Cla21, Cla22, CF09b, DW18, DdG21, DW22, Ell04, ES11, EK19, FS98, FMR04, Gal22, HK07, Hub09a, JPW19, JWW23, JV09, Kön17, Kos07, Mar19a, MZM87, NPP17, OdR03, OdR09, Poz24, SW97b, Ung06, Wil90, Wur93]. group-based [FMR04]. Group-classified [WK91]. Groupoids [JM93, PV05]. Groups [BB92, BE99a, BE99b, BDPR13, BT94, BL96, BC91, But93, CC91, CH97b, CCH97, Cla91, CDO01, CS99, Con90b, CF94, CFTY97, DEP22, DV00, Du 99, EW02, Eic02, EHR91, Geb02, Gla88a, Gla88b, GS90, Höf01, Hol85, Hol91, HRT01, Kem96, KM00b, Lin91a, Lo98a, LO99, LRW97, MO88, Mal87, MO95, NNN98, O’B93, O’B94, Ost99, PW90, PY94, PS97a, Püs02, RS00, Roy87, RT98, SU93a, Sla01, Tri86, Wra88, dGN02, vdP99, AOW23, AE05, BHLGO15, BKW20, BW22, BCI13, Bay03, BM01, BJT22, BV06, Bor22, BJSS89, BCCK20, BFG07, Bri06, Bro03, BC89, CH03, CH04, CELG04, CHU19, CE19, CGS97, CJ22, CVZ21, CM04, CS09, Con90a, CF91b, CO96, DA05, DR23a, DJK05, DF08, DF09, DFO13, DFdG15, Els12, Els17, ES20, FF17]. groups [FNU16, GGM10, GK21, Göb98, GN04, GMP13, HR20, HLN^{+}21, HY23a, HL18, Hul05, Hul13, JPPSG09, Kan91, KS16, Kem22, Kin13, Koh08, KS21, Kud22, LMP19, Led00a, MO85, Mag17, Mar19a, Min98, Mur23, MRD11, NU18, Poz19, RR12, RDU03, SW02, SH17a, Sla86, Sla07, Smi05, Sut15, WO06, Wur93, Yah23, dG09, dG11, vdH07a]. Growth [Sha90a]. guarded [GHS03, dNdR03]. Guessing [BE22, vdH13, KT23]. guest [JMPR04, Ano99c, Ano00b, AJ01, BB93a, Bos01, CH97a, CJS01, CL00, DGPP10, HSW97, Hon96, HL97, KM98, KW13, KR97, MMY00, MNJ94, PZ92, PS95a, SB99, Smo98, WS98]. Guide [McN92, Rei06]. guided [Rob09]. Guillotine [GZ89]. gyroscope [KLR93]. gyroscopes [JSC13].

Habicht [HY96, LR01]. Hadamard [AAFR09, ÁAF^{+}18, KK09, MG94b]. Hahn [AGRZ99, FHR99]. Half [KT02, AC24]. half-canonical [AC24]. Half-Twists [KT02]. Ham [EW86]. Ham-Sandwich [EW86]. Hamiltonian [AMW12, CK99, TM85, VV97]. Handling [GLLZ23]. Hankel [BS17b, CK12c, Gem94, SL92]. Hans [Ano87]. Hard [Bac99, AvW94, Izu16]. hard-to-solve [Izu16]. Harmonic [BCRS89, DMN17, CM09, DJ89]. Harnessing [KC09]. hash [GPGO16]. Hasse [DNS21]. Havas [Van00]. having [CO96, YYZ12]. heap [MP11b]. heat [Maw88]. Hecke [CD87, Deu93, NPP17, Wan94b]. Height [Boy93a, Boy93b, vzGMS10]. heights [CGK^{+}21]. Helical [FGS09a, FGS09b]. Hellman [Tes99]. Helly [DPS16]. help [LM94b]. Helton [AGS18b]. Henneberg [HJA17]. Henneberg-based [HJA17]. Hensel [AAKM21, BDLP22, Leb15, MM88, MT20]. Herbrand [Pic00]. Hermite [AS23, Apé10, Cox21a, Far19, JY17, LE22, RS21b, RS20, Van00]. Hermitian [Key01, KD90, LO09, MPS16, Sch98b]. heterogeneous [Gon17]. Heuristic [CGG89, Fuc00b, Mon92]. Heuristics [QZYC22]. HFE [GMP22]. Hidden [GKM05, LSSW12]. hierarchical [EFRS06]. Hierarchies [SAK89]. Hierarchy [GPWZ02]. High [BB11, Sto03, BBV22, BLV18, Sal08]. High-order [Sto03]. High-precision [BB11]. Higher [AB01, CH85, CH86, GS89, Lug95, QW96, SG89, VRUW99, Adl16, AC04, BCE11, BELP13, BFK18, OS24, RS16, SSS05, WK91]. Higher-Order [Lug95, QW96, AB01, GS89, AC04, BCE11, BELP13, SSS05, WK91]. Highest [dG01, KLR93]. Highest-Weight [dG01]. highly [BNN17]. Hilbert [Fer06a, Ape98, BS92, BLR13, BM04, CI07, DLMM11, Fer88, Fer06b, FH94, JL91, KS21, La 17, LPT20, Lev99, ME21, MOP15, MdCW17, Pop15, ST20, Tra96, WO06]. Hirota [FM24]. History [Buc87, Mos12]. HNF [BFH17]. Hochschild [ES11]. Hodge [AML19]. hodograph [Far19]. hodographs [FGS09a, FGS09b, FG16]. Hoeij [Bel04]. Hofstadter [Fox18]. Hofstadter-like [Fox18]. holomorphic [CS06]. Holonomic [TW01, Zei95, vdH01, BY23, CQ12, Gri22, JPNP23, Mon05, NOT18, Oak13, SK12, WK20, vdH07b]. Holt [But88, BC93]. Hom4PS [CLL17]. Hom4PS-3 [CLL17]. Homogeneous [ASS20, BJT22, Kem99, Kov86, vW95, AHS21, ACS13, DS16, ES18, FEV16, GLLdR19, GLLdR21, HHPS21, HT17, Jou09, Lou21, MR17, Mil93, Nor15, Qi06, SvE21]. Homogenized [OT01, GOT05]. Homological [MV13, AAFR09, CO94, CO96, Lam91]. Homology [BKRW17, AKL17, EL12, GDRV21, Hal13, HSV08, JZ04, RR12]. Homomorphic [BC91]. Homomorphism [GHS01]. Homomorphisms [But85, LGPS91, TW01, MRD11, SS88]. homotopical [BW03]. Homotopies [Ver00, HL16, LY18]. Homotopy [BCvdHS11, GMP13, aZGS05, DEPS11, HESV21, RRS19]. Hong [KMR18, Col16, MR10]. honour [GW11, GP12]. Hope [BCE^{+}94]. Hopf [BK99b, EW00, FGS09a, GES05]. Horn [Gal87, KR91]. HR [Col05]. hulls [JZ04]. human [Sad17]. Hurwitz [Cou00, Stu17].

Hybrid [DE02, AG91, MPH17, MH06, PNM13]. hydrodynamic [CJ90]. HYP [Kra95]. hyperbolic [BGH^{+}04, HLN^{+}21, PV13]. Hyperbolicity [Fit89, Sto20, Mer10]. hypercircles [Tab13]. hyperedge [PS21]. hyperelliptic [Ber98b, Eid23, Har12a, HJ15, LLL08]. Hypergeometric [Ae02, BP99b, BK99a, Chu99, Koe95, Pet92, Zei95, AP04, ABvHP11, ABPS21, CCF^{+}15, CK19, GHS08, IvH17, KY15, Kra95, Liu19, Mat01b, OS24, Rie03, WK20, Zha03]. hypergraphs [Eit94]. hypermetric [DS18a]. hyperplanes [Tab13]. Hyperresolution [GHS03]. hypersurface [Qi06]. Hypersurfaces [ASS97, BS00, ABR17a, BD16, BC05, Koh21, Lee17, Szp22, TCT23]. Hypothesis [von87]. HYPQ [Kra95].

I/O [MMW11]. IB [Sid93]. Ideal [AHLM99, BGG13, BRM01, BW87, CFM96, HKL99, Laz85, Laz92a, Mat01a, Mor99, Pan89, Pri96, del95, Ahn08, BO04, BH23, BDM23, BGMSG07, BCLR13, BJS04, Buc06a, EH21, FGT09, FK11, GTLN17, GSW11, HS21a, HQS19, Hre06, KRK88, KN11, KW24, Kun18, LV14, MM06, MR88, Per04, PS21, TV18, Val11]. Ideal-specific [BGG13]. Ideal-Theoretic [Laz92a]. Ideals [ABKR00, ACOR00, AV00, Bah01, BMNB^{+}11, BLR99, CM97, Fer06a, Fer06b, FGT02, FH94, HS00, Kem02, KM99, KM01, LS00a, LS00b, MS00b, MSY00, Mon02a, Mus00, NY99, Pan89, Pau92a, SST18, SY96, AFT08, ACMB19, AC19, ATY08, AT08, AHS21, BO10, BM88, BT09, BE22, BMQS06, BL12, BL17, BR15, Cer18, Cer21, CR11, CS05c, DS16, DDL^{+}23, DL06, FES11, FGT05, GVHHUE05, GHY17, Gas21, GES05, Gol06, Gol08, HSS18, HHPS21, HOP06, HKPP09, HM09, HP91, HdC13, HdC16, HH04, JFMRS12, Joh15, Jou09, Kem16, KMH89, KW88, LL09, La 17, Lev21, MWZ16, MR17, MR13, MRV21, NT17, NOT18, NT21, NNPZN19, NY04, Pis04, Rou09, Sak21, ST20, Ste13, TN09, UCJ04, WO06, dAM17, GTZ88]. Idempotent [Dav94, GTLN21, HKSS17]. Idempotents [Kon95, BP24, OdR03]. Identifiability [HS21b, Ang18, Ang24, BS22b, CR98, MMS18]. Identifiable [MS14]. Identifying [BDE^{+}21, KT02]. Identities [BH02, Deu93, Ges95, PS95b, ABF09, CS98, GHS08, Kau07, Rad15, Rie03, Sil04, Smo21, Zha03]. Identity [AP93, HJM93, Mul01, HJM94, PP11, Shp14]. Igusa [Sto17]. Igusa-zeta [Sto17]. II [Boy93b, BCGY12, CdG09, CD87, Com98b, DLLP08b, FGS09b, Gri23, HM02a, HLSS15, HLS01a, Kno93, LW03b, LLW03, MS16, MP14, Min02, OP05, Ren92b, Sch17b, Wal02b]. III [BMS20, BC89, DLLP08c, Ren92c]. image [FMR04]. Images [BC91]. Imaginary [Gaá93a, GP96a, HPT02, JTdW18, Rol90, Bus09, Heu06, KT04, Kli90]. imbeddings [JV09]. immersions [MPT20]. Imperative [HCB96, MMW11]. imperfections [JSC13]. implement [Rau11]. Implementation [AK92, ABP96, AM99, GK96b, GKW98, JV09, MNJ94, MF96, NY04, Sho95, Web96, ABM^{+}23, HT91, HH13, HDPS11, JB04, Ree98, RRS19, Smo21]. Implementations [ZSY93]. implemented [BPT12, OdR09]. Implementing [Nie03, Sid93, Sim90b]. implicants [KT90b, KT94]. implicants/implicates [KT90b, KT94]. implicates [KT90b, KT94]. implications [GGdC23]. Implicit [ARS02, MC92, SS98a, BG05, DJO^{+}11, DMR12, DZ09, WRI09]. Implicitization [ABR17a, AS01, BD16, CGZ00, Doh09, FHL96, GC92, Gao03, GV97, Hon97a, Ore01, WC12, CCL05, Chi08, CTY10, HS98, PDS08, RS10, RS11b, Rue11, SS05]. Implicitizing [BC05, LC16, SGD97, Wan04]. implicitly [VL16]. implies [CO94]. Improved [Bro01a, CE96, Els17, GZ89, Jar13, Lec07, McC88, PS20, DJ92, Hre06, KS16, Tsu09]. Improvement [LPS93, Tho02, BPH07]. Improvements [BMS17]. Improving [Gen07, HHT18, MM06]. In-depth [BF20]. in-place [Cox22]. Ince [Che23]. incidence [STW18, LW03a]. incidences [SPZ10]. including [AJGVS09, Sut12]. inclusions [BBV22]. Incomplete [FD93]. Incorporating [ARS02, GHMA13]. Increase [CP00]. Increasing [Pel97]. Incremental [EC95, HAGW12, KT90b, KT94, KÁ21, MU04]. Indefinite [Man93b, Wan94a, Piq91, PS95c]. independence [ÇJM^{+}21, CS21, PSA23]. independent [KW88]. Index [Ano99b, Ano00a, Ano01a, Ano01b, Ano01f, Ano02b, Ano02c, Ano04o, Ano04p, Ano05f, Ano06, GPP93, Gaá95, GP96a, Gaá00, Gau09, LR01, CHSS05, DJO^{+}11, LN13, PR22]. Indexed [Wan94a, LN21]. indices [Abr17b, DMN17, Ung19, Wan18]. Indispensable [ATY08, CTV16]. individual [SS16]. Induced [BH00, AFdCS15, BW03]. Induction [ARS02, Bou97, KNZ91, Str01]. Inductionless [KNZ91]. Inductive [DR93, Fri89, GS22, Pad96, KS12c]. Inequalities [GV88, Str00, Vor92, Bro12, CL20, HJX16, HHPS21, IdW15, Oak13, Pet87]. Inequality [MG94b]. infeasibility [DLMM11]. Inference [BA85, CH95, Pau92b, KW10, MM10]. Inferencing [Bib85]. Infinite [BBB92, Bir98, CP00, Geb02, IZ96, OKK98, PV02, CX09, DF08, DFO13, DW18, Koh08]. infinitesimal [LR98]. infinity [AGS16, BW05, Bod04]. influence [GP12]. Information [Mee94, BD87]. INGRID [DBG89]. Inheritance [DT95, SAK89]. Inhomogeneous [BF91, BCR11, Ede13]. Initial [BM88, HRdWY22, CS05c, HH04, Lem03]. initials [FGLH^{+}23]. injection [MPT20]. injective [HM05]. Injectivity [LS94]. Injectors [Höf01]. inseparability [LW10, Ste05]. insertion [Vat12]. Insights [PdRAEC24]. Instability [EC87]. Installation [GM88]. Instantiation [dB89]. instanton [GS05]. Insurance [AST96]. Integer [CGG89, DSV01, GS02, HM97, KJ96, Lüb02, Pel97, Web96, BLS23, BP11, BSW21b, Bus09, GHLZ21, Har12b, HvdH18, MRH23, Rup04, Wan06]. Integers [Col02, Gem94, Jeb95, Rol86, BV03, BFH17, DF05, FL04, Jam11, JMV18, KY15, Nag11]. Integrability [Adl16, AMW12, AMDW16, GLLdR21]. Integrable [FM02, GZ90, Zha93, BJM17, GSZ85, LW12]. Integral [AF00, AML19, AZ90, Hal01, Mil87, Vas00, YNT94, van94, Ahn08, AGT13, BKSS12, BDLP22, CK90, FD18, IvH17, Mar19a, Mau87, Raa12, Sta18, de 98]. Integrality [DFdG15, Sto03]. Integrals [AB89, Car99, Köl85, Sch85, Bar07, GK21, KS19, KKM15, MCJ21, Oak13, Piq91]. Integrating [Ano01d, AJ01, CTR99]. Integration [Bad06, Bro90a, CS05a, Car99, Che85, Czi95, DTGV01, Jef97, KS19, Kno92, Kno93, KF01, LR90, LS02, Mul97, Tef02, Wan94a, Wol00a, BB11, BLL^{+}16, Bro07, GGAVRC13, MSY21, Qi23, Wol03]. integro [BC24a, RS19]. integro-differential [BC24a, RS19]. Interaction [BC01, Sch94]. Interactive [AGMT98, CC01, FT97, HL98, ST89b]. Interface [Sch96, HPRS11, YW87]. Interfaces [BT98, KM98, KS98, Sch94, LLTPT^{+}11]. intermediate [KN11]. internal [SP10]. interpolants [Far19]. Interpolating [Zip90]. Interpolation [GV96, MR02, MF96, Rob97, AGR16, BH21, BH23, Cox21a, DKS15, GLsL09, HG20, Hua23, JNSV17, KsL03, RS20, RSS10, Sau18].

Interpretation [AB01, BB93b, GSA^{+}12, Zan94]. interpretations [ZWM15]. Interpreter [Hag89b]. intersect [BFMS87]. Intersecting [Gla88b]. intersection [AH13, BEP13, BGMSG07, BE17, BM04, DEPS11, DLLP08a, DLLP08b, DLLP08c, FGVN06, JWC^{+}16, Rod15, Sop13]. Intersections [GS90, Lo98a, MT01, ABC22, BGM15, DLLP08c, FS16, Sta16]. Interval [CJK02, Mer01, PC98, PR22, Sek09]. IntHaar [GK21]. intractable [HYH04]. intransitive [Els12]. Introduce [Bos97]. introducing [Rei06]. Introduction [BFK02, CFG^{+}86, GIK10, MNJ94, Poh87b, Ren92a, GK12a]. Intuitionistic [CH85, CH86, Pau86]. invariance [AT08, DRN24]. Invariant [Cra91, DHS98, DBG89, GG99, Göb95, Hub19, Hul99, JML^{+}13, Kem96, Mor91, SW91b, Whi91b, Wor94, BM88, BDE^{+}16, Cla21, DSW09, DL06, FLE^{+}23, FF17, Fre13, GS22, HJS22, HdC13, HdC16, Kem16, Kin13, KZ10, LN21, Mer10, MAN^{+}10, MV15, PY05, QHL^{+}13, van93, vzGGZ21]. invariant-based [KZ10]. Invariants [BCE^{+}01, CP93, Cre01, Els12, LR23, MS00a, SW02, Ald23, AR06, APS12, Bay03, Bed07, BGLGM17, BR22, BP10a, BP10b, CG23, DZY22, DL88, DNS21, Els15, Els17, FGT15, Göb98, GDRV21, HLO23, HK07, Hub09a, KW10, Kem09, MS03a, RCK07]. Inverse [DR00, Dic92, JKP98, SS99, Tay02, FFP98, HOP06, Lee08, Pom11, WyW93]. Inverses [Sal94]. Inversion [AGRZ99, Kri85, von90c, KMY24, LSSW12]. Investigating [AG91, BENW06]. investigation [Bur03]. involution [BR13a]. Involutive [Ape95, Ape98, GHMA13, HSS02, AH05, CMR15, EW07, HOS23, RZ09, WZ12]. Involving [BFHT85, Köl85, Zip85, Bîl11, CJP22]. IPIA [KT94]. Irrationality [Bee01].

Irreducibility

[Kal85, Kal87, KS21, Mon92, Kal90].

Irreducible [FGT05, GR02, Let01, Pre06, Sho94, Ulm94, GR12, GPMS20, GU21, LMP19, MP89, PS89, Rou09, WK20]. Irregular [BCE^{+}01]. Isochronicity [HR12]. Isoclinism [OUI16]. isogeny [FG08]. Isogroups [CDF92]. Isolated [GLW99, Mou98, FGT15, LZ12, MMS23, NT21, Qur17]. Isolating [CLX^{+}24, XY02, MS11b, Moe05]. isolation [BS17a, BSSY18, BK12b, Bur16, CGY09, CGG12, CJ15, CWZ23, Col16, Col17, EMT21, HLXL18, MSW15, Str12, ST19]. isomers [LMM05]. Isometries [PS97b]. isometry [ACS13]. Isomorph [LT89]. Isomorphic [BP00, LSY07, MS09]. Isomorphism [BL85, Der13, O’B93, O’B94, CH03, GTLN16, MP14]. Isomorphisms [Miy01, Wur93]. isoptic [DP19]. Isotopic [AMT09, BCGY12, DMR12]. isotopies [BS21]. isotropic [ST24a]. isotropy [ASS20]. ISSAC [CL23, JKP12, EY21, ES13, NSW16, Yok17]. Issue [Ano99c, Ano00b, Ano01c, Ano01d, Ano01e, AJ01, BB93a, Bos01, BK90, Buc92, CH97a, CL00, EY21, FLB00, HSW97, Hon96, HL97, JKP12, KM98, KR97, MMY00, MNJ94, PZ92, PS95a, Smo98, WS98, BDT22, BKR19, BBKK15, CFMMP10, CL23, DM09, DKM23, DDM15, ES13, GSSST10, HHK^{+}23, JMPR04, Ker17, KASW05, MMO18, NSW16, SS18, Tra07b, Yok17]. issues [Kad13]. Iterated [For02, LM09, McC99, dC10]. iteration [BSSY18]. Iterations [Cap90, Hen90]. Iterative [Kri85, Izu16, MMS23, YYZ12]. Itself [Dav88]. Iwahori [NPP17]. IZIC [FKM95].

J [AP04, AK06, AP17, CS09, DHM11, FS13, Fer06a, HZ15, HJM94, Hil05b, HdC16, HP08, KT94, KL90, MHXD09, Nor95b, RS11b, Sag89]. J. [Kal90, KMR18]. Jacobi [BKSV20, ES98, Hon04, Sha90b, Ye17, Ye18]. Jacobian [DS16, FS16, FFP98, GR22, HI94]. Jacobians [HLSW16, Sha01]. Jacobson [LS11]. Janko [CH96]. January [Ano24e]. Jeffrey [Szi17]. jet [CS22]. Joachim [GP12]. joinable [AHL03]. jointly [SM18]. joints [KP91]. Jordan [Orl22]. Journal [Ano01b, Ano02c]. JSC [BDT22, Buc92, DKM23, MMO18, PZ92]. July [Ano23i, Ano24f]. Jumps [Mat01b]. June [Ano23h, Ano23k]. Jung [Bec09].

Kähler [HT91]. Kalman [SSV23]. Kaltofen [Vil11]. Kan [BH00, BLW03, CLW95]. KANT [DFK^{+}97b, DFK^{+}97a]. Karatsuba [CM96, Har12b]. Katz [DR00]. Katzman [Yes21]. Kazhdan [SS16]. KdV [GZ90, LL16, Zha93]. KdV-like [GZ90, Zha93]. Kedlaya [Har12a]. Keith [GW11]. Kernel [Mau00, MP89, Wen06]. Kernels [HNS95]. Key [BCE^{+}94, BB10]. Khovanskiı [Ric91]. kind [CJP22]. Kinetics [SME87]. Kirwan [Szi17]. KLEIN [BG01]. Knapsack [FNU16]. knot [IGT15, KPRT18]. knots [KPR10, Sto20]. knotted [EK21]. Knowledge [ABP96, DS96, KF01, LMP89]. Knowledge-Based [DS96]. Knuth [EHR91, KMN88, Sim91]. Kobayashi [Mer10]. Kodaira [MP89]. Kolberg [Rad15]. Kolchin [Lev21, Rei99]. Kolchin-type [Lev21]. Kovacic [UW96, Zha95]. KP [AFMS23]. Krandick [Col15]. Kronecker [BVW18, DFS11, Har09, VK21]. Krull [FD18]. Kruskal [Pue89, Wei94]. Krylov [HNRS21]. Kummer [Dab01, Sut12]. Kutta [PC98, FPT04, Sof94].

labeled [Vat06]. Labelled [GL92]. labellings [XXZZ23]. Lacunary [Mig00, Ave09, CGK^{+}21, GR11, Gre16].

Lagrange [BMS17, Col15, Ren04]. Lagrangian [Wan23]. Laguerre [FHR99]. Lamé [Chu99]. Landau [MG94b]. landscapes [BD17]. Language [BFK02, Hag89b, BGL14, BCP97]. Languages [Boy93c, DT95, FH86, AHH^{+}05, CKKM10]. Large [CH97b, CJMP97, DHK^{+}95, Kal87, KC01, BBK14, CHU19, FG08, STDD16]. Large-Expression [CJMP97]. Large-scale [KC01]. larger [BMNB^{+}11]. Largest [Boy93a, Mig92, AT08]. Lark [Sta89]. last [CG23, HKYY18]. latin [DW18, FMM07]. Lattice [FJN93, HS00, Adl16, BBC^{+}11, BL17, DHTY04, FW14, HM09, HdC13, HdC16, LFD19, LV14, MS03c, Pis04, Sch03a, Sch04, CCT11]. lattice-theoretical [BBC^{+}11]. Lattices [BCS97, PS97b, HM09, JY17, LMR94, SS03a]. Laurent [HdC16, Apé10, HdC13]. Lawrence [PCVT08]. Laws [Fit89, WBM99, BC22, PH11, Wol03]. Layer [ACGR01]. layers [Kin14]. Lazard [MH16, MPP19, Mul97]. Lazard/Rioboo/Trager [Mul97]. Lazy [BO99, MO98, vdH02]. LCR [BS22b]. leading [HJS22]. lean [OB03]. leanCoP [OB03]. Learning [HHK^{+}23, BHM^{+}23, HLO22, HLO23, KMM22, STW18]. Lee [MRG13]. Lefschetz [AS24]. Left [BH00, CLW95, SS94, VL90, BLW03, LGS90]. Left-Shift [SS94]. Legendre [Boy93c, PPR13]. Lehmer [Jeb95]. Leibniz [CGGO09, CIL07, IL09]. Lemma [CR88, MS15]. Lemmas [Fuc00a]. length [BKHG21, KY15]. lengths [Bor22, Far19, Lia13a]. Leopoldt [Kli90]. less [RDU03]. Lesson [Rob97]. Letter [BCE^{+}94]. Letterplace [LL09, LL13, LMZ23]. level [HR19, Ros05]. Levelwise [NÁS^{+}24]. lex [FRR06]. Lexicographic [BR06a, Dah22, Sak21]. Liaison [GNS17]. Libraries [FS95, MNJ94, GK18]. Library [RW94, LMRS11, MRH23]. Lickteig [Lec19]. Lie [AL88, BC06, BdlCRLS19, BJT22, CV00, CdG09, DFdG13, Dra03, DW22, EKP22, Eic10, EG21, FS98, FK89, GK96a, HNVL90, HRT01, Hub09a, KP13, Kol08, LR98, LH17, Roo13, dG09, dGPS09, vdP99]. Liénard [SH17b]. Life [AST96, HKP^{+}06]. lifted [HS17b]. Lifting [LM20, RS16, Sht88, BDLP22, EK11, GNP12, Leb15, MM88, MT20, Sto03]. liftings [BCF19]. Light [WKB86]. Light-Driven [WKB86]. like [BL98a, CH17, DTGV02, Fox18, GZ90, KS18, LS01, Van02, Zha93]. likelihood [ABB^{+}19, BR06b, CS21, Drt06, HR17, RT17]. Limit [LP90, AAKM21, BPD19, HY23b, SH17b]. Limited [RV03, Fas10]. Limits [CMV13, AFMS23, ZX20]. line [AGR16, Ave09, BDS^{+}18, BE17, FKM10, HG20, JS07, JS18, LSY07, Lun16]. Linear [AR13, AC01, BP99a, Bar99, BO99, BBF17, Bro92, Bro00, BEM97, CK99, CF89, Cla21, CSTU02, DHK^{+}95, Die92, DV00, ER95, FT95, Fre04, Gri90, Hen98, KST93, Kov86, LPY01, MM00, Mtn03, Nor99, Pet92, PV05, RS10, RS11b, Sak88, SR07, Sin91, SU93a, SU93b, Sit92, Sof96, Stu00, SD05, Tsa00, Ulm94, UYSA89, Van02, VRUW99, Wei88, Wei90, YNT92, ABvHP11, Abr17b, ABPS21, AMW12, Arr16, BP09a, BCE11, BE13, BELP13, BM19, BSW21a, BJ21, BC24a, Bay03, Ber98a, Ber98b, BF22, BC22, BD16, BLPR15, BD04, BSW21b, BVE21, BR12, BP09b, Cha14, CM10, CGG12, CvH04, CQ12, DS86, DFdG15, DJ07, Dum09, EPY98, EFRS06, FSW10a, FSW10b, Fox18, GKS12, GSZ85, GHL16, GHLZ21, GMW23, GSSV12, HJS18, HM23]. linear [HNE16, HKL24, IvH17, JPP23, Joh15, KS06, LM92, Lem03, LZS11, LST03, LW12, Lin18, LLW24, LNP^{+}21, Mas16, MPT20, MS14, MS04, Mul04, Nak16, Ngu09, PZ12, Rin13, Ros05, RR08, Rue11, Wan06, Wol03, WK20, XZ10, vdH07c, vdP05]. linear-algebraic [BC22]. linear-fractional [GKS12]. linear-recurrent [Fox18]. linearizability [LGM21]. linearization [FKT13]. linearized [Lee08, PWZ18]. Linearizing [HH94]. lines [BDS^{+}18, BLT22]. Lingua [LHK^{+}13]. linkage [HS06, PS18a]. linkages [GNS17, HLSS15, KRV19]. Links [KLZA12, MR87, CN19, KP15]. Liouville [AMDW16, SML91]. Liouvillian [APZ20, Che23, DS86, FSW10a, FSW10b, Kno92, Kno93, Sin91, SU93b, Ulm03, VRUW99]. liquid [ERSG05]. Lisp [MBPLRR10]. List [BL96, DS09, LO08, LO09, MRG17, BB10, DJ15, MRG13]. Liveness [CSS96]. LLL [Poh87a, Ura24]. LLL-Reduction [Poh87a]. Local [AK00, AS07, Alc12, AL10, BO10, DJ96, EMS00, Grä95, Hal01, Her94, Lun16, MC02, Mus00, Ous91, Pau01, Sed02, SJS06, Vor99, ACMB19, Alc08b, ÀL06, BR13c, CJL13, EK21, GNP12, HH07, HM05, JR06, NT17, Nak09, NN10, PP17, PT14, ST20, Sor22, Str16, TN09, WY11, dCR17]. Localization [CM97, Lou08, OTW00, SY96, BDLP22]. localizations [HL21a, HL21b]. locally [BS24]. Locate [NMM90]. Locating [AI90, BR06b]. Location [CS90]. Loci [GKR23, NMM90, BBV22]. locus [BHM^{+}23, BNT18, ELME23, HR17, TBS17]. LODE [AB09]. Loewy [Kin14]. Log [DJ96, hHL21, EFRS06, HZ19, Joh15, Sha90a]. Log-concavity [hHL21]. log-linear [EFRS06, Joh15]. Logarithm [VM14, Gau09]. Logarithmic [AH23, von90c, CJUE06, JPPSG09, KS19, LR90, NT21, Raa12]. Logarithms [Köl85]. Logic [AM89, AV96, Ano96, Bac94b, BB93b, BS87, CSS96, CRAB91, ET96, FT97, FD93, Frü96, FH86, LC96, LO96, LHD96, Lyn97, MGL00, McN92, Pas86, Sch91, YI94, dB89, ARS10, ABFS15, Ant05, BD04, DJS18, Le 89, LW10, Moz89, SA89, ST89a]. Logical [DL93, MMO94]. Logicographic [Nak06]. Logics [BN01, BF95, OS92]. Logspace [DK16]. Long [Jeb95, BP23, HZ04, HZ15]. Looking [Li10]. Loop [CP93, QHL^{+}13, GGL06, MAN^{+}10]. loops [KW10, NV07, RCK07, XL13]. losses [Lia13b]. Low [KST93, EKP22, FGS09b, Shp14]. low-degree [FGS09b]. low-dimensional [Shp14]. Lower [Bus09, WS23, Yap91, vzG13, GHL21a]. lubrication [CJ90]. Lucky [Grä93, Pau92a]. Lusztig [SS16]. Lyapunov [SLX^{+}13]. Lyndon [ES11]. Lyons [Gol01]. Lyubeznik [Sei02].

Macaulay [EK11, FD14, HPS22, MS11a, Pom11, ST20, Sti03]. Macaulay-type [EK11]. Macaulay2 [Ang18]. Macaulayness [DH16]. Macdonald [GG92]. Macdonald-Morris [GG92]. Machine [BHM^{+}23, HHK^{+}23, HLO22, HLO23]. Machine-learning [HLO22]. Machines [BA85, STA94]. MacPherson [Hel16]. MACSYMA [Mag89, PW85, TM85, Mil93, Mos12, Ous91, SR86]. MacWilliam [GS23]. Made [CF91a]. magic [XXZZ23]. Magma [Bos01, BCP97, CP97, Dor21, Key01]. Magnus [KLR93]. Main [Zha93]. Majewski [Van00]. Makes [Hre94]. making [Col05]. Management [AF96, CJMP97, LMP89, BD87, Sid93]. manifold [GV16]. Manifolds [GK94, ACS13, ASS20, EK21, PPR20, aZGS05]. Manin [Sch01]. manipulating [Kau06]. Manipulation [BB92, Boy93c, CD87, Hen90, Mil87, SJA01, SME87, Tri86, Wan91, Kra95]. Manipulator [DBG89]. Many [BF95, Sne98, ARS10, DV21, GKR23, KH23]. many-sorted [ARS10]. Many-valued [BF95]. Map [FOT00, FGS09a, KZ14, MP89].

MAPinsure [AST96]. MAPK [Lic21]. MAPLE [JKP98, PS95c, GHL^{+}00, AST96, Die92, GKLM91, Tef02, Vei97, ACGL04, AV11, AB05, BR09a, CFG^{+}86, Col05, CJ90, DM05, Fit89, GHC92, LMRS11, Pro00, SMB03, Sil04, Ste95, WyW93]. Mapping [Bah01, BBB92, Sod96, MS03a, dC10]. mappings [Win14]. Maps [AK00, YNT92, AB22, BBC^{+}17, Bur04, GDR05, Har13, aZGS05]. Marc [Sza08]. March [HdC16, RS11b, Ano23l, Ano24g]. March-April [Ano23l]. marked [BCLR13, Qi23]. Markoff [GJT13]. Markov [ATY08, AT08, DO06, DE06, HM09, Nor15, Rap06, RS16]. mass [GES05, JSC13]. Massey [BF20, CK12c]. Massey-Sakata [BF20]. Massively [BBRS24]. Matching [BKN87, Bür89, HK95, IZ96, Lav91, Nip91, PS93, RR90a, WKA94, CM21, DT23, DKM21, HYH04, Kut08, YY03]. mate [BvdE03]. materials [PNM13]. Mathematica [AU21, BG01, Fat92, HBN95, Kra95, Nie03, Sit97, ZD02, NP95, PS95b]. Mathematical [BTG02, CC01, DR86, FGT95, FGPGP14, FKM95, GKW98, KF01, Mon97, vdH15]. Mathematics [AGM97, BC01, Ber93, CH85, CH86, HL98, GGAVRC13, GKM05, Par08, SR07]. Mathscape [Bar07]. Matlab [Roq13]. Matrices [CZG02, DE02, DTGV02, EM99, EP02, GSST98, GS02, HLM95, Kon95, Kri85, LS95, Lüb02, SL92, Ste97, Vil95, dG01, AS23, AAFR09, ÁAF^{+}18, BR09a, BCL06, BLV06, BBCM13, Car15, CK04b, CK12c, Dey21, EM98, HNE16, HKS21, Hre06, JFMRS12, KK09, Koz23, Kud22, LLM^{+}13, LE22, LS12, LLW24, LWXZ23, LWX23, MM04a, MS03c, PS18b, WWWX23]. Matrix [BMT21, CFTY97, CW90, DSV01, FZ87, HM97, HRT01, Lin91a, LO99, Ma94, MF90, Mou98, MO95, Ost99, PW90, RT98, Vac17, Vac18, Zha93, AE05, BHLGO15, BLS23, BC89, CHU19, CH17, CL07, DF08, DF09, DFO13, DaZZ04, DPS17, DPS23, Ebe19, EP04, GHL21a, GK21, HvdH18, HYG24, JPS13, KD90, Lab90, LS11, LNP^{+}21, Mil93, Ros23, Vil11, WY11]. Matrix-F5 [Vac17]. matroids [HS21b]. Matthews [Van00]. Maximal [For87a, HLM95, MV15, BELT21, BC24b, CH04, FFP98, HRdWY22, HLSS15, Roo13, ST24a, Sut12, Sut13, Sut16]. Maximally [Bih15]. maximize [Loj13]. Maximum [UC98, ABB^{+}19, BLT22, BR06b, Cla22, Col17, CS21, HR17]. May [SK91, Ano23h, Ano23k]. MBase [KF01]. McLaughlin [HLM95]. mean [GKS12]. meaning [BW05]. Means [Vel00, BMQS06, DE03]. Measure [CMP87]. Measurements [AI90]. Measures [HLS01b]. Measuring [Sor22]. MEAT [LMR94]. MEAT-AXE [LMR94]. Mechanical [AM88b, CP93, DJS18, Hol85, KW10, PH87, Wan91, DJ15]. Mechanics [CJMP97, Cra91, Bar13]. mechanisms [GMS09]. Mechanized [ACGR01]. medial [CN19]. mediated [HRdWY22]. meeting [BR10]. meets [GSA^{+}12]. MEGA [BDT22, DKM23, MMO18]. Membership [Com98a, Com98b, Pri96, SS88]. Memories [MSV22]. Memory [AF96, CM96, CG02, GK96b, STA94]. meshing [BCGY12, DMR12]. Meta [Hag89b, vdH11]. Meta-circular [Hag89b]. Meta-expansion [vdH11]. MetateM [Fis96]. Method [AZ90, Bon96, CZ92, Eck87, FT95, Ges95, Hsi87, KFF88, LW01, NY99, ŌS94, San96, Sch98b, SGD97, Tef02, Tes99, Tra98, Tre92, Ver00, Wan93, YXL99, Zei91, Zha94, ASS07, AABdG21, AAFR09, BW22, Bec09, BGL14, Bîl11, BC22, CX09, CJL13, CJ15, CK12b, Col17, CGK09, DJO^{+}11, DJ92, GLY09, GVYZ09, GLLdR21, GG92, GX04, Hon04, HcL21, HY23b, Izu16, JPP23, KSW13b, KT90b, KT94, KM06, LLZ24, LN21, MMS23, Mao21, MPP19, PT14, Piq91, RS21a, Ric91, Ros05, Sag14, Sal08, Sau18, Sid93, Sza08, Toh10, Wan04, WS09, vdH10]. Methodologies [Bie85]. Methodology [LHD96, Sch91]. Methods [Ano01e, AM99, Boy93c, Bur01, CDO01, Cra91, EP02, EHR91, KT02, KR97, MGL00, MMY00, MMO94, Mou98, PC98, Pau85, SO89, SME87, Sof94, Zha90, BCLR13, BP11, BENW06, DGPP10, DDM15, Els17, FK09, GVHHUE05, GPS09, GGdR^{+}13, GS22, KP13, MM88, MP09, Wan06, dCR17]. Metric [UY15, DZY22, GS23, KP13, Lin18]. metrics [ACS13, ASS20]. MICC [GMMM17]. Microcomputer [Dün94]. microprocessors [VB03]. microstrip [AP90]. middle [Har12b]. Miller [LR15, Sim90b]. Million [BCE^{+}01]. Milne [BR09b]. Milnor [Bod04, DS18b, MS09, NT21]. Minimal [Ae02, Hel00, Kin13, LS98, Lub14, LM94a, MM09, Nor95d, OT01, Ous91, Sak88, ABPR21, AP04, BO04, Bed09, CKR04, CR19, GGMAMF19, HJS16, HKL24, Jam11, JMV23, JNSV17, Mor11, Par22, dCW09]. Minimally [ACOR00, BELT21, HJA17]. minimization [ES18, MH06]. Minimizing [Fie04, CS16, FS10, FS12, FS13, FDS13]. Minimum [Col01, DL88, Gün90, JP10, Toh10]. Minkowski [Fuk04]. Minors [Ma94, BC24b]. MinRank [FES13]. minus [WS09]. mirror [Hie16]. Mixed [BP99b, CLL17, EC95, GLW99, HM02b, Mil92a, Min02, Min03, CCF^{+}15, CK19, CL20, CLX^{+}24, FGG^{+}16, FW15, MRW17]. Mixtures [NMM90, KMM22]. Mizar [RST01]. ML [Hag89a, PMW93]. MLE [EFRS06]. mobile [KS04]. Modal [CRAB91, Kal11]. Model [ABP96, BPT12, Kal11, Pel97, Pel03b, BHLGO15, Bon05, Nor15]. Modeling [TL96, VGT90, BENW06, LZS11]. Modelling [Div91, FH86, PZ96]. Models [CZ92, CP00, Pic00, SLK11, AH23, BPT11, CR98, ÇJM^{+}21, CS21, DG20, EFRS06, GMN22, Har13, MM16, MS14, Pel03a, Pel03b]. Modification [Poh87a]. Modified [Bon96, FK11]. modpn [LMRS11]. Modular [Arn03, BCG10, CD87, Con90b, DV00, EPY98, GAO02, HLM95, IPS11, MU04, Mar02, Mic88, MM10, MF96, NY99, Ohl95, QW96, Sch90a, SW02, VH98, YNT94, Abb17, AH13, CL07, CvH04, Deu93, DFS11, GGdC23, JY17, Kin13, LSSW12, MP11a, MS09, PS89, Rau11, Ren04, Ryb90, SW97b, Wil90, Wur93]. Modularity [Mar96b, TRRK10]. Module [Cha00, BFH17, PS89]. Modules [GO90, Lev00, ŌS94, Rut92, Rut93, Smi00, dG01, ÀL06, Aue05, BW03, CJUE01, CJUE06, DNS21, EG07, Fer06a, Fer06b, GV03, GTLN16, HH07, KZ14, La 17, LS16b, Lia22, Mon05, NOT18, OTW00, OT01, OUI16, QR07, TW01, Wal05, ZW08]. moduli [BBV15, Gen22, HJ15, Lou21, Qi23]. modulo [AB22, BCvdHS11, Con93, Dah22, DMS21, Mar96a]. modulus [PNM13]. Moh [MSS21]. molecular [Bar07]. Molecules [OT87, LMM05]. Moment [LLM^{+}13, BBCM13, HI08]. Moments [PSZ91, KMM22]. Monad [HCB96]. monodromy [DS18b, GGG06, Kön17, dCR17]. Monogenic [BCRS89]. monoid [BLT22, OS04b]. monoidal [BDM^{+}16]. Monoids [MO88, OKK98, Wal02a, Wal02b, GGMAMF19, GMP13, MO85, Mos08, Sch17a]. Monomial [ASS13, EMS00, HKL99, La 17, LS00a, MSY00, Mus00, Püs02, BT09, BD15, BGM06, BGMSG07, BGLGM17, Bro12, Cer18, Cer21, CM21, DJ05, DDL^{+}23, Gas21, HQS19, KK17, NNPZN19, Pis04, Pre06, Rob09, Rou09, dC10, SSS23, Shi04, TV18]. monomials [ATY08, NWW21, Qi23]. monotone [BP23]. monotonic [ET96]. Monster [Nor01]. Moore [MGRR23].

MOPS [DES07]. Morales [AMDW16].

Morley [IKGT11]. Morphisms

[Smi93, DV21]. Morris [GG92]. Morse [CVY17, KKK17]. Moser [BP09a]. Moser- [BP09a]. Most [BE99b, PV02, DEP22, GR12]. Motion [PS18a, VV97, DW07, KLR93, LSS19]. Motions [KS18]. Motzkin [BBCF22]. Moufang [NV07]. Moving [CGZ00, D’A01, Olv03, SGD97, AP22, JWC^{+}16, LC16, Tab13, WC12]. MP [GKW98]. MR [HJM94, Kal90, KT94, KL90]. Much [Hre94]. Mulder [HZ15]. Mulders [HZ04]. Muller [Key01]. Multi [BGK96, CRAB91, DYA97, Mal00, YNT94, AHH^{+}05, BKW20, CKM09, DMN17, ES18, LP03, Ren04, dCW09, TZ21]. multi-branch-point [BKW20]. multi-computational [TZ21]. multi-homogeneous [ES18]. multi-indices [DMN17]. Multi-Modular [YNT94, Ren04]. Multi-Objective [DYA97]. multi-paradigm [AHH^{+}05]. Multi-parameter [Mal00]. multi-Schur [LP03]. multi-state [dCW09]. Multi-Threaded [BGK96]. multi-univariate [CKM09]. Multi-Valued [CRAB91]. Multibasic [BP99b]. Multidimensional [GV97, ABvHP11, AKL17, BBF17, UY15]. multigraded [IK13]. multigrid [PT14]. Multihomogeneous [DE03, EM12, JS07]. Multilinear [EMT21, Whi91a, CGK^{+}21]. multinomials [MO21]. MultInt [Tef02]. multiobjective [BP11]. multipersistence [GDRV21]. Multiple [BBV15, BLS17, DW07, LO96, Mag89, Pet00, San96, Sod96, Tef02, BLV16, BDE^{+}21, CM09, Gon17, JMV18, Lic21, PH11, Rie03, Szp22, XW20]. Multiple-Strategy [Sod96]. Multiplication [CM96, CW90, Pan94, Pon91, Van02, BCvdHS11, CE19, DNS21, Har09, HvdH18, HYG24, LS16a, NPP17, Roc11, vdH07d, vdHL13]. multiplications [CH17]. Multiplicative [Gre00a, SS96b, HL18, Zhe21]. Multiplicities [Bre86, GLLZ23, PS09, ST20]. multiplicity [Cox19, HMS17, HH09, HY21, LZ12]. multipliers [BLS23, EG21]. multiply [HKS21]. multiplying [DPS23]. multipoint [Har09]. MultiPolynomial [MC93]. Multipurpose [PW94]. multiscale [BGL14]. multisequences [KZ08]. Multiseries [SS99]. multistable [HTX15]. multisymmetric [GRW16]. Multivariable [Fit97, UYSA89, Lev07b]. Multivariate [BH21, BEZ23, Cla91, Cza89, DES07, GRS02, Hon98a, MT01, MF96, Pan94, SS92, Vir93, AV11, AGT13, AGR16, AIRR12, BES13, BS10, BH23, BGP09, CK03, CS98, FP09, GLsL09, GHLZ21, Gre16, HG20, KMYZ08, KY16, KLM^{+}21, Lec07, LWX23, MS21, PS09, PS20, Qi06, RS20, Ste05, Tsu09]. MuMath [Tri86]. Mumford [BGM06, HH04, MR15]. MuSimp [Maw88]. MXL [ACFP12]. My [Kal00, Buc06b].

narrow [JPPSG09]. Narrowing [MO98, NRS89, You89, Ant10]. Nash [PSA23]. Natural [MO88, Tak91, Bro03, FF17]. Naturally [CGGO09]. NatureTime [MRS96]. Near [BS17a, BM01, DLLP08a, DLLP08b, DLLP08c, vdH01, AGR95, BSSY18, Tun09]. Near-optimal [DLLP08a, DLLP08b, DLLP08c, BSSY18]. Near-rings [BM01]. near-separated [AGR95]. nearby [GHL21b]. Nearest [LW98, PS09, Sek11]. Nearly [Pan02, PT16, BFPT21]. necessary [GGG06, HP07, HP08]. Necklaces [Bry21]. Need [KMN88]. Negation [Fer96, Fer98]. Negative [Lug95, FT95, JMV18, MES19]. Negative-Conditional [WG94]. Neglected [Kad13]. neighborhoods [vzGM22]. Neighbour [Sch98b]. Neron [PP17]. Nested [Hon97a]. Nesting [BFHT85, OS24]. nets [FK09]. network [Lic21]. networked [SMB03]. networks [BDE^{+}21, GSS05, Gon17, KZ10, MJST22]. Neurons [BHH23]. NEWEUL [HS89]. Newton [ABF09, BSSY18, GV97, Gri23, HL16, HL98, Hil87, LLZ24, Ver00, vdH10]. Newtonian [KS97]. Nie [AGS18b]. Niederreiter [Gen07]. Nilpotence [Sim87]. Nilpotent [HNVL90, Lo98a, LRW97, dGN02, BL06b, CE19, DF08, EKP22, dG11]. nilsoliton [KP13]. Ninth [NSW85]. Ninth-order [NSW85]. No [CGZ00, AS01, HJM94, Kal90, KT94, KL90, Nor95b, Sag89, Tsa23]. nodal [FM24, HMXD07, MHXD09, MPSXD09]. Noether [GHL^{+}00, Rob09]. Noetherian [CMR19, CHKL22, DHS98, FD14, Per04]. nomial [Li10]. Nominal [SSSK18]. Non [ABY90, Baj86, BTBQM00, CS98, CdG09, DS00, ET96, FT95, GP96b, HLM95, KRW90, MS11a, Pas86, PZ12, Pri96, Raj06, Ric92a, Str19, Wid01, Wir12, Abo10, BJT22, BVE21, DMN17, GSZ85, Ger06, GHL21b, GTLN16, GMW23, HHPS21, HJ15, Kin13, Kin14, KH23, KKM15, LL09, Lab90, Lem03, LMZ23, Lev21, Lou21, Mad14, MES19, NW10, Sch19, Sor22]. Non-Algebraic [Ric92a, KKM15]. non-analytic [Lem03]. Non-associative [CdG09, Raj06, Wid01, Ger06]. non-autonomous [NW10]. Non-Cohen [MS11a]. Non-Commutative [Pri96, BTBQM00, CS98, KRW90, GTLN16, Kin14, LL09, Lab90, LMZ23, Sch19]. non-convexity [Sor22]. non-defectivity [Abo10]. Non-general [DS00]. non-homogeneous [HHPS21]. non-hyperelliptic [HJ15]. Non-linear [PZ12, BVE21, GSZ85]. non-modular [Kin13]. Non-monotonic [ET96]. Non-negative [FT95, MES19]. Non-Permutability [Wir12]. non-positive [DMN17]. Non-Principal [HLM95]. non-realizability [GMW23]. non-reflexive [Lev21]. Non-Solvability [Baj86]. non-special [KH23]. Non-Standard [Pas86, GP96b, Str19]. non-symmetric [Mad14]. non-trivial [GHL21b]. nonarchimedean [AGS18a]. Nonassociative [Jac97]. nonclassical [BN04]. Noncommutative [CM09, LPT20, BDM^{+}16, DE06, La 17]. Nonconservative [EC87]. nondegenerate [ELME23, Li04]. nonemptiness [Dum09]. nonexistence [EFRS06]. nonic [BP10b]. Nonlinear [CK99, GH97, Jir97, San96, Tra98, BGH^{+}04, CWZ23, HY23b, LLL13, LGM21, PH11, XZ10]. nonnegative [DIdW18, IdW15, Nie12, Pap23]. nonnegativity [MW23]. nonsingular [BLS23]. Nonsolvable [Mal87]. Norm [AK00, Yam94, BPZ06, Bus09, FMTT13, KS06, KT04, VK21]. Norm-Bounded [Yam94]. norm-Euclidean [Bus09, VK21]. norm-trace [FMTT13]. Normal [Ae02, AF00, BLV06, BC06, CD00, CK99, Drä01, DSV01, Egl96, ES20, Mar96b, San96, Sen02, Sne98, Van00, Vil95, Von98, tW91, vG90, AP04, BJM17, BLS23, BLL^{+}16, BD13, CGO88, CL07, GLLdR19, HR12, JY17, LS11, LLW24, Mon05, MTV21, Pol95, PRR18, Sch04, WY11, XWL23, YY03]. Normalisers [Gla88a]. Normaliz [BS15, BIS16]. Normalization [DHS98, GLS10, BDL^{+}13, LN21, Rob09, Ryb03]. Normalized [Mar96a]. Normalizers [Eic02, GS90, Hol91, Lo98a]. Normalizing [She97b]. Norms [Boy93a, Boy93b, BKV24]. Note [Ano03l, Czi95, Hon04, Lan92, Laz92a, Mul97, RS93, UW96, Ano12m, BR13a, HZ04, HZ15, HKL24]. Notebooks [Mon97, Sit97]. notice [AK06]. notion [HP91]. November [Ano23j]. Nowicki [Kho08]. NP [AvW94, KP99, Lan10]. NP-Complete [KP99]. NP-hard [AvW94]. Nullstellensatz [DLMM11, Dub93]. nullvalues [Pie24]. Number [AF00, BdS01, Bos01, Bru01, BW87, CE95, Ded97, Enc95, GPP93, Gen22, HM02a, LM89, Lim93, Lis95, Mee94, Poh97, Rol86, Rol90, Smi02, Yan99, ZSY93, dM99, AB22, AP11a, Ave09, BELT21, Bel04, BLT22, BFH17, BE17, Col05, DZY22, DL88, EK19, FMM07, Fie04, Har14, Heu06, JPPSG09, KY16, Kau07, KO17, Kli90, KC18, KR23, Kut19, LSSW12, MCJ21, MD24, PZ12, Rob04, SH17b, SvH19, Tsa16, Ura24, VV18]. number-theoretic [Har14]. Numbers [Arn95, CR88, Duv94, Eck87, Ges92, RS90, Str97, Abb17, AH13, AC19, Bas06, BEP13, Bod04, CJP22, DPS16, DEPS11, DGW19, GS05, HMXD07, MHXD09, MPSXD09, Píl07, Rio03, Ryb03, dCW09, Smo21, TCT23, dAM17]. Numerators [KT90a]. Numeric [EP02, KL98b, She97b, WS98, GLsL09, RZ09, WZ12, vdH07a]. Numerical [BL98a, BL98b, Hen90, HSS98, KR97, KL17b, Mro96, NS90, Pan02, SS05, Tra98, BB11, BSC12, CGY09, CWZ23, DZY22, DR23b, EH16, GS03, GHS08, HS17a, IMP17, KS06, MPH17, MS21, Roq13, Rup04, Wan06, Wan86]. Numerically [BL98a, DH16]. numerics [Str06]. Nyström [PC98].

O [MMW11]. Obituary [SSS^{+}11]. Object [KKL92, DW07, SLK11]. object-oriented [SLK11]. Objective [DYA97, FL04]. Objects [BL85, Kau06]. Observability [Sed02]. obstructions [EG07]. obtained [ERSG05]. obtaining [KLR93]. occasion [BDT22, DKM23, MMO18]. occurrence [BDE^{+}21]. Occurring [AB89]. October [Ano24h, Ano23m]. Odd [Dür89]. ODEs [CFRS23, CTR99, FG06, HR12, Man93a, MV15, NW10, NW11, VGW18]. offset [Pet10, SS09]. Offsets [ASS97, Far97, AJGVS09, AS07, Alc08a, Alc08b, Alc12]. OLDT [BB93b]. OLDT-Based [BB93b]. Omega [Ges97]. Omega-Termination [Ges97]. One [Bru01, CJ97, CKS99, SS96b, TA87, Vid99, AGS16, CFRS23, Cou22, Dah22, EGS23, GGSST10, GKO09, LLZ24, Lev21, LZ12, LV14, LM94b, MRG13, MRG17, NNvdPT15, PP17, vdPT15]. one-point [MRG13, MRG17]. One-sided [SS96b, TA87]. ones [MPT20]. online [LS16a]. Only [Bru01, KMN88]. onto [Poz19]. Open [ACGR01, BCE^{+}94, HDHX17, Kal00, BLR13, Dra05, FU17]. OpenMath [CC01]. operands [Mad14]. Operation [GK96b, GO90, McC88, Gas21, NOF10]. Operational [AHH^{+}05, AB99, Har92]. Operations [DJ96, Mon05, Wan96, LNP^{+}21, PWZ18]. operative [LBM98]. Operator [Tsa00, CKM09, CvH04, Fre13, HI08, MKF93, PRR18]. Operators [Ano01e, CSTU02, Dun99, Gri90, ŌS94, Pau86, Pro00, SML91, Van02, van97a, van97b, Ald23, BR13c, CKS16, CHKL22, Che18, Dra03, GHL16, GOT05, GSZ13, Ros05, SS03a, Tra06, Wan94b]. Optics [NSW85]. Optimal [CE85, Kem99, KM99, KM01, Laz88, Pan02, SW97a, Tab11, BS17a, BSSY18, BFPT21, BK12b, DLLP08a, DLLP08b, DLLP08c, EH16, PT16, Pap23]. Optimality [Boy92]. Optimally [WKB86]. Optimization [Wei97, BPH07, BM16b, DIdW18, DGLM^{+}24, GSA^{+}12, GGEZ12, HJX16, KLYZ12, Loj13, MW23]. Optimized [SPZ10, KT94]. optimized-IPIA [KT94]. Oracles [CO01, EFG16, JZ04]. orbifolds [Qur17, Tra06]. Orbit [AB89, Eic02]. Orbit-stabilizer [Eic02]. Orbits [Hel00, Hel96, HY23a, JT03, dG11]. Order [BZ03, BE99b, CH85, CH86, Dom92, Fit97, HS95, Hen98, HKK98, Hre94, Hsi87, Kov86, Lug95, Man93a, MGS89, NNN98, Pau85, QW96, Ren92a, Ren92b, RZAG99, RP89, SV92, Sim90a, SU93a, SU93b, Smo98, Sof96, Tre92, Ulm94, VRUW99, Vid99, Wer98, Adl16, Ald23, AC04, Ano00b, Arr16, AB01, BCE11, BE13, BELP13, CP00, CFS22, CTR99, CK12a, DS18b, Dra03, FMM07, FPT04, FG06, FSW10a, FHR99, GS89, GOP18, HYH04, HW24, IvH17, LL16, MW23, Mil93, MV15, NV07, Nak16, NSW85, NW10, NNvdPT15, OS92, PSS12, Ren92c, ST89a, SSS05, Sla07, SG89, Sto03, SD05, Ulm03, VGW18, WK91, WK20, Wur93, ZL12, vdPT15]. Order-Sorted [Dom92, HS95, HKK98, Smo98, Wer98]. Ordered [Cow92, DS97, CMR19, Pel03b]. Ordering [BP85, MS00a, NR95, Rus87, BLM10, CMR15, CS09, Hre06, PdRAEC24, Smi05, FGLM93]. Orderings [HKL99, Les92, Lev07b, LXZZ23, LLW03]. orderly [Lem03]. Orders [For87a, PZ96, Smi02, Bor22, BGG13, CC24, DEP22, MZ05, Sut12, Sut13, Sut16, Tra07a]. ordinal [ZWM15]. Ordinary [Bro92, Bro00, Gri90, Sch85, AB09, Abr17b, ABPS21, BD12, DP09, FGH08, GLY09, LGM21, MO21, Mil93, Nak16, RR08, ZG09]. Ore [BCL06, CKS16, CS98, GTLN16, GTLN19, HL21a, HL21b, Jar13, LZS11, LS12]. orientable [FGPT03]. Orientational [PSZ91]. Oriented [KKL92, Lyn97, PS89, SLK11]. Origami [IT10, IKGT11, IGT15]. Origin [vKT93]. Orthogonal [Ano99c, FHR99, KM91, RT89, RT98, Sch99, DES07, FKTS12, GS03, Wan23]. orthostochastic [Dey21]. oscillations [LLL13]. Oscillatory [San96]. osculants [Bry21]. O’Sullivan [MRG13]. other [FS23, LGS90]. Otter [Col05]. Outer [You89]. Output [CL07, MPT20, MZ05]. Output-sensitive [CL07]. Oval [Key01]. over-determined [Sza08]. overdetermined [AHS18, KT08]. Overview [Wan96, BD16].

P [HZ19, Lan10]. P-recursive [HZ19]. Package [AST96, BG01, CH91a, MC97, NP95, Pro00, Sch96, Ste95, Tef02, Vei97, AU21, GK21, Kau06, OdR09, Ric92b, Rie03, Sil04, GHL^{+}00]. packages [Kra95]. PACLIB [HNS95]. Padé [CK90, FF92, LOOR^{+}03, RS21b]. Pages [Ano23i, Ano23l, Ano23k, Ano23j, Ano23m]. Pair [BD88, MCJ21, Sto99]. Pair/Completion [Buc87]. pairing [LLL08, LR15]. Pairs [Bec93, GAO02, BM16a, BLM22, BKHG21, CKR04, DW21]. Pairwise [MM00]. papers [Kap06]. Para [Sch96]. Para-Functional [Sch96]. Parabolic [Hel00, KT08]. paradigm [AHH^{+}05]. Parallel [AT96, BA85, BDL^{+}13, CM96, CM17b, CG02, DD90, Hon96, Kal85, Lim93, MM04a, MF96, MG88, PS95c, Pon91, RR90a, STA94, Sch96, Tak89, Tor93, Wan96, Web96, ZSY93, BBRS24, BS22b, Bro07, HM21, Kal90, KC09, Ree98, Ren04, SMB03]. Parallelism [LHD96]. Parallelization [ABM^{+}23, IPS11, KC01, GGL06]. Parallelizing [HCB96]. Parallelogram [DDD95]. Parameter [DTGV01, Mu08, AR13, GG92, Mal00, Pet10]. parameterization [DLLP08a, DLLP08b, DLLP08c, HS98, WC12]. parameterizations [CK12d]. Parameterized [BPZ06, Hem02, Arr16, KZ10, Tak93]. Parameterizing [AJGVS09, DLLP08c]. Parameters [Kem99, KM00a, Mon02b, Arr16, CHM12, DHM11, MW10, NT17]. Parametric [ACOR00, AHKY09, ASS97, DH17, DTGV01, FHL96, GC92, GC93, GV97, HSW97, MC92, MMO94, Ore01, Sit92, Yan99, AHW05, BDE^{+}21, Bry21, CES23, CCL05, CJL13, FL11, Gao03, GKS12, KSW13a, KLM^{+}21, KRTZ23, LR07, LE22, LJ09, NT17, PDS03, PD07, PT98, Wal03, WWWX23]. parametrizable [NW10]. Parametrization [LSW01, Sch92, Sch00, SW91a, SW97a, ARST09, BD15, CR19, GSHPBS12, RSV09, Sch98a, SPZ10, SJS06]. Parametrizations [PP97, Sen02, van97c, FL21, Sch03a, SS11]. parametrize [RSS13]. Parametrized [BCR15, GS07a, Lou08]. Parametrizing [dGPS09, LSY07]. Paramodulation [NR97, PP91a]. Parser [Mer01]. Part [CJ97, BBRS24, CFSGL21, KR23, KW24, LR90, Nag21a, Raa12, VM14, BMS20, Com98a, Com98b, DV21, HLS01b, HLS01a, Kno92, Kno93, MS15, MS16, Ren92a, Ren92b, Ren92c, Wal02a, Wal02b]. Partial [CH91b, FK04, LS01, SS09, Sch99, Wol02, dv96, BVE21, CQ12, DW18, FGLH^{+}23, GHL16, Liu19, PPR20, PH11, vdH07c]. Partially [JM93, Wra88, Min06]. particle [Loj13]. particular [Oll88]. Partition [Mul90, AU21, Hem18, LFD19, MOP15, RS21a, Smo21]. Partitioning [Gün90]. Partitions [GZ89, Leo91]. Past [Cav86]. Patching [Cou00]. Path [Bac94b, Rus87, KS04, Lea06]. paths [BBCF22]. Pattern [Lav91, PS93, RR90a, WKA94, BBCF22, KMY24]. pattern-avoiding [BBCF22, KMY24]. Patterns [MS00a, BP23, CKKM10]. payoff [GKS12]. PBW [NOT18, RR05, dG01]. PBW-Type [dG01]. PDAE [WRI09]. PDE [BC22, Lem03]. PDEs [AC01, BGH^{+}04, KT08, LST03, Wol00a, Wol03]. Peakword [LMR94]. Peano [ZWM15]. Pellet [BSSY18]. Pellikaan [Toh10]. Pencils [Sch92, DLLP08b]. Perfect [BRM01, GH05a, JWW23, JM95, GR12, GR11, MBC^{+}10]. Performance [CM96, HLS01b]. Period [GSST98]. Periodic [Sad16, JT03, LH98, LLL13]. periodics [GKR23]. Permanental [LS00b]. Permutability [Wir12]. Permutation [BL96, BC91, CH97b, CCH97, CCH01, CF94, CFTY97, Göb95, GO90, Hol91, Leo91, LRW97, Mal87, Sim90a, Tri86, BP98, BC89, CHU19, CJ22, CF91b, Els17, Göb98, HdC13, HdC16, HV16, Hul05, JPW19, JWW23, Kan91, Koh08, Min98, Poz24, RDU03, Vat12]. Permutation-Invariant [Göb95]. Permutations [BP23, Top14, Vat06].

Permutative [SS89b, Fer98]. persistence [BD17]. persistent [AKL17, BKRW17]. personal [Mos12]. perspective [BBC^{+}11, LO08]. Perspectives [Gre00b]. Perturbation [CJMP97, KC01, Rat02, TBS17, Lam91, AG91]. perturbation-Galerkin [AG91]. perturbations [EGB12, vzGMS10]. Perturbed [VV97, BM19, PY05, Rue11]. Peter [And95]. Pethö [TU18]. Petkovsek [CS05b, PS95d]. Petri [FK09]. Pezzo [GSHPBS12, dGPS09]. Pfaffian [BJM17, MM16]. Phase [Fit89, GKLM91, WK91]. Phat [BKRW17]. PhD [Buc06a, Buc06b]. Phenomena [GKLM91]. Philippe [SSS^{+}11]. Phylogenetic [AR06, CFSGL21]. phylogenetics [HS21b]. Physics [CF91a, Vor89]. PID [GTLN16]. Piecewise [Von98, Sek11]. Pisot [VV18]. place [AGS16, Cox22]. placement [Pal13]. plain [BLV16]. Plan [Rob97]. Planar [LP02, Mul90, ZS01, CC24, CVY17, GMS09, NW11, PS18a]. Plane [BE02, GSST98, Ric92a, Sen02, Alc08a, BD15, BGI18, BE17, BBN18, BS21, CGL07, CWL08, CP10, DS18b, Els15, FFP98, GGG06, Har13, HH09, HH13, HMXD07, KH23, MP89, Mic13, MHXD09, NS90, PD07, PV13, SS09, Zen06]. planes [LC16]. Planning [LBM98, DW07, KS04, PS18a]. Plans [HLS01b, HLS01a]. Plates [Eis90, Roq13]. Player [BCGR92]. Plus [WS09]. Pochhammer [GKsL03]. Pohlig [Tes99]. Pohlig-Hellman [Tes99]. Poincaré [CIL07]. Point [BdS01, CE85, CS90, Cuy97, EPW90, MV15, Vid99, WKA94, BKW20, BPD19, DHTY04, EGS23, GS12, Hub09b, MRG13, MRG17, MV10, Ros05]. pointer [LCQ^{+}10]. Points [ABKR00, AH01, AS01, CGZ00, HI98, MO95, SF90, AFT08, AB09, AHW05, ACMB19, AAKM21, BE17, CWL08, CS05c, Fas10, FLE^{+}23, GP20, HHS23, KH23, SMJ19, SS24, Szp22, dCR17].

Poisson [ASJ97, DD90, SJA01].

Polarimetry [BG01]. pole [DS18b, Pal13]. poles [GLLZ23]. policies [LMA11]. Polly [CCT11]. poly [HLXL18]. poly-powers [HLXL18]. Pólya [CPR09, CPR11]. PolyBoRi [BD09]. Polycyclic [Eic02, Geb02, Lo98b, LO99, MR98, Ost99, Sim90b, AE05, CELG04, NU18, SH17a]. polycyclic-by-finite [SH17a]. Polygon [Lis95, Sch03a]. polygonal [GMF13, HC12, IGT15, PS18a]. Polygons [Gün90, WW94, BK13, Hil87]. Polyhedra [AC01, Aur87]. Polyhedral [CS89, EFRS06, BKV24, IS10, LY18]. polylogarithms [DMN17]. polymake [GP24]. Polynome [MZM87]. Polynomial [AP10, ACOR00, AM88a, AM99, Bah01, BZ85, Bea92, BTW93, BF22, Boy93a, Boy93b, CMP87, CGG89, Chi96, CM97, Col01, CSTU02, CKS99, Ded97, Dic92, FGT02, GLW99, GH02, Gie98, GGdC23, GV88, HLN^{+}21, HMZ21, HS97, Kal94, Kal98, Klü99, Klü00, KL89, KL90, Kri85, LM89, LS00a, Ley01, MM16, McC97, Mig92, MG94a, Mil96, MT01, MF96, Pan89, Pan94, Pan96, Pet92, Pri96, RS93, Rob04, Roj99, Sad17, SML91, SS92, Sau93, Sch06, Sch85, Sed02, SL92, SY96, SvE21, Sho95, Str00, Tun02, UC98, Ver00, Vil95, Wan93, Wan96, Wan98, Wan99, YNT94, Zha96, AB09, AB22, AHS18, AC19, AHKY09, AGR16, Ave09, BV03, BC24a, BNN17, BMS17, BS24, BLV06, BCG10, Bih15, BP11, BPH07, BC22, BU09, BVE21, Bry21]. polynomial [BM16b, Buc06a, BR06b, BK16, CES23, CJP22, CMR19, CM12, CL20, CLX^{+}24, CL07, CGG12, CJ15, CG06, CK03, CK04a, CGK09, DIdW18, DHS22, ELME23, ES18, EYZ21, EMT20, EMT21, FGS09a, FGS09b, FS10, FS13, FG06, FGT05, FD18, GLY09, GVYZ09, GLLdR19, GTZ88, Göb98, GPGO16, GTLN19, Guo20, GP20, HR12, HJS16, Har09, HSS18, HHPS21, HS21a, HL17, HOP06, HKPP09, HR19, HJS13, HV16, HM21, HcL21, Hua23, HYG24, Hub19, IK21, Jam11, Jar13, JLR03, JP10, JPP23, KLYZ12, KRK88, KSW13a, KW88, KW24, LL13, LR07, LE22, Lec07, Li04, LMRS11, LJ09, LWXZ23, LWX23, LNP^{+}21, MM04a, MRW17, MR13, MW12, MS11b, MS21, MPT20, Mtn03, MP11b, MT20, MW10, MD24, MP09, MTV21, MS03c, Mur23, Nag11, NP20, NY04, Oak13, PT16]. polynomial [PZ12, Qi06, RZ09, Roc11, RCK07, Sek11, SLX^{+}13, SvE14, SH17b, The06, TBS17, TUÖ05, Tun09, WWWX23, WY11, WZ12, XLY15, XWL23, YYZ12, Zhe21, vdHL13, vzGMS10, Laz09]. Polynomial-division-based [BF22]. Polynomial-Time [Chi96, YNT94, AM88a, HLN^{+}21, MM16]. polynomial-transcendental [MW12]. Polynomials [Ano99c, BS90a, Baj86, Bea92, BL98a, BL98b, BCRS89, Can90, CE19, Cla91, CE96, CR90, DE02, DTGV01, DTGV02, Enc95, For02, FHR99, Gao01, GK00, Göb95, Hon98a, HM02b, Kal87, KT90a, KL98a, KM00a, Koh92, LSW01, Led00b, Lev99, Lev00, Ley01, LW98, Mv90, Mal87, Mal00, MZM87, Mig00, Min02, Mon92, NG93, Pan02, PS00, Pau01, Pon91, Pro00, RR90b, RZAG99, Sch99, Sho94, SW91b, Tho02, Vei97, Vir93, VH98, Vor92, WS98, Yam94, Yan98, Yan99, YNT94, Zip90, von90a, von90b, vzGP01, ABPR21, AV11, AGT13, AGR95, Ang15, AIRR12, BES13, BS10, BCL06, BDM23, BKSV20, BU14, BEZ23, BR09b, BD09, BD13, BMT21, CL17, CGK^{+}21, CKM09, DSW09, Dah22, DGS10, DP24, DNS21, DIdW18, DES07, DMS21, Eid23, EGB12]. polynomials [FDS13, Fas10, FP09, FES11, FW15, Fer06a, Fer06b, FKTS12, FKT13, GVHHUE05, GR12, Gal13a, Gal13b, GKL04, Gen07, Ger19, Ger06, GKsL03, GLsL09, GHL21a, GHLZ21, GS03, GPMS20, GTLN16, GRW16, Gre16, GSW11, GGEZ12, GK16, Gri23, GX04, GNP12, GU21, HL04, HHK17, HY21, HG20, IdW15, Jar13, JTdW18, KL19, KL21, KMYZ08, KLM^{+}21, KP15, Lee08, Lev07b, Lev21, LSS19, Lou08, MES19, MMS18, Mic13, Min03, Min06, MRV21, Nag21b, Nie12, OP05, Pap23, Poh05, PS20, PWZ18, Rup04, SM16, Sch19, Sch05, Sek09, SvE21, Sut15, TCT23, Tsu09, Wei13, WK91, WY20, YYZ12, ZWM15, ZG09, ZW08, vzG13, vzGGZ21]. Polyominoes [DDD95]. polytope [DS18a, Sal08]. polytopes [BM88, DHTY04, EFG16, Fuk04, Kuo06, Lun16]. Polytopical [Aur87]. polytropes [BEZ23]. polyzetas [BDM17]. Pommaret [AFdCS15]. Pose [VGT90]. Position [Kal01b, NS90, CJL13, CJ15]. Positive [ARE02, CES23, HLXL18, Kem02, Lip93, Lug95, Mat01a, WG94, BBRS24, Bih15, DMN17, HH16, JP10, MWZ16, Ste05, vdPT15]. positive-dimensional [MWZ16]. Positive/Negative [WG94]. Positive/Negative-Conditional [WG94]. positively [EKP22, GS22]. Positiveness [Hon98a, BS10, KMR18, MR10, PS20]. positivity [GRW16, Lun16]. possible [BP98]. PoSSo [AF96]. postulation [BDS^{+}18]. Potential [Kid02, HBN95]. Power [Ave86, Bec90, Bec93, BCGR92, CK90, Kal02, Koe92, Ous91, Sne98, van97b, ABK15, Ber98a, BIS16, GKsL03, Hir89, LS16a, MRW17, MJK17, NNN98, Shp14, SK12, vdH07c]. power-trigonometric [MJK17]. Powering [Pon91]. Powers [Gas21, Köl85, Ryb01, DDL^{+}23, GMKP21, GR11, HLXL18]. Practical [LO99, MP14, Ost99, PV00, BHLGO15, BD16, DET09]. practice [LMS09]. Pre [AB99]. Pre-defined [AB99]. precise [Mil93]. precision [BB11, Fas10, Lia13b, MRH23, Vac17]. Predicate [Pau85]. predicates [CCG06]. Predicting [Kal02]. predictions [Hie16]. predictor [HL16]. predictor-corrector [HL16]. Preface [BZ03, CGR05, CLM09, GK12b, Gao18, GP03, HM06, Kap06, KG03, Vil02]. Prefix [Mil92a, JKKK20]. prefix-assignment [JKKK20]. Preliminaries [Ren92a, Ren92b]. Prelle [Man93a, MM97]. preprocessing [AV11]. prescribed [Far19]. presence [LPR17]. presentation [EG21]. Presentations [HKK98, HH99, Le 86, MO88, AE05, CF91b]. Presented [GK96a, Let01, Lin91a, NÓ89, OKK98, PS97a, CdG09, Lab90, MO85]. Preserving [BH21, Egl96, PG86, Hul13]. PRESS [SBB^{+}89]. Preunification [QW96]. Primality [CO01]. Primary [HS00, Laz85, MS00b, Mon02a, Rut92, Rut93, Sau01, SY96, Ste05, CHKL22, Dur09, GTZ88, KN11, PS21]. Prime [AF00, BL98a, BL98b, Bru01, Kal94, KMN88, NNN98, Ulm94, del95, AB22, FSW10a, GS12, Gol06, KT90b, KT94, Lev21, Mar19a, NY04, Sut15, Wan18]. Prime-power [NNN98]. Primes [BR87, BCE^{+}01, Grä93, DiP16, DEP22, Jam11]. Primitive [BL96, Gue18, Mal87, YNT89, Ang15, BP24, HY23a, OdR03, RDU03]. Principal [AHLM99, BW87, HLM95, Pan89, EH21, Koz23]. Priori [Bea92]. Priority [Lav91]. PRIZ [MT88]. Probabilistic [BLV16, HJS16, HJS22, Sed02, RT17]. probabilities [Mor13]. probability [BP07, CM09, DG20, FW14]. Problem [BL93, BL85, Der13, DR00, EHR91, GV96, HI94, JM93, KFK97, KP99, MSKO93, OZ94, PH87, PZ96, PY94, Pri96, Ren92a, Ren92b, Sta89, Tak92, Tay02, Wid01, Wra88, YP91, Yap91, AP10, AHKY09, AvW94, CC24, CJ90, DKS15, DMW17, FES13, FU17, GGG06, Gau09, GGM10, GV16, KD90, LSSW12, MO85, MPH17, Min98, MH06, NU18, Oll88, Sch17a, SS90, Sha12, TU18, VM14]. Problems [AGRZ99, BKN87, CE85, CS89, CL89, ER95, Eic02, Fer96, HK95, HLS01a, Kal00, KC01, KR91, Laz92a, Meg90, Mer01, Mic90, Mil92b, OT87, San96, SO89, SME87, Stu00, TA87, Vel00, Wei88, Wei90, Zha90, ZBH96, BR06a, BR12, Bus09, Cla22, Con93, DLMM11, DGLM^{+}24, Fit85, FNU16, HJX16, HYH04, Kap86, LWX23, MM16, MW12, PY05, PT98, Pic03, QSGB19, RW90, Ros05, RR08, Rou08, UY15, Yah23]. Procedure [Buc87, CLW95, Fri89, Gan91, KMN88, Man93a, Pau92b, Sim91, BN04, GMF13, Ye17]. Procedures [ARS02, BS96, SV92, SS03a, VB03]. Proceedings [Bos01]. Process [DHK^{+}95]. Processes [DS96]. Processing [DD90, KL98b, Sod96, FMR04, JMPR04, Par04]. Processor [CM93]. Processors [ASJ97]. produced [CK04a]. producing [KN11]. Product [Hre94, Mig92, BMT21, CJ22, FMR04, GGdC23, HZ04, HZ15, Har12b, BLG12]. Products [Bac94a, Bea92, Shp14, BNRW22, Cha14, DK16, FW15, FNU16, OS24, PV05, Piq91, RS16]. profile [DPS17, JPS13]. profiles [DKLP21]. Program [AB00a, FD93, FLOR00, Mag89, ZD02, BSC12, JGF09, ST89a, Vis05]. programmatic [BKHG21]. Programming [AM89, AV96, Ant10, BFK02, BB93a, Bie85, BS87, Frü96, FH86, Gib87, LC96, LO96, Lyn97, SJG96, Sch96, Ant05, DIdW18, GKMW21, GKS12, GMW23, GL05, LR23, MW23, MT88, SA89, TM89]. Programs [BB93b, DR93, HCB96, LHD96, Nie94b, PMW93, Sch91, Sch93, Sod96, Tak91, Tra89, YI94, dB89, AGS18a, AGR16, BP11, HNE21, HG20, JS07, JS18, LCQ^{+}10, MMW11, Nal18, SJ12, Wan86, XZ10, SD05]. Progress [BB11]. progression [GJT13]. Progressions [CW90, BPZ06, Smo21]. Projection [Aur87, Bro01a, McC88, Wer12, CKM09, HJX16, Kun18, MH16, Mor11]. Projections [Ass94, BGI18, JTdW18, Str16]. Projective [BE02, Cha00, Con90b, DV00, GHL^{+}00, HJ18, Stu91, Whi91b, Alu03, BO04, BMNB^{+}11, BGM06, BGLGM17, BCF19, Brå24, DH16, Eng10, FGT15, GV03, Har17, Hel16, JWC^{+}16, JV09, LW03a, LW03b, SSS23, Stu17, Wib07, GHL^{+}00]. Projectively [SW91b]. Projectors [Höf01]. prompter [Izu16]. Prony [Sau18]. Proof [BC01, CK19, Dub93, Gol01, JL91, LBM98, Mao21, Mar19b, Zha90, Bec03, But88, CL20, Col15, DHKS07, FU17, GKMW21, GG92, GK18, IKGT11, KS12c, LW03a, LW03b, Liu19, MPP19, Sil04]. Proofs [AP93, Bon96, CH95, CO01, DT23, Fuc00b, GAO02, Tak91, Tre92, Bel03, HLN^{+}21, KKK^{+}16, Ley04, MU04]. Propagating [Gal16]. Proper [Sch00, BJT22, KY15]. Properness [Mor11]. Properties [Cha00, DCC95, Ede85, Kal98, LH98, Ohl95, ST89b, Zha92, AS24, BD04, DSW09, Ebe19, El 05, GIL88, JGF09, LLTPT^{+}11]. property [CO96]. Propositional [ZBH96]. Prospector [BLG12]. Protocol [GKW98]. Protocols [KM98, BD04, KKK^{+}16, LNP^{+}21]. provable [BGP09]. Prove [PP91a, Ye18]. Prover [BH95, ML92, ST89b, ZBH96, Win06]. Provers [BT98, Str01]. proves [CS98]. Providing [LLTPT^{+}11, GLLdR19]. Proving [AGMT98, Ano00b, AB00b, ACS13, BF95, Baj86, BZ03, Bon96, Bou97, FT97, GHS08, HJX16, Hsi87, LBM98, MR87, NR95, Pad96, PS95b, Pet00, Soc91, Wal02a, Wal02b, Zha94, AHL03, Bon05, DJ15, KS86, OB03, PSS12, PS95e, RV03, Rie03, Rus91, Sid93, Zha03]. Prüfer [CLQ10]. PSATO [ZBH96]. PSD [Koz23]. Pseudo [MO88, MD24]. Pseudo-Natural [MO88]. pseudo-polynomial [MD24]. pseudogroups [LR98]. Pseudonatural [MO85]. Pseudoprime [BB00]. Pseudoprimes [Arn95]. PSPCLink [ASJ97]. PTOPO [KRTZ23]. Public [BCE^{+}94]. Publisher [Ano03l, Ano12m]. Puiseux [CFS22, CFRS23, Gri23, Hil87, PR12]. pullback [Ald23]. pure [BL12, Par08, TZ21]. pursuit [YY03]. pushout [BGS11]. Putinar [ME21]. puzzle [RW90]. pyramid [BIS16]. Pythagorean [AK06, AK04, BL06b, FGS09a, FGS09b, FG16, Far19]. Pythagorean-hodograph [Far19]. Pythagoreanization [AK06, AK04].

qFunctions [AU21]. qMultiSum [Rie03]. QUAD [BGP09]. quadrant [FU17]. Quadratic [BdlCRLS19, Gaá93a, GP96a, HPT02, Rol90, Bus09, FDS13, FGG^{+}16, GV16, Heu06, Hub09b, KT04, Kop08, KC18, KR23, Kut19, LR23, Sad16, Sak21, Sta16, VK21]. quadri [Mad14]. quadri-algebras [Mad14]. Quadrics [CGZ00, AJGVS09, DLLP08a, DLLP08b, DLLP08c, GP13, HDPS11, JWC^{+}16, LC16, Sta16, UY15]. Quality [HH16]. quandles [HMN06]. Quantified [Rat02, SD05]. Quantifier [Arn88a, AM88b, CH97a, CM16, CH91b, DH88, DS97, DYA97, EW00, GV96, GGL06, HL97, HLS97, Jir97, Laz88, PS00, Ren92b, Ren92c, Wei97, XLY15, EH16, HE12, SD05]. Quantifier-free [DS97]. quantitative [CPR09]. Quantized [dG01]. Quantum [DJK05, KS16, KKK^{+}16, Sag88, Sag89]. Quarks [BH87]. Quartic [BLT22, CcK02, GPP93, PSV11, Bri06, Els15, GP24, PT98]. quartics [CP10, CF09a, Vil23]. Quasi [CS21, Hub09b, Les92, BDM^{+}16, CGGO09, GLLdR19, GLLdR21, Jou09, Lou21, Wan18]. quasi-filiform [CGGO09]. quasi-homogeneous [GLLdR19, GLLdR21, Jou09, Lou21]. Quasi-independence [CS21]. Quasi-Orderings [Les92]. quasi-prime [Wan18]. Quasi-quadratic [Hub09b]. quasi-symmetric [BDM^{+}16]. Quasigroup [ZBH96]. quasiseparable [PS18b]. quaternion [FDS13, FGG^{+}16, Kut19, FGS09a]. quaternionic [DGS10]. queries [CES23]. question [Mar19a]. Questions [AP93, KSD16]. Quillen [LS00a]. Quillen-Suslin [LS00a]. quintic [FS12, FDS13, Far19, Heu98]. Quintuple [Che92]. quiver [Kos07]. Quotient [HNVL90, Lo98b, Nie94a, Ple87, Sim90b, Mou05]. Quotients [LS00a, MBC^{+}10, CMV13, FC04, IvH17, Kem22, Smo21].

R [Tri86]. R. [Gol01]. Race [AK92]. Radar [BG01]. radial [KMH89]. radiation [AP90]. Radical [CcK02, Kem02, Mat01a, SS11, Gol08, LLM^{+}13, MWZ16, Sut12]. radicality [HT23]. Radicals [FGT02, Höf01, Kal94, Zip85, EYZ21, Gir21, Har13, JFMRS12, Ryb03]. radii [BR10]. Radu [Smo21]. RAM [PH87]. Ramanujan [Gar95, Hem18, MSZ09, Rad15, Sil04]. ramification [Abr17b]. Ramified [Bru01, HM02a]. Ramis [AMDW16]. Random [BS90a, Car15, CF94, CR90, DPS16, Mul01, AP22]. Randomized [KT02]. Range [SO89]. Rank [FZ87, JPS13, Mat01b, Orl22, Qi06, ABK15, BBV22, BGI11, BR13b, BDPR13, BR13c, CP10, DPS17, DKLP21, GHL21a, GS23, HNR24, LN13, Mos08, Nal18, Top14, WS23]. rank-2 [CP10]. rank-constrained [Nal18]. rank-metric [GS23]. Rank-profile [JPS13]. ranked [DE06]. ranking [GKO09]. Ranks [Sei20]. rate [HR11]. Rational [Ae02, AP04, ACOR00, AS01, Bar99, CE95, CGZ00, Czi95, DP24, DTGV01, DGW19, FS10, FS13, FGH08, FHL96, GC92, GK00, GO91, GO00, GS02, GLLZ23, GRS02, HH98, Hie16, HW24, HK07, Jef97, Kri85, LS01, Lim93, MS95, MM97, MC92, MF90, Mul97, MQS99, NW10, NW11, NY99, PP97, Pet10, PS95d, Sch98a, SGD97, WZ23, ZS01, van97a, van97c, Abb17, AP10, ABPS21, AK04, AK06, Alc08b, AHM18, AGR95, AE05, AF08, BC24a, BGI18, BD16, BC05, CW03, CCL05, CS05b, CWL08, CDWZ21, CJL13, CS16, CR19, CS21, DKS15, DHTY04, DHH^{+}04, Doh09, DDL^{+}23, FS12, FDS13, FG06, FM24, FG08, FL21, GLSV21, Gao03, GHLZ22, GSPB17, GEL05, Gri20, HJ18, JWG10, KLYZ12, Kön17, KPR10, KS21, KH23, LC16, LR90, MS03a, MS21, OdR03]. rational [Píl07, PS95c, RSV09, Sek11, SPD14, SJG13, Tab11, VW08, VGW18, VL16, Wan04, ZX20, BLPR15, OK08]. Rationalizability [BF21]. rationally [AJGVS09]. Ratliff [Eli04, Gas21]. rays [BS17b]. RC [SM18]. RC-constructibility [SM18]. RComp [NP95]. Reachability [LPY01, SJ12]. Reaction [GH02, Gon17, Lic21]. Reactions [SME87]. reactive [SJ12]. Real [Arn88a, AM88a, ARE02, CR88, DH88, EH16, FB93, GVGC99, HNE16, Mee94, RS97, Ris88, RS90, Sch00, Str12, Tra98, XY02, Yan99, AMT09, AS05, AABdG21, AAKM21, ADCZ21, BELT21, BS17a, BHM^{+}23, BE17, BJT22, BMS20, BS21, BVE21, BK12b, CMV13, CCG06, CLX^{+}24, CGY09, CJL13, CJ15, CWZ23, Col15, Col16, Col17, DH07, DFdG13, DdG21, EYZ21, El 08, EGS23, EPY98, FGPT03, FGL04, FGP05, FGT15, Gal13a, GP20, KPT15, KSD16, LLM^{+}13, MWZ16, MS15, MS16, MS11b, Mic13, PT16, PES24, Qi05, Ren17, Rio03, RSV09, SM16, Sek09, Sor22, ST19, VK21, WS23, Zen06]. real-root [BK12b]. RealAlgebraic [BX97]. realizability [GMW23]. Realization [KM00b, Led00a, Ous91]. Realizations [Nor95d]. Realized [Die92, Vei97]. Reals [Ren92a, Ren92b, Ren92c, DET09, LE22]. reason [Kap86]. Reasoning [ACGR01, CCM95, ET96, FGT95, FOT00, KRV19, LS02, BKR19, CS05a]. Recitations [Zei95]. recognisable [GIL88, HKL24]. Recognising [Aur87]. Recognition [BP00, CC91, Tak92, Bro03]. Recognize [Ric97]. Recognizing [DFO13, LSY07, VL16]. Reconstruction [ABY90, Bon96, CE95, GLSV21, GMKP21, GLLZ23, WW94, Abb17, AB05, HV22]. reconstructions [BNT18]. record [Sal08]. record-high [Sal08]. Recording [DS96]. Recovering [GIJ14]. recovery [KY16]. Rectangular [Chi01, GZ89, EM98]. Rectifying [Jef97]. Rectilinear [Gün90]. Recurrence [CM93, Tak95, BF22]. Recurrences [Nor99, Pet92, Fox18, JPP23, MZ05]. recurrent [Fox18, Mtn03]. Recurring [Sak88]. Recursion [Ful90, Pau86]. Recursive [HOS23, NP95, Rus87, BBF17, HZ19, hHL21, MS10]. Recursively [MG88, Tor93]. REDUCE [Sag89, BBB92, Fit85, LP90, Ng89, Nor95a, RT89, Sag88, SV92, dos89]. REDUCE-Procedures [SV92]. Reduced [FD14, GS98, Kon95, KM99, KM01, Lev00, Sny93, AMW12, EP10, MP04]. Reducibility [McC97, NOF10]. Reducible [Ber02, CS99, DW22, Ren17]. Reducing [AKL17, Gir21]. Reduction [CvHKK18, Che18, DCC95, GL05, GHC92, Kid02, Nau98, OS04b, Poh87a, RS93, Sti87, Tak89, Wol02, AAFR09, BP09a, BELP13, BSW21a, BCL06, Bon05, CQ12, Con93, DJO^{+}11, DFS11, HP91, HMZ21, JKKK20, LLW24, LY05, LWXZ23, MS03c, PR12, RZ09, Rei06, TV18, WZ23, WZ12, OK08]. Reduction-based [CvHKK18]. reductions [AB22, AH05, CMR19, ST20]. reductive [DH07]. Redundancy [BL00, Pic00, Tak91]. Redundancy-elimination [BL00]. Redundancy-free [Tak91]. redundant [KN11]. Reed [BB10, Key01, LO08]. Rees [BD15]. Refined [EMSS16, BSW21a, Sch08]. Refinement [Her94, PT16]. refinements [Mag17]. refiners [JPW19, JWW23]. Reflection [DR23a, KS18]. Reflexive [MAN^{+}10, Lev21]. Refutational [Pau92b, Wal02a, Wal02b]. Refutations [CZ92, MOP15]. region [BGG13]. Regions [Fit89, GMF13, Gon17]. regressions [Drt06]. Regular [AF88, Bac94b, BP99a, HY23a, MN02, MM04b, NS90, Nip91, Wan00, Yel87, vdH01, Alc08b, BCE11, Bel03, BLPM19, CM16, DW21, HKL24, Joh15, KR22, Mon05, NWW21, Poz15, SH17b, Vat12, Wal05, Wei06]. Regularity [CCD^{+}09, AHS21, BMNB^{+}11, BGM06, Dum09, HH04]. regularization [STVvR24]. regulatory [HTX15]. Rejection [LT89]. Related [Hes02, Wal02a, Wal02b, CG23, Cla22, FMM07, Kau06, RW90]. relating [MGRR23, Smo21]. Relation [Dün94, Soc91, ACFP12, Coo09, FK09, HKL24, KLV10]. relation-algebraic [FK09]. Relations [BdS01, Com98c, Ful90, LM90, Mor13, MR02, Sak88, Tak95, BE22, BF22, CI07, HR19, JML^{+}13, KZ08, Kos07, MAN^{+}10, OS04b, Rei06, Ye17, Ye18, Zhe21]. Relationship [KR89]. Relative [AV00, Bel04, Dön13, Els17, ZW08]. Relatively [BL98a, BL98b]. Relax [vdH02]. Relaxation [CDM^{+}13b, BM16b]. relaxations [MWZ16]. Relaxed [Leb15, vdH07d]. reliability [MdCW17, dCW09]. relocation [RW90]. remainder [CK90, Jar13]. Remark [IZ96]. Remarks [DS00, Hen90, SS98b, LOOR^{+}03, LWX23]. Removing [BJ21]. Renormalization [BK99b, Cap90]. reparametrization [Tab11]. Reparametrizations [RS97, MS14]. Replacement [BLG12]. Representation [CH96, Dra05, Drä01, HPS97, KH23, Lip93, MC92, Mic88, Mic90, NY99, OS24, PPR13, SML91, BFG07, Bro03, BDM17, CS05b, CGG12, CF09b, FL04, Kud22, LMM05, MCMMPR14, Mor20, PS95d, XWL23]. Representation-finite [Drä01]. Representations [Cla91, CFTY97, Die92, DV00, Drä01, DR86, FKTS12, Kal93, Let01, Lin91a, LO99, PW90, Pic00, PS97a, Püs02, Sch90a, dGN02, AABdG21, BG05, CHU19, DA05, Dey21, FGS09a, FS98, Fie04, HKL24, Kol08, Lab90, ME21, MM04b, Naw16, PV13, Pre06, Ryb90, BLPR15]. representatives [dG11]. Representing [KF01]. Research [Bos97, GP12]. Residuals [Höf01, Mur23]. Residue [MPS02, YNT92, ZSY93, Buc06a, FD18, Hul13, KKM15]. residues [Szi17]. residuosity [DF05]. resilient [PNM13]. Resolution [BF95, BL00, BS00, BS01, Gaá93b, GPP93, GP96a, Gaá02, OS92, PP91a, Rus91, Soc91, SS03a, AFdCS15, AHS21, DNV03, Pel03a, Pel03b, Pet87, RV03, dNdR03]. Resolution-based [BF95, SS03a]. Resolutions [LS98, Lam91, OT01, Sei02, AS24, Ell04, Gal16, Gal22, GS07b, HM05, dCW09]. Resolvents [AV00, Ren04]. resource [RV03]. Respect [PZ96, RZAG99, Wan94a, BK16, Lev07b]. respectively [Kra95]. restricted [GGEZ12, Oll88, Vat06]. Restriction [CC07, Fri89]. Restrictions [SO89, BMQS06]. Result [Bea92, Boy92, Pau92b, Bur03, KS16]. Resultant [BU99, BK16, CZG02, DE02, EC95, EP02, Gaá02, HM02b, MC93, Min02, CK04b, DE03, EM12, FC04, IMP17, Khe03, LYG15, Min03, MKF93, Pal13, PDS08, Rue11, SJG13]. resultant-based [PDS08]. Resultants [BEM00, CK04a, D’A01, HS06, HKSS17, McC99, Min06, BFSS06, BG05, BMT21, Chi01, CK03, EK11, Hil05a, Hil05b, JS07, JS18, OK08, QZYC22, RS10, RS11b, TUÖ05, YZ21]. Results [Lab95, Lug95, PP91b, AB92, PES24, Sch03b, TBS17]. Retention [Mil92b]. Retraction [AK06]. Retrieval [CE85]. retrieving [MCMMPR14]. Revealing [Gon17, JPS13]. Reverse [BT09, Sak21]. reversible [STDD16]. Review [Fat92]. Revised [BPT12, AP11b, AP17]. Revisited [NRS89, Pic98, Rus87, Sch90b, Tor93, Bec09, GS89, HY23b, IKGT11, Pom11, SG89].

Revisiting [CW03]. Revolution [CJ97]. Rewrite [Bir98, Com98a, Com98b, FJN93, Hsi87, Les92, CLS91, LA96, RV05]. rewrite-based [CLS91]. Rewriting [Ave86, BP85, Boo87, BH00, Der87b, Dur94, Ges97, GAO02, HKL99, HH94, Kah95, KKSd96, KM91, Lav91, Mar96a, Mar96b, Ohl95, OKK98, PY94, PP91b, Smo98, Sny93, Wer98, Wid01, You89, Zan95, Zha92, ABFS15, AR03, AB05, BGHW06, Bur03, Che18, Der87a, DS15, EW07, GL05, HK07, IT10, Kap87, NOF10, Wir09, Zan94]. Reynolds [ZGG23]. Reznick [ME21]. Rham [Wal00]. RIA [LLTPT^{+}11]. Riccati [BTG02, LS01]. Riccati-like [LS01]. Riemann [BCI13, Chu99, GSST98, Hes02, HI94]. Right [Gre00a, La 17]. rigid [BELT21, BR13c, HJA17, Hub09b, SPZ10]. rigidity [STW18]. Rigorous [Mro96]. Ring [DHS98, Mil96, Pro00, Sne98, Zha90, Zha94, AP10, AC19, Ber98a, BFH17, Buc06a, DF05, DGS10, EG21, GRV17, Mou05, Qi23, Sch16, Sht88, TUÖ05, van93]. ringed [FG16]. Rings [Ano01e, Bec90, Bec93, BL93, BF01, Gie98, Göb95, GHS01, HNVL90, HS90, JL91, Kal94, Kal98, Kem96, LS00a, MR98, Mil96, ŌS94, Pri96, Sch90a, Sti87, Von98, Wid01, YNT92, AS24, AB92, ACMB19, BP24, BL06b, BM01, BU09, BJSS89, Bus09, CM17a, CdG09, CLQ10, EH21, EG21, ES11, FF17, FD14, FD18, GLS10, HH07, HM05, Hul13, Jam11, Kin13, LL13, Mar19a, MRW17, Mor20, NP20, OS24, Pau07, Per04, PP17, PRR18, Rei06, ST20, Wei06]. Rioboo [Mul97]. Riquier [WRI09]. Risch [Bro90b]. roadmaps [PES24]. Roberts [NT21]. Robot [AI90, HS89, VGT90, KS04]. robotics [CES23]. Robust [BR15, DYA97, Sch93]. Roch [Hes02, HI94]. Rogers [MSZ09, Sil04]. Role [CC01]. roles [GGdR^{+}13]. Root [CGG12, Col01, EMT21, Koi19, Pan02, BS17a, BSSY18, BK12b, Bur16, CJ15, CWZ23, Col15, Col16, Col17, EH16, GG92, HNE16, HH16, HHT18, HLXL18, LJ09, MSW15, PS09, Sch06, Str12, ST19]. Root-finding [Pan02]. Roots [BF91, BFHT85, For02, Mig92, Mig00, Mou98, MR02, Yan99, dM99, AS23, AHKY09, Ave09, BMS17, BF21, CLX^{+}24, CGY09, Drt06, FDS13, Gal13b, GR11, Hon04, KO17, MMS23, MS11b, Mtn03, PT16, SM16, YYZ12, Zhe21]. Rosenberger [GJT13]. Rosenfeld [GKMO08]. Rossi [ASS13]. rotation [CS16, FS10, FS12, FS13, FDS13]. rotation-minimizing [CS16, FS10, FS12, FS13, FDS13]. rotational [OT87]. Routines [WBM99]. row [BELP13, BCL06]. rows [LY05]. Roy [Lec19]. RRtools [Sil04]. Rubik [KC09]. rule [CKKM10, Vis05]. rule-based [CKKM10, Vis05]. ruled [BEG09, CW03, Doh09, FGVN06, GLSV21, SPD14]. Rules [AB00a, Com98a, Lav91, Wan94a, dB89, Ber04, GGdR^{+}13]. Runge [FPT04, PC98, Sof94]. Rush [Eli04, Gas21].

S [YW87]. SACLIB [HNS95]. SACLIB/PACLIB [HNS95]. safety [LCQ^{+}10, MM10]. SAGBI [AHLM99, BR22, Gat03, Göb98, Nor02, TUÖ05, BC24b, Kha14]. SAGBI-bases [Gat03]. SAGBI-Gröbner [AHLM99]. Sakata [BF20]. samba [Hem18]. same [DZY22]. sampling [DPS16]. Samuel [ST20]. Sandwich [EW86]. SAT [BKG21, BKHG21]. Satellite [VV97, Pal13]. Satisfaction [LC96]. Satisfiability [Fer96, DEG^{+}21, VB03]. Satisfied [FHR99]. Satisfies [JJ01]. Sato [Bah01, BO10, HLO22, Ley01, NOT18, UCJ04]. saturated [Pel03b]. saturation [DMY16]. Saturations [BR22]. Scalable [CG02]. Scalar [FDS13, BF20]. Scalar-FGLM [BF20]. Scalar-vector [FDS13]. scale [KC01]. scaled [EM12]. Scales [San96].

Scene [Cra91]. Scheduling [DHK^{+}95, MG88, Tor93]. Schemas [BM00, FLOR00, FLB00, Fuc00b]. Schemata [AB00a]. scheme [BLR13, CS22, DH16, JLW13]. Schemes [GV99, Miy01, AKR05, Alu03, BCLR13, BCF19, Fer88, Guo20, KLZA12, MP04, Wib07]. Schinzel [SS90]. Schlosser [DT23]. Schmidt [Rei99]. Schmidt-Kolchin [Rei99]. Schreier [Sut16, DS12, DMW17, MO95, Ros93, Sut13]. Schreier-Sims [MO95]. Schubert [CFS24, CM21, HS17b, HSS98, Koh92, Vei97, ZD02]. Schur [EG21, GK16, Koh92, LP03, Pic98, Sau93, Sta23, Ung19]. Schützenberger [GMP13]. Schwartz [Hel16]. science [BBKK15, DK18, Kut10]. Scientific [CJ97, DR86]. Scientist [BCE^{+}94]. scope [Wer12]. scope-determined [Wer12]. SCSCP [LHK^{+}13]. SDP [LR23]. Search [CZ92, LT89, Rob88, UC98, BT09, Bec03, Bon05, BKHG21, JPW19, MH06]. searches [MSZ09]. Searching [SO89]. Second [CTR99, Hen98, Hol85, Kov86, SV92, SU93a, SU93b, Ald23, Arr16, CJP22, HYH04, IvH17, LL16, MW23, Mil93, MV15, Nak16, Bos01]. Second-Order [SV92, CTR99, Arr16, HYH04, LL16, MW23, MV15, Nak16]. secrecy [BD04]. Sections [GO00, OKK98]. security [BGP09, KKK^{+}16, LMA11, Ran12]. security-aware [Ran12]. Seeley [GK94]. seemingly [Drt06]. Segre [Abo10, Har17, Hel16, HT23]. Segrè-driven [HT23]. Selected [CH85, CH86]. Selecting [Lia13a, MO95, PdRAEC24]. selection [Ebe19, PZ12]. Self [BU14, HK10]. Self-dual [BU14, HK10]. Semantic [Wer98, vdH15]. Semantics [ABP96, Fis96, Har92, SAK89, AHH^{+}05]. Semi [BG05, CR88, KKK^{+}16, Liu19, Ren92a, Rup04, Sod96, XY02, Ald23, Bas06, BR10, CDM^{+}13b, CDM^{+}13a, HHS23, KBRV24, MS21, OS04a, WRI09].

Semi-Algebraic [CR88, Ren92a, XY02, Bas06, BR10, CDM^{+}13b, CDM^{+}13a, HHS23, KBRV24]. Semi-automated [KKK^{+}16, Liu19]. Semi-Automatic [Sod96]. semi-discretizations [WRI09]. Semi-implicit [BG05]. semi-invariants [Ald23]. Semi-numerical [Rup04, MS21]. semi-unification [OS04a]. Semialgebraic [ADCZ21, Vor99]. semidefinite [AGS18a, GKMW21, GX04, HNE21, MWZ16, Nal18]. Semigroup [Kon95, HM05]. Semigroups [CH91a, KM01, LM90, LPRR02, NÓ89, ABMN10, EENMP19, GGMFVT13, PCVT08]. semilattice [MJST22]. semilinear [DW07]. semirings [HKSS17, KL19, KL21, OS04b]. Semisimple [BR87, MM04b, OdR09]. semisymmetric [KLZA12]. Semiunification [DR92]. sensitive [CL07]. sentence [GTLN17]. sentence-ambient [GTLN17]. separability [GTLN17]. separable [EG04, MM04b]. separated [AGR95]. Separating [BLPR15, IdW15, Kem09]. Separation [Col01, Ded97, EMT20, KT90a, HHT18, Koi19, Sch06]. September [Ano23m, Ano24h]. Septic [Bru01]. Sequence [Nor95d, She92, BY23, ES11, Kut07, SSS23, KLV10]. Sequences [LR01, NP95, NG93, Nor95c, ABvHP11, APZ20, BGLGM17, BBF17, BE22, CK90, CELG04, CLM16, Gri22, HKL24, HZ19, hHL21, Jar13, JPNP23, KZ08, KMY24, Kuo06, MS10, Mtn03, MGRR23, Nor95b, RRS06, VM14]. Sequent [CCM95]. Sequential [Dur94, DGLM^{+}24]. Sequentiality [KM91]. Series [ASJ97, Bec90, Bec93, CH97b, DD90, FH94, Kal02, Koe92, LW01, Ous91, SJA01, Sne98, Zei95, van97b, AU21, AB09, ABK15, Apé10, BBV15, Ber98a, BM04, BS22b, BS15, CK90, CI07, CFS22, CFRS23, CX09, HKL24, JSC13, Kra95, La 17, LPT20, LS16a, Liu19, MRW17, MJK17, MdCW17, PR12, SK12, WZ23, vdH07c, vdHL13]. series-parallel [BS22b]. Serre [CQ12, ES11, MGRR23]. Server [FKM95]. service [BCR15]. services [BPT11, Ran12]. Set [BCGR92, Bou97, GIM07, Hul99, JWG10, Ley01, LM94a, Mar96a, PH87, Pue89, Sak88, Sny93, Str01, Vor99, BV03, EGS23, GLY09, GVYZ09, GH12, HNE21, HcL21, Hub09a, LFD19, Win06]. Set-theoretic [GIM07, JWG10]. Sets [AM99, ALM99, BCGR92, CR88, EPW90, HH94, KNZ91, OPP93, Ren92a, SG89, ASS07, ADCZ21, Bas06, BR10, Bel03, BP23, BCvdHS11, BLM10, Bur04, CKR04, CDM^{+}13b, Fas10, GS89, GS22, Gol06, HHS23, HRdWY22, HM09, HJS13, HL18, Hub19, JMV23, JLW13, Kin13, KBRV24, KW88, Leb15, LMS09, Nie03, Pel03b, PS13, PES24, Sch03b]. Setting [BTBQM00, LPS93]. Several [Arn95, DTGV02, GVGC99, CFRS23, EGB12, Lev07b, Sau18]. Sextactic [SS24]. Sextic [Gaá95, GP96a, KH23]. SFA [Pro00]. Shallow [Wir09]. Shanks [KT04]. shape [AS07, Alc12]. shaped [BGG13]. Shapes [ERSG05]. Shared [GK96b, Sch91]. sharing [DZY22]. Sharp [Bea92, BTW93, MZ05, BE17]. Sheaf [Bac99]. Shift [SS94]. Shifted [vzGM22, Shp14]. Shifting [Nie94b]. shifts [GKsL03]. Shirshov [BV06, Bok08, GSZ13, LG21, Mad14, ZGG23]. Shoda [BM16a]. Short [DHH^{+}04, CF91b, FU17, GMMM17, HZ04, HZ15, LW03a, LW03b]. Shortest [Nor99, Rol90, VK16]. shuffle [BDM17]. Sibirsky [JLR03]. sided [SS96b, TA87]. Sieve [VM14, PZ12]. sift [QZYC22]. Sign [AZ90]. signal [JMPR04, Par04]. Signature [HV22, Ede13, EF17, ELME23, Lai24, Mal21, MdCW17]. signature-based [Ede13, EF17, ELME23, Mal21]. Signatures [DR23b]. Significant [Tri86]. silhouettes [GLSV21]. Similarity [AHM18, GTLN16]. Simó [AMDW16]. Simple [BCE11, Bro01b, Gol01, HLM95, Pan94, SW95, Wan98, WW94, APS12, BE13, BLM22, BJT22, Eic10, LS16a, MS15, Mar19a, Píl07, RCK07, Wan04, XXZZ23]. simplest [YY03]. simplex [JP10]. Simplicial [BT94, CFS07, BGM15, GDR05, RdC13, DRN24]. Simplicity [Mic88]. Simplification [DS97, GDR05, Sch03a, Zip85, ARST09, BBK14, BVE21, GR98, HS98, Sto11]. simplifications [Bro12]. simplified [HJX16]. Simplifies [Chi08]. simplify [Ebe19]. Simplifying [KR22, Kap87]. simply [aZGS05]. Sims [MO95]. Simulation [MRS96, SS96a, Wei97, BCR15]. Simultaneous [CZ92, DFS11, GVGC99, GLLZ23, Ang24, BELP13, RS21b]. sine [GR98]. sine-cosine [GR98]. Singer [Man93a, MM97]. Single [Col04, EK11, GNP12, Bas06, BK15, EG15, FL11, NÁS^{+}24]. Single-factor [Col04, GNP12]. Single-lifting [EK11]. single-parametric [FL11]. Singular [LW98, SF90, Vid99, AHS18, BSW21a, BP07, BCGY12, CWL08, CK12c, DLLP08c, HR17, KMYZ08, LLZ24, LZ12, MMS23, ST24b, WZ12, vdH13, BMS20, LMZ23, MS15, MS16]. Singularities [BGI18, BS00, JH21, vdH01, BJ21, BMS20, CGL07, CKLZ19, DS16, FGT15, FK04, GS05, MS15, MS16, PD07, RSV09, Sha12, SJG13]. Singularity [BS01, Cou22, MP89, NT21, WC12]. singularly [BM19]. singularly-perturbed [BM19]. SIS [BD87]. six [CC24]. six-body [CC24]. Size [Wol02, DP24, JMV23]. skeletal [Nie03]. skeleton [EFG16]. skeletons [HC12]. Skew [Gie98, HYG24, Koh92, LL13, BU09, BU14, CL17, Li10, NP20]. Skew-Polynomial [Gie98]. Skew-polynomial-sparse [HYG24]. Slice [Rou09]. sliding [GGM10]. Slope [Mil92b]. Smale [CVY17]. Small [BK13, Bru01, DW18, Dün94, Gaá93a, Led00a, MSKO93, Mil87, DFS11, EGB12, Gau09, HR20, HY23b, KY16, KT04, MS11a, MPS16, NY04, Roo13, vzGMS10, vzGMS10]. small-weight [MPS16]. Smallest [Boy93b, MG94a]. Smith [BLS23, DSV01, GHL21b, LLW24, LWXZ23, Vil95, WWWX23, WY11]. Smooth [HHS23, Ahn08, AKS12, BJS04, DEPS11, EGS23, GP24, Gen22, GGEZ12, Lun16, PES24]. smoothness [BFK18]. SMPs [Wan96]. Smullyan [Sta89]. Socle [LW01]. Socles [CH97b]. Software [EW00, KF01, Sch94, Ang18, BBKK15, DK18, Ker17, Kut10, LHK^{+}13]. Software-component [EW00]. Solid [Vor89]. solitons [AFMS23]. Solomon [BB10, BDPR13, LO08]. Soluble [Con90b, Gla88a, Gla88b, GS90, Höf01, Nie94a, Ple87, Sla01, CELG04]. Solution [BF91, CF89, CJ90, FGG^{+}16, NS90, OT87, PV00, Sin91, Tra98, Wol00b, Zha96, ZWH11, AP08, EG15, Har13, HJS13, HTZ04, Lem03, LST03, LZ12, Pra13]. Solutions [AC01, BP99a, Bar99, Bro92, Bro00, BEM97, CE85, Die92, FT95, Gaá93a, HH98, Hv95, KST93, Laz88, LS01, Man93a, Pet92, San96, Sin90, SU93b, Tun02, VRUW99, Vel00, XY02, YNT92, Zha95, dv96, van97b, AB09, ABvHP11, Abr17b, ABPS21, AHS18, BGH^{+}04, BCE11, BJM17, BM19, BC24a, BS24, BS22a, BD12, BR06b, CFS22, CFRS23, Cha14, Che23, CvH04, CS06, Cou18, DS86, FG06, FGH08, FSW10a, FSW10b, Fox18, GIM07, HL17, Heu06, HW24, IvH17, JPP23, Kal11, Mil93, MRV21, NW10, NW11, PDS03, Sau18, SvE14, SvE21, Tun09, Ulm03, VGW18, WK20, WZ12, vdH07c]. Solvability [AK00, Baj86, SSS02, Ngu09]. Solvable [EW02, LM94a, Püs02, Sim90a, Bor22, CM04, DFdG15, HY23a, KRW90, Kem22, MM16, Poz15, XL13]. Solve [EHR91, Mil87, BB10, Izu16, Kho08, KC09, LM94b, Wan06]. Solver [AF96, BVE21]. Solvers [KR94, BKG21]. Solving [AP89, AK92, AGS18a, AP11a, AC04, AM99, ARE02, BGK86, CZ92, CG23, Com98b, Con93, Cza89, DCC95, DH00, Fit85, Fit97, Gaá00, GGM10, Gem94, GV88, HESV21, HS99, HJA97, KFK97, KFF88, Kov86, Kut07, Laz92b, LR07, LE22, Mas16, Maw88, Mil92b, MT01, MR02, Nal18, Pel97, PV02, RZ09, Roj99, RR08, Rou08, SME87, She92, Sit92, Sma96, SBB^{+}89, Str00, Sza08, Wid01, BP11, BSW21b, CM10, DET09, ES18, GH12, HJX16, HcL21, Min98, MP09, MTV21, MS04, Mul04, Ros05, RSS10, Laz09]. Some [AB92, AP93, Bec90, BP11, BGK86, BF11, Cha00, CR90, DS00, DPS23, Eic10, FGP05, Hav91, Hen90, Key01, Lab95, LOOR^{+}03, LLL19, LRW97, ML92, Meg90, Mic90, OT87, PP91b, Poz15, Ris88, RT89, Sal94, Wan94a, Wei94, Yam94, Zha92, AS24, BV06, BDE^{+}21, DS18a, FRR06, GU21, MPS16, MJK17, MPSXD09, MS03b, PS21, SS03a, TZ21, Win14]. SONC [MW23, Pap23]. Sorted [Dom92, HS95, HKK98, MGS89, Smo98, Wer98, ARS10]. Sorting [Tra89, DJS18]. SP [Vei97]. Space [ACM88, GO91, McC88, Alc16, AS05, AHM18, CJL13, CS16, El 08, FS10, FS12, FS13, FGT09, FK09, JWG10, JWC^{+}16, MSS21, PS18b, PH11, Qi23, RSV09, RSS13, SJG13, SS06, UY15, WK91]. spaced [Roc11]. Spaces [Cou00, GHS01, Hes02, AP08, BBV15, BJT22, DH07, DPS16, HJS18, KL17a, LST03, Lou21, MV13, NNPZN19, Rin13, Shp14]. Spacetime [Rd91]. span [ST24b]. spans [HJS18]. Sparse [DE02, DHK^{+}95, DSV01, EC95, FM17, GH02, HM02b, Hua23, JS18, KY16, LS95, LYG15, Min02, MP11b, MF96, Roj99, AGT13, AGR16, AIRR12, BE17, EK11, GLsL09, HJS13, HM21, HG20, HYG24, HNRS21, KsL03, Mas16, MT20, Mul04, PPR13, Sal08, YZ21, vdHL13]. sparse-FGLM [HNRS21]. sparsest [GKsL03]. Sparsification [EM98]. spatial [FG16, Far19, GLLdR19]. Specht [MRV21]. Special [Ano99c, Ano00b, Ano01c, Ano01d, Ano01e, AJ01, BDT22, BB93a, Bos01, BKR19, BBKK15, BK90, Buc92, Bür89, CFMMP10, CH97a, CELG04, CL00, Che85, CL23, DM09, DKM23, DDM15, EY21, ES13, FLB00, GSSST10, HHK^{+}23, HSW97, Hon96, HL97, HTX15, JMPR04, JKP12, KM98, Ker17, KR97, MMY00, MNJ94, NSW16, PZ92, PS95a, Smo98, Tra07b, WS98, Yok17, Zha92, AJGVS09, BFSS06, GS03, KASW05, KH23, MMO18, Piq91, SS18, Sta16]. Specializations [Kal97a]. specific [BGG13]. Specification [Ott91, Smi93]. Specifications [AB99, CSS96, FD93, Pad96, PH87, WG94, BD04, NOF10, Tak93]. spectrahedra [BS17b, Vil23]. Spectral [Boy93c, AHKY09, DJ89, ES11, GDRV21, MGRR23, RRS06]. Spectroscopy [OT87]. Speeding [Pic00]. Spencer [MP89]. spheres [GMW23, Pet10]. spherical [DJ89]. Spinor [BG01]. spirit [Gir21]. Spline [BX97, DiP16, MV13]. split [GTLN17, KS12a, Roo13, SS88]. Splitting [Kut19, MS15, Pan96, DT06, FS23, HAGW12]. Sporadic [Gol01, HLM95]. SqFreeEVAL [BK12b]. Square [BFHT85, EH16, BF21, DHS22, HL17, HS17b, Nag21a, Sla07]. square-free [Nag21a, Sla07]. Squarefree [DTGV01, XWL23, HOP06]. Squares [LSW01, AOW23, DP24, DR23a, DW18, FMM07, GKMW21, GS24, GGEZ12, GIJ14, IdW15, KLYZ12, MES19]. Stability [GV99, HLS97, Kal97a, PC98, Sof96, WKB86, HTX15, SS05]. Stabilizability [YXL99]. Stabilization [SS98b]. stabilizer [Eic02]. Stable [AFT08, BL98a, GLW99, BCLR13, CR11, MP11a, Poz19, Qi23]. Stafford [HS01, Ley04]. Staged [GMN22]. staggered [HM23]. Standard [BM88, Bec90, Bec93, CM21, Cla91, Cow92, EPP21, HH07, JL91, Lüb23, Mar08, MRW17, Pas86, BR88, FK04, GP96b, JP10, NT17, NWW21, Str19, SD05]. Stanley [JV09]. star [BGG13]. star-shaped [BGG13]. Stars [Sof96]. State [FK09, Vor89, BM88, Nor15, dCW09]. states [BDE^{+}21, Gon17, Lic21]. Static [HGKV11]. stationary [AR06]. statistics [BD17]. steady [BDE^{+}21, Gon17, Lic21]. Steiner [BH87, WC12]. step [BS18, BdlCRLS19, LLZ24]. Steps [Wir12, HRdWY22]. stereo [LMM05]. stereo-isomers [LMM05]. Stickelberger [Sch10]. Stictenoth [DS09]. Stiefel [ASS20]. stiffness [Loj13]. Stirling [CJP22, Kau07]. STO [AvW94, KP99]. STO-problem [AvW94]. stochastic [AGS18a, CFSGL21]. storage [PNM13]. Story [MN89]. straight [AGR16, HG20, JS07, JS18]. straight-line [AGR16, HG20, JS07, JS18]. Straightening [MW91]. Strands [Bur01]. Strassen [CH17]. Strassen-like [CH17]. Strategies [LS98, Ant05, Bon05, GL05, LGS90, RV05, Vis05]. Strategy [AT96, BGK96, Sod96, RV03]. Strategy-Accurate [AT96]. Stratification [Alc16, LMP89, NN10, Ric99, Wal05]. stratifications [Wal03]. Stratified [DNV03, YP91, vdPT15]. Stream [LT22, BGP09]. Stream/block [LT22]. Strict [BBV22, Str00, CCG06, FK11]. String [Bir98, BGHW06, OKK98, Zha92]. String-Rewriting [OKK98, Zha92]. Strings [Sim91]. Strong [ABL93, Arn95, BB00, Fri89, YXL99, BM16a, CF91b, DW21]. Stronger [GMP22]. Strongly [Dur94, Hag89b, Per04, BCLR13, CR11, DMW17, LLL13]. Structural [BS22b, Cra91, TL96, CHU19, Rap06]. Structure [Bro07, BDM17, Egl96, EP02, GC93, LRD00, PG86, Rón90, Tes99, Yan98, AC04, DH07, GMN22, HMS17, HY21, LZ12, LXZZ23, LR98, MS16, Mou05, RS19, Rua09]. Structure-Preserving [Egl96]. Structured [Sod96, BNN17, BE22]. Structures [AB99, DR86, FS95, Lia13b, Rob86, Wal02a, Wal02b, HOS23, Loj13, MM88, Nie03, Rei06].

Students [Bos97]. Study

[AGMT98, BB92, PH87, SV92, BR12, BEG09, DP19, EH16, KLR93, KD90, PdRAEC24]. studying [AB22, FK04]. stuffle [BDM17]. Sturm [Moe05, ZX20]. style [MBL21]. Sub [GV88, Sau93, LNP^{+}21]. Sub-Exponential [GV88]. sub-linear [LNP^{+}21]. Sub-Transforms [Sau93]. subalgebra [Kha14]. subalgebras [AGSM17, BR22, DFdG13, Roo13]. Subanalytic [ABvHP11]. subbilateral [NP20]. Subdivision [MP09, BS17a, BSSY18, BCGY12, BGT20]. Subexponential [CDO97]. Subfield [GP96a]. Subfields [CFM96, Dab97a, Gaá00, KP97a, Dab97b, EK19, KP97b, SvH19, vHKN13]. subgoal [Bon05]. subgoal-reduction [Bon05]. Subgroup [HH99]. Subgroups [BC91, CCH97, CCH01, Chu99, EW02, Gla88b, Hul99, BJT22, BC89, CH04, CHSS05, FG08, Hel00, LMP19, Lüb23]. submatrices [Koz23]. submodule [LMR94]. submodules [BL12, DMY16]. submonoids [GGMAMF19]. subnormality [Mur23]. subprogram [MM10]. Subquadratic [Tho02]. Subresultant [She92, El 05, Lec19, Sza08]. Subresultants [Apé10, BKSV20, DJ05, DKS15, HY96, Hon97b, LRD00, Mul97, Vil95, DHKS07, El 03, LP03, PR22, PB07, RS11a, RS20]. Subring [Sne98]. Subroutines [SR86]. subsequences [BP23]. Subset [NU18, CLS91]. subsets [Mic13]. subspaces [AH13, vzGGZ21]. Substitute [Sim91]. substitution [DFS11, Har09]. Substitutions [Ede85, KFK97]. Substring [Rob88]. Subsumption [DR92]. subsystems [DHS22]. Subterms [Rus87]. subtractive [BLV18]. Subtyping [DT95]. subvariance [Bel03]. subvarieties [KS12a]. Sum [AP93, GS24, LSW01, AHKY09, GIJ14, JPP23, MES19, NU18, Pap23]. Sum-of-squares [GS24]. SumCracker [Kau06]. summands [ASS20]. Summation [BK99a, Kar85, Kau07, Koe95, MS95, Pau95, Sch17b, ACGL04, BKSS12, PS95c, PS95d, Rie03, Sch08, Sch16, vdH07b]. Summations [Man93b, Wan94a]. Sums [GO90, Tak95, BLS17, CM09, DHKS07, DHKS09, DKS15, DP24, DR23a, DMN17, FW15, GKMW21, GMKP21, GGEZ12, IdW15, KLYZ12, Kau06, Kra95, KS12c, LP03, Liu19, MV10, Naw16, RS11a, RS20, XW20]. sums-of-squares [DP24, GKMW21, KLYZ12]. Super [Ges92, BP09a]. super-reduction [BP09a]. Superalgebras [GK96a, Kol08]. supercongruence [Mao21]. supercongruences [HMZ21, Liu19]. Superfluous [Bec93]. Superposition [GN04, Rus91]. Superpositions [KMN88, SK91]. superscalar [VB03]. supersingular [Pie24]. supersymmetric [Qi05, Qi06]. Supersymmetry [dos89]. Support [Nor95a, AJGVS09, ÀL06, EM12, Mon05, dC10]. supported [AJGVS09, Bih15]. supporting [ASS13]. Supports [EMS00]. Surface [GKLM91, WW94, AMT09, BNT18, CW03, CCL05, CR19, DMR12, DZ09, FGL04, Sch03a, SS05]. Surfaces [AS01, BX97, CGZ00, D’A01, DS00, FHL96, FKM95, GSST98, GV97, LSW01, MC92, PP97, Rie93, Sch00, AHW05, AKS12, AJGVS09, ASS07, Alc08b, BBV15, BCI13, Bec09, BLT22, Bri06, BG05, BEG09, Doh09, FG16, FGVN06, FGPT03, FGT15, GLSV21, GS07a, GSHPBS12, GSPB17, GEL05, GP13, HS09, HC12, LC16, Lub14, Par22, PDS03, PDS08, Pet10, RSTV16, Sch98a, Sei20, SPD14, SJS06, VL16, Wan04, WC12, WG18, dGPS09]. surjective [SS88]. Surprising [Ber93]. Surprisingly [SW95]. Survey [KS98, vzGP01, EF17, Top14, Vis05]. Suslin [LS00a, LY05]. suspension [KLR93]. Suzuki [Kos07]. swarm [Loj13]. Sweeping [NS90]. Sweeping-plane [NS90]. swell [NPD09]. swung [RSTV16]. syllogistic [CGO88]. Sylow [BC89, BC91, CCH97, FF17, LMP19]. Sylvester [BST16, CK04b, DHKS07, DHKS09, DKS15, KS12c, LP03, LR01, Mul01, RS11a, RS20]. Sylvester-Habicht [LR01]. Sylvester-type [CK04b]. Symb [AP17, HZ15, HdC16, KMR18]. Symbol [ES98, Sha90b, Tri86]. Symbolic [ASJ97, AP04, AK06, Ano01b, Ano02c, AB89, BGH^{+}04, Bar13, BFK02, BB92, BCGR92, Bur92, Car99, CV00, CL00, CCM95, CS09, CD87, CN07, Cuy97, DHM11, DK18, DEG^{+}21, DT95, DR86, Edi85, Eis90, EY21, EC87, EP02, FPT04, FS13, Fer06a, Fit89, Gar95, GLsL09, GH97, GL92, GK94, HMN06, Hen90, HJM94, Her94, Hil05b, HJA97, Hon96, HP08, Hug90, Kal90, Kal00, KT94, KR94, KL98b, KS04, KL90, Kri85, Kut10, LPY01, LL16, LFD19, Lic21, LLL13, MJK17, MHXD09, Mro96, NSW16, NS90, NR97, Nor95b, Ore11, Par04, PSZ91, Pau95, PS95a, PdRAEC24, PH11, RT85, Roq13, RS11b, Sag89, SS98a, SS99, SJA01, Sau18, Sav90, SME87, SW91a, Sha12, She97b, Sod96, Sof94, SBB^{+}89, Tra98, TL96, UYSA89, VGT90, Vor89, WS98]. Symbolic [Wol00a, XL13, Yap90, Yok17, Zha96, dv96, ACGL04, ACS13, BBK14, BD87, BGL14, BKSS12, BBKK15, CC24, DM09, DZY22, EG21, ES13, ERSG05, Fox18, GS03, Gue18, HKP^{+}06, JMV18, Kau06, LHK^{+}13, MPH17, MSY21, MPT20, MP04, MKF93, Naw16, PT14, Piq91, PNM13, RS21a, RZ09, Ros05, Sch08, Sch16, Wan86, WZ12, Ye17, YW87, vdH07a, Buc92, CL23, JKP12, PZ92]. Symbolic-Numeric [KL98b, RZ09, WZ12]. Symbolic-Numerical [Tra98, MPH17]. Symbolically [Mil93, DES07, Maw88]. symbols [Nak06, NOF10]. Symmetric [Cla91, CH96, CF09b, Hel00, KL19, KL21, KKL92, MRV21, PB07, Pro00, Ste95, BS18, BGI11, BR09b, BDM^{+}16, BK16, Cha14, Cla21, DH07, GMF13, KS16, Koz23, Mad14, MS03b, NPP17, Poz24, RS20, Sei20, SvE21, Ste13, WS23, WY20]. SYMMETRICA [KKL92]. symmetrically [BFMS87]. Symmetries [CV00, Hem02, MC97, SV92, BN04, CGK09, FK89, HJ18]. symmetrising [BCCK20]. symmetrization [Hub19]. Symmetrized [Ryb01]. Symmetrizing [CIM17]. Symmetry [BH23, EP04, LP02, BH21, Hie16, JKKK20]. Symmetry-based [EP04]. symplectic [BKW20]. symplectically [BR13c]. Syntactical [Bur01]. syntax [SR07, SP10]. Synthesis [CM93, DR93, DJ15, FB93, FD93, FLOR00, PMW93, Tra89, DJS18, EH16, QHL^{+}13, ST89a, STDD16]. Synthetic [SW91b]. System [AK92, AGM97, BP99a, Ber93, BCGR92, CP97, Ded97, Die92, Dün94, EC87, GP96b, HS95, Hen90, Jir97, KKL92, Laz09, MM00, Mil87, MT01, MR02, PMW93, Pro00, RST01, San95, Sch94, Tra98, Tri86, Vei97, BV03, BD87, BS24, Bed07, Bed09, BCP97, Bur03, BK16, DJ07, ELME23, ES18, GG92, KSW13a, Khe03, Lem03, Mas16, MT88, MS04, Mul04, PS09, dCW09, Sid93, SH17b, Wan86]. Systematic [DH00, Cox19]. Systems [Ano96, ACGR01, AM99, ARE02, Ave86, BP85, BC01, Bar99, Bir98, BF91, BGK86, Boo87, BEM97, BH00, Che92, CD00, CK99, CCM95, Com98a, Com98b, DT95, DHK^{+}95, Dur94, Ebe01, EG15, Fit89, FJN93, GV99, GC93, GLW99, GH02, Gem94, GZ90, Ges97, GH97, GV88, Har92, HKL99, HH94, HLS01b, HH99, Kah95, KS98, Kem99, KM91, KFF88, KF01, Laz92b, LA96, LS02, LP02, MSKO93, Mat01b, MT93, Mid94, Nau98, Nie94b, Ohl95, OKK98, Ous91, Roj99, SS96a, Sch85, Sit92, Sny93, Str00, TL96, UYSA89, Vor92, Wan91, Wan93, Wan98, Wan99, Wan00, Wol02, XY02, Yap91, YXL99, YNT92, You89, Zei95, Zha92, ZSY93, Zha93, AP08, ABK15, Abr17b, AHS18, Ang15, AMW12, BGLHR12, BP09a, BCE11, BE13, BELP13, BJM17, BM19, BSW21a]. systems [BJ21, BNN17, BW05, Ber98b, Bih15, BPH07, BR13c, BS22b, BLPR15, BKG21, BGHW06, BR06b, CFRS23, CES23, CS05a, CJP22, CM10, CM12, CDM^{+}13a, CM17b, CKLZ19, Che18, CGY09, CGG12, CJ15, CWZ23, CK03, CK04a, CQ12, CGK09, CDSS09, DJO^{+}11, DET09, DHS22, Dum09, EYZ21, EM12, EMT20, EMT21, EW07, FEV16, FLE^{+}23, GLY09, GVYZ09, GLLdR19, GLLdR21, GES05, GV16, GPGO16, GMP22, GDRV21, HBN95, HR12, HT91, HL17, HESV21, HOP06, HJS13, HTX15, HKYY18, HcL21, HY23b, Hub19, JLR03, KS06, Kap87, KKM15, LMA11, LR07, LE22, LLZ24, LST03, Li04, LW12, Lin18, LR98, LH17, MM09, MS21, MPT20, MGRR23, MW10, MS03b, MTV21, MRV21, NOT18, NOF10, NW11, Pom11, RH18, RZ09, SLK11, SLX^{+}13, SPZ10, STW18, Str11, Sza08, TM85, Vis05, Wan06, Wan18]. systems [Wir09, Wol03, WZ12, XWL23, YW87]. Syzygies [BS88, DS16, HNR24, WG18, AHW05, BDM23, BD16, EMSS16, HV22, Möl88, RR05, Wol03, Hub09a]. Syzygy [HT17]. Sz [Kos07].

Table

[Sch90b, Ber04, BDE^{+}16, TV18, Ung06]. Tableau [AGRZ99, AB01, Fuc00a, MGL00]. Tableaux [Cla91, Wil95, Bec03, CIM17, CM21]. tables [DO06, Gal13a]. Tactics [ACGR01]. Tame [Zie16, von90a, Sch05, Wen06]. Tamely [HM02a]. Tameness [NNPZN19]. Tangent [GOT05, Nak16]. tangents [CK12d, Zen06]. Tarski [Bro12, Gri88, KBRV24]. Tate [HLO22]. Taxonomies [QSGB19]. Taylor [Sei02, Ye17]. Teach [Bos97, Mon97]. Teaching [Kal97b]. technical [SWF11]. Technique [FF92, AG91, JKKK20, LOOR^{+}03, PNM13, YY03]. Techniques [AB00b, ABM^{+}23, BGH93, BS01, BTBQM00, CP00, CE96, Mil87, Mil92b, BCvdHS11, CDM^{+}13b, DJS18, DP09, Dur09, FGPGP14, GDR05, HESV21, HNRS21, MV13]. Technology [GGAVRC13]. telescopers [CCF^{+}15, CDWZ21, KY15]. Telescoping [ACGL04, Zei91, CK12a, CvHKK18, GHLZ22]. Temporal [AM89, Ano96, CSS96, ET96, FT97, Fis96, Frü96, GP96b, LO96, MRS96]. Ten [Sto11]. Tensor [Bac94a, BG01, KMM22, Ryb01, DRN24, BBCM13, BMT21, PRR18, Qi05, Qi06, ST24b]. tensor-product [BMT21]. tensors [BGI11, HHLQ13, MMS18, OO13, SSV23, WS23]. Term [Ave86, BGK96, CMR15, Dur94, Fit97, FJN93, Ges97, HH94, Kah95, KM91, Lav91, MGL00, Ohl95, PZ96, PY94, PP91b, PS93, You89, Zan95, BY23, GG92, Hre06, Kap87, LLW03, NOF10, Tra07a, Wir09, Zan94]. Term-ordering [CMR15]. Term-Rewriting [Kah95]. Terminating [Ges97, BSW21b]. Termination [BP85, Der87a, Der87b, Ges97, GAO02, XZ10, Zan94, Zan95, KsL03, Kap87, MU04, MO21, XL13]. Terms [Ae02, BN01, Boy93a, Boy93b, Che85, HS99, Kar85, MS00a, Pel97, AP04, Bad06, Cha14, CCF^{+}15, CK19, DS15, KY15, KS19, WK20]. ternary [BS17b, Vil23]. Terrains [CS89]. tessellation [HS09]. Test [Bou97, HH94, KNZ91, Mon92, Sed02, Adl16, BSSY18, BFK18, CF91b, GTLN16, KK17, MP04, vdHS06]. Testing [BW87, HLS97, Kal85, Kal87, McC97, O’B93, O’B94, RR90b, CH03, Gal87, GRV17, HT23, Kal90, Mic13, Shp14]. Tests [BB00, Car01, BGT20]. tetrahedra [DZY22, Tsa16]. Tetrahedral [Her94]. Tetrahedrizing [EPW90]. th [BY23]. Their [Bro92, Fuc00b, GSST98, KT90a, Zip90, Ald23, BP09a, BCE11, BE13, BBRS24, BM01, Bro00, Bro03, CV11, FS16, GLSV21, GR11, HH07, HKL24, LLL19, LWXZ23, Lüb23, MBC^{+}10, MS03b, Naw16, PZ96, Piq91, PSV11, PWZ18, Sch17a, TS24, WRI09, dG09]. Them [Mon97, BB10]. Theorem [AGMT98, AL10, Ano00b, AB00b, BF95, BZ03, BT98, Bon96, Bou97, CR90, FT97, FD14, GC93, HS01, Hsi87, JL91, LS00a, LRD00, LBM98, ML92, MR87, NSW85, NR95, Pad96, Pet00, Pue89, Rus91, Soc91, ST89b, Wal02a, Wal02b, Wei94, Wil95, Zha94, AMDW16, AHL03, Bon05, BLPM19, CIL07, CFS24, FGLH^{+}23, IKGT11, KS86, OB03, PSS12, PS95e, RV03, Ric91, Sid93, VK21, ZX20, CPR09, CPR11, GS23, Sch10]. Theorem-Prover [ST89b]. Theorem-proving [Rus91, Bon05]. Theorema [Win06]. Theorems [CJUE01, DTGV02, Bro07, HdC13, HdC16, Ley04, LW03a, LW03b]. Theoretic [Cra91, Laz92a, PH87, GIM07, Har14, JWG10]. Theoretical [Gre95, BBC^{+}11, GVHHUE05, SA89]. Theories [ALM99, Baa89, BS96, BS86, BHSS89, Gar95, JM95, KR91, NR97, Pet00, SS89a, SS89b, Tha93, Tre92, Yel87, DKM21, Fer98, LM94b, TRRK10]. Theory [Ape98, AB89, BE10, Bos01, Bos97, CH85, CH86, Cow92, DTL10, DS02, EM99, FJN93, GG99, Hsi87, Kar85, Leo91, MMY00, Mic90, MR87, Pau86, Ren92a, Ren92b, Ren92c, Rob86, SS96b, Sch17b, She97a, Sie89, Sof96, YX95, vdP99, AU21, AKL17, BM88, BVE21, CFMMP10, CM17a, CM21, Col05, DEPS11, DT06, Dra05, El 03, FS23, Fre13, GH05b, GES05, Giu88, HT91, HS21a, HOS23, HLSS15, Hir89, HHLQ13, Kem16, KKK17, KD90, Lai24, Li04, LMS09, Li10, Mer10, MS03b, Mor91, NP20, Per04, PWZ18, Sch07, Sch08, Sch16, SS03a, Win06, vdP05]. thesis [Buc06a, Buc06b]. Theta [Pie24, Ye17, Ye18]. Third [Nak16, SU93a, SU93b, SW97b, Ulm03, WK20]. Third-order [Nak16, WK20]. Thirty [Laz09]. Thom [CR88]. Thomas [BGLHR12, HPT02, Heu06]. Those [BCE^{+}94]. thousand [Rou08]. thousand-digit [Rou08]. Threaded [BGK96]. Three [ACM88, Bur01, EPW90, McC88, SS92, Sha90b, VRUW99, Zei95, BDPR13, CDWZ21, Cou22, DO06, Eng10, FS16, Nor15]. Three-Dimensional [ACM88, McC88]. three-state [Nor15]. three-way [DO06]. threefolds [Hie16]. threshold [HT17]. Thue [Boo87, Heu98, HPT02, HTZ04, Heu06, Pet87, Yap91]. Tietze [Rob88]. tight [HJS18]. Time [AV96, ACOR00, Chi96, CKS99, Dic92, FB93, GV88, LO96, PS18b, Sed02, YNT94, AM88a, Bas06, Ber98a, Cla21, CK12b, Col17, FG06, HLN^{+}21, MM16, Mur23, RH18]. Timed [SJG96]. Todd [CLW95]. tolerant [Abb17]. Too [vdH02]. tool [FK04, GMMM17, Pra13]. toolbox [BD17, BKRW17]. Tools [CH95, GVGC99]. top [MBL21]. top-down [MBL21]. Topics [Hes02]. Topological [AM88a, CGP23, BD17, FGT15, GS22, Hel16]. topologically [Lou21]. Topology [CR88, El 08, HPRS11, Ric92a, AS05, FGPT03, FGL04, GDR05, IMP17, KRTZ23, KS12b, Ker17]. topos [Hir89, Nie03]. toral [Roo13]. tori [Gal16, PY05]. Toric [CV11, CM97, CDSS09, EMS00, GES05, Sop13, STVvR24, Ver00, ABB^{+}19, ATY08, AT08, BGMSG07, BGM15, BE11, BR15, BK13, CM21, CC07, DHH^{+}04, EGW09, GMN22, GMS09, Lun16, Nor15, OK08, RS16, Rua09, SS06, VJ07, BLR99]. Torsion [dGN02, CE19, FG08]. Torsion-free [dGN02, CE19]. Total [Zan95]. Totally [Gaá95, Ges97]. tower [CCQ18, DS09]. Towers [HM02a, DS12]. Trace [MMW11, BCCK20, FMTT13]. Trace-based [MMW11]. traces [JFMRS12]. Tracking [vKT93, HL16]. tractability [GSSST10]. Tractable [HYH04]. Trading [CK12a, vdH10]. Trager [Mul97]. trajectory [Pal13]. transcendence [BDM17]. Transcendental [Kno92, Kno93, Bro90b, Gue18, MW12, Raa12, Str11]. Transducer [Du 99]. transform [AK04, AK06, Cox22, FK11, KS16, RH18]. Transformation [LM90, LPRR02, YI94, dB89, BGL14, GKO09, SLK11, Vis05]. Transformations [BB93b, CD87, Jef97, Rob88, She97b, SG89, Bîl11, Deu93, GS89, Nak16, Pra13, Sta16, WyW93, WS09]. Transforming [BR12, LW12]. Transforms [Sau93, CM04, Cla21, Cox21b, Har14]. Transitive [But93, CC91, KM00b, Roy87, RP89, Cla22, Els12, HR20, Hul05, MAN^{+}10, PSV13]. Transitivity [ABL93, BW22]. Translation [Boy92, PH87, PG86, Buc06b, Lev21]. translational [WG18]. Translations [Egl96]. transportation [BR06a]. transseries [vdH11]. transversal [Eit94]. Treatment [Yap90, Izu16, Sag88, Sag89, Sch07]. Tree [BH87, Ger06, KFK97, LM94b, RR90a, RV05, Sod96, DG20, GR10, GMN22, HJA17, Qi23, Wil93]. tree-based [Qi23]. tree-decomposable [HJA17]. Tree-Structured [Sod96]. Trees [GL92, Lab92, Pue89, CFS07, Coo09, DV21, DJS18, KMY24, Vat06]. Triangular [Ang15, AM99, ALM99, CDM^{+}13a, GSSV12, Kal93, Sta18, Wan00, ABM^{+}23, BCvdHS11, CM12, CGY09, DKLP21, GPGO16, LPR17, Leb15, LMS09, MRSW07, MBL21, MV13, PS13, Sch03b]. triangulation [AMT09]. Trigonometric [GHC92, Jef97, PS00, CL20, CLX^{+}24, HS98, MJK17]. trigonometric-polynomial [CL20, CLX^{+}24]. trinomials [KO17, Koi19]. Triple [HLM95]. triples [GJT13]. trivial [GHL21b]. Trivializing [Píl07]. Tropical [AV11, GKS12, Gri23, AFMS23, AGS18b, BBRS24, BJS^{+}07, Coo09, DV21, FGLH^{+}23, GP24, GK16, Gri20, Gri22, GR22, HJS18, KL19, LM20, Ren17, Rin13, The06, Vac18, VVY21]. Tropicalization [WY20]. true [AGS18b]. Truncated [MTV21, Cox22]. Truth [BDE^{+}16]. Trying [She97a]. Tsun [GK12a]. Tubular [Sch00, RSTV16]. tuples [ST24b]. Tutorial [Bie85, CFG^{+}86, Ant10]. Twelve [Roy87]. Twin [Abb12]. Twin-float [Abb12]. Twists [KT02]. Two [BL98b, BFHS92, BS01, BGS11, CS22, Chu99, CD87, CJMP97, EW86, Fit89, FMR04, LSW01, Laz85, Laz88, LLZ24, Lev99, PV02, Rut93, Sak88, SS98a, SSS02, Tay02, Vid99, Wid01, ACMB19, Apé10, ASS20, BDM23, BGM06, BFMS87, BM10, Brå24, Cox21a, Cox21b, FGVN06, GVHHUE05, GG92, GIJ14, HT17, HQS19, HSV08, JWC^{+}16, KLZA12, KP15, Ley04, Nor15, PY05, Pet10, PT98, Ros05, SS90, Sto20]. two-body [PY05]. two-bridge [KP15]. Two-Dimensional [Sak88, FMR04, ACMB19, HSV08, SS90]. two-parameter [Pet10]. two-parametric [PT98]. Two-Phase [Fit89]. two-point [Ros05]. Two-step [LLZ24]. two-variable [HT17]. Type [AM88a, BP99b, CH85, CH86, DS00, Gaá02, Har92, HRT01, Pau86, San95, YX95, BL06b, CK04b, CLS91, CO94, CO96, EK11, GMP13, GSZ13, HJ15, Hir89, KRW90, KK09, Lev21, Sil04, Zan94, dG01]. Typed [Hag89b, HKK98]. Types [MMO94, MdCW17, CS16, CGP23, EL12, GP20].

UML [BPT11]. Unavoidable [Pue89]. Unbounded [Meg90, PES24]. Uncurried [KKSd96]. Undecidability [Tre92]. Undecidable [Ges97, Ott91, SS89b, Zan95]. Undergraduates [She97a]. Unexpected [Szp22]. Unfold [BB93b]. unfoldings [AK86]. Unification [AK92, Baa89, BS96, BN01, BO99, BZ93, BS86, BJSS89, Bou93, BHSS89, Bür89, Dom92, Fag87, For87b, IZ96, KFK97, KR89, LC89, MN89, MGS89, Mil92a, Pau92b, SS89a, SS89b, SS96b, Sie89, SAK89, TA87, Yel87, BL06b, Con93, DJ92, GS89, GGSST10, Kap87, Le 89, OS04a, SSS05, SSSK18, SG89, KLV10].

Unifications [Ede85]. Unified [CK99, Bon05, Cla22, MM88]. uniform [Guo20, OS04a]. uniformity [MP04]. unifying [Bel03]. Unimodal [BMS20, MS16]. unimodular [LY05, SS06]. Union [BS96, Ore01, Gen22]. unions [Pis04]. unique [MS11a, VK16]. Uniqueness [Bec93]. Unirational [BEM00, FL21, GS06]. Unit [CDO97, SS96b, Sma96, Zha93, CVZ21, VV18]. Unitals [Key01]. Unitary [Tha93, GR12]. Univariate [BLPR15, CE96, DTGV01, DTGV02, For02, KL98a, Mon92, NY99, OK08, Pan02, SvE14, ST19, TCT23, CGG12, CKM09, Dah22, EGB12, Gal13a, HHK17, HY21, MES19, Nag21b, PT16, PDS08, TUÖ05, WWWX23, vzG13]. Universal [Gol08, ST24a, AK86, FS98, Gol06, HP91, IL09, Sau18, SS03a]. Universally [Kol08]. unknown [LCQ^{+}10]. Unmixed [BRM01, HM02b, Min02, CK03, CK04a, EK11, Khe03]. Unmixed-dimensional [BRM01]. unnecessary [AH05]. unrelated [Drt06]. unsatisfiability [Gal87]. Unusual [DR86]. unwinding [MCJ21]. Upgraded [BCLR13]. Upper [Laz92a, HS21a, KL21, MZ05]. upper-bound [KL21]. Use [BCE^{+}94, Bos97, CO01, EHR91, Fuc00a, Fuc00b, Hav91, LBM98, Mee94, MNJ94, NMM90, CHU19, FK11, Loj13, VB03]. User [AGMT98, BT98, KM98, KS98, BCP97, HPRS11, LLTPT^{+}11, YW87]. Uses [CF91a]. Using [AV00, BS90a, BBB92, Ber93, BB93b, BH00, BC91, CP97, CDF92, CJMP97, CGK09, Ebe01, Fit89, GKLM91, GAO02, GV97, GL92, GHC92, HH94, JSC13, KFK97, KT02, Kap86, Kem16, Lab95, Man93a, Mil87, MT01, PP91a, Pue89, Raa12, RT85, SM18, SS88, SJG13, SR86, TUÖ05, dos89, van97c, von90c, ABPR21, AK04, AK06, Ald23, AGS18a, AHKY09, AG91, AK86, AHL03, BSW21a, BMS20, BP00, BC05, CK03, CK04b, CK04a, Col16, Col05, DS09, DM05, ERSG05, FS23, FGVN06, Fox18, GKMW21, GVGC99, Gon17, GMS09, GGEZ12, Hal13, HI08, HESV21, HS21b, Hub09b, IvH17, JS07, KMYZ08, KNZ91, LC16, LO09, LS11, LZS11, LS12, LHK^{+}13, LW01, MM06, MS15, MS16, Mas16, MH16, Möl88, MP11b, Ng89, PT14, PNM13, RH18, Roq13]. using [Rou08, Sag88, Sag89, Sag14, Sek11, SL92, Sid93, STDD16, Sti03, Str06, Str16, SH17b, Szi17, Tsa16, Wan06, WC12, Wei13, Wol03, dGPS09]. utilization [Kad13].

V4 [DFK^{+}97b, DFK^{+}97a]. Vahlen [VK21]. Valiant [von87]. Validated [KS06, KR97, Poh97, Str06]. Validation [HS97]. Validity [CGZ00, MPP19]. valuation [DMY16, Vac17]. Valuations [MM00, MS02, PV02, Mos08, Oki23]. Value [Mil92b, Mos08, IK21, KMYZ08, Ros05]. Valued [CRAB91, Stu00, BF95, OS92]. Values [BR87, Zip90, Bod04, JMV18, XW20]. vanishing [Fas10, GSW11]. Variable [CKS99, Eis90, Sch91, Bec03, GGSST10, HAGW12, HT17, LXZZ23, PdRAEC24, Str11]. Variables [CD87, Laz85, Lev99, Rut93, SS98a, SSS02, Sne98, Wan94a, CFRS23, CDWZ21, GHL16, HQS19, Kut07, Sau18, SSSK18, Shp14]. Variadic [DKM21]. Variant [HE12, EP10]. variants [BF20]. Variational [Mil87]. variations [JWC^{+}16]. Varieties [AH01, Bur92, BEM00, Chi96, EMS00, Kal93, Ore01, Wal00, ZD02, Abo10, AH13, AML19, ABB^{+}19, AP22, BKV24, BL06b, BP07, BBRS24, BGM06, BGM15, BE11, BJS^{+}07, BS09, CFSGL21, CRSW22, CFS24, CM21, CC07, DEPS11, El 05, FM24, Gau09, Giu88, Har17, Hel16, Hel00, LR15, Lun16, Pie24, PW06, Qur17, Sch07, SSV23, VJ07, vzGM22]. Variety [GHL^{+}00, BJS04, FKO18, GGEZ12, HMXD07, Hor24, JLR03, MHXD09, MPSXD09, Mor11, SS06, Stu17, van93].

Vasconcelos [BST16]. Vector [LPY01, Tho02, Wor94, BR09b, FDS13, JT03, NT21, Pos18, YY03]. Vectorization [HCB96]. Vegas [BCG10]. Verification [KL98b, LNP^{+}21, BPT11, BD04, GGdC23, GKM05, KZ10, KKK^{+}16, MMW11, Ran12, TUÖ05, VB03]. verifications [GHS08]. verified [MBPLRR10, MRH23]. Verifying [Hie16, LCQ^{+}10, SWF11, Sim87, RS21a]. Veronese [Abo10]. versal [MP89]. Version [HS01, PS95b, GR22]. Versus [Cuy97, Lan10]. Vertex [RP89, HR20, PSV13]. Vertex-Transitive [RP89, HR20, PSV13]. vertices [DS18a, KLZA12, PSV13]. Via [Sma96, AHW05, APS12, Ang18, AB05, BGI18, BL12, BDLP22, BD16, BKHG21, BDM^{+}16, Bur03, BST16, CW90, DZY22, Dey21, DEPS11, DV00, DIdW18, DHS22, DR23b, FS98, FG08, FFP98, GS24, Gal13b, GLW99, GLLdR21, GG92, GDRV21, Guo20, HJX16, Har09, JKP98, KLYZ12, KLZA12, Lam91, LXZZ23, MW23, Mao21, MM16, MS21, MG94b, Mro96, Nak16, Nie03, OK08, Pic00, PRR18, Pos18, RZ09, Sei02, WW94, WZ12, XW20, dCR17]. Vibration [OT87]. Vibration-rotational [OT87]. vibrations [Sag88, Sag89]. vibratory [JSC13]. View [AB00b, BdS01, Rie93, BNT18, MM88]. viewpoint [Hir89]. views [Brå24]. violator [DPS16]. Virtual [CN19]. virtually [BJT22]. Visibility [CS89]. vision [FKO18, NPD09]. Visualization [FKM95, BS21]. Visualizations [AGM97]. Viterbi [Kuo06]. Vizing [GKMW21, GS24]. Vladimir [MSV22]. VLIW [VB03]. Volume [Ano99a, Ano99b, Ano00a, Ano01a, Ano01b, Ano01f, Ano02b, Ano02c, BFHS92, EC95, BFMS87, BEZ23, BR09b, Tsa16, Ano06]. Volumes [Ano04b, Ano04p, GLW99, BBV15]. Voronoi [AH23, BKV24, CRSW22]. vortex [Tsa23]. voting [MRG17]. vs [IZ96, RSTV16].

Walk [CKM97, Aue05, FJLT07, Kha14]. walks [BDS17]. Wall [CG02]. Waring [Ang18, Ang24, OO13]. Wasserstein [ÇJM^{+}21]. wavelet [LS04]. Waves [Div91]. Way [BF91, DO06, RR12]. ways [HKS21]. Weak [ABL93, Ric99, HDHX17]. weakly [Li04, LLW24]. web [KBRV24, HGKV11, SWF11]. Web-based [SWF11]. WebDSL [HGKV11]. Wedderburn [OdR03, OdR09]. wedderga [OdR09]. Weight [Bre86, Joh15, dG01, BM10, MPS16]. Weighted [Rob88, Coo09, DP24, DS16, FEV16, MES19, Qur17]. Weights [MS00a, Gal16]. Weil [HKYY18]. Weingarten [GK21]. Weispfenning [CM17a, NP20]. Well [Les92, CMR19]. Wen [GK12a]. Wen-Tsun [GK12a]. Weyl [DdG21, HSS02, QR07, Tsa00]. Where [Hre94, DMY16]. whether [BGMSG07, BM04]. Which [Arn95, Bru01, LLW03, Ous91]. Whitehead [MH06]. Whittaker [Che23]. Who [BCE^{+}94]. whose [BFMS87, DEP22, HLSW16]. WhyMP [MRH23]. width [Roc22]. Wiedemann [HJS16, HJS22, Tho02]. Wild [von90b, vzG13]. Wilf [CK19, Mao21]. Williamson [BKG21, KK09]. Wilson [FKT13]. within [BFK02]. without [Bec03, CGK^{+}21, KN11, ZWH11]. witness [vdH06]. Witt [CIL07, Sut16]. Word [AP89, EHR91, JM93, KR91, Sta89, Wid01, Wra88, Yap91, MO85]. words [DS15]. Worksheets [Mon97]. world [AKR11]. Worst [Sha90b, KS12b]. worst-case [KS12b]. wreath [BNRW22, FMR04, PV05]. Write [SR86]. Wronskian [KPT15]. Wu [GK12a, Ric91, Ric99]. WZ [CX09, Ges95, Tef02].

XL [MO21]. XYZ [Sch94].

Yang [GIM07]. Yau [BR13c, Hie16]. years [Laz09]. yielding [Tsa23]. Yoshida [GLLdR21]. Young [Wil95].

Zacharias [NP20, Mor20]. Zassenhaus [Ano87]. Zeilberger [CHM05, CHM12, CK19, GG92, Mao21, MZ05, PS95b]. Zermelo [Win06]. Zero [Chi96, FGLM93, GC93, Kal02, Laz92b, Mon02a, NY99, PV00, PV02, Ric97, Tak92, AKR05, AP11a, Buc06a, CGY09, CGG12, CJ15, Dur09, HOP06, HKPP09, HKYY18, KMH89, Li04, MRSW07, MP04, Mos08, NT17, PS13, Qi23, Sek11, ST20, TBS17, Wal03, XWL23, vdHS06]. Zero-Characteristic [Chi96]. Zero-Dimensional [FGLM93, Laz92b, Mon02a, NY99, AKR05, CGY09, CGG12, CJ15, Dur09, HOP06, HKPP09, HKYY18, KMH89, MP04, Mos08, NT17, PS13, ST20, XWL23]. zero-locus [TBS17]. zero-test [vdHS06]. Zeros [GLW99, HS97, Wor94, Yam94, BM10, CPR09, CPR11, GS03, LLZ24, Lou08, Rap06, XWL23]. zeta [BM10, JMV18, Sto17, XW20]. Zhang [Yes21]. Zindler [Roc22]. Zippel [Lan92]. zonotope [Fuk04]. zur [GP12].

References

Ali:2018:EAA

[AAB^{+}18] Sajid Ali, Hassan Azad, Indranil Biswas, Ryad Ghanam, and M. T. Mustafa. Embedding algorithms and applications to differential equations. Journal of Symbolic Computation, 86(??):166–188, May/ June 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                                     //www.sciencedirect.com/
                                                     science/article/pii/S0747717117300512.

Ali:2021:CMD

[AABdG21] Sajid Ali, Hassan Azad, Indranil Biswas, and Willem A. de Graaf. A constructive method for decomposing real representations. Journal of Symbolic Computation, 104 (??):328–342, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300547.

Alvarez:2018:GBC

[ÁAF^{+}18] Víctor Álvarez, José Andrés Armario, Raúl M. Falcón, María Dolores Frau, and Félix Gudiel. Gröbner bases and cocyclic Hadamard matrices. Journal of Symbolic Computation, 89(??): 26–40, November/December 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117301098.

Alvarez:2009:HRM

[AAFR09] V. Alvarez, J. A. Armario, M. D. Frau, and P. Real. The homological reduction method for computing cocyclic Hadamard matrices. Journal of Symbolic Computation, 44(5): 558–570, May 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alvandi:2021:EHC

[AAKM21] Parisa Alvandi, Masoud Ataei, Mahsa Kazemi, and Marc Moreno Maza. On the extended Hensel construction and its application to the computation of real limit points. Journal of Symbolic Computation, 98(??):120–162, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300744.

Autin:1989:SEI

[AB89] B. Autin and J. Bengtsson. Symbolic evaluation of integrals occurring in accelerator orbit theory. Journal of Symbolic Computation, 7(2):183– 188 (or 183–187??), February 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Adams:1992:SRG

[AB92] William W. Adams and Ann K. Boyle. Some results on Gröbner bases over commutative rings. Journal of Symbolic Computation, 13(5):473–484, May 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Avenhaus:1999:FOE

[AB99] J. Avenhaus and K. Becker. A framework for operational equational specifications with pre-defined structures. Journal of Symbolic Computation, 27(3):271–310, March 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anderson:2000:PDS

[AB00a] Penny Anderson and David Basin. Program development schemata as derived rules. Journal of Symbolic Computation, 30(1):5–36, July 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0346;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0346/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0346/ref.

Aravindan:2000:TPT

[AB00b] Chandrabose Aravindan and Peter Baumgartner. Theorem proving techniques for view deletion in databases. Journal of Symbolic Computation, 29(2):119–147, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0358;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0358/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0358/ref.

Ayari:2001:HOI

[AB01] Abdelwaheb Ayari and David Basin. A higher-order interpretation of deductive tableau. Journal of Symbolic Computation, 31(5):487– 520, May 1, 2001. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0444;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0444/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0444/ref.

Armando:2005:REM

[AB05] Alessandro Armando and Clemens Ballarin. A reconstruction and extension of Maple’s assume facility via constraint contextual rewriting. Journal of Symbolic Computation, 39(5):503–521, May 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Abramov:2009:DSS

[AB09] S. A. Abramov and M. A. Barkatou. D’Alembertian series solutions at ordinary points of LODE with polynomial coefficients. Journal of Symbolic Computation, 44(1): 48–59, January 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Adamus:2022:ASP

[AB22] Elzbieta Adamus and Paweł Bogdan. Algorithm for studying polynomial maps and reductions modulo prime number. Journal of Symbolic Computation, 112(??):1– 24, September/October 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000821.

Abbott:2012:TFA

[Abb12]  John Abbott.  Twin-float

arithmetic. Journal of Symbolic Computation, 47(5): 536–551, May 2012. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001970.

Abbott:2013:BFI

[Abb13] John Abbott. Bounds on factors in Z[x]. Journal of Symbolic Computation, 50 (??):532–563, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001691.

Abbott:2017:FTM

[Abb17] John Abbott. Fault-tolerant modular reconstruction of rational numbers. Journal of Symbolic Computation, 80 (part 3)(??):707–718, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300773.

Amendola:2019:MLD

[ABB^{+}19] Carlos Améndola, Nathan Bliss, Isaac Burke, Courtney R. Gibbons, Martin Helmer, Serkan Hoşten, Evan D. Nash, Jose Israel Rodriguez, and Daniel Smolkin. The maximum likelihood degree of toric varieties. Journal of Symbolic Computation, 92 (??):222–242, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300476.

Angelini:2022:CIC

[ABC22] Elena Angelini, Cristiano Bocci, and Luca Chiantini. Catalecticant intersections and confinement of decompositions of forms. Journal of Symbolic Computation, 109 (??):220–237, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300584.

Augot:2009:DBC

[ABF09] Daniel Augot, Magali Bardet, and Jean-Charles Faugère. On the decoding of binary cyclic codes with the Newton identities. Journal of Symbolic Computation, 44(12):1608– 1625, December 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alpuente:2015:ECR

[ABFS15] M. Alpuente, D. Ballis, F. Frechina, and J. Sapiña. Exploring conditional rewriting logic computations. Journal of Symbolic Computation, 69(??):3–39, July/August 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000960.

Abramov:2015:FRD

[ABK15] S. A. Abramov, M. A. Barkatou, and D. E. Khmelnov. On full rank differential systems with power series coefficients. Journal of Symbolic Computation, 68 (part 1) (??):120–137, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000601.

Abbott:2000:CIP

[ABKR00] J. Abbott, A. Bigatti, M. Kreuzer, and L. Robbiano. Computing ideals of points. Journal of Symbolic Computation, 30(4):341–356, October 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0410;   http://www.
idealibrary.com/links/doi/

                                        10.1006/jsco.1999.0410/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0411;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.0411/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0411/ref.

Adams:1993:TWS

[ABL93] W. W. Adams, A. Boyle, and P. Loustaunau. Transitivity for weak and strong Gröbner bases. Journal of Symbolic Computation, 15(1):49–66 (or 49–65??), January 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Asadi:2023:PTD

[ABM^{+}23] Mohammadali Asadi, Alexander Brandt, Robert H. C. Moir, Marc Moreno Maza, and Yuzhen Xie. Parallelization of triangular decompositions: Techniques and implementation. Journal of Symbolic Computation, 115(??): 371–406, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000815.

Araujo:2010:CAS

[ABMN10] J. Araújo, P. V. Bünau, J. D. Mitchell, and M. Neunhöffer. Computing automorphisms of semigroups. Journal of Symbolic Computation, 45(3):373– 392, March 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Abo:2010:NDC

[Abo10] Hirotachi Abo. On nondefectivity of certain Segre– Veronese varieties. Journal of Symbolic Computation, 45(12):1254–1269, December 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Andreoli:1996:CBK

[ABP96] Jean-Marc Andreoli, Uwe M.

Borghoff, and Remo Pareschi. The constraint-based knowledge broker model: Semantics, implementation and analysis. Journal of Symbolic Computation, 21(4/5/6):635– 668 (or 635–667??), April, May & June 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Abbott:2021:CUM

[ABPR21] John Abbott, Anna Maria Bigatti, Elisa Palezzato, and Lorenzo Robbiano. Computing and using minimal polynomials. Journal of Symbolic Computation, 100 (??):137–163, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300884.

Abramov:2021:RHS

[ABPS21] Sergei A. Abramov, Manuel Bronstein, Marko Petkovsek, and Carsten Schneider. On rational and hypergeometric solutions of linear ordinary difference equations in ΠΣ^{∗}-field extensions. Journal of Symbolic Computation, 107(??): 23–66, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000080.

Abbott:2017:IH

[ABR17a] John Abbott, Anna Maria Bigatti, and Lorenzo Robbiano. Implicitization of hypersurfaces. Journal of Symbolic Computation, 81(??): 20–40, July/August 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301213.

Abramov:2017:RIF

[Abr17b] S. A. Abramov. On ramification indices of formal solutions of constructive linear ordinary differential systems. Journal of Symbolic Computation, 79 (part 2)(??): 475–481, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300013.

Abramov:2011:SSL

[ABvHP11] S. A. Abramov, M. A. Barkatou, M. van Hoeij, and M. Petkovsek. Subanalytic solutions of linear difference equations and multidimensional hypergeometric sequences. Journal of Symbolic Computation, 46(11): 1205–1228, November 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001155.

Alevizos:1990:NCC

[ABY90] P. Alevizos, J. D. Boissonnat, and M. Yvinec. Non convex contour reconstruction. Journal of Symbolic Computation, 10(3–4): 225–252, September/October 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Aroca:2001:FSL

[AC01] F. Aroca and J. Cano. Formal solutions of linear PDEs and convex polyhedra. Journal of Symbolic Computation, 32(6):717–737, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0492;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0492/
pdf;   http://www.idealibrary.

                                        com/links/doi/10.1006/jsco.
                                        2001.0492/ref.

Andrei:2004:SCH

[AC04]  Ştefan Andrei and Wei-Ngan

Chin. Solving a class of higher-order equations over a group structure. Journal of Symbolic Computation, 37(3): 329–341, March 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Amata:2019:CGI

[AC19] Luca Amata and Marilena Crupi. Computation of graded ideals with given extremal Betti numbers in a polynomial ring. Journal of Symbolic Computation, 93(??): 120–132, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300506.

Ablett:2024:DHC

[AC24] Patience Ablett and Stephen Coughlan. Deformations of half-canonical Gorenstein curves in codimension four. Journal of Symbolic Computation, 121(??):Article 102251, March/April 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000652.

Albrecht:2012:RBM

[ACFP12] Martin R. Albrecht, Carlos Cid, Jean-Charles Faugère, and Ludovic Perret. On the relation between the MXL family of algorithms and Gröbner basis algorithms. Journal of Symbolic Computation, 47(8):926–941, August 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711200003X.

Abramov:2004:TCS

[ACGL04] S. A. Abramov, J. J. Carette, K. O. Geddes, and H. Q. Le. Telescoping in the context of symbolic summation in Maple. Journal of Symbolic Computation, 38 (4):1303–1326, October 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Armando:2001:CLO

[ACGR01] Alessandro Armando, Alessandro Coglio, Fausto Giunchiglia, and Silvio Ranise. The control layer in open mechanized reasoning systems: Annotations and tactics. Journal of Symbolic Computation, 32(4):305–332, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0464;   http://www.
idealibrary.com/links/doi/

                                        10.1006/jsco.2000.0464/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0464/ref.

Arnon:1988:AAC

[ACM88] Dennis S. Arnon, George E. Collins, and Scott McCallum. An adjacency algorithm for cylindrical algebraic decompositions of threedimensional space. Journal of Symbolic Computation, 5 (1–2):163–188 (or 163–187??), February/April 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alberich-Carraminana:2019:ECB

[ACMB19] Maria Alberich-Carramiñana, Josep Àlvarez Montaner, and Guillem Blanco. Effective computation of base points of ideals in two-dimensional local rings. Journal of Symbolic Computation, 92 (??):93–109, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300014.

Albano:2000:MGI

[ACOR00] G. Albano, F. Cioffi, F. Orecchia, and I. Ramella. Minimally generating ideals of rational parametric curves in polynomial time. Journal of Symbolic Computation, 30(2):137–149, August 2000. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0354;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0354/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0354/ref.

Arvanitoyeorgos:2013:PIH

[ACS13]  Andreas  Arvanitoyeorgos,

Ioannis Chrysikos, and Yusuke Sakane. Proving isometry for homogeneous Einstein metrics on flag manifolds by symbolic computation. Journal of Symbolic Computation, 55(??):59–71, August 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000497.

Ansola:2021:SSR

[ADCZ21] M. Ansola, A. Díaz-Cano, and M. A. Zurro. Semialgebraic sets and real binary forms decompositions. Journal of Symbolic Computation, 107(??):209–220, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000201.

Adler:2016:ITE

[Adl16] V. E. Adler. Integrability test for evolutionary lattice equations of higher order. Journal of Symbolic Computation, 74(??):125–139, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000565.

Abramov:2002:RNF

[Ae02] S. A. Abramov and M. Petkov ek. Rational normal forms and minimal decompositions of hypergeometric terms. Journal of Symbolic Computation, 33(5):521–543, May 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See erratum [AP04].

Assmann:2005:CPP

[AE05] Björn Assmann and Bettina Eick. Computing polycyclic presentations for polycyclic rational matrix groups. Journal of Symbolic Computation, 40(6):1269–1284, December 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arnborg:1988:ADR

[AF88] Stefan Arnborg and Huichun Feng. Algebraic decomposition of regular curves. Journal of Symbolic Computation, 5(1–2):131–140, February/April 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Attardi:1996:MMP

[AF96] Giuseppe Attardi and Tito Flagella. Memory management in the PoSSo solver. Journal of Symbolic Computation, 21(3):293–312 (or 293– 311??), March 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Acciaro:2000:FNI

[AF00] Vincenzo Acciaro and Claus Fieker. Finding normal integral bases of cyclic number fields of prime degree. Journal of Symbolic Computation, 30(2):129–136, August 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0335;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0335/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0335/ref.

Ayad:2008:DRF

[AF08] Mohamed Ayad and Peter Fleischmann. On the decomposition of rational functions. Journal of Symbolic Computation, 43(4):259–274, April 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Albert:2015:FRI

[AFdCS15] Mario Albert, Matthias Fetzer, Eduardo Sáenz de Cabezón, and Werner M. Seiler. On the free resolution induced by a Pommaret basis. Journal of Symbolic Computation, 68 (part 2)(??):4–26, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000765.

Agostini:2023:KST

[AFMS23] Daniele Agostini, Claudia Fevola, Yelena Mandelshtam, and Bernd Sturmfels. KP solitons from tropical limits. Journal of Symbolic Computation, 114(??):282– 301, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000293.

Augot:2009:F

[AFP09] Daniel Augot, Jean-Charles Faugère, and Ludovic Perret. Foreword. Journal of Symbolic Computation, 44(12):1605– 1607, December 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Abbott:2008:SBB

[AFT08] John Abbott, Claudia Fassino, and Maria-Laura Torrente. Stable border bases for ideals of points. Journal of Symbolic Computation, 43(12): 883–894, December 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Andersen:1991:IHP

[AG91] Carl M. Andersen and James F.

Geer. Investigating a hybrid perturbation-Galerkin technique using computer algebra. Journal of Symbolic Computation, 12(6):695–714, December 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Amrhein:1997:VMC

[AGM97] B. Amrhein, O. Gloor, and R. E. Maeder. Visualizations for mathematics courses based on a computer algebra system. Journal of Symbolic Computation, 23(5–6):447–452, May/ June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Aitken:1998:ITP

[AGMT98] J. S. Aitken, P. Gray, T. Melham, and M. Thomas. Interactive theorem proving: An empirical study of user activity. Journal of Symbolic Computation, 25(2):263– 284, February 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alonso:1995:RFD

[AGR95] Cesar Alonso, Jaime Gutierrez, and Tomas Recio. A rational function decomposition algorithm by near-separated polynomials. Journal of Symbolic Computation, 19(6): 527–544, June 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arnold:2016:FSM

[AGR16] Andrew Arnold, Mark Giesbrecht, and Daniel S. Roche. Faster sparse multivariate polynomial interpolation of straight-line programs. Journal of Symbolic Computation, 75(??):4–24, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001042.

Area:1999:IPH

[AGRZ99] I. Area, E. Godoy, A. Ronveaux, and A. Zarzo. Inversion problems in the q-Hahn tableau. Journal of Symbolic Computation, 28 (6):767–776, December 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0339/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0339/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0339/production/ref;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0340/production;
http://www.idealibrary.

com/links/doi/10.1006/jsco.
1998.0340/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0340/production/ref;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0338/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0338/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0338/production/ref.

Assi:2016:ACO

[AGS16] Abdallah Assi and Pedro A.

García-Sánchez. Algorithms for curves with one place at infinity. Journal of Symbolic Computation, 74(??): 475–492, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000905.

Allamigeon:2018:SGN

[AGS18a] Xavier Allamigeon, Stéphane Gaubert, and Mateusz Skomra. Solving generic nonarchimedean semidefinite programs using stochastic game algorithms. Journal of Symbolic Computation, 85(??): 25–54, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300664.

Allamigeon:2018:TAH

[AGS18b] Xavier Allamigeon, Stéphane Gaubert, and Mateusz Skomra. The tropical analogue of the Helton–Nie conjecture is true. Journal of Symbolic Computation, 91(??):129–148, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300828.

Assi:2017:BSI

[AGSM17] A. Assi, P. A. García-Sánchez, and V. Micale. Bases of subalgebras of K[[x]] and K[x]. Journal of Symbolic Computation, 79 (part 1) (??):4–22, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300815.

Allem:2013:ESF

[AGT13] Luiz Emilio Allem, Shuhong Gao, and Vilmar Trevisan. Extracting sparse factors from multivariate integral polynomials. Journal of Symbolic Computation, 52(??):3–16, May 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001265.

Atkinson:1986:CGC

[AH86] M. D. Atkinson and R. A. Hassan. On the computation of group characters. Journal of Symbolic Computation, 2 (1):45–50, March 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Adleman:2001:CPC

[AH01] Leonard M. Adleman and Ming-Deh Huang. Counting points on curves and Abelian varieties over finite fields. Journal of Symbolic Computation, 32 (3):171–189, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0470;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0470/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0470/ref.

Apel:2005:DUR

[AH05] Joachim Apel and Ralf Hemmecke. Detecting unnecessary reductions in an involutive basis computation. Journal of Symbolic Computation, 40(4–5):1131–1149, October/November 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Agashe:2013:CIN

[AH13] Amod Agashe and Randy Heaton. Computing intersection numbers between abelian varieties associated to subspaces of modular forms. Journal of Symbolic Computation, 57(??):70–77, October 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000692.

Alexandr:2023:LVC

[AH23] Yulia Alexandr and Serkan Hosten. Logarithmic Voronoi cells for Gaussian models. Journal of Symbolic Computation, 122(??):Article 102256, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000706.

Albert:2005:OSD

[AHH^{+}05] Elvira Albert, Michael Hanus, Frank Huch, Javier Oliver, and Germán Vidal. Operational semantics for declarative multi-paradigm languages. Journal of Symbolic Computation, 40(1):795–829, July 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anai:2009:PPS

[AHKY09] Hirokazu Anai, Shinji Hara, Masaaki Kanno, and Kazuhiro Yokoyama. Parametric polynomial spectral factorization using the sum of roots and its application to a control design problem. Journal of Symbolic Computation, 44(7):703–725, July 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Avenhaus:2003:UGJ

[AHL03] J. Avenhaus, Th. Hillenbrand, and B. Löchner. On using ground joinable equations in equational theorem proving. Journal of Symbolic Computation, 36(1–2): 217–233, July/August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Adams:1999:SSG

[AHLM99] W. W. Adams, S. Hoşten, P. Loustaunau, and J. L. Miller. SAGBI and SAGBI-Gröbner bases over principal ideal domains. Journal of Symbolic Computation, 27(1): 31–48, January 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alcazar:2018:SDR

[AHM18] Juan Gerardo Alcázar, Carlos Hermoso, and Georg Muntingh. Similarity detection of rational space curves. Journal of Symbolic Computation, 85 (??):4–24, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300652.

Ahn:2008:DCD

[Ahn08] Jeaman Ahn. The degreecomplexity of the defining ideal of a smooth integral curve. Journal of Symbolic Computation, 43(6–7):422– 441, June/July 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Akoglu:2018:CSO

[AHS18] Tulay Ayyildiz Akoğlu, Jonathan D. Hauenstein, and Agnes Szanto. Certifying solutions to overdetermined and singular polynomial systems over Q. Journal of Symbolic Computation, 84(??):147–171, January/February 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300329.

Aramideh:2021:CRR

[AHS21] Nasibeh Aramideh, Amir Hashemi, and Werner M. Seiler. Computing the resolution regularity of bihomogeneous ideals. Journal of Symbolic Computation, 103 (??):141–156, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711930149X.

Adkins:2005:EPS

[AHW05] William A. Adkins, J. William Hoffman, and Hao Hao Wang.

Equations of parametric surfaces with base points via syzygies. Journal of Symbolic Computation, 39(1):73– 101, January 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Avis:1990:LRA

[AI90] David Avis and Hiroshi Imai.

Locating a robot with angle measurements. Journal of Symbolic Computation, 10(3–4):311–326, September/ October 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Avendano:2012:FAF

[AIRR12] Martín Avendaño, Ashraf Ibrahim, J. Maurice Rojas, and Korben Rusek. Faster padic feasibility for certain multivariate sparse polynomials. Journal of Symbolic Computation, 47(4):454–479, April 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001428.

Armando:2001:SIC

[AJ01] Alessandro Armando and Tudor Jebelean. Special issue on Calculemus-99: Integrating computation and deduction foreword of the Guest Editors. Journal of Symbolic Computation, 32 (4):303–304, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0467;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0467/
                                        pdf.

Aigner:2009:PSC

[AJGVS09] Martin Aigner, Bert Jüttler, Laureano Gonzalez-Vega, and Josef Schicho. Parameterizing surfaces with certain special support functions, including offsets of quadrics and rationally supported surfaces. Journal of Symbolic Computation, 44(2):180–191, February 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Armbruster:1986:CUU

[AK86] D. Armbruster and H. Kredel. Constructing universal unfoldings using Gröbner bases. Journal of Symbolic Computation, 2(4):383– 388, December 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Adi:1992:AUR

[AK92] Mohamed Adi and Claude Kirchner. AC-unification race: The system solving approach, implementation and benchmarks. Journal of Symbolic Computation, 14(1):51–70, July 1992. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Acciaro:2000:CLA

[AK00] Vincenzo Acciaro and Jürgen Klüners. Computing local Artin maps, and solvability of norm equations. Journal of Symbolic Computation, 30(3):239–252, September 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0361;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0361/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0361/ref.

Ahn:2004:CPC

[AK04] Min-Ho Ahn and Gwang-Il Kim. Characterization of Pythagorean curves and Pythagoreanization using a rational transform. Journal of Symbolic Computation, 37(3): 377–389, March 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See retraction [AK06].

Ahn:2006:RNC

[AK06] Min-Ho Ahn and Gwang-Il Kim. Retraction notice to “Characterization of Pythagorean curves and Pythagoreanization using a rational transform” [J. Symbolic Comput. 37 (3) (2004) 377–389]. Journal of Symbolic Computation, 41(1):122, January 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [AK04].

Allili:2017:RCM

[AKL17] Madjid Allili, Tomasz Kaczynski, and Claudia Landi. Reducing complexes in multidimensional persistent homology theory. Journal of Symbolic Computation, 78(??):61– 75, January/February 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300086.

Abbott:2005:CZD

[AKR05] J. Abbott, M. Kreuzer, and L. Robbiano. Computing zero-dimensional schemes. Journal of Symbolic Computation, 39(1):31–49, January 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arnold:2011:BBW

[AKR11] Elizabeth Arnold, Ilias Kotsireas, and Markus Rosenkranz. Bruno Buchberger and the world of Gröbner bases. Journal of Symbolic Computation, 46(5):495–497, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001707.

Ahn:2012:DCS

[AKS12] Jeaman Ahn, Sijong Kwak, and Yeongseok Song. The degree complexity of smooth surfaces of codimension 2. Journal of Symbolic Computation, 47(5):568–581, May 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000028.

Apel:1988:EBA

[AL88] J. Apel and W. Lassner. An extension of Buchberger’s algorithm and calculations in enveloping fields of Lie algebras. Journal of Symbolic Computation, 6(2–3):361–370, October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

AlvarezMontaner:2006:CSL

[ÀL06] Josep Àlvarez Montaner and Anton Leykin. Computing the support of local cohomology modules. Journal of Symbolic Computation, 41(12):1328–1344, December 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alonso:2010:LBT

[AL10] M. Emilia Alonso and Henri Lombardi. Local Bézout theorem. Journal of Symbolic Computation, 45(10):

975–985, October 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alcazar:2008:GGB

[Alc08a] Juan Gerardo Alcazar. Good global behavior of offsets to plane algebraic curves. Journal of Symbolic Computation, 43(9):659–680, September 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alcazar:2008:GLB

[Alc08b] Juan Gerardo Alcazar. Good local behavior of offsets to rational regular algebraic surfaces. Journal of Symbolic Computation, 43(12): 845–857, December 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alcazar:2012:LSG

[Alc12]  Juan G. Alcazar.  Local

shape of generalized offsets to algebraic curves. Journal of Symbolic Computation, 47(3):327–341, March 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001933.

Alcantara:2016:SSF

[Alc16] Claudia R. Alcántara. Stratification of the space of foliations on CP^{2}. Journal of Symbolic Computation, 72(??):

147–160, January/February 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000127.

Aldossari:2023:CPF

[Ald23] Shayea Aldossari. Computing pullback function of second order differential operators by using their semiinvariants. Journal of Symbolic Computation, 119(??): 38–49, November/December 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000160.

Aubry:1999:TTS

[ALM99] Philippe Aubry, Daniel Lazard, and Marc Moreno Maza. On the theories of triangular sets. Journal of Symbolic Computation, 28 (1–2):105–124, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0269;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0269/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0269/ref.

Aluffi:2003:CCC

[Alu03] Paolo Aluffi. Computing characteristic classes of projective schemes. Journal of Symbolic Computation, 35(1): 3–19, January 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arnon:1988:PTA

[AM88a] Dennis S. Arnon and Scott McCallum. A polynomialtime algorithm for the topological type of a real algebraic curve. Journal of Symbolic Computation, 5(1–2):213–236, February/April 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arnon:1988:MQE

[AM88b] Dennis S. Arnon and Maurice Mignotte. On mechanical quantifier elimination for elementary algebra and geometry. Journal of Symbolic Computation, 5(1–2):237–260 (or 237–259??), February/April 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Abadi:1989:TLP

[AM89] Martin Abadi and Zohar Manna. Temporal logic programming. Journal of Symbolic Computation, 8(3):277– 296 (or 277–295??), September 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Aubry:1999:TSS

[AM99] Philippe Aubry and Marc Moreno Maza. Triangular sets for solving polynomial systems: a comparative implementation of four methods. Journal of Symbolic Computation, 28 (1–2):125–154, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0270;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0270/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0270/ref.

Aparicio-Monforte:2016:LIE

[AMDW16] A. Aparicio-Monforte, T. Dreyfus, and J.-A. Weil. Liouville integrability: an effective Morales–Ramis–Simó theorem. Journal of Symbolic Computation, 74(??): 537–560, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000942.

Aljovin:2019:IHC

[AML19] Enzo Aljovin, Hossein Movasati, and Roberto Villaflor Loyola. Integral Hodge conjecture for Fermat varieties. Journal of Symbolic Computation, 95(??):177–184, November/December 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300173.

Alberti:2009:ITR

[AMT09] Lionel Alberti, Bernard Mourrain, and Jean-Pierre Técourt. Isotopic triangulation of a real algebraic surface. Journal of Symbolic Computation, 44(9): 1291–1310, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Aparicio-Monforte:2012:RFL

[AMW12] Ainhoa Aparicio-Monforte and Jacques-Arthur Weil. A reduced form for linear differential systems and its application to integrability of Hamiltonian systems. Journal of Symbolic Computation, 47(2):192–213, February 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001465.

Andrews:1995:CPB

[And95] George E. Andrews. On a conjecture of Peter Borwein. Journal of Symbolic Computation, 20(5–6):487–502 (or 487– 501??), November/December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Angermuller:2015:TSG

[Ang15] Gerhard Angermüller. Triangular systems and a generalization of primitive polynomials. Journal of Symbolic Computation, 68 (part 1) (??):316–325, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711400090X.

Angelini:2018:WDI

[Ang18] Elena Angelini. Waring decompositions and identifiability via Bertini and Macaulay2 software. Journal of Symbolic Computation, 91(??):200–212, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300865.

Angelini:2024:CCS

[Ang24] Elena Angelini. A counterexample to a conjecture on simultaneous Waring identifiability. Journal of Symbolic Computation, 120(??):Article 102223, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000305.

Anonymous:1987:BHZ

[Ano87]  Anonymous.  Bibliography

of Hans Zassenhaus. Journal of Symbolic Computation, 4(1):129–135, 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:1996:ETL

[Ano96] Anonymous. Executable temporal logic systems. Journal of Symbolic Computation, 22(5–6):721–735, November/ December 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:1999:CV

[Ano99a] Anonymous. Contents of volume 28. Journal of Symbolic Computation, 28(6):i– ii, December 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0329/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0329/production/pdf.

Anonymous:1999:IV

[Ano99b] Anonymous. Index to volume 28. Journal of Symbolic Computation, 28(6):iii– viii, December 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0331/production;
http://www.idealibrary.

com/links/doi/10.1006/jsco.
1999.0331/production/pdf.

Anonymous:1999:SIO

[Ano99c] Anonymous. Special issue on orthogonal polynomials and computer algebra foreword of the Guest Editors. Journal of Symbolic Computation, 28(6):737–738, December 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0300/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0300/production/pdf.

Anonymous:2000:CIV

[Ano00a] Anonymous. Contents and index of volume 30. Journal of Symbolic Computation, 30(6):1–7, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0432;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0432/
pdf.

Anonymous:2000:SIA

[Ano00b] Anonymous. Special issue on advances in first-order theorem proving foreword of the Guest Editors. Journal of Symbolic Computation, 29(2):117–118, February 2000. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0357;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0357/
                                        pdf.

Anonymous:2001:CIV

[Ano01a] Anonymous. Contents and index to volume 31. Journal of Symbolic Computation, 31(6):i–vii, June 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0475;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0475/
                                        pdf.

Anonymous:2001:JSC

[Ano01b] Anonymous. Journal of Symbolic Computation, volume 32 contents and author index. Journal of Symbolic Computation, 32(6):i–vii, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0507;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0507/
                                        pdf.

Anonymous:2001:SI

[Ano01c] Anonymous. Special issue. Journal of Symbolic Computation, 32(1–2), July/August 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0458;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0458/
pdf.

Anonymous:2001:SIC

[Ano01d] Anonymous. Special issue on Calculemus-99: Integrating computation and deduction. Journal of Symbolic Computation, 32(4):??, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2001:SIE

[Ano01e] Anonymous. Special issue on effective methods in rings of differential operators. Journal of Symbolic Computation, 32 (6):??, December 1, 2001. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2001:VCA

[Ano01f] Anonymous. Volume 32 contents and author index. Journal of Symbolic Computation, 32(6):??, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2002:A

[Ano02a] Anonymous. Applications.

Journal of Symbolic Computation, 33(3):??, March 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2002:CIV

[Ano02b] Anonymous. Contents and index of volume 34. Journal of Symbolic Computation, 34 (6):i–v, December 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2002:JSC

[Ano02c] Anonymous. Journal of Symbolic Computation, volume 33 contents and author index. Journal of Symbolic Computation, 34(1):i–vii, July 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:A

[Ano03a] Anonymous. Advert. Journal of Symbolic Computation, 35(4):CO3, April 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:EBa

[Ano03b] Anonymous. Editorial Board. Journal of Symbolic Computation, 35(1):CO2, January 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:EBb

[Ano03c] Anonymous. Editorial Board. Journal of Symbolic Computation, 35(2):CO2, February 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:EBc

[Ano03d] Anonymous. Editorial Board.

Journal of Symbolic Computation, 35(3):CO2, March 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:EBd

[Ano03e] Anonymous. Editorial Board.

Journal of Symbolic Computation, 35(4):CO2, April 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:EBe

[Ano03f] Anonymous. Editorial Board.

Journal of Symbolic Computation, 35(5):CO2, May 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:EBf

[Ano03g] Anonymous. Editorial Board.

Journal of Symbolic Computation, 35(6):CO2, June 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:EBg

[Ano03h] Anonymous. Editorial Board.

Journal of Symbolic Computation, 36(1–2):CO2, July/ August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:EBh

[Ano03i] Anonymous. Editorial Board. Journal of Symbolic Computation, 36(3–4):CO2, September/October 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:EBi

[Ano03j] Anonymous. Editorial Board. Journal of Symbolic Computation, 36(5):CO2, November 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:EBj

[Ano03k] Anonymous. Editorial Board. Journal of Symbolic Computation, 36(6):CO2, December 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2003:PN

[Ano03l] Anonymous. Publisher’s note. Journal of Symbolic Computation, 35(1):1–2, January 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:C

[Ano04a] Anonymous. Contents. Journal of Symbolic Computation, 37(1):121–125, January 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:CV

[Ano04b] Anonymous. Contents of volumes 37 and 38. Journal of Symbolic Computation, 38 (6):??, December 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBa

[Ano04c] Anonymous. Editorial Board.

Journal of Symbolic Computation, 37(1):1, January 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBb

[Ano04d] Anonymous. Editorial Board.

Journal of Symbolic Computation, 37(2):CO2, February 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBc

[Ano04e] Anonymous. Editorial Board.

Journal of Symbolic Computation, 37(3):CO2, March 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBd

[Ano04f] Anonymous. Editorial Board.

Journal of Symbolic Computation, 37(4):CO2, April 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBe

[Ano04g] Anonymous. Editorial Board. Journal of Symbolic Computation, 37(5):CO2, May 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBf

[Ano04h] Anonymous. Editorial Board. Journal of Symbolic Computation, 37(6):CO2, June 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBg

[Ano04i] Anonymous. Editorial Board. Journal of Symbolic Computation, 38(1):CO2, July 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBh

[Ano04j] Anonymous. Editorial Board. Journal of Symbolic Computation, 38(2):CO2, August 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBi

[Ano04k] Anonymous. Editorial Board. Journal of Symbolic Computation, 38(3):CO2, September 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBj

[Ano04l] Anonymous. Editorial Board.

Journal of Symbolic Computation, 38(4):CO2, October 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBk

[Ano04m] Anonymous. Editorial Board.

Journal of Symbolic Computation, 38(5):CO2, November 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:EBl

[Ano04n] Anonymous. Editorial Board.

Journal of Symbolic Computation, 38(6):CO2, December 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:I

[Ano04o] Anonymous. Index. Journal of Symbolic Computation, 37(1):127–132, January 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2004:IV

[Ano04p] Anonymous. Index to volumes 37 and 38. Journal of Symbolic Computation, 38 (6):??, December 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2005:EBa

[Ano05a] Anonymous. Editorial Board. Journal of Symbolic Computation, 39(1):CO2, January 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2005:EBb

[Ano05b] Anonymous. Editorial Board. Journal of Symbolic Computation, 39(2):CO2, February 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2005:EBc

[Ano05c] Anonymous. Editorial Board. Journal of Symbolic Computation, 39(3–4):CO2, March/ April 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2005:EBd

[Ano05d] Anonymous. Editorial Board. Journal of Symbolic Computation, 39(5):CO2, May 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2005:EBe

[Ano05e] Anonymous. Editorial Board. Journal of Symbolic Computation, 39(6):CO2, June 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2005:I

[Ano05f] Anonymous. Index. Journal of Symbolic Computation, 40 (6):vi–x, December 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2006:IV

[Ano06] Anonymous. Index to Volume 41. Journal of Symbolic Computation, 41(12):??, December 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2011:EBa

[Ano11a] Anonymous. Editorial Board.

Journal of Symbolic Computation, 46(1):??, January 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001823.

Anonymous:2011:EBb

[Ano11b] Anonymous. Editorial Board.

Journal of Symbolic Computation, 46(2):??, February 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001938.

Anonymous:2011:EBc

[Ano11c] Anonymous. Editorial Board.

Journal of Symbolic Computation, 46(3):??, March 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110002026.

Anonymous:2011:EBd

[Ano11d] Anonymous. Editorial Board. Journal of Symbolic Computation, 46(4):??, April 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000034.

Anonymous:2011:EBe

[Ano11e] Anonymous. Editorial Board. Journal of Symbolic Computation, 46(5):??, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000289.

Anonymous:2011:EBf

[Ano11f] Anonymous. Editorial Board. Journal of Symbolic Computation, 46(6):??, June 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000447.

Anonymous:2011:EBg

[Ano11g] Anonymous. Editorial Board. Journal of Symbolic Computation, 46(7):??, July 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000526.

Anonymous:2011:EBh

[Ano11h] Anonymous. Editorial Board.

Journal of Symbolic Computation, 46(8):??, August 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000666.

Anonymous:2011:EBi

[Ano11i] Anonymous. Editorial Board.

Journal of Symbolic Computation, 46(9):??, September 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000897.

Anonymous:2011:EBj

[Ano11j] Anonymous. Editorial Board.

Journal of Symbolic Computation, 46(10):??, October 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001039.

Anonymous:2011:EBk

[Ano11k] Anonymous. Editorial Board.

Journal of Symbolic Computation, 46(11):??, November 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001490.

Anonymous:2011:EBl

[Ano11l] Anonymous. Editorial Board.

Journal of Symbolic Computation, 46(12):??, December 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001593.

Anonymous:2012:EBa

[Ano12a] Anonymous. Editorial Board. Journal of Symbolic Computation, 47(1):??, January 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001696.

Anonymous:2012:EBb

[Ano12b] Anonymous. Editorial Board. Journal of Symbolic Computation, 47(2):??, February 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001805.

Anonymous:2012:EBc

[Ano12c] Anonymous. Editorial Board. Journal of Symbolic Computation, 47(3):??, March 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002367.

Anonymous:2012:EBd

[Ano12d] Anonymous. Editorial Board. Journal of Symbolic Computation, 47(4):??, April 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000077.

Anonymous:2012:EBe

[Ano12e] Anonymous. Editorial Board.

Journal of Symbolic Computation, 47(5):??, May 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000193.

Anonymous:2012:EBf

[Ano12f] Anonymous. Editorial Board.

Journal of Symbolic Computation, 47(6):??, June 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000284.

Anonymous:2012:EBg

[Ano12g] Anonymous. Editorial Board.

Journal of Symbolic Computation, 47(7):??, July 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000405.

Anonymous:2012:EBh

[Ano12h] Anonymous. Editorial Board.

Journal of Symbolic Computation, 47(8):??, August 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000491.

Anonymous:2012:EBi

[Ano12i] Anonymous. Editorial Board. Journal of Symbolic Computation, 47(9):??, September 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112000570.

Anonymous:2012:EBj

[Ano12j] Anonymous. Editorial Board. Journal of Symbolic Computation, 47(10):??, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112000703.

Anonymous:2012:EBk

[Ano12k] Anonymous. Editorial Board. Journal of Symbolic Computation, 47(11):??, November 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112000855.

Anonymous:2012:EBl

[Ano12l] Anonymous. Editorial Board. Journal of Symbolic Computation, 47(12):??, December 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112000971.

Anonymous:2012:PN

[Ano12m] Anonymous. Publisher’s note. Journal of Symbolic Computation, 47(2):??, February 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001866.

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial Board.

Journal of Symbolic Computation, 48(??):??, January 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001447.

Anonymous:2013:EBb

[Ano13b] Anonymous. Editorial Board.

Journal of Symbolic Computation, 49(??):??, February 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001605.

Anonymous:2013:EBc

[Ano13c] Anonymous. Editorial Board.

Journal of Symbolic Computation, 50(??):??, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001800.

Anonymous:2013:EBd

[Ano13d] Anonymous. Editorial Board.

Journal of Symbolic Computation, 51(??):??, April 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001940.

Anonymous:2013:EBe

[Ano13e] Anonymous. Editorial Board. Journal of Symbolic Computation, 52(??):??, May 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000072.

Anonymous:2013:EBf

[Ano13f] Anonymous. Editorial Board. Journal of Symbolic Computation, 53(??):??, June 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000321.

Anonymous:2013:EBg

[Ano13g] Anonymous. Editorial Board. Journal of Symbolic Computation, 54(??):??, July 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000394.

Anonymous:2013:EBh

[Ano13h] Anonymous. Editorial Board. Journal of Symbolic Computation, 55(??):??, August 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000527.

Anonymous:2013:EBi

[Ano13i] Anonymous. Editorial Board.

Journal of Symbolic Computation, 56(??):??, September 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000746.

Anonymous:2013:EBj

[Ano13j] Anonymous. Editorial Board.

Journal of Symbolic Computation, 57(??):??, October 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000862.

Anonymous:2013:EBk

[Ano13k] Anonymous. Editorial Board.

Journal of Symbolic Computation, 58(??):??, November 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001004.

Anonymous:2013:EBl

[Ano13l] Anonymous. Editorial Board.

Journal of Symbolic Computation, 59(??):??, December 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001119.

Anonymous:2015:EBa

[Ano15a] Anonymous. Editorial Board.

Journal of Symbolic Computation, 66(??):ifc, January/February 2015. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000431.

Anonymous:2015:EBb

[Ano15b] Anonymous. Editorial Board. Journal of Symbolic Computation, 67(??):ifc, March/April 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000637.

Anonymous:2015:EBc

[Ano15c] Anonymous. Editorial Board. Journal of Symbolic Computation, 68 (part 1)(??):ifc, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001138.

Anonymous:2015:EBd

[Ano15d] Anonymous. Editorial Board. Journal of Symbolic Computation, 68 (part 2)(??):ifc, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001205.

Anonymous:2015:EBe

[Ano15e] Anonymous. Editorial Board. Journal of Symbolic Computation, 69(??):ifc, July/August 2015. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000036.

Anonymous:2015:EBf

[Ano15f] Anonymous. Editorial Board.

Journal of Symbolic Computation, 70(??):ifc, September/October 2015. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000218.

Anonymous:2015:EBg

[Ano15g] Anonymous. Editorial Board.

Journal of Symbolic Computation, 71(??):ifc, November/December 2015. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000346.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial Board.

Journal of Symbolic Computation, 72(??):ifc, January/February 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000620.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial Board.

Journal of Symbolic Computation, 73(??):ifc, March/April 2016. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000802.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial Board. Journal of Symbolic Computation, 74(??):ifc, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115001194.

Anonymous:2016:EBd

[Ano16d] Anonymous. Editorial Board. Journal of Symbolic Computation, 75(??):ifc, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115001339.

Anonymous:2016:EBe

[Ano16e] Anonymous. Editorial Board. Journal of Symbolic Computation, 76(??):ifc, September/October 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116000250.

Anonymous:2016:EBf

[Ano16f] Anonymous. Editorial Board. Journal of Symbolic Computation, 77(??):ifc, November/December 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300153.

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial Board.

Journal of Symbolic Computation, 78(??):ifc, January/February 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711630061X.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial Board.

Journal of Symbolic Computation, 79 (part 1)(??): ifc, March/April 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300967.

Anonymous:2017:EBc

[Ano17c] Anonymous. Editorial Board.

Journal of Symbolic Computation, 79 (part 2)(??): ifc, March/April 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301055.

Anonymous:2017:EBd

[Ano17d] Anonymous. Editorial Board.

Journal of Symbolic Computation, 79 (part 3)(??): ifc, March/April 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301122.

Anonymous:2017:EBe

[Ano17e] Anonymous. Editorial Board. Journal of Symbolic Computation, 80 (part 2)(??):ifc, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301481.

Anonymous:2017:EBf

[Ano17f] Anonymous. Editorial Board. Journal of Symbolic Computation, 80 (part 3)(??):ifc, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301559.

Anonymous:2017:EBg

[Ano17g] Anonymous. Editorial Board. Journal of Symbolic Computation, 81(??):ifc, July/August 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300032.

Anonymous:2017:EBh

[Ano17h] Anonymous. Editorial Board. Journal of Symbolic Computation, 82(??):ifc, September/October 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300160.

Anonymous:2017:EBi

[Ano17i] Anonymous. Editorial Board.

Journal of Symbolic Computation, 83(??):ifc, November/December 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300408.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board.

Journal of Symbolic Computation, 84(??):ifc, January/February 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711730086X.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board.

Journal of Symbolic Computation, 85(??):ifc, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300962.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board.

Journal of Symbolic Computation, 86(??):ifc, May/June 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117301049.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Journal of Symbolic Computation, 87(??):ii, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300051.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. Journal of Symbolic Computation, 88(??):ii, September/October 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300208.

Anonymous:2018:EBf

[Ano18f] Anonymous. Editorial Board. Journal of Symbolic Computation, 89(??):ii, November/December 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300543.

Anonymous:2018:EBg

[Ano18g] Anonymous. Editorial board. Journal of Symbolic Computation, 90(??):ii, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300609.

Anonymous:2018:EBh

[Ano18h] Anonymous. Editorial Board.

Journal of Symbolic Computation, 91(??):ii, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118301147.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board.

Journal of Symbolic Computation, 92(??):ii, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711830124X.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board.

Journal of Symbolic Computation, 93(??):ii, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300045.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board.

Journal of Symbolic Computation, 94(??):ii, September/October 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300343.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. Journal of Symbolic Computation, 95(??):ii, November/December 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300446.

Anonymous:2019:EBe

[Ano19e] Anonymous. Editorial Board. Journal of Symbolic Computation, 97(??):ii, ???? 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301002.

Anonymous:2020:EBa

[Ano20a] Anonymous. Editorial Board. Journal of Symbolic Computation, 96(??):ii, January/ February 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300501.

Anonymous:2020:EBc

[Ano20b] Anonymous. Editorial Board. Journal of Symbolic Computation, 99(??):ii, July/August 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300043.

Anonymous:2020:EBb

[Ano20c] Anonymous. Editorial Board.

Journal of Symbolic Computation, 101(??):ii, November/December 2020. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300377.

Anonymous:2021:EBh

[Ano21a] Anonymous. Editorial Board.

Journal of Symbolic Computation, 98(??):ii, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301427.

Anonymous:2021:EBa

[Ano21b] Anonymous. Editorial Board.

Journal of Symbolic Computation, 100(??):ii, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300134.

Anonymous:2021:EBb

[Ano21c] Anonymous. Editorial Board.

Journal of Symbolic Computation, 102(??):ii, January/February 2021. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712030078X.

Anonymous:2021:EBc

[Ano21d] Anonymous. Editorial Board. Journal of Symbolic Computation, 103(??):ii, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771712030105X.

Anonymous:2021:EBd

[Ano21e] Anonymous. Editorial Board. Journal of Symbolic Computation, 104(??):ii, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120301231.

Anonymous:2021:EBe

[Ano21f] Anonymous. Editorial Board. Journal of Symbolic Computation, 105(??):ii, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000043.

Anonymous:2021:EBf

[Ano21g] Anonymous. Editorial Board. Journal of Symbolic Computation, 106(??):ii, September/October 2021. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000171.

Anonymous:2021:EBg

[Ano21h] Anonymous. Editorial Board.

Journal of Symbolic Computation, 107(??):ii, November/December 2021. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000341.

Anonymous:2022:EBf

[Ano22a] Anonymous. Editorial Board.

Journal of Symbolic Computation, 108(??):ii, January/February 2022. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000481.

Anonymous:2022:EBg

[Ano22b] Anonymous. Editorial Board.

Journal of Symbolic Computation, 109(??):ii, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712100064X.

Anonymous:2022:EBh

[Ano22c] Anonymous. Editorial Board.

Journal of Symbolic Computation, 110(??):ii, May/June 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000766.

Anonymous:2022:EBi

[Ano22d] Anonymous. Editorial Board. Journal of Symbolic Computation, 111(??):ii, July/August 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000870.

Anonymous:2022:EBj

[Ano22e] Anonymous. Editorial Board. Journal of Symbolic Computation, 112(??):ii, September/October 2022. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000141.

Anonymous:2022:EBk

[Ano22f] Anonymous. Editorial Board. Journal of Symbolic Computation, 113(??):ii, November/December 2022. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712200044X.

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial Board. Journal of Symbolic Computation, 114(??):ii, January/February 2023. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000566.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial Board.

Journal of Symbolic Computation, 115(??):ii, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000918.

Anonymous:2023:EBc

[Ano23c] Anonymous. Editorial Board.

Journal of Symbolic Computation, 116(??):ii, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122001110.

Anonymous:2023:EBd

[Ano23d] Anonymous. Editorial Board.

Journal of Symbolic Computation, 117(??):ii, July/August 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000081.

Anonymous:2023:EBe

[Ano23e] Anonymous. Editorial Board.

Journal of Symbolic Computation, 118(??):ii, September/October 2023. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000238.

Anonymous:2023:EBf

[Ano23f] Anonymous. Editorial Board. Journal of Symbolic Computation, 119(??):ii, November/December 2023. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000421.

Anonymous:2023:EBg

[Ano23g] Anonymous. Editorial Board. Journal of Symbolic Computation, 122(??):Article 102282, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000962.

Anonymous:2023:MJ

[Ano23h] Anonymous. May–June 2024. Journal of Symbolic Computation, 122(??):??, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2023:PJA

[Ano23i] Anonymous. Pages 1–118 (July–August 2023). Journal of Symbolic Computation, 117 (??):1–118, July/August 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2023:PND

[Ano23j] Anonymous. Pages 1–144 (November–December 2023).

Journal of Symbolic Computation, 119(??):1–144, November/December 2023. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2023:PMJ

[Ano23k] Anonymous. Pages 1–426 (May–June 2023). Journal of Symbolic Computation, 116 (??):1–426, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2023:PMA

[Ano23l] Anonymous. Pages 1–518 (March-April 2023). Journal of Symbolic Computation, 115(??):1–518, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2023:PSO

[Ano23m]  Anonymous.  Pages 1–94

(September–October 2023). Journal of Symbolic Computation, 118(??):1–94, September/October 2023. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2024:EBa

[Ano24a] Anonymous. Editorial Board.

Journal of Symbolic Computation, 120(??):Article 102245, January/February 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000597.

Anonymous:2024:EBb

[Ano24b] Anonymous. Editorial Board. Journal of Symbolic Computation, 121(??):Article 102267, March/April 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000810.

Anonymous:2024:EBc

[Ano24c] Anonymous. Editorial Board. Journal of Symbolic Computation, 123(??):??, July/August 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717124000075.

Anonymous:2024:EBd

[Ano24d] Anonymous. Editorial Board. Journal of Symbolic Computation, 124(??):??, September/October 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717124000233.

Anonymous:2024:JF

[Ano24e] Anonymous. January–February 2024. Journal of Symbolic Computation, 120(??):??, January/February 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2024:JA

[Ano24f]  Anonymous.  July–August

2024. Journal of Symbolic Computation, 123(??): ??, July/August 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2024:MA

[Ano24g]  Anonymous.  March–April

2024. Journal of Symbolic Computation, 121(??): ??, March/April 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anonymous:2024:SO

[Ano24h]  Anonymous.  September–

October 2024. Journal of Symbolic Computation, 124 (??):??, September/October 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Antoy:2005:ESF

[Ant05] Sergio Antoy. Evaluation strategies for functional logic programming. Journal of Symbolic Computation, 40(1): 875–903, July 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Antoy:2010:PNT

[Ant10] Sergio Antoy. Programming with narrowing: a tutorial.

Journal of Symbolic Computation, 45(5):501–522, May 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arvasi:2023:CDG

[AOW23] Z. Arvasi, A. Odabas, and C. D. Wensley. Computing 3-dimensional groups: Crossed squares and cat ^{2}- groups. Journal of Symbolic Computation, 114(??): 267–281, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000402.

Abdulrab:1989:SWE

[AP89] Habib Abdulrab and Jean-Pierre Pécuchet. Solving word equations. Journal of Symbolic Computation, 8(5):499– 522 (or 499–521??), November 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Abouzahra:1990:CAA

[AP90] M. D. Abouzahra and R. Pavelle. Computer algebra applied to radiation from microstrip discontinuities. Journal of Symbolic Computation, 10(5):525– 528, November 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Andrews:1993:SQC

[AP93] G. E. Andrews and P. Paule.

Some questions concerning computer-generated proofs of a binomial double-sum identity. Journal of Symbolic Computation, 16(2):147–154 (or 147–153??), August 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Abramov:2004:ERN

[AP04] S. A. Abramov and M. Petkovšek.

Erratum to “Rational normal forms and minimal decompositions of hypergeometric terms” [J. Symbolic Comput. 33 (2002) 521–543]. Journal of Symbolic Computation, 38(3):1165, September 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [Ae02].

Abramov:2008:DSS

[AP08] S. A. Abramov and M. Petkovšek.

Dimensions of solution spaces of H-systems. Journal of Symbolic Computation, 43(5): 377–394, May 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Abramov:2010:PRA

[AP10] S. A. Abramov and M. Petkovšek.

Polynomial ring automorphisms, rational (w, σ)-canonical forms, and the assignment problem. Journal of Symbolic Computation, 45(6):684–708, June 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alvanos:2011:SGZ

[AP11a] Paraskevas Alvanos and Dimitrios Poulakis. Solving genus zero Diophantine equations over number fields. Journal of Symbolic Computation, 46(1):54–69, January 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001513.

Arri:2011:FCR

[AP11b] Alberto Arri and John Perry. The F5 criterion revised. Journal of Symbolic Computation, 46(9):1017–1029, September 2011. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000642.

See corrigendum [AP17].

Arri:2017:CFC

[AP17] Alberto Arri and John Perry. Corrigendum to “The F5 criterion revised” [j. symb. comput. 46 (2) (2011) 1017– 1029]. Journal of Symbolic Computation, 82(??): 164–165, September/October 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/

science/article/pii/S074771711630164X.

See [AP11b].

Amendola:2022:AVM

[AP22] Carlos Améndola and Viet Son Pham. Autocovariance varieties of moving average random fields. Journal of Symbolic Computation, 109(??): 202–219, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300572.

Apel:1995:GAI

[Ape95] Joachim Apel. A Gröbner approach to involutive bases. Journal of Symbolic Computation, 19(5):441–458 (or 441– 457??), April 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Apel:1998:TID

[Ape98] Joachim Apel. The theory of involutive divisions and an application to Hilbert function computations. Journal of Symbolic Computation, 25 (6):683–704, June 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Apery:2010:STH

[Apé10] François Apéry. Subresultants of two Hermite–Laurent series. Journal of Symbolic Computation, 45(4):

443–461, April 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Amato:2012:DIS

[APS12] Gianluca Amato, Maurizio Parton, and Francesca Scozzari. Discovering invariants via simple component analysis. Journal of Symbolic Computation, 47(12): 1533–1560, December 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002550.

Abramov:2020:CLS

[APZ20] Sergei A. Abramov, Marko Petkovsek, and Helena Zakrajsek. Convolutions of Liouvillian sequences. Journal of Symbolic Computation, 101(??):73–89, November/December 2020. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300616.

Armando:2003:CCR

[AR03] Alessandro Armando and Silvio Ranise. Constraint contextual rewriting. Journal of Symbolic Computation, 36 (1–2):193–216, July/August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Allman:2006:PIS

[AR06] Elizabeth S. Allman and John A. Rhodes. Phylogenetic invariants for stationary base composition. Journal of Symbolic Computation, 41(2): 138–150, February 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Abramov:2013:LID

[AR13] S. A. Abramov and A. A.

Ryabenko. Linear q-difference equations depending on a parameter. Journal of Symbolic Computation, 49 (??):65–77, February 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002100.

Aubry:2002:RSP

[ARE02] Philippe Aubry, Fabrice Rouillier, and Mohab Safey El Din. Real solving for positive dimensional systems. Journal of Symbolic Computation, 34(6):543–560, December 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arnon:1988:BQE

[Arn88a] Dennis S. Arnon. A bibliography of quantifier elimination for real closed fields. Journal of Symbolic Computation, 5(1–2):267–274, February/April 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arnon:1988:CBC

[Arn88b] Dennis S. Arnon. A clusterbased cylindrical algebraic decomposition algorithm. Journal of Symbolic Computation, 5(1–2):189–212, February/April 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arnault:1995:CCN

[Arn95]  François Arnault.  Con-

structing Carmichael numbers which are strong pseudoprimes to several bases. Journal of Symbolic Computation, 20(2):151–162 (or 151– 161??), August 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arnold:2003:MAC

[Arn03] Elizabeth A. Arnold. Modular algorithms for computing Gröbner bases. Journal of Symbolic Computation, 35 (4):403–419, April 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arreche:2016:CPD

[Arr16] Carlos E. Arreche. On the computation of the parameterized differential Galois group for a second-order linear differential equation with differential parameters. Journal of Symbolic Computation, 75 (??):25–55, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001054.

Armando:2002:IDP

[ARS02] Alessandro Armando, Michaël Rusinowitch, and Sorin Stratulat. Incorporating decision procedures in implicit induction. Journal of Symbolic Computation, 34(4):241– 258, October 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Abadi:2010:DFM

[ARS10] Aharon Abadi, Alexander Rabinovich, and Mooly Sagiv. Decidable fragments of manysorted logic. Journal of Symbolic Computation, 45(2):153– 172, February 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Andradas:2009:SCP

[ARST09] Carlos Andradas, Tomas Recio, J. Rafael Sendra, and Luis Felipe Tabera. On the simplification of the coefficients of a parametrization. Journal of Symbolic Computation, 44(2):192–210, February 2009. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Aries:2001:IAR

[AS01] Franck Aries and Rachid Senoussi. An implicitization algorithm for rational surfaces with no base points. Journal of Symbolic Computation, 31(4):357–365, April 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0436;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0436/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0436/ref.

Alcazar:2005:CTR

[AS05] Juan Gerardo Alcázar and J. Rafael Sendra. Computation of the topology of real algebraic space curves. Journal of Symbolic Computation, 39 (6):719–744, June 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Alcazar:2007:LSO

[AS07] Juan Gerardo Alcazar and Juan Rafael Sendra. Local shape of offsets to algebraic curves. Journal of Symbolic Computation, 42(3): 338–351, March 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Akoglu:2023:CHM

[AS23] Tulay Ayyildiz Akoglu and Agnes Szanto. Certified Hermite matrices from approximate roots. Journal of Symbolic Computation, 117(??): 101–118, July/August 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122001171.

Abdallah:2024:FRL

[AS24] Nancy Abdallah and Hal Schenck. Free resolutions and Lefschetz properties of some Artin Gorenstein rings of codimension four. Journal of Symbolic Computation, 121(??): Article 102257, March/April 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000718.

Abad:1997:PCB

[ASJ97] Alberto Abad and Felix San-Juan. PSPCLink: a cooperation between general symbolic and Poisson series processors. Journal of Symbolic Computation, 24(1):113–122, July 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Assi:1994:FGP

[Ass94] Abdallah Assi. On flatness of generic projections. Journal of Symbolic Computation, 18(5):

447–462, November 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arrondo:1997:PGO

[ASS97] Enrique Arrondo, Juana Sendra, and J. Rafael Sendra. Parametric generalized offsets to hypersurfaces. Journal of Symbolic Computation, 23 (2–3):267–286 (or 267–285??), February/March 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Alcazar:2007:DBM

[ASS07] Juan Gerardo Alcazar, Josef Schicho, and Juan Rafael Sendra. A delineability-based method for computing critical sets of algebraic surfaces. Journal of Symbolic Computation, 42(6):678–691, June 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arslan:2013:MCF

[ASS13] Feza Arslan, Neslihan Sipahi, and Nil Sahin. Monomial curve families supporting Rossi’s conjecture. Journal of Symbolic Computation, 55(??):10–18, August 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000461.

Arvanitoyeorgos:2020:HEM

[ASS20]  Andreas  Arvanitoyeorgos,

Yusuke Sakane, and Marina Statha. Homogeneous Einstein metrics on Stiefel manifolds associated to flag manifolds with two isotropy summands. Journal of Symbolic Computation, 101(??):189– 201, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300926.

Aschenwald:1996:MMP

[AST96] D. Aschenwald, T. Siegl, and R. F. Tichy. MAPinsure

— a MAPLE package for life insurance. Journal of Symbolic Computation, 22(2): 227–234, August 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Attardi:1996:SAP

[AT96] Giuseppe Attardi and Carlo Traverso. Strategy-accurate parallel Buchberger algorithms. Journal of Symbolic Computation, 21(4/5/6):411– 426 (or 411–425??), April, May & June 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Aoki:2008:LGI

[AT08] Satoshi Aoki and Akimichi Takemura. The largest group of invariance for Markov bases and toric ideals. Journal of Symbolic Computation, 43(5): 342–358, May 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Aoki:2008:IMT

[ATY08] Satoshi Aoki, Akimichi Takemura, and Ruriko Yoshida. Indispensable monomials of toric ideals and Markov bases. Journal of Symbolic Computation, 43(6–7):490–507, June/ July 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ablinger:2021:QMP

[AU21] Jakob Ablinger and Ali Kemal Uncu. qFunctions — a Mathematica package for q-series and partition theory applications. Journal of Symbolic Computation, 107(??):145– 166, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000134.

Auerbach:2005:GFG

[Aue05]  Ruth L. Auerbach.  The

Gröbner fan and Gröbner walk for modules. Journal of Symbolic Computation, 39(2): 127–153, February 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Aurenhammer:1987:RPC

[Aur87] F. Aurenhammer. Recognising polytopical cell complexes and constructing projection polyhedra. Journal of Symbolic Computation, 3 (3):249–255, June 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ahmed:1996:DTL

[AV96] M. Ahmed and G. Venkatesh.

Dense time logic programming. Journal of Symbolic Computation, 22(5–6):585– 614 (or 585–613??), November/December 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Aubry:2000:UGI

[AV00] Philippe Aubry and Annick Valibouze. Using Galois ideals for computing relative resolvents. Journal of Symbolic Computation, 30 (6):635–651, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0376;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0376/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0376/ref.

Adrovic:2011:TAG

[AV11] Danko Adrovic and Jan Verschelde. Tropical algebraic geometry in Maple: a preprocessing algorithm for finding common factors for multivariate polynomials with approximate coefficients. Journal of Symbolic Computation, 46(7):755–772, July 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001410.

Avenhaus:1986:DPT

[Ave86] Jürgen Avenhaus. On the descriptive power of term rewriting systems. Journal of Symbolic Computation, 2(2): 109–122, June 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Avendano:2009:NRL

[Ave09]  Martín Avendaño.  The

number of roots of a lacunary bivariate polynomial on a line. Journal of Symbolic Computation, 44(9): 1280–1284, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Apt:1994:SPN

[AvW94] Krzysztof R. Apt, Peter van Emde Boas, and Angelo Welling. The STO-problem is NP-hard. Journal of Symbolic Computation, 18(5):489– 495, November 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Almkvist:1990:MDU

[AZ90] Gert Almkvist and Doron Zeilberger. The method of differentiating under the integral sign. Journal of Symbolic Computation, 10(6):571– 592 (or 571–591??), December 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhao:2005:HCM

[aZGS05] Xu an Zhao, Hongzhu Gao, and Xiaole Su. Homotopy classification of maps between simply connected four manifolds. Journal of Symbolic Computation, 39(6):631–642, June 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bibel:1985:BPI

[BA85] W. Bibel and K. Aspetsberger. A bibliography on parallel inference machines. Journal of Symbolic Computation, 1(1):115–118, March 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Baader:1989:UCT

[Baa89] Franz Baader. Unification in commutative theories. Journal of Symbolic Computation, 8(5):479–498 (or 479– 497??), November 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bach:1994:TPC

[Bac94a] Eric Bach. Tensor products and computability. Journal of Symbolic Computation, 18(6): 585–593, December 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Backofen:1994:RPE

[Bac94b] Rolf Backofen. Regular path expressions in feature logic. Journal of Symbolic Computation, 17(5):421–455, May 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bach:1999:SCP

[Bac99] E. Bach. Sheaf cohomology is #P-hard. Journal of Symbolic Computation, 27(4): 429–433, April 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Baddoura:2006:IFT

[Bad06] Jamil Baddoura. Integration in finite terms with elementary functions and dilogarithms. Journal of Symbolic Computation, 41(8):909– 942, August 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bahloul:2001:ACB

[Bah01] Rouchdi Bahloul. Algorithm for computing Bernstein–Sato ideals associated with a polynomial mapping. Journal of Symbolic Computation, 32 (6):643–662, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0487;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0487/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0487/ref.

Bajaj:1986:PGA

[Baj86]  Chanderjit Bajaj.  Prov-

ing geometric algorithm nonsolvability: An application of factoring polynomials. Journal of Symbolic Computation, 2(1):99–102, March 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Barkatou:1999:RSS

[Bar99] Moulay A. Barkatou. On rational solutions of systems of linear differential equations. Journal of Symbolic Computation, 28(4–5): 547–567, October/November 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0314/production;

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0314/production/pdf;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0314/production/ref.

Barnett:2007:MMI

[Bar07] Michael P. Barnett. Mathscape and molecular integrals. Journal of Symbolic Computation, 42(3):265–289, March 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Barbosa:2013:SCA

[Bar13] Joaquim Infante Barbosa. Symbolic computation in applied computational mechanics. Journal of Symbolic Computation, 61–62(??):1–2, ???? 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001235.

Basu:2006:CFF

[Bas06] Saugata Basu. Computing the first few Betti numbers of semi-algebraic sets in single exponential time. Journal of Symbolic Computation, 41(10):1125–1154, October 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bauch:2015:GCG

[Bau15] Jens-Dietrich Bauch. Genus computation of global function fields. Journal of Symbolic Computation, 66(??):8– 20, January/February 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000327.

Bayer:2003:ACI

[Bay03] Thomas Bayer. An algorithm for computing invariants of linear actions of algebraic groups up to a given degree. Journal of Symbolic Computation, 35(4):441–449, April 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bergeron:1992:SMS

[BB92] François Bergeron and Nantel Bergeron. Symbolic manipulation for the study of the descent algebra of finite Coxeter groups. Journal of Symbolic Computation, 14(2–3): 127–140 (or 127–139??), August/September 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bibel:1993:SIA

[BB93a] W. Bibel and A. W. Biermann. Special issue: Automatic programming — foreword of the Guest Editors. Journal of Symbolic Computation, 15(5–6):463–466, May/ June 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boulanger:1993:DFU

[BB93b] Dmitri Boulanger and Maurice Bruynooghe. Deriving fold/unfold transformations of logic programs using extended OLDT-based abstract interpretation. Journal of Symbolic Computation, 15(5–6): 495–522 (or 495–521??), May/ June 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Berrizbeitia:2000:GSP

[BB00] Pedro Berrizbeitia and T. G. Berry. Generalized strong pseudoprime tests and applications. Journal of Symbolic Computation, 30(2):151– 160, August 2000. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0343;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0343/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0343/ref.

Beelen:2010:KEL

[BB10] Peter Beelen and Kristian Brander. Key equations for list decoding of Reed– Solomon codes and how to solve them. Journal of Symbolic Computation, 45(7): 773–786, July 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bailey:2011:HPN

[BB11] D. H. Bailey and J. M. Borwein. High-precision numerical integration: Progress and challenges. Journal of Symbolic Computation, 46(7):741–754, July 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001409.

Barbier:1992:AGM

[BBB92] Christine Barbier, Peter Bettess, and Jacqueline A. Bettess. Automatic generation of mapping functions for infinite elements using REDUCE. Journal of Symbolic Computation, 14(5):523–534, November 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Baldan:2011:LTP

[BBC^{+}11] Paolo Baldan, Filippo Bonchi, Andrea Corradini, Tobias Heindel, and Barbara König. A lattice-theoretical perspective on adhesive categories. Journal of Symbolic Computation, 46(3):222–245, March 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001653.

Botbol:2017:ECB

[BBC^{+}17] Nicolás Botbol, Laurent Busé, Marc Chardin, Seyed Hamid Hassanzadeh, Aron Simis, and Quang Hoa Tran. Effective criteria for bigraded birational maps. Journal of Symbolic Computation, 81 (??):69–87, July/August 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301626.

Bean:2022:GFP

[BBCF22] Christian Bean, Antonio Bernini, Matteo Cervetti, and Luca Ferrari. On the generating functions of patternavoiding Motzkin paths. Journal of Symbolic Computation, 113(??):126–138, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000189.

Bernardi:2013:GTD

[BBCM13] A. Bernardi, J. Brachat, P. Comon, and B. Mourrain. General tensor decomposition, moment matrices and applications. Journal of Symbolic Computation, 52(??):51–71, May 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001290.

Berthomieu:2017:LAC

[BBF17] Jérémy Berthomieu, Brice Boyer, and Jean-Charles Faugère. Linear algebra for computing Gröbner bases of linear recursive multidimensional sequences. Journal of Symbolic Computation, 83(??):36–67, November/December 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301249.

Bailey:2014:ASL

[BBK14] David H. Bailey, Jonathan M.

Borwein, and Alexander D. Kaiser. Automated simplification of large symbolic expressions. Journal of Symbolic Computation, 60 (??):120–136, January 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711300117X.

Bouhoula:2015:SIS

[BBKK15] Adel Bouhoula, Bruno Buchberger, Laura Kovács, and Temur Kutsia. Special issue on symbolic computation in software science. Journal of Symbolic Computation, 69(??):1–2, July/August 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000959.

Boutry:2018:FAE

[BBN18] Pierre Boutry, Gabriel Braun, and Julien Narboux. Formalization of the arithmetization of Euclidean plane geometry and applications. Journal of Symbolic Computation, 90(??):149–168, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300361.

Bendle:2024:MPC

[BBRS24] Dominik Bendle, Janko Böhm, Yue Ren, and Benjamin Schröter. Massively parallel computation of tropical varieties, their positive part, and tropical Grassmannians. Journal of Symbolic Computation, 120(??):Article 102224, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000317.

Baldoni:2015:MBS

[BBV15] Velleda Baldoni, Arzu Boysal, and Michèle Vergne. Multiple Bernoulli series and volumes of moduli spaces of flat bundles over surfaces. Journal of Symbolic Computation, 68 (part 2)(??):27–60, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000704.

Ballico:2022:SIH

[BBV22] Edoardo Ballico, Alessandra Bernardi, and Emanuele Ventura. Strict inclusions of high rank loci. Journal of Symbolic Computation, 109(??): 238–249, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300596.

Butler:1989:CPM

[BC89] Gregory Butler and John Cannon. Computing in permutation and matrix groups. III: Sylow subgroups. Journal of Symbolic Computation, 8 (3):241–252, September 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Butler:1991:CSS

[BC91] Gregory Butler and John Cannon. Computing Sylow subgroups of permutation groups using homomorphic images of centralizers. Journal of Symbolic Computation, 12(4–5): 443–458 (or 443–457??), October/November 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Butler:1993:HA

[BC93] Greg Butler and John J. Cannon. On Holt’s algorithm. Journal of Symbolic Computation, 15(2):229–234 (or 229– 233??), February 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Barendregt:2001:ECM

[BC01] Henk Barendregt and Arjeh M. Cohen. Electronic communication of mathematics and the interaction of computer algebra systems and proof assistants. Journal of Symbolic Computation, 32(1–2):3–22, July/August 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0455;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0455/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0455/ref.

Buse:2005:IRH

[BC05] Laurent Busé and Marc Chardin. Implicitizing rational hypersurfaces using approximation complexes. Journal of Symbolic Computation, 40(4–5):1150–1168, October/November 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bendersky:2006:NFC

[BC06] Martin Bendersky and Richard C. Churchill. Normal forms in a cyclically graded Lie algebra. Journal of Symbolic Computation, 41(6):633–662, June 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boreale:2022:LAM

[BC22] Michele Boreale and Luisa Collodi. A linear-algebraic method to compute polynomial PDE conservation laws. Journal of Symbolic Computation, 108(??):55–72, January/February 2022. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000390.

Barkatou:2024:CRS

[BC24a] Moulay Barkatou and Thomas Cluzeau. On the computation of rational solutions of linear integro-differential equations with polynomial coefficients. Journal of Symbolic Computation, 121(??): Article 102252, March/April 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000664.

Bruns:2024:SCM

[BC24b] Winfried Bruns and Aldo Conca. Sagbi combinatorics of maximal minors and a Sagbi algorithm. Journal of Symbolic Computation, 120(??):Article 102237, January/February 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000512.

Boura:2020:BST

[BCCK20] Christina Boura, Eirini Chavli, Maria Chlouveraki, and Konstantinos Karvounis. The BMM symmetrising trace conjecture for groups G_{4}, G_{5}, G_{6}, G_{7}, G_{8}. Journal of Symbolic Computation, 96(??):62– 84, January/February 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300239.

Barkee:1994:WYC

[BCE^{+}94] Boo Barkee, Deh Cac Can, Julia Ecks, Theo Moriarty, and R. F. Ree. Why you cannot even hope to use Gröbner bases in public key cryptography: An open letter to a scientist who failed and a challenge to those who have not yet failed. Journal of Symbolic Computation, 18(6):497– 502 (or 497–501??), December 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Buhler:2001:IPC

[BCE^{+}01] Joe Buhler, Richard Crandall, Reijo Ernvall, Tauno Metsänkylä, and M. Amin Shokrollahi. Irregular primes and cyclotomic invariants to 12 million. Journal of Symbolic Computation, 31(1– 2):89–96, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1011;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.1011/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1011/ref.

Barkatou:2011:SFH

[BCE11] Moulay A. Barkatou, Thomas Cluzeau, and Carole El Bacha. Simple forms of higher-order linear differential systems and their applications in computing regular solutions. Journal of Symbolic Computation, 46(6):633–658, June 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000137.

Bertone:2019:FLP

[BCF19] Cristina Bertone, Francesca Cioffi, and Davide Franco. Functors of liftings of projective schemes. Journal of Symbolic Computation, 94(??): 105–125, September/October 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300907.

Bertone:2010:MVA

[BCG10] Cristina Bertone, Guillaume Chèze, and André Galligo. Modular Las Vegas algorithms for polynomial absolute factorization. Journal of Symbolic Computation, 45(12): 1280–1295, December 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Berque:1992:SPS

[BCGR92] David Berque, Ronald Cecchini, Mark Goldberg, and Reid Rivenburgh. The set player system for symbolic computation on power sets. Journal of Symbolic Computation, 14(6):645–662, December 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Burr:2012:CSA

[BCGY12] Michael Burr, Sung Woo Choi, Ben Galehouse, and Chee K. Yap. Complete subdivision algorithms, II: Isotopic meshing of singular algebraic curves. Journal of Symbolic Computation, 47(2):131–152, February 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001337.

Bartolini:2013:AGC

[BCI13] Gabriel Bartolini, Antonio F. Costa, and Milagros Izquierdo. On automorphisms groups of cyclic pgonal Riemann surfaces. Journal of Symbolic Computation, 57(??):61–69, October 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000680.

Beckermann:2006:FFR

[BCL06] Bernhard Beckermann, Howard Cheng, and George Labahn. Fraction-free row reduction of matrices of Ore polynomials. Journal of Symbolic Computation, 41(5):513–543, May 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bertone:2013:UME

[BCLR13] Cristina Bertone, Francesca Cioffi, Paolo Lella, and Margherita Roggero. Upgraded methods for the effective computation of marked schemes on a strongly stable ideal. Journal of Symbolic Computation, 50(??): 263–290, March 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001241.

Bosma:1997:MAS

[BCP97] Wieb Bosma, John J. Cannon, and Catherine Playoust. The Magma algebra system. I. The user language. Journal of Symbolic Computation, 24(3–4): 235–265, September/October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Bigatti:2011:CIG

[BCR11] A. M. Bigatti, M. Caboara, and L. Robbiano. Computing inhomogeneous Gröbner bases. Journal of Symbolic Computation, 46(5):498–510, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001719.

Belkhir:2015:PAS

[BCR15] Walid Belkhir, Yannick Chevalier, and Michael Rusinowitch. Parametrized automata simulation and application to service composition. Journal of Symbolic Computation, 69 (??):40–60, July/August 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000972.

Brackx:1989:HMD

[BCRS89] F. Brackx, D. Constales, A. Ronveaux, and H. Serras. On the harmonic and monogenic decomposition of polynomials. Journal of Symbolic Computation, 8(3):297– 304, September 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bosma:1997:LCE

[BCS97] Wieb Bosma, John J. Cannon, and Allan Steel. Lattices of compatibly embedded finite fields. Journal of Symbolic Computation, 24(3–4):351– 370 (or 351–369??), September/October 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Bostan:2011:HTM

[BCvdHS11] Alin Bostan, Muhammad F. I.

Chowdhury, Joris van der Hoeven, and Éric Schost. Homotopy techniques for multiplication modulo triangular sets. Journal of Symbolic Computation, 46(12): 1378–1402, December 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001271.

Bandyopadhyay:1987:SSI

[BD87] S. Bandyopadhyay and J. S.

Devitt. SIS: a symbolic information management system. Journal of Symbolic Computation, 4(3):397–408, December 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bachmair:1988:CPC

[BD88] Leo Bachmair and Nachum Dershowitz. Critical pair criteria for completion. Journal of Symbolic Computation, 6(1):1–18, August 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bozzano:2004:AVS

[BD04] Marco Bozzano and Giorgio Delzanno. Automatic verification of secrecy properties for linear logic specifications of cryptographic protocols. Journal of Symbolic Computation, 38(5): 1375–1415, November 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Brickenstein:2009:PFG

[BD09] Michael Brickenstein and Alexander Dreyer. PolyBoRi: framework for Gröbner-basis computations with Boolean polynomials. Journal of Symbolic Computation, 44(9): 1326–1345, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bluman:2012:NSO

[BD12] George Bluman and Raouf Dridi. New solutions for ordinary differential equations. Journal of Symbolic Computation, 47(1):76–88, January 2012. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001301.

Brickenstein:2013:GFN

[BD13] Michael Brickenstein and Alexander Dreyer. Gröbnerfree normal forms for Boolean polynomials. Journal of Symbolic Computation, 48 (??):37–53, January 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000624.

Benitez:2015:RAM

[BD15] Teresa Cortadellas Benítez and Carlos D’Andrea. The Rees Algebra of a monomial plane parametrization. Journal of Symbolic Computation, 70(??):71–105, September/October 2015. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000947.

Botbol:2016:IRH

[BD16] Nicolás Botbol and Alicia Dickenstein. Implicitization of rational hypersurfaces via linear syzygies: a practical overview. Journal of Symbolic Computation, 74 (??):493–512, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000917.

Bubenik:2017:PLT

[BD17] Peter Bubenik and Pawel Dlotko. A persistence landscapes toolbox for topological statistics. Journal of Symbolic Computation, 78(??):91– 114, January/February 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300104.

Bradford:2016:TTI

[BDE^{+}16] Russell Bradford, James H. Davenport, Matthew England, Scott McCallum, and David Wilson. Truth table invariant cylindrical algebraic decomposition. Journal of Symbolic Computation, 76(??):1–35, September/October 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115001005.

Bradford:2021:IPO

[BDE^{+}21] Russell Bradford, James H. Davenport, Matthew England, Hassan Errami, Vladimir Gerdt, Dima Grigoriev, Charles Hoyt, Marek Košta, Ovidiu Radulescu, Thomas Sturm, and Andreas Weber. Identifying the parametric occurrence of multiple steady states for some biological networks. Journal of Symbolic Computation, 98(??):84–119, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300732.

Bohm:2013:PAN

[BDL^{+}13] Janko Böhm, Wolfram Decker, Santiago Laplagne, Gerhard Pfister, Andreas Steenpaß, and Stefan Steidel. Parallel algorithms for normalization. Journal of Symbolic Computation, 51(??):99–114, April 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001186.

Benito:2019:QSL

[BdlCRLS19] Pilar Benito, Daniel de-la Concepción, Jorge Roldán-López, and Iciar Sesma. Quadratic 2-step Lie algebras: Computational algorithms and classification. Journal of Symbolic Computation, 94(??):70–89, September/October 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300889.

Bohm:2022:CIB

[BDLP22] Janko Böhm, Wolfram Decker, Santiago Laplagne, and Gerhard Pfister. Computing integral bases via localization and Hensel lifting. Journal of Symbolic Computation, 109 (??):283–324, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300626.

Bui:2016:DBN

[BDM^{+}16] V. C. Bui, G. H. E. Duchamp, V. Hoang Ngoc Minh, L. Kane, and C. Tollu. Dual bases for noncommutative symmetric and quasi-symmetric functions via monoidal factorization. Journal of Symbolic Computation, 75(??):56–73, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115001066.

Bui:2017:SPE

[BDM17] V. C. Bui, G. H. E. Duchamp, and V. Hoang Ngoc Minh. Structure of polyzetas and explicit representation on transcendence bases of shuffle and stuffle algebras. Journal of Symbolic Computation, 83(??):93–111, November/December 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301262.

Benitez:2023:BDS

[BDM23] Teresa Cortadellas Benítez, Carlos D’Andrea, and M. Eulàlia Montoro. Bounds for degrees of syzygies of polynomials defining a grade two ideal. Journal of Symbolic Computation, 115(??): 124–141, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000700.

Bishop:2013:CCA

[BDPR13] Marcus Bishop, J. Matthew Douglass, Götz Pfeiffer, and Gerhard Röhrle. Computations for Coxeter arrangements and Solomon’s descent algebra: Groups of rank three and four. Journal of Symbolic Computation, 50(??):139–158, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001095.

Bosma:2001:CNR

[BdS01] Wieb Bosma and Bart de Smit.

Class number relations from a computational point of view. Journal of Symbolic Computation, 31(1–2): 97–112, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1016;   http://www.
idealibrary.com/links/doi/

                                        10.1006/jsco.1999.1016/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.1016/ref.

Bostan:2017:ADW

[BDS17] Alin Bostan, Louis Dumont, and Bruno Salvy. Algebraic diagonals and walks: Algorithms, bounds, complexity. Journal of Symbolic Computation, 83(??): 68–92, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301250.

Bauer:2018:PLF

[BDS^{+}18] Thomas Bauer, Sandra Di Rocco, David Schmitz, Tomasz Szemberg, and Justyna Szpond. On the postulation of lines and a fat line. Journal of Symbolic Computation, 91(??):3–16, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300750.

Bernardi:2022:FSI

[BDT22] Alessandra Bernardi, Carlos D’Andrea, and Thorsten Theobald. Foreword: Special issue of JSC on the occasion of MEGA 2019. Journal of Symbolic Computation, 109(??): 199–201, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300560.

Besche:1999:CFG

[BE99a] H. U. Besche and B. Eick.

Construction of finite groups. Journal of Symbolic Computation, 27(4):387–404, April 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Besche:1999:GOM

[BE99b] H. U. Besche and B. Eick. The groups of order at most 1000 except 512 and 768. Journal of Symbolic Computation, 27 (4):405–414, April 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bouziane:2002:CDP

[BE02] Driss Bouziane and Mhammed El Kahoui. Computation of the dual of a plane projective curve. Journal of Symbolic Computation, 34(2): 105–117, August 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bonacina:2010:TDD

[BE10] Maria Paola Bonacina and Mnacho Echenim. Theory decision by decomposition. Journal of Symbolic Computation, 45(2):229–260, February 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Blanco:2011:EDT

[BE11] Rocío Blanco and Santiago Encinas. Embedded desingularization of toric varieties. Journal of Symbolic Computation, 46(11): 1229–1241, November 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001179.

Barkatou:2013:SFF

[BE13] Moulay A. Barkatou and Carole El Bacha. On ksimple forms of first-order linear differential systems and their computation. Journal of Symbolic Computation, 54(??):36–58, July 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000059.

Bihan:2017:SBN

[BE17] Frédéric Bihan and Boulos El Hilany. A sharp bound on the number of real intersection points of a sparse plane curve with a line. Journal of Symbolic Computation, 81 (??):88–96, July/August 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301651.

Berthomieu:2022:GGB

[BE22] Jérémy Berthomieu and Mohab Safey El Din. Guessing Gröbner bases of structured ideals of relations of sequences. Journal of Symbolic Computation, 111(??): 1–26, July/August 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000717.

Beauzamy:1992:PPP

[Bea92] Bernard Beauzamy. Products of polynomials and a priori estimates for coefficients in polynomial decompositions: a sharp result. Journal of Symbolic Computation, 13(5): 463–472, May 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Becker:1990:SBS

[Bec90] Thomas Becker. Standard bases and some computations in rings of power series. Journal of Symbolic Computation, 10(2):165–178, August 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Becker:1993:SBP

[Bec93] Thomas Becker. Standard bases in power series rings: Uniqueness and superfluous critical pairs. Journal of Symbolic Computation, 15(3):251– 266 (or 251–265??), March 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Beckert:2003:DFP

[Bec03] Bernhard Beckert. Depthfirst proof search without backtracking for free-variable clausal tableaux. Journal of Symbolic Computation, 36 (1–2):117–138, July/August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Beck:2009:FDS

[Bec09] Tobias Beck. Formal desingularization of surfaces: The Jung method revisited. Journal of Symbolic Computation, 44(2):131–160, February 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bedratyuk:2007:CSI

[Bed07] Leonid Bedratyuk. On complete system of invariants for the binary form of degree 7. Journal of Symbolic Computation, 42(10):935–947, October 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bedratyuk:2009:CMS

[Bed09] Leonid Bedratyuk. A complete minimal system of covariants for the binary form of degree 7. Journal of Symbolic Computation, 44(2):211– 220, February 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Beeson:2001:ADI

[Bee01]  Michael Beeson.  Auto-

matic derivation of the irrationality of e. Journal of Symbolic Computation, 32 (4):333–349, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0465;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0465/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0465/ref.

Buse:2009:CSR

[BEG09] Laurent Busé, Mohamed Elkadi, and André Galligo. A computational study of ruled surfaces. Journal of Symbolic Computation, 44(3): 232–241, March 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Belinfante:2003:CPA

[Bel03] Johan Gijsbertus Frederik Belinfante. Computer proofs about finite and regular sets: the unifying concept of subvariance. Journal of Symbolic Computation, 36(1–2): 271–285, July/August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Belabas:2004:RVH

[Bel04] Karim Belabas. A relative van Hoeij algorithm over number fields. Journal of Symbolic Computation, 37(5):641–668, May 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Barkatou:2013:SRC

[BELP13] Moulay A. Barkatou, Carole El Bacha, George Labahn, and Eckhard Pflügel. On simultaneous row and column reduction of higher-order linear differential systems. Journal of Symbolic Computation, 49(??):45–64, February 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002094.

Bartzos:2021:MNR

[BELT21] Evangelos Bartzos, Ioannis Z. Emiris, Jan Legerský, and Elias Tsigaridas. On the maximal number of real embeddings of minimally rigid graphs in R_{2}, R_{3} and S^{2}. Journal of Symbolic Computation, 102(??): 189–208, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301257.

Brown:1997:NAC

[BEM97] B. M. Brown, M. S. P. Eastham, and D. K. R. McCormack. A new algorithm for computing the asymptotic solutions of a class of linear differential systems. Journal of Symbolic Computation, 23(1): 119–131, January 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Buse:2000:GRU

[BEM00] Laurent Busé, Mohamed Elkadi, and Bernard Mourrain. Generalized resultants over unirational algebraic varieties. Journal of Symbolic Computation, 29(4–5):515–526, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0304;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0304/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0304/ref.

Brown:2006:AMI

[BENW06] Christopher W. Brown, M’hammed El Kahoui, Dominik Novotni, and Andreas Weber. Algorithmic methods for investigating equilibria in epidemic modeling. Journal of Symbolic Computation, 41(11):1157– 1173, November 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bates:2013:CIN

[BEP13] Daniel J. Bates, David Eklund, and Chris Peterson. Computing intersection numbers of Chern classes. Journal of Symbolic Computation, 50(??):493–507, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001587.

Bergeron:1993:SMU

[Ber93] François Bergeron. Surprising mathematics using a computer algebra system. Journal of Symbolic Computation, 15 (3):365–370, March 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bernstein:1998:CPS

[Ber98a] Daniel J. Bernstein. Composing power series over a finite ring in essentially linear time. Journal of Symbolic Computation, 26(3):339– 341, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Berry:1998:CLS

[Ber98b] T. G. Berry. Construction of linear systems on hyperelliptic curves. Journal of Symbolic Computation, 26(3):315– 327, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Berman:2002:CGG

[Ber02] Peter Berman. Calculating the Galois group of Y′ = AY + B, Y′ = AY completely reducible. Journal of Symbolic Computation, 33(6):887–898, June 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bernstein:2004:CCR

[Ber04] Dan Bernstein. The computational complexity of rules for the character table of S_{n}. Journal of Symbolic Computation, 37(6):727–748, June 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ballico:2013:EMP

[BES13] E. Ballico, M. Elia, and M. Sala. On the evaluation of multivariate polynomials over finite fields. Journal of Symbolic Computation, 50(??):255–262, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711200123X.

Brandenburg:2023:MVE

[BEZ23] Marie-Charlotte Brandenburg, Sophia Elia, and Leon Zhang. Multivariate volume, Ehrhart, and h^{∗}-polynomials of polytropes. Journal of Symbolic Computation, 114(??): 209–230, January/February 2023. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000311.

Bjorck:1991:FWC

[BF91] Göran Björck and Ralf Fröberg. A faster way to count the solution of inhomogeneous systems of algebraic equations, with applications to cyclic n-roots. Journal of Symbolic Computation, 12 (3):329–336, September 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Baumgartner:1993:CFC

[BF93] Peter Baumgartner and Ulrich Furbach. Consolution as a framework for comparing calculi. Journal of Symbolic Computation, 16(5):445– 477, November 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Baaz:1995:RBT

[BF95] Matthias Baaz and Christian G. Fermüller. Resolutionbased theorem proving for many-valued logics. Journal of Symbolic Computation, 19 (4):353–391, April 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Byrne:2001:GBG

[BF01] Eimear Byrne and Patrick Fitzpatrick. Gröbner bases over Galois rings with an application to decoding alternant codes. Journal of Symbolic Computation, 31 (5):565–584, May 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0442;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0442/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0442/ref.

Bretto:2011:CGG

[BF11] Alain Bretto and Alain Faisant. Cayley graphs and G-graphs: Some applications. Journal of Symbolic Computation, 46(12): 1403–1412, December 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001283.

Berthomieu:2020:DCB

[BF20] Jérémy Berthomieu and Jean-Charles Faugère. In-depth comparison of the Berlekamp– Massey-sakata and the Scalar-FGLM algorithms: the adaptive variants. Journal of Symbolic Computation, 101(??):270–303, November/ December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300963.

Besier:2021:RSR

[BF21] Marco Besier and Dino Festi. Rationalizability of square roots. Journal of Symbolic Computation, 106(??): 48–67, September/October 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120301292.

Berthomieu:2022:PDB

[BF22] Jérémy Berthomieu and Jean-Charles Faugère. Polynomialdivision-based algorithms for computing linear recurrence relations. Journal of Symbolic Computation, 109(??): 1–30, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000432.

Bretto:2007:GNR

[BFG07] Alain Bretto, Alain Faisant, and Luc Gillibert. Ggraphs: a new representation of groups. Journal of Symbolic Computation, 42(5):549–560, May 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Biasse:2017:CHM

[BFH17] Jean-François Biasse, Claus Fieker, and Tommy Hofmann. On the computation of the HNF of a module over the ring of integers of a number field. Journal of Symbolic Computation, 80 (part 3) (??):581–615, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300736.

Beyer:1992:VCT

[BFHS92] W. A. Beyer, L. R. Fawcett, L. P. Harten, and B. K. Swartz. The volume common to two congruent circular cylinders. Journal of Symbolic Computation, 13(2):221– 230, February 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Borodin:1985:DND

[BFHT85] Allan Borodin, Ronald Fagin, John E. Hopcroft, and Martin Tompa. Decreasing the nesting depth of expressions involving square roots. Journal of Symbolic Computation, 1(2):169–188, June 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bauer:2002:IGF

[BFK02] Christian Bauer, Alexander Frink, and Richard Kreckel. Introduction to the GiNaC framework for symbolic computation within the C++ programming language. Journal of Symbolic Computation, 33(1):1–12, January 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0494;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0494/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0494/ref.

Bohm:2018:STH

[BFK18] Janko Böhm and Anne Frühbis-Krüger. A smoothness test for higher codimensions. Journal of Symbolic Computation, 86(??): 153–165, May/June 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300500.

Beyer:1987:VCT

[BFMS87] W. A. Beyer, L. R. Fawcett, R. D. Mauldin, and B. K. Swartz. The volume common to two congruent circular cones whose axes intersect symmetrically. Journal of Symbolic Computation, 4(3): 381–390, December 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bender:2021:NOA

[BFPT21] Matías R. Bender, Jean-Charles Faugère, Ludovic Perret, and Elias Tsigaridas.

A nearly optimal algorithm to decompose binary forms. Journal of Symbolic Computation, 105(??):71–96, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300444.

Bardet:2015:CIG

[BFS15] Magali Bardet, Jean-Charles Faugère, and Bruno Salvy. On the complexity of the F_{5} Gröbner basis algorithm. Journal of Symbolic Computation, 70(??):49–70, September/October 2015. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000935.

Bostan:2006:FCS

[BFSS06] Alin Bostan, Philippe Flajolet, Bruno Salvy, and Éric Schost. Fast computation of special resultants. Journal of Symbolic Computation, 41 (1):1–29, January 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bebbington:2001:KMP

[BG01] David Bebbington and Manfred Göbel. KLEIN: a Mathematica package for radar polarimetry based on spinor and tensor algebra. Journal of Symbolic Computation, 31(6):745–751, June 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0460;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.0460/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0460/ref.

Buse:2005:SIR

[BG05] Laurent Busé and André Galligo. Semi-implicit representations of surfaces in P^{3}, resultants and applications. Journal of Symbolic Computation, 39(3–4):317– 329, March/April 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bosse:2013:ISE

[BGG13] Hartwig Bosse, Christine Gärtner, and Oleg Golubitsky. Ideal-specific elimination orders form a starshaped region. Journal of Symbolic Computation, 56 (??):69–79, September 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711300062X.

Bennett:1993:CAT

[BGH93] J. P. Bennett, M. Grinfeld, and J. Hubble. Computer algebra techniques in affinity binding equations: the dimer case. Journal of Symbolic Computation, 15(1):79–84 (or 79–83??), January 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Baldwin:2004:SCE

[BGH^{+}04] D. Baldwin, Ü. Göktaš, W. Hereman, L. Hong, R. S. Martino, and J. C. Miller. Symbolic computation of exact solutions expressible in hyperbolic and elliptic functions for nonlinear PDEs. Journal of Symbolic Computation, 37 (6):669–705, June 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Brown:2006:SRD

[BGHW06] Ronald Brown, Neil Ghani, Anne Heyworth, and Christopher D. Wensley. String rewriting for double coset systems. Journal of Symbolic Computation, 41(5):573–590, May 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bernardi:2011:CSR

[BGI11] Alessandra Bernardi, Alessandro Gimigliano, and Monica Idà. Computing symmetric rank for symmetric tensors. Journal of Symbolic Computation, 46(1):34–53, January 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001240.

Bernardi:2018:SPR

[BGI18] Alessandra Bernardi, Alessandro Gimigliano, and Monica Idà. Singularities of plane rational curves via projections. Journal of Symbolic Computation, 86(??):189–214, May/ June 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300524.

Boege:1986:SES

[BGK86] W. Boege, R. Gebauer, and H. Kredel. Some examples for solving systems of algebraic equations by calculating Gröbner bases. Journal of Symbolic Computation, 2 (1):83–98, March 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bundgen:1996:SCM

[BGK96] Reinhard Bündgen, Manfred Göbel, and Wolfgang Küchlin. Strategy compliant multithreaded term completion. Journal of Symbolic Computation, 21(4/5/6):475–506 (or 475–505??), April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Belkhir:2014:STL

[BGL14] Walid Belkhir, Alain Giorgetti, and Michel Lenczner. A symbolic transformation language and its application to a multiscale method. Journal of Symbolic Computation, 65(??):49–78, November 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000145.

Bermejo:2017:AIP

[BGLGM17] Isabel Bermejo, Eva García-Llorente, and Ignacio García-Marco. Algebraic invariants of projective monomial curves associated to generalized arithmetic sequences. Journal of Symbolic Computation, 81 (??):1–19, July/August 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301201.

Bachler:2012:ATD

[BGLHR12] Thomas Bächler, Vladimir Gerdt, Markus Lange-Hegermann, and Daniel Robertz. Algorithmic Thomas decomposition of algebraic and differential systems. Journal of Symbolic Computation, 47(10):1233–1266, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711100246X.

Bermejo:2006:CMR

[BGM06] Isabel Bermejo, Philippe Gimenez, and Marcel Morales. Castelnuovo–Mumford regularity of projective monomial varieties of codimension two. Journal of Symbolic Computation, 41(10):1105–1124, October 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bermejo:2015:CIS

[BGM15] Isabel Bermejo and Ignacio García-Marco. Complete intersections in simplicial toric varieties. Journal of Symbolic Computation, 68 (part 1) (??):265–286, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000881.

Bermejo:2007:ACW

[BGMSG07] Isabel Bermejo, Ignacio García-Marco, and Juan José Salazar-González. An algorithm for checking whether the toric ideal of an affine monomial curve is a complete intersection. Journal of Symbolic Computation, 42(10): 971–991, October 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Berbain:2009:QMS

[BGP09] Côme Berbain, Henri Gilbert, and Jacques Patarin. QUAD: a multivariate stream cipher with provable security. Journal of Symbolic Computation, 44(12):1703–1723, December 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Braatz:2011:HDC

[BGS11] Benjamin Braatz, Ulrike Golas, and Thomas Soboll. How to delete categorically

— two pushout complement constructions. Journal of Symbolic Computation, 46 (3):246–271, March 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001665.

Burr:2020:CSD

[BGT20] Michael Burr, Shuhong Gao, and Elias Tsigaridas. The complexity of subdivision for diameter–distance tests. Journal of Symbolic Computation, 101(??):1–27, November/December 2020. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300574.

Beyer:1987:STA

[BH87] W. A. Beyer and L. Heller.

A Steiner tree associated with tree quarks. Journal of Symbolic Computation, 3(3): 283–289, June 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bonacina:1995:DDC

[BH95] Maria Paola Bonacina and Jieh Hsiang. Distributed deduction by clause-diffusion: Distributed contraction and the Aquarius prover. Journal of Symbolic Computation, 19(1/2/3):245–268 (or 245–267??), January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Brown:2000:URS

[BH00] Ronald Brown and Anne Heyworth. Using rewriting systems to compute left Kan extensions and induced actions of categories. Journal of Symbolic Computation, 29(1):5–31, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0294;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0294/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0294/ref;   http://
www.idealibrary.com/links/
doi/10.1006/jsco.1999.0295;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.

1999.0295/pdf;   http://
www.idealibrary.com/links/
doi/10.1006/jsco.1999.0295/
ref.

Bremner:2002:IAA

[BH02] Murray Bremner and Irvin Hentzel. Identities for the associator in alternative algebras. Journal of Symbolic Computation, 33(3): 255–273, March 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0510;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0510/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0510/ref.

Bazan:2021:MIP

[BH21] Erick Rodriguez Bazan and Evelyne Hubert. Multivariate interpolation: Preserving and exploiting symmetry. Journal of Symbolic Computation, 107(??):1–22, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000109.

Bazan:2023:SMI

[BH23] Erick Rodriguez Bazan and Evelyne Hubert. Symmetry in multivariate ideal interpolation. Journal of Symbolic Computation, 115(??):

174–200, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000803.

Bao:2023:NA

[BHH23] Jiakang Bao, Yang-Hui He, and Edward Hirst. Neurons on amoebae. Journal of Symbolic Computation, 116 (??):1–38, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000840.

Baarnhielm:2015:PMC

[BHLGO15] Henrik Bäärnhielm, Derek Holt, C. R. Leedham-Green, and E. A. O’Brien. A practical model for computation with matrix groups. Journal of Symbolic Computation, 68 (part 1)(??):27–60, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711400056X.

Bernal:2023:MLR

[BHM^{+}23] Edgar A. Bernal, Jonathan D. Hauenstein, Dhagash Mehta, Margaret H. Regan, and Tingting Tang. Machine learning the real discriminant locus. Journal of Symbolic Computation, 115(??): 409–426, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000670.

Burckert:1989:ETU

[BHSS89] Hans-Jürgen Bürckert, Alexander Herold, and Manfred Schmidt-Schauss. On equational theories, unification, and (un)decidability. Journal of Symbolic Computation, 8(1–2):3–50 (or 3–49??), July/August 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bibel:1985:AI

[Bib85] W. Bibel. Automated inferencing. Journal of Symbolic Computation, 1(3):245– 260, September 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Biermann:1985:APT

[Bie85] A. W. Biermann. Automatic programming: a tutorial on formal methodologies. Journal of Symbolic Computation, 1(2):119–142, June 1985.

Bihan:2015:MPP

[Bih15] Frédéric Bihan. Maximally positive polynomial systems supported on circuits. Journal of Symbolic Computation, 68 (part 2)(??):61–74, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000716.

Bila:2011:NMF

[Bîl11] Nicoleta Bîlǎ. On a new method for finding generalized equivalence transformations for differential equations involving arbitrary functions. Journal of Symbolic Computation, 46(6):659–671, June 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111000149.

Birget:1998:ISR

[Bir98] Jean-Camille Birget. Infinite string rewrite systems and complexity. Journal of Symbolic Computation, 25(6): 759–793, June 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bruns:2016:PPD

[BIS16] Winfried Bruns, Bogdan Ichim, and Christof Söger. The power of pyramid decomposition in Normaliz. Journal of Symbolic Computation, 74(??):513–536, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000930.

Barkatou:2021:RAS

[BJ21] Moulay A. Barkatou and Maximilian Jaroschek. Removing apparent singularities of linear difference systems. Journal of Symbolic Computation, 102(??): 86–107, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711930118X.

Barkatou:2017:FSC

[BJM17] Moulay A. Barkatou, Maximilian Jaroschek, and Suzy S. Maddah. Formal solutions of completely integrable Pfaffian systems with normal crossings. Journal of Symbolic Computation, 81(??): 41–68, July/August 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301468.

Blanco:2004:CGI

[BJS04] Cristina Blanco, Gabriela Jeronimo, and Pablo Solernó. Computing generators of the ideal of a smooth affine algebraic variety. Journal of Symbolic Computation, 38(1): 843–872, July 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bogart:2007:CTV

[BJS^{+}07] T. Bogart, A. N. Jensen, D. Speyer, B. Sturmfels, and R. R. Thomas. Computing tropical varieties. Journal of Symbolic Computation, 42(1– 2):54–73, January/February 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boudet:1989:UBR

[BJSS89] Alexandre Boudet, Jean-Pierre P. Jouannaud, and Manfred Schmidt-Schauss. Unification in Boolean rings and Abelian groups. Journal of Symbolic Computation, 8(5):449–478 (or 449– 477??), November 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bochenski:2022:HSR

[BJT22] Maciej Bocheński, Piotr Jastrzebski, and Aleksy Tralle. Homogeneous spaces of real simple Lie groups with proper actions of non virtually abelian discrete subgroups: a computational approach. Journal of Symbolic Computation, 113(??):171–180, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000244.

Buchberger:1990:SIC

[BK90] B. Buchberger and E. Kaltofen. Special issue: Computational algebraic complexity— editorial. Journal of Symbolic Computation, 9(3):225– 228, March 1990.

Boing:1999:AHS

[BK99a] Harald Böing and Wolfram Koepf. Algorithms for qhypergeometric summation in computer algebra. Journal of Symbolic Computation, 28(6):777–799, December 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0339/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0339/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0339/production/ref.

Broadhurst:1999:RAH

[BK99b] D. J. Broadhurst and D. Kreimer.

Renormalization automated by Hopf algebra. Journal of Symbolic Computation, 27(6):581–600, June 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0283/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0283/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0283/production/ref.

Ballis:2011:F

[BK11] Demis Ballis and Temur Kutsia. Foreword. Journal of Symbolic Computation, 46 (2):93–94, February 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001331.

Bjorner:2012:F

[BK12a] Nikolaj Bjørner and Laura Kovács. Foreword. Journal of Symbolic Computation, 47(12):1413–1415, December 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002501.

Burr:2012:SAO

[BK12b] Michael A. Burr and Felix Krahmer. SqFreeEVAL: an (almost) optimal real-root isolation algorithm. Journal of Symbolic Computation, 47(2):153–166, February 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001349.

Brown:2013:SPT

[BK13] Gavin Brown and Alexander M. Kasprzyk. Small polygons and toric codes. Journal of Symbolic Computation, 51(??):55–62, April 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001150.

Brown:2015:CSC

[BK15] Christopher W. Brown and Marek Kosta. Constructing a single cell in cylindrical algebraic decomposition. Journal of Symbolic Computation, 70(??):14– 48, September/October 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000923.

Buse:2016:REP

[BK16] Laurent Busé and Anna Karasoulou. Resultant of an equivariant polynomial system with respect to the symmetric group. Journal of Symbolic Computation, 76(??): 142–157, September/October 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001315.

Bright:2021:ACA

[BKG21] Curtis Bright, Ilias Kotsireas, and Vijay Ganesh. Applying computer algebra systems with SAT solvers to the Williamson conjecture. Journal of Symbolic Computation, 100(??):187–209, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300902.

Bright:2021:CGP

[BKHG21] Curtis Bright, Ilias Kotsireas, Albert Heinle, and Vijay Ganesh. Complex Golay pairs up to length 28: a search via computer algebra and programmatic SAT. Journal of Symbolic Computation, 102(??):153– 172, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711930121X.

Benanav:1987:CMP

[BKN87] Dan Benanav, Deepak Kapur, and Paliath Narendran. Complexity of matching problems. Journal of Symbolic Computation, 3(1–2):203–216, February/April 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Rewriting techniques and applications (Dijon, 1985).

Botana:2019:SID

[BKR19] Francisco Botana, Zoltán Kovács, and Tomas Recio. Special issue on dynamic geometry and automated reasoning. Journal of Symbolic Computation, 97(??):1–2, ???? 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118301275.

Bauer:2017:PPH

[BKRW17] Ulrich Bauer, Michael Kerber, Jan Reininghaus, and Hubert Wagner. Phat — Persistent Homology Algorithms Toolbox. Journal of Symbolic Computation, 78(??):76– 90, January/February 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300098.

Blumlein:2012:SSA

[BKSS12] Johannes Blümlein, Sebastian Klein, Carsten Schneider, and Flavia Stan. A symbolic summation approach to Feynman integral calculus. Journal of Symbolic Computation, 47(10):1267–1289, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002471.

Bostan:2020:SJP

[BKSV20] A. Bostan, T. Krick, A. Szanto, and M. Valdettaro. Subresultants of (x−α)^{m} and (x−β)^{n}, Jacobi polynomials and complexity. Journal of Symbolic Computation, 101(??):330– 351, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301063.

Becedas:2024:VDA

[BKV24] Adrian Becedas, Kathlén Kohn, and Lorenzo Venturello. Voronoi diagrams of algebraic varieties under polyhedral norms. Journal of Symbolic Computation, 120(??):Article 102229, January/February 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000366.

Barth:2020:ACF

[BKW20] Dominik Barth, Joachim König, and Andreas Wenz. An approach for computing families of multi-branch-point covers and applications for symplectic Galois groups. Journal of Symbolic Computation, 101(??):352–366, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301087.

Butler:1985:GBA

[BL85] G. Butler and C. W. H. Lam. A general backtrack algorithm for the isomorphism problem of combinatorial objects. Journal of Symbolic Computation, 1(4):363–382 (or 363– 381??), December 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Burris:1993:EPF

[BL93] Stanley Burris and John Lawrence. The equivalence problem for finite rings. Journal of Symbolic Computation, 15(1):67–72 (or 67–71??), January 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Buekenhout:1996:LFP

[BL96] Francis Buekenhout and Dimitri Leemans. On the list of finite primitive permutation groups of degree ≤ 50. Journal of Symbolic Computation, 22(2):215–226 (or 215– 225??), August 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Beckermann:1998:FNS

[BL98a] B. Beckermann and G. Labahn.

A fast and numerically stable Euclidean-like algorithm for detecting relatively prime numerical polynomials. Journal of Symbolic Computation, 26 (6):691–714, December 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Beckermann:1998:WTN

[BL98b] B. Beckermann and G. Labahn.

When are two numerical polynomials relatively prime? Journal of Symbolic Computation, 26(6):677–690, December 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Baaz:2000:CER

[BL00] Matthias Baaz and Alexander Leitsch. Cut-elimination and redundancy-elimination by resolution. Journal of Symbolic Computation, 29 (2):149–177, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0359;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0359/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0359/ref.

Baaz:2006:TCA

[BL06a] Matthias Baaz and Alexander Leitsch. Towards a clausal analysis of cut-elimination. Journal of Symbolic Computation, 41(3–4):381–410, March/April 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Belitskii:2006:UTP

[BL06b] G. Belitskii and R. Lipyanski. The unification type of the Pythagorean equation in varieties of nilpotent rings. Journal of Symbolic Computation, 41(1):67–79, January 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boffi:2012:CGB

[BL12] Giandomenico Boffi and Alessandro Logar. Computing Gröbner bases of pure binomial ideals via submodules of Z^{n}. Journal of Symbolic Computation, 47(10): 1297–1308, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002495.

Boffi:2017:BBL

[BL17] Giandomenico Boffi and Alessandro Logar. Border bases for lattice ideals. Journal of Symbolic Computation, 79 (part 1)(??):43–56, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300840.

Baarnhielm:2012:PRP

[BLG12] Henrik Bäärnhielm and C. R.

Leedham-Green. The Product Replacement Prospector. Journal of Symbolic Computation, 47(1):64–75, January 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001295.

Boulier:2016:ANF

[BLL^{+}16] François Boulier, François Lemaire, Joseph Lallemand, Georg Regensburger, and Markus Rosenkranz. Additive normal forms and integration of differential fractions. Journal of Symbolic Computation, 77(??):16–38, November/December 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116000031.

Boulier:2010:CDC

[BLM10] François Boulier, François Lemaire, and Marc Moreno Maza. Computing differential characteristic sets by change of ordering. Journal of Symbolic Computation, 45(1):124– 149, January 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bian:2022:GAS

[BLM22] Ke-Xin Bian, Yue Liu, and Yan-Ping Mu. q-Gosper algorithm and simple Bailey pairs. Journal of Symbolic Computation, 113(??): 39–52, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000086.

Boulier:2019:ETR

[BLPM19] François Boulier, François Lemaire, Adrien Poteaux, and Marc Moreno Maza. An equivalence theorem for regular differential chains. Journal of Symbolic Computation, 93 (??):34–55, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300415.

Bouzidi:2015:SLF

[BLPR15] Yacine Bouzidi, Sylvain Lazard, Marc Pouget, and Fabrice Rouillier. Separating linear forms and Rational Univariate Representations of bivariate systems. Journal of Symbolic Computation, 68 (part 1)(??):84–119, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000595.

Bigatti:1999:CTI

[BLR99] A. M. Bigatti, R. La Scala, and L. Robbiano. Computing Toric ideals. Journal of Symbolic Computation, 27(4): 351–366, April 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bertone:2013:BOC

[BLR13] Cristina Bertone, Paolo Lella, and Margherita Roggero. A Borel open cover of the Hilbert scheme. Journal of Symbolic Computation, 53(??):119–135, June 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000023.

Bostan:2017:MBS

[BLS17] Alin Bostan, Pierre Lairez, and Bruno Salvy. Multiple binomial sums. Journal of Symbolic Computation, 80 (part 2)(??):351–386, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711630030X.

Birmpilis:2023:FAC

[BLS23] Stavros Birmpilis, George Labahn, and Arne Storjohann. A fast algorithm for computing the Smith normal form with multipliers for a nonsingular integer matrix. Journal of Symbolic Computation, 116(??):146–182, May/ June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000955.

Beltrametti:2022:QMS

[BLT22] Mauro Carlo Beltrametti, Alessandro Logar, and Maria Laura Torrente. Quartic monoid surfaces with maximum number of lines. Journal of Symbolic Computation, 109(??): 250–258, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300602.

Beckermann:2006:NFG

[BLV06] Bernhard Beckermann, George Labahn, and Gilles Villard. Normal forms for general polynomial matrices. Journal of Symbolic Computation, 41 (6):708–737, June 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Berthe:2016:PAP

[BLV16] Valérie Berthé, Loïck Lhote, and Brigitte Vallée. Probabilistic analyses of the plain multiple gcd algorithm. Journal of Symbolic Computation, 74(??):425–474, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000899.

Berthe:2018:BGA

[BLV18] Valérie Berthé, Loïck Lhote, and Brigitte Vallée. The Brun gcd algorithm in high dimensions is almost always subtractive. Journal of Symbolic Computation, 85(??): 72–107, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300688.

Bush:2003:CLK

[BLW03] M. R. Bush, M. Leeming, and R. F. C. Walters. Computing left Kan extensions.

Journal of Symbolic Computation, 35(2):107–126, February 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bayer:1988:SBG

[BM88] David Bayer and Ian Morrison. Standard bases and geometric invariant theory. I. initial ideals and state polytopes. Journal of Symbolic Computation, 6(2–3):209–218 (or 209– 217??), October/December 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Brien:2000:CS

[BM00] S. M. Brien and A. P. Martin. A calculus for schemas in Z. Journal of Symbolic Computation, 30(1):63–91, July 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0347;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0347/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0347/ref.

Binder:2001:AFN

[BM01] Franz Binder and Peter Mayr. Algorithms for finite near-rings and their Ngroups. Journal of Symbolic Computation, 32(1–2):

23–38, July/August 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0449;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0449/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0449/ref.

Bogvad:2004:ACW

[BM04] Rikard Bøgvad and Thomas Meyer. On algorithmically checking whether a Hilbert series comes from a complete intersection. Journal of Symbolic Computation, 38(6): 1487–1506, December 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boston:2010:WDC

[BM10] Nigel Boston and Gary McGuire. The weight distributions of cyclic codes with two zeros and zeta functions. Journal of Symbolic Computation, 45(7):723–733, July 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bakshi:2016:ESS

[BM16a] Gurmeet K. Bakshi and Sugandha Maheshwary. Extremely strong Shoda pairs with GAP. Journal of Symbolic Computation, 76(??): 97–106, September/October 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115001285.

Bucero:2016:BBR

[BM16b] Marta Abril Bucero and Bernard Mourrain. Border basis relaxation for polynomial optimization. Journal of Symbolic Computation, 74 (??):378–399, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000772.

Barkatou:2019:FSS

[BM19] Moulay A. Barkatou and Suzy S. Maddah. Formal solutions of singularlyperturbed linear differential systems. Journal of Symbolic Computation, 94(??): 183–209, September/October 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300944.

Beder:2011:ILP

[BMNB^{+}11] Jesse Beder, Jason McCullough, Luis Núñez-Betancourt, Alexandra Seceleanu, Bart Snapp, and Branden Stone. Ideals with larger projective dimension and regularity. Journal of Symbolic Computation, 46(10):1105–1113, October 2011. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000824.

Beth:2006:CRI

[BMQS06] Thomas Beth, Jörn Müller-Quade, and Rainer Steinwandt. Computing restrictions of ideals in finitely generated k-algebras by means of Buchberger’s algorithm. Journal of Symbolic Computation, 41(3–4):372–380, March/April 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Batra:2017:ILB

[BMS17] Prashant Batra, Maurice Mignotte, and Doru Stefanescu. Improvements of Lagrange’s bound for polynomial roots. Journal of Symbolic Computation, 82(??):19– 25, September/October 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301675.

Bohm:2020:CRS

[BMS20] Janko Böhm, Magdaleen S.

Marais, and Andreas Steenpaß. The classification of real singularities using Singular. Part III: Unimodal singularities of corank 2. Journal of Symbolic Computation, 99(??):250–282, July/August 2020. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300562.

Buse:2021:MFR

[BMT21] Laurent Busé, Angelos Mantzaflaris, and Elias Tsigaridas. Matrix formulæ for resultants and discriminants of bivariate tensor-product polynomials. Journal of Symbolic Computation, 98(??):65–83, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300720.

Baader:2001:UCT

[BN01] Franz Baader and Paliath Narendran. Unification of concept terms in description logics. Journal of Symbolic Computation, 31 (3):277–305, March 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0426;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.0426/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0426/ref.

Bila:2004:NPF

[BN04] Nicoleta Bîlă and Jitse Niesen. On a new procedure for finding nonclassical symmetries. Journal of Symbolic Computation, 38(6): 1523–1533, December 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bates:2017:DHS

[BNN17] Daniel J. Bates, Andrew J.

Newell, and Matthew E. Niemerg. Decoupling highly structured polynomial systems. Journal of Symbolic Computation, 79 (part 3)(??): 508–515, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300530.

Bernhardt:2022:CCC

[BNRW22] Dominik Bernhardt, Alice C. Niemeyer, Friedrich Rober, and Lucas Wollenhaupt. Conjugacy classes and centralisers in wreath products. Journal of Symbolic Computation, 113(??): 97–125, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000177.

Bertolini:2018:BSC

[BNT18] Marina Bertolini, Roberto Notari, and Cristina Turrini. The Bordiga surface as critical locus for 3-view reconstructions. Journal of Symbolic Computation, 91(??):74–97, ???? 2018. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300798.

Bellia:1999:LLC

[BO99] M. Bellia and M. E. Occhiuto. Lazy linear combinatorial unification. Journal of Symbolic Computation, 27(2): 185–206, February 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ballico:2004:CMG

[BO04] E. Ballico and F. Orecchia. Computing minimal generators of the ideal of a general projective curve. Journal of Symbolic Computation, 37(3): 295–304, March 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bahloul:2010:LBS

[BO10] Rouchdi Bahloul and Toshinori Oaku. Local Bernstein– Sato ideals: Algorithm and examples. Journal of Symbolic Computation, 45(1):46– 59, January 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bodin:2004:CMN

[Bod04] Arnaud Bodin. Computation of Milnor numbers and critical values at infinity. Journal of Symbolic Computation, 38(5): 1417–1427, November 2004.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bokut:2008:GSB

[Bok08]  L. A. Bokut.  Gröbner–

Shirshov basis for the braid group in the Artin–Garside generators. Journal of Symbolic Computation, 43(6–7): 397–405, June/July 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bonacina:1996:RPD

[Bon96] Maria Paola Bonacina. On the reconstruction of proofs in distributed theorem proving: a modified clause-diffusion method. Journal of Symbolic Computation, 21(4/5/6):507– 522, April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Bonacina:2005:TUM

[Bon05] Maria Paola Bonacina. Towards a unified model of search in theorem-proving: subgoal-reduction strategies. Journal of Symbolic Computation, 39(2):209–255, February 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Book:1987:TSR

[Boo87]  Ronald V. Book.  Thue

systems as rewriting systems. Journal of Symbolic Computation, 3(1–2):39–68, February/April 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Rewriting techniques and applications (Dijon, 1985).

Bors:2022:COC

[Bor22] Alexander Bors. Computation of orders and cycle lengths of automorphisms of finite solvable groups. Journal of Symbolic Computation, 108(??): 117–136, January/February 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300195.

Boston:1997:UCT

[Bos97] N. Boston. A use of computers to teach group theory and introduce students to research. Journal of Symbolic Computation, 23(5–6):453–458, May/ June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bosma:2001:SIC

[Bos01] Wieb Bosma. Special issue on computational algebra and number theory: Proceedings of the Second Magma Conference: Foreword of the Guest Editor. Journal of Symbolic Computation, 31(1–2):1, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0435;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0435/
pdf.

Boudet:1993:CUA

[Bou93] Alexandre Boudet. Combining unification algorithms. Journal of Symbolic Computation, 16(6):597–626, December 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bouhoula:1997:ATP

[Bou97] Adel Bouhoula. Automated theorem proving by test set induction. Journal of Symbolic Computation, 23(1):47–78 (or 47–77??), January 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

BoydelaTour:1992:ORC

[Boy92] Thierry Boy de la Tour. An optimality result for clause form translation. Journal of Symbolic Computation, 14(4): 283–302 (or 283–301??), October 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boyd:1993:BHFa

[Boy93a] David W. Boyd. Bounds for the height of a factor of a polynomial in terms of Bombieri’s norms: I. the largest factor. Journal of Symbolic Computation, 16(2):115–130, August 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boyd:1993:BHFb

[Boy93b] David W. Boyd. Bounds for the height of a factor of a polynomial in terms of Bombieri’s norms: II. the smallest factor. Journal of Symbolic Computation, 16(2):131–145, August 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boyd:1993:CLS

[Boy93c] John P. Boyd. Chebyshev and Legendre spectral methods in algebraic manipulation languages. Journal of Symbolic Computation, 16(4):377– 399, October 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bachmair:1985:TOA

[BP85] Leo Bachmair and David A. Plaisted. Termination ordering for associative commutative rewriting systems. Journal of Symbolic Computation, 1(4):329–350 (or 329– 349??), December 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Breuer:1998:FPP

[BP98] Thomas Breuer and Götz Pfeiffer. Finding possible permutation characters. Journal of Symbolic Computation, 26(3):343–354, 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Barkatou:1999:ACR

[BP99a] Moulay Barkatou and Eckhard Pflügel. An algorithm computing the regular formal solutions of a system of linear differential equations. Journal of Symbolic Computation, 28(4–5): 569–587, October/November 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0315/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0315/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0315/production/ref.

Bauer:1999:MMH

[BP99b] Andrej Bauer and Marko Petkovšek. Multibasic and mixed hypergeometric Gospertype algorithms. Journal of Symbolic Computation, 28(4–5):711–736, October/November 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0321/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.

                                        1999.0321/production/pdf;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0321/production/ref.

Bratus:2000:FCR

[BP00] Sergey Bratus and Igor Pak. Fast constructive recognition of a black box group isomorphic to S_{n} or A_{n} using Goldbach’s conjecture. Journal of Symbolic Computation, 29(1):33–57, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0295;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0295/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0295/ref.

Beltran:2007:PDS

[BP07] Carlos Beltrán and Luis Miguel Pardo. On the probability distribution of singular varieties of given corank. Journal of Symbolic Computation, 42 (1–2):4–29, January/February 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Barkatou:2009:MSR

[BP09a] Moulay A. Barkatou and Eckhard Pflügel. On the Moser- and super-reduction algorithms of systems of linear differential equations and their complexity. Journal of Symbolic Computation, 44(8): 1017–1036, August 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bulygin:2009:BDD

[BP09b] Stanislav Bulygin and Ruud Pellikaan. Bounded distance decoding of linear errorcorrecting codes with Gröbner bases. Journal of Symbolic Computation, 44(12):1626– 1643, December 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Brouwer:2010:IBD

[BP10a] Andries E. Brouwer and Mihaela Popoviciu. The invariants of the binary decimic. Journal of Symbolic Computation, 45(8):837–843, August 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Brouwer:2010:IBN

[BP10b] Andries E. Brouwer and Mihaela Popoviciu. The invariants of the binary nonic. Journal of Symbolic Computation, 45(6):709–720, June 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Blanco:2011:SAM

[BP11] Víctor Blanco and Justo Puerto. Some algebraic methods for solving multiobjective polynomial integer programs. Journal of Symbolic Computation, 46(5):511–533, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001720.

Bona:2023:PAS

[BP23] Miklós Bóna and Jay Pantone. Permutations avoiding sets of patterns with long monotone subsequences. Journal of Symbolic Computation, 116(??):130–138, May/ June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000943.

Barcau:2024:CPI

[BP24] Mugurel Barcau and Vicentiu Pasol. Computing primitive idempotents in finite commutative rings and applications. Journal of Symbolic Computation, 123(??):??, July/August 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123001086.

Blasco:2019:LPF

[BPD19] Angel Blasco and Sonia Pérez-Díaz. The limit point and the T -function. Journal of Symbolic Computation, 94(??):30– 51, September/October 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300737.

Bleylevens:2007:EID

[BPH07] Ivo Bleylevens, Ralf Peeters, and Bernard Hanzon. Efficiency improvement in an nD systems approach to polynomial optimization. Journal of Symbolic Computation, 42(1– 2):30–53, January/February 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Banti:2011:AVE

[BPT11] Federico Banti, Rosario Pugliese, and Francesco Tiezzi. An accessible verification environment for UML models of services. Journal of Symbolic Computation, 46 (2):119–149, February 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001355.

Baumgartner:2012:MEE

[BPT12] Peter Baumgartner, Björn Pelzer, and Cesare Tinelli. Model evolution with equality — revised and implemented. Journal of Symbolic Computation, 47(9): 1011–1045, September 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002240.

Berczes:2006:PNF

[BPZ06] A. Bérczes, A. Pethő, and V. Ziegler. Parameterized norm form equations with arithmetic progressions. Journal of Symbolic Computation, 41(7):790–810, July 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boffgen:1987:CDP

[BR87] R. Böffgen and M. A. Reichert. Computing the decomposition of primes p and p-adic absolute values in semisimple algebras over Q. Journal of Symbolic Computation, 4(1):3–10, August 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Brundu:1988:CGS

[BR88] Michela Brundu and Fabio Rossi. On the computation of generalised standard bases. Journal of Symbolic Computation, 6(2–3):323–344 (or 323– 343??), October/December 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Boffi:2006:LGB

[BR06a] Giandomenico Boffi and Fabio Rossi. Lexicographic Gröbner bases for transportation problems of format r × 3 × 3. Journal of Symbolic Computation, 41(3–4):336–356, March/April 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Buot:2006:CLS

[BR06b] Max-Louis G. Buot and Donald St. P. Richards. Counting and locating the solutions of polynomial systems of maximum likelihood equations, I. Journal of Symbolic Computation, 41(2):234–244, February 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Barakat:2009:CCC

[BR09a] Mohamed Barakat and Daniel Robertz. conley: Computing connection matrices in Maple. Journal of Symbolic Computation, 44(5):540–557, May 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Briand:2009:MVF

[BR09b] Emmanuel Briand and Mercedes Rosas. Milne’s volume function and vector symmetric polynomials. Journal of Symbolic Computation, 44(5): 583–590, May 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Basu:2010:BRB

[BR10] Saugata Basu and Marie-Françoise Roy. Bounding the radii of balls meeting every connected component of semi-algebraic sets. Journal of Symbolic Computation, 45(12):1270–1279, December 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Buchberger:2012:TPA

[BR12] Bruno Buchberger and Markus Rosenkranz. Transforming problems from analysis to algebra: a case study in linear boundary problems. Journal of Symbolic Computation, 47(6):589–609, June 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711100215X.

Ballantyne:2013:NCI

[BR13a] John Ballantyne and Peter Rowley. A note on computing involution centralizers. Journal of Symbolic Computation, 54(??):1–8, July 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000035.

Bernardi:2013:CRC

[BR13b] Alessandra Bernardi and Kristian Ranestad. On the cactus rank of cubics forms. Journal of Symbolic Computation, 50(??):291–297, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001332.

Bogner:2013:SRL

[BR13c] Michael Bogner and Stefan Reiter. On symplectically rigid local systems of rank four and Calabi–Yau operators. Journal of Symbolic Computation, 48(??):64–100, January 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000788.

Boocher:2015:RTI

[BR15] Adam Boocher and Elina Robeva. Robust toric ideals. Journal of Symbolic Computation, 68 (part 1) (??):254–264, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711400087X.

Bigatti:2022:SSS

[BR22] Anna Maria Bigatti and Lorenzo Robbiano. Saturations of subalgebras, SAGBI bases, and U-invariants. Journal of Symbolic Computation, 109(??):259–282, March/April 2022. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300614.

Braatelund:2024:CCT

[Brå24]  Martin Bråtelund.  Criti-

cal configurations for two projective views, a new approach. Journal of Symbolic Computation, 120(??):Article 102226, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000330.

Bremner:1986:FCW

[Bre86] Murray R. Bremner. Fast computation of weight multiplicities. Journal of Symbolic Computation, 2(4):357– 362, December 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bright:2006:BGD

[Bri06] Martin Bright. Brauer groups of diagonal quartic surfaces. Journal of Symbolic Computation, 41(5):544–558, May 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bouziane:2001:UDD

[BRM01] Driss Bouziane, Abdelilah Kandri Rody, and Hamid Maârouf. Unmixed-dimensional decomposition of a finitely generated perfect differential ideal. Journal of Symbolic Computation, 31(6):631–649, June 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1562;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.1562/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1562/ref.

Bronstein:1990:IEF

[Bro90a] Manuel Bronstein. Integration of elementary functions. Journal of Symbolic Computation, 9(2):117–174 (or 117– 173??), February 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bronstein:1990:TRD

[Bro90b] Manuel Bronstein. The transcendental Risch differential equation. Journal of Symbolic Computation, 9(1):49– 60, January 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bronstein:1992:SLO

[Bro92] Manuel Bronstein. On solutions of linear ordinary differential equations in their coefficient field. Journal of Symbolic Computation, 13(4): 413–440 (or 413–439??), April 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bronstein:2000:SLO

[Bro00] Manuel Bronstein. On solutions of linear ordinary difference equations in their coefficient field. Journal of Symbolic Computation, 29(6):841–877, June 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0368;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0368/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0368/ref.

Brown:2001:IPC

[Bro01a] Christopher W. Brown. Improved projection for cylindrical algebraic decomposition. Journal of Symbolic Computation, 32(5): 447–465, November 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0463;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0463/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0463/ref.

Brown:2001:SCC

[Bro01b] Christopher W. Brown. Simple CAD construction and its applications. Journal of Symbolic Computation, 31(5):521–547, May 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0394;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0394/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0394/ref.

Brooksbank:2003:CRC

[Bro03] Peter A. Brooksbank. Constructive recognition of classical groups in their natural representation. Journal of Symbolic Computation, 35(2):195– 239, February 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bronstein:2007:STP

[Bro07] Manuel Bronstein. Structure theorems for parallel integration. Journal of Symbolic Computation, 42(7):757–769, July 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Brown:2012:FST

[Bro12] Christopher W. Brown. Fast simplifications for Tarski formulas based on monomial inequalities. Journal of Symbolic Computation, 47(7): 859–882, July 2012. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002045.

Brueggeman:2001:SNF

[Bru01] Sharon Brueggeman. Septic number fields which are ramified only at one small prime. Journal of Symbolic Computation, 31(5):549–555, May 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0440;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0440/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0440/ref.

Brysiewicz:2021:NCP

[Bry21] Taylor Brysiewicz. Necklaces count polynomial parametric osculants. Journal of Symbolic Computation, 103(??): 95–107, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301385.

Book:1986:UET

[BS86] R. V. Book and J. H. Siekmann. On unification: Equational theories are not bounded. Journal of Symbolic Computation, 2(4):317– 324, December 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Buttner:1987:EBE

[BS87] Wolfram Büttner and Helmut Simonis. Embedding Boolean expressions into logic programming. Journal of Symbolic Computation, 4(2): 191–205, October 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bayer:1988:CCS

[BS88] David Bayer and Michael Stillman. On the complexity of computing syzygies. Journal of Symbolic Computation, 6(2–3):135–148 (or 135– 147??), October/December 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Bach:1990:FPU

[BS90a] Eric Bach and Victor Shoup.

Factoring polynomials using fewer random bits. Journal of Symbolic Computation, 9(3):229–240 (or 229– 239??), March 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Beckmann:1990:CGA

[BS90b] P. Beckmann and J. Stückrad.

The concept of Gröbner algebras. Journal of Symbolic Computation, 10(5):465– 479 (or 465–480??), November 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bayer:1992:CHF

[BS92] Dave Bayer and Mike Stillman. Computation of Hilbert functions. Journal of Symbolic Computation, 14(1):31–50, July 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Baader:1996:UUD

[BS96] Franz Baader and Klaus U. Schulz. Unification in the union of disjoint equational theories: Combining decision procedures. Journal of Symbolic Computation, 21(2):211– 244 (or 211–243??), February 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bodnar:2000:ARS

[BS00] Gábor Bodnár and Josef Schicho. Automated resolution of singularities for hypersurfaces. Journal of Symbolic Computation, 30 (4):401–428, October 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0414;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0414/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0414/ref.

Bodnar:2001:TCT

[BS01] Gábor Bodnár and Josef Schicho. Two computational techniques for singularity resolution. Journal of Symbolic Computation, 32(1–2):39–54, July/August 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0452;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0452/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0452/ref.

Burgisser:2009:CCC

[BS09] Peter Bürgisser and Peter Scheiblechner. On the complexity of counting components of algebraic varieties. Journal of Symbolic Computation, 44(9):1114–1136, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Batra:2010:BAP

[BS10] Prashant Batra and Vikram Sharma. Bounds on absolute positiveness of multivariate polynomials. Journal of Symbolic Computation, 45(6): 617–628, June 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bruns:2015:CGE

[BS15] Winfried Bruns and Christof Söger. The computation of generalized Ehrhart series in Normaliz. Journal of Symbolic Computation, 68 (part 2)(??):75–86, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000728.

Batra:2017:NOS

[BS17a] Prashant Batra and Vikram Sharma. Near optimal subdivision algorithms for real root isolation. Journal of Symbolic Computation, 83(??): 4–35, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301237.

Blekherman:2017:ERH

[BS17b] Grigoriy Blekherman and Rainer Sinn. Extreme rays of Hankel spectrahedra for ternary forms. Journal of Symbolic Computation, 79 (part 1)(??):23–42, March/April 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300827.

Bach:2018:BSG

[BS18] Eric Bach and Bryce Sandlund. Baby-step giant-step algorithms for the symmetric group. Journal of Symbolic Computation, 85(??): 55–71, March/April 2018.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300676.

Brodmann:2021:FBR

[BS21] Markus Brodmann and Peter Schenzel. Families of blowups of the real affine plane: Classification, isotopies and visualization. Journal of Symbolic Computation, 104 (??):874–898, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301140.

Bilski:2022:EAS

[BS22a] Marcin Bilski and Peter Scheiblechner. Effective approximation of the solutions of algebraic equations. Journal of Symbolic Computation, 109 (??):144–176, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000560.

Bortner:2022:SIS

[BS22b] Cashous Bortner and Seth Sullivant. Structural identifiability of series-parallel LCR systems. Journal of Symbolic Computation, 112(??): 79–104, September/October 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000025.

Becker:2024:CSP

[BS24] Ruben Becker and Michael Sagraloff. Counting solutions of a polynomial system locally and exactly. Journal of Symbolic Computation, 120(??):Article 102222, January/February 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000299.

Bouissou:2012:AAF

[BSC12] Olivier Bouissou, Yassamine Seladji, and Alexandre Chapoutot. Acceleration of the abstract fixpoint computation in numerical program analysis. Journal of Symbolic Computation, 47(12): 1479–1511, December 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002537.

Becker:2018:NOS

[BSSY18] Ruben Becker, Michael Sagraloff, Vikram Sharma, and Chee Yap. A near-optimal subdivision algorithm for complex root isolation based on the Pellet test and Newton iteration. Journal of Symbolic Computation, 86 (??):51–96, May/June 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300378.

Burity:2016:CVS

[BST16] Ricardo Burity, Aron Simis, and Stefan O. Tohăneanu. On a conjecture of Vasconcelos via Sylvester forms. Journal of Symbolic Computation, 77(??):39–62, November/December 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000067.

Barkatou:2021:FRS

[BSW21a] Moulay A. Barkatou, Joelle Saade, and Jacques-Arthur Weil. Formal reduction of singular linear differential systems using eigenrings: a refined approach. Journal of Symbolic Computation, 102(??):231–258, January/February 2021. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301282.

Bromberger:2021:CTA

[BSW21b] Martin Bromberger, Thomas Sturm, and Christoph Weidenbach. A complete and terminating approach to linear integer solving. Journal of Symbolic Computation, 100(??):102–136, ????

2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300872.

Brown:1994:CSC

[BT94] Ronald Brown and Andrew Tonks. Calculations with simplicial and cubical groups in AXIOM. Journal of Symbolic Computation, 17(2):159– 180 (or 159–179??), February 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bertot:1998:GAB

[BT98] Y. Bertot and L. Théry. A generic approach to building user interfaces for theorem provers. Journal of Symbolic Computation, 25(2):161– 194, February 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bayer:2009:RSM

[BT09] Dave Bayer and Amelia Taylor. Reverse search for monomial ideals. Journal of Symbolic Computation, 44(10): 1477–1486, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Borges-Trenard:2000:CGB

[BTBQM00] M. A. Borges-Trenard, M. Borges-Quintana, and T. Mora. Computing gröbner bases by FGLM techniques in a noncommutative setting. Journal of Symbolic Computation, 30(4):429–449, October 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0415;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0415/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0415/ref.

Boyle:2002:REM

[BTG02] P. P. Boyle, W. Tian, and Fred Guan. The Riccati equation in mathematical finance. Journal of Symbolic Computation, 33(3):343–355, March 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0508;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0508/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0508/ref.

Beauzamy:1993:PFS

[BTW93] Bernard Beauzamy, Vilmar Trevisan, and Paul S. Wang. Polynomial factorization: Sharp bounds, efficient algorithms. Journal of Symbolic Computation, 15(4): 393–414 (or 393–413??), April 1993. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-

855X (electronic).

Bikker:1999:BCR

[BU99] P. Bikker and A. Yu. Uteshev. On the Bézout construction of the resultant. Journal of Symbolic Computation, 28(1–2):45–88, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0267;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0267/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0267/ref.

Boucher:2009:CSP

[BU09] Delphine Boucher and Felix Ulmer. Coding with skew polynomial rings. Journal of Symbolic Computation, 44(12):1644–1656, December 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boucher:2014:SDS

[BU14] Delphine Boucher and Felix Ulmer. Self-dual skew codes and factorization of skew polynomials. Journal of Symbolic Computation, 60(??):47–61, January 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001223.

Buchberger:1987:HBF

[Buc87] Bruno Buchberger. History and basic features of the critical-pair/completion procedure. Journal of Symbolic Computation, 3(1–2): 3–38, February/April 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Rewriting techniques and applications (Dijon, 1985).

Buchberger:1992:SIJ

[Buc92] B. Buchberger. Special issue of JSC on “Symbolic Computation in Combinatorics”

— foreward of the Editorin-Chief. Journal of Symbolic Computation, 14(2–3): 121–122, August/September 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Buchberger:2006:BBP

[Buc06a] Bruno Buchberger. Bruno Buchberger’s PhD thesis 1965: An algorithm for finding the basis elements of the residue class ring of a zero dimensional polynomial ideal. Journal of Symbolic Computation, 41(3–4):475–511, March/April 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Buchberger:2006:CTM

[Buc06b] Bruno Buchberger. Comments on the translation of my PhD thesis. Journal of Symbolic Computation, 41(3–4): 471–474, March/April 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Burckert:1989:MSC

[Bür89] Hans-Jürgen J. Bürckert. Matching — a special case of unification? Journal of Symbolic Computation, 8(5):523– 536, November 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Burris:1992:DVS

[Bur92] Stanley Burris. Discriminator varieties and symbolic computation. Journal of Symbolic Computation, 13(2):175– 208 (or 175–207??), February 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Burckel:2001:SMB

[Bur01]  Serge Burckel.  Syntac-

tical methods for braids of three strands. Journal of Symbolic Computation, 31(5):557–564, May 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0473;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0473/
pdf;   http://www.idealibrary.

com/links/doi/10.1006/jsco.
2000.0473/ref.

Burckel:2003:RBI

[Bur03] Serge Burckel. A result on braids via the investigation of a rewriting system. Journal of Symbolic Computation, 35(2): 169–175, February 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Burckel:2004:EDA

[Bur04] Serge Burckel. Elementary decompositions of arbitrary maps over finite sets. Journal of Symbolic Computation, 37 (3):305–310, March 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Burr:2016:CAE

[Bur16] Michael A. Burr. Continuous amortization and extensions: With applications to bisection-based root isolation. Journal of Symbolic Computation, 77(??): 78–126, November/December 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000080.

Busch:2009:LBD

[Bus09] J. Busch. Lower bounds for decision problems in imaginary, norm-euclidean quadratic integer rings. Journal of Symbolic Computation, 44(6):683–699, June 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Butler:1985:ECG

[But85] Gregory Butler. Effective computations with group homomorphisms. Journal of Symbolic Computation, 1(2): 143–158 (or 143–157??), June 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Butler:1988:PHA

[But88] Gregory Butler. A proof of Holt’s algorithm. Journal of Symbolic Computation, 5(3):275–284 (or 275–283??), June 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Butler:1993:TGD

[But93] Greg Butler. The transitive groups of degree fourteen and fifteen. Journal of Symbolic Computation, 16(5):413– 422, November 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Baines:2003:ACS

[BV03] Rosemary Baines and Peter Vámos. An algorithm to compute the set of characteristics of a system of polynomial equations over the integers. Journal of Symbolic Computation, 35(3):269–279, March 2003. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Bokut:2006:GSB

[BV06] Leonid Bokut and Andrei Vesnin. Gröbner–Shirshov bases for some braid groups. Journal of Symbolic Computation, 41(3–4):357–371, March/April 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Berson:2003:AFC

[BvdE03] Joost Berson and Arno van den Essen. An algorithm to find a coordinate’s mate. Journal of Symbolic Computation, 36(6):835–843, December 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Brown:2021:SPT

[BVE21] Christopher W. Brown and Fernando Vale-Enriquez. From simplification to a partial theory solver for non-linear real polynomial constraints. Journal of Symbolic Computation, 100(??):72–101, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300860.

Baldoni:2018:CDK

[BVW18] V. Baldoni, M. Vergne, and M. Walter. Computation of dilated Kronecker coefficients. Journal of Symbolic Computation, 84(??): 113–146, January/February 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300330.

Blankertz:2013:CCD

[BvzGZ13] Raoul Blankertz, Joachim von zur Gathen, and Konstantin Ziegler. Compositions and collisions at degree p^{2}. Journal of Symbolic Computation, 59(??):113–145, December 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000928.

Buchmann:1987:PIT

[BW87] Johannes B. Buchmann and H. C. Williams. On principal ideal testing in algebraic number fields. Journal of Symbolic Computation, 4(1):11–19, August 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Brown:2003:CHA

[BW03] Ronald Brown and Christopher D. Wensley. Computation and homotopical applications of induced crossed modules. Journal of Symbolic Computation, 35(1):59– 72, January 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Beeson:2005:MIC

[BW05] Michael Beeson and Freek Wiedijk. The meaning of infinity in calculus and computer algebra systems. Journal of Symbolic Computation, 39(5):523–538, May 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Barth:2022:EMB

[BW22] Dominik Barth and Andreas Wenz. On Elkies’ method for bounding the transitivity degree of Galois groups. Journal of Symbolic Computation, 108 (??):17–22, January/February 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000304.

Bajaj:1997:SAR

[BX97] Chandrajit L. Bajaj and Guoliang Xu. Spline approximations of RealAlgebraic surfaces. Journal of Symbolic Computation, 23(2–3):315– 333, February/March 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Bostan:2023:FCT

[BY23] Alin Bostan and Sergey Yurkevich. Fast computation of the N-th term of a qholonomic sequence and applications. Journal of Symbolic Computation, 115(??): 96–123, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000669.

Barton:1985:PDA

[BZ85] David R. Barton and Richard Zippel. Polynomial decomposition algorithms. Journal of Symbolic Computation, 1 (2):159–168, June 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Berarducci:1993:GU

[BZ93] Alessandro Berarducci and Marisa Venturini Zilli. Generalizations of unification. Journal of Symbolic Computation, 16(5):478–492 (or 479– 491??), November 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Baumgartner:2003:PFO

[BZ03] Peter Baumgartner and Hantao Zhang. Preface to first order theorem proving. Journal of Symbolic Computation, 36 (1–2):1–3, July/August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Canny:1990:GCP

[Can90] John Canny. Generalised characteristic polynomials.

Journal of Symbolic Computation, 9(3):241–250, March 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Caprasse:1990:RGF

[Cap90] H. Caprasse. Renormalization group, function iterations and computer algebra. Journal of Symbolic Computation, 9(1): 61–72, January 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Carlson:1999:TSI

[Car99]  B. C. Carlson.  Toward

symbolic integration of elliptic integrals. Journal of Symbolic Computation, 28 (6):739–753, December 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0336/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0336/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0336/production/ref.

Carlson:2001:CGC

[Car01] Jon F. Carlson. Calculating group cohomology: Tests for completion. Journal of Symbolic Computation, 31(1–2): 229–242, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.1003;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.1003/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.1003/ref.

Caruso:2015:RMD

[Car15] Xavier Caruso. Random matrices over a DVR and LU factorization. Journal of Symbolic Computation, 71(??): 98–123, November/December 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711400128X.

Caviness:1986:CAP

[Cav86] B. F. Caviness. Computer algebra: Past and future. Journal of Symbolic Computation, 2(3):217–236, September 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Caviness:2000:E

[Cav00] B. F. Caviness. Editorial. Journal of Symbolic Computation, 29(1):1, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0351;   http://www.
                                        idealibrary.com/links/doi/

10.1006/jsco.2000.0351/
pdf.

Cameron:1991:FRD

[CC91] P. J. Cameron and J. Cannon. Fast recognition of doubly transitive groups. Journal of Symbolic Computation, 12(4–5):459–474, October/November 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Caprotti:2001:ROI

[CC01] O. Caprotti and A. M. Cohen. On the role of Open-Math in interactive mathematical documents. Journal of Symbolic Computation, 32(4):351–364, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0466;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0466/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0466/ref.

Curran:2007:RDD

[CC07] Raymond Curran and Eduardo Cattani. Restriction of A-discriminants and dual defect toric varieties. Journal of Symbolic Computation, 42(1–2):115–135, January/February 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chang:2024:TFC

[CC24] Ke-Ming Chang and Kuo-Chang Chen. Toward finiteness of central configurations for the planar six-body problem by symbolic computations. (I) Determine diagrams and orders. Journal of Symbolic Computation, 123(??):??, July/August 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000913.

Cesaratto:2009:REA

[CCD^{+}09] Eda Cesaratto, Julien Clément, Benoît Daireaux, Loïck Lhote, Véronique Maume-Deschamps, and Brigitte Vallée. Regularity of the Euclid Algorithm; application to the analysis of fast GCD algorithms. Journal of Symbolic Computation, 44 (7):726–767, July 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2015:ETM

[CCF^{+}15] Shaoshi Chen, Frédéric Chyzak, Ruyong Feng, Guofeng Fu, and Ziming Li. On the existence of telescopers for mixed hypergeometric terms. Journal of Symbolic Computation, 68 (part 1)(??):1–26, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000558.

Cantone:2006:DAF

[CCG06] Domenico Cantone, Gianluca Cincotti, and Giovanni Gallo. Decision algorithms for fragments of real analysis. I. Continuous functions with strict convexity and concavity predicates. Journal of Symbolic Computation, 41(7):763–789, July 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cannon:1997:CSS

[CCH97] John J. Cannon, Bruce C.

Cox, and Derek F. Holt. Computing Sylow subgroups in permutation groups. Journal of Symbolic Computation, 24(3–4):303–316, September/ October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Cannon:2001:CSP

[CCH01] John J. Cannon, Bruce C.

Cox, and Derek F. Holt. Computing the subgroups of a permutation group. Journal of Symbolic Computation, 31(1–2):149–161, January/February 2001. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.

                                        com/links/doi/10.1006/jsco.
                                        2000.1012;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.1012/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.1012/ref.

Chu:2002:QFR

[CcK02] Huah Chu and Ming chang Kang. Quartic fields and radical extensions. Journal of Symbolic Computation, 34 (1):83–89, July 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2005:BIR

[CCL05] Falai Chen, David Cox, and Yang Liu. The μ-basis and implicitization of a rational parametric surface. Journal of Symbolic Computation, 39 (6):689–706, June 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cioni:1995:SCA

[CCM95] Gianna Cioni, Attilio Colagrossi, and Alfonso Miola. A sequent calculus for automated reasoning in symbolic computation systems. Journal of Symbolic Computation, 19(1/2/3):175–200 (or 175–199??), January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Carvalho:2018:ECC

[CCQ18] Cícero Carvalho, María Chara, and Luciane Quoos. On evaluation codes coming from a tower of function fields. Journal of Symbolic Computation, 89(??):121–128, November/December 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117301153.

Caboara:2011:LPC

[CCT11] Massimo Caboara, Fabrizio Caruso, and Carlo Traverso. Lattice Polly Cracker cryptosystems. Journal of Symbolic Computation, 46(5): 534–549, May 2011. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001732.

Coppersmith:1985:AF

[CD85] D. Coppersmith and J. H.

Davenport. An application of factoring. Journal of Symbolic Computation, 1(2): 241–243, June 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cohn:1987:ASM

[CD87] Harvey Cohn and Jesse Ira Deutsch. Application of symbolic manipulation to the Hecke transformations of modular forms in two variables, II. Journal of Symbolic Computation, 4(1):35–40, August 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2000:ACN

[CD00] Guoting Chen and Jean Della Dora. An algorithm for computing a new normal form for dynamical systems. Journal of Symbolic Computation, 29(3):393–418, March 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0305;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0305/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0305/ref.

Carminati:1992:IDE

[CDF92] John Carminati, John S. Devitt, and Greg J. Fee. Isogroups of differential equations using algebraic computing. Journal of Symbolic Computation, 14(1):103–120, July 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cicalo:2009:NAG

[CdG09] Serena Cicalò and Willem A. de Graaf. Non-associative Gröbner bases, finitely-presented Lie rings and the Engel condition, II. Journal of Symbolic Computation, 44(7):786–800, July 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2013:TDS

[CDM^{+}13a] Changbo Chen, James H.

Davenport, John P. May, Marc Moreno Maza, Bican Xia, and Rong Xiao. Triangular decomposition of semialgebraic systems. Journal of Symbolic Computation, 49(??):3–26, February 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002070.

Chen:2013:CSA

[CDM^{+}13b] Changbo Chen, James H.

Davenport, Marc Moreno Maza, Bican Xia, and Rong Xiao. Computing with semi-algebraic sets: Relaxation techniques and effective boundaries. Journal of Symbolic Computation, 52(??):72–96, May 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001307.

Cohen:1997:SAC

[CDO97] H. Cohen, F. Diaz y Diaz, and M. Olivier. Subexponential algorithms for class group and unit computations.

Journal of Symbolic Computation, 24(3–4):433–442 (or 433– 441??), September/October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Cohen:2001:AMF

[CDO01] Henri Cohen, Francisco Diaz y Diaz, and Michel Olivier. Algorithmic methods for finitely generated Abelian groups. Journal of Symbolic Computation, 31(1–2): 133–147, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.1014;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.1014/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.1014/ref.

Craciun:2009:TDS

[CDSS09] Gheorghe Craciun, Alicia Dickenstein, Anne Shiu, and Bernd Sturmfels. Toric dynamical systems. Journal of Symbolic Computation, 44 (11):1551–1565, November 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2021:ETR

[CDWZ21] Shaoshi Chen, Lixin Du, Rong-Hua Wang, and Chaochao Zhu. On the existence of telescopers for rational functions in three variables. Journal of Symbolic Computation, 104(??):494–522, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300870.

Chazelle:1985:OSC

[CE85] B. Chazelle and H. Edelsbrunner. Optimal solutions for a class of point retrieval problems. Journal of Symbolic Computation, 1(1): 47–56, March 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Collins:1995:ERN

[CE95] George E. Collins and Mark J.

Encarnación. Efficient rational number reconstruction. Journal of Symbolic Computation, 20(3):287–298 (or 287– 297??), September 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Collins:1996:ITF

[CE96] George E. Collins and Mark J.

Encarnación. Improved techniques for factoring univariate polynomials. Journal of Symbolic Computation, 21(3):313– 328 (or 313–327??), March 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cant:2019:PDM

[CE19] Alexander Cant and Bettina Eick. Polynomials describing the multiplication in finitely generated torsion-free nilpotent groups. Journal of Symbolic Computation, 92 (??):203–210, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300440.

Cannon:2004:SPG

[CELG04] John J. Cannon, Bettina Eick, and Charles R. Leedham-Green. Special polycyclic generating sequences for finite soluble groups. Journal of Symbolic Computation, 38(5): 1445–1460, November 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ceria:2018:BCM

[Cer18] Michela Ceria. Bar code for monomial ideals. Journal of Symbolic Computation, 91(??):30–56, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300774.

Ceria:2021:CDM

[Cer21] Michela Ceria. Combinatorial decompositions for monomial ideals. Journal of Symbolic Computation, 104(??):

630–652, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300948.

Capco:2023:PDP

[CES23] Jose Capco, Mohab Safey El Din, and Josef Schicho. Positive dimensional parametric polynomial systems, connectivity queries and applications in robotics. Journal of Symbolic Computation, 115(??): 320–345, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000748.

Clausen:1989:ESL

[CF89] Michael Clausen and Albrecht Fortenbacher. Efficient solution of linear Diophantine equations. Journal of Symbolic Computation, 8(1–2): 201–216, July/August 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cohen:1991:UMC

[CF91a] H. I. Cohen and J. P. Fitch.

Uses made of computer algebra in physics. Journal of Symbolic Computation, 11(3): 291–305, March 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cooperman:1991:SGT

[CF91b] Gene Cooperman and Larry Finkelstein. A strong generating test and short presentations for permutation groups. Journal of Symbolic Computation, 12(4–5): 475–498 (or 475–497??), October/November 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Cooperman:1994:RBC

[CF94] Gene Cooperman and Larry Finkelstein. A random base change algorithm for permutation groups. Journal of Symbolic Computation, 17(6): 513–528, June 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cremona:2009:EBQ

[CF09a] J. E. Cremona and T. A. Fisher. On the equivalence of binary quartics. Journal of Symbolic Computation, 44 (6):673–682, June 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Curtis:2009:SRE

[CF09b] R. T. Curtis and B. T. Fairbairn. Symmetric representation of the elements of the Conway group _{.}0. Journal of Symbolic Computation, 44(8): 1044–1067, August 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Char:1986:TIM

[CFG^{+}86] Bruce W. Char, G. J. Fee, Keith O. Geddes, Gaston H. Gonnet, and Michael B. Monagan. A tutorial introduction to Maple. Journal of Symbolic Computation, 2(2): 179–200, June 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Casperson:1996:IDS

[CFM96] David Casperson, David Ford, and John McKay. Ideal decompositions and subfields. Journal of Symbolic Computation, 21(2):133–138 (or 133– 137??), February 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Campillo:2010:SIA

[CFMMP10] Antonio Campillo, Patrick Fitzpatrick, Edgar Martínez-Moro, and Ruud Pellikaan. Special issue algebraic coding theory and applications. Journal of Symbolic Computation, 45(7):721–722, July 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cano:2023:APS

[CFRS23] José Cano, Sebastian Falkensteiner, Daniel Robertz, and J. Rafael Sendra. Algebraic and Puiseux series solutions of systems of autonomous algebraic ODEs of dimension one in several variables. Journal of Symbolic Computation, 114 (??):1–17, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000323.

Caboara:2007:SCC

[CFS07] Massimo Caboara, Sara Faridi, and Peter Selinger. Simplicial cycles and the computation of simplicial trees. Journal of Symbolic Computation, 42(1– 2):74–88, January/February 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cano:2022:ECP

[CFS22] José Cano, Sebastian Falkensteiner, and J. Rafael Sendra. Existence and convergence of Puiseux series solutions for autonomous first order differential equations. Journal of Symbolic Computation, 108(??):137–151, January/February 2022. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300535.

Cioffi:2024:EDT

[CFS24] Francesca Cioffi, Davide Franco, and Carmine Sessa. An effective decomposition theorem for Schubert varieties. Journal of Symbolic Computation, 121(??):Article 102238, March/April 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000524.

Casanellas:2021:DSP

[CFSGL21] Marta Casanellas, Jesús Fernández-Sánchez, and Marina Garrote-López. Distance to the stochastic part of phylogenetic varieties. Journal of Symbolic Computation, 104(??):653–682, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300936.

Cooperman:1997:CPR

[CFTY97] Gene Cooperman, Larry Finkelstein, Michael Tselman, and Bryant York. Constructing permutation representations for matrix groups. Journal of Symbolic Computation, 24(3–4):471–488, September/ October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Cooperman:2002:SPC

[CG02] Gene Cooperman and Victor Grinberg. Scalable parallel coset enumeration: Bulk definition and the memory wall. Journal of Symbolic Computation, 33(5):563–585, May 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cheze:2006:AEA

[CG06] Guillaume Chèze and André Galligo. From an approximate to an exact absolute polynomial factorization. Journal of Symbolic Computation, 41 (6):682–696, June 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Caminata:2023:SDL

[CG23] Alessio Caminata and Elisa Gorla. Solving degree, last fall degree, and related invariants. Journal of Symbolic Computation, 114(??): 322–335, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000487.

Char:1989:GHP

[CGG89] Bruce W. Char, Keith O. Geddes, and Gaston H. Gonnet. GCDHEU: Heuristic polynomial GCD algorithm based on integer GCD computation. Journal of Symbolic Computation, 7(1):31–48, January 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cheng:2012:RIZ

[CGG12] Jin-San Cheng, Xiao-Shan Gao, and Leilei Guo. Root isolation of zero-dimensional polynomial systems with linear univariate representation. Journal of Symbolic Computation, 47(7):843–858, July 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002033.

Camacho:2009:NGQ

[CGGO09] L. M. Camacho, J. R. Gómez, A. J. González, and B. A. Omirov. Naturally graded quasi-filiform Leibniz algebras. Journal of Symbolic Computation, 44(5):527–539, May 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Corless:2009:USE

[CGK09] Robert M. Corless, Karin Gatermann, and Ilias S. Kotsireas. Using symmetries in the eigenvalue method for polynomial systems. Journal of Symbolic Computation, 44(11):1536–1550, November 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chattopadhyay:2021:CMF

[CGK^{+}21] Arkadev Chattopadhyay, Bruno Grenet, Pascal Koiran, Natacha Portier, and Yann Strozecki. Computing the multilinear factors of lacunary polynomials without heights. Journal of Symbolic Computation, 104(??):183–206, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300286.

Campillo:2007:ECP

[CGL07] Antonio Campillo, Gert-Martin Greuel, and Christoph Lossen. Equisingular calculations for plane curve singularities. Journal of Symbolic Computation, 42(1–2):89–114, January/February 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cantone:1988:ASS

[CGO88] Domenico Cantone, Susanna Ghelfo, and Eugenio Omodeo. The automation of syllogistic. I. syllogistic normal forms. Journal of Symbolic Computation, 6(1):83–98 (or 82– 98??), August 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Conti:2023:TTA

[CGP23] Diego Conti, Alessandro Ghigi, and Roberto Pignatelli. Topological types of actions on curves. Journal of Symbolic Computation, 118(??): 17–31, September/October 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000020.

Cohen:2005:P

[CGR05] Arjeh M. Cohen, Gert-Martin Greuel, and Marie-Françoise Roy. Preface. Journal of Symbolic Computation, 39(3–4): 257–258, March/April 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Carlson:1997:CEA

[CGS97] Jon F. Carlson, Edward L.

Green, and Gerhard J. A. Schnieder. Computing Ext algebras for finite groups. Journal of Symbolic Computation, 24(3–4):317–326 (or 317– 325??), September/October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Cohen:2004:E

[CGT04] Arjeh M. Cohen, Xiao-Shan Gao, and Nobuki Takayama. Editorial. Journal of Symbolic Computation, 38(4): 1167–1168, October 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cheng:2009:CNI

[CGY09] Jin-San Cheng, Xiao-Shan Gao, and Chee-Keng Yap. Complete numerical isolation of real roots in zerodimensional triangular systems. Journal of Symbolic Computation, 44(7):768–785, July 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cox:2000:VIM

[CGZ00] David Cox, Ronald Goldman, and Ming Zhang. On the validity of implicitization by moving quadrics for rational surfaces with no base points. Journal of Symbolic Computation, 29(3):419–440, March 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0325;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0325/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0325/ref.

Coquand:1985:SBC

[CH85] Thierry Coquand and Gérard A. Huet. A selected bibliography on constructive mathematics, intuitionistic type theory and higher order deduction. Journal of Symbolic Computation, 1(3):323–328, September 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Coquand:1986:SBC

[CH86] T. Coquand and G. P. Huet. A selected bibliography on constructive mathematics, intuitionistic type theory and higher order deduction. Journal of Symbolic Computation, 3(1):??, ???? 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Champarnaud:1991:ACP

[CH91a] J. M. Champarnaud and G. Hansel. AUTOMATE, a computing package for automata and finite semigroups. Journal of Symbolic Computation, 12(2):197–220, August 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Collins:1991:PCA

[CH91b] George E. Collins and Hoon Hong. Partial cylindrical algebraic decomposition for quantifier elimination. Journal of Symbolic Computation, 12 (3):299–328, September 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Caferra:1995:GGF

[CH95] R. Caferra and M. Herment.

A generic graphic framework for combining inference tools and editing proofs and formulae. Journal of Symbolic Computation, 19(1/2/3):217– 244 (or 217–243??), January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Curtis:1996:SRE

[CH96] R. T. Curtis and Z. Hasan. Symmetric representation of the elements of the Janko group J_{1}. Journal of Symbolic Computation, 22(2):201– 214, August 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cannon:1997:SIA

[CH97a] J. J. Cannon and D. F. Holt. Special issue on applications of quantifier elimination: Foreword of the Guest Editors. Journal of Symbolic Computation, 24(3–4): 233–234, September/October 1997.

Cannon:1997:CCS

[CH97b] John J. Cannon and Derek F. Holt. Computing chief series, composition series and socles in large permutation groups. Journal of Symbolic Computation, 24(3–4):285–302 (or 285– 301??), September/October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Cannon:2003:AGC

[CH03] John J. Cannon and Derek F. Holt. Automorphism group computation and isomorphism testing in finite groups. Journal of Symbolic Computation, 35(3):241–267, March 2003. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Cannon:2004:CMS

[CH04] John Cannon and Derek F.

Holt. Computing maximal subgroups of finite groups. Journal of Symbolic Computation, 37(5):589–609, May 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cenk:2017:ACS

[CH17] Murat Cenk and M. Anwar Hasan. On the arithmetic complexity of Strassen-like matrix multiplications. Journal of Symbolic Computation, 80 (part 2)(??):484–501, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300359.

Chardin:2000:ASP

[Cha00]  Marc Chardin.  Applica-

tions of some properties of the canonical module in computational projective algebraic geometry. Journal of Symbolic Computation, 29(4–5):527–544, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0330;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0330/
pdf;   http://www.idealibrary.

                                        com/links/doi/10.1006/jsco.
                                        1999.0330/ref.

Cha:2014:CFS

[Cha14] Yongjae Cha. Closed form solutions of linear difference equations in terms of symmetric products. Journal of Symbolic Computation, 60 (??):62–77, January 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001211.

Cherry:1985:IFT

[Che85] G. W. Cherry. Integration in finite terms with special functions: the error function. Journal of Symbolic Computation, 1(3):283–302, September 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chee:1992:GQS

[Che92] Yeow Meng Chee. On graphical quintuple systems. Journal of Symbolic Computation, 13(6):677–682 (or 677–681??), June 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chenavier:2018:ROC

[Che18] Cyrille Chenavier. Reduction operators and completion of rewriting systems. Journal of Symbolic Computation, 84 (??):57–83, January/February 2018. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300238.

Cheung:2023:LSW

[Che23] Tsz Yung Cheung. Liouvillian solutions of Whittaker–Ince equation. Journal of Symbolic Computation, 115(??): 18–38, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000608.

Chistov:1996:PTC

[Chi96] Alexander L. Chistov. Polynomialtime computation of the dimension of algebraic varieties in zero-characteristic. Journal of Symbolic Computation, 22(1):1–26 (or 1–25??), July 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chionh:2001:RCC

[Chi01]  Eng-Wee Chionh.  Rect-

angular corner cutting and Dixon A-resultants. Journal of Symbolic Computation, 31(6):651–669, June 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0448;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0448/
pdf;   http://www.idealibrary.

                                        com/links/doi/10.1006/jsco.
                                        2001.0448/ref.

Chionh:2008:SI

[Chi08] Eng-Wee Chionh. 0/0 simplifies implicitization. Journal of Symbolic Computation, 43(6– 7):475–489, June/July 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2022:NOP

[CHKL22] Justin Chen, Marc Härkönen, Robert Krone, and Anton Leykin. Noetherian operators and primary decomposition. Journal of Symbolic Computation, 110(??):1–23, May/ June 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000675.

Chen:2005:AAZ

[CHM05] William Y. C. Chen, Qing-Hu Hou, and Yan-Ping Mu. Applicability of the q-analogue of Zeilberger’s algorithm. Journal of Symbolic Computation, 39(2):155–170, February 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2012:EZA

[CHM12] William Y. C. Chen, Qing-Hu Hou, and Yan-Ping Mu. The extended Zeilberger algorithm with parameters. Journal of Symbolic Computation, 47(6):643–654, June 2012.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002173.

Cannon:2005:CSB

[CHSS05] John J. Cannon, Derek F.

Holt, Michael Slattery, and Allan K. Steel. Computing subgroups of bounded index in a finite group. Journal of Symbolic Computation, 40(2): 1013–1022, August 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Churchill:1999:TGS

[Chu99] R. C. Churchill. Two generator subgroups of SL(2, C) and the hypergeometric, Riemann, and Lamé equations. Journal of Symbolic Computation, 28(4–5): 521–545, October/November 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0313/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0313/production/pdf;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0313/production/ref.

Cannon:2019:UPR

[CHU19] John J. Cannon, Derek F.

Holt, and William R. Unger. The use of permutation representations in structural computations in large finite matrix groups. Journal of Symbolic Computation, 95(??): 26–38, November/December 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300993.

Cameron:2007:GRH

[CI07] Peter Cameron and Natalia Iyudu. Graphs of relations and Hilbert series. Journal of Symbolic Computation, 42(11–12):1066–1078, November/December 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Casas:2007:PBW

[CIL07] José Manuel Casas, Manuel A. Insua, and Manuel Ladra. Poincaré–Birkhoff–Witt theorem for Leibniz n. Journal of Symbolic Computation, 42(11–12):1052–1065, November/December 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cheung:2017:STC

[CIM17] Man-Wai Cheung, Christian Ikenmeyer, and Sevak Mkrtchyan. Symmetrizing tableaux and the 5th case of the Foulkes conjecture. Journal of Symbolic Computation, 80 (part 3)(??):833–843, May/ June 2017. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300931.

Cipu:2008:GBD

[Cip08] Mihai Cipu. Gröbner bases and Diophantine analysis. Journal of Symbolic Computation, 43(10):681–687, October 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Corless:1990:SHL

[CJ90] R. M. Corless and D. J.

Jeffrey. Solution of a hydrodynamic lubrication problem with Maple. Journal of Symbolic Computation, 9(4): 503–514 (or 503–513??), April 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Corless:1997:SCO

[CJ97] R. M. Corless and D. J. Jeffrey. Scientific computing: One part of the revolution. Journal of Symbolic Computation, 23(5–6):485–496, May/ June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cheng:2015:GPB

[CJ15] Jin-San Cheng and Kai Jin. A generic position based method for real root isolation of zerodimensional polynomial systems. Journal of Symbolic Computation, 68 (part 1) (??):204–224, May/June 2015.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000856.

Chang:2022:DDP

[CJ22] Mun See Chang and Christopher Jefferson. Disjoint direct product decompositions of permutation groups. Journal of Symbolic Computation, 108 (??):1–16, January/February 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000274.

Castro-Jimenez:2009:FE

[CJGV09] Francisco J. Castro-Jimenez and Laureano Gonzalez-Vega. Foreword from the editors. Journal of Symbolic Computation, 44(5):419–420, May 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Collins:2002:IAC

[CJK02] George E. Collins, Jeremy R. Johnson, and Werner Krandick. Interval arithmetic in cylindrical algebraic decomposition. Journal of Symbolic Computation, 34(2):145–157, August 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cheng:2013:CRP

[CJL13] Jin-San Cheng, Kai Jin, and Daniel Lazard. Certified rational parametric approximation of real algebraic space curves with local generic position method. Journal of Symbolic Computation, 58 (??):18–40, November 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000953.

Celik:2021:WDI

[ÇJM^{+}21] Türkü Özlüm Çelik, Asgar Jamneshan, Guido Montúfar, Bernd Sturmfels, and Lorenzo Venturello. Wasserstein distance to independence models. Journal of Symbolic Computation, 104(??):855– 873, May/June 2021. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301152.

Corless:1997:TPC

[CJMP97] R. M. Corless, D. J. Jeffrey, M. B. Monagan, and Pratibha. Two perturbation calculations in fluid mechanics using large-expression management. Journal of Symbolic Computation, 23(4): 427–443, April 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Castro-Jimenez:2022:CPS

[CJP22] F. J. Castro-Jiménez and H. Cobo Pablos. On certain polynomial systems involving Stirling numbers of second kind. Journal of Symbolic Computation, 109(??): 325–350, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300638.

Castro-Jimenez:2001:FGE

[CJS01] F. J. Castro-Jiménez and J. Rafael Sendra. Foreword by the Guest Editors. Journal of Symbolic Computation, 32(6):573–574, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0483;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0483/
pdf.

Castro-Jimenez:2001:ECT

[CJUE01] F. J. Castro-Jiménez and J. M. Ucha-Enríquez. Explicit comparison theorems for D-modules. Journal of Symbolic Computation, 32 (6):677–685, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0489;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0489/
pdf;   http://www.idealibrary.

com/links/doi/10.1006/jsco.
2001.0489/ref.

Castro-Jimenez:2006:GBL

[CJUE06] F. J. Castro-Jiménez and J. M. Ucha-Enríquez. Gröbner bases and logarithmic Dmodules. Journal of Symbolic Computation, 41(3–4): 317–335, March/April 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cabay:1990:PSR

[CK90] S. Cabay and P. Kossowski.

Power series remainder sequences and Padé fractions over an integral domain. Journal of Symbolic Computation, 10(2):138–164 (or 139– 163??), August 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Churchill:1999:UAL

[CK99] R. C. Churchill and M. A.

Kummer. A unified approach to linear and nonlinear normal forms for Hamiltonian systems. Journal of Symbolic Computation, 27(1):49– 131, January 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chtcherba:2003:ERC

[CK03] Arthur D. Chtcherba and Deepak Kapur. Exact resultants for corner-cut unmixed multivariate polynomial systems using the Dixon formulation. Journal of Symbolic Computation, 36(3–4): 289–315, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chtcherba:2004:RUB

[CK04a] A. D. Chtcherba and D. Kapur. Resultants for unmixed bivariate polynomial systems produced using the Dixon formulation. Journal of Symbolic Computation, 38(2):915– 958, August 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chtcherba:2004:CST

[CK04b] Arthur D. Chtcherba and Deepak Kapur. Constructing Sylvester-type resultant matrices using the Dixon formulation. Journal of Symbolic Computation, 38(1):777–814, July 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2012:TOD

[CK12a] Shaoshi Chen and Manuel Kauers. Trading order for degree in creative telescoping. Journal of Symbolic Computation, 47(8):968–995, August 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000168.

Collins:2012:CTC

[CK12b] George E. Collins and Werner Krandick. On the computing time of the continued fractions method. Journal of Symbolic Computation, 47(11):1372–1412, November 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000387.

Comer:2012:BMA

[CK12c] Matthew T. Comer and Erich L. Kaltofen. On the Berlekamp/Massey algorithm and counting singular Hankel matrices over a finite field. Journal of Symbolic Computation, 47(4):480–491, April 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711100143X.

Couveignes:2012:GFT

[CK12d] Jean-Marc Couveignes and Jean-Gabriel Kammerer. The geometry of flex tangents to a cubic curve and its parameterizations. Journal of Symbolic Computation, 47(3):266–281, March 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001891.

Chen:2019:PWZ

[CK19] Shaoshi Chen and Christoph Koutschan. Proof of the Wilf–Zeilberger conjecture for mixed hypergeometric terms. Journal of Symbolic Computation, 93(??): 133–147, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300671.

Cirstea:2010:APR

[CKKM10] Horatiu Cirstea, Claude Kirchner, Radu Kopetz, and Pierre-Etienne Moreau. Antipatterns for rule-based languages. Journal of Symbolic Computation, 45(5):523–550, May 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2019:ASF

[CKLZ19] Shaoshi Chen, Manuel Kauers, Ziming Li, and Yi Zhang. Apparent singularities of D-finite systems. Journal of Symbolic Computation, 95(??):217–237, November/December 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300203.

Collart:1997:CBG

[CKM97] S. Collart, M. Kalkbrener, and D. Mall. Converting bases with the Gröbner walk.

Journal of Symbolic Computation, 24(3–4):465–470 (or 465– 469??), September/October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Chtcherba:2009:CDP

[CKM09] Arthur D. Chtcherba, Deepak Kapur, and Manfred Minimair. Cayley–Dixon projection operator for multiunivariate composed polynomials. Journal of Symbolic Computation, 44(8):972– 999, August 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Caboara:2004:ECM

[CKR04] M. Caboara, M. Kreuzer, and L. Robbiano. Efficiently computing minimal sets of critical pairs. Journal of Symbolic Computation, 38 (4):1169–1190, October 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cucker:1999:PTA

[CKS99] F. Cucker, P. Koiran, and S. Smale. A polynomial time algorithm for Diophantine equations in one variable. Journal of Symbolic Computation, 27(1):21–30, January 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2016:DOO

[CKS16] Shaoshi Chen, Manuel Kauers, and Michael F. Singer. Desingularization of Ore operators. Journal of Symbolic Computation, 74(??):617–626, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000991.

Comon:1989:EPD

[CL89] Hubert Comon and Pierre Lescanne. Equational problems and disunification. Journal of Symbolic Computation, 7(3–4):371–425 (or 371– 426??), March/April 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cattani:2000:SIS

[CL00] Eduardo Cattani and Reinhard C. Laubenbacher. Special issue on symbolic computation in algebra, analysis, and geometry foreword of the Guest Editors. Journal of Symbolic Computation, 29(4–5):483, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0371;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0371/
                                        pdf.

Cheng:2007:OSM

[CL07] Howard Cheng and George Labahn. Output-sensitive modular algorithms for polynomial matrix normal forms. Journal of Symbolic Computation, 42(7):733–750, July 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Caruso:2017:NFA

[CL17] Xavier Caruso and Jérémy Le Borgne. A new faster algorithm for factoring skew polynomials over finite fields. Journal of Symbolic Computation, 79 (part 2)(??): 411–443, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000328.

Chen:2020:APM

[CL20] Shiping Chen and Zhong Liu.

Automated proof of mixed trigonometric-polynomial inequalities. Journal of Symbolic Computation, 101(??): 318–329, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301051.

Chyzak:2023:SIS

[CL23] Frédéric Chyzak and George Labahn. Special issue on Symbolic and Algebraic Computation: ISSAC 2021. Journal of Symbolic Computation, 116(??):425–426, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712200102X.

Clausen:1991:MPS

[Cla91]  Michael Clausen.  Multi-

variate polynomials, standard tableaux and representations of symmetric groups. Journal of Symbolic Computation, 11(5–6):483–522, May/ June 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Invarianttheoretic algorithms in geometry (Minneapolis, MN, 1987).

Clausen:2021:LTF

[Cla21] Michael Clausen. Linear time Fourier transforms of S_{n−k}- invariant functions on the symmetric group S_{n}. Journal of Symbolic Computation, 98(??):319–357, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300823.

Clausen:2022:UFB

[Cla22] Michael Clausen. A unified FFT-based approach to maximum assignment problems related to transitive finite group actions. Journal of Symbolic Computation, 109(??):88–115, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000523.

Chen:2017:MCC

[CLL17] Tianran Chen, Tsung-Lin Lee, and Tien-Yien Li. Mixed cell computation in Hom4PS-

3. Journal of Symbolic Computation, 79 (part 3)(??): 516–534, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300542.

Corless:2009:P

[CLM09] Robert M. Corless, Reiner Lauterbach, and Hans-Michael Möller. Preface. Journal of Symbolic Computation, 44(11):1533–1535, November 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cioffi:2016:CDF

[CLM16] Francesca Cioffi, Paolo Lella, and Maria Grazia Marinari. A combinatorial description of finite O-sequences and aCM genera. Journal of Symbolic Computation, 73(??): 104–119, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000310.

Coquand:2010:CCP

[CLQ10] Thierry Coquand, Henri Lombardi, and Claude Quitté. Curves and coherent Prüfer rings. Journal of Symbolic Computation, 45(12):1378– 1390, December 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Comon:1991:RBT

[CLS91] H. Comon, D. Lugiez, and Ph. Schnoebelen. A rewrite-based type discipline for a subset of computer algebra. Journal of Symbolic Computation, 11 (4):349–368, April 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Carmody:1995:TCP

[CLW95] S. Carmody, M. Leeming, and R. F. C. Walters. The Todd– Coxeter procedure and left Kan extensions. Journal of Symbolic Computation, 19(5): 459–488, April 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2024:IAR

[CLX^{+}24] Rizeng Chen, Haokun Li, Bican Xia, Tianqi Zhao, and Tao Zheng. Isolating all the real roots of a mixed trigonometricpolynomial. Journal of Symbolic Computation, 121(??): Article 102250, March/April 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000640.

Clauss:1993:SAD

[CM93] Philippe Clauss and Catherine Mongenet. Synthesis aspects in the design of efficient processor arrays from affine recurrence equations. Journal of Symbolic Computation, 15(5–6):547–570 (or 547– 569??), May/June 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cesari:1996:PAP

[CM96] Giovanni Cesari and Roman Maeder. Performance analysis of the parallel Karatsuba multiplication algorithm for distributed memory architectures. Journal of Symbolic Computation, 21(4/5/6):467– 474 (or 467–473??), April, May & June 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Collart:1997:TDP

[CM97] Stéphane Collart and Daniel Mall. Toric degenerations of polynomial ideals and geometric localization of fans. Journal of Symbolic Computation, 24(3–4):443–464, September/ October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Clausen:2004:GFF

[CM04] M. Clausen and M. Müller. Generating fast Fourier transforms of solvable groups. Journal of Symbolic Computation, 37(2):137–156, February 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Costermans:2009:NAM

[CM09] Christian Costermans and Hoang Ngoc Minh. Noncommutative algebra, multiple harmonic sums and applications in discrete probability. Journal of Symbolic Computation, 44(7):801–817, July 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2010:ASL

[CM10] Liang Chen and Michael Monagan. Algorithms for solving linear systems over cyclotomic fields. Journal of Symbolic Computation, 45(9):902– 917, September 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2012:ACT

[CM12] Changbo Chen and Marc Moreno Maza. Algorithms for computing triangular decomposition of polynomial systems.

Journal of Symbolic Computation, 47(6):610–642, June 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002161.

Chen:2016:QEC

[CM16] Changbo Chen and Marc Moreno Maza. Quantifier elimination by cylindrical algebraic decomposition based on regular chains. Journal of Symbolic Computation, 75(??): 74–93, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001078.

Ceria:2017:BWT

[CM17a] Michela Ceria and Teo Mora.

Buchberger–Weispfenning theory for effective associative rings. Journal of Symbolic Computation, 83(??):112–146, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301274.

Chen:2017:PDC

[CM17b] Tianran Chen and Dhagash Mehta. Parallel degree computation for binomial systems. Journal of Symbolic Computation, 79 (part 3)(??): 535–558, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300554.

Clarke:2021:SMT

[CM21] Oliver Clarke and Fatemeh Mohammadi. Standard monomial theory and toric degenerations of Schubert varieties from matching field tableaux. Journal of Symbolic Computation, 104(??):683–723, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300973.

Cerlienco:1987:CMP

[CMP87] L. Cerlienco, M. Mignotte, and F. Piras. Computing the measure of a polynomial. Journal of Symbolic Computation, 4(1):21–33, August 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ceria:2015:TOF

[CMR15] Michela Ceria, Teo Mora, and Margherita Roggero. Term-ordering free involutive bases. Journal of Symbolic Computation, 68 (part 2) (??):87–108, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711400073X.

Ceria:2019:GFN

[CMR19] Michela Ceria, Teo Mora, and Margherita Roggero. A general framework for Noetherian well ordered polynomial reductions. Journal of Symbolic Computation, 95(??):100–133, November/December 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300124.

Cadavid:2013:LQB

[CMV13] C. Cadavid, S. Molina, and J. D. Vélez. Limits of quotients of bivariate real analytic functions. Journal of Symbolic Computation, 50 (??):197–207, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001204.

Creel:2007:SCF

[CN07] Conrad Creel and Sam Nelson. Symbolic computation with finite biquandles. Journal of Symbolic Computation, 42(10):992–1000, October 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chien:2019:VLF

[CN19] Julien Chien and Sam Nelson. Virtual links with finite medial bikei. Journal of Symbolic Computation, 92 (??):211–221, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300452.

Cremanns:1994:FDT

[CO94] Robert Cremanns and Friedrich Otto. Finite derivation type implies the homological finiteness condition FP_{3}. Journal of Symbolic Computation, 18 (2):91–112, August 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cremanns:1996:GPH

[CO96] Robert Cremanns and Friedrich Otto. For groups the property of having finite derivation type is equivalent to the homological finiteness condition FP_{3}. Journal of Symbolic Computation, 22(2):155– 178 (or 155–177??), August 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Caprotti:2001:FEP

[CO01] Olga Caprotti and Martijn Oostdijk. Formal and efficient primality proofs by use of computer algebra oracles. Journal of Symbolic Computation, 32(1–2): 55–70, July/August 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.

com/links/doi/10.1006/jsco.
2001.0457;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0457/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0457/ref.

Collins:2001:PMR

[Col01] George E. Collins. Polynomial minimum root separation. Journal of Symbolic Computation, 32(5): 467–473, November 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0481;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0481/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0481/ref.

Collins:2002:FEA

[Col02] George E. Collins. A fast Euclidean algorithm for Gaussian integers. Journal of Symbolic Computation, 33(4): 385–392, April 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Collins:2004:SFC

[Col04] George E. Collins. Singlefactor coefficient bounds. Journal of Symbolic Computation, 38(6):1507–1521, December 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Colton:2005:ACM

[Col05] Simon Colton. Automated conjecture making in number theory using HR, Otter and Maple. Journal of Symbolic Computation, 39(5):593–615, May 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Collins:2015:KPL

[Col15] George E. Collins. Krandick’s proof of Lagrange’s real root bound claim. Journal of Symbolic Computation, 70(??): 106–111, September/October 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114001060.

Collins:2016:CFR

[Col16] George E. Collins. Continued fraction real root isolation using the Hong root bound. Journal of Symbolic Computation, 72(??):21– 54, January/February 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114001308.

Collins:2017:MCT

[Col17] George E. Collins. On the maximum computing time of the bisection method for real root isolation. Journal of Symbolic Computation, 79 (part 2)(??):444– 456, March/April 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711600033X.

Comon:1998:CRSa

[Com98a] Hubert Comon. Completion of rewrite systems with membership constraints. Part I: Deduction rules. Journal of Symbolic Computation, 25(4): 397–420 (or 397–419??), April 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Comon:1998:CRSb

[Com98b] Hubert Comon. Completion of rewrite systems with membership constraints. Part II: Constraint solving. Journal of Symbolic Computation, 25(4): 421–454 (or 421–453??), April 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Compoint:1998:DEA

[Com98c] Elie Compoint. Differential equations and algebraic relations. Journal of Symbolic Computation, 25(6):705–725, June 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Conlon:1990:CCG

[Con90a]  S. B. Conlon.  Calculat-

ing characters of p-groups. Journal of Symbolic Computation, 9(5–6):535–550, May/ June 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Conlon:1990:CMP

[Con90b]  S. B. Conlon.  Comput-

ing modular and projective character degrees of soluble groups. Journal of Symbolic Computation, 9(5–6): 551–570, May/June 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Contejean:1993:SPM

[Con93] Evelyne Contejean. Solving ∗- problems modulo distributivity by a reduction to AC1unification. Journal of Symbolic Computation, 16(5):493– 521, November 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cools:2009:RBW

[Coo09] Filip Cools. On the relation between weighted trees and tropical Grassmannians. Journal of Symbolic Computation, 44(8):1079–1086, August 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Couveignes:2000:BHS

[Cou00] Jean-Marc Couveignes. Boundary of Hurwitz spaces and explicit patching. Journal of Symbolic Computation, 30(6):739–759, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0381;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0381/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0381/ref.

Coutinho:2018:BDS

[Cou18] S. C. Coutinho. Bounding the degree of solutions of differential equations. Journal of Symbolic Computation, 89(??): 9–25, November/December 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117301025.

Coutinho:2022:CFD

[Cou22] S. C. Coutinho. On the classification of foliations of degree three with one singularity. Journal of Symbolic Computation, 112(??): 62–78, September/October 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000013.

Cowell:1992:AOS

[Cow92] R. G. Cowell. Application of ordered standard bases to catastrophe theory. Journal of Symbolic Computation, 13(1): 101–115, January 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Coxon:2019:FSE

[Cox19] Nicholas Coxon. Fast systematic encoding of multiplicity codes. Journal of Symbolic Computation, 94(??): 234–254, September/October 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300968.

Coxon:2021:FHI

[Cox21a] Nicholas Coxon. Fast Hermite interpolation and evaluation over finite fields of characteristic two. Journal of Symbolic Computation, 98(??):270–283, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300793.

Coxon:2021:FTF

[Cox21b] Nicholas Coxon. Fast transforms over finite fields of characteristic two. Journal of Symbolic Computation, 104 (??):824–854, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301127.

Coxon:2022:PTF

[Cox22] Nicholas Coxon. An in-place truncated Fourier transform. Journal of Symbolic Computation, 110(??):66–80, May/ June 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000705.

Chadha:1993:MDL

[CP93] Ritu Chadha and David A.

Plaisted. On the mechanical derivation of loop invariants. Journal of Symbolic Computation, 15(5–6):705–744, May/ June 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cannon:1997:UMC

[CP97] J. Cannon and C. Playoust.

Using the Magma computer algebra system in abstract algebra courses. Journal of Symbolic Computation, 23(5– 6):459–484, May/June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Caferra:2000:CED

[CP00] Ricardo Caferra and Nicolas Peltier. Combining enumeration and deductive techniques in order to increase the class of constructible infinite models.

Journal of Symbolic Computation, 29(2):177–211, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0360;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0360/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0360/ref.

Coles:2010:DRB

[CP10] Drue Coles and Emma Previato. Decoding by rank-2 bundles over plane quartics. Journal of Symbolic Computation, 45(7):757–772, July 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Castle:2009:QPT

[CPR09] Mari Castle, Victoria Powers, and Bruce Reznick. A quantitative Pólya’s Theorem with zeros. Journal of Symbolic Computation, 44(9): 1285–1290, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Castle:2011:PTZ

[CPR11] Mari Castle, Victoria Powers, and Bruce Reznick. Pólya’s Theorem with zeros. Journal of Symbolic Computation, 46(9):1039–1048, September 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000770.

Cluzeau:2012:SRL

[CQ12] Thomas Cluzeau and Alban Quadrat. Serre’s reduction of linear partial differential systems with holonomic adjoints. Journal of Symbolic Computation, 47(10):1192–1213, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002343.

Coste:1988:TLC

[CR88] M. Coste and M. F. Roy.

Thom’s lemma, the coding of real algebraic numbers and the computation of the topology of semi-algebraic sets. Journal of Symbolic Computation, 5(1–2):121–130 (or 121– 129??), February/April 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cucker:1990:TRP

[CR90] F. Cucker and M. F. Roy. A theorem on random polynomials and some consequences in average complexity. Journal of Symbolic Computation, 10(5):405–410 (or 405– 409??), November 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Caboara:1998:ACA

[CR98] Massimo Caboara and Eva Riccomagno. An algebraic computational approach to the identifiability of Fourier models. Journal of Symbolic Computation, 26(2):245– 260, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cioffi:2011:FFS

[CR11] Francesca Cioffi and Margherita Roggero. Flat families by strongly stable ideals and a generalization of Gröbner bases. Journal of Symbolic Computation, 46(9): 1070–1084, September 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111000800.

Cid-Ruiz:2019:BDM

[CR19] Yairon Cid-Ruiz. Bounding the degrees of a minimal μ-basis for a rational surface parametrization. Journal of Symbolic Computation, 95(??):134–150, November/December 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300148.

Crapo:1991:ITM

[Cra91] Henry Crapo. Invariant theoretic methods in scene analysis and structural mechanics.

Journal of Symbolic Computation, 11(5–6):523–548, May/ June 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Invarianttheoretic algorithms in geometry (Minneapolis, MN, 1987).

Chazarain:1991:MVL

[CRAB91] J. Chazarain, A. Riscos, J. A.

Alonso, and E. Briales. Multivalued logic and Gröbner bases with applications to modal logic. Journal of Symbolic Computation, 11(3): 181–194, March 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cremona:2001:CID

[Cre01] J. E. Cremona. Classical invariants and 2-descent on elliptic curves. Journal of Symbolic Computation, 31(1– 2):71–87, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.1004;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1998.1004/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.1004/ref.

Cifuentes:2022:VCV

[CRSW22] Diego Cifuentes, Kristian Ranestad, Bernd Sturmfels, and Madeleine Weinstein. Voronoi cells of varieties. Journal of Symbolic Computation, 109(??): 351–366, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771712030064X.

Cole:1989:VPP

[CS89] Richard Cole and Micha Sharir. Visibility problems for polyhedral terrains. Journal of Symbolic Computation, 7 (1):11–30, January 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chazelle:1990:AGP

[CS90] Bernard Chazelle and Micha Sharir. An algorithm for generalized point location and its applications. Journal of Symbolic Computation, 10(3–4):281–310 (or 281– 309??), September/October 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chyzak:1998:NCE

[CS98] Frédéric Chyzak and Bruno Salvy. Non-commutative elimination in Ore algebras proves multivariate identities. Journal of Symbolic Computation, 26(2):187–227, 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Compoint:1999:CGG

[CS99] Elie Compoint and Michael F.

Singer. Computing Galois groups of completely reducible differential equations. Journal of Symbolic Computation, 28(4–5): 473–494, October/November 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0311/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0311/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0311/production/ref.

Caprotti:2005:IAR

[CS05a] Olga Caprotti and Volker Sorge. Integration of automated reasoning and computer algebra systems. Journal of Symbolic Computation, 39(5):501–502, May 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2005:GPR

[CS05b] William Y. C. Chen and Husam L. Saad. On the Gosper–Petkovšek representation of rational functions. Journal of Symbolic Computation, 40(2):955–963, August 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Conca:2005:GII

[CS05c] Aldo Conca and Jessica Sidman. Generic initial ideals of points and curves. Journal of Symbolic Computation, 40(3): 1023–1038, September 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Coutinho:2006:ASH

[CS06] S. C. Coutinho and L. Menasché Schechter. Algebraic solutions of holomorphic foliations: An algorithmic approach. Journal of Symbolic Computation, 41 (5):603–618, May 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Clay:2009:COF

[CS09] Adam Clay and Lawrence H. Smith. Corrigendum to: “On ordering free groups” [J. Symbolic Comput. 40 (2005) 1285–1290]. Journal of Symbolic Computation, 44(10): 1529–1532, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [Smi05].

Cheng:2016:NTR

[CS16] Charles Ching-An Cheng and Takis Sakkalis. On new types of rational rotationminimizing frame space curves. Journal of Symbolic Computation, 74(??):400–407, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000784.

Coons:2021:QIM

[CS21] Jane Ivy Coons and Seth Sullivant. Quasi-independence models with rational maximum likelihood estimator. Journal of Symbolic Computation, 104(??):917–941, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301164.

Canon:2022:TAC

[CS22] Mario Morán Cañón and Julien Sebag. Two algorithms for computing the general component of jet scheme and applications. Journal of Symbolic Computation, 113(??): 74–96, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000098.

Caleiro:1996:DLG

[CSS96] C. Caleiro, G. Saake, and A. Sernadas. Deriving liveness goals from temporal logic specifications. Journal of Symbolic Computation, 22(5– 6):521–554 (or 521–553??), November/December 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cormier:2002:LDO

[CSTU02] Olivier Cormier, Michael F. Singer, Barry M. Trager, and Felix Ulmer. Linear differential operators for polynomial equations. Journal of Symbolic Computation, 34(5):355– 398, November 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cheb-Terrab:1999:IFS

[CTR99] E. S. Cheb-Terrab and A. D. Roche. Integrating factors for second-order ODEs. Journal of Symbolic Computation, 27 (5):501–519, May 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Charalambous:2016:BFI

[CTV16] Hara Charalambous, Apostolos Thoma, and Marius Vladoiu. Binomial fibers and indispensable binomials. Journal of Symbolic Computation, 74(??):578–591, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000966.

Cueto:2010:ICB

[CTY10] María Angélica Cueto, Enrique A. Tobis, and Josephine Yu. An implicitization challenge for binary factor analysis. Journal of Symbolic Computation, 45(12):1296– 1315, December 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cuyt:1997:FPV

[Cuy97] Annie Cuyt. Floating-point versus symbolic computations in the QD-algorithm. Journal of Symbolic Computation, 24(6):695–704 (or 695– 703??), December 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Carminati:2000:SCD

[CV00] John Carminati and Khai Vu. Symbolic computation and differential equations: Lie symmetries. Journal of Symbolic Computation, 29(1):95–116, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0299;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0299/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0299/ref.

Castryck:2011:TFE

[CV11] Wouter Castryck and Frederik Vercauteren. Toric forms of elliptic curves and their arithmetic. Journal of Symbolic Computation, 46 (8):943–966, August 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111000368.

Cluzeau:2004:MAC

[CvH04] Thomas Cluzeau and Mark van Hoeij. A modular algorithm for computing the exponential solutions of a linear differential operator. Journal of Symbolic Computation, 38(3):1043–1076, September 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2018:RBC

[CvHKK18] Shaoshi Chen, Mark van Hoeij, Manuel Kauers, and Christoph Koutschan. Reductionbased creative telescoping for fuchsian D-finite functions. Journal of Symbolic Computation, 85(??): 108–127, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711730069X.

Chattopadhyay:2017:CCP

[CVY17] Amit Chattopadhyay, Gert Vegter, and Chee K. Yap. Certified computation of planar Morse–Smale complexes. Journal of Symbolic Computation, 78(??):3–40, January/February 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300062.

Chen:2021:CUG

[CVZ21] Justin Chen, Sameera Vemulapalli, and Leon Zhang. Computing unit groups of curves. Journal of Symbolic Computation, 104(??):236–255, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300316.

Coppersmith:1990:MMA

[CW90] Don Coppersmith and Shmuel Winograd. Matrix multiplication via arithmetic progressions. Journal of Symbolic Computation, 9(3):251– 280, March 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2003:RBR

[CW03] Falai Chen and Wenping Wang. Revisiting the μ-basis of a rational ruled surface. Journal of Symbolic Computation, 36(5):699–716, November 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chen:2008:CSP

[CWL08] Falai Chen, Wenping Wang, and Yang Liu. Computing singular points of plane rational curves. Journal of Symbolic Computation, 43(2):92– 117, February 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cheng:2023:CNR

[CWZ23] Jin-San Cheng, Junyi Wen, and Bingwei Zhang. Certified numerical real root isolation for bivariate nonlinear systems. Journal of Symbolic Computation, 114(??): 149–171, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000256.

Chen:2009:WMI

[CX09] William Y. C. Chen and Ernest X. W. Xia. The q-WZ method for infinite series. Journal of Symbolic Computation, 44(8):960–971, August 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Caferra:1992:MSS

[CZ92] Ricardo Caferra and Nicolas Zabel. A method for simultaneous search for refutations and models by equational constraint solving. Journal of Symbolic Computation, 13(6): 613–642 (or 613–641??), June 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Czapor:1989:SAE

[Cza89] S. R. Czapor. Solving algebraic equations: Combining Buchberger’s algorithm with multivariate factorization. Journal of Symbolic Computation, 7(1):49–54 (or 49–53??), January 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chionh:2002:FCB

[CZG02] Eng-Wee Chionh, Ming Zhang, and Ronald N. Goldman. Fast computation of the Bezout and Dixon resultant matrices. Journal of Symbolic Computation, 33(1):13–29, January 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0462;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0462/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0462/ref.

Czichowski:1995:NGB

[Czi95] Günter Czichowski. A note on Gröbner bases and integration of rational functions. Journal of Symbolic Computation, 20(2):163–168 (or 163– 167??), August 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DAndrea:2001:RMS

[D’A01]  Carlos D’Andrea.  Resul-

tants and moving surfaces. Journal of Symbolic Computation, 31(5):585–602, May 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0443;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0443/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0443/ref.

Dabbaghian-Abdoly:2005:ACR

[DA05] Vahid Dabbaghian-Abdoly. An algorithm for constructing representations of finite groups. Journal of Symbolic Computation, 39(6):671–688, June 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Daberkow:1997:CS

[Dab97a]  M. Daberkow.  Comput-

ing with subfields. Journal of Symbolic Computation, 24(3–4):371–384, September/ October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Daberkow:1997:CSC

[Dab97b]  M. Daberkow.  Comput-

ing with subfields. Journal of Symbolic Computation, 24 (3–4):371–384, 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Daberkow:2001:CKE

[Dab01] Mario Daberkow. On computations in Kummer extensions. Journal of Symbolic Computation, 31(1–2): 113–131, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.1013;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.1013/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.1013/ref.

Dahan:2022:LGB

[Dah22] Xavier Dahan. Lexicographic Gröbner bases of bivariate polynomials modulo a univariate one. Journal of Symbolic Computation, 110 (??):24–65, May/June 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000699.

deAlba:2017:BNB

[dAM17] Hernán de Alba and Marcel Morales. Betti numbers of binomial ideals. Journal of Symbolic Computation, 80 (part 2)(??):387–402, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300311.

Davenport:1988:CAA

[Dav88] J. H. Davenport. Computer algebra applied to itself. Journal of Symbolic Computation, 6(1):127–132, August 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Davis:1994:ICF

[Dav94] Richard A. Davis. Idempotent computation over finite fields. Journal of Symbolic Computation, 17(3):237–258, March 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Davenport:2002:ECA

[Dav02] James H. Davenport. Equality in computer algebra and beyond. Journal of Symbolic Computation, 34(4):259– 270, October 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Duan:2004:CME

[DaZZ04] Haibao Duan, Xu an Zhao, and Xuezhi Zhao. The Cartan matrix and enumerative calculus. Journal of Symbolic Computation, 38(3): 1119–1144, September 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

deSchreye:1989:TLP

[dB89] Danny de Schreye and Maurice Bruynooghe. On the transformation of logic programs with instantiation based computation rules. Journal of Symbolic Computation, 7(2):125–154, February 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dutton:1989:IGI

[DBG89] R. D. Dutton, R. C. Brigham, and F. Gomez. INGRID: a graph invariant manipulator. Journal of Symbolic Computation, 7(2):163–178 (or 163– 177??), February 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Saenz-de-Cabezon:2010:CSM

[dC10] Eduardo Sáenz de Cabezón.

Computing the support of monomial iterated mapping cones. Journal of Symbolic Computation, 45(10): 953–964, October 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dauchet:1995:ARP

[DCC95] Max Dauchet, Anne-Cécile Caron, and Jean-Luc Coquidé. Automata for reduction properties solving. Journal of Symbolic Computation, 20(2):215–233, August 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

delCampo:2017:CPM

[dCR17] Abraham Martín del Campo and Jose Israel Rodriguez. Critical points via monodromy and local methods. Journal of Symbolic Computation, 79 (part 3)(??): 559–574, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300566.

Saenz-de-Cabezon:2009:BNM

[dCW09] Eduardo Sáenz de Cabezón and Henry P. Wynn. Betti numbers and minimal free resolutions for multi-state system reliability bounds. Journal of Symbolic Computation, 44(9): 1311–1325, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Deprit:1990:PPS

[DD90] T. Deprit and E. Deprit. Processing Poisson series in parallel. Journal of Symbolic Computation, 10(2):179– 202 (or 179–201??), August 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Delest:1995:PPC

[DDD95] M. Delest, J. P. Dubernard, and I. Dutour. Parallelogram polyominoes and corners. Journal of Symbolic Computation, 20(5–6):503– 516 (or 503–515??), November/December 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Dietrich:2021:CRW

[DdG21] Heiko Dietrich and Willem A.

de Graaf. Computing the real Weyl group. Journal of Symbolic Computation, 104 (??):1–14, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712030016X.

Dickenstein:2013:FE

[DDHS13] Alicia Dickenstein, Sandra Di Rocco, Evelyne Hubert, and Josef Schicho. Foreword from the Editors. Journal of Symbolic Computation, 51(??):1–2, April 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001113.

Dongre:2023:CRP

[DDL^{+}23] Pratik Dongre, Benjamin Drabkin, Josiah Lim, Ethan Partida, Ethan Roy, Dylan Ruff, Alexandra Seceleanu, and Tingting Tang. Computing rational powers of monomial ideals. Journal of Symbolic Computation, 116 (??):39–57, May/June 2023.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000852.

Dickenstein:2015:SIE

[DDM15] Alicia Dickenstein, Jan Draisma, and Bernard Mourrain. Special issue on effective methods in algebraic computation. Journal of Symbolic Computation, 68 (part 2) (??):1–3, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000698.

Dumas:2011:CEC

[DDR11] Jean-Guillaume Dumas, Dominique Duval, and Jean-Claude Reynaud. Cartesian effect categories are Freyd-categories. Journal of Symbolic Computation, 46(3):272–293, March 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001677.

deJong:1998:ACI

[de 98] Theo de Jong. An algorithm for computing the integral closure. Journal of Symbolic Computation, 26(3):273– 277, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DAndrea:2002:HSR

[DE02] Carlos D’Andrea and Ioannis Z. Emiris. Hybrid sparse resultant matrices for bivariate polynomials. Journal of Symbolic Computation, 33(5): 587–608, May 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dickenstein:2003:MRF

[DE03] Alicia Dickenstein and Ioannis Z. Emiris. Multihomogeneous resultant formulae by means of complexes. Journal of Symbolic Computation, 36(3–4):317–342, September/ October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Diaconis:2006:MBN

[DE06] Persi Diaconis and Nicholas Eriksson. Markov bases for noncommutative Fourier analysis of ranked data. Journal of Symbolic Computation, 41(2): 182–195, February 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dedieu:1997:ESN

[Ded97] Jean-Pierre Dedieu. Estimations for the separation number of a polynomial system. Journal of Symbolic Computation, 24(6):683–693, December 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Davenport:2021:SCS

[DEG^{+}21] James H. Davenport, Matthew England, Alberto Griggio, Thomas Sturm, and Cesare Tinelli. Symbolic computation and satisfiability checking. Journal of Symbolic Computation, 100(??):1–10, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300835.

Dehon:1994:CGC

[Deh94] Michel Dehon. Classifying geometries with Cayley. Journal of Symbolic Computation, 17 (3):259–276, March 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

delCorso:1995:FPI

[del95] Ilaria del Corso. Factorization of prime ideal extensions in Dedekind domains. Journal of Symbolic Computation, 19(5): 435–440 (or 435–439??), April 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dietrich:2022:GWO

[DEP22] Heiko Dietrich, Bettina Eick, and Xueyu Pan. Groups whose orders factorise into at most four primes. Journal of Symbolic Computation, 108 (??):23–40, January/February 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000298.

DiRocco:2011:CNS

[DEPS11] Sandra Di Rocco, David Eklund, Chris Peterson, and Andrew J. Sommese. Chern numbers of smooth varieties via homotopy continuation and intersection theory. Journal of Symbolic Computation, 46(1):23–33, January 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001148.

Dershowitz:1987:CTR

[Der87a] Nachum Dershowitz. Corrigendum: “Termination of rewriting”. Journal of Symbolic Computation, 4(3):409– 410, 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [Der87b].

Dershowitz:1987:TR

[Der87b] Nachum Dershowitz. Termination of rewriting. Journal of Symbolic Computation, 3 (1–2):69–115, February/April 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Rewriting techniques and applications (Dijon, 1985). See corrigendum [Der87a].

Derksen:2013:GIP

[Der13] Harm Derksen. The graph isomorphism problem and approximate categories. Journal of Symbolic Computation, 59(??):81–112, December 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711300093X.

Dumitriu:2007:MMO

[DES07] Ioana Dumitriu, Alan Edelman, and Gene Shuman. MOPS: Multivariate orthogonal polynomials (symbolically). Journal of Symbolic Computation, 42(6):587–620, June 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Diochnos:2009:APC

[DET09] Dimitrios I. Diochnos, Ioannis Z. Emiris, and Elias P. Tsigaridas. On the asymptotic and practical complexity of solving bivariate systems over the reals. Journal of Symbolic Computation, 44 (7):818–835, July 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Deutsch:1993:IAH

[Deu93] Jesse Ira Deutsch. Identities arising from Hecke transformations of modular forms over Q(\sqrt{2}) and Q(\sqrt{3}). Journal of Symbolic Computation, 15(3):315–324 (or 315– 323??), March 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dey:2021:DDR

[Dey21] Papri Dey. Definite determinantal representations via orthostochastic matrices. Journal of Symbolic Computation, 104(??):15–37, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300201.

Damgaard:2005:EAG

[DF05] Ivan Bjerre Damgård and Gudmund Skovbjerg Frandsen. Efficient algorithms for the gcd and cubic residuosity in the ring of Eisenstein integers. Journal of Symbolic Computation, 39(6):643–652, June 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Detinko:2008:ACN

[DF08] A. S. Detinko and D. L. Flannery. Algorithms for computing with nilpotent matrix groups over infinite domains. Journal of Symbolic Computation, 43(1):8–26, January 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Detinko:2009:DFM

[DF09] A. S. Detinko and D. L. Flannery. On deciding finiteness of matrix groups. Journal of Symbolic Computation, 44(8): 1037–1043, August 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dietrich:2013:CRL

[DFdG13] Heiko Dietrich, Paolo Faccin, and Willem A. de Graaf. Computing with real Lie algebras: Real forms, Cartan decompositions, and Cartan subalgebras. Journal of Symbolic Computation, 56 (??):27–45, September 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000709.

Detinko:2015:IAS

[DFdG15] A. S. Detinko, D. L. Flannery, and W. A. de Graaf. Integrality and arithmeticity of solvable linear groups. Journal of Symbolic Computation, 68 (part 1)(??):138–145, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000613.

Daberkow:1997:KVC

[DFK^{+}97a] M. Daberkow, C. Fieker, J. Klüners, M. Pohst, K. Roegner, M. Schörnig, and K. Wildanger. KANT V4. Journal of Symbolic Computation, 24 (3–4):267–283, 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Daberkow:1997:KV

[DFK^{+}97b] M. Daberkow, C. Fieker, J. Klüners, M. Pohst, K. Roegner, M. Schörnig, and K. Wildanger. KANT V4. Journal of Symbolic Computation, 24(3–4):267–284, September/ October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Detinko:2013:RFM

[DFO13] A. S. Detinko, D. L. Flannery, and E. A. O’Brien. Recognizing finite matrix groups over infinite fields. Journal of Symbolic Computation, 50(??):100–109, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001071.

Dumas:2011:SMR

[DFS11] Jean-Guillaume Dumas, Laurent Fousse, and Bruno Salvy. Simultaneous modular reduction and Kronecker substitution for small finite fields. Journal of Symbolic Computation, 46(7):823–840, July 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001458.

deGraaf:2001:CQE

[dG01] W. A. de Graaf. Computing with quantized enveloping algebras: PBW-Type bases, highest-weight modules and R-matrices. Journal of Symbolic Computation, 32(5):475–490, November 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0479;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0479/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0479/ref.

deGraaf:2009:CAG

[dG09] Willem A. de Graaf. Constructing algebraic groups from their Lie algebras. Journal of Symbolic Computation, 44(9):1223–1233, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

deGraaf:2011:CRN

[dG11] Willem A. de Graaf. Computing representatives of nilpotent orbits of θ-groups. Journal of Symbolic Computation, 46(4):438–458, April 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001914.

Dehornoy:2014:AGC

[DG14] Patrick Dehornoy and Volker Gebhardt. Algorithms for Garside calculus. Journal of Symbolic Computation, 63(??):68–116, May 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001648.

Duarte:2020:EDP

[DG20] Eliana Duarte and Christiane Görgen. Equations defining probability tree models. Journal of Symbolic Computation, 99(??):127–146, July/August 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300379.

Dressler:2024:AOS

[DGLM^{+}24] Mareike Dressler, Marina Garrote-López, Guido Montúfar, Johannes Müller, and Kemal Rose. Algebraic optimization of sequential decision problems. Journal of Symbolic Computation, 121(??): Article 102241, March/April 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712300055X.

deGraaf:2002:CFR

[dGN02] Willem A. de Graaf and Werner Nickel. Constructing faithful representations of finitely-generated torsionfree nilpotent groups. Journal of Symbolic Computation, 33(1):31–41, January 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0497;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0497/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0497/ref.

DAndrea:2010:EMA

[DGPP10] Carlos D’Andrea, Marc Giusti, Luis M. Pardo, and Ragni Piene. Effective methods in algebraic geometry 2009: Barcelona. Guest editors’ foreword. Journal of Symbolic Computation, 45(12):1251– 1253, December 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

deGraaf:2009:PDP

[dGPS09] Willem A. de Graaf, Jana Pílniková, and Josef Schicho. Parametrizing Del Pezzo surfaces of degree 8 using Lie algebras. Journal of Symbolic Computation, 44(1):1– 14, January 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dickenstein:2007:FE

[DGR07] Alicia Dickenstein, Patrizia Gianni, and Tomás Recio. Foreword from the Editors. Journal of Symbolic Computation, 42(1–2):1–3, January/February 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dalmas:1996:DCE

[DGS96] S. Dalmas, M. Gaetano, and A. Sausse. A distributed and cooperative environment for computer algebra. Journal of Symbolic Computation, 21(4/5/6):427–440 (or 427– 439??), April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Damiano:2010:CRQ

[DGS10] Alberto Damiano, Graziano Gentili, and Daniele Struppa. Computations in the ring of quaternionic polynomials. Journal of Symbolic Computation, 45(1):38–45, January 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Du:2019:RCG

[DGW19] Juan Du, Ron Goldman, and Xuhui Wang. Rational curves over generalized complex numbers. Journal of Symbolic Computation, 93 (??):56–84, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300397.

Davenport:1988:RQE

[DH88] James H. Davenport and Joos Heintz. Real quantifier elimination is doubly exponential. Journal of Symbolic Computation, 5(1–2):29–36 (or 29– 35??), February/April 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Durand:2000:SFS

[DH00] Cassiano Durand and Christoph M. Hoffmann. A systematic framework for solving geometric constraints analytically. Journal of Symbolic Computation, 30(5):493–519, October 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0392;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0392/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0392/ref.

Daniel:2007:CFS

[DH07] Jennifer R. Daniel and Aloysius G. Helminck. Computing the fine structure of real reductive symmetric spaces. Journal of Symbolic Computation, 42(5):497–510, May 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Daleo:2016:NDA

[DH16] Noah S. Daleo and Jonathan D.

Hauenstein. Numerically deciding the arithmetically Cohen–Macaulayness of a projective scheme. Journal of Symbolic Computation, 72(??):128–146, January/February 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000115.

Darmian:2017:PFA

[DH17] Mahdi Dehghani Darmian and Amir Hashemi. Parametric FGLM algorithm. Journal of Symbolic Computation, 82(??):38–56, September/October 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301705.

DeLoera:2004:SRF

[DHH^{+}04] J. A. De Loera, D. Haws, R. Hemmecke, P. Huggins, B. Sturmfels, and R. Yoshida. Short rational functions for toric algebra and applications. Journal of Symbolic Computation, 38(2):959–973, August 2004. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Diaz:1995:PSD

[DHK^{+}95] A. Diaz, M. Hitz, E. Kaltofen, A. Lobo, and T. Valente. Process scheduling in DSC and the large sparse linear systems challenge. Journal of Symbolic Computation, 19(1/2/3):269–282, January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DAndrea:2007:EPS

[DHKS07] Carlos D’Andrea, Hoon Hong, Teresa Krick, and Agnes Szanto. An elementary proof of Sylvester’s double sums for subresultants. Journal of Symbolic Computation, 42(3): 290–297, March 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DAndrea:2009:SDS

[DHKS09] Carlos D’Andrea, Hoon Hong, Teresa Krick, and Agnes Szanto. Sylvester’s double sums: The general case. Journal of Symbolic Computation, 44(9):1164–1175, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Darmian:2011:ENA

[DHM11] Mahdi Dehghani Darmian, Amir Hashemi, and Antonio Montes. Erratum to “A new algorithm for discussing Gröbner bases with parameters” [J. Symbolic Comput. 33 (1–2) (2002) 183–208]. Journal of Symbolic Computation, 46(10):1187–1188, October 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000629.

See [Mon02b].

Decker:1998:GNN

[DHS98] Wolfram Decker, Agnes Eileen Heydtmann, and Frank-Olaf Schreyer. Generating a Noetherian normalization of the invariant ring of a finite group. Journal of Symbolic Computation, 25(6):727–731, June 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Duff:2022:CPS

[DHS22] Timothy Duff, Nickolas Hein, and Frank Sottile. Certification for polynomial systems via square subsystems. Journal of Symbolic Computation, 109(??): 367–385, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300651.

DeLoera:2004:ELP

[DHTY04] Jesús A. De Loera, Raymond Hemmecke, Jeremiah Tauzer, and Ruriko Yoshida. Effective lattice point counting in rational convex polytopes. Journal of Symbolic Computation, 38(4):1273–1302, October 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dickerson:1992:IAP

[Dic92] Matthew Dickerson. The inverse of an automorphism in polynomial time. Journal of Symbolic Computation, 13(2): 209–220, February 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dressler:2018:ACP

[DIdW18] Mareike Dressler, Sadik Iliman, and Timo de Wolff. An approach to constrained polynomial optimization via nonnegative circuit polynomials and geometric programming. Journal of Symbolic Computation, 91(??):149–172, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711830083X.

Dietrich:1992:EAC

[Die92] Volker Dietrich. ELISE, an algorithm to compute asymptotic representations for solutions of linear differential equations, realized with the computer algebra system MAPLE. Journal of Symbolic Computation, 14 (1):85–92, July 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DiPasquale:2016:APS

[DiP16] Michael DiPasquale. Associated primes of spline complexes. Journal of Symbolic Computation, 76(??): 158–199, September/October 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000055.

Diver:1991:MWC

[Div91]  D. A. Diver.  Modelling

waves with computer algebra. Journal of Symbolic Computation, 11(3):275–290 (or 275– 289??), March 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dudley:1989:CAD

[DJ89] M. L. Dudley and R. W.

James. Computer-aided derivation of spherical harmonic spectral equations in astro-geophysics. Journal of Symbolic Computation, 8(4): 423–427, October 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dougherty:1992:IGU

[DJ92] Daniel J. Dougherty and Patricia Johann. An improved general E-unification method.

Journal of Symbolic Computation, 14(4):303–320, October 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DeBosschere:1996:EFL

[DJ96] Koen De Bosschere and Jean-Marie Jacquet. Extending the μlog framework with local and conditional blackboard operations. Journal of Symbolic Computation, 21(4/5/6):669– 698 (or 669–697??), April, May & June 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

DAndrea:2005:SGM

[DJ05] Carlos D’Andrea and Gabriela Jeronimo. Subresultants and generic monomial bases. Journal of Symbolic Computation, 39(3–4):259–277, March/April 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dumnicki:2007:NEB

[DJ07] Marcin Dumnicki and Witold Jarnicki. New effective bounds on the dimension of a linear system in P^{2}. Journal of Symbolic Computation, 42 (6):621–635, June 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dramnesc:2015:SLA

[DJ15] Isabela Dramnesc and Tudor Jebelean. Synthesis of list algorithms by mechanical proving. Journal of Symbolic Computation, 69 (??):61–92, July/August 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000984.

Derksen:2005:QAA

[DJK05] Harm Derksen, Emmanuel Jeandel, and Pascal Koiran. Quantum automata and algebraic groups. Journal of Symbolic Computation, 39(3–4): 357–371, March/April 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DAlfonso:2011:GIR

[DJO^{+}11] L. D’Alfonso, G. Jeronimo, F. Ollivier, A. Sedoglavic, and P. Solernó. A geometric index reduction method for implicit systems of differential algebraic equations. Journal of Symbolic Computation, 46(10):1114–1138, October 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000836.

Dramnesc:2018:MSS

[DJS18] Isabela Dramnesc, Tudor Jebelean, and Sorin Stratulat. Mechanical synthesis of sorting algorithms for binary trees by logic and combinatorial techniques. Journal of Symbolic Computation, 90(??):3–41, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300312.

Diekert:2016:LCG

[DK16] Volker Diekert and Jonathan Kausch. Logspace computations in graph products. Journal of Symbolic Computation, 75(??):94–109, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711500108X.

Davenport:2018:SCS

[DK18] James Harold Davenport and Temur Kutsia. Symbolic computation in software science. Journal of Symbolic Computation, 90(??):1–2, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300300.

Dumas:2021:EBC

[DKLP21] Jean-Guillaume Dumas, Erich Kaltofen, David Lucas, and Clément Pernet. Eliminationbased certificates for triangular equivalence and rank profiles. Journal of Symbolic Computation, 98(??):246–269, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300781.

Dundua:2021:VEM

[DKM21] Besik Dundua, Temur Kutsia, and Mircea Marin. Variadic equational matching in associative and commutative theories. Journal of Symbolic Computation, 106(??): 78–109, September/October 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000079.

DAndrea:2023:SIJ

[DKM23] Carlos D’Andrea, Kaie Kubjas, and Fatemeh Mohammadi. Special issue of JSC on the occasion of MEGA 2021. Journal of Symbolic Computation, 114(??): 99–101, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000281.

DAndrea:2015:SSS

[DKS15] Carlos D’Andrea, Teresa Krick, and Agnes Szanto. Subresultants, Sylvester sums and the rational interpolation problem. Journal of Symbolic Computation, 68 (part 1)(??):72–83, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000583.

Dixmier:1988:MNF

[DL88] J. Dixmier and D. Lazard. Minimum number of fundamental invariants for the binary form of degree 7. Journal of Symbolic Computation, 6(1):113–116 (or 113– 115??), August 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dershowitz:1993:LD

[DL93] N. Dershowitz and Y.-L. Lee. Logical debugging. Journal of Symbolic Computation, 15(5–6):745–774 (or 745– 773??), May/June 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Drensky:2006:GBI

[DL06] Vesselin Drensky and Roberto La Scala. Gröbner bases of ideals invariant under endomorphisms. Journal of Symbolic Computation, 41(7): 835–846, July 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dupont:2008:NOPa

[DLLP08a] Laurent Dupont, Daniel Lazard, Sylvain Lazard, and Sylvain Petitjean. Near-optimal parameterization of the intersection of quadrics: I. The generic algorithm. Journal of Symbolic Computation, 43(3): 168–191, March 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dupont:2008:NOPb

[DLLP08b] Laurent Dupont, Daniel Lazard, Sylvain Lazard, and Sylvain Petitjean. Near-optimal parameterization of the intersection of quadrics: II. A classification of pencils. Journal of Symbolic Computation, 43 (3):192–215, March 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dupont:2008:NOPc

[DLLP08c] Laurent Dupont, Daniel Lazard, Sylvain Lazard, and Sylvain Petitjean. Near-optimal parameterization of the intersection of quadrics: III. Parameterizing singular intersections. Journal of Symbolic Computation, 43(3):216– 232, March 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DeLoera:2011:CIC

[DLMM11] Jesús A. De Loera, Jon Lee, Peter N. Malkin, and Susan Margulies. Computing infeasibility certificates for combinatorial problems through Hilbert’s Nullstellensatz. Journal of Symbolic Computation, 46(11): 1260–1283, November 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001192.

deMan:1999:GFN

[dM99] Ronald de Man. The generating function for the number of roots of a Coxeter group. Journal of Symbolic Computation, 27(6):535–541, June 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0280/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0280/production/pdf;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0280/production/ref.

Delahaye:2005:DAE

[DM05] David Delahaye and Micaela Mayero. Dealing with algebraic expressions over a field in Coq using Maple. Journal of Symbolic Computation, 39 (5):569–592, May 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DAndrea:2009:SIS

[DM09] Carlos D’Andrea and Bernard Mourrain. Special issue on symbolic and algebraic computation. Journal of Symbolic Computation, 44(7):701–702, July 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Duchamp:2017:HSP

[DMN17] G. H. E. Duchamp, V. Hoang Ngoc Minh, and Quoc Hoan Ngo. Harmonic sums and polylogarithms at non-positive multiindices. Journal of Symbolic Computation, 83(??):166–186, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301298.

Diatta:2012:IMA

[DMR12] Daouda Niang Diatta, Bernard Mourrain, and Olivier Ruatta. On the isotopic meshing of an algebraic implicit surface. Journal of Symbolic Computation, 47(8):903–925, August 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001453.

Dwivedi:2021:EFP

[DMS21] Ashish Dwivedi, Rajat Mittal, and Nitin Saxena. Efficiently factoring polynomials modulo p^{4}. Journal of Symbolic Computation, 104 (??):805–823, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120301115.

Diekert:2017:ASG

[DMW17] Volker Diekert, Alexei G. Myasnikov, and Armin Weiß. Amenability of Schreier graphs and strongly generic algorithms for the conjugacy problem. Journal of Symbolic Computation, 83(??):147–165, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301286.

Ducos:2016:CSF

[DMY16] Lionel Ducos, Samiha Monceur, and Ihsen Yengui. Computing the V -saturation of finitely-generated submodules of V [X]^{m} where V is a valuation domain. Journal of Symbolic Computation, 72(??): 196–205, January/February 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000152.

deNivelle:2003:DGF

[dNdR03] Hans de Nivelle and Maarten de Rijke. Deciding the guarded fragments by resolution. Journal of Symbolic Computation, 35(1):21– 58, January 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Doliskani:2021:DMC

[DNS21] Javad Doliskani, Anand Kumar Narayanan, and Éric Schost. Drinfeld modules with complex multiplication, Hasse invariants and factoring polynomials over finite fields. Journal of Symbolic Computation, 105(??):199–213, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300493.

Degtyarev:2003:SR

[DNV03] Anatoli Degtyarev, Robert Nieuwenhuis, and Andrei Voronkov. Stratified resolution. Journal of Symbolic Computation, 36(1–2):79–99, July/August 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DeLoera:2006:MBT

[DO06] Jesús A. De Loera and Shmuel Onn. Markov bases of threeway tables are arbitrarily complicated. Journal of Symbolic Computation, 41(2):173– 181, February 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dohm:2009:IRR

[Doh09] Marc Dohm. Implicitization of rational ruled surfaces with μ-bases. Journal of Symbolic Computation, 44(5):479–489, May 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Domenjoud:1992:AUT

[Dom92] Eric Domenjoud. AC unification through order-sorted AC1 unification. Journal of Symbolic Computation, 14(6): 537–556, December 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Donch:2013:CRG

[Dön13] Christian Dönch. Characterization of relative Gröbner bases. Journal of Symbolic Computation, 55(??): 19–29, August 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000473.

Draisma:2017:F

[DOR17] Jan Draisma, Giorgio Ottaviani, and Fabrice Rouillier. Foreword. Journal of Symbolic Computation, 79 (part 1)(??):1–3, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300803.

Doris:2021:EAC

[Dor21] Christopher Doris. Exact padic computation in Magma. Journal of Symbolic Computation, 104(??):476–493, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300869.

dosSantos:1989:URS

[dos89] Renato P. dos Santos. Using REDUCE in supersymmetry. Journal of Symbolic Computation, 7(5):523–525, May 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dridi:2009:NCT

[DP09] Raouf Dridi and Michel Petitot. New classification techniques for ordinary differential equations. Journal of Symbolic Computation, 44(7): 836–851, July 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dana-Picard:2019:ASI

[DP19] Thierry Dana-Picard. An automated study of isoptic curves of an astroid. Journal of Symbolic Computation, 97(??):56–68, ???? 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118301317.

Davis:2024:RDC

[DP24] Maria M. Davis and Dávid Papp. Rational dual certificates for weighted sumsof-squares polynomials with boundable bit size. Journal of Symbolic Computation, 121(??):Article 102254, March/April 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000688.

DeLoera:2016:RSC

[DPS16] Jesús A. De Loera, Sonja Petrović, and Despina Stasi. Random sampling in computational algebra: Helly numbers and violator spaces. Journal of Symbolic Computation, 77(??):1–15, November/December 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711600002X.

Dumas:2017:FCR

[DPS17] Jean-Guillaume Dumas, Clément Pernet, and Ziad Sultan. Fast computation of the rank profile matrix and the generalized Bruhat decomposition. Journal of Symbolic Computation, 83(??):187–210, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301304.

Dumas:2023:SFA

[DPS23] Jean-Guillaume Dumas, Clément Pernet, and Alexandre Sedoglavic. Some fast algorithms multiplying a matrix by its adjoint. Journal of Symbolic Computation, 115(??): 285–315, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712200075X.

Duncan:1986:RUM

[DR86] Anthony Duncan and Ralph Roskies. Representations of unusual mathematical structures in scientific applications of symbolic computation. Journal of Symbolic Computation, 2(2):201–206 (or 201–207??), June 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dorre:1992:SSF

[DR92] Jochen Dörre and William C.

Rounds. On subsumption and semiunification in feature algebras. Journal of Symbolic Computation, 13(4): 441–461, April 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dershowitz:1993:DIS

[DR93] Nachum Dershowitz and Uday S. Reddy. Deductive and inductive synthesis of equational programs. Journal of Symbolic Computation, 15(5–6):467–494, May/June 1993. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-

855X (electronic).

Dettweiler:2000:AKA

[DR00] Michael Dettweiler and Stefan Reiter. An algorithm of Katz and its application to the inverse Galois problem. Journal of Symbolic Computation, 30(6):761–798, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0382;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.0382/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0382/ref.

Debus:2023:RGC

[DR23a] Sebastian Debus and Cordian Riener. Reflection groups and cones of sums of squares. Journal of Symbolic Computation, 119(??):112–144, November/December 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000263.

Duff:2023:SAC

[DR23b] Timothy Duff and Michael Ruddy. Signatures of algebraic curves via numerical algebraic geometry. Journal of Symbolic Computation, 115 (??):452–477, March/April 2023. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000694.

Draxler:2001:NFR

[Drä01]  Peter Dräxler.  Normal

forms for representations of representation-finite algebras. Journal of Symbolic Computation, 32(5): 491–497, November 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0480;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0480/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0480/ref.

Draisma:2003:CLA

[Dra03] Jan Draisma. Constructing Lie algebras of first order differential operators. Journal of Symbolic Computation, 36(5): 685–698, November 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Draisma:2005:RTO

[Dra05]  Jan Draisma.  Represen-

tation theory on the open Bruhat cell. Journal of Symbolic Computation, 39(3–4): 279–303, March/April 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

lasCuevas:2024:TDS

[DRN24] Gemma De las Cuevas, Matt Hoogsteder Riera, and Tim Netzer. Tensor decompositions on simplicial complexes with invariance. Journal of Symbolic Computation, 124 (??):??, September/October 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717124000038.

Drton:2006:CAR

[Drt06] Mathias Drton. Computing all roots of the likelihood equations of seemingly unrelated regressions. Journal of Symbolic Computation, 41(2): 245–254, February 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Davenport:1986:ELS

[DS86] J. H. Davenport and M. F. Singer. Elementary and Liouvillian solutions of linear differential equations. Journal of Symbolic Computation, 2 (3):237–260, September 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Denzinger:1996:RAK

[DS96] J. Denzinger and S. Schulz. Recording and analysing knowledge-based distributed deduction processes. Journal of Symbolic Computation, 21(4/5/6):523–541 (or 523– 542??), April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dolzmann:1997:SQF

[DS97] Andreas Dolzmann and Thomas Sturm. Simplification of quantifier-free formulae over ordered fields. Journal of Symbolic Computation, 24 (2):209–231, August 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Applications of quantifier elimination (Albuquerque, NM, 1995). This paper is not a part of the special issue, but is on a related topic.

Decker:2000:NGT

[DS00] Wolfram Decker and Frank-Olaf Schreyer. Non-general type surfaces in P^{4}: Some remarks on bounds and constructions. Journal of Symbolic Computation, 29 (4–5):545–582, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0323;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0323/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0323/ref.

Dickenstein:2002:ETC

[DS02] Alicia Dickenstein and Bernd Sturmfels. Elimination theory in codimension 2. Journal of Symbolic Computation, 34(2): 119–135, August 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dolzmann:2006:E

[DS06] Andreas Dolzmann and Thomas Sturm. Editorial. Journal of Symbolic Computation, 41(11):1155–1156, November 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Das:2009:LDC

[DS09] M. Prem Laxman Das and Kripasindhu Sikdar. List decoding codes on Garcia– Stictenoth tower using Gröbner basis. Journal of Symbolic Computation, 44(12):1657– 1661, December 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DeFeo:2012:FAA

[DS12] Luca De Feo and Éric Schost. Fast arithmetics in Artin– Schreier towers over finite fields. Journal of Symbolic Computation, 47(7):771–792, July 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002008.

Durand:2015:BRW

[DS15] I. Durand and G. Sénizergues.

Bottom-up rewriting for words and terms. Journal of Symbolic Computation, 67(??):93–121, March/April 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000418.

Dimca:2016:SJI

[DS16] Alexandru Dimca and Gabriel Sticlaru. Syzygies of Jacobian ideals and weighted homogeneous singularities. Journal of Symbolic Computation, 74(??):627–634, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001017.

Deza:2018:HCP

[DS18a] Michel Deza and Mathieu Dutour Sikirić. The hypermetric cone and polytope on eight vertices and some generalizations. Journal of Symbolic Computation, 88(??):67– 84, September/October 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300154.

Dimca:2018:CMP

[DS18b] Alexandru Dimca and Gabriel Sticlaru. Computing the monodromy and pole order filtration on Milnor fiber cohomology of plane curves. Journal of Symbolic Computation, 91(??):98–115, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300804.

Dumas:2001:ESI

[DSV01] Jean-Guillaume Dumas, B. David Saunders, and Gilles Villard. On efficient sparse integer matrix Smith Normal Form computations. Journal of Symbolic Computation, 32(1–2):71–99, July/August 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0451;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0451/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0451/ref.

Dahan:2009:EPI

[DSW09] Xavier Dahan, Éric Schost, and Jie Wu. Evaluation properties of invariant polynomials. Journal of Symbolic Computation, 44(11):1592– 1604, November 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DiBlasio:1995:SIA

[DT95] Paolo Di Blasio and Marco Temperini. Subtyping inheritance and its application in languages for symbolic computation systems. Journal of Symbolic Computation, 19 (1/2/3):39–64 (or 39–63??), January, February, March 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Diaz-Toca:2006:GTS

[DT06] Gema M. Diaz-Toca. Galois theory, splitting fields and computer algebra. Journal of Symbolic Computation, 41(11):1174–1186, November 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Du:2023:PFC

[DT23] Julia Q. D. Du and Dazhao Tang. Proofs of five conjectures on matching coefficients of Baruah, Das and Schlosser by an algorithmic approach. Journal of Symbolic Computation, 116(??):213–242, May/ June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000979.

Diaz-Toca:2001:SDU

[DTGV01] Gema M. Diaz-Toca and Laureano Gonzalez-Vega. Squarefree decomposition of univariate polynomials depending on a parameter. application to the integration of parametric rational functions. Journal of Symbolic Computation, 32(3):191–209, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0471;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0471/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0471/ref.

Diaz-Toca:2002:BTA

[DTGV02] Gema M. Diaz-Toca and Laureano Gonzalez-Vega. Barnett’s theorems about the greatest common divisor of several univariate polynomials through Bezout-like matrices. Journal of Symbolic Computation, 34(1):59–81, July 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Diaz-Toca:2010:DGT

[DTL10] G. M. Diaz-Toca and H. Lombardi. Dynamic Galois theory. Journal of Symbolic Computation, 45(12):1316– 1329, December 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

DuCloux:1999:TAC

[Du 99] F. Du Cloux. A transducer approach to Coxeter groups. Journal of Symbolic Computation, 27(3):311–324, March 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dube:1993:CPE

[Dub93] Thomas W. Dubé. A combinatorial proof of the effective Nullstellensatz. Journal of Symbolic Computation, 15 (3):277–296, March 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dumnicki:2009:ABR

[Dum09] Marcin Dumnicki. An algorithm to bound the regularity and nonemptiness of linear systems in P^{n}. Journal of Symbolic Computation, 44(10):1448–1462, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Duntsch:1994:MBS

[Dün94] Ivo Düntsch. A microcomputer based system for small relation algebras. Journal of Symbolic Computation, 18 (1):83–86, July 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dunkl:1999:CDD

[Dun99] Charles F. Dunkl. Computing with differential-difference operators. Journal of Symbolic Computation, 28(6): 819–826, December 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1997.0341/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1997.0341/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1997.0341/production/ref.

Dur:1989:CCF

[Dür89] Arne Dür. On computing the canonical form for a binary form of odd degree. Journal of Symbolic Computation, 8(4): 327–334 (or 327–333??), October 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Durand:1994:BSS

[Dur94]  Iréne Durand.  Bounded,

strongly sequential and forwardbranching term rewriting systems. Journal of Symbolic Computation, 18(4):319– 352, October 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Durvye:2009:ETZ

[Dur09]  Clémence Durvye.  Eval-

uation techniques for zerodimensional primary decomposition. Journal of Symbolic Computation, 44(9):

1089–1113, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Duval:1994:ANE

[Duv94]  Dominique Duval.  Alge-

braic numbers: An example of dynamic evaluation. Journal of Symbolic Computation, 18(5):429–446 (or 429– 445??), November 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

deJager:1996:SSC

[dv96] Bram de Jager and Bram van Asch. Symbolic solutions for a class of partial differential equations. Journal of Symbolic Computation, 22(4): 459–468, October 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dieulefait:2000:PLG

[DV00] Luis Dieulefait and Núria Vila. Projective linear groups as Galois groups over Q via modular representations. Journal of Symbolic Computation, 30 (6):799–810, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0383;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0383/

                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0383/ref.

Draisma:2021:CMT

[DV21] Jan Draisma and Alejandro Vargas. Catalan-many tropical morphisms to trees; Part I: Constructions. Journal of Symbolic Computation, 104 (??):580–629, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300961.

Dolzmann:2007:MOS

[DW07] Andreas Dolzmann and Volker Weispfenning. Multiple object semilinear motion planning. Journal of Symbolic Computation, 42(3):324–337, March 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dietrich:2018:SPL

[DW18] Heiko Dietrich and Ian M. Wanless. Small partial latin squares that embed in an infinite group but not into any finite group. Journal of Symbolic Computation, 86 (??):142–152, May/June 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300494.

Dong:2021:CSR

[DW21] Rina Dong and Dongming Wang. Computing strong regular characteristic pairs with Gröbner bases. Journal of Symbolic Computation, 104(??):312–327, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300559.

Dreyfus:2022:CLA

[DW22] Thomas Dreyfus and Jacques-Arthur Weil. Computing the Lie algebra of the differential Galois group: the reducible case. Journal of Symbolic Computation, 112(??): 122–163, September/October 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000062.

Dorato:1997:RMO

[DYA97] Peter Dorato, Wei Yang, and Chaouki Abdallah. Robust multi-objective feedback design by quantifier elimination. Journal of Symbolic Computation, 24(2):153–160 (or 153–159??), August 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Applications of quantifier elimination (Albuquerque, NM, 1995).

Dohm:2009:IEC

[DZ09] Marc Dohm and Severinas Zube. The implicit equation of a canal surface. Journal of Symbolic Computation, 44(2): 111–130, February 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dehbi:2022:NTS

[DZY22] Lydia Dehbi, Zhenbing Zeng, and Lu Yang. The number of tetrahedra sharing the same metric invariants via symbolic and numerical computations. Journal of Symbolic Computation, 108 (??):41–54, January/February 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000389.

England:2021:CAD

[EBD21] Matthew England, Russell Bradford, and James H. Davenport. Cylindrical algebraic decomposition with equational constraints. Journal of Symbolic Computation, 100(??):38–71, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300859.

Ebert:2001:CFG

[Ebe01]  G. L. Ebert.  Construc-

tions in finite geometry using computer algebra systems. Journal of Symbolic Computation, 31(1– 2):55–70, January/February 2001. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.1006;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1998.1006/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.1006/ref.

Eberly:2019:AAS

[Ebe19] Wayne Eberly. Automating algorithm selection: Checking for matrix properties that can simplify computations. Journal of Symbolic Computation, 94(??):1– 29, September/October 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300713.

Elishakoff:1987:ASA

[EC87] Isaac Elishakoff and Brian Couch. Application of symbolic algebra to the instability of a nonconservative system. Journal of Symbolic Computation, 4(3):391–396, December 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Emiris:1995:EIA

[EC95] Ioannis Z. Emiris and John F.

Canny. Efficient incremental algorithms for the sparse resultant and the mixed volume. Journal of Symbolic Computation, 20(2):117–150 (or 117– 149??), August 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eckhardt:1987:CCN

[Eck87] Carsten Eckhardt. Computation of class numbers by an analytic method. Journal of Symbolic Computation, 4(1): 41–52, August 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eder:1985:PSU

[Ede85] Elmar Eder. Properties of substitutions and unifications. Journal of Symbolic Computation, 1(1):31–46, March 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eder:2013:AIS

[Ede13] Christian Eder. An analysis of inhomogeneous signaturebased Gröbner basis computations. Journal of Symbolic Computation, 59(??):21– 35, December 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001089.

Editors:1985:SCE

[Edi85] Editors. Symbolic computation (an editorial). Journal of Symbolic Computation, 1(1): 1–6, March 1985.

East:2019:CFS

[EENMP19] James East, Attila Egri-Nagy, James D. Mitchell, and Yann Péresse. Computing finite semigroups. Journal of Symbolic Computation, 92 (??):110–155, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300087.

Eder:2017:SSB

[EF17] Christian Eder and Jean-Charles Faugère. A survey on signature-based algorithms for computing Gröbner bases. Journal of Symbolic Computation, 80 (part 3) (??):719–784, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300785.

Emiris:2016:EES

[EFG16] Ioannis Z. Emiris, Vissarion Fisikopoulos, and Bernd Gärtner. Efficient edgeskeleton computation for polytopes defined by oracles. Journal of Symbolic Computation, 73(??): 139–152, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000425.

Eriksson:2006:PCN

[EFRS06] Nicholas Eriksson, Stephen E. Fienberg, Alessandro Rinaldo, and Seth Sullivant. Polyhedral conditions for the nonexistence of the MLE for hierarchical log-linear models. Journal of Symbolic Computation, 41(2):222–233, February 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eberly:2000:EDA

[EG00] W. Eberly and M. Giesbrecht. Efficient decomposition of associative algebras over finite fields. Journal of Symbolic Computation, 29(3):441–458, March 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0308;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0308/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0308/ref.

Eberly:2004:EDS

[EG04] W. Eberly and M. Giesbrecht. Efficient decomposition of separable algebras. Journal of Symbolic Computation, 37(1): 35–81, January 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eriksen:2007:COE

[EG07] Eivind Eriksen and Trond Stølen Gustavsen. Computing obstructions for existence of connections on modules. Journal of Symbolic Computation, 42 (3):313–323, March 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Esterov:2015:SES

[EG15] Alexander Esterov and Gleb Gusev. Systems of equations with a single solution. Journal of Symbolic Computation, 68 (part 2)(??):116–130, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000753.

Eick:2021:CSM

[EG21] Bettina Eick and Taleea Jalaeeyan Ghorbanzadeh. Computing the Schur multipliers of the Lie p-rings in the family defined by a symbolic Lie p-ring presentation. Journal of Symbolic Computation, 106(??): 68–77, September/October 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301309.

Elkadi:2012:AGS

[EGB12] Mohamed Elkadi, André Galligo, and Thang Luu Ba. Approximate GCD of several univariate polynomials with small degree perturbations. Journal of Symbolic Computation, 47(4):410–421, April 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001404.

Egly:1996:DSP

[Egl96]  Uwe Egly.  On different

structure-preserving translations to normal form. Journal of Symbolic Computation, 22(2):121–142, August 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Elliott:2023:BCC

[EGS23] Jesse Elliott, Mark Giesbrecht, and Éric Schost. Bit complexity for computing one point in each connected component of a smooth real algebraic set. Journal of Symbolic Computation, 116 (??):72–97, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000761.

Elkadi:2009:TTA

[EGW09] M. Elkadi, A. Galligo, and M. Weimann. Towards toric absolute factorization. Journal of Symbolic Computation, 44(9):1194–1211, September 2009. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Erascu:2016:RQE

[EH16] Madalina Erascu and Hoon Hong. Real quantifier elimination for the synthesis of optimal numerical algorithms (case study: Square root computation). Journal of Symbolic Computation, 75(??): 110–126, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001091.

Eder:2021:EGB

[EH21] Christian Eder and Tommy Hofmann. Efficient Gröbner bases computation over principal ideal rings. Journal of Symbolic Computation, 103 (??):1–13, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301336.

Epstein:1991:UKB

[EHR91] D. B. A. Epstein, D. F.

Holt, and S. E. Rees. The use of Knuth–Bendix methods to solve the word problem in automatic groups. Journal of Symbolic Computation, 12(4–5):397–414, October/November 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Eick:2002:OSP

[Eic02] Bettina Eick. Orbit-stabilizer problems and computing normalizers for polycyclic groups. Journal of Symbolic Computation, 34(1):1–19, July 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eick:2010:SNS

[Eic10] Bettina Eick. Some new simple Lie algebras in characteristic 2. Journal of Symbolic Computation, 45(9):943– 951, September 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eid:2023:ECC

[Eid23] Elie Eid. Efficient computation of Cantor’s division polynomials of hyperelliptic curves over finite fields. Journal of Symbolic Computation, 117(??):68–100, July/August 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122001031.

Eisenberger:1990:ASA

[Eis90] Moshe Eisenberger. Application of symbolic algebra to the analysis of plates on variable elastic foundation. Journal of Symbolic Computation, 9(2):207–214 (or 207– 213??), February 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eiter:1994:ETH

[Eit94]  Thomas Eiter.  Exact

transversal hypergraphs and application to Boolean μfunctions. Journal of Symbolic Computation, 17(3):215– 226 (or 215–225??), March 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Emiris:2011:SLM

[EK11] Ioannis Z. Emiris and Christos Konaxis. Single-lifting Macaulay-type formulae of generalized unmixed sparse resultants. Journal of Symbolic Computation, 46(8):919– 942, August 2011. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000356.

Elsenhans:2019:CSN

[EK19] Andreas-Stephan Elsenhans and Jürgen Klüners. Computing subfields of number fields and applications to Galois group computations. Journal of Symbolic Computation, 93 (??):1–20, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300439.

Ellis:2021:CLC

[EK21] Graham Ellis and Kelvin Killeen. Cohomology with local coefficients and knotted manifolds. Journal of Symbolic Computation, 107(??):299–321, November/ December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000286.

Edwards:2022:CPG

[EKP22] John Edwards, Cameron Krome, and Tracy L. Payne. Computation of positively graded filiform nilpotent Lie algebras in low dimensions. Journal of Symbolic Computation, 108(??):73–90, January/February 2022. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000377.

ElKahoui:2003:EAS

[El 03] Mhammed El Kahoui. An elementary approach to subresultants theory. Journal of Symbolic Computation, 35(3): 281–292, March 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

ElKahoui:2005:BPG

[El 05] M’hammed El Kahoui. Birational properties of the gap subresultant varieties. Journal of Symbolic Computation, 39 (1):61–71, January 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

ElKahoui:2008:TRA

[El 08] M’hammed El Kahoui. Topology of real algebraic space curves. Journal of Symbolic Computation, 43(4): 235–258, April 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ellis:2012:CHT

[EL12] Graham Ellis and Le Van Luyen. Computational homology of n-types. Journal of Symbolic Computation, 47(11):1309–1317, November 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711200017X.

Elias:2004:CRR

[Eli04] Juan Elias. On the computation of the Ratliff–Rush closure. Journal of Symbolic Computation, 37(6):717–725, June 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ellis:2004:CGR

[Ell04] Graham Ellis. Computing group resolutions. Journal of Symbolic Computation, 38(3): 1077–1118, September 2004.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eder:2023:SBA

[ELME23] Christian Eder, Pierre Lairez, Rafael Mohr, and Mohab Safey El Din. A signaturebased algorithm for computing the nondegenerate locus of a polynomial system. Journal of Symbolic Computation, 119(??):1–21, November/December 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000111.

Elsenhans:2012:ICI

[Els12] Andreas-Stephan Elsenhans. Invariants for the computation of intransitive and transitive Galois groups. Journal of Symbolic Computation, 47(3):315–326, March 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001921.

Elsenhans:2015:ECI

[Els15] Andreas-Stephan Elsenhans. Explicit computations of invariants of plane quartic curves. Journal of Symbolic Computation, 68 (part 2) (??):109–115, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000741.

Elsenhans:2017:IMC

[Els17] Andreas-Stephan Elsenhans.

Improved methods for the construction of relative invariants for permutation groups. Journal of Symbolic Computation, 79 (part 2)(??): 211–231, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000146.

Egner:1998:SRM

[EM98] Sebastian Egner and Torsten Minkwitz. Sparsification of rectangular matrices. Journal of Symbolic Computation, 26(2):135–149, 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Emiris:1999:MET

[EM99] Ioannis Z. Emiris and Bernard Mourrain. Matrices in elimination theory. Journal of Symbolic Computation, 28(1–2):3–43, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0265;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1998.0265/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.

                                        1998.0266;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1998.0266/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1998.0266/ref.

Emiris:2012:MRF

[EM12] Ioannis Z. Emiris and Angelos Mantzaflaris. Multihomogeneous resultant formulae for systems with scaled support. Journal of Symbolic Computation, 47(7):820–842, July 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002021.

Eisenbud:2000:CTV

[EMS00] David Eisenbud, Mircea Mustaţa, and Mike Stillman. Cohomology on toric varieties and local cohomology with monomial supports. Journal of Symbolic Computation, 29(4–5):583–600, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0326;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0326/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0326/ref.

Erocal:2016:RAC

[EMSS16] Burçin Eröcal, Oleksandr Motsak, Frank-Olaf Schreyer, and Andreas Steenpaß. Refined algorithms to compute syzygies. Journal of Symbolic Computation, 74(??): 308–327, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000735.

Emiris:2020:SBP

[EMT20] Ioannis Emiris, Bernard Mourrain, and Elias Tsigaridas. Separation bounds for polynomial systems. Journal of Symbolic Computation, 101(??):128–151, November/ December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300641.

Emiris:2021:MPS

[EMT21] Ioannis Z. Emiris, Angelos Mantzaflaris, and Elias P. Tsigaridas. Multilinear polynomial systems: Root isolation and bit complexity. Journal of Symbolic Computation, 105(??):145–164, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712030047X.

Encarnacion:1995:CGP

[Enc95] Mark J. Encarnación. Computing GCDs of polynomials over algebraic number fields.

Journal of Symbolic Computation, 20(3):299–314 (or 299– 313??), September 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Engheta:2010:BPD

[Eng10] Bahman Engheta. A bound on the projective dimension of three cubics. Journal of Symbolic Computation, 45(1): 60–73, January 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Emiris:2002:SNM

[EP02] Ioannis Z. Emiris and Victor Y. Pan. Symbolic and numeric methods for exploiting structure in constructing resultant matrices. Journal of Symbolic Computation, 33 (4):393–413, April 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Egner:2004:SBM

[EP04] Sebastian Egner and Markus Püschel. Symmetry-based matrix factorization. Journal of Symbolic Computation, 37(2): 157–186, February 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eder:2010:FVF

[EP10] Christian Eder and John Perry. F5C: a variant of Faugère’s F5 algorithm with reduced Gröbner bases. Journal of Symbolic Computation, 45(12):1442–1458, December 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eder:2021:SBE

[EPP21] Christian Eder, Gerhard Pfister, and Adrian Popescu. Standard bases over Euclidean domains. Journal of Symbolic Computation, 102 (??):21–36, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301130.

Edelsbrunner:1990:TPS

[EPW90] H. Edelsbrunner, F. P.

Preparata, and D. B. West. Tetrahedrizing point sets in three dimensions. Journal of Symbolic Computation, 10(3–4):335–348 (or 335– 347??), September/October 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Emiris:1998:MAL

[EPY98] Ioannis Z. Emiris, Victor Y.

Pan, and Yanqiang Yu. Modular arithmetic for linear algebra computations in the real field. Journal of Symbolic Computation, 26(1):71–87, July 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eaves:1995:LPL

[ER95] B. Curtis Eaves and Uriel G. Rothblum. Linear problems and linear algorithms. Journal of Symbolic Computation, 20(2):207–214, August 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Er-Riani:2005:SLD

[ERSG05] Mustapha Er-Riani and Olivier Sero-Guillaume. Shapes of liquid drops obtained using symbolic computation. Journal of Symbolic Computation, 40(6): 1340–1360, December 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Eikenberry:1998:EAC

[ES98] S. Meyer Eikenberry and J. P. Sorenson. Efficient algorithms for computing the Jacobi symbol. Journal of Symbolic Computation, 26(4):509– 523, October 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ellis:2011:CGC

[ES11] Graham Ellis and Paul Smith. Computing group cohomology rings from the Lyndon– Hochschild–Serre spectral sequence. Journal of Symbolic Computation, 46(4): 360–370, April 2011. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001501.

Emiris:2013:SIS

[ES13] Ioannis Z. Emiris and Éric Schost. Special issue on symbolic and algebraic computation: Foundations, algorithmics and applications: ISSAC 2011. Journal of Symbolic Computation, 52 (??):1–2, May 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001253.

ElDin:2018:BCM

[ES18] Mohab Safey El Din and Éric Schost. Bit complexity for multi-homogeneous polynomial system solving — application to polynomial minimization. Journal of Symbolic Computation, 87(??): 176–206, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300937.

Ezome:2020:NBD

[ES20] Tony Ezome and Mohamadou Sall. Normal bases from 1dimensional algebraic groups. Journal of Symbolic Computation, 101(??):152–169, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300653.

ElManssour:2023:CDA

[ES23] Rida Ait El Manssour and Anna-Laura Sattelberger. Combinatorial differential algebra of x^{p}. Journal of Symbolic Computation, 114(??): 193–208, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771712200030X.

Engelfriet:1996:ETL

[ET96] Joeri Engelfriet and Jan Treur. Executable temporal logic for non-monotonic reasoning. Journal of Symbolic Computation, 22(5–6):615– 626 (or 615–625??), November/December 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Edelsbrunner:1986:CHS

[EW86] H. Edelsbrunner and R. Waupotitsch. Computing a hamsandwich cut in two dimensions. Journal of Symbolic Computation, 2(2):171–178, June 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

ElKahoui:2000:DHB

[EW00] M’hammed El Kahoui and Andreas Weber. Deciding Hopf bifurcations by quantifier elimination in a software-component architecture. Journal of Symbolic Computation, 30(2):161– 179, August 2000. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0353;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0353/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0353/ref.

Eick:2002:CSE

[EW02] Bettina Eick and Charles R. B. Wright. Computing subgroups by exhibition in finite solvable groups. Journal of Symbolic Computation, 33(2):129–143, February 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0503;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0503/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0503/ref.

Evans:2007:CIR

[EW07] Gareth A. Evans and Christopher D. Wensley. Complete involutive rewriting systems. Journal of Symbolic Computation, 42(11–12):1034–1051, November/December 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

ElDin:2021:SIS

[EY21] Mohab Safey El Din and Chee Yap. Special issue on symbolic and algebraic computation: ISSAC 2017. Journal of Symbolic Computation, 98(??):1–2, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300665.

ElDin:2021:CRR

[EYZ21] Mohab Safey El Din, Zhi-Hong Yang, and Lihong Zhi. Computing real radicals and S-radicals of polynomial systems. Journal of Symbolic Computation, 102(??):259– 278, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301294.

Fages:1987:ACU

[Fag87] François Fages. Associativecommutative unification. Journal of Symbolic Computation, 3(3):257–275, June 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Farouki:1997:CAC

[Far97] Rida T. Farouki. Conic approximation of conic offsets.

Journal of Symbolic Computation, 23(2–3):301–314 (or 301–313??), February/March 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Farouki:2019:EPH

[Far19] Rida T. Farouki. Existence of Pythagorean-hodograph quintic interpolants to spatial G^{1} Hermite data with prescribed arc lengths. Journal of Symbolic Computation, 95(??):202–216, November/December 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300197.

Fassino:2010:AVP

[Fas10] Claudia Fassino. Almost vanishing polynomials for sets of limited precision points. Journal of Symbolic Computation, 45(1):19–37, January 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fateman:1992:RM

[Fat92] Richard J. Fateman. A review of Mathematica. Journal of Symbolic Computation, 13 (5):545–579, May 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fahmy:1993:SRT

[FB93] Amr F. Fahmy and Alan W. Biermann. Synthesis of real time acceptors. Journal of Symbolic Computation, 15(5– 6):807–842, May/June 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Foo:2004:CEC

[FC04] Mao-Ching Foo and Eng-Wee Chionh. Corner edge cutting and Dixon A-resultant quotients. Journal of Symbolic Computation, 37(1):101– 119, January 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Flener:1993:LPS

[FD93] Pierre Flener and Yves Deville. Logic program synthesis from incomplete specifications. Journal of Symbolic Computation, 15(5–6): 775–806 (or 775–805??), May/ June 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Francis:2014:RGB

[FD14] Maria Francis and Ambedkar Dukkipati. Reduced Gröbner bases and Macaulay– Buchberger basis theorem over Noetherian rings. Journal of Symbolic Computation, 65(??):1–14, November 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711400011X.

Francis:2018:GBK

[FD18] Maria Francis and Ambedkar Dukkipati. On Gröbner bases and Krull dimension of residue class rings of polynomial rings over integral domains. Journal of Symbolic Computation, 86(??):1–19, May/June 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300251.

Farouki:2013:SVA

[FDS13] Rida T. Farouki, Petroula Dospra, and Takis Sakkalis. Scalar-vector algorithm for the roots of quadratic quaternion polynomials, and the characterization of quintic rational rotation-minimizing frame curves. Journal of Symbolic Computation, 58 (??):1–17, November 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000965.

Ferro:1988:GBH

[Fer88] Giuseppa Carrà Ferro. Gröbner bases and Hilbert schemes. I. Journal of Symbolic Computation, 6(2–3):219–230, October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Fernandez:1996:ACP

[Fer96] Maribel Fernández. AC complement problems: Satisfiability and negation elimination. Journal of Symbolic Computation, 22(1):49–82, July 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fernandez:1998:NEE

[Fer98] Maribel Fernández. Negation elimination in empty or permutative theories. Journal of Symbolic Computation, 26 (1):97–133, July 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ferro:2006:EIB

[Fer06a] Giuseppa Carrá Ferro. Erratum to “Ideals, bifiltered modules and bivariate Hilbert polynomials” [J. Symbolic Comput. 41 (2006) 112–121]. Journal of Symbolic Computation, 41(5):619, May 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [Fer06b].

Ferro:2006:IBM

[Fer06b] Giuseppa Carrá Ferro. Ideals, bifiltered modules and bivariate Hilbert polynomials. Journal of Symbolic Computation, 41(1):112–121, January 2006. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). See erratum [Fer06a].

Faugere:2011:GBB

[FES11] Jean-Charles Faugère, Mohab Safey El Din, and Pierre-Jean Spaenlehauer. Gröbner bases of bihomogeneous ideals generated by polynomials of bidegree (1, 1): Algorithms and complexity. Journal of Symbolic Computation, 46(4):406–437, April 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001902.

Faugere:2013:CGM

[FES13] Jean-Charles Faugère, Mohab Safey El Din, and Pierre-Jean Spaenlehauer. On the complexity of the generalized MinRank problem. Journal of Symbolic Computation, 55(??):30–58, August 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000485.

Faugere:2016:CCG

[FEV16] Jean-Charles Faugère, Mohab Safey El Din, and Thibaut Verron. On the complexity of computing Gröbner bases for weighted homogeneous systems. Journal of Symbolic Computation, 76(??):107–141, September/October 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115001273.

Fitzpatrick:1992:GBT

[FF92] Patrick Fitzpatrick and John Flynn. A Gröbner basis technique for Padé approximation. Journal of Symbolic Computation, 13(2):133–138, February 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ferreira:2017:IRS

[FF17] Jorge N. M. Ferreira and Peter Fleischmann. The invariant rings of the Sylow groups of GU(3, q^{2}), GU(4, q^{2}), Sp(4, q) and O^{+}(4, q) in the natural characteristic. Journal of Symbolic Computation, 79 (part 2)(??):356– 371, March/April 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116000225.

Fournie:1998:CMD

[FFP98] M. Fournié, J.-Ph. Furter, and D. Pinchon. Computation of the maximal degree of the inverse of a cubic automorphism of the affine plane with Jacobian 1 via Gröbner bases. Journal of Symbolic Computation, 26(3):381–386, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Feng:2006:PTA

[FG06] Ruyong Feng and Xiao-Shan Gao. A polynomial time algorithm for finding rational general solutions of first order autonomous ODEs. Journal of Symbolic Computation, 41 (7):739–762, July 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Flynn:2008:DIE

[FG08] E. V. Flynn and C. Grattoni.

Descent via isogeny on elliptic curves with large rational torsion subgroups. Journal of Symbolic Computation, 43 (4):293–303, April 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Farouki:2016:GRS

[FG16] Rida T. Farouki and Robert Gutierrez. Geometry of the ringed surfaces in R^{4} that generate spatial Pythagorean hodographs. Journal of Symbolic Computation, 73(??): 87–103, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000309.

Farouki:2016:SQQ

[FGG^{+}16] Rida T. Farouki, Graziano Gentili, Carlotta Giannelli, Alessandra Sestini, and Caterina Stoppato. Solution of a quadratic quaternion equation with mixed coefficients. Journal of Symbolic Computation, 74(??):140–151, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000577.

Feng:2008:RSO

[FGH08] Ruyong Feng, Xiao-Shan Gao, and Zhenyu Huang. Rational solutions of ordinary difference equations. Journal of Symbolic Computation, 43 (10):746–763, October 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fortuna:2004:ADT

[FGL04] E. Fortuna, P. Gianni, and D. Luminati. Algorithmical determination of the topology of a real algebraic surface. Journal of Symbolic Computation, 38(6):1551– 1567, December 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Falkensteiner:2023:IFT

[FGLH^{+}23] Sebastian Falkensteiner, Cristhian Garay-López, Mercedes Haiech, Marc Paul Noordman, François Boulier, and Zeinab Toghani. On initials and the fundamental theorem of tropical partial differential algebraic geometry. Journal of Symbolic Computation, 115(??): 53–73, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000712.

Faugere:1993:ECZ

[FGLM93] J. C. Faugere, P. Gianni, D. Lazard, and T. Mora. Efficient Computation of Zero-Dimensional Gröbner Bases by Change of Ordering. Journal of Symbolic Computation, 16(4):329–344, October 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fortuna:2005:SCR

[FGP05] E. Fortuna, P. Gianni, and P. Parenti. Some constructions for real algebraic curves. Journal of Symbolic Computation, 40(4–5):1169–1179, October/November 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Faugere:2014:MCA

[FGPGP14] Jean-Charles Faugère, Domingo Gómez-Pérez, Jaime Gutierrez, and Ludovic Perret. Mathematical and computer algebra techniques in cryptology. Journal of Symbolic Computation, 64(??):1–2, August 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001673.

Fortuna:2003:ACT

[FGPT03] E. Fortuna, P. Gianni, P. Parenti, and C. Traverso. Algorithms to compute the topology of orientable real algebraic surfaces. Journal of Symbolic Computation, 36(3–4): 343–364, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fleming:2003:DC

[FGR03] Michael Fleming, Ryan Gunther, and Robert Rosebrugh. A database of categories. Journal of Symbolic Computation, 35(2):127–135, February 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Farouki:2009:HPCa

[FGS09a] Rida T. Farouki, Carlotta Giannelli, and Alessandra Sestini. Helical polynomial curves and double Pythagorean hodographs I. Quaternion and Hopf map representations. Journal of Symbolic Computation, 44(2):161–179, February 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Farouki:2009:HPCb

[FGS09b] Rida T. Farouki, Carlotta Giannelli, and Alessandra Sestini. Helical polynomial curves and double Pythagorean hodographs II. Enumeration of low-degree curves. Journal of Symbolic Computation, 44 (4):307–332, April 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Farmer:1995:CMR

[FGT95] William M. Farmer, Joshua D.

Guttman, and F. Javier Thayer. Contexts in mathematical reasoning and computation. Journal of Symbolic Computation, 19(1/2/3): 201–216, January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Fortuna:2002:DRP

[FGT02] E. Fortuna, P. Gianni, and B. Trager. Derivations and radicals of polynomial ideals over fields of arbitrary characteristic. Journal of Symbolic Computation, 33(5):609–625, May 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fortuna:2005:IDP

[FGT05] E. Fortuna, P. Gianni, and B. Trager. Irreducible decomposition of polynomial ideals. Journal of Symbolic Computation, 39(3–4):305– 316, March/April 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fortuna:2009:GIA

[FGT09] E. Fortuna, P. Gianni, and B. Trager. Generators of the ideal of an algebraic space curve. Journal of Symbolic Computation, 44(9): 1234–1254, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fortuna:2015:CTI

[FGT15] E. Fortuna, P. Gianni, and B. M. Trager. Computation of topological invariants for real projective surfaces with isolated singularities. Journal of Symbolic Computation, 68 (part 2)(??):131–166, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000777.

Fioravanti:2006:CIT

[FGVN06] Mario Fioravanti, Laureano Gonzalez-Vega, and Ioana Necula. Computing the intersection of two ruled surfaces by using a new algebraic approach. Journal of Symbolic Computation, 41(11):1187– 1205, November 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Furbach:1986:MCF

[FH86] U. Furbach and S. Hölldobler.

Modelling the combination of functional and logic programming languages. Journal of Symbolic Computation, 2(2): 123–138, June 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Froberg:1994:HSI

[FH94] Ralf Fröberg and Joachim Hollman. Hilbert series for ideals generated by generic forms. Journal of Symbolic Computation, 17(2):149– 158 (or 149–157??), February 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fix:1996:IRP

[FHL96] George Fix, Chih-Ping Hsu, and Tie Luo. Implicitization of rational parametric surfaces. Journal of Symbolic Computation, 21(3):329– 336, March 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Foupouagnigni:1999:FOD

[FHR99] Mama Foupouagnigni, M. Norbert Hounkonnou, and André Ronveaux. The fourth-order difference equation satisfied by the associated orthogonal polynomials of the Δ-Laguerre-Hahn class. Journal of Symbolic Computation, 28(6):801–818, December 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0340/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0340/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0340/production/ref.

Fieker:2004:MRN

[Fie04] Claus Fieker. Minimizing representations over number fields. Journal of Symbolic Computation, 38(1):833–842, July 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fisher:1996:TSC

[Fis96] M. Fisher. Temporal semantics for concurrent MetateM. Journal of Symbolic Computation, 22(5–6):627–648, November/December 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fitch:1985:SAP

[Fit85] John P. Fitch. Solving algebraic problems with RE-DUCE. Journal of Symbolic Computation, 1(2):211– 228 (or 211–227??), June 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fitt:1989:SCH

[Fit89] A. D. Fitt. Symbolic computation of hyperbolicity regions for systems of two-phase flow conservation laws using Maple. Journal of Symbolic Computation, 8(3):305– 308, September 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fitzpatrick:1997:SMC

[Fit97] Patrick Fitzpatrick. Solving a multivariable congruence by change of term order. Journal of Symbolic Computation, 24(5):575–590 (or 575– 589??), November 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fukuda:2007:GGW

[FJLT07] K. Fukuda, A. N. Jensen, N. Lauritzen, and R. Thomas. The generic Gröbner walk. Journal of Symbolic Computation, 42(3):298–312, March 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Freese:1993:TRS

[FJN93] Ralph Freese, J. Ježek, and J. B. Nation. Term rewrite systems for lattice theory. Journal of Symbolic Computation, 16(3):279–288, September 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fushchich:1989:CAA

[FK89] W. I. Fushchich and V. V. Kornyak. Computer algebra application for determining Lie and Lie–Bäcklund symmetries of differential equations. Journal of Symbolic Computation, 7(6):611–619, June 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fruhbis-Kruger:2004:PSB

[FK04] Anne Frühbis-Krüger. Partial standard bases as a tool for studying families of singularities. Journal of Symbolic Computation, 38(4): 1191–1205, October 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fronk:2009:SSA

[FK09] Alexander Fronk and Britta Kehden. State space analysis of Petri nets with relationalgebraic methods. Journal of Symbolic Computation, 44 (1):15–47, January 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fruhbis-Kruger:2011:MCI

[FK11]  Anne Frühbis-Krüger.  A

modified coefficient ideal for use with the strict transform. Journal of Symbolic Computation, 46(5):550–560, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001744.

Fournier:1995:VMS

[FKM95] Robert Fournier, Norbert Kajler, and Bernard Mourrain. Visualization of mathematical surfaces: the IZIC server approach. Journal of Symbolic Computation, 19(1/2/3):159– 174 (or 159–173??), January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Fish:2010:BCL

[FKM10] W. Fish, J. D. Key, and E. Mwambene. Binary codes from the line graph of the ncube. Journal of Symbolic Computation, 45(7):800–812, July 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Floystad:2018:CFE

[FKO18] Gunnar Fløystad, Joe Kileel, and Giorgio Ottaviani. The Chow form of the essential variety in computer vision. Journal of Symbolic Computation, 86(??):97–119, May/ June 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711730038X.

Foupouagnigni:2013:CLC

[FKT13] M. Foupouagnigni, W. Koepf, and D. D. Tcheutia. Connection and linearization coefficients of the Askey– Wilson polynomials. Journal of Symbolic Computation, 53(??):96–118, June 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001927.

Foupouagnigni:2012:ROP

[FKTS12] M. Foupouagnigni, W. Koepf, D. D. Tcheutia, and P. Njionou Sadjang. Representations of qorthogonal polynomials. Journal of Symbolic Computation, 47(11):1347–1371, November 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112000375.

Fiore:2004:ORG

[FL04] Marcelo Fiore and Tom Leinster. An objective representation of the Gaussian integers. Journal of Symbolic Computation, 37(6):707–716, June 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Faugere:2011:ADS

[FL11] Jean-Charles Faugère and Ye Liang. Artificial discontinuities of single-parametric Gröbner bases. Journal of Symbolic Computation, 46(4):459–466, April 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001999.

Fu:2021:UDC

[FL21] Lei Fu and Wei Li. Unirational differential curves and differential rational parametrizations. Journal of Symbolic Computation, 104(??): 539–562, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300900.

Flener:2000:FSI

[FLB00] Pierre Flener, Kung-Kiu Lau, and Wolfgang Bibel. Foreword to the special issue on schemas. Journal of Symbolic Computation, 30(1):1–3, July 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0419;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0419/
pdf.

Faugere:2023:CCP

[FLE^{+}23] Jean-Charles Faugère, George Labahn, Mohab Safey El Din, Éric Schost, and Thi Xuan Vu. Computing critical points for invariant algebraic systems.

Journal of Symbolic Computation, 116(??):365–399, May/ June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000992.

Flener:2000:AFC

[FLOR00] Pierre Flener, Kung-Kiu Lau, Mario Ornaghi, and Julian Richardson. An abstract formalization of correct schemas for program synthesis. Journal of Symbolic Computation, 30(1):93–127, July 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0348;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0348/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0348/ref.

Foursov:2002:CAC

[FM02] Mikhail V. Foursov and Marc Moreno Maza. On computer-assisted classification of coupled integrable equations. Journal of Symbolic Computation, 33(5): 647–660, May 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Faugere:2017:SFA

[FM17] Jean-Charles Faugère and Chenqi Mou. Sparse FGLM algorithms. Journal of Symbolic Computation, 80 (part 3) (??):538–569, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300700.

Fevola:2024:HVR

[FM24] Claudia Fevola and Yelena Mandelshtam. Hirota varieties and rational nodal curves. Journal of Symbolic Computation, 120(??):Article 102239, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000536.

Falcon:2007:GBN

[FMM07] R. M. Falcón and J. Martín-Morales. Gröbner bases and the number of Latin squares related to autotopisms of order ≤ 7. Journal of Symbolic Computation, 42 (11–12):1142–1154, November/December 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Foote:2004:TDW

[FMR04] Richard Foote, Gagan Mirchandani, and Daniel Rockmore. Two-dimensional wreath product group-based image processing. Journal of Symbolic Computation, 37(2): 187–207, February 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Farran:2013:GBN

[FMTT13] J. I. Farrán, C. Munuera, G. Tizziotti, and F. Torres. Gröbner basis for normtrace codes. Journal of Symbolic Computation, 48 (??):54–63, January 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112000685.

Frenkel:2016:KPP

[FNU16] Elizaveta Frenkel, Andrey Nikolaev, and Alexander Ushakov. Knapsack problems in products of groups. Journal of Symbolic Computation, 74(??):96–108, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000541.

Ford:1987:CMO

[For87a] David J. Ford. The construction of maximal orders over a Dedekind domain. Journal of Symbolic Computation, 4(1):69–75, August 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fortenbacher:1987:AAU

[For87b] Albrecht Fortenbacher. An algebraic approach to unification under associativity and commutativity. Journal of Symbolic Computation, 3(3): 217–229, June 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fortune:2002:IEA

[For02] Steven Fortune. An iterated eigenvalue algorithm for approximating roots of univariate polynomials. Journal of Symbolic Computation, 33(5): 627–646, May 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Formisano:2000:GBA

[FOT00] Andrea Formisano, Eugenio G. Omodeo, and Marco Temperini. Goals and benchmarks for automated map reasoning. Journal of Symbolic Computation, 29 (2):259–297, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0362;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0362/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0362/ref.

Fox:2018:DLR

[Fox18]  Nathan Fox.  Discover-

ing linear-recurrent solutions to Hofstadter-like recurrences using symbolic computation. Journal of Symbolic Computation, 87(??): 99–126, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300627.

Faugere:2009:EAD

[FP09] Jean-Charles Faugère and Ludovic Perret. An efficient algorithm for decomposing multivariate polynomials and its applications to cryptography. Journal of Symbolic Computation, 44(12):1676– 1689, December 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Famelis:2004:SDR

[FPT04] I. Th. Famelis, S. N. Papakostas, and Ch. Tsitouras. Symbolic derivation of Runge–Kutta order conditions. Journal of Symbolic Computation, 37(3):311– 327, March 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Faugere:2009:F

[FR09] Jean-Charles Faugère and Fabrice Rouillier. Foreword. Journal of Symbolic Computation, 44(3):221, March 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fredet:2004:LDE

[Fre04] Anne Fredet. Linear differential equations in exponential extensions. Journal of Symbolic Computation, 38(2): 975–1002, August 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Freudenburg:2013:FIT

[Fre13] Gene Freudenburg. Foundations of invariant theory for the down operator. Journal of Symbolic Computation, 57(??):19–47, October 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000667.

Fribourg:1989:SRI

[Fri89] Laurent Fribourg. A strong restriction of the inductive completion procedure. Journal of Symbolic Computation, 8(3):253–276, September 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Felszeghy:2006:LGS

[FRR06] Bálint Felszeghy, Balázs Ráth, and Lajos Rónyai. The lex game and some applications. Journal of Symbolic Computation, 41(6):663–681, June 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fruhwirth:1996:TAC

[Frü96]  T. Frühwirth.  Temporal

annotated constraint logic programming. Journal of Symbolic Computation, 22(5– 6):555–584 (or 555–583??), November/December 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Flajolet:1995:CAL

[FS95] Philippe Flajolet and Bruno Salvy. Computer algebra libraries for combinatorial structures. Journal of Symbolic Computation, 20(5– 6):653–672 (or 653–671??), November/December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Feinsilver:1998:CRL

[FS98] Philip Feinsilver and René Schott. Computing representations of a Lie group via the universal enveloping algebra. Journal of Symbolic Computation, 26(3):329–338, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Farouki:2010:RRM

[FS10] Rida T. Farouki and Takis Sakkalis. Rational rotationminimizing frames on polynomial space curves of arbitrary degree. Journal of Symbolic Computation, 45(8): 844–856, August 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See corrigendum [FS13].

Farouki:2012:CCQ

[FS12] Rida T. Farouki and Takis Sakkalis. A complete classification of quintic space curves with rational rotationminimizing frames. Journal of Symbolic Computation, 47(2):214–226, February 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001477.

Farouki:2013:CSR

[FS13] Rida T. Farouki and Takis Sakkalis. Corrigendum to “Rational rotation-minimizing frames on polynomial space curves of arbitrary degree” [J. Symbolic Comput. 45(8) (2010) 844–856]. Journal of Symbolic Computation, 58 (??):99–102, November 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000825.

See [FS10].

Feng:2016:CIT

[FS16] Ruyong Feng and Li-Yong Shen. Computing the intersections of three conics according to their Jacobian curve. Journal of Symbolic Computation, 73(??):175–191, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000450.

Fieker:2023:CSF

[FS23] Claus Fieker and Nicole Sutherland. Computing splitting fields using Galois theory and other Galois constructions. Journal of Symbolic Computation, 116(??): 243–262, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000980.

Feng:2010:ACL

[FSW10a] Ruyong Feng, Michael F. Singer, and Min Wu. An algorithm to compute Liouvillian solutions of prime order linear difference–differential equations. Journal of Symbolic Computation, 45(3):306– 323, March 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Feng:2010:LSL

[FSW10b] Ruyong Feng, Michael F. Singer, and Min Wu. Liouvillian solutions of linear difference–differential equations. Journal of Symbolic Computation, 45(3):287– 305, March 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Filgueiras:1995:FMF

[FT95] Miguel Filgueiras and Ana Paula Tomás. A fast method for finding the basis of nonnegative solutions to a linear Diophantine equation. Journal of Symbolic Computation, 19(6):507–526, June 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Felty:1997:ITP

[FT97] Amy Felty and Laurent Théry.

Interactive theorem proving with temporal logic. Journal of Symbolic Computation, 23(4):367–398 (or 367– 397??), April 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fernando:2017:SPO

[FU17] José F. Fernando and Carlos Ueno. A short proof for the open quadrant problem. Journal of Symbolic Computation, 79 (part 1)(??): 57–64, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300839.

Fuchs:2000:CUC

[Fuc00a] Marc Fuchs. Controlled use of clausal lemmas in connection tableau calculi. Journal of Symbolic Computation, 29(2):299–341, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0363;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0363/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0363/ref.

Fuchs:2000:PST

[Fuc00b] Matthias Fuchs. Proofs as schemas and their heuristic use. Journal of Symbolic Computation, 30(1):37–61, July 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0349;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0349/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0349/ref.

Fukuda:2004:ZCM

[Fuk04] Komei Fukuda. From the zonotope construction to the Minkowski addition of convex polytopes. Journal of Symbolic Computation, 38 (4):1261–1272, October 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fulling:1990:AAR

[Ful90] S. A. Fulling. The analytic approach to recursion relations. Journal of Symbolic Computation, 9(1):73–86 (or 73–85??), January 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fontein:2014:PGL

[FW14] Felix Fontein and Pawel Wocjan. On the probability of generating a lattice. Journal of Symbolic Computation, 64(??):3–15, August 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001685.

Feng:2015:ACM

[FW15] Lei Feng and Weiping Wang.

An algorithm for computing mixed sums of products of Bernoulli polynomials and Euler polynomials. Journal of Symbolic Computation, 66 (??):84–97, January/February 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000182.

Folz:1987:WRD

[FZ87] H. G. Folz and H. G. Zimmer. What is the rank of the Demjanenko matrix? Journal of Symbolic Computation, 4(1):53–67, August 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gaal:1993:FAF

[Gaá93a] István Gaál. A fast algorithm for finding “small” solutions of F (x, y) = G(x, y) over imaginary quadratic fields. Journal of Symbolic Computation, 16(4):321–328, October 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gaal:1993:R

[Gaá93b] István Gaál. On the resolution of F (x, y) = G(x, y). Journal of Symbolic Computation, 16(3):295–304 (or 295– 303??), September 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gaal:1995:CEG

[Gaá95] István Gaál. Computing elements of given index in totally complex cyclic sextic fields. Journal of Symbolic Computation, 20(1):61–70 (or 61–69??), July 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gaal:2000:SIF

[Gaá00] István Gaál. Solving index form equations in fields of degree 9 with cubic subfields. Journal of Symbolic Computation, 30(2):181–193, August 2000. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0356;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0356/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0356/ref.

Gaal:2002:RRT

[Gaá02] István Gaál. On the resolution of resultant type equations. Journal of Symbolic Computation, 34(2):137–144, August 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gallier:1987:FAT

[Gal87] Jean H. Gallier. Fast algorithms for testing unsatisfiability of ground Horn clauses with equations. Journal of Symbolic Computation, 4(2): 233–254, October 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Galligo:2013:BTR

[Gal13a] André Galligo. Budan tables of real univariate polynomials. Journal of Symbolic Computation, 53(??):64–80, June 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001903.

Galligo:2013:DRP

[Gal13b]  André Galligo.  Deforma-

tion of roots of polynomials via fractional derivatives. Journal of Symbolic Computation, 52(??):35–50, May 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001289.

Galetto:2016:PWT

[Gal16]  Federico Galetto.  Propa-

gating weights of tori along free resolutions. Journal of Symbolic Computation, 74 (??):1–45, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000504.

Galetto:2022:FGC

[Gal22]  Federico Galetto.  Finite

group characters on free resolutions. Journal of Symbolic Computation, 113(??): 29–38, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000074.

Ganzinger:1991:CPC

[Gan91] Harald Ganzinger. A completion procedure for conditional equations. Journal of Symbolic Computation, 11(1– 2):51–82 (or 51–81??), January/February 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gao:2001:DCF

[Gao01] Shuhong Gao. On the deterministic complexity of factoring polynomials. Journal of Symbolic Computation, 31(1– 2):19–36, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1001;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.1001/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1001/ref.

Giesl:2002:MTP

[GAO02] Jürgen Giesl, Thomas Arts, and Enno Ohlebusch. Modular termination proofs for rewriting using dependency pairs. Journal of Symbolic Computation, 34(1):21–58, July 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gao:2003:IDR

[Gao03] Xiao-Shan Gao. Implicitization of differential rational parametric equations. Journal of Symbolic Computation, 36 (5):811–824, November 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gao:2018:P

[Gao18]  Xiao-Shan Gao.  Preface.

Journal of Symbolic Computation, 85(??):1–3, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300640.

Garvan:1995:RTE

[Gar95] Frank G. Garvan. Ramanujan’s theories of elliptic functions to alternative bases — a symbolic excursion. Journal of Symbolic Computation, 20(5–6):517–536, November/ December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Gasanova:2021:PMI

[Gas21] Oleksandra Gasanova. Powers of monomial ideals and the Ratliff–Rush operation. Journal of Symbolic Computation, 104(??):69–89, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300171.

Gatermann:2003:ASB

[Gat03] Karin Gatermann. Applications of SAGBI-bases in dynamics. Journal of Symbolic Computation, 35(5):543–575, May 2003. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Gaudry:2009:ICA

[Gau09] Pierrick Gaudry. Index calculus for abelian varieties of small dimension and the elliptic curve discrete logarithm problem. Journal of Symbolic Computation, 44(12):1690– 1702, December 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gao:1992:IRP

[GC92] Xiao Shan Gao and Shang-Ching Chou. Implicitization of rational parametric equations. Journal of Symbolic Computation, 14(5):459– 470, November 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gao:1993:ZST

[GC93] Xiao Shan Gao and Shang-Ching Chou. A zero structure theorem for differential parametric systems. Journal of Symbolic Computation, 16(6):585–596 (or 585– 595??), December 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gonzalez-Diaz:2005:STM

[GDR05] Rocio Gonzalez-Diaz and Pedro Real. Simplification techniques for maps in simplicial topology. Journal of Symbolic Computation, 40(4–5):1208–1224, October/November 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Guidolin:2021:CIM

[GDRV21] Andrea Guidolin, Jose Divasón, Ana Romero, and Francesco Vaccarino. Computing invariants for multipersistence via spectral systems and effective homology. Journal of Symbolic Computation, 104(??):724–753, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301097.

Gebhardt:2002:ECI

[Geb02] Volker Gebhardt. Efficient collection in infinite polycyclic groups. Journal of Symbolic Computation, 34(3):213– 228, September 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

GrafvBothmer:2005:NFR

[GEL05] Hans-Christian Graf v. Bothmer, Cord Erdenberger, and Katharina Ludwig. A new family of rational surfaces in P^{4}. Journal of Symbolic Computation, 39(1):51–60, January 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gemignani:1994:SHS

[Gem94] Luca Gemignani. Solving Hankel systems over the integers. Journal of Symbolic Computation, 18(6):573– 584, December 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Genovese:2007:IAB

[Gen07] Giulio Genovese. Improving the algorithms of Berlekamp and Niederreiter for factoring polynomials over finite fields. Journal of Symbolic Computation, 42(1–2):159–177, January/February 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Genzmer:2022:NMU

[Gen22]  Yohann Genzmer.  Num-

ber of moduli for a union of smooth curves in (C_{2}, 0). Journal of Symbolic Computation, 113(??):148–170, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000207.

Gerritzen:2006:TPN

[Ger06] Lothar Gerritzen. Tree polynomials and non-associative Gröbner bases. Journal of Symbolic Computation, 41 (3–4):297–316, March/April 2006. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-

855X (electronic).

Gerling:2019:BCP

[Ger19]  Melanie Gerling.  Bivari-

ate chromatic polynomials in computer algebra. Journal of Symbolic Computation, 93(??):183–199, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300701.

Gessel:1992:SBN

[Ges92] Ira M. Gessel. Super ballot numbers. Journal of Symbolic Computation, 14(2–3): 179–194, August/September 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gessel:1995:FIW

[Ges95] Ira M. Gessel. Finding identities with the WZ method. Journal of Symbolic Computation, 20(5–6):537–566, November/December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Geser:1997:OTU

[Ges97]  Alfons Geser.  Omega-

termination is undecidable for totally terminating term rewriting systems. Journal of Symbolic Computation, 23(4): 399–412 (or 399–411??), April 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gatermann:2005:TIG

[GES05] Karin Gatermann, Markus Eiswirth, and Anke Sensse. Toric ideals and graph theory to analyze Hopf bifurcations in mass action systems. Journal of Symbolic Computation, 40(6):1361–1382, December 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Garvan:1992:PTP

[GG92] Frank G. Garvan and Gaston H. Gonnet. A proof of the two parameter q-cases of the Macdonald-Morris constant term root system conjecture for S(F_{4}) and S(F_{4})^{∨} via Zeilberger’s method. Journal of Symbolic Computation, 14 (2–3):141–178 (or 141–177??), August/September 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gatermann:1999:GBI

[GG99] Karin Gatermann and Frédéric Guyard. Gröbner bases, invariant theory and equivariant dynamics. Journal of Symbolic Computation, 28 (1–2):275–302, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.

                                        1998.0277;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1998.0277/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1998.0277/ref.

Galan-Garcia:2013:TIM

[GGAVRC13] José Luis Galán-García, Gabriel Aguilera-Venegas, and Pedro Rodríguez-Cielos. Technology integration into mathematics education and applications. Journal of Symbolic Computation, 61–62(??):66–69, ???? 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001302.

Giorgi:2023:PMP

[GGdC23] Pascal Giorgi, Bruno Grenet, and Armelle Perret du Cray. Polynomial modular product verification and its implications. Journal of Symbolic Computation, 116(??): 98–129, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000773.

Garcia:2013:CAM

[GGdR^{+}13] Alfonsa García, Francisco García, Ángel Martín del Rey, Gerardo Rodríguez, and Agustín de la Villa. Changing assessment methods: New rules, new roles. Journal of Symbolic Computation, 61– 62(??):70–84, ???? 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001314.

Greuet:2012:GOP

[GGEZ12] Aurélien Greuet, Feng Guo, Mohab Safey El Din, and Lihong Zhi. Global optimization of polynomials restricted to a smooth variety using sums of squares. Journal of Symbolic Computation, 47(5):503–518, May 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001957.

Garcia:2006:NCM

[GGG06] Isaac A. García, Jaume Giné, and Maite Grau. A necessary condition in the monodromy problem for analytic differential equations on the plane. Journal of Symbolic Computation, 41(9):943–958, September 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Grosslinger:2006:QEA

[GGL06] Armin Größlinger, Martin Griebl, and Christian Lengauer. Quantifier elimination in automatic loop parallelization. Journal of Symbolic Computation, 41(11):1206– 1221, November 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gebhardt:2010:SCP

[GGM10] Volker Gebhardt and Juan González-Meneses. Solving the conjugacy problem in Garside groups by cyclic sliding. Journal of Symbolic Computation, 45(6):629–656, June 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Garcia-Garcia:2019:DCS

[GGMAMF19] J. I. García-García, D. Marín-Aragón, and M. A. Moreno-Frías. On divisor-closed submonoids and minimal distances in finitely generated monoids. Journal of Symbolic Computation, 93(??): 230–245, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300725.

Garcia-Garcia:2013:DSA

[GGMFVT13] J. I. García-García, M. A. Moreno-Frías, and A. Vigneron-Tenorio. On decomposable semigroups and applications. Journal of Symbolic Computation, 58(??):103–116, November 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000722.

Gascon:2010:CUO

[GGSST10] Adrià Gascón, Guillem Godoy, Manfred Schmidt-Schauß, and Ashish Tiwari. Context unification with one context variable. Journal of Symbolic Computation, 45(2):173– 193, February 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Goktas:1997:SCC

[GH97] Ünal Göktaş and Willy Hereman. Symbolic computation of conserved densities for systems of nonlinear evolution equations. Journal of Symbolic Computation, 24(5):591– 621, November 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gatermann:2002:FSP

[GH02] Karin Gatermann and Birkett Huber. A family of sparse polynomial systems arising in chemical reaction systems. Journal of Symbolic Computation, 33(3):275–305, March 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0512;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0512/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0512/ref.

Gann:2005:PBD

[GH05a] Sebastian Gann and Herwig Hauser. Perfect bases for differential equations. Journal of Symbolic Computation, 40(2): 979–997, August 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gatermann:2005:CAB

[GH05b] Karin Gatermann and Serkan Hosten. Computational algebra for bifurcation theory. Journal of Symbolic Computation, 40(4–5):1180–1207, October/November 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gao:2012:CSA

[GH12] Xiao-Shan Gao and Zhenyu Huang. Characteristic set algorithms for equation solving in finite fields. Journal of Symbolic Computation, 47(6):655–679, June 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002185.

Guimaraes:1992:DRT

[GHC92] J. E. F. Guimaraes, G. R. Heppler, and S. R. Czapor. On the derivation and reduction of C^{1} trigonometric basis functions using Maple. Journal of Symbolic Computation, 13(6):643–676 (or 643–675??), June 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Giusti:2000:PNM

[GHL^{+}00] Marc Giusti, Klemens Hägele, Grégoire Lecerf, Joël Marchand, and Bruno Salvy. The Projective Noether Maple Package: Computing the dimension of a projective variety. Journal of Symbolic Computation, 30(3): 291–307, September 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0369;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0369/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0369/ref.

Giesbrecht:2016:FLP

[GHL16] Mark Giesbrecht, Albert Heinle, and Viktor Levandovskyy. Factoring linear partial differential operators in n variables. Journal of Symbolic Computation, 75(??): 127–148, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001108.

Giesbrecht:2021:CLR

[GHL21a] Mark Giesbrecht, Joseph Haraldson, and George Labahn.

Computing lower rank approximations of matrix polynomials. Journal of Symbolic Computation, 98(??):225–245, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711930077X.

Giesbrecht:2021:CNN

[GHL21b] Mark Giesbrecht, Joseph Haraldson, and George Labahn. Computing nearby non-trivial Smith forms. Journal of Symbolic Computation, 102(??): 304–327, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301312.

Giesbrecht:2021:EIL

[GHLZ21] Mark Giesbrecht, Hui Huang, George Labahn, and Eugene Zima. Efficient qinteger linear decomposition of multivariate polynomials. Journal of Symbolic Computation, 107(??):122–144, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000110.

Giesbrecht:2022:ERC

[GHLZ22] Mark Giesbrecht, Hui Huang, George Labahn, and Eugene Zima. Efficient rational creative telescoping. Journal of Symbolic Computation, 109(??):57–87, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000535.

Gerdt:2013:IBA

[GHMA13] Vladimir P. Gerdt, Amir Hashemi, and Benyamin M.- Alizadeh. Involutive bases algorithm incorporating F_{5} criterion. Journal of Symbolic Computation, 59(??):1– 20, December 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001090.

Green:2001:CHS

[GHS01] Edward L. Green, Lenwood S.

Heath, and Craig A. Struble. Constructing homomorphism spaces and endomorphism rings. Journal of Symbolic Computation, 32 (1–2):101–117, July/August 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0454;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0454/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0454/ref.

Georgieva:2003:HGF

[GHS03] Lilia Georgieva, Ullrich Hustadt, and Renate A. Schmidt. Hyperresolution for guarded formulae. Journal of Symbolic Computation, 36(1–2): 163–192, July/August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Guo:2008:PHI

[GHS08] Qiang-Hui Guo, Qing-Hu Hou, and Lisa H. Sun. Proving hypergeometric identities by numerical verifications. Journal of Symbolic Computation, 43(12):895–907, December 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gao:2017:BDI

[GHY17] Xiao-Shan Gao, Zhang Huang, and Chun-Ming Yuan. Binomial difference ideals. Journal of Symbolic Computation, 80 (part 3)(??):665–706, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300761.

Gibert:1987:FPC

[Gib87] Jacek Gibert. Functional programming with combinators. Journal of Symbolic Computation, 4(3):269–293, December 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Giesbrecht:1998:FSP

[Gie98] M. Giesbrecht. Factoring in skew-polynomial rings over finite fields. Journal of Symbolic Computation, 26(4):463– 486, October 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gutierrez:2014:RST

[GIJ14] Jaime Gutierrez, Álvar Ibeas, and Antoine Joux. Recovering a sum of two squares decomposition. Journal of Symbolic Computation, 64 (??):16–21, August 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001697.

Giese:2010:I

[GIK10] Martin Giese, Andrew Ireland, and Laura Kovács. Introduction. Journal of Symbolic Computation, 45(11): 1097–1100, November 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gateva-Ivanova:1988:RPA

[GIL88] Tatiana Gateva-Ivanova and Victor Latyshev. On recognisable properties of associative algebras. Journal of Symbolic Computation, 6(2–3):371–388, October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Gateva-Ivanova:2007:STS

[GIM07] Tatiana Gateva-Ivanova and Shahn Majid. Set-theoretic solutions of the Yang–Baxter equation, graphs and computations. Journal of Symbolic Computation, 42(11– 12):1079–1112, November/ December 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Girstmair:2021:RRS

[Gir21] Kurt Girstmair. Reducing radicals in the spirit of Euclid. Journal of Symbolic Computation, 104(??):356–365, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300742.

Giusti:1988:CDT

[Giu88] Marc Giusti. Combinatorial dimension theory of algebraic varieties. Journal of Symbolic Computation, 6 (2–3):249–266 (or 249–265??), October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Gonzalez-Jimenez:2013:MRT

[GJT13] Enrique González-Jiménez and José M. Tornero. Markoff– Rosenberger triples in arithmetic progression. Journal of Symbolic Computation, 53(??):53–63, June 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001897.

Gusynin:1994:SCD

[GK94] V. P. Gusynin and V. V. Kornyak. Symbolic computation of DeWitt-Seeley-Gilkey coefficients on curved manifolds. Journal of Symbolic Computation, 17(3):283–294, March 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gerdt:1996:CFP

[GK96a] Vladimir P. Gerdt and Vladimir V.

Kornyak. Construction of finitely presented Lie algebras and superalgebras. Journal of Symbolic Computation, 21(3):337–350 (or 337– 349??), March 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gladitz:1996:SMI

[GK96b] K. Gladitz and H. Kuchen.

Shared memory implementation of the gamma-operation. Journal of Symbolic Computation, 21(4/5/6):577–592 (or 577–591??), April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Geissler:2000:GGC

[GK00] Katharina Geissler and Jürgen Klüners. Galois group computation for rational polynomials. Journal of Symbolic Computation, 30(6): 653–674, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0377;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.0377/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0377/ref.

Gao:2012:BIW

[GK12a] Xiao-Shan Gao and Deepak Kapur. A brief introduction to Wen-Tsun Wu’s academic career. Journal of Symbolic Computation, 47(6):586–588, June 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002148.

Gao:2012:P

[GK12b] Xiao-Shan Gao and Deepak Kapur. Preface. Journal of Symbolic Computation, 47(6):583–585, June 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002136.

Grigoriev:2016:CTS

[GK16] Dima Grigoriev and Gleb Koshevoy. Complexity of tropical Schur polynomials. Journal of Symbolic Computation, 74(??):46–54, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000516.

Gauthier:2018:ACA

[GK18] Thibault Gauthier and Cezary Kaliszyk. Aligning concepts across proof assistant libraries. Journal of Symbolic Computation, 90(??):89–123, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300348.

Ginory:2021:WCI

[GK21] Alejandro Ginory and Jongwon Kim. Weingarten calculus and the IntHaar package for integrals over compact matrix groups. Journal of Symbolic Computation, 103(??): 178–200, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301518.

Gao:2004:DDD

[GKL04] Shuhong Gao, Erich Kaltofen, and Alan G. B. Lauder. Deterministic distinct-degree factorization of polynomials over finite fields. Journal of Symbolic Computation, 38(6): 1461–1470, December 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Garbey:1991:UMA

[GKLM91] Marc Garbey, Hans G. Kaper, Gary K. Leaf, and Bernard J. Matkowsky. Using MAPLE for the analysis of bifurcation phenomena in condensedphase surface combustion. Journal of Symbolic Computation, 12(1):89–114 (or 89– 113??), July 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gottliebsen:2005:HVC

[GKM05] Hanne Gottliebsen, Tom Kelsey, and Ursula Martin. Hidden verification for computational mathematics. Journal of Symbolic Computation, 39(5):539–567, May 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Golubitsky:2008:BRG

[GKMO08] Oleg Golubitsky, Marina Kondratieva, Marc Moreno Maza, and Alexey Ovchinnikov. A bound for the Rosenfeld– Gröbner algorithm. Journal of Symbolic Computation, 43(8): 582–610, August 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gaar:2021:TCP

[GKMW21] Elisabeth Gaar, Daniel Krenn, Susan Margulies, and Angelika Wiegele. Towards a computational proof of Vizing’s conjecture using semidefinite programming and sumsof-squares. Journal of Symbolic Computation, 107(??): 67–105, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000092.

Golubitsky:2009:ATD

[GKO09] Oleg Golubitsky, Marina Kondratieva, and Alexey Ovchinnikov. Algebraic transformation of differential characteristic decompositions from one ranking to another. Journal of Symbolic Computation, 44 (4):333–357, April 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Garcia:2023:LPE

[GKR23] Ronaldo Garcia, Jair Koiller, and Dan Reznik. Loci of 3periodics in an Elliptic Billiard: Why so many ellipses? Journal of Symbolic Computation, 114(??): 336–358, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000505.

Gaubert:2012:TLF

[GKS12] Stéphane Gaubert, Ricardo D. Katz, and Sergei Sergeev. Tropical linear-fractional programming and parametric mean payoff games. Journal of Symbolic Computation, 47(12):1447–1478, December 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002525.

Giesbrecht:2003:ACS

[GKsL03] Mark Giesbrecht, Erich Kaltofen, and Wen shin Lee. Algorithms for computing sparsest shifts of polynomials in power, Chebyshev, and Pochhammer bases. Journal of Symbolic Computation, 36(3–4): 401–424, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gray:1998:DIM

[GKW98] S. Gray, N. Kajler, and P. S. Wang. Design and implementation of MP, a protocol for efficient exchange of mathematical expressions. Journal of Symbolic Computation, 25(2): 213–238, February 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Grossman:1992:SCD

[GL92] Robert Grossman and Richard G. Larson. Symbolic computation of derivations using labelled trees. Journal of Symbolic Computation, 13(5):511– 524 (or 511–523??), May 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gramlich:2005:RSR

[GL05] Bernhard Gramlich and Salvador Lucas. Reduction strategies in rewriting and programming. Journal of Symbolic Computation, 40(1): 745–747, July 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Glasby:1988:CNF

[Gla88a] S. P. Glasby. Constructing normalisers in finite soluble groups. Journal of Symbolic Computation, 5(3): 285–294, June 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Glasby:1988:ISF

[Gla88b] S. P. Glasby. Intersecting subgroups of finite soluble groups. Journal of Symbolic Computation, 5(3):295–302 (or 295– 301??), June 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Garcia:2019:APN

[GLLdR19] Belén García, Jaume Llibre, Antón Lombardero, and Jesús S. Pérez del Río. An algorithm for providing the normal forms of spatial quasihomogeneous polynomial differential systems. Journal of Symbolic Computation, 95(??):1–25, November/December 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300981.

Garcia:2021:AIQ

[GLLdR21] Belén García, Jaume Llibre, Antón Lombardero, and Jesús S. Pérez del Río. Analytic integrability of quasihomogeneous systems via the Yoshida method. Journal of Symbolic Computation, 104 (??):960–980, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301188.

Guerrini:2023:SRF

[GLLZ23] Eleonora Guerrini, Kamel Lairedj, Romain Lebreton, and Ilaria Zappatore. Simultaneous rational function reconstruction with errors: Handling multiplicities and poles. Journal of Symbolic Computation, 116(??):345–364, May/ June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122001043.

Greuel:2010:NR

[GLS10] Gert-Martin Greuel, Santiago Laplagne, and Frank Seelisch. Normalization of rings. Journal of Symbolic Computation, 45(9):887–901, September 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Giesbrecht:2009:SNS

[GLsL09] Mark Giesbrecht, George Labahn, and Wen shin Lee. Symbolic–numeric sparse interpolation of multivariate polynomials. Journal of Symbolic Computation, 44(8):943– 959, August 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gallet:2021:RRR

[GLSV21] Matteo Gallet, Niels Lubbes, Josef Schicho, and Jan Vrsek. Reconstruction of rational ruled surfaces from their silhouettes. Journal of Symbolic Computation, 104(??): 366–380, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300833.

Gao:1999:FAI

[GLW99] Tangan Gao, T. Y. Li, and Xiaoshen Wang. Finding all isolated zeros of polynomial systems in C_{n} via stable mixed volumes. Journal of Symbolic Computation, 28 (1–2):187–212, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0272;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1998.0272/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0272/ref;   http://
www.idealibrary.com/links/
doi/10.1006/jsco.1999.0273;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0273/pdf;   http://
www.idealibrary.com/links/
doi/10.1006/jsco.1999.0273/
ref.

Gao:2009:CSMa

[GLY09] Xiao-Shan Gao, Yong Luo, and Chunming Yuan. A characteristic set method for ordinary difference polynomial systems. Journal of Symbolic Computation, 44(3):242– 260, March 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gebauer:1988:IBA

[GM88] Rüdiger Gebauer and H. Michael Möller. On an installation of Buchberger’s algorithm. Journal of Symbolic Computation, 6(2–3):275–286, October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Gaspar:2013:PCS

[GMF13] Miguel Belbut Gaspar and Nelson Martins-Ferreira. A procedure for computing the symmetric difference of regions defined by polygonal curves. Journal of Symbolic Computation, 61– 62(??):53–65, ???? 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001272.

Garcia-Marco:2021:RAS

[GMKP21] Ignacio García-Marco, Pascal Koiran, and Timothée Pecatte. Reconstruction algorithms for sums of affine powers. Journal of Symbolic Computation, 98(??):284–318, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300811.

Glenn:2017:MTC

[GMMM17] Paul Glenn, William W.

Menasco, Kayla Morrell, and Matthew J. Morse. MICC: a tool for computing short distances in the curve complex. Journal of Symbolic Computation, 78(??): 115–132, January/February 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300116.

Gorgen:2022:STM

[GMN22] Christiane Görgen, Aida Maraj, and Lisa Nicklasson. Staged tree models with toric structure. Journal of Symbolic Computation, 113(??):242– 268, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000268.

Gray:2013:HBF

[GMP13] R. Gray, A. Malheiro, and S. J. Pride. Homotopy bases and finite derivation type for Schützenberger groups of monoids. Journal of Symbolic Computation, 50(??): 50–78, March 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001058.

Gorla:2022:SBC

[GMP22] Elisa Gorla, Daniela Mueller, and Christophe Petit. Stronger bounds on the cost of computing Gröbner bases for HFE systems. Journal of Symbolic Computation, 109(??): 386–398, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300663.

Gosselin:2009:DBP

[GMS09] Clément M. Gosselin, Brian Moore, and Josef Schicho. Dynamic balancing of planar mechanisms using toric geometry. Journal of Symbolic Computation, 44(9): 1346–1358, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gouveia:2023:GNR

[GMW23] João Gouveia, Antonio Macchia, and Amy Wiebe. General non-realizability certificates for spheres with linear programming. Journal of Symbolic Computation, 114(??):172–192, January/February 2023. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000335.

Godoy:2004:SCB

[GN04] Guillem Godoy and Robert Nieuwenhuis. Superposition with completely builtin Abelian groups. Journal of Symbolic Computation, 37 (1):1–33, January 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Guardia:2012:SFL

[GNP12] Jordi Guàrdia, Enric Nart, and Sebastian Pauli. Singlefactor lifting and factorization of polynomials over local fields. Journal of Symbolic Computation, 47(11): 1318–1346, November 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000363.

Gallet:2017:LL

[GNS17] Matteo Gallet, Georg Nawratil, and Josef Schicho. Liaison linkages. Journal of Symbolic Computation, 79 (part 1)(??): 65–98, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300852.

Gollan:1990:OCS

[GO90] H. W. Gollan and T. W. Ostermann. Operation of class sums on permutation modules. Journal of Symbolic Computation, 9(1):39–48 (or 39–47??), January 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gentili:1991:RGS

[GO91] Graziano Gentili and Michael A. O’Connor. Rational geometry in space. Journal of Symbolic Computation, 11(3):213– 230 (or 213–229??), March 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gentili:2000:RGC

[GO00] Graziano Gentili and Michael A.

O’Connor. On rational geometry of conic sections. Journal of Symbolic Computation, 29(3):459–470, March 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0332;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0332/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0332/ref.

Gobel:1995:CBR

[Göb95] Manfred Göbel. Computing bases for rings of permutationinvariant polynomials. Journal of Symbolic Computation, 19(4):285–292 (or 285– 291??), April 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gobel:1998:CDS

[Göb98] Manfred Göbel. A constructive description of SAGBI bases for polynomial invariants of permutation groups. Journal of Symbolic Computation, 26(3):261–272, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gollan:2001:NEP

[Gol01] Holger W. Gollan. A new existence proof for Ly, the sporadic simple group of R. Lyons. Journal of Symbolic Computation, 31(1–2): 203–209, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.1010;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.1010/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.1010/ref.

Golubitsky:2006:GFU

[Gol06] Oleg Golubitsky. Gröbner fan and universal characteristic sets of prime differential ideals. Journal of Symbolic Computation, 41(10):1091– 1104, October 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Golubitsky:2008:UCD

[Gol08] Oleg Golubitsky. Universal characteristic decomposition of radical differential ideals. Journal of Symbolic Computation, 43(1):27–45, January 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gonzalez:2017:RRM

[Gon17] José Martín Méndez González. Revealing regions of multiple steady states in heterogeneous catalytic chemical reaction networks using Gröbner basis. Journal of Symbolic Computation, 80 (part 3) (??):521–537, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300694.

Gustavson:2018:NOB

[GOP18] Richard Gustavson, Alexey Ovchinnikov, and Gleb Pogudin. New order bounds in differential elimination algorithms. Journal of Symbolic Computation, 85(??): 128–147, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300706.

Granger:2005:TCA

[GOT05] Michel Granger, Toshinori Oaku, and Nobuki Takayama. Tangent cone algorithm for homogenized differential operators. Journal of Symbolic Computation, 39(3–4): 417–431, March/April 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gaal:1996:RIF

[GP96a] István Gaál and Michael Pohst. On the resolution of index form equations in sextic fields with an imaginary quadratic subfield. Journal of Symbolic Computation, 22(4): 425–434, October 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gagne:1996:NST

[GP96b] J.-R. Gagne and J. Plaice. A non-standard temporal deductive database system. Journal of Symbolic Computation, 22(5–6):649–664, November/ December 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Giusti:2003:P

[GP03] Marc Giusti and Luis M. Pardo. Preface. Journal of Symbolic Computation, 36(3– 4):287, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Giesbrecht:2012:HRI

[GP12] Mark Giesbrecht and Daniel Panario. In honour of the research and influence of Joachim von zur Gathen at

60. Journal of Symbolic Computation, 47(4):355–357, April 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001362.

Gruber:2013:CSQ

[GP13] David Gruber and Martin Peternell. Conchoid surfaces of quadrics. Journal of Symbolic Computation, 59 (??):36–53, December 2013.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000989.

Guo:2020:TDC

[GP20] Feng Guo and Tiêo’n Pha.m.

On types of degenerate critical points of real polynomial functions. Journal of Symbolic Computation, 99(??): 108–126, July/August 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300306.

Geiger:2024:CTB

[GP24] Alheydis Geiger and Marta Panizzut. Computing tropical bitangents to smooth quartic curves in polymake. Journal of Symbolic Computation, 120(??):Article 102225, January/February 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000329.

Gomez-Perez:2016:CCT

[GPGO16] Domingo Gómez-Pérez, Jaime Gutierrez, and Alina Ostafe. Common composites of triangular polynomial systems and hash functions. Journal of Symbolic Computation, 72(??):182–195, January/February 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000140.

Gomez-Perez:2020:CEC

[GPMS20] Domingo Gómez-Pérez, László Mérai, and Igor E. Shparlinski. On the complexity of exact counting of dynamically irreducible polynomials. Journal of Symbolic Computation, 99(??):231–241, July/August 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300549.

Gaal:1993:RIF

[GPP93] István Gaál, Attila Pethö, and Michael Pohst. On the resolution of index form equations in quartic number fields. Journal of Symbolic Computation, 16(6):563–584, December 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Galligo:2009:EMA

[GPS09] André Galligo, Luis Miguel Pardo, and Josef Schicho. Effective methods in algebraic geometry. Journal of Symbolic Computation, 44(9): 1087–1088, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Geuvers:2002:CAH

[GPWZ02] Herman Geuvers, Randy Pollack, Freek Wiedijk, and Jan Zwanenburg. A constructive algebraic hierarchy in Coq. Journal of Symbolic Computation, 34(4):271–286, October 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gutierrez:1998:ASS

[GR98] Jaime Gutierrez and Tomas Recio. Advances on the simplification of sine-cosine equations. Journal of Symbolic Computation, 26(1):31–70, July 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Griess:2001:EPP

[GR01] Robert L. Griess, Jr. and A. J. E. Ryba. Embeddings of PSL(2, 41) and PSL(2, 49) in E_{8}(C). Journal of Symbolic Computation, 31(1–2): 211–227, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1000;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.1000/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1000/ref.

Galligo:2002:IDC

[GR02] André Galligo and David Rupprecht. Irreducible decomposition of curves. Journal of Symbolic Computation, 33(5):661–677, May 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Genet:2010:EAT

[GR10] Thomas Genet and Vlad Rusu. Equational approximations for tree automata completion. Journal of Symbolic Computation, 45(5):574–597, May 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Giesbrecht:2011:DLP

[GR11] Mark Giesbrecht and Daniel S. Roche. Detecting lacunary perfect powers and computing their roots. Journal of Symbolic Computation, 46(11): 1242–1259, November 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001180.

Gallardo:2012:AUP

[GR12] Luis H. Gallardo and Olivier Rahavandrainy. All unitary perfect polynomials over F_{2} with at most four distinct irreducible factors. Journal of Symbolic Computation, 47(4):492–502, April 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001441.

Grigoriev:2022:TVJ

[GR22] Dima Grigoriev and Danylo Radchenko. On a tropical version of the Jacobian conjecture. Journal of Symbolic Computation, 109(??): 399–403, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300675.

Grabe:1993:LP

[Grä93] Hans-Gert Gräbe. On lucky primes. Journal of Symbolic Computation, 15(2):199– 210 (or 199–209??), February 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Grabe:1995:ALA

[Grä95] Hans-Gert Gräbe. Algorithms in local algebra. Journal of Symbolic Computation, 19(6): 545–558 (or 545–557??), June 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Greenhill:1995:TEC

[Gre95] Catherine S. Greenhill. Theoretical and experimental comparison of efficiency of finite field extensions. Journal of Symbolic Computation, 20(4): 419–430 (or 419–429??), October 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Green:2000:MBG

[Gre00a] Edward L. Green. Multiplicative bases, gröbner bases, and right gröbner bases. Journal of Symbolic Computation, 29(4–5):601–623, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0324;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0324/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0324/ref.

Greuel:2000:CAA

[Gre00b] Gert-Martin Greuel. Computer algebra and algebraic geometry — achievements and perspectives. Journal of Symbolic Computation, 30 (3):253–289, September 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0362;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0362/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0362/ref.

Grenet:2016:BDF

[Gre16] Bruno Grenet. Boundeddegree factors of lacunary multivariate polynomials. Journal of Symbolic Computation, 75(??):171–192, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001121.

Grigorev:1988:CDT

[Gri88] D. Yu. Grigor′ev. Complexity of deciding Tarski algebra. Journal of Symbolic Computation, 5(1–2):65–108, February/April 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Grigorev:1990:CFC

[Gri90] D. Yu. Grigor′ev. Complexity of factoring and calculating the GCD of linear ordinary differential operators. Journal of Symbolic Computation, 10(1):7–37, July 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Grigoriev:2020:DTR

[Gri20] Dima Grigoriev. Decomposing tropical rational functions. Journal of Symbolic Computation, 101(??): 61–72, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300604.

Grigoriev:2022:ETH

[Gri22] Dima Grigoriev. Entropy of tropical holonomic sequences. Journal of Symbolic Computation, 108(??):91– 97, January/February 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000407.

Grigoriev:2023:TNP

[Gri23] Dima Grigoriev. Tropical Newton–Puiseux polynomials II. Journal of Symbolic Computation, 115(??): 316–319, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000827.

Gutierrez:2002:MRF

[GRS02] Jaime Gutierrez, Rosario Rubio, and David Sevilla. On multivariate rational function decomposition. Journal of Symbolic Computation, 33(5): 545–562, May 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gitler:2017:CRG

[GRV17] Isidoro Gitler, Enrique Reyes, and Juan A. Vega. CIO and ring graphs: Deficiency and testing. Journal of Symbolic Computation, 79 (part 2)(??):249–268, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711600016X.

Gorlach:2016:DPM

[GRW16] Paul Görlach, Cordian Riener, and Tillmann Weißer. Deciding positivity of multisymmetric polynomials. Journal of Symbolic Computation, 74(??):603–616, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711500098X.

Gallier:1989:HOU

[GS89] J. Gallier and W. Snyder.

Higher-order unification revisited: complete sets of transformations. Journal of Symbolic Computation, 8(1– 2):101–140, ???? 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Glasby:1990:CIN

[GS90] S. P. Glasby and Michael C.

Slattery. Computing intersections and normalizers in soluble groups. Journal of Symbolic Computation, 9(5–6):637–652 (or 637– 651??), May/June 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Gutierrez:1998:RGB

[GS98] J. Gutierrez and R. R. San Miguel. Reduced Gröbner bases under composition. Journal of Symbolic Computation, 26(4):433–444, October 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Giesbrecht:2002:CRF

[GS02] Mark Giesbrecht and Arne Storjohann. Computing rational forms of integer matrices. Journal of Symbolic Computation, 34(3):157–172, September 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gil:2003:CSN

[GS03] Amparo Gil and Javier Segura. A combined symbolic and numerical algorithm for the computation of zeros of orthogonal polynomials and special functions. Journal of Symbolic Computation, 35(5): 465–485, May 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gasparim:2005:CIN

[GS05] Elizabeth Gasparim and Irena Swanson. Computing instanton numbers of curve singularities. Journal of Symbolic Computation, 40(2):965– 978, August 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gutierrez:2006:CUF

[GS06] Jaime Gutierrez and David Sevilla. Computation of unirational fields. Journal of Symbolic Computation, 41(11): 1222–1244, November 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Galligo:2007:GPB

[GS07a] A. Galligo and M. Stillman. On the geometry of parametrized bicubic surfaces. Journal of Symbolic Computation, 42(1–2):136–158, January/February 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Green:2007:AAR

[GS07b] Edward L. Green and Øyvind Solberg. An algorithmic approach to resolutions. Journal of Symbolic Computation, 42(11–12):1012–1033, November/December 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gaudry:2012:GPC

[GS12] Pierrick Gaudry and Éric Schost. Genus 2 point counting over prime fields. Journal of Symbolic Computation, 47(4):368–400, April 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001386.

Gustavson:2018:EBC

[GS18] Richard Gustavson and Omar León Sánchez. Effective bounds for the consistency of differential equations. Journal of Symbolic Computation, 89(??): 41–72, November/December 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117301104.

Ghorbal:2022:CPI

[GS22] Khalil Ghorbal and Andrew Sogokon. Characterizing positively invariant sets: Inductive and topological methods. Journal of Symbolic Computation, 113(??):1–28, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000049.

Gorla:2023:MET

[GS23] Elisa Gorla and Flavio Salizzoni. MacWilliams’ Extension Theorem for rankmetric codes. Journal of Symbolic Computation, 122(??): Article 102263, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000779.

Gaar:2024:SSC

[GS24] Elisabeth Gaar and Melanie Siebenhofer. Sum-of-squares certificates for Vizing’s conjecture via determining Gröbner bases. Journal of Symbolic Computation, 120(??):Article 102236, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000500.

Gawlitza:2012:AIM

[GSA^{+}12] Thomas Martin Gawlitza, Helmut Seidl, Assalé Adjé, Stéphane Gaubert, and Éric Goubault. Abstract interpretation meets convex optimization. Journal of Symbolic Computation, 47(12): 1416–1446, December 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002513.

Gonzalez-Sanchez:2012:ADP

[GSHPBS12] Jon González-Sánchez, Michael Harrison, Irene Polo-Blanco, and Josef Schicho. Algorithms for Del Pezzo surfaces of degree 5 (construction, parametrization). Journal of Symbolic Computation, 47(3):342–353, March 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001945.

Gonzalez-Sanchez:2017:CAR

[GSPB17] Jon González-Sánchez and Irene Polo-Blanco. Construction algorithms for rational cubic surfaces. Journal of Symbolic Computation, 79 (part 2)(??):309– 326, March/April 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116000195.

Garcia:2005:AGB

[GSS05] Luis David Garcia, Michael Stillman, and Bernd Sturmfels. Algebraic geometry of Bayesian networks. Journal of Symbolic Computation, 39 (3–4):331–355, March/April 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ghilardi:2010:SIA

[GSSST10] Silvio Ghilardi, Viorica Sofronie-Stokkermans, Ulrike Sattler, and Ashish Tiwari. Special issue on automated deduction: Decidability, complexity, tractability. Journal of Symbolic Computation, 45(2): 151–152, February 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gianni:1998:RSP

[GSST98] P. Gianni, M. Seppälä, R. Silhol, and B. Trager. Riemann surfaces, plane algebraic curves and their period matrices. Journal of Symbolic Computation, 26(6):789– 803, December 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gupta:2012:TBD

[GSSV12] Somit Gupta, Soumojit Sarkar, Arne Storjohann, and Johnny Valeriote. Triangular x-basis decompositions and derandomization of linear algebra algorithms over K[x]. Journal of Symbolic Computation, 47(4):422–453, April 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001416.

Greuel:2011:GBI

[GSW11] Gert-Martin Greuel, Frank Seelisch, and Oliver Wienand. The Gröbner basis of the ideal of vanishing polynomials. Journal of Symbolic Computation, 46(5):561–570, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001756.

Gerdt:1985:CAA

[GSZ85] V. P. Gerdt, A. B. Shvachka, and A. Yu. Zharkov. Computer algebra application for classification of integrable non-linear evolution equations. Journal of Symbolic Computation, 1(1):101– 107, March 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Guo:2013:DTO

[GSZ13] Li Guo, William Y. Sit, and Ronghua Zhang. Differential type operators and Gröbner– Shirshov bases. Journal of Symbolic Computation, 52(??):97–123, May 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001319.

Gomez-Torrecillas:2016:ITM

[GTLN16] José Gómez-Torrecillas, F. J. Lobillo, and Gabriel Navarro. An isomorphism test for modules over a non-commutative PID. Applications to similarity of Ore polynomials. Journal of Symbolic Computation, 75(??):149–170, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711500111X.

Gomez-Torrecillas:2017:CSE

[GTLN17] José Gómez-Torrecillas, F. J. Lobillo, and Gabriel Navarro. Computing separability elements for the sentenceambient algebra of split ideal codes. Journal of Symbolic Computation, 83(??):211–227, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301316.

Gomez-Torrecillas:2019:CBO

[GTLN19] José Gómez-Torrecillas, F. J.

Lobillo, and Gabriel Navarro. Computing the bound of an Ore polynomial. Applications to factorization. Journal of Symbolic Computation, 92 (??):269–297, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711830049X.

Gomez-Torrecillas:2021:CDI

[GTLN21] José Gómez-Torrecillas, F. J.

Lobillo, and Gabriel Navarro. Cyclic distances of idempotent convolutional codes. Journal of Symbolic Computation, 102(??):37–62, January/February 2021. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301142.

Gianni:1988:GBP

[GTZ88] Patrizia Gianni, Barry Trager, and Gail Zacharias. Gröbner bases and primary decomposition of polynomial ideals. Journal of Symbolic Computation, 6(2–3):149–168 (or 149– 167??), October/December 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Gutierrez:2021:SCI

[GU21] Jaime Gutierrez and Jorge Jiménez Urroz. On some classes of irreducible polynomials. Journal of Symbolic Computation, 105(??):64–70, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300432.

Guevara:2018:PTF

[Gue18] Douglas Navarro Guevara. Primitive transcendental functions and symbolic computation. Journal of Symbolic Computation, 87(??): 28–53, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300573.

Gunther:1990:MPR

[Gün90] Oliver Günther. Minimum K-partitioning of rectilinear polygons. Journal of Symbolic Computation, 9(4):457– 484 (or 457–483??), April 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Guo:2020:DPF

[Guo20]  Zeyu Guo.  Deterministic

polynomial factoring over finite fields: a uniform approach via P -schemes. Journal of Symbolic Computation, 96(??):22–61, January/February 2020. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300227.

Grigorev:1988:SSP

[GV88] D. Yu. Grigor′ev and N. N.

Vorobjov, Jr. Solving systems of polynomial inequalities in sub-exponential time. Journal of Symbolic Computation, 5(1–2):37–64, February/April 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gonzalez-Vega:1996:AQE

[GV96] Laureano Gonzalez-Vega. Applying quantifier elimination to the Birkhoff interpolation problem. Journal of Symbolic Computation, 22(1):83–103, July 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gonzalez-Vega:1997:IPC

[GV97] Laureano Gonzalez-Vega. Implicitization of parametric curves and surfaces by using multidimensional Newton formulae. Journal of Symbolic Computation, 23(2– 3):137–152 (or 137–151??), February/March 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Ganzha:1999:ACA

[GV99] Victor G. Ganzha and Evgenii V. Vorozhtsov. Application of computer algebra systems for stability analysis of difference schemes on curvilinear grids. Journal of Symbolic Computation, 28(3):401–433, September 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0289/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0289/production/pdf;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0289/production/ref.

Gago-Vargas:2003:BPM

[GV03] Jesús Gago-Vargas. Bases for projective modules in A_{n}(k). Journal of Symbolic Computation, 36(6):845–853, December 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gine:2016:CPC

[GV16] Jaume Giné and Claudia Valls. Center problem in the center manifold for quadratic differential systems in R^{3}. Journal of Symbolic Computation, 73(??): 250–267, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000498.

Gonzalez-Vega:1999:SEU

[GVGC99] Laureano Gonzalez-Vega and Neila Gonzalez-Campos. Simultaneous elimination by using several tools from real algebraic geometry. Journal of Symbolic Computation, 28(1–2):89–103, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0268;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1998.0268/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0268/ref.

Gago-Vargas:2005:CTC

[GVHHUE05] J. Gago-Vargas, M. I. Hartillo-Hermoso, and J. M. Ucha-Enríquez. Comparison of theoretical complexities of two methods for computing annihilating ideals of polynomials. Journal of Symbolic Computation, 40(3): 1076–1086, September 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gao:2000:AEF

[GvPS00] Shuhong Gao, Joachim von zur Gathen, Daniel Panario, and Victor Shoup. Algorithms for exponentiation in finite fields. Journal of Symbolic Computation, 29(6):879–889, June 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0309;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0309/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0309/ref.

Gao:2009:CSMb

[GVYZ09] X. S. Gao, J. Van der Hoeven, C. M. Yuan, and G. L. Zhang. Characteristic set method for differential–difference polynomial systems. Journal of Symbolic Computation, 44(9): 1137–1163, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Giesbrecht:2011:HKG

[GW11] Mark W. Giesbrecht and Stephen M. Watt. In honour of Keith Geddes on his 60th birthday. Journal of Symbolic Computation, 46(7):735–740, July 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001392.

Guangxing:2004:EDM

[GX04] Zeng Guangxing and Zeng Xiaoning. An effective decision method for semidefinite polynomials. Journal of Symbolic Computation, 37(1):

83–99, January 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gonzalez:1989:IBR

[GZ89] Teofilo Gonzalez and Si-Qing Zheng. Improved bounds for rectangular and guillotine partitions. Journal of Symbolic Computation, 7(6): 591–610, June 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gerdt:1990:CCI

[GZ90] V. P. Gerdt and A. Yu.

Zharkov. Computer classification of integrable coupled KdV-like systems. Journal of Symbolic Computation, 10(2): 203–208 (or 203–207??), August 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hashemi:2010:EC

[HA10] Amir Hashemi and Gwénolé Ars. Extended F_{5} criteria. Journal of Symbolic Computation, 45(12):1330– 1340, December 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hagino:1989:CM

[Hag89a] Tatsuya Hagino. Codatatypes in ML. Journal of Symbolic Computation, 8(6):629– 650, December 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hagiya:1989:MCI

[Hag89b]  Masami Hagiya.  Meta-

circular interpreter for a strongly typed language. Journal of Symbolic Computation, 8(6):651–680, December 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hansen:2012:IVS

[HAGW12] Christian Mahesh Hansen, Roger Antonsen, Martin Giese, and Arild Waaler. Incremental variable splitting. Journal of Symbolic Computation, 47(9): 1046–1065, September 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002252.

Hallouin:2001:CLI

[Hal01] Emmanuel Hallouin. Computing local integral closures. Journal of Symbolic Computation, 32(3): 211–230, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0447;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0447/
pdf;   http://www.idealibrary.

com/links/doi/10.1006/jsco.
2001.0447/ref.

Hall:2013:CHU

[Hal13] Becky Eide Hall. Computing homology using generalized Gröbner bases. Journal of Symbolic Computation, 54(??):59–71, July 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000138.

Harper:1992:CTS

[Har92] Robert Harper. Constructing type systems over an operational semantics. Journal of Symbolic Computation, 14 (1):71–84, July 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Harvey:2009:FPM

[Har09] David Harvey. Faster polynomial multiplication via multipoint Kronecker substitution. Journal of Symbolic Computation, 44(10):1502–1510, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Harrison:2012:EKA

[Har12a] Michael C. Harrison. An extension of Kedlaya’s algorithm for hyperelliptic curves. Journal of Symbolic Computation, 47(1):89–101, January 2012. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001313.

Harvey:2012:KIM

[Har12b] David Harvey. The Karatsuba integer middle product. Journal of Symbolic Computation, 47(8):954–967, August 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000053.

Harrison:2013:ESR

[Har13] Michael Harrison. Explicit solution by radicals, gonal maps and plane models of algebraic curves of genus 5 or

6. Journal of Symbolic Computation, 51(??):3–21, April 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001125.

Harvey:2014:FAN

[Har14] David Harvey. Faster arithmetic for number-theoretic transforms. Journal of Symbolic Computation, 60 (??):113–119, January 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001181.

Harris:2017:CSC

[Har17] Corey Harris. Computing Segrè classes in arbitrary projective varieties. Journal of Symbolic Computation, 82(??):26–37, September/October 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301687.

Havel:1991:SEU

[Hav91] Timothy F. Havel. Some examples of the use of distances as coordinates for Euclidean geometry. Journal of Symbolic Computation, 11(5–6): 579–594 (or 579–593??), May/ June 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Invarianttheoretic algorithms in geometry (Minneapolis, MN, 1987).

Haager:1995:ADP

[HBN95] Gernot Haager, Gerd Baumann, and Theo F. Nonnenmacher. An algorithm to determine potential systems in Mathematica. Journal of Symbolic Computation, 20(2): 179–196, August 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hubert:2012:CSB

[HC12] Evelyne Hubert and Marie-Paule Cani. Convolution surfaces based on polygonal curve skeletons. Journal of Symbolic Computation, 47(6):680–699, June 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002197.

Hill:1996:PIF

[HCB96] Jonathan M. D. Hill, Keith M. Clarke, and Richard Bornat. Parallelizing imperative functional programs: the vectorization monad. Journal of Symbolic Computation, 21 (4/5/6):561–576, April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Huang:2021:ESB

[HcL21] Zhenyu Huang, Yao Sun (c), and Dongdai Lin. On the efficiency of solving Boolean polynomial systems with the characteristic set method. Journal of Symbolic Computation, 103(??): 66–94, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301373.

Hillar:2013:FTA

[HdC13] Christopher J. Hillar and Abraham Martín del Campo. Finiteness theorems and algorithms for permutation invariant chains of Laurent lattice ideals. Journal of Symbolic Computation, 50 (??):314–334, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001368.

See corrigendum [HdC16].

Hillar:2016:CFT

[HdC16] Christopher J. Hillar and Abraham Martín del Campo. Corrigendum to “Finiteness theorems and algorithms for permutation invariant chains of Laurent lattice ideals” [J. Symb. Comput. 50 (March 2013) 314–334]. Journal of Symbolic Computation, 74 (??):650–652, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000929.

See [HdC13].

Han:2017:OWC

[HDHX17] Jingjun Han, Liyun Dai, Hoon Hong, and Bican Xia. Open weak CAD and its applications. Journal of Symbolic Computation, 80 (part 3) (??):785–816, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300797.

Hemmer:2011:CEE

[HDPS11] Michael Hemmer, Laurent Dupont, Sylvain Petitjean, and Elmar Schömer. A complete, exact and efficient implementation for computing the edge-adjacency graph of an arrangement of quadrics. Journal of Symbolic Computation, 46(4):467–494, April 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110002002.

Hong:2012:VQE

[HE12] Hoon Hong and Mohab Safey El Din. Variant quantifier elimination. Journal of Symbolic Computation, 47(7):883–901, July 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002057.

Helminck:1996:CO

[Hel96] A. G. Helminck. Computing B-orbits on G/H. Journal of Symbolic Computation, 21(2):169–210 (or 169– 209??), February 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Helminck:2000:COM

[Hel00] Aloysius G. Helminck. Computing orbits of minimal parabolic k-subgroups acting on symmetric k-varieties. Journal of Symbolic Computation, 30(5):521–553, October 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0395;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0395/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0395/ref.

Helmer:2016:ACT

[Hel16] Martin Helmer. Algorithms to compute the topological Euler characteristic, Chern– Schwartz–MacPherson class and Segrè class of projective varieties. Journal of Symbolic Computation, 73(??): 120–138, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000322.

Hemmecke:2002:CPS

[Hem02] Ralf Hemmecke. Continuously parameterized symmetries and Buchberger’s algorithm. Journal of Symbolic Computation, 33(1):43–55, January 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0478;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0478/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0478/ref.

Hemmecke:2018:DSR

[Hem18] Ralf Hemmecke. Dancing samba with Ramanujan partition congruences. Journal of Symbolic Computation, 84 (??):14–24, January/February 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300147.

Heng:1990:SRN

[Hen90] Aik-Koan K. Heng. Some remarks on numerical iterations in a symbolic manipulation system. Journal of Symbolic Computation, 10(2): 209–221, August 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hendriks:1998:ADD

[Hen98] P. A. Hendriks. An algorithm determining the difference Galois group of second order linear difference equations. Journal of Symbolic Computation, 26(4):445–462, October 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Herbert:1994:SLR

[Her94] D. J. Herbert. Symbolic local refinement of tetrahedral grids. Journal of Symbolic Computation, 17(5):457–472, May 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hess:2002:CRR

[Hes02] F. Hess. Computing Riemann– Roch spaces in algebraic function fields and related topics. Journal of Symbolic Computation, 33(4):425–445, April 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hauenstein:2021:SDS

[HESV21] Jon D. Hauenstein, Mohab Safey El Din, Éric Schost, and Thi Xuan Vu. Solving determinantal systems using homotopy techniques. Journal of Symbolic Computation, 104(??):754–804, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301103.

Heuberger:1998:FQT

[Heu98] Clemens Heuberger. On a family of quintic Thue equations. Journal of Symbolic Computation, 26(2):173–185, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Heuberger:2006:AST

[Heu06] Clemens Heuberger. All solutions to Thomas’ family of Thue equations over imaginary quadratic number fields. Journal of Symbolic Computation, 41(9):980–998, September 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Huang:2020:FIA

[HG20] Qiao-Long Huang and Xiao-Shan Gao. Faster interpolation algorithms for sparse multivariate polynomials given by straight-line programs. Journal of Symbolic Computation, 101(??):367– 386, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301105.

Hemel:2011:SCC

[HGKV11] Zef Hemel, Danny M. Groenewegen, Lennart C. L. Kats, and Eelco Visser. Static consistency checking of Web applications with WebDSL. Journal of Symbolic Computation, 46(2):150–182, February 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001367.

Hofbauer:1994:LTR

[HH94] Dieter Hofbauer and Maria Huber. Linearizing term rewriting systems using test sets. Journal of Symbolic Computation, 17(1):91– 129, January 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Conditional term rewriting systems (Pont-à-Mousson, 1992).

Hajdu:1998:EBS

[HH98] L. Hajdu and T. Herendi. Explicit bounds for the solutions of elliptic equations with rational coefficients. Journal of Symbolic Computation, 25(3): 361–366, March 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Holt:1999:CAC

[HH99] D. F. Holt and D. F. Hurt.

Computing automatic coset systems and subgroup presentations. Journal of Symbolic Computation, 27(1):1– 20, January 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hoa:2004:CMR

[HH04] Lê Tuân Hoa and Eero Hyry.

Castelnuovo–Mumford regularity of initial ideals. Journal of Symbolic Computation, 38(5):1327–1341, November 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hefez:2007:SBL

[HH07] A. Hefez and M. E. Hernandes. Standard bases for local rings of branches and their modules of differentials. Journal of Symbolic Computation, 42(1–2):178–191, January/February 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hefez:2009:ACP

[HH09] Abramo Hefez and Marcelo Escudeiro Hernandes. Analytic classification of plane branches up to multiplicity

4. Journal of Symbolic Computation, 44(6):626–634, June 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hefez:2013:AIA

[HH13] Abramo Hefez and Marcelo Escudeiro Hernandes. Algorithms for the implementation of the analytic classification of plane branches. Journal of Symbolic Computation, 50(??):308–313, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001356.

Herman:2016:QPR

[HH16] Aaron Herman and Hoon Hong. Quality of positive root bounds. Journal of Symbolic Computation, 74 (??):592–602, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000978.

Hong:2017:ACB

[HHK17] Hoon Hong, Zachary Hough, and Irina A. Kogan. Algorithm for computing μbases of univariate polynomials. Journal of Symbolic Computation, 80 (part 3) (??):844–874, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300943.

Hauenstein:2023:SIA

[HHK^{+}23] Jonathan D. Hauenstein, Yang-Hui He, Ilias Kotsireas, Dhagash Mehta, and Tingting Tang. Special issue on algebraic geometry and machine learning. Journal of Symbolic Computation, 118(??): 93–94, September/October 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122001006.

Hou:2021:LCR

[hHL21] Qing hu Hou and Guojie Li. Log-concavity of P - recursive sequences. Journal of Symbolic Computation, 107(??):251–268, November/ December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000237.

Hu:2013:DET

[HHLQ13] Shenglong Hu, Zheng-Hai Huang, Chen Ling, and Liqun Qi. On determinants and eigenvalue theory of tensors. Journal of Symbolic Computation, 50(??):508–531, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711200168X.

Hashemi:2021:BIN

[HHPS21] Amir Hashemi, Joos Heintz, Luis M. Pardo, and Pablo Solernó. On Bézout inequalities for non-homogeneous polynomial ideals. Journal of Symbolic Computation, 106(??):1– 22, September/October 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771712030119X.

Harris:2023:SPS

[HHS23] Katherine Harris, Jonathan D. Hauenstein, and Agnes Szanto. Smooth points on semialgebraic sets. Journal of Symbolic Computation, 116 (??):183–212, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000967.

Herman:2018:IRS

[HHT18] Aaron Herman, Hoon Hong, and Elias Tsigaridas. Improving root separation bounds. Journal of Symbolic Computation, 84(??):25–56, January/February 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300226.

Huang:1994:EAR

[HI94] Ming-Deh D. Huang and Doug Ierardi. Efficient algorithms for the Riemann– Roch problem and for addition in the Jacobian of a curve. Journal of Symbolic Computation, 18(6):519–540 (or 519– 539??), December 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Huang:1998:CPC

[HI98] Ming-Deh D. Huang and Doug Ierardi. Counting points on curves over finite fields. Journal of Symbolic Computation, 25(1):1–22 (or 1– 21??), January 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hashiguchi:2008:CFM

[HI08] Hiroki Hashiguchi and Toshiya Iwashita. Calculation of a formal moment generating function by using a differential operator. Journal of Symbolic Computation, 43(6–7):508– 514, June/July 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hiep:2016:RCC

[Hie16] Dang Tuan Hiep. Rational curves on Calabi–Yau threefolds: Verifying mirror symmetry predictions. Journal of Symbolic Computation, 76(??):65–83, September/October 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115001303.

Hilali:1987:ADN

[Hil87]  A. Hilali.  On the alge-

braic and differential Newton– Puiseux polygons. Journal of Symbolic Computation, 4(3): 335–349, December 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hillar:2005:CR

[Hil05a] Christopher J. Hillar. Cyclic resultants. Journal of Symbolic Computation, 39(6): 653–669, June 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See erratum [Hil05b].

Hillar:2005:ECR

[Hil05b] Christopher J. Hillar. Erratum to “Cyclic resultants” [J. Symbolic Comput. 39 (6) (2005) 653–669]. Journal of Symbolic Computation, 40(3): 1126–1127, September 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [Hil05a].

Hiroyuki:1989:CBC

[Hir89] Sato Hiroyuki. E-CCC: between CCC and topos — its expressive power from the viewpoint of data type theory. Journal of Symbolic Computation, 8(6):681–695, December 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hidalgo:2015:FMG

[HJ15] Ruben A. Hidalgo and Pilar Johnson. Field of moduli of generalized Fermat curves of type (k, 3) with an application to non-hyperelliptic dessins d’enfants. Journal of Symbolic Computation, 71(??): 60–72, November/December 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001102.

Hauer:2018:PAS

[HJ18] Michael Hauer and Bert Jüttler. Projective and affine symmetries and equivalences of rational curves in arbitrary dimension. Journal of Symbolic Computation, 87 (??):68–86, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300603.

Hoffmann:1997:SCC

[HJA97] Christoph M. Hoffmann and Robert Joan-Arinyo. Symbolic constraints in constructive geometric constraint solving. Journal of Symbolic Computation, 23(2–3):287–300 (or 287–299??), February/March 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Hidalgo:2017:HBA

[HJA17] Marta R. Hidalgo and Robert Joan-Arinyo. A Hennebergbased algorithm for generating tree-decomposable minimally rigid graphs. Journal of Symbolic Computation, 79 (part 2)(??):232– 248, March/April 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116000158.

Hentzel:1993:EI

[HJM93] Irvin Roy Hentzel, David P. Jacobs, and Sekhar V. Muddana. Experimenting with the identity (xy)z = y(zx). Journal of Symbolic Computation, 16(3):289–294 (or 289– 293??), September 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See errata [HJM94].

Hentzel:1994:EEI

[HJM94] I. R. Hentzel, D. P. Jacobs, and S. V. Muddana. Erratum: “Experimenting with the identity (xy)z = y(zx)” [J. Symbolic Comput. 16 (1993), no. 3, 289– 293, MR 94i:20101]. Journal of Symbolic Computation, 17(2):213, 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [HJM93].

Herrero:2013:ASS

[HJS13] María Isabel Herrero, Gabriela Jeronimo, and Juan Sabia. Affine solution sets of sparse polynomial systems. Journal of Symbolic Computation, 51(??):34–54, April 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001149.

Harrison:2016:PAW

[HJS16] Gavin Harrison, Jeremy Johnson, and B. David Saunders. Probabilistic analysis of Wiedemann’s algorithm for minimal polynomial computation. Journal of Symbolic Computation, 74 (??):55–69, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000528.

Hampe:2018:ATS

[HJS18] Simon Hampe, Michael Joswig, and Benjamin Schröter. Algorithms for tight spans and tropical linear spaces. Journal of Symbolic Computation, 91(??):116–128, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300816.

Harrison:2022:PAB

[HJS22] Gavin Harrison, Jeremy Johnson, and B. David Saunders. Probabilistic analysis of block Wiedemann for leading invariant factors. Journal of Symbolic Computation, 108(??): 98–116, January/February 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000419.

Han:2016:PIS

[HJX16] Jingjun Han, Zhi Jin, and Bican Xia. Proving inequalities and solving global optimization problems via simplified CAD projection. Journal of Symbolic Computation, 72(??):206–230, January/February 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000164.

Hermann:1995:CCP

[HK95] Miki Hermann and Phokion G. Kolaitis. The complexity of counting problems in equational matching. Journal of Symbolic Computation, 20 (3):343–362, September 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hubert:2007:RIG

[HK07] Evelyne Hubert and Irina A.

Kogan. Rational invariants of a group action. Construction and rewriting. Journal of Symbolic Computation, 42(1–2):203–217, January/February 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Han:2010:FSD

[HK10] Sunghyu Han and Jon-Lark Kim. Formally self-dual additive codes over F_{4}. Journal of Symbolic Computation, 45 (7):787–799, July 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Horacek:2021:CCA

[HK21] Jan Horácek and Martin Kreuzer. On conversions from CNF to ANF. Journal of Symbolic Computation, 100(??):164–186, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300896.

Hintermeier:1998:DTC

[HKK98] Claus Hintermeier, Claude Kirchner, and Hélène Kirchner. Dynamically typed computations for order-sorted equational presentations. Journal of Symbolic Computation, 25(4):455–526, April 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hermiller:1999:MOR

[HKL99] S. M. Hermiller, X. H. Kramer, and R. C. Laubenbacher. Monomial orderings, rewriting systems, and Gröbner bases for the commutator ideal of a free algebra. Journal of Symbolic Computation, 27(2):133–142, February 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Heuberger:2024:NRB

[HKL24] Clemens Heuberger, Daniel Krenn, and Gabriel F. Lipnik. A note on the relation between recognisable series and regular sequences, and their minimal linear representations. Journal of Symbolic Computation, 123 (??):??, July/August 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123001098.

Hong:2006:BBL

[HKP^{+}06] Hoon Hong, Deepak Kapur, Peter Paule, Franz Winkler, and Faculty of RISC-Linz. Bruno Buchberger — a life devoted to symbolic computation. Journal of Symbolic Computation, 41(3–4): 255–258, March/April 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Heldt:2009:ACZ

[HKPP09] Daniel Heldt, Martin Kreuzer, Sebastian Pokutta, and Hennie Poulisse. Approximate computation of zerodimensional polynomial ideals. Journal of Symbolic Computation, 44(11):1566– 1591, November 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Heule:2021:NWM

[HKS21] Marijn J. H. Heule, Manuel Kauers, and Martina Seidl. New ways to multiply 3 × 3matrices. Journal of Symbolic Computation, 104(??): 899–916, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301139.

Hong:2017:RCI

[HKSS17] Hoon Hong, Yonggu Kim, Georgy Scholten, and J. Rafael Sendra. Resultants over commutative idempotent semirings I: Algebraic aspect. Journal of Symbolic Computation, 79 (part 2)(??): 285–308, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116000183.

Huang:2018:LFD

[HKYY18] Ming-Deh A. Huang, Michiel Kosters, Yun Yang, and Sze Ling Yeo. On the last fall degree of zero-dimensional Weil descent systems. Journal of Symbolic Computation, 87(??):207–226, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300949.

Hong:1997:SIA

[HL97] Hoon Hong and Richard Liska. Special issue on applications of quantifier elimination: Foreword of the Guest Editors. Journal of Symbolic Computation, 24(2):123, August 1997.

Hayden:1998:NIE

[HL98] M. B. Hayden and E. A. Lamagna. Newton: An interactive environment for exploring mathematics. Journal of Symbolic Computation, 25(2): 195–212, February 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Heath:2004:NAG

[HL04] Lenwood S. Heath and Nicholas A. Loehr. New algorithms for generating Conway polynomials over finite fields.

Journal of Symbolic Computation, 38(2):1003–1024, August 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hauenstein:2016:CPC

[HL16] Jonathan D. Hauenstein and Alan C. Liddell, Jr. Certified predictor-corrector tracking for Newton homotopies. Journal of Symbolic Computation, 74(??):239–254, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711500070X.

Hauenstein:2017:CSS

[HL17] Jonathan D. Hauenstein and Viktor Levandovskyy. Certifying solutions to square systems of polynomial–exponential equations. Journal of Symbolic Computation, 79 (part 3)(??):575–593, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300578.

Huang:2018:GSM

[HL18] Ming-Deh Huang and Lian Liu. Generating sets for the multiplicative groups of algebras over finite fields and expander graphs. Journal of Symbolic Computation, 85 (??):170–187, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711730072X.

Hoffmann:2021:CAOa

[HL21a] Johannes Hoffmann and Viktor Levandovskyy. Constructive arithmetics in Ore localizations enjoying enough commutativity. Journal of Symbolic Computation, 102(??): 209–230, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301269.

Hoffmann:2021:CAOb

[HL21b] Johannes Hoffmann and Viktor Levandovskyy. Constructive arithmetics in Ore localizations of domains. Journal of Symbolic Computation, 98(??):23–46, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300690.

Hiss:1995:MDM

[HLM95] Gerhard Hiss, Klaus Lux, and Jürgen Müller. The 2-modular decomposition matrices of the non-principal blocks of maximal defect of the triple cover of the sporadic simple McLaughlin group. Journal of Symbolic Computation, 19 (6):585–600, June 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Holt:2021:PTP

[HLN^{+}21] Derek Holt, Stephen Linton, Max Neunhöffer, Richard Parker, Markus Pfeiffer, and Colva M. Roney-Dougal. Polynomial-time proofs that groups are hyperbolic. Journal of Symbolic Computation, 104(??):419–475, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300845.

He:2022:MLS

[HLO22] Yang-Hui He, Kyu-Hwan Lee, and Thomas Oliver. Machinelearning the Sato–Tate conjecture. Journal of Symbolic Computation, 111(??): 61–72, July/August 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000729.

He:2023:MLI

[HLO23] Yang-Hui He, Kyu-Hwan Lee, and Thomas Oliver. Machine learning invariants of arithmetic curves. Journal of Symbolic Computation, 115(??): 478–491, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000839.

Hong:1997:TSQ

[HLS97] Hoon Hong, Richard Liska, and Stanly Steinberg. Testing stability by quantifier elimination. Journal of Symbolic Computation, 24(2):161– 188 (or 161–187??), August 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Applications of quantifier elimination (Albuquerque, NM, 1995).

Hoffman:2001:DPGb

[HLS01a] Christoph M. Hoffman, Andrew Lomonosov, and Meera Sitharam. Decomposition plans for geometric constraint problems, Part II: New algorithms. Journal of Symbolic Computation, 31(4):409–427, April 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0403;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0403/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0403/ref.

Hoffman:2001:DPGa

[HLS01b] Christoph M. Hoffman, Andrew Lomonosov, and Meera Sitharam. Decomposition plans for geometric constraint systems, Part I: Performance measures for CAD. Journal of Symbolic Computation, 31(4):367–408, April 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0402;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0402/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0402/ref.

Hegedus:2015:TBI

[HLSS15] Gábor Hegedüs, Zijia Li, Josef Schicho, and Hans-Peter Schröcker. The theory of bonds II: Closed 6R linkages with maximal genus. Journal of Symbolic Computation, 68 (part 2)(??):167–180, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001035.

Hoffman:2016:GCW

[HLSW16] J. William Hoffman, Zhibin Liang, Yukiko Sakai, and Haohao Wang. Genus 3 curves whose Jacobians have endomorphisms by Q(ζ_{7} + ζ̄_{7}). Journal of Symbolic Computation, 74(??):561–577, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000954.

Huang:2018:PRI

[HLXL18] Cheng-Chao Huang, Jing-Cao Li, Ming Xu, and Zhi-Bin Li. Positive root isolation for poly-powers by exclusion and differentiation. Journal of Symbolic Computation, 85(??):148–169, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300718.

Havas:1997:IMD

[HM97] George Havas and Bohdan S. Majewski. Integer matrix diagonalization. Journal of Symbolic Computation, 24 (3–4):399–408, September/ October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Hajir:2002:TRT

[HM02a] Farshid Hajir and Christian Maire. Tamely ramified towers and discriminant bounds for number fields—II. Journal of Symbolic Computation, 33 (4):415–423, April 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hong:2002:SRC

[HM02b] Hoon Hong and Manfred Minimair. Sparse resultant of composed polynomials I mixed–unmixed case. Journal of Symbolic Computation, 33 (4):447–465, April 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Helm:2005:AGI

[HM05] David Helm and Ezra Miller.

Algorithms for graded injective resolutions and local cohomology over semigroup rings. Journal of Symbolic Computation, 39(3–4): 373–395, March/April 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hosten:2006:P

[HM06] Serkan Hoşten and Christopher Meek. Preface. Journal of Symbolic Computation, 41(2):123–124, February 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hemmecke:2009:CGS

[HM09] Raymond Hemmecke and Peter N. Malkin. Computing generating sets of lattice ideals and Markov bases of lattices. Journal of Symbolic Computation, 44(10):1463–1476, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hu:2021:FPS

[HM21] Jiaxiong Hu and Michael Monagan. A fast parallel sparse polynomial GCD algorithm. Journal of Symbolic Computation, 105(??):28–63, ???? 2021. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300420.

Hashemi:2023:NAC

[HM23] Amir Hashemi and H. Michael Möller. A new algorithm for computing staggered linear bases. Journal of Symbolic Computation, 117(??): 1–14, July/August 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712200116X.

Henderson:2006:SCF

[HMN06] Richard Henderson, Todd Macedo, and Sam Nelson. Symbolic computation with finite quandles. Journal of Symbolic Computation, 41(7): 811–817, July 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hauenstein:2017:DMS

[HMS17] Jonathan D. Hauenstein, Bernard Mourrain, and Agnes Szanto. On deflation and multiplicity structure. Journal of Symbolic Computation, 83(??):228–253, November/December 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711630133X.

Hernandez:2007:CCN

[HMXD07] Xavier Hernández, Josep M.

Miret, and Sebastià Xambó-Descamps. Computing the characteristic numbers of the variety of nodal plane cubics in P^{3}. Journal of Symbolic Computation, 42(1–2): 192–202, January/February 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See corrigendum [MHXD09].

Hou:2021:PRS

[HMZ21] Qing-Hu Hou, Yan-Ping Mu, and Doron Zeilberger. Polynomial reduction and supercongruences. Journal of Symbolic Computation, 103(??): 127–140, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301488.

Henrion:2016:RRF

[HNE16] Didier Henrion, Simone Naldi, and Mohab Safey El Din. Real root finding for determinants of linear matrices. Journal of Symbolic Computation, 74(??):205–238, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000607.

Henrion:2021:EAS

[HNE21] Didier Henrion, Simone Naldi, and Mohab Safey El Din. Exact algorithms for semidefinite programs with degenerate feasible set. Journal of Symbolic Computation, 104 (??):942–959, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301176.

Harkonen:2024:SCR

[HNR24] Marc Härkönen, Lisa Nicklasson, and Bogdan Raita. Syzygies, constant rank, and beyond. Journal of Symbolic Computation, 123(??): ??, July/August 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000883.

Hyun:2021:BKT

[HNRS21] Seung Gyu Hyun, Vincent Neiger, Hamid Rahkooy, and Éric Schost. Block-Krylov techniques in the context of sparse-FGLM algorithms. Journal of Symbolic Computation, 98(??):163–191, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300756.

Hong:1995:DSP

[HNS95] Hoon Hong, Andreas Neubacher, and Wolfgang Schreiner. The design of the SACLIB/ PACLIB kernels. Journal of Symbolic Computation, 19(1/2/3):111–132, January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Havas:1990:NQA

[HNVL90] George Havas, M. F. Newman, and M. R. Vaughan-Lee. A nilpotent quotient algorithm for graded Lie rings. Journal of Symbolic Computation, 9(5–6):653–664, May/ June 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 1.

Hofling:2001:CPI

[Höf01]  Burkhard Höfling.  Com-

puting projectors, injectors, residuals and radicals of finite soluble groups. Journal of Symbolic Computation, 32(5):499–511, November 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0477;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0477/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0477/ref.

Holt:1985:MCF

[Hol85] D. F. Holt. The mechanical computation of first and second cohomology groups. Journal of Symbolic Computation, 1(4):351–362 (or 351– 361??), December 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Holt:1991:CNP

[Hol91] D. F. Holt. The computation of normalizers in permutation groups. Journal of Symbolic Computation, 12(4–5): 499–516, October/November 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Hong:1996:SIP

[Hon96] H. Hong. Special issue on parallel symbolic computation: Foreword of the Guest Editor. Journal of Symbolic Computation, 21(4/5/6):377, April, May & June 1996.

Hong:1997:INC

[Hon97a]  Hoon Hong.  Implicitiza-

tion of nested circular curves. Journal of Symbolic Computation, 23(2–3):177–190 (or 177–189??), February/March 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Hong:1997:SUC

[Hon97b] Hoon Hong. Subresultants under composition. Journal of Symbolic Computation, 23(4): 355–366 (or 355–365??), April 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hong:1998:BAP

[Hon98a] Hoon Hong. Bounds for absolute positiveness of multivariate polynomials. Journal of Symbolic Computation, 25(5):571–586 (or 571–585??), May 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hong:1998:GBU

[Hon98b] Hoon Hong. Gröbner basis under composition I. Journal of Symbolic Computation, 25(5):643–664 (or 643–663??), May 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hong:2000:E

[Hon00] Hoon Hong. Editorial. Journal of Symbolic Computation, 29(1):3–4, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0345;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0345/
pdf.

Hong:2004:NJM

[Hon04] Hoon Hong. Note on Jacobi’s method for approximating dominant roots. Journal of Symbolic Computation, 37 (4):449–453, April 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Heiss:2006:ISS

[HOP06] Werner Heiß, Ulrich Oberst, and Franz Pauer. On inverse systems and squarefree decomposition of zerodimensional polynomial ideals. Journal of Symbolic Computation, 41(3–4):261– 284, March/April 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Horobet:2024:CCD

[Hor24] Emil Horobet. The critical curvature degree of an algebraic variety. Journal of Symbolic Computation, 121(??): Article 102259, March/April 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000731.

Hashemi:2023:RSI

[HOS23] Amir Hashemi, Matthias Orth, and Werner M. Seiler. Recursive structures in involutive bases theory. Journal of Symbolic Computation, 118(??):32–68, September/October 2023. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000032.

Herfort:1991:NNR

[HP91] W. Herfort and H. Penz.

A new notion of reduction: generating universal Gröbner bases of ideals in K(x, y). Journal of Symbolic Computation, 12(6):585–606 (or 585– 605??), December 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hong:2007:BCN

[HP07] Hoon Hong and John Perry.

Are Buchberger’s criteria necessary for the chain condition? Journal of Symbolic Computation, 42(7):717–732, July 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See corrigendum [HP08].

Hong:2008:CBC

[HP08] Hoon Hong and John Perry.

Corrigendum to “Are Buchberger’s criteria necessary for the chain condition?” [J. Symbolic Comput. 42 (2007) 717–732]. Journal of Symbolic Computation, 43(3): 233, March 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [HP07].

Heras:2011:FUI

[HPRS11] J. Heras, V. Pascual, J. Rubio, and F. Sergeraert. fKenzo: a user interface for computations in algebraic topology. Journal of Symbolic Computation, 46(6):685–698, June 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111000174.

Holt:1997:CRG

[HPS97] D. F. Holt, W. Plesken, and B. Souvignier. Constructing a representation of the group (2,3,7;11). Journal of Symbolic Computation, 24(3–4): 489–492, September/October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Hashemi:2022:CMC

[HPS22] Amir Hashemi, Hossein Parnian, and Werner M. Seiler. Computation of Macaulay constants and degree bounds for Gröbner bases. Journal of Symbolic Computation, 111(??):44–60, July/August 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771712100081X.

Heuberger:2002:TFT

[HPT02] Clemens Heuberger, Attila Peth, and Robert F. Tichy.

Thomas’ family of Thue equations over imaginary quadratic fields. Journal of Symbolic Computation, 34(5): 437–449, November 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Herzog:2019:FCM

[HQS19] Jürgen Herzog, Ayesha Asloob Qureshi, and Maryam Mohammadi Saem. The fiber cone of a monomial ideal in two variables. Journal of Symbolic Computation, 94(??):52– 69, September/October 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300877.

Hebisch:2011:ERM

[HR11] Waldemar Hebisch and Martin Rubey. Extended rate, more GFUN. Journal of Symbolic Computation, 46 (8):889–903, August 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000162.

Han:2012:INF

[HR12] Maoan Han and Valery G.

Romanovski. Isochronicity and normal forms of polynomial systems of ODEs. Journal of Symbolic Computation, 47(10):1163–1174, October 2012. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711100232X.

Horobet:2017:MLD

[HR17] Emil Horobet and Jose Israel Rodriguez. The maximum likelihood data singular locus. Journal of Symbolic Computation, 79 (part 1)(??):99–107, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300864.

Hemmecke:2019:CAP

[HR19] Ralf Hemmecke and Silviu Radu. Construction of all polynomial relations among Dedekind eta functions of level N. Journal of Symbolic Computation, 95(??): 39–52, November/December 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118301196.

Holt:2020:CST

[HR20] Derek Holt and Gordon Royle. A census of small transitive groups and vertex-transitive graphs. Journal of Symbolic Computation, 101(??): 51–60, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300598.

Hartzer:2022:ISC

[HRdWY22] Jacob Hartzer, Olivia Röhrig, Timo de Wolff, and Oguzhan Yürük. Initial steps in the classification of maximal mediated sets. Journal of Symbolic Computation, 109(??): 404–425, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300687.

Hreinsdottir:1994:CWC

[Hre94] Freyja Hreinsdóttir. A case where choosing a product order makes the calculations of a Gröbner basis much faster. Journal of Symbolic Computation, 18(4):373–378, October 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hreinsdottir:2006:ITO

[Hre06] Freyja Hreinsdóttir. An improved term ordering for the ideal of commuting matrices. Journal of Symbolic Computation, 41(9):999–1003, September 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Howlett:2001:MGE

[HRT01] R. B. Howlett, L. J. Rylands, and D. E. Taylor. Matrix generators for exceptional groups of Lie type. Journal of Symbolic Computation, 31(4):429–445, April 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0431;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0431/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0431/ref.

Hirschberg:1989:ANR

[HS89] W. Hirschberg and D. Schramm. Application of NEWEUL in robot dynamics. Journal of Symbolic Computation, 7(2): 199–204, February 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hunt:1990:CEC

[HS90] H. B. Hunt III and R. E. Stearns. The complexity of equivalence for commutative rings. Journal of Symbolic Computation, 10(5):411– 436, November 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hearn:1995:CAS

[HS95] Anthony C. Hearn and Eberhard Schrüfer. A computer algebra system based on ordersorted algebra. Journal of Symbolic Computation, 19 (1/2/3):65–79 (or 65–77??), January, February, March 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Hribernig:1997:DVC

[HS97] V. Hribernig and H. J. Stetter.

Detection and validation of clusters of polynomial zeros. Journal of Symbolic Computation, 24(6):667–681, December 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hong:1998:ATC

[HS98] Hoon Hong and Josef Schicho. Algorithms for trigonometric curves (simplification, implicitization, parameterization). Journal of Symbolic Computation, 26(3):279–300, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hendriks:1999:SDE

[HS99] P. A. Hendriks and M. F.

Singer. Solving difference equations in finite terms. Journal of Symbolic Computation, 27(3):239–260, March 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hosten:2000:PDL

[HS00] Serkan Hoşten and Jay Shapiro. Primary decomposition of lattice basis ideals.

Journal of Symbolic Computation, 29(4–5):625–639, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0397;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0397/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0397/ref.

Hillebrand:2001:TEV

[HS01] Andre Hillebrand and Wiland Schmale. Towards an effective version of a theorem of Stafford. Journal of Symbolic Computation, 32(6):699–716, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0491;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0491/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0491/ref.

Hallgrimsdottir:2006:RGL

[HS06] Ingileif B. Hallgrímsdóttir and Bernd Sturmfels. Resultants in genetic linkage analysis. Journal of Symbolic Computation, 41(2):125–137, February 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hanson:2009:TFS

[HS09] Andrew J. Hanson and Ji-Ping Sha. A tessellation for Fermat surfaces in CP^{3}. Journal of Symbolic Computation, 44(6):591–605, June 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hauenstein:2017:WNA

[HS17a] Jonathan D. Hauenstein and Andrew J. Sommese. What is numerical algebraic geometry? Journal of Symbolic Computation, 79 (part 3)(??):499–507, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300529.

Hein:2017:LSF

[HS17b] Nickolas Hein and Frank Sottile. A lifted square formulation for certifiable Schubert calculus. Journal of Symbolic Computation, 79 (part 3)(??):594–608, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711630058X.

Hashemi:2021:DDD

[HS21a] Amir Hashemi and Werner M.

Seiler. Dimension and depth dependent upper bounds in polynomial ideal theory. Journal of Symbolic Computation, 98(??):47–64, ???? 2021.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300707.

Hollering:2021:IPU

[HS21b] Benjamin Hollering and Seth Sullivant. Identifiability in phylogenetics using algebraic matroids. Journal of Symbolic Computation, 104(??): 142–158, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300274.

Hsiang:1987:RMT

[Hsi87] Jieh Hsiang. Rewrite method for theorem proving in first order theory with equality. Journal of Symbolic Computation, 3(1–2):133–151, February/April 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Rewriting techniques and applications (Dijon, 1985).

Huber:1998:NSC

[HSS98] B. Huber, F. Sottile, and B. Sturmfels. Numerical Schubert calculus. Journal of Symbolic Computation, 26(6):767– 788, December 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hausdorf:2002:IBW

[HSS02] Marcus Hausdorf, Werner M.

Seiler, and Rainer Steinwandt. Involutive bases in the Weyl algebra. Journal of Symbolic Computation, 34(3):181– 198, September 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hashemi:2018:DGP

[HSS18] Amir Hashemi, Michael Schweinfurter, and Werner M. Seiler. Deterministic genericity for polynomial ideals. Journal of Symbolic Computation, 86(??):20–50, May/ June 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300366.

Hong:2008:HTD

[HSV08] Jooyoun Hong, Aron Simis, and Wolmer V. Vasconcelos. On the homology of two-dimensional elimination. Journal of Symbolic Computation, 43(4):275–292, April 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hoffmann:1997:SIP

[HSW97] C. M. Hoffmann, J. R. Sendra, and F. Winkler. Special issue on parametric algebraic curves and applications: Foreword by the Guest Editors. Journal of Symbolic Computation, 23(2–3):133–135, February/March 1997.

Hartley:1991:CIC

[HT91] David Hartley and Robin W. Tucker. A constructive implementation of the Cartan– Kähler theory of exterior differential systems. Journal of Symbolic Computation, 12(6):655–668 (or 655– 667??), December 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hartley:1995:GBC

[HT95] David Hartley and Philip Tuckey. Gröbner bases in Clifford and Grassmann algebras. Journal of Symbolic Computation, 20(2):197–206 (or 197– 205??), August 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hernandez:2017:TFS

[HT17] Daniel J. Hernández and Pedro Teixeira. F -threshold functions: Syzygy gap fractals and the two-variable homogeneous case. Journal of Symbolic Computation, 80 (part 2)(??):451–483, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300347.

Helmer:2023:SDR

[HT23] Martin Helmer and Elias Tsigaridas. Segrè-driven radicality testing. Journal of Symbolic Computation, 122(??): Article 102262, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000767.

Hong:2015:SAS

[HTX15] Hoon Hong, Xiaoxian Tang, and Bican Xia. Special algorithm for stability analysis of multistable biological regulatory systems. Journal of Symbolic Computation, 70(??): 112–135, September/October 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001072.

Heuberger:2004:ASF

[HTZ04] Clemens Heuberger, Alain Togbé, and Volker Ziegler. Automatic solution of families of Thue equations and an example of degree 8. Journal of Symbolic Computation, 38(3): 1145–1163, September 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Huang:2023:SPI

[Hua23] Qiao-Long Huang. Sparse polynomial interpolation based on derivatives. Journal of Symbolic Computation, 114(??):359–375, January/ February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000517.

Hubert:1999:ECA

[Hub99]  Evelyne Hubert.  Essen-

tial components of an algebraic differential equation. Journal of Symbolic Computation, 28(4–5): 657–680, October/November 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0319/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0319/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0319/production/ref.

Hubert:2000:FFD

[Hub00] Evelyne Hubert. Factorizationfree decomposition algorithms in differential algebra. Journal of Symbolic Computation, 29(4–5):641–662, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0344;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0344/

pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0344/ref.

Hubert:2009:DIL

[Hub09a] Evelyne Hubert. Differential invariants of a Lie group action: Syzygies on a generating set. Journal of Symbolic Computation, 44(4):382–416, April 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hubrechts:2009:QQE

[Hub09b] Hendrik Hubrechts. Quasiquadratic elliptic curve point counting using rigid cohomology. Journal of Symbolic Computation, 44(9): 1255–1267, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hubert:2019:IAS

[Hub19]  Evelyne Hubert.  Invari-

ant algebraic sets and symmetrization of polynomial systems. Journal of Symbolic Computation, 95(??): 53–67, November/December 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118301202.

Hughes:1990:SCF

[Hug90]  D. I. Hughes.  Symbolic

computation with Fermions. Journal of Symbolic Computation, 10(6):657–664, December 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hulpke:1999:CSI

[Hul99] A. Hulpke. Computing subgroups invariant under a set of automorphisms. Journal of Symbolic Computation, 27 (4):415–428, April 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hulpke:2005:CTP

[Hul05]  Alexander Hulpke.  Con-

structing transitive permutation groups. Journal of Symbolic Computation, 39(1): 1–30, January 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hulpke:2013:CGG

[Hul13] Alexander Hulpke. Computing generators of groups preserving a bilinear form over residue class rings. Journal of Symbolic Computation, 50(??):298–307, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001344.

Hendriks:1995:GAS

[Hv95] Peter A. Hendriks and Marius van der Put. Galois action on solutions of a differential equation. Journal of Symbolic Computation, 19(6):559–576, June 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Homberger:2016:EAE

[HV16] Cheyne Homberger and Vincent Vatter. On the effective and automatic enumeration of polynomial permutation classes. Journal of Symbolic Computation, 76(??):84– 96, September/October 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001297.

Hofstadler:2022:SGB

[HV22] Clemens Hofstadler and Thibaut Verron. Signature Gröbner bases, bases of syzygies and cofactor reconstruction in the free algebra. Journal of Symbolic Computation, 113(??):211–241, November/ December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000219.

Harvey:2018:CIM

[HvdH18] David Harvey and Joris van der Hoeven. On the complexity of integer matrix multiplication. Journal of Symbolic Computation, 89(??):1– 8, November/December 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117301013.

Hou:2024:RSF

[HW24] Qing-Hu Hou and Yarong Wei. Rational solutions to the first order difference equations in the bivariate difference field. Journal of Symbolic Computation, 124(??): ??, September/October 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717124000129.

Ho:1996:HAS

[HY96] Chung-Jen Ho and Chee Keng Yap. The Habicht approach to subresultants. Journal of Symbolic Computation, 21(1): 1–14, January 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hong:2021:CMS

[HY21] Hoon Hong and Jing Yang. A condition for multiplicity structure of univariate polynomials. Journal of Symbolic Computation, 104(??): 523–538, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300882.

Holt:2023:ROF

[HY23a] Derek Holt and Yong Yang. Regular orbits of finite primitive solvable groups, the final classification. Journal of Symbolic Computation, 116 (??):139–145, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000864.

Huang:2023:AAS

[HY23b] Bo Huang and Chee Yap. An algorithmic approach to small limit cycles of nonlinear differential systems: the averaging method revisited. Journal of Symbolic Computation, 115 (??):492–517, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300894.

Huang:2024:SPS

[HYG24] Qiao-Long Huang, Ke Ye, and Xiao-Shan Gao. Skewpolynomial-sparse matrix multiplication. Journal of Symbolic Computation, 121(??): Article 102240, March/April 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000548.

Hirata:2004:TIS

[HYH04] Kouichi Hirata, Keizo Yamada, and Masateru Harao. Tractable and intractable second-order matching problems. Journal of Symbolic Computation, 37(5):611–628, May 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Hanrot:2004:LNM

[HZ04] G. Hanrot and P. Zimmermann. A long note on Mulders’ short product. Journal of Symbolic Computation, 37 (3):391–401, March 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See corrigendum [HZ15].

Hanrot:2015:CLN

[HZ15] G. Hanrot and P. Zimmermann. Corrigendum to “A long note on Mulders’ short product” [J. Symb. Comput. 37 (3) (2004) 391– 401]. Journal of Symbolic Computation, 66(??): 111–112, January/February 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000212.

See [HZ04].

Hou:2019:ALC

[HZ19] Qing-Hu Hou and Zuo-Ru Zhang. Asymptotic r-log-convexity and p-recursive sequences. Journal of Symbolic Computation, 93(??): 21–33, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300427.

Iliman:2015:SIN

[IdW15] Sadik Iliman and Timo de Wolff. Separating inequalities for nonnegative polynomials that are not sums of squares. Journal of Symbolic Computation, 68 (part 2) (??):181–194, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000789.

Ida:2015:FPK

[IGT15] Tetsuo Ida, Fadoua Ghourabi, and Kazuko Takahashi. Formalizing polygonal knot origami. Journal of Symbolic Computation, 69(??):93–108, July/ August 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000996.

Ilten:2013:CGM

[IK13] Nathan Owen Ilten and Lars Kastner. Calculating generators of multigraded algebras. Journal of Symbolic Computation, 51(??):22–33, April 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001137.

Imamoglu:2021:CDC

[IK21] Erdal Imamoglu and Erich L. Kaltofen. On computing the degree of a Chebyshev polynomial from its value. Journal of Symbolic Computation, 104(??):159–167, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300262.

Ida:2011:MTR

[IKGT11] Tetsuo Ida, Asem Kasem, Fadoua Ghourabi, and Hidekazu Takahashi. Morley’s theorem revisited: Origami construction and automated proof. Journal of Symbolic Computation, 46(5):571–583, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001768.

Insua:2009:GBU

[IL09] Manuel A. Insua and Manuel Ladra. Gröbner bases in universal enveloping algebras of Leibniz algebras. Journal of Symbolic Computation, 44(5): 517–526, May 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Imbach:2017:CNA

[IMP17] Rémi Imbach, Guillaume Moroz, and Marc Pouget. A certified numerical algorithm for the topology of resultant and discriminant curves. Journal of Symbolic Computation, 80 (part 2)(??):285–306, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300128.

Idrees:2011:PMA

[IPS11] Nazeran Idrees, Gerhard Pfister, and Stefan Steidel. Parallelization of modular algorithms. Journal of Symbolic Computation, 46(6):672–684, June 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000150.

Ilten:2010:AGC

[IS10] Nathan Owen Ilten and Hendrik Süß. Algebraic geometry codes from polyhedral divisors. Journal of Symbolic Computation, 45(7):734–756, July 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ida:2010:OFA

[IT10] Tetsuo Ida and Hidekazu Takahashi. Origami fold as algebraic graph rewriting. Journal of Symbolic Computation, 45(4):393–413, April 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Imamoglu:2017:CHS

[IvH17] Erdal Imamoğlu and Mark van Hoeij. Computing hypergeometric solutions of second order linear differential equations using quotients of formal solutions and integral bases. Journal of Symbolic Computation, 83(??):254–271, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301353.

Intrigila:1996:RIM

[IZ96] Benedetto Intrigila and Marisa Venturini Zilli. A remark on infinite matching vs. infinite unification. Journal of Symbolic Computation, 21(3): 289–292, March 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Izumi:2016:PMT

[Izu16] Hideaki Izumi. The prompter method: a treatment for hardto-solve iterative functional equations. Journal of Symbolic Computation, 75(??): 193–208, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115001133.

Jacobs:1997:CCN

[Jac97] D. P. Jacobs. A course in computational nonassociative algebra. Journal of Symbolic Computation, 23(5– 6):497–502, May/June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jambor:2011:CMA

[Jam11]  Sebastian Jambor.  Com-

puting minimal associated primes in polynomial rings over the integers. Journal of Symbolic Computation, 46(10):1098–1104, October 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000812.

Jaroschek:2013:IPR

[Jar13] Maximilian Jaroschek. Improved polynomial remainder sequences for Ore polynomials. Journal of Symbolic Computation, 58(??):64– 76, November 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000849.

Johnson:2004:ADI

[JB04] Jeremy R. Johnson and Anthony F. Breitzman. Automatic derivation and implementation of fast convolution algorithms. Journal of Symbolic Computation, 37(2):261– 293, February 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jebelean:1993:AED

[Jeb93] Tudor Jebelean. An algorithm for exact division. Journal of Symbolic Computation, 15(2): 169–180, February 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jebelean:1995:DDL

[Jeb95] Tudor Jebelean. A doubledigit Lehmer–Euclid algorithm for finding the GCD of long integers. Journal of Symbolic Computation, 19(1/2/3):145–158 (or 145–157??), January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Jeffrey:1997:RTI

[Jef97] D. J. Jeffrey. Rectifying transformations for the integration of rational trigonometric functions. Journal of Symbolic Computation, 24(5):563– 574 (or 563–573??), November 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Janovitz-Freireich:2012:CMT

[JFMRS12] Itnuit Janovitz-Freireich, Bernard Mourrain, Lajos Rónyai, and Ágnes Szántó. On the computation of matrices of traces and radicals of ideals. Journal of Symbolic Computation, 47(1):102–122, January 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001325.

Jorge:2009:CPE

[JGF09] J. Santiago Jorge, Victor M.

Gulias, and Jose L. Freire. Certifying properties of an efficient functional program for computing Gröbner bases. Journal of Symbolic Computation, 44(5):571–582, May 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jiang:2021:SGE

[JH21] Yuhan Jiang and Weiqiao Han. Singularities and genus of the k-ellipse. Journal of Symbolic Computation, 104 (??):343–355, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300511.

Jirstrand:1997:NCS

[Jir97]  Mats Jirstrand.  Nonlin-

ear control system design by quantifier elimination. Journal of Symbolic Computation, 24(2):137–152, August 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Applications of quantifier elimination (Albuquerque, NM, 1995).

Jara:2001:SBD

[JJ01] P. Jara and J. Jódar. Uq(sl(2)) satisfies a Bernstein duality. Journal of Symbolic Computation, 32(6): 687–698, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0490;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0490/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0490/ref.

Junttila:2020:APA

[JKKK20] Tommi Junttila, Matti Karppa, Petteri Kaski, and Jukka Kohonen. An adaptive prefixassignment technique for symmetry reduction. Journal of Symbolic Computation, 99 (??):21–49, July/August 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300288.

Jones:1998:CAG

[JKP98] Jon Jones, N. P. Karampetakis, and A. C. Pugh. The computation and application of the generalized inverse via ∥MAPLE∥. Journal of Symbolic Computation, 25(1):99– 124, January 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Johnson:2012:SIS

[JKP12] Jeremy R. Johnson, Erich Kaltofen, and Hyungju Park. Special issue on Symbolic and Algebraic Computation Foundations, Algorithmics and Applications: ISSAC 2009. Journal of Symbolic Computation, 47(7):751, July 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001982.

Jacobsson:1991:SBG

[JL91] Carl Jacobsson and Clas Löfwall. Standard bases for general coefficient rings and a new constructive proof of Hilbert’s basis theorem. Journal of Symbolic Computation, 12(3):337–372 (or 337– 371??), September 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jarrah:2003:SCC

[JLR03] Abdul Salam Jarrah, Reinhard Laubenbacher, and Valery Romanovski. The Sibirsky component of the center variety of polynomial differential systems. Journal of Symbolic Computation, 35 (5):577–589, May 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jin:2013:NAS

[JLW13] Meng Jin, Xiaoliang Li, and Dongming Wang. A new algorithmic scheme for computing characteristic sets. Journal of Symbolic Computation, 50(??):431–449, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001423.

Jacobs:1993:WPF

[JM93] David P. Jacobs and Sekhar V. Muddana. The word problem for free partially commutative, partially associative groupoids. Journal of Symbolic Computation, 16(6):557– 562, December 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jezek:1995:PBE

[JM95] Jaroslav Ježek and George F. McNulty. Perfect bases for equational theories. Journal of Symbolic Computation, 19 (5):489–505, April 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jilani:2013:IFI

[JML^{+}13] Lamia Labed Jilani, Olfa Mraihi, Asma Louhichi, Wided Ghardallou, Khaled Bsaies, and Ali Mili. Invariant functions and invariant relations: an alternative to invariant assertions. Journal of Symbolic Computation, 48(??):1–36, January 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000557.

Johnson:2004:SIC

[JMPR04] Jeremy R. Johnson, José M. F. Moura, Markus Püschel, and Daniel Rockmore. Special issue on computer algebra and signal processing: forward by the guest editors. Journal of Symbolic Computation, 37(2): 133–135, February 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jiu:2018:SAM

[JMV18] Lin Jiu, Victor H. Moll, and Christophe Vignat. A symbolic approach to multiple zeta values at negative integers. Journal of Symbolic Computation, 84(??):1– 13, January/February 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300093.

Janota:2023:CGS

[JMV23] Mikolás Janota, António Morgado, and Petr Vojtechovský. Computing generating sets of minimal size in finite algebras. Journal of Symbolic Computation, 119(??): 50–63, November/December 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000147.

Jeannerod:2017:CMI

[JNSV17] Claude-Pierre Jeannerod, Vincent Neiger, Éric Schost, and Gilles Villard. Computing minimal interpolation bases. Journal of Symbolic Computation, 83(??):272–314, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301365.

Jeannerod:2021:FCA

[JNV21] Claude-Pierre Jeannerod, Vincent Neiger, and Gilles Villard. Fast computation of approximant bases in canonical form. Journal of Symbolic Computation, 98(??):192–224, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300768.

Johnson:2015:WIA

[Joh15] Jeremiah W. Johnson. Weight ideals associated to regular and log-linear arrays. Journal of Symbolic Computation, 67 (??):1–15, March/April 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000546.

Jouanolou:2009:EDQ

[Jou09] Jean-Pierre Jouanolou. An explicit duality for quasihomogeneous ideals. Journal of Symbolic Computation, 44 (7):864–871, July 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jeronimo:2010:MPP

[JP10] Gabriela Jeronimo and Daniel Perrucci. On the minimum of a positive polynomial over the standard simplex. Journal of Symbolic Computation, 45 (4):434–442, April 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jimenez-Pastor:2023:EHS

[JPNP23] Antonio Jiménez-Pastor, Philipp Nuspl, and Veronika Pillwein. An extension of holonomic sequences: C^{2}-finite sequences. Journal of Symbolic Computation, 116(??):400–424, May/ June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122001055.

Jimenez-Pastor:2019:CED

[JPP19] Antonio Jiménez-Pastor and Veronika Pillwein. A computable extension for D-finite functions: DD-finite functions. Journal of Symbolic Computation, 94(??): 90–104, September/October 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300890.

Jimenez-Pastor:2023:FBM

[JPP23] Antonio Jiménez-Pastor and Marko Petkovsek. The factorial-basis method for finding definite-sum solutions of linear recurrences with polynomial coefficients. Journal of Symbolic Computation, 117(??):15–50, July/August 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122001158.

Jaulent:2009:CGN

[JPPSG09] Jean-François Jaulent, Sebastian Pauli, Michael E. Pohst, and Florence Soriano-Gafiuk. Computation of 2-groups of narrow logarithmic divisor classes of number fields. Journal of Symbolic Computation, 44(7):852–863, July 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jeannerod:2013:RPR

[JPS13] Claude-Pierre Jeannerod, Clément Pernet, and Arne Storjohann. Rank-profile revealing Gaussian elimination and the CUP matrix decomposition. Journal of Symbolic Computation, 56(??):46–68, September 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000631.

Jefferson:2019:NRP

[JPW19] Christopher Jefferson, Markus Pfeiffer, and Rebecca Waldecker. New refiners for permutation group search. Journal of Symbolic Computation, 92 (??):70–92, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711730127X.

Jones:2006:DLF

[JR06] John W. Jones and David P.

Roberts. A database of local fields. Journal of Symbolic Computation, 41(1):80– 97, January 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jeronimo:2007:CMR

[JS07] Gabriela Jeronimo and Juan Sabia. Computing multihomogeneous resultants using straight-line programs. Journal of Symbolic Computation, 42(1–2):218–235, January/February 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jeronimo:2018:SRS

[JS18] Gabriela Jeronimo and Juan Sabia. Sparse resultants and straight-line programs. Journal of Symbolic Computation, 87(??):14–27, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300561.

Joubert:2013:UFS

[JSC13] S. V. Joubert, M. Y. Shatalov, and C. E. Coetzee. Using Fourier series to analyse mass imperfections in vibratory gyroscopes. Journal of Symbolic Computation, 61–62(??):116–127, ???? 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711300134X.

Jacquemard:2003:CAP

[JT03] A. Jacquemard and M. A. Teixeira. Computer analysis of periodic orbits of discontinuous vector fields. Journal of Symbolic Computation, 35 (5):617–636, May 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jorgens:2018:IPP

[JTdW18] Thorsten Jörgens, Thorsten Theobald, and Timo de Wolff. Imaginary projections of polynomials. Journal of Symbolic Computation, 91(??):181–199, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300853.

Joshua:2009:ISA

[JV09] Roy Joshua and Shaun Van Ault. Implementation of Stanley’s algorithm for projective group imbeddings. Journal of Symbolic Computation, 44 (6):655–672, June 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jia:2016:CDV

[JWC^{+}16] Xiaohong Jia, Wenping Wang, Yi-King Choi, Bernard Mourrain, and Changhe Tu. Continuous detection of the variations of the intersection curve of two moving quadrics in 3-dimensional projective space. Journal of Symbolic Computation, 73(??): 221–243, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000474.

Jia:2010:STG

[JWG10] Xiaohong Jia, Haohao Wang, and Ron Goldman. Settheoretic generators of rational space curves. Journal of Symbolic Computation, 45(4): 414–433, April 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jefferson:2023:PRP

[JWW23] Christopher Jefferson, Rebecca Waldecker, and Wilf A. Wilson. Perfect refiners for permutation group backtracking algorithms. Journal of Symbolic Computation, 114 (??):18–36, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771712200027X.

Jing:2017:MAC

[JY17] Rui-Juan Jing and Chun-Ming Yuan. A modular algorithm to compute the generalized Hermite normal form for Z[x]-lattices. Journal of Symbolic Computation, 81(??):97–118, July/August 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301699.

Joswig:2004:CHO

[JZ04] Michael Joswig and Günter M. Ziegler. Convex hulls, oracles, and homology. Journal of Symbolic Computation, 38 (4):1247–1259, October 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kremer:2021:FIC

[KÁ21] Gereon Kremer and Erika Ábrahám. Fully incremental cylindrical algebraic decomposition. Journal of Symbolic Computation, 100(??):11–37, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300847.

Kadijevich:2013:NCI

[Kad13] Djordje M. Kadijevich. Neglected critical issues of effective CAS utilization. Journal of Symbolic Computation, 61–62(??):85–99, ???? 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001326.

Kahrs:1995:CCT

[Kah95] Stefan Kahrs. Confluence of curried term-rewriting systems. Journal of Symbolic Computation, 19(6):601–623, June 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kaltofen:1985:FPA

[Kal85] Erich Kaltofen. Fast parallel absolute irreducibility testing. Journal of Symbolic Computation, 1(1):57–67, March 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See errata [Kal90].

Kaltofen:1987:DIT

[Kal87] Erich Kaltofen. Deterministic irreducibility testing of polynomials over large finite fields. Journal of Symbolic Computation, 4(1):77–82, August 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kaltofen:1990:EFP

[Kal90] Erich Kaltofen. Errata to: “Fast parallel absolute irreducibility testing” [J. Symbolic Comput. 1 (1985), no. 1, 57–67, MR 87b:12002]. Journal of Symbolic Computation, 9(3):320, 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [Kal85].

Kalkbrener:1993:GEA

[Kal93] Michael Kalkbrener. A generalized Euclidean algorithm for computing triangular representations of algebraic varieties. Journal of Symbolic Computation, 15(2):143– 168 (or 143–167??), February 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kalkbrener:1994:PDR

[Kal94] Michael Kalkbrener. Prime decompositions of radicals in polynomial rings. Journal of Symbolic Computation, 18(4): 365–372, October 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kalkbrener:1997:SGB

[Kal97a] Michael Kalkbrener. On the stability of Gröbner bases under specializations. Journal of Symbolic Computation, 24 (1):51–58, July 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kaltofen:1997:TCA

[Kal97b] E. Kaltofen. Teaching computational abstract algebra. Journal of Symbolic Computation, 23(5–6):503–516, May/ June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kalkbrener:1998:APP

[Kal98] M. Kalkbrener. Algorithmic properties of polynomial rings. Journal of Symbolic Computation, 26(5):525– 582, November 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kalkbrener:1999:CGB

[Kal99]  Michael Kalkbrener.  On

the complexity of Gröbner bases conversion. Journal of Symbolic Computation, 28 (1–2):265–273, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0276;   http://www.
idealibrary.com/links/doi/

10.1006/jsco.1998.0276/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0276/ref.

Kaltofen:2000:CSC

[Kal00] Erich Kaltofen. Challenges of symbolic computation: My favorite open problems. Journal of Symbolic Computation, 29(6):891–919, June 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0370;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0370/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0370/ref.

Kalorkoti:2001:CGB

[Kal01a]  K. Kalorkoti.  Counting

and Gröbner bases. Journal of Symbolic Computation, 31(3):307–313, March 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.1575;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.1575/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.1575/ref.

Kalorkoti:2001:DAC

[Kal01b] K. Kalorkoti. Detecting algebraic curves in bad position.

Journal of Symbolic Computation, 31(6):671–690, June 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0424;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0424/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0424/ref.

Kalorkoti:2002:PZC

[Kal02] K. Kalorkoti. Predicting zero coefficients in formal power series computations. Journal of Symbolic Computation, 33(3):307–320, March 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0495;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0495/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0495/ref.

Kalorkoti:2011:MCM

[Kal11] K. Kalorkoti. Model checking in the modal μ-calculus and generic solutions. Journal of Symbolic Computation, 46(5):584–594, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711000177X.

Kaltofen:2021:F

[Kal21] Erich L. Kaltofen. Foreword. Journal of Symbolic Computation, 105(??):1–3, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300213.

Kantor:1991:FCF

[Kan91] William M. Kantor. Finding composition factors of permutation groups of degree n ≤ 10^{6}. Journal of Symbolic Computation, 12(4–5): 517–526, October/November 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Kapur:1986:UGB

[Kap86]  Deepak Kapur.  Using

Gröbner bases to reason about geometry problems. Journal of Symbolic Computation, 2 (4):399–408, December 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kaplan:1987:SCT

[Kap87] Stéphane Kaplan. Simplifying conditional term rewriting systems: unification, termination and confluence. Journal of Symbolic Computation, 4(3):295–334, December 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kapur:2006:PCP

[Kap06] Deepak Kapur. Preface on the contributed papers. Journal of Symbolic Computation, 41 (3–4):259–260, March/April 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Karr:1985:TSF

[Kar85] Michael Karr. Theory of summation in finite terms. Journal of Symbolic Computation, 1(3):303–316 (or 303– 315??), September 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kotsireas:2005:FSI

[KASW05] Ilias S. Kotsireas, Alkiviadis G.

Akritas, Stanly L. Steinberg, and Michael J. Wester. Foreword to the special issue on applications of computer algebra. Journal of Symbolic Computation, 40 (4–5):1129–1130, October/ November 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kauers:2006:SPM

[Kau06] Manuel Kauers. SumCracker:

a package for manipulating symbolic sums and related objects. Journal of Symbolic Computation, 41(9): 1039–1057, September 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kauers:2007:SAS

[Kau07]  Manuel Kauers.  Summa-

tion algorithms for Stirling number identities. Journal of Symbolic Computation, 42 (10):948–970, October 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kovacs:2024:CTF

[KBRV24] Zoltán Kovács, Christopher Brown, Tomás Recio, and Róbert Vajda. Computing with Tarski formulas and semi-algebraic sets in a web browser. Journal of Symbolic Computation, 120(??):Article 102235, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000494.

Khanin:2001:PPA

[KC01] Raya Khanin and Matthew Cartmell. Parallelization of perturbation analysis: Application to large-scale engineering problems. Journal of Symbolic Computation, 31(4):461–473, April 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0434;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0434/
pdf;   http://www.idealibrary.

com/links/doi/10.1006/jsco.
1999.0434/ref.

Kunkle:2009:HPD

[KC09] Daniel Kunkle and Gene Cooperman. Harnessing parallel disks to solve Rubik’s cube. Journal of Symbolic Computation, 44(7):872–890, July 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Koprowski:2018:CQF

[KC18] Przemyslaw Koprowski and Alfred Czogala. Computing with quadratic forms over number fields. Journal of Symbolic Computation, 89(??):129–145, November/December 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117301165.

Kwong:1990:CAS

[KD90] Man Kam Kwong and Mary Beth Dever. Computer-aided study of a problem in Hermitian matrix theory. Journal of Symbolic Computation, 9(1): 87–112, January 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kemper:1996:CIR

[Kem96] Gregor Kemper. Calculating invariant rings of finite groups over arbitrary fields. Journal of Symbolic Computation, 21 (3):351–366, March 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kemper:1999:ACO

[Kem99] G. Kemper. An algorithm to calculate optimal homogeneous systems of parameters. Journal of Symbolic Computation, 27(2):171–184, February 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kemper:2002:CRI

[Kem02] Gregor Kemper. The calculation of radical ideals in positive characteristic. Journal of Symbolic Computation, 34 (3):229–238, September 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kemper:2009:SI

[Kem09] Gregor Kemper. Separating invariants. Journal of Symbolic Computation, 44(9): 1212–1222, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kemper:2016:UED

[Kem16] Gregor Kemper. Using extended Derksen ideals in computational invariant theory. Journal of Symbolic Computation, 72(??): 161–181, January/February 2016. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000139.

Kemper:2022:CQC

[Kem22] Gregor Kemper. Computing quotients by connected solvable groups. Journal of Symbolic Computation, 109(??): 426–440, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300699.

Kerber:2017:SIA

[Ker17] Michael Kerber. Special issue on algorithms and software for computational topology. Journal of Symbolic Computation, 78(??):1– 2, January/February 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300050.

Key:2001:SAM

[Key01] J. D. Key. Some applications of Magma in designs and codes: Oval designs, Hermitian unitals and generalized Reed–Muller codes. Journal of Symbolic Computation, 31(1– 2):37–53, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.

com/links/doi/10.1006/jsco.
1999.1007;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.1007/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1007/ref.

Kohlhase:2001:MRK

[KF01] Michael Kohlhase and Andreas Franke. MBase: Representing knowledge and context for the integration of mathematical software systems. Journal of Symbolic Computation, 32(4): 365–402, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0468;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0468/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0468/ref.

Kobayashi:1988:SSA

[KFF88] Hidetsune Kobayashi, Tetsuro Fujise, and Akio Furukawa. Solving systems of algebraic equations by a general elimination method. Journal of Symbolic Computation, 5(3): 303–320, June 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kaji:1997:SUP

[KFK97] Yuichi Kaji, Toru Fujiwara, and Tadao Kasami. Solving a unification problem under constrained substitutions using tree automata. Journal of Symbolic Computation, 23 (1):79–118 (or 79–114??), January 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Koepf:2003:PCA

[KG03] Wolfram Koepf and Karin Gatermann. Preface to computer algebra and computer analysis. Journal of Symbolic Computation, 35(5):463–464, May 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kudo:2023:RNS

[KH23] Momonari Kudo and Shushi Harashita. Representation of non-special curves of genus 5 as plane sextic curves and its application to finding curves with many rational points. Journal of Symbolic Computation, 122(??):Article 102272, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712300086X.

Khan:2014:CSB

[Kha14] Junaid Alam Khan. Converting subalgebra bases with the Sagbi walk. Journal of Symbolic Computation, 60 (??):78–93, January 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711300120X.

Khetan:2003:RUB

[Khe03] Amit Khetan. The resultant of an unmixed bivariate system. Journal of Symbolic Computation, 36(3–4): 425–442, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Khoury:2008:GBA

[Kho08] Joseph Khoury. A Gröbner basis approach to solve a Conjecture of Nowicki. Journal of Symbolic Computation, 43 (12):908–922, December 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kida:2002:PGR

[Kid02] Masanari Kida. Potential good reduction of elliptic curves. Journal of Symbolic Computation, 34(3):173– 180, September 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

King:2013:MGS

[Kin13]  Simon A. King.  Mini-

mal generating sets of nonmodular invariant rings of finite groups. Journal of Symbolic Computation, 48 (??):101–109, January 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711200079X.

King:2014:NCI

[Kin14] Simon A. King. A noncommutative F_{5} algorithm with an application to the computation of Loewy layers. Journal of Symbolic Computation, 65(??):111–129, November 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000169.

Krandick:1996:BEI

[KJ96] Werner Krandick and Tudor Jebelean. Bidirectional exact integer division. Journal of Symbolic Computation, 21(4/5/6):441–456 (or 441– 455??), April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Kotsireas:2009:HMW

[KK09] Ilias S. Kotsireas and Christos Koukouvinos. Hadamard matrices of Williamson type: a challenge for Computer Algebra. Journal of Symbolic Computation, 44(3): 271–279, March 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Keicher:2017:TMC

[KK17] Simon Keicher and Thomas Kremer. A test for monomial containment. Journal of Symbolic Computation, 82(??):74– 90, September/October 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300019.

Kubota:2016:SAV

[KKK^{+}16] Takahiro Kubota, Yoshihiko Kakutani, Go Kato, Yasuhito Kawano, and Hideki Sakurada. Semi-automated verification of security proofs of quantum cryptographic protocols. Journal of Symbolic Computation, 73(??): 192–220, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000462.

King:2017:BDD

[KKK17] Henry King, Kevin Knudson, and Neza Mramor Kosta. Birth and death in discrete Morse theory. Journal of Symbolic Computation, 78(??):41– 60, January/February 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300074.

Kerber:1992:SOO

[KKL92] Adalbert Kerber, Axel Kohnert, and Alain Lascoux. SYM-METRICA, an object oriented computer-algebra system for the symmetric group. Journal of Symbolic Computation, 14(2–3):195–204 (or 195– 203??), August/September 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kytmanov:2015:FRI

[KKM15] A. A. Kytmanov, A. M. Kytmanov, and E. K. Myshkina. Finding residue integrals for systems of non-algebraic equations in C^{n}. Journal of Symbolic Computation, 66(??):98– 110, January/February 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000170.

Kennaway:1996:CCU

[KKSd96] Richard Kennaway, Jan Willem Klop, Ronan Sleep, and Fer-Jan de Vries. Comparing curried and uncurried rewriting. Journal of Symbolic Computation, 21(1):15–40 (or 15– 39??), January 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kozen:1989:PDA

[KL89] Dexter Kozen and Susan Landau. Polynomial decomposition algorithms. Journal of Symbolic Computation, 7(5): 445–456, May 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See errata [KL90].

Kozen:1990:EPD

[KL90] D. Kozen and S. Landau. Errata: “Polynomial decomposition algorithms” [J. Symbolic Comput. 7 (1989), no. 5, 445–456, MR 91c:13022]. Journal of Symbolic Computation, 10(5):529, 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [KL89].

Karmarkar:1998:AGU

[KL98a] N. K. Karmarkar and Y. N. Lakshman. On approximate GCDs of univariate polynomials. Journal of Symbolic Computation, 26(6):653– 666, December 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kocbach:1998:GVA

[KL98b] Ladislav Kocbach and Richard Liska. Generation and verification of algorithms for symbolic-numeric processing. Journal of Symbolic Computation, 25(3):367–382, March 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Krone:2017:EDS

[KL17a] Robert Krone and Anton Leykin. Eliminating dual spaces. Journal of Symbolic Computation, 79 (part 3)(??):609–622, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300591.

Krone:2017:NAD

[KL17b] Robert Krone and Anton Leykin. Numerical algorithms for detecting embedded components. Journal of Symbolic Computation, 82(??):1– 18, September/October 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301663.

Kalisnik:2019:SPT

[KL19] Sara Kalisnik and Davorin Lesnik. Symmetric polynomials in tropical algebra semirings. Journal of Symbolic Computation, 93(??): 100–119, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300373.

Kalisnik:2021:SPU

[KL21] Sara Kalisnik and Davorin Lesnik. Symmetric polynomials in upper-bound semirings. Journal of Symbolic Computation, 103(??): 280–299, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300092.

Klingen:1990:LCI

[Kli90] Norbert Klingen. Leopoldt’s conjecture for imaginary Galois number fields. Journal of Symbolic Computation, 10(6):531–546 (or 531– 545??), December 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kapur:2021:ACG

[KLM^{+}21] Deepak Kapur, Dong Lu, Michael Monagan, Yao Sun, and Dingkang Wang. Algorithms for computing greatest common divisors of parametric multivariate polynomials. Journal of Symbolic Computation, 102(??):3– 20, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301129.

Klimov:1993:SMG

[KLR93] D. M. Klimov, V. V. Leonov, and V. M. Rudenko. The study of motion for a gyroscope with gimbal suspension: obtaining the highest approximations for a drift of Magnus. Journal of Symbolic Computation, 15(1):73–78, January 1993. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Kluners:1999:PD

[Klü99] J. Klüners. On polynomial decompositions. Journal of Symbolic Computation, 27(3): 261–270, March 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kluners:2000:PGG

[Klü00] Jürgen Klüners. A polynomial with Galois group SL_{2}(11). Journal of Symbolic Computation, 30(6): 733–737, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0380;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0380/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0380/ref.

Kutsia:2010:RBC

[KLV10] Temur Kutsia, Jordi Levy, and Mateu Villaret. On the relation between Context and Sequence Unification. Journal of Symbolic Computation, 45 (1):74–95, January 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kaltofen:2012:ECG

[KLYZ12] Erich L. Kaltofen, Bin Li, Zhengfeng Yang, and Lihong Zhi. Exact certification in global polynomial optimization via sums-of-squares of rational functions with rational coefficients. Journal of Symbolic Computation, 47(1):1–15, January 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001143.

Kozen:1996:DAF

[KLZ96] Dexter Kozen, Susan Landau, and Richard Zippel. Decomposition of algebraic functions. Journal of Symbolic Computation, 22(3):235– 246, September 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Klin:2012:LBT

[KLZA12] Mikhail Klin, Josef Lauri, and Matan Ziv-Av. Links between two semisymmetric graphs on 112 vertices via association schemes. Journal of Symbolic Computation, 47(10):1175–1191, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002331.

Klop:1991:SOT

[KM91] Jan Willem Klop and Aart Middeldorp. Sequentiality in orthogonal term rewriting systems. Journal of Symbolic Computation, 12(2):161– 196 (or 161–195??), August 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kajler:1998:SIG

[KM98] Norbert Kajler and Michael Monagan. Special issue on graphical user interfaces and protocols foreword of the Guest Editors. Journal of Symbolic Computation, 25(2): 125, February 1998.

Koppenhagen:1999:OAC

[KM99] Ulla Koppenhagen and Ernst W.

Mayr. An optimal algorithm for constructing the reduced Gröbner basis of binomial ideals. Journal of Symbolic Computation, 28(3):317–338, September 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0285/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0285/production/pdf;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0285/production/ref.

Kemper:2000:GPF

[KM00a] Gregor Kemper and Elena Mattig. Generic polynomials with few parameters. Journal of Symbolic Computation, 30(6):843–857, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0385;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0385/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0385/ref.

Kluners:2000:EGR

[KM00b] Jürgen Klüners and Gunter Malle. Explicit Galois realization of transitive groups of degree up to 15. Journal of Symbolic Computation, 30(6):675–716, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0378;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0378/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0378/ref.

Koppenhagen:2001:OAC

[KM01] Ulla Koppenhagen and Ernst W. Mayr. An optimal algorithm for constructing the reduced Gröbner basis of binomial ideals, and applications to commutative semigroups. Journal of Symbolic Computation, 31(1–2): 259–276, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1015;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.1015/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1015/ref.

Krandick:2006:NBD

[KM06] Werner Krandick and Kurt Mehlhorn. New bounds for the Descartes method. Journal of Symbolic Computation, 41(1): 49–66, January 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kobayashi:1989:RZD

[KMH89] Hidetsune Kobayashi, Shuichi Moritsugu, and Robert W. Hogan. On radial zerodimensional ideals. Journal of Symbolic Computation, 8(6): 545–552, December 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Khouja:2022:TDL

[KMM22] Rima Khouja, Pierre-Alexandre Mattei, and Bernard Mourrain. Tensor decomposition for learning Gaussian mixtures from moments. Journal of Symbolic Computation, 113(??):193–210, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000232.

Kapur:1988:OPS

[KMN88] Deepak Kapur, David R. Musser, and Paliath Narendran. Only prime superpositions need be considered in the Knuth–Bendix completion procedure. Journal of Symbolic Computation, 6(1): 19–36, August 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Koprowski:2018:CFA

[KMR18] Przemysław Koprowski, Kurt Mehlhorn, and Saurabh Ray. Corrigendum to “Faster algorithms for computing Hong’s bound on absolute positiveness” [J. Symb. Comput. 45 (2010) 677–683]. Journal of Symbolic Computation, 87(??):238–241, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300597.

See [MR10].

Kotsireas:2024:AAB

[KMY24] Ilias Kotsireas, Toufik Mansour, and Gökhan Yildirim. An algorithmic approach based on generating trees for enumerating pattern-avoiding inversion sequences. Journal of Symbolic Computation, 120(??):Article 102231, January/February 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712300038X.

Kaltofen:2008:AFM

[KMYZ08] Erich Kaltofen, John P. May, Zhengfeng Yang, and Lihong Zhi. Approximate factorization of multivariate polynomials using singular value decomposition. Journal of Symbolic Computation, 43(5): 359–376, May 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kawazoe:2011:ACP

[KN11] Taro Kawazoe and Masayuki Noro. Algorithms for computing a primary ideal decomposition without producing intermediate redundant components. Journal of Symbolic Computation, 46(10): 1158–1172, October 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711100085X.

Knowles:1992:ICT

[Kno92] Paul H. Knowles. Integration of a class of transcendental Liouvillian functions with error-functions, Part I. Journal of Symbolic Computation, 13(5):525–544 (or 525–543??), May 1992. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-

855X (electronic).

Knowles:1993:ICT

[Kno93] Paul H. Knowles. Integration of a class of transcendental Liouvillian functions with error-functions, Part II. Journal of Symbolic Computation, 16(3):227–242 (or 227– 241??), September 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kapur:1991:AII

[KNZ91] Deepak Kapur, Paliath Narendran, and Hantao Zhang. Automating inductionless induction using test sets. Journal of Symbolic Computation, 11(1–2):83–112 (or 83–111??), January/February 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kelley:2017:ENR

[KO17] Zander Kelley and Sean W. Owen. Estimating the number of roots of trinomials over finite fields. Journal of Symbolic Computation, 79 (part 1)(??):108–118, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300876.

Koepf:1992:PSC

[Koe92] Wolfram Koepf. Power series in computer algebra. Journal of Symbolic Computation, 13(6):581–604 (or 581–603??), June 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Koepf:1995:AFH

[Koe95] Wolfram Koepf. Algorithms for m-fold hypergeometric summation. Journal of Symbolic Computation, 20(4):399– 418 (or 399–417??), October 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kohnert:1992:SPS

[Koh92]  Axel Kohnert.  Schubert

polynomials and skew Schur functions. Journal of Symbolic Computation, 14(2–3): 205–210, August/September 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kohl:2008:ACI

[Koh08] Stefan Kohl. Algorithms for a class of infinite permutation groups. Journal of Symbolic Computation, 43(8):545– 581, August 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kohn:2021:CHG

[Koh21] Kathlén Kohn. Coisotropic hypersurfaces in Grassmannians. Journal of Symbolic Computation, 103(??): 157–177, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301506.

Koiran:2019:RST

[Koi19] Pascal Koiran. Root separation for trinomials. Journal of Symbolic Computation, 95(??):151–161, November/December 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711930015X.

Kolbig:1985:EEC

[Köl85] K. S. Kölbig. Explicit evaluation of certain definite integrals involving powers of logarithms. Journal of Symbolic Computation, 1(1):109– 114, March 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kolesnikov:2008:UDR

[Kol08] Pavel Kolesnikov. Universally defined representations of Lie conformal superalgebras. Journal of Symbolic Computation, 43(6–7):406– 421, June/July 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Konieczny:1995:RIS

[Kon95] Janusz Konieczny. Reduced idempotents in the semigroup of Boolean matrices. Journal of Symbolic Computation, 20(4):471–482, October 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Konig:2017:RFM

[Kön17] Joachim König. On rational functions with monodromy group M_{11}. Journal of Symbolic Computation, 79 (part 2)(??):372–383, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000237.

Koprowski:2008:AQF

[Kop08] Przemysław Koprowski. Algorithms for quadratic forms. Journal of Symbolic Computation, 43(2):140–152, February 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Koshita:2007:ERE

[Kos07] Hitoshi Koshita. An example of relations on the Ext-quiver for the Suzuki group Sz(8) in characteristic 2. Journal of Symbolic Computation, 42(4): 429–442, April 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kauers:2021:F

[KOS21] Manuel Kauers, Alexey Ovchinnikov, and Eric Schost. Foreword. Journal of Symbolic Computation, 102(??): 1–2, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301117.

Kovacic:1986:ASS

[Kov86] Jerald J. Kovacic. An algorithm for solving second order linear homogeneous differential equations. Journal of Symbolic Computation, 2 (1):3–43, March 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kozhasov:2023:ESM

[Koz23]  Khazhgali Kozhasov.  On

eigenvalues of symmetric matrices with PSD principal submatrices. Journal of Symbolic Computation, 119(??): 90–100, November/December 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000159.

Krantz:1991:AED

[KP91] Steven G. Krantz and William H. Paulsen. Asymptotic eigenfrequency distributions for the N-beam Euler–Bernoulli coupled beam equation with dissipative joints. Journal of Symbolic Computation, 11(4): 369–418, April 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kluners:1997:CS

[KP97a] J. Klüners and M. Pohst. On computing subfields. Journal of Symbolic Computation, 24(3–4):385–398, September/ October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kluners:1997:CSC

[KP97b] Jürgen Klüners and Michael Pohst. On computing subfields. Journal of Symbolic Computation, 24(3–4):385– 397, 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Krysta:1999:SPN

[KP99] P. Krysta and L. Pacholski.

The STO problem is NPcomplete. Journal of Symbolic Computation, 27(2):207– 220, February 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kadioglu:2013:CMN

[KP13] Hülya Kadioglu and Tracy L.

Payne. Computational methods for nilsoliton metric Lie algebras I. Journal of Symbolic Computation, 50 (??):350–373, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001381.

Koseleff:2015:ACP

[KP15] Pierre-Vincent Koseleff and Daniel Pecker. On Alexander– Conway polynomials of twobridge links. Journal of Symbolic Computation, 68 (part 2) (??):215–229, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000790.

Koseleff:2010:FRC

[KPR10] P.-V. Koseleff, D. Pecker, and F. Rouillier. The first rational Chebyshev knots. Journal of Symbolic Computation, 45(12):1341–1358, December 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Koseleff:2018:CCK

[KPRT18] P.-V. Koseleff, D. Pecker, F. Rouillier, and C. Tran. Computing Chebyshev knot diagrams. Journal of Symbolic Computation, 86(??): 120–141, May/June 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300482.

Koiran:2015:WAR

[KPT15] Pascal Koiran, Natacha Portier, and Sébastien Tavenas. A Wronskian approach to the real τ-conjecture. Journal of Symbolic Computation, 68 (part 2)(??):195–214, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001047.

Kanellakis:1989:RCC

[KR89] Paris C. Kanellakis and Peter Z. Z. Revesz. On the relationship of congruence closure and unification. Journal of Symbolic Computation, 7(3–4):427–444, March/April 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kounalis:1991:WPH

[KR91] Emmanuel Kounalis and Michael Rusinowitch. On word problems in Horn theories. Journal of Symbolic Computation, 11(1–2):113– 128 (or 113–127??), January/February 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kirchner:1994:CSC

[KR94] Hélène Kirchner and Christophe Ringeissen. Combining symbolic constraint solvers on algebraic domains. Journal of Symbolic Computation, 18(2): 113–156 (or 113–155??), August 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Krandick:1997:SIV

[KR97] W. Krandick and S. Rump. Special issue on validated numerical methods and computer algebra: Foreword of the Guest Editors. Journal of Symbolic Computation, 24(6): 625–626, December 1997.

Kahrs:2022:SRE

[KR22] Stefan Kahrs and Colin Runciman. Simplifying regular expressions further. Journal of Symbolic Computation, 109(??):124–143, March/April 2022. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000572.

Koprowski:2023:APQ

[KR23] Przemysław Koprowski and Beata Rothkegel. The anisotropic part of a quadratic form over a number field. Journal of Symbolic Computation, 115(??):39–52, March/ April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771712200061X.

Krattenthaler:1995:HHM

[Kra95] C. Krattenthaler. HYP and HYPQ: Mathematica packages for the manipulation of binomial sums and hypergeometric series, respectively q-binomial sums and basic hypergeometric series. Journal of Symbolic Computation, 20(5–6):737–744, November/December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Krishnamurthy:1985:SIA

[Kri85] E. V. Krishnamurthy. Symbolic iterative algorithm for generalized inversion of rational polynomial matrices. Journal of Symbolic Computation, 1(3):271–282 (or 271– 281??), September 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kandri-Rody:1988:CGB

[KRK88] Abdelilah Kandri-Rody and Deepak Kapur. Computing a Gröbner basis of a polynomial ideal over a Euclidean domain. Journal of Symbolic Computation, 6(1):37–58 (or 37–57??), August 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Katsamaki:2023:PCG

[KRTZ23] Christina Katsamaki, Fabrice Rouillier, Elias Tsigaridas, and Zafeirakis Zafeirakopoulos. PTOPO: Computing the geometry and the topology of parametric curves. Journal of Symbolic Computation, 115 (??):427–451, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000785.

Kovacs:2019:RAL

[KRV19] Zoltán Kovács, Tomás Recio, and M. Pilar Vélez. Reasoning about linkages with dynamic geometry. Journal of Symbolic Computation, 97(??):16–30, ???? 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118301299.

Kandri-Rody:1990:NCG

[KRW90] A. Kandri-Rody and V. Weispfenning. Non-commutative Gröbner bases in algebras of solvable type. Journal of Symbolic Computation, 9(1): 1–26, January 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kutzler:1986:ABA

[KS86] B. Kutzler and S. Stifter. On the application of Buchberger’s algorithm to automated geometry theorem proving. Journal of Symbolic Computation, 2(4):389– 397, December 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kozen:1997:CNG

[KS97] Dexter Kozen and Kjartan Stefánsson. Computing the Newtonian graph. Journal of Symbolic Computation, 24(2):125–136, August 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Applications of quantifier elimination (Albuquerque, NM, 1995).

Kajler:1998:SUI

[KS98] N. Kajler and N. Soiffer.

A survey of user interfaces for computer algebra systems. Journal of Symbolic Computation, 25(2):127–160, February 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Koussoulas:2004:SCM

[KS04] Nick T. Koussoulas and Photis Skiadas. Symbolic computation for mobile robot path planning. Journal of Symbolic Computation, 37(6):761–775, June 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kanno:2006:VNC

[KS06] Masaaki Kanno and Malcolm C. Smith. Validated numerical computation of the L_{∞}-norm for linear dynamical systems. Journal of Symbolic Computation, 41(6):697–707, June 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Katzman:2012:ACC

[KS12a] Mordechai Katzman and Karl Schwede. An algorithm for computing compatibly Frobenius split subvarieties. Journal of Symbolic Computation, 47(8):996–1008, August 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112000260.

Kerber:2012:WCB

[KS12b] Michael Kerber and Michael Sagraloff. A worst-case bound for topology computation of algebraic curves. Journal of Symbolic Computation, 47(3):239–258, March 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001775.

Krick:2012:SDS

[KS12c] Teresa Krick and Agnes Szanto. Sylvester’s double sums: an inductive proof of the general case. Journal of Symbolic Computation, 47(8):942–953, August 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112000041.

Kawano:2016:QFT

[KS16] Yasuhito Kawano and Hiroshi Sekigawa. Quantum Fourier transform over symmetric groups — improved result. Journal of Symbolic Computation, 75(??): 219–243, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001157.

Kitson:2018:MGL

[KS18] D. Kitson and B. Schulze.

Motions of grid-like reflection frameworks. Journal of Symbolic Computation, 88(??):47– 66, September/October 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300142.

Kaur:2019:IFT

[KS19] Yashpreet Kaur and Varadharaj R. Srinivasan. Integration in finite terms with dilogarithmic integrals, logarithmic integrals and error functions. Journal of Symbolic Computation, 94(??): 210–233, September/October 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300956.

Krumm:2021:GGR

[KS21] David Krumm and Nicole Sutherland. Galois groups over rational function fields and explicit Hilbert irreducibility. Journal of Symbolic Computation, 103(??): 108–126, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301476.

Kosta:2016:BAR

[KSD16] Marek Kosta, Thomas Sturm, and Andreas Dolzmann. Better answers to real questions. Journal of Symbolic Computation, 74(??):255–275, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000711.

Kaltofen:2003:ETS

[KsL03] Erich Kaltofen and Wen shin Lee. Early termination in sparse interpolation algorithms. Journal of Symbolic Computation, 36(3–4): 365–400, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kalkbrener:1993:LDS

[KST93] M. Kalkbrener, M. Sweedler, and L. Taylor. Low degree solutions to linear equations with K(x) coefficients. Journal of Symbolic Computation, 16(1):75–81, July 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kapur:2013:EAC

[KSW13a] Deepak Kapur, Yao Sun, and Dingkang Wang. An efficient algorithm for computing a comprehensive Gröbner system of a parametric polynomial system. Journal of Symbolic Computation, 49 (??):27–44, February 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002082.

Kapur:2013:EMC

[KSW13b] Deepak Kapur, Yao Sun, and Dingkang Wang. An efficient method for computing comprehensive Gröbner bases. Journal of Symbolic Computation, 52(??):124–142, May 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001320.

Kaltofen:1990:CPG

[KT90a] Erich Kaltofen and Barry M.

Trager. Computing with polynomials given by black boxes for their evaluations: Greatest common divisors, factorization, separation of numerators and denominators. Journal of Symbolic Computation, 9(3):300–320 (or 301– 320??), March 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kean:1990:IMG

[KT90b] Alex Kean and George Tsiknis. An incremental method for generating prime implicants/implicates. Journal of Symbolic Computation, 9(2): 185–206, February 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See corrigendum [KT94].

Kean:1994:COI

[KT94] Alex Kean and George Tsiknis. A corrigendum for the optimized-IPIA: “An incremental method for generating prime implicants/ implicates” [J. Symbolic Comput. 9 (1990), no. 2, 185– 206, MR 91h:68071]. Journal of Symbolic Computation, 17(2):181–188 (or 181– 187??), February 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [KT90b].

Kaplan:2002:IHT

[KT02] S. Kaplan and M. Teicher. Identifying half-twists using randomized algorithm methods. Journal of Symbolic Computation, 34(2):91– 103, August 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kirschenhofer:2004:ESN

[KT04] Peter Kirschenhofer and Jörg M. Thuswaldner. Elements of small norm in Shanks’ cubic extensions of imaginary quadratic fields. Journal of Symbolic Computation, 38(6):1471–1486, December 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Krupchyk:2008:COP

[KT08] Katsiaryna Krupchyk and Jukka Tuomela. Completion of overdetermined parabolic PDEs. Journal of Symbolic Computation, 43(3):153– 167, March 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Krityakierne:2023:CGG

[KT23] Tipaluck Krityakierne and Thotsaporn Aek Thanatipanonda. The card guessing game: a generating function approach. Journal of Symbolic Computation, 115 (??):1–17, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000529.

Kudo:2022:CRM

[Kud22] Momonari Kudo. Computing representation matrices for the action of Frobenius on cohomology groups. Journal of Symbolic Computation, 109 (??):441–464, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300705.

Kunkle:2018:FEI

[Kun18] Thomas Kunkle. Flat extension and ideal projection. Journal of Symbolic Computation, 89(??):109–120, November/December 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117301141.

Kuo:2006:VSP

[Kuo06] Eric H. Kuo. Viterbi sequences and polytopes. Journal of Symbolic Computation, 41(2):151–163, February 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kutsia:2007:SES

[Kut07] Temur Kutsia. Solving equations with sequence variables and sequence functions. Journal of Symbolic Computation, 42(3):352–388, March 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kutsia:2008:FM

[Kut08] Temur Kutsia. Flat matching. Journal of Symbolic Computation, 43(12):858–873, December 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kutsia:2010:SCS

[Kut10]  Temur Kutsia.  Symbolic

computation in software science: Foreword from the Editor. Journal of Symbolic Computation, 45(5):499–500, May 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kutas:2019:SQA

[Kut19]  Péter Kutas.  Splitting

quaternion algebras over quadratic number fields. Journal of Symbolic Computation, 94(??):173–182, September/October 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300932.

Kredel:1988:CDI

[KW88] Heinz Kredel and Volker Weispfenning. Computing dimension and independent sets for polynomial ideals. Journal of Symbolic Computation, 6 (2–3):231–248 (or 231–247??), October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Kauer:2010:MII

[KW10] Stefan Kauer and Jürgen F. H. Winkler. Mechanical inference of invariants for FORloops. Journal of Symbolic Computation, 45(11):1101– 1113, November 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Koepf:2013:GEF

[KW13] Wolfram Koepf and Stephen M. Watt. Guest Editors’ foreword. Journal of Symbolic Computation, 49(??): 1–2, February 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002069.

Kreuzer:2024:CBP

[KW24] Martin Kreuzer and Florian Walsh. Computing the binomial part of a polynomial ideal. Journal of Symbolic Computation, 124(??): ??, September/October 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717124000026.

Kauers:2015:LIT

[KY15] Manuel Kauers and Lily Yen. On the length of integers in telescopers for proper hypergeometric terms. Journal of Symbolic Computation, 66 (??):21–33, January/February 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000224.

Kaltofen:2016:SMF

[KY16] Erich L. Kaltofen and Zhengfeng Yang. Sparse multivariate function recovery with a small number of evaluations. Journal of Symbolic Computation, 75(??):209–218, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001145.

Kauers:2008:CAR

[KZ08] Manuel Kauers and Burkhard Zimmermann. Computing the algebraic relations of C-finite sequences and multisequences. Journal of Symbolic Computation, 43(11):787–803, November 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Konnov:2010:IBA

[KZ10] Igor V. Konnov and Vladimir A.

Zakharov. An invariant-based approach to the verification of asynchronous parameterized networks. Journal of Symbolic Computation, 45(11):1144– 1162, November 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Katzman:2014:AAM

[KZ14] Mordechai Katzman and Wenliang Zhang. Annihilators of Artinian modules compatible with a Frobenius map. Journal of Symbolic Computation, 60(??):29–46, January 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001284.

LaScala:2017:MRI

[La 17] Roberto La Scala. Monomial right ideals and the Hilbert series of noncommutative modules. Journal of Symbolic Computation, 80 (part 2)(??):403–415, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300323.

Levy:1996:BRS

[LA96] Jordi Levy and Jaume Agustí. Bi-rewrite systems. Journal of Symbolic Computation, 22 (3):279–314, September 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Labonte:1990:ACM

[Lab90] Gilles Labonté. An algorithm for the construction of matrix representations for finitely presented non-commutative algebras. Journal of Symbolic Computation, 9(1):27– 38, January 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Labelle:1992:CAE

[Lab92] Gilbert Labelle. Counting asymmetric enriched trees.

Journal of Symbolic Computation, 14(2–3):211–242, August/September 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Labelle:1995:SCR

[Lab95] Gilbert Labelle. Some combinatorial results first found using computer algebra. Journal of Symbolic Computation, 20(5–6):567–594, November/ December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Lairez:2024:ATS

[Lai24] Pierre Lairez. Axioms for a theory of signature bases. Journal of Symbolic Computation, 123(??):??, July/August 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000895.

Lambe:1991:RHP

[Lam91] Larry A. Lambe. Resolutions via homological perturbation. Journal of Symbolic Computation, 12(1):71–88 (or 71–87??), July 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lambe:1997:E

[Lam97]  L. A. Lambe.  Editorial.

Journal of Symbolic Computation, 23(5–6):445–446, May/ June 1997. Special Issue on Symbolic Computation in Upper Level Undergraduate Education.

Landau:1992:NZD

[Lan92] Susan Landau. A note on “Zippel denesting”. Journal of Symbolic Computation, 13 (1):41–46 (or 41–45??), January 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Landsberg:2010:PVN

[Lan10] J. M. Landsberg. P versus NP and geometry. Journal of Symbolic Computation, 45(12):1359–1377, December 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Laville:1991:CPR

[Lav91] Alain Laville. Comparison of priority rules in pattern matching and term rewriting. Journal of Symbolic Computation, 11(4):321–348 (or 321– 347??), April 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lazard:1985:IBP

[Laz85] D. Lazard. Ideal bases and primary decomposition: Case of two variables. Journal of Symbolic Computation, 1 (3):261–270, September 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lazard:1988:QEO

[Laz88] Daniel Lazard. Quantifier elimination: Optimal solutions for two classical examples. Journal of Symbolic Computation, 5(1–2):261–266, February/April 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lazard:1992:NUB

[Laz92a] D. Lazard. A note on upper bounds for ideal-theoretic problems. Journal of Symbolic Computation, 13(3):231– 234 (or 231–233??), March 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lazard:1992:SZD

[Laz92b] D. Lazard. Solving zerodimensional algebraic systems. Journal of Symbolic Computation, 13(2):117– 132 (or 117–131??), February 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lazard:2009:TYP

[Laz09] Daniel Lazard. Thirty years of Polynomial System Solving, and now? Journal of Symbolic Computation, 44(3):

222–231, March 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lowe:1998:UPP

[LBM98] H. Lowe, A. Bundy, and D. McLean. The use of proof planning for co-operative theorem proving. Journal of Symbolic Computation, 25(2): 239–262, February 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lincoln:1989:AAC

[LC89] Patrick Lincoln and Jim Christian. Adventures in associative-commutative unification. Journal of Symbolic Computation, 8(1–2): 217–240, July/August 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Also appears in Unification, edited by Claude Kirchner (Academic, 1990), pages 393–416.

Leung:1996:CSD

[LC96] Ho-Fung Leung and Keith L. Clark. Constraint satisfaction in distributed concurrent logic programming. Journal of Symbolic Computation, 21(4/5/6):699–714, April, May & June 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Lai:2016:IRS

[LC16] Yisheng Lai and Falai Chen.

Implicitizing rational surfaces using moving quadrics constructed from moving planes. Journal of Symbolic Computation, 77(??):127–161, November/December 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000092.

Luo:2010:VPS

[LCQ^{+}10] Chenguang Luo, Florin Craciun, Shengchao Qin, Guanhua He, and Wei-Ngan Chin. Verifying pointer safety for programs with unknown calls. Journal of Symbolic Computation, 45(11):1163–1183, November 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

LeChenadec:1986:CCG

[Le 86] Philippe Le Chenadec. A catalogue of complete group presentations. Journal of Symbolic Computation, 2(4):363– 381, December 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

LeChenadec:1989:LU

[Le 89] Philippe Le Chenadec. On the logic of unification. Journal of Symbolic Computation, 8(1–2):141–199, July/August 1989. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-

855X (electronic).

Le:2022:SPS

[LE22] Huu Phuoc Le and Mohab Safey El Din. Solving parametric systems of polynomial equations over the reals through Hermite matrices. Journal of Symbolic Computation, 112(??): 25–61, September/October 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000833.

Leamer:2006:GFP

[Lea06] Micah J. Leamer. Gröbner finite path algebras. Journal of Symbolic Computation, 41(1): 98–111, January 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lebreton:2015:RHL

[Leb15] Romain Lebreton. Relaxed Hensel lifting of triangular sets. Journal of Symbolic Computation, 68 (part 2) (??):230–258, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000807.

Lecerf:2007:IDM

[Lec07] Grégoire Lecerf. Improved dense multivariate polynomial factorization algorithms.

Journal of Symbolic Computation, 42(4):477–494, April 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lecerf:2019:CLR

[Lec19] Grégoire Lecerf. On the complexity of the Lickteig– Roy subresultant algorithm. Journal of Symbolic Computation, 92(??):243–268, May/ June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300488.

Ledet:2000:GER

[Led00a] Arne Ledet. Generic and explicit realization of small p-groups. Journal of Symbolic Computation, 30(6): 859–865, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0386;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0386/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0386/ref.

Ledet:2000:GEG

[Led00b] Arne Ledet. Generic extensions and generic polynomials. Journal of Symbolic Computation, 30(6): 867–872, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0387;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.0387/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0387/ref.

Lee:2008:ADI

[Lee08] Dong Hoon Lee. Algebraic degree of the inverse of linearized polynomials. Journal of Symbolic Computation, 43(6–7):442–451, June/ July 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lee:2017:EDD

[Lee17] Hwangrae Lee. The Euclidean distance degree of Fermat hypersurfaces. Journal of Symbolic Computation, 80 (part 2)(??):502–510, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300396.

Lemaire:2003:OLP

[Lem03] François Lemaire. An orderly linear PDE system with analytic initial conditions with a non-analytic solution. Journal of Symbolic Computation, 35 (5):487–498, May 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Leon:1991:PGA

[Leo91] Jeffrey S. Leon. Permutation group algorithms based on partitions, I: Theory and algorithms. Journal of Symbolic Computation, 12(4–5): 533–583, October/November 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Lescanne:1992:WRO

[Les92] Pierre Lescanne. Well rewrite orderings and well quasiorderings. Journal of Symbolic Computation, 14(5):419– 436 (or 419–435??), November 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Letzter:2001:CIR

[Let01] Edward S. Letzter. Constructing irreducible representations of finitely presented algebras. Journal of Symbolic Computation, 32 (3):255–262, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0445;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0445/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0445/ref.

Levin:1999:CHP

[Lev99] Alexander Levin. Computation of Hilbert polynomials in two variables. Journal of Symbolic Computation, 28(4–5): 681–710, October/November 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0320/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0320/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0320/production/ref.

Levin:2000:RGB

[Lev00]  Alexander Levin.  Re-

duced gröbner bases, free difference-differential modules and difference-differential dimension polynomials. Journal of Symbolic Computation, 30(4):357–382, October 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0412;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0412/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0412/ref.

Levandovskyy:2007:FE

[Lev07a] Viktor Levandovskyy. Foreword from the Editor. Journal of Symbolic Computation, 42(11–12):1001–1002, November/December 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Levin:2007:GBR

[Lev07b] Alexander B. Levin. Gröbner bases with respect to several orderings and multivariable dimension polynomials. Journal of Symbolic Computation, 42(5):561–578, May 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Levin:2021:BKT

[Lev21]  Alexander Levin.  Bivari-

ate Kolchin-type dimension polynomials of non-reflexive prime difference-differential ideals. The case of one translation. Journal of Symbolic Computation, 102(??): 173–188, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301245.

Leykin:2001:CSP

[Ley01] Anton Leykin. Constructibility of the set of polynomials with a fixed Bernstein– Sato polynomial: an algorithmic approach. Journal of Symbolic Computation, 32(6):663–675, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0488;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0488/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0488/ref.

Leykin:2004:APT

[Ley04] Anton Leykin. Algorithmic proofs of two theorems of Stafford. Journal of Symbolic Computation, 38(6): 1535–1550, December 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Li:2019:SCP

[LFD19] Yongbin Li, Qiuju Fu, and Renbin Deng. Symbolic computation on the partition lattice of an n-set. Journal of Symbolic Computation, 95(??):185–201, November/December 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300185.

Li:2021:CFD

[LG21] Yunnan Li and Li Guo. Construction of free differential algebras by extending Gröbner–Shirshov bases. Journal of Symbolic Computation, 107(??):167–189, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000213.

Lyakhov:2021:ALN

[LGM21] Dmitry A. Lyakhov, Vladimir P.

Gerdt, and Dominik L. Michels. On the algorithmic linearizability of nonlinear ordinary differential equations. Journal of Symbolic Computation, 98(??):3–22, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300689.

Leedham-Green:1991:CGH

[LGPS91] Charles R. Leedham-Green, Cheryl E. Praeger, and Leonard H. Soicher. Computing with group homomorphisms. Journal of Symbolic Computation, 12(4–5): 527–532, October/November 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Leedham-Green:1990:CLO

[LGS90] C. R. Leedham-Green and L. H. Soicher. Collection from the left and other strategies. Journal of Symbolic Computation, 9(5–6):665–676 (or 665– 675??), May/June 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Liu:1998:PGB

[LH98] Mulan Liu and Lei Hu. Properties of Gröbner bases and applications to doubly periodic arrays. Journal of Symbolic Computation, 26(3):301– 314, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lisle:2017:ACL

[LH17] Ian G. Lisle and S.-L. Tracy Huang. Algorithmic calculus for Lie determining systems. Journal of Symbolic Computation, 79 (part 2)(??): 482–498, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300025.

Levandovskyy:2018:FAA

[LH18] Viktor Levandovskyy and Albert Heinle. A factorization algorithm for G-algebras and its applications. Journal of Symbolic Computation, 85(??):188–205, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300731.

Lopez:1996:MGB

[LHD96] P. Lopez, M. Hermenegildo, and S. Debray. A methodology for granularity-based control of parallelism in logic programs. Journal of Symbolic Computation, 21(4/5/6):715– 734, April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Linton:2013:ECS

[LHK^{+}13] S. Linton, K. Hammond, A. Konovalov, C. Brown, P. W. Trinder, H.-W. Loidl, P. Horn, and D. Roozemond. Easy composition of symbolic computation software using SCSCP: a new Lingua Franca for symbolic computation. Journal of Symbolic Computation, 49(??):95–119, February 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002124.

Li:2004:ATW

[Li04] Yong-Bin Li. Applications of the theory of weakly nondegenerate conditions to zero decomposition for polynomial systems. Journal of Symbolic Computation, 38(1):815–832, July 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Li:2010:LGB

[Li10]  Huishi Li.  Looking for

Gröbner basis theory for (almost) skew 2-nomial algebras. Journal of Symbolic Computation, 45(9):918–942, September 2010. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-

855X (electronic).

Liang:2013:SLF

[Lia13a] Ye Liang. Selecting lengths of floats for the computation of approximate Gröbner bases. Journal of Symbolic Computation, 53(??):40–52, June 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001885.

Liang:2013:SPL

[Lia13b] Ye Liang. Structures of precision losses in computing approximate Gröbner bases. Journal of Symbolic Computation, 53(??):81–95, June 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001915.

Liang:2022:DBG

[Lia22] Yihui Liang. Degree bounds for Gröbner bases of modules. Journal of Symbolic Computation, 111(??): 27–43, July/August 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000808.

Lichtblau:2021:SAM

[Lic21] Daniel Lichtblau. Symbolic analysis of multiple steady states in a MAPK chemical reaction network. Journal of Symbolic Computation, 105(??):118–144, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300468.

Limongelli:1993:EAB

[Lim93] Carla Limongelli. On an efficient algorithm for big rational number computations by parallel p-adics. Journal of Symbolic Computation, 15(2):181–198 (or 181– 197??), February 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Linton:1991:CMR

[Lin91a] S. A. Linton. Constructing matrix representations of finitely presented groups. Journal of Symbolic Computation, 12(4–5):427–438, October/November 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Linton:1991:DCE

[Lin91b] Stephen A. Linton. Double coset enumeration. Journal of Symbolic Computation, 12(4–5):415–426, October/November 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Lin:2018:CLS

[Lin18] Bo Lin. Computing linear systems on metric graphs. Journal of Symbolic Computation, 87(??):54–67, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300585.

Lippok:1993:RBC

[Lip93] Frank Lippok. On the representation of 1 by binary cubic forms of positive discriminant. Journal of Symbolic Computation, 15(3):297–314 (or 297– 313??), March 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lisonek:1995:CFN

[Lis95] Petr Lisoněk. Closed forms for the number of polygon dissections. Journal of Symbolic Computation, 20(5–6):595– 602 (or 595–601??), November/December 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Liu:2019:SAP

[Liu19] Ji-Cai Liu. Semi-automated proof of supercongruences on partial sums of hypergeometric series. Journal of Symbolic Computation, 93(??): 221–229, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300683.

Liang:2009:ACC

[LJ09] Songxin Liang and David J.

Jeffrey. Automatic computation of the complete root classification for a parametric polynomial. Journal of Symbolic Computation, 44(10): 1487–1501, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

LaScala:2009:LIN

[LL09] Roberto La Scala and Viktor Levandovskyy. Letterplace ideals and non-commutative Gröbner bases. Journal of Symbolic Computation, 44(10):1374–1393, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

LaScala:2013:SPR

[LL13] Roberto La Scala and Viktor Levandovskyy. Skew polynomial rings, Gröbner bases and the letterplace embedding of the free associative algebra. Journal of Symbolic Computation, 48(??):110–131, January 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112000934.

Lee:2016:SCS

[LL16] C. T. Lee and C. C. Lee. Symbolic computation on a second-order KdV equation. Journal of Symbolic Computation, 74(??):70–95, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711500053X.

Leykin:2023:F

[LL23] Anton Leykin and Pierre Lairez. Foreword. Journal of Symbolic Computation, 115 (??):407–408, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000621.

Lee:2008:EPC

[LLL08] Eunjeong Lee, Hyang-Sook Lee, and Yoonjin Lee. Eta pairing computation on general divisors over hyperelliptic curves y^{2} = x^{p} − x + d. Journal of Symbolic Computation, 43(6–7):452–474, June/ July 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Liu:2013:SCS

[LLL13] Y. P. Liu, S. J. Liao, and Z. B. Li. Symbolic computation of strongly nonlinear periodic oscillations. Journal of Symbolic Computation, 55(??):72–95, August 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000503.

Liu:2019:SCG

[LLL19] Jinwang Liu, Dongmei Li, and Weijun Liu. Some criteria for Gröbner bases and their applications. Journal of Symbolic Computation, 92 (??):15–21, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117301232.

Lasserre:2013:MMB

[LLM^{+}13] Jean-Bernard Lasserre, Monique Laurent, Bernard Mourrain, Philipp Rostalski, and Philippe Trébuchet. Moment matrices, border bases and real radical computation. Journal of Symbolic Computation, 51(??):63–85, April 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001162.

Linaje:2011:PRU

[LLTPT^{+}11] Marino Linaje, Adolfo Lozano-Tello, Miguel A. Perez-Toledano, Juan Carlos Preciado, Roberto Rodriguez-Echeverria, and Fernando Sanchez-Figueroa. Providing RIA user interfaces with accessibility properties. Journal of Symbolic Computation, 46(2):207–217, February 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001380.

Liu:2003:TOW

[LLW03] Jinwang Liu, Zhuojun Liu, and Mingsheng Wang. The term orderings which are compatible with composition II. Journal of Symbolic Computation, 35(2):153–168, February 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Liu:2024:SNF

[LLW24] Jinwang Liu, Dongmei Li, and Tao Wu. The Smith normal form and reduction of weakly linear matrices. Journal of Symbolic Computation, 120(??):Article 102232, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000469.

Lee:2024:TSN

[LLZ24] Kisun Lee, Nan Li, and Lihong Zhi. Two-step Newton’s method for deflationone singular zeros of analytic systems. Journal of Symbolic Computation, 123 (??):??, July/August 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000925.

Langemyr:1989:CPG

[LM89] Lars Langemyr and Scott Mc-Callum. The computation of polynomial greatest common divisors over an algebraic number field. Journal of Symbolic Computation, 8(5):429– 448, November 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lallement:1990:DGR

[LM90] Gerard Lallement and Robert McFadden. On the determination of Green’s relations in finite transformation semigroups. Journal of Symbolic Computation, 10(5):481– 498, November 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lassez:1992:CFG

[LM92] Jean-Louis L. Lassez and Ken McAloon. A canonical form for generalized linear constraints. Journal of Symbolic Computation, 13(1):1– 24, January 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Also appears as IBM Technical Report RC 15004.

Lucchini:1994:CSG

[LM94a] Andrea Lucchini and Federico Menegazzo. Computing a set of generators of minimal cardinality in a solvable group. Journal of Symbolic Computation, 17(5):409–420, May 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lugiez:1994:TAH

[LM94b] D. Lugiez and J. L. Moysset. Tree automata help one to solve equational formulae in AC-theories. Journal of Symbolic Computation, 18(4): 297–318, October 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lazard:2009:ID

[LM09] Daniel Lazard and Scott Mc-Callum. Iterated discriminants. Journal of Symbolic Computation, 44(9): 1176–1193, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Len:2020:LTB

[LM20] Yoav Len and Hannah Markwig. Lifting tropical bitangents. Journal of Symbolic Computation, 96(??): 122–152, January/February 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300264.

Langar:2011:FES

[LMA11] Mahjoub Langar, Mohamed Mejri, and Kamel Adi. Formal enforcement of security policies on concurrent systems. Journal of Symbolic Computation, 46(9):997–1016, September 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000630.

Lygeros:2005:ERS

[LMM05] Nik Lygeros, Paul-Valère Marchand, and Marc Massot. Enumeration and 3D representation of the stereoisomers of alkane molecules. Journal of Symbolic Computation, 40(4–5):1225–1241, October/November 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lassez:1989:SKB

[LMP89] Catherine Lassez, Ken McAloon, and Graeme Port. Stratification and knowledge base management. Journal of Symbolic Computation, 7(5):509–522, May 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Le:2019:ICS

[LMP19] Tung Le, Kay Magaard, and Alessandro Paolini. The irreducible characters of the Sylow p-subgroups of the Chevalley groups D_{6}(p^{f}) and E_{6}(p^{f}). Journal of Symbolic Computation, 95(??): 68–99, November/December 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300112.

Lemaire:2011:WDE

[LMPX11] François Lemaire, Marc Moreno Maza, Wei Pan, and Yuzhen Xie. When does ⟨T ⟩ equal sat(T)? Journal of Symbolic Computation, 46(12): 1291–1305, December 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001222.

Lux:1994:PCS

[LMR94] Klaus Lux, Jürgen Müller, and Michael Ringe. Peakword condensation and submodule lattices: an application of the MEAT-AXE. Journal of Symbolic Computation, 17(6): 529–544, June 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Li:2011:MLB

[LMRS11] Xin Li, Marc Moreno Maza, Raqeeb Rasheed, and Éric Schost. The modpn library: Bringing fast polynomial arithmetic into Maple. Journal of Symbolic Computation, 46(7):841–858, July 2011. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711000146X.

Li:2009:FAT

[LMS09] Xin Li, Marc Moreno Maza, and Éric Schost. Fast arithmetic for triangular sets: From theory to practice. Journal of Symbolic Computation, 44(7):891–907, July 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Levandovskyy:2023:CFN

[LMZ23] Viktor Levandovskyy, Tobias Metzlaff, and Karim Abou Zeid. Computing free non-commutative Gröbner bases over Z with Singular:Letterplace. Journal of Symbolic Computation, 115 (??):201–222, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000736.

Lange:2013:BRC

[LN13] H. Lange and P. E. Newstead. Bundles of rank 3 on curves of Clifford index

3. Journal of Symbolic Computation, 57(??):3–18, October 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000655.

Liu:2021:DIM

[LN21] Jiang Liu and Feng Ni. Distance invariant method for normalization of indexed differentials. Journal of Symbolic Computation, 104(??): 256–275, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300304.

Lucas:2021:VPS

[LNP^{+}21] David Lucas, Vincent Neiger, Clément Pernet, Daniel S. Roche, and Johan Rosenkilde. Verification protocols with sub-linear communication for polynomial matrix operations. Journal of Symbolic Computation, 105(??):165–198, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300481.

Liu:1996:DMG

[LO96] Chuchang Liu and M. A. Orgun. Dealing with multiple granularity of time in temporal logic programming. Journal of Symbolic Computation, 22(5–6):699–720, November/December 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lo:1998:FIN

[Lo98a] Eddie H. Lo. Finding intersections and normalizers in finitely generated nilpotent groups. Journal of Symbolic Computation, 25(1):45–60 (or 45–59??), January 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lo:1998:PQA

[Lo98b] Eddie H. Lo. A polycyclic quotient algorithm. Journal of Symbolic Computation, 25 (1):61–98 (or 61–97??), January 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lo:1999:PAF

[LO99] Eddie H. Lo and Gretchen Ostheimer. A practical algorithm for finding matrix representations for polycyclic groups. Journal of Symbolic Computation, 28(3): 339–360, September 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0286/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0286/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0286/production/ref.

Lee:2008:LDR

[LO08] Kwankyu Lee and Michael E.

O’Sullivan. List decoding of Reed–Solomon codes from a Gröbner basis perspective.

Journal of Symbolic Computation, 43(9):645–658, September 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lee:2009:LDH

[LO09] Kwankyu Lee and Michael E. O’Sullivan. List decoding of Hermitian codes using Gröbner bases. Journal of Symbolic Computation, 44(12):1662–1675, December 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Loja:2013:UPS

[Loj13] M. A. R. Loja. On the use of particle swarm optimization to maximize bending stiffness of functionally graded structures. Journal of Symbolic Computation, 61– 62(??):12–30, ???? 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001259.

Little:2003:SRF

[LOOR^{+}03] John B. Little, David Ortiz, Ricardo Ortiz-Rosado, Rebecca Pablo, and Karen Ríos-Soto. Some remarks on Fitzpatrick and Flynn’s Gröbner basis technique for Padé approximation. Journal of Symbolic Computation, 35 (4):451–461, April 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Louboutin:2008:LCZ

[Lou08] Stéphane R. Louboutin. Localization of the complex zeros of parametrized families of polynomials. Journal of Symbolic Computation, 43(4): 304–309, April 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Loubani:2021:DMS

[Lou21] Jinan Loubani. The dimension of the moduli spaces of curves defined by topologically non quasi-homogeneous functions. Journal of Symbolic Computation, 104(??): 207–235, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300298.

Lloyd:1990:RBL

[LP90] N. G. Lloyd and J. M. Pearson. REDUCE and the bifurcation of limit cycles. Journal of Symbolic Computation, 9(2):215–224, February 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lloyd:2002:SPD

[LP02] N. G. Lloyd and J. M. Pearson. Symmetry in planar dynamical systems. Journal of Symbolic Computation, 33(3):357–366, March 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0502;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0502/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0502/ref.

Lascoux:2003:DSS

[LP03] Alain Lascoux and Piotr Pragacz. Double Sylvester sums for subresultants and multi-Schur functions. Journal of Symbolic Computation, 35(6): 689–710, June 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Levandovskyy:2012:FE

[LPPR12] Viktor Levandovskyy, Dusan Pagon, Marko Petkovsek, and Valery Romanovski. Foreword from the Editors. Journal of Symbolic Computation, 47(10):1137–1139, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711100229X.

Lazard:2017:BTD

[LPR17] Sylvain Lazard, Marc Pouget, and Fabrice Rouillier. Bivariate triangular decompositions in the presence of asymptotes. Journal of Symbolic Computation, 82(??): 123–133, September/October 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300111.

Linton:2002:CTS

[LPRR02] S. A. Linton, G. Pfeiffer, E. F. Robertson, and N. Ruškuc. Computing transformation semigroups. Journal of Symbolic Computation, 33(2):145–162, February 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0406;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0406/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0406/ref.

Lisonek:1993:IDS

[LPS93] Petr Lisoněk, Peter Paule, and Volker Strehl. Improvement of the degree setting in Gosper’s algorithm. Journal of Symbolic Computation, 16 (3):243–258, September 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

LaScala:2020:NAC

[LPT20] Roberto La Scala, Dmitri Piontkovski, and Sharwan K. Tiwari. Noncommutative algebras, context-free grammars and algebraic Hilbert series. Journal of Symbolic Computation, 101(??): 28–50, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300586.

Lafferriere:2001:SRC

[LPY01] Gerardo Lafferriere, George J. Pappas, and Sergio Yovine. Symbolic reachability computation for families of linear vector fields. Journal of Symbolic Computation, 32 (3):231–253, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0472;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0472/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0472/ref.

Lazard:1990:IRF

[LR90] D. Lazard and R. Rioboo. Integration of rational functions: rational computation of the logarithmic part. Journal of Symbolic Computation, 9(2):113–116 (or 113– 115??), February 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lisle:1998:GSL

[LR98] Ian G. Lisle and Gregory J.

Reid. Geometry and structure of Lie pseudogroups from infinitesimal defining systems. Journal of Symbolic Computation, 26(3):355–379, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lickteig:2001:SHS

[LR01] Thomas Lickteig and Marie-Françoise Roy. Sylvester-Habicht sequences and fast Cauchy index computation. Journal of Symbolic Computation, 31(3):315–341, March 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0427;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0427/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0427/ref.

Lazard:2007:SPP

[LR07] Daniel Lazard and Fabrice Rouillier. Solving parametric polynomial systems. Journal of Symbolic Computation, 42 (6):636–667, June 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lubicz:2015:GMA

[LR15] David Lubicz and Damien Robert. A generalisation of Miller’s algorithm and applications to pairing computations on abelian varieties. Journal of Symbolic Computation, 67(??): 68–92, March/April 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000510.

Lindberg:2023:ISE

[LR23] Julia Lindberg and Jose Israel Rodriguez. Invariants of SDP exactness in quadratic programming. Journal of Symbolic Computation, 122(??): Article 102258, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712300072X.

Lombardi:2000:NST

[LRD00] Henri Lombardi, Marie-Françoise Roy, and Mohab Safey El Din. New structure theorem for subresultants. Journal of Symbolic Computation, 29(4–5):663–689, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0322;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0322/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0322/ref.

Luks:1997:SAN

[LRW97] Eugene M. Luks, Ferenc Rákóczi, and Charles R. B. Wright. Some algorithms for nilpotent permutation groups. Journal of Symbolic Computation, 23(4):335–354, April 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Larsen:1994:ICF

[LS94] Kim S. Larsen and Michael I.

Schwartzbach. Injectivity of composite functions. Journal of Symbolic Computation, 17 (5):393–408, May 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lee:1995:FFG

[LS95] Hong R. Lee and B. David Saunders. Fraction free Gaussian elimination for sparse matrices. Journal of Symbolic Computation, 19(5): 393–402, April 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

LaScala:1998:SCM

[LS98] R. La Scala and M. Stillman. Strategies for computing minimal free resolutions. Journal of Symbolic Computation, 26(4):409–432, October 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Laubenbacher:2000:AQS

[LS00a] Reinhard Laubenbacher and Karen Schlauch. An algorithm for the Quillen-Suslin theorem for quotients of polynomial rings by monomial ideals. Journal of Symbolic Computation, 30(5):555–571, October 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0367;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0367/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0367/ref.

Laubenbacher:2000:PI

[LS00b] Reinhard C. Laubenbacher and Irena Swanson. Permanental ideals. Journal of Symbolic Computation, 30(2):195–205, August 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0363;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0363/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0363/ref.

Li:2001:RSR

[LS01] Ziming Li and Fritz Schwarz. Rational solutions of Riccatilike partial differential equations. Journal of Symbolic Computation, 31(6): 691–716, June 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0461;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0461/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0461/ref.

Linton:2002:EIA

[LS02] Steve Linton and Roberto Sebastiani. Editorial: The integration of automated reasoning and computer algebra systems. Journal of Symbolic Computation, 34(4): 239, October 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lebrun:2004:GBW

[LS04] Jérôme Lebrun and Ivan Selesnick. Gröbner bases and wavelet design. Journal of Symbolic Computation, 37(2): 227–259, February 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Levandovskyy:2011:CDF

[LS11] Viktor Levandovskyy and Kristina Schindelar. Computing diagonal form and Jacobson normal form of a matrix using Gröbner bases. Journal of Symbolic Computation, 46(5):595–608, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001781.

Levandovskyy:2012:FFA

[LS12] Viktor Levandovskyy and Kristina Schindelar. Fractionfree algorithm for the computation of diagonal forms matrices over Ore domains using Gröbner bases. Journal of Symbolic Computation, 47(10):1214–1232, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002458.

Lebreton:2016:SFO

[LS16a] Romain Lebreton and Éric Schost. A simple and fast online power series multiplication and its analysis. Journal of Symbolic Computation, 72(??):231–251, January/February 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000176.

Lenstraand:2016:DCF

[LS16b] H. W. Lenstra, Jr. and A. Silverberg. Determining cyclicity of finite modules. Journal of Symbolic Computation, 73(??):153–156, March/April 2016. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000437.

Li:2019:FMP

[LSS19] Zijia Li, Josef Schicho, and Hans-Peter Schröcker. Factorization of motion polynomials. Journal of Symbolic Computation, 92(??):190–202, May/ June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300166.

Ling:2012:MIH

[LSSW12] San Ling, Igor E. Shparlinski, Ron Steinfeld, and Huaxiong Wang. On the modular inversion hidden number problem. Journal of Symbolic Computation, 47(4):358–367, April 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001374.

Li:2003:FSL

[LST03] Ziming Li, Fritz Schwarz, and Serguei P. Tsarev. Factoring systems of linear PDEs with finite-dimensional solution spaces. Journal of Symbolic Computation, 36(3–4): 443–471, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Landsmann:2001:PCS

[LSW01] Günter Landsmann, Josef Schicho, and Franz Winkler. The parametrization of canal surfaces and the decomposition of polynomials into a sum of two squares. Journal of Symbolic Computation, 32 (1–2):119–132, July/August 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0453;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0453/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0453/ref.

Lam:2007:RPC

[LSY07] Chi-Ming Lam, Vladimir Shpilrain, and Jie-Tai Yu. Recognizing and parametrizing curves isomorphic to a line. Journal of Symbolic Computation, 42(7):751–756, July 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lam:1989:BSI

[LT89] C. W. H. Lam and L. Thiel. Backtrack search with isomorph rejection and consistency check. Journal of Symbolic Computation, 7(5):473– 486 (or 473–485??), May 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Scala:2022:SBC

[LT22] Roberto La Scala and Sharwan K. Tiwari. Stream/block ciphers, difference equations and algebraic attacks. Journal of Symbolic Computation, 109 (??):177–198, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000584.

Lubeck:2002:CED

[Lüb02] Frank Lübeck. On the computation of elementary divisors of integer matrices. Journal of Symbolic Computation, 33(1):57–65, January 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0430;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0430/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0430/ref.

Lubbes:2014:MFC

[Lub14] Niels Lubbes. Minimal families of curves on surfaces. Journal of Symbolic Computation, 65(??):29–48, November 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000133.

Lubeck:2023:SGF

[Lüb23] Frank Lübeck. Standard generators of finite fields and their cyclic subgroups. Journal of Symbolic Computation, 117(??):51–67, July/August 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122001146.

Lugiez:1995:PNR

[Lug95] D. Lugiez. Positive and negative results for higher-order disunification. Journal of Symbolic Computation, 20(4): 431–470, October 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lundman:2016:LPL

[Lun16] Anders Lundman. Local positivity of line bundles on smooth toric varieties and Cayley polytopes. Journal of Symbolic Computation, 74 (??):109–124, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115000553.

Lopez:2014:CDL

[LV14] Hiram H. López and Rafael H. Villarreal. Computing the degree of a lattice ideal of dimension one. Journal of Symbolic Computation, 65 (??):15–28, November 2014.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000121.

Lihong:1998:NSP

[LW98] Z. Lihong and W. Wenda.

Nearest singular polynomials. Journal of Symbolic Computation, 26(6):667–676, December 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lux:2001:DSS

[LW01] Klaus Lux and Markus Wiegelmann. Determination of Socle series using the condensation method. Journal of Symbolic Computation, 31(1–2):163–178, January/February 2001. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1009;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.1009/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1009/ref.

Li:2003:ASPa

[LW03a] Hongbo Li and Yihong Wu.

Automated short proof generation for projective geometric theorems with Cayley and bracket algebras: I. Incidence geometry. Journal of Symbolic Computation, 36(5):717– 762, November 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Li:2003:ASPb

[LW03b] Hongbo Li and Yihong Wu. Automated short proof generation for projective geometric theorems with Cayley and bracket algebras: II. Conic geometry. Journal of Symbolic Computation, 36(5):763– 809, November 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lutz:2010:DIC

[LW10] Carsten Lutz and Frank Wolter. Deciding inseparability and conservative extensions in the description logic EL. Journal of Symbolic Computation, 45(2):194– 228, February 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Li:2012:TLF

[LW12] Ziming Li and Min Wu. Transforming linear functional systems into fully integrable systems. Journal of Symbolic Computation, 47(6):711–732, June 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002215.

Lu:2023:NRF

[LWX23] Dong Lu, Dingkang Wang, and Fanghui Xiao. New remarks on the factorization and equivalence problems for a class of multivariate polynomial matrices. Journal of Symbolic Computation, 115 (??):266–284, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000633.

Lu:2023:ERB

[LWXZ23] Dong Lu, Dingkang Wang, Fanghui Xiao, and Xiaopeng Zheng. Equivalence and reduction of bivariate polynomial matrices to their Smith forms. Journal of Symbolic Computation, 118(??):1– 16, September/October 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000019.

Li:2023:CBV

[LXZZ23] Haokun Li, Bican Xia, Huiying Zhang, and Tao Zheng. Choosing better variable orderings for cylindrical algebraic decomposition via exploiting chordal structure. Journal of Symbolic Computation, 116(??):324–344, May/ June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122001067.

Lombardi:2005:SAR

[LY05] Henri Lombardi and Ihsen Yengui. Suslin’s algorithms for reduction of unimodular rows. Journal of Symbolic Computation, 39(6):707–717, June 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Leykin:2018:BPH

[LY18] Anton Leykin and Josephine Yu. Beyond polyhedral homotopies. Journal of Symbolic Computation, 91(??):173–180, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300841.

Li:2015:SDR

[LYG15] Wei Li, Chun-Ming Yuan, and Xiao-Shan Gao. Sparse difference resultant. Journal of Symbolic Computation, 68 (part 1)(??):169–203, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000844.

Lynch:1997:OEL

[Lyn97]  Christopher Lynch.  Ori-

ented equational logic programming is complete. Journal of Symbolic Computation, 23(1):23–46 (or 23–45??), January 1997. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Li:2012:CMS

[LZ12] Nan Li and Lihong Zhi. Computing the multiplicity structure of an isolated singular solution: Case of breadth one. Journal of Symbolic Computation, 47(6):700–710, June 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002203.

Levandovskyy:2011:ELM

[LZS11] Viktor Levandovskyy, Eva Zerz, and Kristina Schindelar. Exact linear modeling using Ore algebras. Journal of Symbolic Computation, 46(11):1189–1204, November 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001131.

Ma:1994:MDG

[Ma94] Yonghao Ma. On the minors defined by a generic matrix. Journal of Symbolic Computation, 18(6):503–518, December 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Madariaga:2014:GSB

[Mad14] Sara Madariaga. Gröbner– Shirshov bases for the nonsymmetric operands of dendriform algebras and quadrialgebras. Journal of Symbolic Computation, 60(??): 1–14, January 2014. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001351.

Magnan:1989:MPM

[Mag89] J. F. Magnan. A MACSYMA program for the multiple bifurcation analysis of doublediffusive convection. Journal of Symbolic Computation, 7(2):189–198 (or 189– 197??), February 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Maglione:2017:ECR

[Mag17] Joshua Maglione. Efficient characteristic refinements for finite groups. Journal of Symbolic Computation, 80 (part 2) (??):511–520, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300426.

Malle:1987:PPN

[Mal87] Gunter Malle. Polynomials for primitive nonsolvable permutation groups of degree d ≤ 15. Journal of Symbolic Computation, 4(1):83–92, August 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Malle:2000:MPP

[Mal00]  Gunter Malle.  Multi-

parameter polynomials with given Galois group. Journal of Symbolic Computation, 30(6):717–731, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0379;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0379/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0379/ref.

Maletzky:2021:GEF

[Mal21]  Alexander Maletzky.  A

generic and executable formalization of signature-based Gröbner basis algorithms. Journal of Symbolic Computation, 106(??):23–47, September/October 2021. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120301280.

Man:1993:CCFb

[Man93a] Yiu-Kwong Man. Computing closed form solutions of first order ODEs using the Prelle– Singer procedure. Journal of Symbolic Computation, 16(5):423–444 (or 423– 443??), November 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Man:1993:CCFa

[Man93b] Yiu-Kwong Man. On computing closed forms for indefinite summations. Journal of Symbolic Computation, 16(4): 355–376, October 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mili:2010:RTI

[MAN^{+}10] Ali Mili, Shir Aharon, Chaitanya Nadkarni, Lamia Labed Jilani, Asma Louhichi, and Olfa Mraihi. Reflexive transitive invariant relations: a basis for computing loop functions. Journal of Symbolic Computation, 45(11):1114– 1143, November 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mao:2021:PSW

[Mao21]  Guo-Shuai Mao.  Proof

of a supercongruence via the Wilf–Zeilberger method. Journal of Symbolic Computation, 107(??):269–278, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717121000250.

Marche:1996:NRA

[Mar96a]  Claude Marché.  Normal-

ized rewriting: An alternative to rewriting Modulo a set of equations. Journal of Symbolic Computation, 21(3): 253–288, March 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Marchiori:1996:MNF

[Mar96b] Massimo Marchiori. On the modularity of normal forms in rewriting. Journal of Symbolic Computation, 22(2): 143–154, August 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Martin:2002:ACF

[Mar02]  Bernd Martin.  Algorith-

mic computation of flattenings and of modular deformations. Journal of Symbolic Computation, 34(3):199– 212, September 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Markwig:2008:SB

[Mar08] Thomas Markwig. Standard bases in K[[t_{1}, …, t_{m}]][x_{1}, …, x_{n}]^{s}. Journal of Symbolic Computation, 43(11):765–786, November 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Margolis:2019:PGQ

[Mar19a] Leo Margolis. On the prime graph question for integral group rings of Conway simple groups. Journal of Symbolic Computation, 95(??):162–176, November/December 2019.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300161.

Marin:2019:PBC

[Mar19b] Ivan Marin. Proof of the BMR conjecture for G_{20} and G_{21}. Journal of Symbolic Computation, 92(??):1–14, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117301220.

Massri:2016:SSS

[Mas16]  César Massri.  Solving a

sparse system using linear algebra. Journal of Symbolic Computation, 73(??): 157–174, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000449.

Matsumoto:2001:CRI

[Mat01a] Ryutaroh Matsumoto. Computing the radical of an ideal in positive characteristic. Journal of Symbolic Computation, 32(3): 263–271, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0446;   http://www.
idealibrary.com/links/doi/

10.1006/jsco.2001.0446/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0446/ref.

Matusevich:2001:RJC

[Mat01b] Laura Felicia Matusevich.

Rank jumps in codimension 2A-hypergeometric systems. Journal of Symbolic Computation, 32(6): 619–641, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0486;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0486/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0486/ref.

Maus:1987:CIB

[Mau87] Eckart Maus. Computation of integral bases in certain S_{n} extensions of Q. Journal of Symbolic Computation, 4(1): 99–102, August 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Maubach:2000:ACK

[Mau00] Stefan Maubach. An algorithm to compute the kernel of a derivation up to a certain degree. Journal of Symbolic Computation, 29(6):959–970, June 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0334;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0334/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0334/ref.

Mawata:1988:SHE

[Maw88]  C. P. Mawata.  Solving

the heat equation symbolically with MuSimp. Journal of Symbolic Computation, 5 (3):385–389, June 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Martinez:2010:QGG

[MBC^{+}10] C. Martínez, R. Beivide, C. Camarero, E. Stafford, and E. M. Gabidulin. Quotients of Gaussian graphs and their application to perfect codes. Journal of Symbolic Computation, 45(7):813–824, July 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mou:2021:CGT

[MBL21] Chenqi Mou, Yang Bai, and Jiahua Lai. Chordal graphs in triangular decomposition in top-down style. Journal of Symbolic Computation, 102(??):108–131, January/February 2021. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301191.

Medina-Bulo:2010:VCL

[MBPLRR10] Inmaculada Medina-Bulo, Francisco Palomo-Lozano, and José-Luis Ruiz-Reina. A verified Common Lisp. Journal of Symbolic Computation, 45(1):96–123, January 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Manocha:1992:IRR

[MC92] Dinesh Manocha and John F.

Canny. Implicit representation of rational parametric surfaces. Journal of Symbolic Computation, 13(5):485–510, May 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Manocha:1993:MRA

[MC93] Dinesh Manocha and John F.

Canny. MultiPolynomial resultant algorithms. Journal of Symbolic Computation, 15(2): 99–122, February 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mansfield:1997:ADA

[MC97] Elizabeth L. Mansfield and Peter A. Clarkson. Applications of the differential algebra package diffgrob2 to classical symmetries of differential equations. Journal of Symbolic Computation, 23(5–6): 517–534 (or 517–533??), May/ June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

McCallum:2002:LBA

[MC02] Scott McCallum and George E. Collins. Local box adjacency algorithms for cylindrical algebraic decompositions. Journal of Symbolic Computation, 33(3):32–342, March 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0499;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0499/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0499/ref.

McCallum:1988:IPO

[McC88] Scott McCallum. An improved projection operation for cylindrical algebraic decomposition of threedimensional space. Journal of Symbolic Computation, 5 (1–2):141–162 (or 141–161??), February/April 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

McCallum:1997:TBP

[McC97] Scott McCallum. On testing a bivariate polynomial for analytic reducibility. Journal of Symbolic Computation, 24(5):509–536 (or 509– 535??), November 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

McCallum:1999:FIR

[McC99] S. McCallum. Factors of iterated resultants and discriminants. Journal of Symbolic Computation, 27(4): 367–386, April 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Moir:2021:UNP

[MCJ21] Robert H. C. Moir, Robert M.

Corless, and David J. Jeffrey. An unwinding number pair for continuous expressions of integrals. Journal of Symbolic Computation, 105(??):97–117, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300456.

Marquez-Corbella:2014:CAR

[MCMMPR14] Irene Márquez-Corbella, Edgar Martínez-Moro, Ruud Pellikaan, and Diego Ruano. Computational aspects of retrieving a representation of an algebraic geometry code. Journal of Symbolic Computation, 64(??):67–87, August 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001740.

McNulty:1992:FGE

[McN92]  George F. McNulty.  A

field guide to equational logic. Journal of Symbolic Computation, 14(4):371–398 (or 371– 397??), October 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Morales:2024:PPA

[MD24] Marcel Morales and Nguyen Thi Dung. A “pseudo-polynomial” algorithm for the Frobenius number and Gröbner basis. Journal of Symbolic Computation, 120(??):Article 102233, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000470.

Mohammadi:2017:TSA

[MdCW17] Fatemeh Mohammadi, Eduardo Sáenz de Cabezón, and Henry P. Wynn. Types of signature analysis in reliability based on Hilbert series. Journal of Symbolic Computation, 79 (part 1)(??): 140–155, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711630089X.

Magron:2021:ERH

[ME21] Victor Magron and Mohab Safey El Din. On exact Reznick, Hilbert–Artin and Putinar’s representations. Journal of Symbolic Computation, 107(??):221–250, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000249.

Meer:1994:RNC

[Mee94] Klaus Meer. Real number computations: On the use of information. Journal of Symbolic Computation, 18(3):199– 206, September 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Megiddo:1990:CSG

[Meg90] Nimrod Megiddo. On the complexity of some geometric problems in unbounded dimension. Journal of Symbolic Computation, 10(3–4): 327–334, September/October 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Merlet:2001:PIE

[Mer01] J.-P. Merlet. A parser for the interval evaluation of analytical functions and its application to engineering problems. Journal of Symbolic Computation, 31(4):475–486, April 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.

                                        2000.0429;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.0429/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0429/ref.

Merker:2010:ACI

[Mer10] Joël Merker. Application of computational invariant theory to Kobayashi hyperbolicity and to Green–Griffiths algebraic degeneracy. Journal of Symbolic Computation, 45 (10):986–1074, October 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Magron:2019:AWS

[MES19] Victor Magron, Mohab Safey El Din, and Markus Schweighofer. Algorithms for weighted sum of squares decomposition of non-negative univariate polynomials. Journal of Symbolic Computation, 93(??): 200–220, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300695.

Mathieu:1990:ACR

[MF90] Marie-Hélène H. Mathieu and David Ford. On p-adic computation of the rational form of a matrix. Journal of Symbolic Computation, 10(5):453– 464, November 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Murao:1996:MAS

[MF96] Hirokazu Murao and Tetsuro Fujise. Modular algorithm for sparse multivariate polynomial interpolation and its parallel implementation. Journal of Symbolic Computation, 21(4/5/6):377– 396, April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Myers:1988:PSR

[MG88] Thomas J. Myers and Maya B.

Gokhale. Parallel scheduling of recursively defined arrays. Journal of Symbolic Computation, 6(1):65–82 (or 65–81??), August 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mignotte:1994:SDP

[MG94a] Maurice Mignotte and Philippe Glesser. On the smallest divisor of a polynomial. Journal of Symbolic Computation, 17 (3):277–282, March 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mignotte:1994:LIH

[MG94b] Maurice Mignotte and Philippe H.

Glesser. Landau’s inequality via Hadamard’s. Journal of Symbolic Computation, 18(4):

379–384 (or 379–383??), October 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Martin:2000:TML

[MGL00] P. J. Martin, A. Gavilanes, and J. Leach. Tableau methods for a logic with term declarations. Journal of Symbolic Computation, 29(2):343–372, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0364;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0364/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0364/ref.

Miguel:2023:ESS

[MGRR23] Daniel Miguel, Andrea Guidolin, Ana Romero, and Julio Rubio. Effective spectral systems relating Serre and Eilenberg–Moore spectral sequences. Journal of Symbolic Computation, 114(??): 122–148, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000347.

Meseguer:1989:OSU

[MGS89] José Meseguer, Joseph A. Goguen, and Gert Smolka. Order sorted unification.

Journal of Symbolic Computation, 8(4):383–414 (or 383– 413??), October 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Preliminary version appeared as Report CSLI-87-86, Center for the Study of Language and Information, Stanford University, March 1987.

Myasnikov:2006:HSA

[MH06] A. D. Myasnikov and R. M.

Haralick. A hybrid search algorithm for the Whitehead minimization problem. Journal of Symbolic Computation, 41(7):818–834, July 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

McCallum:2016:ULP

[MH16] Scott McCallum and Hoon Hong. On using Lazard’s projection in CAD construction. Journal of Symbolic Computation, 72(??):65– 81, January/February 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000097.

Miret:2009:CCC

[MHXD09] Josep M. Miret, Xavier Hernández, and Sebastià Xambó-Descamps. Corrigendum to: “Computing the characteristic numbers of the variety of nodal plane cubics in P^{3}” [J. Symbolic Comput. 42 (1–2) (2007) 192–202]. Journal of Symbolic Computation, 44(4): 417–418, April 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [HMXD07].

Michler:1988:ADS

[Mic88] Gerhard O. Michler. An algorithm for determining the simplicity of a modular group representation. Journal of Symbolic Computation, 6(1): 105–112 (or 105–111??), August 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Michler:1990:SPC

[Mic90] Gerhard O. Michler. Some problems in computational representation theory. Journal of Symbolic Computation, 9(5–6):571–582, May/ June 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Michalska:2013:CTB

[Mic13]  Maria Michalska.  Curves

testing boundedness of polynomials on subsets of the real plane. Journal of Symbolic Computation, 56(??): 107–124, September 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000606.

Middeldorp:1994:CCC

[Mid94] Aart Middeldorp. Completeness of combinations of conditional constructor systems. Journal of Symbolic Computation, 17(1):3–21, January 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Conditional term rewriting systems (Pont-à-Mousson, 1992).

Mignotte:1992:PLR

[Mig92] Maurice Mignotte. On the product of the largest roots of a polynomial. Journal of Symbolic Computation, 13 (6):605–612 (or 605–611??), June 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mignotte:2000:BRL

[Mig00] Maurice Mignotte. Bounds for the roots of lacunary polynomials. Journal of Symbolic Computation, 30 (3):325–327, September 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0373;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0373/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0373/ref.

Mills:1987:USA

[Mil87]  R. D. Mills.  Using a

small algebraic manipulation system to solve differential and integral equations by variational and approximation techniques. Journal of Symbolic Computation, 3(3): 291–301, June 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Miller:1992:UUM

[Mil92a] Dale Miller. Unification under a mixed prefix. Journal of Symbolic Computation, 14(4): 321–358, October 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mills:1992:SRT

[Mil92b] R. D. Mills. Slope retention techniques for solving boundary-value problems in differential equations. Journal of Symbolic Computation, 13 (1):59–80, January 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mills:1993:SPS

[Mil93] Harry S. D. Mills. Symbolically precise solutions to a homogeneous second order matrix ordinary differential equation with Macsyma. Journal of Symbolic Computation, 15 (1):91–98, January 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Miller:1996:AGB

[Mil96] J. Lyn Miller. Analogs of Gröbner bases in polynomial rings over a ring. Journal of Symbolic Computation, 21(2):139–154 (or 139– 153??), February 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Minkwitz:1998:ASF

[Min98] Torsten Minkwitz. An algorithm for solving the factorization problem in permutation groups. Journal of Symbolic Computation, 26 (1):89–95, July 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Minimair:2002:SRC

[Min02] Manfred Minimair. Sparse resultant of composed polynomials II unmixed–mixed case. Journal of Symbolic Computation, 33(4):467–478, April 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Minimair:2003:DRC

[Min03] Manfred Minimair. Dense resultant of composed polynomials: Mixed–mixed case. Journal of Symbolic Computation, 36(6):825–834, December 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Minimair:2006:RPC

[Min06] Manfred Minimair. Resultants of partially composed polynomials. Journal of Symbolic Computation, 41(5): 591–602, May 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Miyamoto:2001:CIA

[Miy01]  Izumi Miyamoto.  Com-

puting isomorphisms of association schemes and its application. Journal of Symbolic Computation, 32 (1–2):133–141, July/August 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0450;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0450/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0450/ref.

Masjed-Jamei:2017:SCS

[MJK17] Mohammad Masjed-Jamei and Wolfram Koepf. Symbolic computation of some powertrigonometric series. Journal of Symbolic Computation, 80 (part 2)(??):273–284, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300049.

Malekbala:2022:DSN

[MJST22] Ghazaleh Malekbala, Leila Musavizadeh Jazaeri, Leila Sharifan, and Maryam Taha. On the dynamics of semilattice networks. Journal of Symbolic Computation, 113(??): 53–73, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000104.

Murao:1993:FSR

[MKF93] Hirokazu Murao, Hidetsune Kobayashi, and Tetsuro Fujise. On factorizing the symbolic U-resultant— application of the ddet operator. Journal of Symbolic Computation, 15(2):123– 142, February 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Martin:1992:SEC

[ML92] Ursula Martin and Michael Lai. Some experiments with a completion theorem prover. Journal of Symbolic Computation, 13(1):81–100, January 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Miola:1988:CLG

[MM88] A. Miola and T. Mora. Constructive lifting in graded structures: a unified view of Buchberger and Hensel methods. Journal of Symbolic Computation, 6(2–3):305–322, October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Man:1997:RAP

[MM97] Yiu-Kwong Man and Malcolm A. H. MacCallum. A rational approach to the Prelle– Singer algorithm. Journal of Symbolic Computation, 24(1): 31–44 (or 31–43??), July 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Matsumoto:2000:FBL

[MM00] Ryutaroh Matsumoto and Shinji Miura. Finding a basis of a linear system with pairwise distinct discrete valuations on an algebraic curve. Journal of Symbolic Computation, 30 (3):309–323, September 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0372;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0372/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0372/ref.

Marco:2004:PCD

[MM04a] Ana Marco and José-Javier Martínez. Parallel computation of determinants of matrices with polynomial entries. Journal of Symbolic Computation, 37(6):749–760, June 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Martinez-Moro:2004:RRF

[MM04b] Edgar Martínez-Moro. Regular representations of finitedimensional separable semisimple algebras and Gröbner bases. Journal of Symbolic Computation, 37(5):575–587, May 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Manubens:2006:IDA

[MM06] Montserrat Manubens and Antonio Montes. Improving the DISPGB algorithm using the discriminant ideal. Journal of Symbolic Computation, 41(11):1245–1263, November 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Manubens:2009:MCC

[MM09] Montserrat Manubens and Antonio Montes. Minimal canonical comprehensive Gröbner systems. Journal of Symbolic Computation, 44 (5):463–478, May 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Moy:2010:MIS

[MM10] Yannick Moy and Claude Marché. Modular inference of subprogram contracts for safety checking. Journal of Symbolic Computation, 45(11):1184–1211, November 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Margulies:2016:PTS

[MM16] S. Margulies and J. Morton. Polynomial-time solvable #CSP problems via algebraic models and Pfaffian circuits. Journal of Symbolic Computation, 74(??):152–180, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000589.

Miglioli:1994:APC

[MMO94] Pierangelo Miglioli, Ugo Moscato, and Mario Ornaghi. Abstract parametric classes and abstract data types defined by classical and constructive logical methods. Journal of Symbolic Computation, 18(1):41–82 (or 41–81??), July 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Markwig:2018:FSI

[MMO18] Hannah Markwig, Bernard Mourrain, and Giorgio Ottaviani. Foreword on the special issue of JSC on the occasion of MEGA 2017. Journal of Symbolic Computation, 91(??):1–2, ???? 2018.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300749.

Massarenti:2018:EIC

[MMS18] Alex Massarenti, Massimiliano Mella, and Giovanni Staglianò. Effective identifiability criteria for tensors and polynomials. Journal of Symbolic Computation, 87(??):227–237, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711730113X.

Mantzaflaris:2023:CIM

[MMS23] Angelos Mantzaflaris, Bernard Mourrain, and Agnes Szanto. A certified iterative method for isolated singular roots. Journal of Symbolic Computation, 115(??):223–247, March/April 2023. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000724.

Malecha:2011:TBV

[MMW11] Gregory Malecha, Greg Morrisett, and Ryan Wisnesky. Trace-based verification of imperative programs with I/O. Journal of Symbolic Computation, 46(2):95–118, February 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001343.

Matzat:2000:SIA

[MMY00] B. H. Matzat, J. McKay, and K. Yokoyama. Special issue on algorithmic methods in Galois theory: Foreword of the Guest Editors. Journal of Symbolic Computation, 30(6):631–633, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0389;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.0389/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0389/ref.

Martin:1989:BUS

[MN89] Ursula Martin and Tobias Nipkow. Boolean unification — the story so far. Journal of Symbolic Computation, 7(3–4):275–294 (or 275– 293??), March/April 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Maansson:2002:RGB

[MN02] Jonas Månsson and Patrik Nordbeck. Regular Gröbner bases. Journal of Symbolic Computation, 33(2):163–181, February 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0500;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0500/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0500/ref.

Mehlhorn:1994:SIA

[MNJ94] K. Mehlhorn, S. Näher, and Nievergelt J. Special issue on algorithms: Implementation, libraries and use— introduction of the Guest Editors. Journal of Symbolic Computation, 17(4):295, April 1994.

Mnuk:1997:AAC

[Mňu97]  Michal Mňuk.  An alge-

braic approach to computing adjoint curves. Journal of Symbolic Computation, 23(2– 3):229–240, February/March 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Madlener:1985:PAW

[MO85] Klaus Madlener and Friedrich Otto. Pseudonatural algorithms for the word problem for finitely presented monoids and groups. Journal of Symbolic Computation, 1(4):383– 418, December 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Madlener:1988:PNA

[MO88] Klaus Madlener and Friedrich Otto. Pseudo-natural algorithms for finitely generated presentations of monoids and groups. Journal of Symbolic Computation, 5(3): 339–358, June 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Murray:1995:SBP

[MO95] Scott H. Murray and E. A. O’Brien. Selecting base points for the Schreier-Sims algorithm for matrix groups. Journal of Symbolic Computation, 19(6):577–584, June 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Middeldorp:1998:DLN

[MO98] Aart Middeldorp and Satoshi Okui. A deterministic lazy narrowing calculus. Journal of Symbolic Computation, 25 (6):733–757, June 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

McGuire:2021:TGX

[MO21] Gary McGuire and Daniela O’Hara. On the termination of the general XL algorithm and ordinary multinomials. Journal of Symbolic Computation, 104(??):90–104, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300225.

Moeckel:2005:SAI

[Moe05] Richard Moeckel. Sturm’s algorithm and isolating blocks. Journal of Symbolic Computation, 40(4–5):1242–1255, October/November 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Moller:1988:CGB

[Möl88]  H. Michael Möller.  On

the construction of Gröbner bases using syzygies. Journal of Symbolic Computation, 6 (2–3):345–360 (or 345–359??), October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Monagan:1992:HIT

[Mon92] Michael B. Monagan. A heuristic irreducibility test for univariate polynomials. Journal of Symbolic Computation, 13(1):47–58 (or 47–57??), January 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Monagan:1997:WNC

[Mon97]  M. B. Monagan.  Work-

sheets and notebooks: Can we teach mathematical algorithms with them? Journal of Symbolic Computation, 23(5– 6):535–550, May/June 1997.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Monico:2002:CPD

[Mon02a] Chris Monico. Computing the primary decomposition of zero-dimensional ideals. Journal of Symbolic Computation, 34(5):451–459, November 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Montes:2002:NAD

[Mon02b] Antonio Montes. A new algorithm for discussing Gröbner bases with parameters. Journal of Symbolic Computation, 33(2):183–208, February 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0504;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0504/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0504/ref. See erratum

[DHM11].

Montaner:2005:ORH

[Mon05] Josep Àlvarez Montaner. Operations with regular holonomic D-modules with support a normal crossing. Journal of Symbolic Computation, 40(2):999–1012, August 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Margulies:2015:CHR

[MOP15] S. Margulies, S. Onn, and D. V. Pasechnik. On the complexity of Hilbert refutations for partition. Journal of Symbolic Computation, 66 (??):70–83, January/February 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000194.

Morgenstern:1991:IGA

[Mor91] Jacques Morgenstern. Invariant and geometric aspects of algebraic complexity theory. I. Journal of Symbolic Computation, 11(5–6):455–470 (or 455– 469??), May/June 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Invariant-theoretic algorithms in geometry (Minneapolis, MN, 1987).

Morrison:1999:DI

[Mor99] Sally Morrison. The differential ideal [P]: M^{∞}. Journal of Symbolic Computation, 28(4–5):631–656, October/November 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0318/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0318/production/pdf;
 http://www.idealibrary.

                                        com/links/doi/10.1006/jsco.
                                        1999.0318/production/ref.

Moroz:2011:PDP

[Mor11] Guillaume Moroz. Properness defects of projection and minimal discriminant variety. Journal of Symbolic Computation, 46(10):1139–1157, October 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111000848.

Morton:2013:RAC

[Mor13]  Jason Morton.  Relations

among conditional probabilities. Journal of Symbolic Computation, 50(??): 478–492, March 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001575.

Mora:2020:ZRE

[Mor20] Teo Mora. Zacharias representation of effective associative rings. Journal of Symbolic Computation, 99(??): 147–188, July/August 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300392.

Mosteig:2008:VMZ

[Mos08]  Edward Mosteig.  Value

monoids of zero-dimensional valuations of rank 1. Journal of Symbolic Computation, 43 (10):688–725, October 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Moses:2012:MPH

[Mos12] Joel Moses. Macsyma: a personal history. Journal of Symbolic Computation, 47 (2):123–130, February 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001483.

Mourrain:1998:CIR

[Mou98] B. Mourrain. Computing the isolated roots by matrix methods. Journal of Symbolic Computation, 26(6):715– 738, December 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mourrain:2005:BQR

[Mou05]  B. Mourrain.  Bezoutian

and quotient ring structure. Journal of Symbolic Computation, 39(3–4):397–415, March/April 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mozes:1989:DDB

[Moz89]  Eyal Mozes.  A deduc-

tive database based on Aristotelian logic. Journal of Symbolic Computation, 7(5):487– 508 (or 487–507??), May 1989.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Martin:1989:KKS

[MP89] Bernd Martin and Gerhard Pfister. The kernel of the Kodaira–Spencer map of the versal μ-constant deformation of an irreducible plane curve singularity with C^{r}-action. Journal of Symbolic Computation, 7(6):527–532 (or 527– 531??), June 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Migliore:2004:STU

[MP04] Juan Migliore and Chris Peterson. A symbolic test for (i, j)-uniformity in reduced zero-dimensional schemes. Journal of Symbolic Computation, 37(3):403–413, March 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mourrain:2009:SMS

[MP09] B. Mourrain and J. P. Pavone. Subdivision methods for solving polynomial equations. Journal of Symbolic Computation, 44(3):292–306, March 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Maggiolo:2011:GSM

[MP11a] Stefano Maggiolo and Nicola Pagani. Generating stable modular graphs. Journal of Symbolic Computation, 46(10):1087–1097, October 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000794.

Monagan:2011:SPD

[MP11b] Michael Monagan and Roman Pearce. Sparse polynomial division using a heap. Journal of Symbolic Computation, 46(7):807–822, July 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001446.

McKay:2014:PGI

[MP14] Brendan D. McKay and Adolfo Piperno. Practical graph isomorphism, II. Journal of Symbolic Computation, 60(??):94–112, January 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001193.

Mahdi:2017:HSN

[MPH17] Adam Mahdi, Claudio Pessoa, and Jonathan D. Hauenstein. A hybrid symbolic-numerical approach to the center-focus problem. Journal of Symbolic Computation, 82(??):57– 73, September/October 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116301717.

McCallum:2019:VPL

[MPP19] Scott McCallum, Adam Parusiński, and Laurentiu Paunescu. Validity proof of Lazard’s method for CAD construction. Journal of Symbolic Computation, 92(??):52– 69, May/June 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117301268.

Meier:2002:CAE

[MPS02] Andreas Meier, Martin Pollet, and Volker Sorge. Comparing approaches to the exploration of the domain of residue classes. Journal of Symbolic Computation, 34(4): 287–306, October 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Marcolla:2016:SWC

[MPS16] Chiara Marcolla, Marco Pellegrini, and Massimiliano Sala. On the small-weight codewords of some Hermitian codes. Journal of Symbolic Computation, 73(??): 27–45, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711500019X.

Miret:2009:CSF

[MPSXD09] Josep M. Miret, Jordi Pujolàs, Kumar Saurav, and Sebastià Xambó-Descamps. Computing some fundamental numbers of the variety of nodal cubics in P^{3}. Journal of Symbolic Computation, 44(10):1425–1447, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Menini:2020:SAC

[MPT20] Laura Menini, Corrado Possieri, and Antonio Tornambè. A symbolic algorithm to compute immersions of polynomial systems into linear ones up to an output injection. Journal of Symbolic Computation, 99(??): 1–20, July/August 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300276.

Muller-Quade:1999:BAR

[MQS99] J. Müller-Quade and R. Steinwandt. Basic algorithms for rational function fields. Journal of Symbolic Computation, 27(2):143–170, February 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Muller-Quade:2000:GBA

[MQS00] Jörn Müller-Quade and Rainer Steinwandt. Gröbner bases applied to finitely generated field extensions. Journal of Symbolic Computation, 30(4):469–490, October 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0417;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0417/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0417/ref;   http://
www.idealibrary.com/links/
doi/10.1006/jsco.2000.0419;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0419/pdf.

Murray:1987:TLA

[MR87] Neil V. Murray and Erik Rosenthal. Theory links: Applications to automated theorem proving. Journal of Symbolic Computation, 4(2): 173–190, October 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mora:1988:GFI

[MR88] Teo Mora and Lorenzo Robbiano. The Gröbner Fan of an ideal. Journal of Symbolic Computation, 6(2–3):183–208, October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Madlener:1998:GGB

[MR98] Klaus Madlener and Birgit Reinert. A generalization of Gröbner basis algorithms to polycyclic group rings. Journal of Symbolic Computation, 25(1):23–44 (or 23–43??), January 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mourrain:2002:RBR

[MR02] Bernard Mourrain and Olivier Ruatta. Relations between roots and coefficients, interpolation and application to system solving. Journal of Symbolic Computation, 33(5): 679–699, May 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mehlhorn:2010:FAC

[MR10] Kurt Mehlhorn and Saurabh Ray. Faster algorithms for computing Hong’s bound on absolute positiveness. Journal of Symbolic Computation, 45 (6):677–683, June 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See corrigendum [KMR18].

Mayr:2013:DDB

[MR13] Ernst W. Mayr and Stephan Ritscher. Dimension-dependent bounds for Gröbner bases of polynomial ideals. Journal of Symbolic Computation, 49(??):78–94, February 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002112.

Morrison:2015:AMC

[MR15] Ralph Morrison and Qingchun Ren. Algorithms for Mumford curves. Journal of Symbolic Computation, 68 (part 2) (??):259–284, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000819.

Markwig:2017:GFH

[MR17] Thomas Markwig and Yue Ren. Gröbner fans of x-homogeneous ideals in R[[t]][x]. Journal of Symbolic Computation, 83(??):315–341, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301377.

Murray:2011:CHC

[MRD11] Scott H. Murray and Colva M. Roney-Dougal. Constructive homomorphisms for classical groups. Journal of Symbolic Computation, 46(4): 371–384, April 2011. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001525.

Matsumoto:2013:GLO

[MRG13] Ryutaroh Matsumoto, Diego Ruano, and Olav Geil. Generalization of the Lee– O’Sullivan list decoding for one-point AG codes. Journal of Symbolic Computation, 55(??):1–9, August 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711300045X.

Matsumoto:2017:LDA

[MRG17] Ryutaroh Matsumoto, Diego Ruano, and Olav Geil. List decoding algorithm based on voting in Gröbner bases for general one-point AG codes. Journal of Symbolic Computation, 79 (part 2)(??): 384–410, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000316.

Melquiond:2023:WFV

[MRH23] Guillaume Melquiond and Raphaël Rieu-Helft. WhyMP, a formally verified arbitraryprecision integer library. Journal of Symbolic Computation, 115(??):74–95, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000657.

Mrozek:1996:REA

[Mro96] Marian Mrozek. Rigorous error analysis of numerical algorithms via symbolic computations. Journal of Symbolic Computation, 22(4):435– 458, October 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mota:1996:NTG

[MRS96] Edjard Mota, David Robertson, and Alan Smaill. Nature-Time: Temporal granularity in simulation of ecosystems. Journal of Symbolic Computation, 22(5–6):665–698, November/December 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Maza:2007:ATD

[MRSW07] Marc Moreno Maza, Greg Reid, Robin Scott, and Wenyuan Wu. On approximate triangular decompositions in dimension zero. Journal of Symbolic Computation, 42(7):693–716, July 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Moustrou:2021:SIS

[MRV21] Philippe Moustrou, Cordian Riener, and Hugues Verdure. Symmetric ideals, Specht polynomials and solutions to symmetric systems of equations. Journal of Symbolic Computation, 107(??):106– 121, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000122.

Markwig:2017:SBM

[MRW17] Thomas Markwig, Yue Ren, and Oliver Wienand. Standard bases in mixed power series and polynomial rings over rings. Journal of Symbolic Computation, 79 (part 1)(??):119–139, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300888.

Malm:1995:SRF

[MS95] D. E. G. Malm and T. N. Subramaniam. The summation of rational functions by an extended Gosper algorithm. Journal of Symbolic Computation, 19(4):293–304, April 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Martin:2000:IPW

[MS00a] Ursula Martin and Duncan Shand. Invariants, patterns and weights for ordering terms. Journal of Symbolic Computation, 29(6):921–957, June 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.

                                        com/links/doi/10.1006/jsco.
                                        1999.0333;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0333/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0333/ref.

MartinezdeCastilla:2000:CBI

[MS00b] Ignacio Ojeda Martínez de Castilla and Ramón Peidra Sánchez. Cellular binomial ideals. primary decomposition of binomial ideals. Journal of Symbolic Computation, 30(4):383–400, October 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0413;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0413/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0413/ref.

Mosteig:2002:VF

[MS02] Edward Mosteig and Moss Sweedler. Valuations and filtrations. Journal of Symbolic Computation, 34(5):399– 435, November 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Matera:2003:FCD

[MS03a] G. Matera and A. Sedoglavic. Fast computation of discrete invariants associated to a differential rational mapping. Journal of Symbolic Computation, 36(3–4): 473–499, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mora:2003:GBS

[MS03b] Teo Mora and Massimiliano Sala. On the Gröbner bases of some symmetric systems and their application to coding theory. Journal of Symbolic Computation, 35(2):177– 194, February 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mulders:2003:LRP

[MS03c] T. Mulders and A. Storjohann. On lattice reduction for polynomial matrices. Journal of Symbolic Computation, 35 (4):377–401, April 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mulders:2004:CDL

[MS04] T. Mulders and A. Storjohann. Certified dense linear system solving. Journal of Symbolic Computation, 37(4): 485–510, April 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Martin:2009:MAC

[MS09] Bernd Martin and Hendrik Süß. Milnor algebras could be isomorphic to modular algebras. Journal of Symbolic Computation, 44(9): 1268–1279, September 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mezzarobba:2010:EBR

[MS10] Marc Mezzarobba and Bruno Salvy. Effective bounds for P -recursive sequences. Journal of Symbolic Computation, 45(10):1075–1096, October 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Marcelo:2011:NCM

[MS11a] Agustin Marcelo and Peter Schenzel. Non-Cohen– Macaulay unique factorization domains in small dimensions. Journal of Symbolic Computation, 46(5):609–621, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001793.

Mehlhorn:2011:DAI

[MS11b] Kurt Mehlhorn and Michael Sagraloff. A deterministic algorithm for isolating real roots of a real polynomial. Journal of Symbolic Computation, 46(1):70–90, January 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001537.

Meshkat:2014:IRL

[MS14] Nicolette Meshkat and Seth Sullivant. Identifiable reparametrizations of linear compartment models. Journal of Symbolic Computation, 63(??):46–67, May 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711300165X.

Marais:2015:CRS

[MS15] Magdaleen S. Marais and Andreas Steenpaß. The classification of real singularities using Singular. Part I: Splitting lemma and simple singularities. Journal of Symbolic Computation, 68 (part 1)(??):61–71, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000571.

Marais:2016:CRS

[MS16] Magdaleen S. Marais and Andreas Steenpaß. The classification of real singularities using Singular. Part II: the structure of the equivalence classes of the unimodal singularities. Journal of Symbolic Computation, 74(??):346–366, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000759.

Melczer:2021:ECA

[MS21] Stephen Melczer and Bruno Salvy. Effective coefficient asymptotics of multivariate rational functions via seminumerical algorithms for polynomial systems. Journal of Symbolic Computation, 103 (??):234–279, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300079.

Muratore:2022:ECA

[MS22] Giosuè Muratore and Csaba Schneider. Effective computations of the Atiyah–Bott formula. Journal of Symbolic Computation, 112(??): 164–181, September/October 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000050.

Madlener:1993:PGS

[MSKO93] Klaus Madlener, Andrea Sattler-Klein, and Friedrich Otto. On the problem of generating small convergent systems. Journal of Symbolic Computation, 16(2):167–188 (or 167–187??), August 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mehta:2021:MEA

[MSS21] Ranjana Mehta, Joydip Saha, and Indranath Sengupta.

Moh’s example of algebroid space curves. Journal of Symbolic Computation, 104 (??):168–182, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300250.

Mayr:2022:MVG

[MSV22] Ernst W. Mayr, Werner M.

Seiler, and Evgenii V. Vorozhtsov. Memories on Vladimir Gerdt. Journal of Symbolic Computation, 109(??):50–56, March/ April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000444.

Mehlhorn:2015:AFR

[MSW15] Kurt Mehlhorn, Michael Sagraloff, and Pengming Wang. From approximate factorization to root isolation with application to cylindrical algebraic decomposition. Journal of Symbolic Computation, 66(??):34– 69, January/February 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000200.

Miller:2000:GCM

[MSY00] Ezra Miller, Bernd Sturmfels, and Kohji Yanagawa. Generic and cogeneric monomial ideals. Journal of Symbolic Computation, 29(4–5):691–708, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0290;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0290/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0290/ref.

Malykh:2021:SIA

[MSY21] M. D. Malykh, L. A. Sevastianov, and Y. Yu. On symbolic integration of algebraic functions. Journal of Symbolic Computation, 104 (??):563–579, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300924.

McLaughlin:2009:RRC

[MSZ09] James McLaughlin, Andrew V. Sills, and Peter Zimmer. Rogers–Ramanujan computer searches. Journal of Symbolic Computation, 44(8): 1068–1078, August 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mints:1988:PSP

[MT88] G. Mints and E. Tyugu. The programming system PRIZ. Journal of Symbolic Computation, 5(3):359–376 (or 359– 375??), June 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Middeldorp:1993:CCC

[MT93] Aart Middeldorp and Yoshihito Toyama. Completeness of combinations of constructor systems. Journal of Symbolic Computation, 15(3): 331–348, March 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Moller:2001:MPS

[MT01] H. Michael Möller and Ralf Tenberg. Multivariate polynomial system solving using intersections of eigenspaces. Journal of Symbolic Computation, 32(5):513–531, November 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0476;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0476/pdf;   http://
www.idealibrary.com/links/
doi/10.1006/jsco.2001.0476/
ref.

Monagan:2020:CSH

[MT20] Michael Monagan and Baris Tuncer. The complexity of sparse Hensel lifting and sparse polynomial factorization. Journal of Symbolic Computation, 99(??): 189–230, July/August 2020.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300409.

Mignotte:2003:LRS

[Mtn03] Maurice Mignotte and Doru tef nescu. Linear recurrent sequences and polynomial roots. Journal of Symbolic Computation, 35(6):637–649, June 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mourrain:2021:TNF

[MTV21] Bernard Mourrain, Simon Telen, and Marc Van Barel. Truncated normal forms for solving polynomial systems: Generalized and efficient algorithms. Journal of Symbolic Computation, 102(??):63– 85, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301166.

Marche:2004:MIP

[MU04] Claude Marché and Xavier Urbain. Modular and incremental proofs of ACtermination. Journal of Symbolic Computation, 38(1): 873–897, July 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mu:2008:PAQ

[Mu08] Yan-Ping Mu. Parameter augmentation and the q-Gosper algorithm. Journal of Symbolic Computation, 43(12): 874–882, December 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mulmuley:1990:FPP

[Mul90] Ketan Mulmuley. A fast planar partition algorithm, I. Journal of Symbolic Computation, 10(3–4):253–280, September/October 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mulders:1997:NSL

[Mul97] Thom Mulders. A note on subresultants and the Lazard/ Rioboo/Trager formula in rational function integration. Journal of Symbolic Computation, 24(1):45–50, July 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mulders:2001:GSI

[Mul01] Thom Mulders. A generalized Sylvester identity and fraction-free random Gaussian elimination. Journal of Symbolic Computation, 31(4):447–460, April 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.

com/links/doi/10.1006/jsco.
2000.0409;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0409/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0409/ref.

Mulders:2004:CSL

[Mul04] Thom Mulders. Certified sparse linear system solving. Journal of Symbolic Computation, 38(5):1343– 1373, November 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Murashka:2023:FFG

[Mur23] Viachaslau I. Murashka. Formations of finite groups in polynomial time: F -residuals and F -subnormality. Journal of Symbolic Computation, 122(??):Article 102271, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000858.

Mustata:2000:LCM

[Mus00] Mircea Mustaţa. Local cohomology at monomial ideals. Journal of Symbolic Computation, 29(4–5):709–720, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0302;   http://www.
idealibrary.com/links/doi/

10.1006/jsco.1999.0302/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0302/ref.

Ma:1990:AEA

[Mv90] Keju Ma and Joachim von zur Gathen. Analysis of Euclidean algorithms for polynomials over finite fields. Journal of Symbolic Computation, 9(4):429–456 (or 429– 455??), April 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mihailescu:2010:EGS

[MV10] Preda Mihăilescu and Victor Vuletescu. Elliptic Gauss sums and applications to point counting. Journal of Symbolic Computation, 45(8): 825–836, August 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Mourrain:2013:HTA

[MV13] Bernard Mourrain and Nelly Villamizar. Homological techniques for the analysis of the dimension of triangular spline spaces. Journal of Symbolic Computation, 50 (??):564–577, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001708.

Milson:2015:PES

[MV15] Robert Milson and Francis Valiquette. Point equivalence of second-order ODEs: Maximal invariant classification order. Journal of Symbolic Computation, 67(??): 16–41, March/April 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114000534.

McMillan:1991:DSA

[MW91] Tim McMillan and Neil L. White. The dotted straightening algorithm. Journal of Symbolic Computation, 11(5–6):471–482, May/June 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Invarianttheoretic algorithms in geometry (Minneapolis, MN, 1987).

Montes:2010:GBP

[MW10] Antonio Montes and Michael Wibmer. Gröbner bases for polynomial systems with parameters. Journal of Symbolic Computation, 45(12):1391– 1425, December 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

McCallum:2012:DPT

[MW12] Scott McCallum and Volker Weispfenning. Deciding polynomial-transcendental problems. Journal of Symbolic Computation, 47(1):16– 31, January 2012. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001167.

Magron:2023:SOE

[MW23] Victor Magron and Jie Wang.

SONC optimization and exact nonnegativity certificates via second-order cone programming. Journal of Symbolic Computation, 115(??): 346–370, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000682.

Ma:2016:CSR

[MWZ16] Yue Ma, Chu Wang, and Lihong Zhi. A certificate for semidefinite relaxations in computing positivedimensional real radical ideals. Journal of Symbolic Computation, 72(??):1– 20, January/February 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001291.

Mohammed:2005:SUB

[MZ05] Mohamud Mohammed and Doron Zeilberger. Sharp upper bounds for the orders of the recurrences output by the Zeilberger and q-Zeilberger algorithms. Journal of Symbolic Computation, 39(2):201– 207, February 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Matzat:1987:PGG

[MZM87] B. Heinrich Matzat and Andreas Zeh-Marschke. Polynome mit der Galoisgruppe M_{11} über Q. (German) [polynomials with the Galois group M_{11} over Q]. Journal of Symbolic Computation, 4(1): 93–97, August 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nagasaka:2011:APG

[Nag11] Kosaku Nagasaka. Approximate polynomial GCD over integers. Journal of Symbolic Computation, 46(12): 1306–1317, December 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001234.

Nagasaka:2021:ASF

[Nag21a] Kosaku Nagasaka. Approximate square-free part and decomposition. Journal of Symbolic Computation, 104 (??):402–418, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300857.

Nagasaka:2021:TBA

[Nag21b] Kosaku Nagasaka. Toward the best algorithm for approximate GCD of univariate polynomials. Journal of Symbolic Computation, 105(??):4–27, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300419.

Nakagawa:2006:LS

[Nak06]  Koji Nakagawa.  Logico-

graphic symbols. Journal of Symbolic Computation, 41 (3–4):411–434, March/April 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nakayama:2009:ACL

[Nak09] Hiromasa Nakayama. Algorithm computing the local b function by an approximate division algorithm in D̂. Journal of Symbolic Computation, 44(5):449–462, May 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nakpim:2016:TOO

[Nak16] Warisa Nakpim. Third-order ordinary differential equations equivalent to linear secondorder ordinary differential equations via tangent transformations. Journal of Symbolic Computation, 77(??): 63–77, November/December 2016. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000079.

Naldi:2018:SRC

[Nal18]  Simone Naldi.  Solving

rank-constrained semidefinite programs in exact arithmetic. Journal of Symbolic Computation, 85(??): 206–223, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300743.

Nalbach:2024:LCS

[NÁS^{+}24] Jasper Nalbach, Erika Ábrahám, Philippe Specht, Christopher W. Brown, James H. Davenport, and Matthew England. Levelwise construction of a single cylindrical algebraic cell. Journal of Symbolic Computation, 123 (??):??, July/August 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123001025.

Nauheim:1998:SAE

[Nau98] Raphael Nauheim. Systems of algebraic equations with bad reduction. Journal of Symbolic Computation, 25(5):619– 642 (or 619–641??), May 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nawalaniec:2016:ACS

[Naw16] Wojciech Nawalaniec. Algorithms for computing symbolic representations of basic e-sums and their application to composites. Journal of Symbolic Computation, 74(??):328–345, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000747.

Ng:1989:CCU

[Ng89] Tze Beng Ng. Computation of the cohomology of BŜO_{n}⟨16⟩ for 23 ≤ n ≤ 26 using RE-DUCE. Journal of Symbolic Computation, 7(1):93–100 (or 93–99??), January 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Niederreiter:1993:FPF

[NG93] Harald Niederreiter and Rainer Göttfert. Factorization of polynomials over finite fields and characteristic sequences. Journal of Symbolic Computation, 16(5):401–412, November 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nguyen:2009:SLD

[Ngu09] An Khuong Nguyen. On dsolvability for linear differential equations. Journal of Symbolic Computation, 44(5): 421–434, May 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Niemeyer:1994:FSQ

[Nie94a] Alice C. Niemeyer. A finite soluble quotient algorithm. Journal of Symbolic Computation, 18(6):541–562 (or 541– 561??), December 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nievergelt:1994:CAP

[Nie94b] J. Nievergelt. Complexity, algorithms, programs, systems: The shifting focus. Journal of Symbolic Computation, 17 (4):297–310, April 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Niefield:2003:IFS

[Nie03] Susan B. Niefield. Implementing finite structures in Mathematica via a skeletal topos of finite sets. Journal of Symbolic Computation, 35(2):137– 151, February 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nie:2012:DNP

[Nie12] Jiawang Nie. Discriminants and nonnegative polynomials. Journal of Symbolic Computation, 47(2):167–191, February 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001350.

Nipkow:1991:CMA

[Nip91]  Tobias Nipkow.  Combin-

ing matching algorithms: The regular case. Journal of Symbolic Computation, 12(6):633– 654 (or 633–653??), December 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nielsen:1990:UCA

[NMM90] Glen C. Nielsen, Mark O.

McLinden, and Graham Morrison. Use of computer algebra to locate critical loci in fluid mixtures. Journal of Symbolic Computation, 10(5): 499–508, November 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nishiyama:2010:SAL

[NN10] Kenta Nishiyama and Masayuki Noro. Stratification associated with local b-functions. Journal of Symbolic Computation, 45(4):462–480, April 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Newman:1998:DGP

[NNN98] M. F. Newman, Werner Nickel, and Alice C. Niemeyer. Descriptions of groups of prime-power order. Journal of Symbolic Computation, 25 (5):665–682, May 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nejad:2019:BAS

[NNPZN19] Abbas Nasrollah Nejad, Ashkan Nikseresht, Ali Akbar Yazdan Pour, and Rashid Zaare-Nahandi. On the blowup of affine spaces along monomial ideals: Tameness. Journal of Symbolic Computation, 93 (??):85–99, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300385.

Ngo:2015:EDE

[NNvdPT15] L. X. Chau Ngo, K. A. Nguyen, M. van der Put, and J. Top. Equivalence of differential equations of order one. Journal of Symbolic Computation, 71(??): 47–59, November/December 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001096.

Narendran:1989:CFP

[NÓ89] Paliath Narendran and Colm Ó’Dúnlaing. Cancellativity in finitely presented semigroups. Journal of Symbolic Computation, 7(5):457–472, May 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nakamura:2010:ROS

[NOF10] Masaki Nakamura, Kazuhiro Ogata, and Kokichi Futatsugi. Reducibility of operation symbols in term rewriting systems and its application to behavioral specifications. Journal of Symbolic Computation, 45 (5):551–573, May 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Norton:1990:AAE

[Nor90]  G. H. Norton.  On the

asymptotic analysis of the Euclidean algorithm. Journal of Symbolic Computation, 10 (1):53–58, July 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Norman:1995:CDS

[Nor95a] A. C. Norman. Compact delivery support for REDUCE. Journal of Symbolic Computation, 19(1/2/3):133–143 (or 133–144??), January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Norton:1995:CDS

[Nor95b]  Graham Norton.  Corri-

genda: “On n-dimensional sequences. I” [J. Symbolic Comput. 20 (1995), no. 1, 71– 92, 1 374 225]. Journal of Symbolic Computation, 20 (5–6):769–770, 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [Nor95c].

Norton:1995:DS

[Nor95c]  Graham Norton.  On n-

dimensional sequences. I. Journal of Symbolic Computation, 20(1):71–92, July 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See corrigenda [Nor95b].

Norton:1995:MRF

[Nor95d] Graham Norton. On the minimal realizations of a finite sequence. Journal of Symbolic Computation, 20(1): 93–115, July 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Norton:1999:SLR

[Nor99] G. H. Norton. On shortest linear recurrences. Journal of Symbolic Computation, 27(3): 325–349, March 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Norton:2001:CM

[Nor01] Simon Norton. Computing in the monster. Journal of Symbolic Computation, 31(1–2): 193–201, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1008;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.1008/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1008/ref.

Nordbeck:2002:SBU

[Nor02] Patrik Nordbeck. SAGBI bases under composition. Journal of Symbolic Computation, 33(1):67–76, January 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0498;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0498/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0498/ref.

Noren:2015:TST

[Nor15] Patrik Norén. The three-state toric homogeneous Markov chain model has Markov degree two. Journal of Symbolic Computation, 68 (part 2) (??):285–296, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000820.

Nabeshima:2018:CGS

[NOT18] Katsusuke Nabeshima, Katsuyoshi Ohara, and Shinichi Tajima. Comprehensive Gröbner systems in PBW algebras, Bernstein–Sato ideals and holonomic D-modules. Journal of Symbolic Computation, 89(??):146–170, November/December 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117301177.

Nemes:1995:RMP

[NP95] István Nemes and Marko Petkovšek. RComp: a mathematica package for computing with recursive sequences. Journal of Symbolic Computation, 20(5–6):745–754 (or 745– 753??), November/December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Nguefack:2020:EBZ

[NP20] Bertrand Nguefack and Emmanuel Pola. Effective Buchberger–Zacharias–Weispfenning theory of skew polynomial extensions of subbilateral coherent rings. Journal of Symbolic Computation, 99(??): 50–107, July/August 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711930029X.

Neut:2009:ECG

[NPD09] Sylvain Neut, Michel Petitot, and Raouf Dridi. Élie Cartan’s geometrical vision or how to avoid expression swell. Journal of Symbolic Computation, 44(3):261–270, March 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Niemeyer:2017:CMI

[NPP17] Alice C. Niemeyer, Götz Pfeiffer, and Cheryl E. Praeger. On the complexity of multiplication in the Iwahori– Hecke algebra of the symmetric group. Journal of Symbolic Computation, 80 (part 3) (??):817–832, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711630092X.

Nieuwenhuis:1995:TPO

[NR95] Robert Nieuwenhuis and Albert Rubio. Theorem proving with ordering and equality constrained clauses. Journal of Symbolic Computation, 19(4):321–352 (or 321– 351??), April 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nieuwenhuis:1997:PBA

[NR97] Robert Nieuwenhuis and Albert Rubio. Paramodulation with built-in AC-theories and symbolic constraints. Journal of Symbolic Computation, 23(1):1–22 (or 1–21??), January 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nutt:1989:BNR

[NRS89] Werner Nutt, Pierre Réty, and Gert Smolka. Basic narrowing revisited. Journal of Symbolic Computation, 7(3–4):295–318 (or 295–317??), March/April 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nef:1990:CSP

[NS90] Walter Nef and Peter-Michael M. Schmidt. Computing a sweeping-plane in regular (“general”) position: a numerical and a symbolic solution. Journal of Symbolic Computation, 10(6):633– 646, December 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Navarro-Saad:1985:AFT

[NSW85] Miguel Navarro-Saad and Kurt Bernardo Wolf. Applications of a factorization theorem for ninth-order aberration optics. Journal of Symbolic Computation, 1(2): 235–240 (or 235–239??), June 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Nagasaka:2016:SIC

[NSW16] Kosaku Nagasaka, Agnes Szanto, and Franz Winkler.

Special issue on the conference ISSAC 2014: Symbolic computation and computer algebra. Journal of Symbolic Computation, 75 (??):1–3, July/August 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001030.

Nabeshima:2017:ALC

[NT17] Katsusuke Nabeshima and Shinichi Tajima. Algebraic local cohomology with parameters and parametric standard bases for zero-dimensional ideals. Journal of Symbolic Computation, 82(??): 91–122, September/October 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711730010X.

Nabeshima:2021:NAC

[NT21] Katsusuke Nabeshima and Shinichi Tajima. A new algorithm for computing logarithmic vector fields along an isolated singularity and Bruce–Roberts Milnor ideals. Journal of Symbolic Computation, 107(??):190–208, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000225.

Nikolaev:2018:SSP

[NU18] Andrey Nikolaev and Alexander Ushakov. Subset sum problem in polycyclic groups. Journal of Symbolic Computation, 84(??):84–94, January/February 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300354.

Nagy:2007:MLO

[NV07] Gábor P. Nagy and Petr Vojtěchovský. The Moufang loops of order 64 and 81. Journal of Symbolic Computation, 42(9):871–883, September 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ngo:2010:RGS

[NW10] L. X. Châu Ngô and Franz Winkler. Rational general solutions of first order nonautonomous parametrizable ODEs. Journal of Symbolic Computation, 45(12):1426– 1441, December 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ngo:2011:RGS

[NW11] L. X. Châu Ngô and Franz Winkler. Rational general solutions of planar rational systems of autonomous ODEs. Journal of Symbolic Computation, 46(10):1173–1186, October 2011. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000861.

Nishida:2021:CAC

[NWW21] Yuki Nishida, Sennosuke Watanabe, and Yoshihide Watanabe. Combinatorial algorithm for the computation of cyclically standard regular bracket monomials. Journal of Symbolic Computation, 103(??):46–65, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301361.

Noro:1999:MMC

[NY99] Masayuki Noro and Kazuhiro Yokoyama. A modular method to compute the rational univariate representation of zero-dimensional ideals. Journal of Symbolic Computation, 28(1–2): 243–264, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0275;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0275/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0275/ref.

Noro:2004:IPD

[NY04] Masayuki Noro and Kazuhiro Yokoyama. Implementation of prime decomposition of polynomial ideals over small finite fields. Journal of Symbolic Computation, 38 (4):1227–1246, October 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Oaku:2013:AIH

[Oak13]  Toshinori Oaku.  Algo-

rithms for integrals of holonomic functions over domains defined by polynomial inequalities. Journal of Symbolic Computation, 50(??):1–27, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112000946.

OBrien:1990:GGA

[O’B90]  E. A. O’Brien.  The p-

group generation algorithm. Journal of Symbolic Computation, 9(5–6):677–698, May/ June 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

OBrien:1993:ITG

[O’B93] E. A. O’Brien. Isomorphism testing for p-groups. Journal of Symbolic Computation, 16 (3):305–320, September 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

OBrien:1994:ITG

[O’B94]  E. A. O’Brien.  Isomor-

phism testing for p-groups. Journal of Symbolic Computation, 17(2):133–148 (or 133– 147??), February 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Otten:2003:LLC

[OB03] Jens Otten and Wolfgang Bibel. leanCoP: lean connectionbased theorem proving. Journal of Symbolic Computation, 36(1–2):139–161, July/ August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Olivieri:2003:ACP

[OdR03] Aurora Olivieri and Ángel del Río. An algorithm to compute the primitive central idempotents and the Wedderburn decomposition of a rational group algebra. Journal of Symbolic Computation, 35(6): 673–687, June 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Olteanu:2009:ACW

[OdR09] Gabriela Olteanu and Ángel del Río. An algorithm to compute the Wedderburn decomposition of semisimple group algebras implemented in the GAP package wedderga. Journal of Symbolic Computation, 44(5): 507–516, May 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ohlebusch:1995:MPC

[Ohl95] Enno Ohlebusch. Modular properties of composable term rewriting systems. Journal of Symbolic Computation, 20(1): 1–42 (or 1–41??), July 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ouchi:2008:RUR

[OK08] Koji Ouchi and John Keyser. Rational Univariate Reduction via toric resultants. Journal of Symbolic Computation, 43(11):811–844, November 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Oki:2023:CVD

[Oki23] Taihei Oki. Computing valuations of the Dieudonné determinants. Journal of Symbolic Computation, 116 (??):284–323, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122001079.

Otto:1998:ICS

[OKK98] F. Otto, M. Katsura, and Y. Kobayashi. Infinite convergent string-rewriting systems and cross-sections for finitely presented monoids. Journal of Symbolic Computation, 26(5): 621–648, November 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ollongren:1988:PRF

[Oll88] Alexander Ollongren. On a particular restricted five-body problem: an analysis with computer algebra. Journal of Symbolic Computation, 6(1): 117–127 (or 117–126??), August 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Olver:2003:MF

[Olv03]  Peter J. Olver.  Moving

frames. Journal of Symbolic Computation, 36(3–4): 501–512, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Oeding:2013:ETA

[OO13] Luke Oeding and Giorgio Ottaviani. Eigenvectors of tensors and algorithms for Waring decomposition. Journal of Symbolic Computation, 54(??):9–35, July 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000047.

Omana:2005:FPG

[OP05] José Méndez Omaña and Michael E. Pohst. Factoring polynomials over global fields II. Journal of Symbolic Computation, 40(6): 1325–1339, December 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Omodeo:1993:DAE

[OPP93] Eugenio G. Omodeo, Franco Parlamento, and Alberto Policriti. A derived algorithm for evaluating, ϵ-expressions over abstract sets. Journal of Symbolic Computation, 15(5– 6):673–704, May/June 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Orecchia:2001:IGU

[Ore01]  F. Orecchia.  Implicitiza-

tion of a general union of parametric varieties. Journal of Symbolic Computation, 31(3):343–356, March 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0428;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0428/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0428/ref.

Orejas:2011:SGA

[Ore11]  Fernando Orejas.  Sym-

bolic graphs for attributed graph constraints. Journal of Symbolic Computation, 46(3):294–315, March 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001689.

Orlitzky:2022:RCE

[Orl22] Michael Orlitzky. Rank computation in Euclidean Jordan algebras. Journal of Symbolic Computation, 113(??):181– 192, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000220.

OHearn:1992:RFF

[OS92] Peter W. O’Hearn and Zbigniew Stachniak. A resolution framework for finitely-valued first-order logics. Journal of Symbolic Computation, 13(3): 235–254, March 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Oaku:1994:GBM

[ŌS94] Toshinori Ōaku and Takeshi Shimoyama. A Gröbner basis method for modules over rings of differential operators. Journal of Symbolic Computation, 18(3):223–248, September 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Oliart:2004:FAU

[OS04a] Alberto Oliart and Wayne Snyder. Fast algorithms for uniform semi-unification. Journal of Symbolic Computation, 37(4):455–484, April 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Otto:2004:RRM

[OS04b] Friedrich Otto and Olga Sokratova. Reduction relations for monoid semirings. Journal of Symbolic Computation, 37(3):343–376, March 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ocansey:2024:RHP

[OS24] Evans Doe Ocansey and Carsten Schneider. Representation of hypergeometric products of higher nesting depths in difference rings. Journal of Symbolic Computation, 120(??):Article 102220, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000275.

Ostheimer:1999:PAP

[Ost99] Gretchen Ostheimer. Practical algorithms for polycyclic matrix groups. Journal of Symbolic Computation, 28(3):361–379, September 1999. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0287/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0287/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0287/production/ref.

Ogilvie:1987:ASC

[OT87] J. F. Ogilvie and R. H. Tipping. On the analytic solution by computer algebra of some problems in the vibrationrotational spectroscopy of diatomic molecules. Journal of Symbolic Computation, 3(3): 277–281, June 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Oaku:2001:MFR

[OT01] Toshinori Oaku and Nobuki Takayama. Minimal free resolutions of homogenized D-modules. Journal of Symbolic Computation, 32 (6):575–595, December 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0484;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0484/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0484/ref.

Oyono:2013:GAC

[OT13] Roger Oyono and Nicolas Thériault. Group arithmetic in C_{3,5} curves. Journal of Symbolic Computation, 56 (??):1–26, September 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000710.

Otto:1991:WES

[Ott91] Friedrich Otto. When is an extension of a specification consistent? decidable and undecidable cases. Journal of Symbolic Computation, 12(3):255– 274 (or 255–273??), September 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Oaku:2000:LAM

[OTW00] Toshinori Oaku, Nobuki Takayama, and Uli Walther. A localization algorithm for D-modules. Journal of Symbolic Computation, 29 (4–5):721–728, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0398;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0398/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0398/ref.

Odabas:2016:ICM

[OUI16] A. Odabas, E. Ö. Uslu, and E. Ilgaz. Isoclinism of crossed modules. Journal of Symbolic Computation, 74(??): 408–424, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000887.

Oussous:1991:MCL

[Ous91] Nour Eddine Oussous. Macsyma computation of local minimal realization of dynamical systems of which generating power series are finite. Journal of Symbolic Computation, 12(1):115–126, July 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Orr:1994:CAA

[OZ94] Craig R. Orr and Doron Zeilberger. A computer algebra approach to the discrete Dirichlet problem. Journal of Symbolic Computation, 18 (1):87–90, July 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Padawitz:1996:ITP

[Pad96] Peter Padawitz. Inductive theorem proving for design specifications. Journal of Symbolic Computation, 21(1): 41–100 (or 41–99??), January 1996. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-

855X (electronic).

Palancz:2013:ADR

[Pal13]  B. Paláncz.  Application

of Dixon resultant to satellite trajectory control by pole placement. Journal of Symbolic Computation, 50(??):79–99, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711200106X.

Pan:1989:BPI

[Pan89] Luquan Pan. On the D-basis of polynomial ideals over principal ideal domains. Journal of Symbolic Computation, 7 (1):55–70 (or 55–69??), January 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pan:1994:SMP

[Pan94] Victor Y. Pan. Simple multivariate polynomial multiplication. Journal of Symbolic Computation, 18(3):183– 186, September 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pan:1996:CAS

[Pan96] Victor Y. Pan. Computing x_{m} mod p(x) and an application to splitting a polynomial into factors over a fixed disc. Journal of Symbolic Computation, 22(4):377–380, October 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pan:2002:UPN

[Pan02]  Victor Y. Pan.  Univari-

ate polynomials: Nearly optimal algorithms for numerical factorization and rootfinding. Journal of Symbolic Computation, 33(5):701–733, May 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Papp:2023:DSN

[Pap23] Dávid Papp. Duality of sum of nonnegative circuit polynomials and optimal SONC bounds. Journal of Symbolic Computation, 114(??): 246–266, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000359.

Park:2004:SCS

[Par04]  Hyungju Park.  Symbolic

computation and signal processing. Journal of Symbolic Computation, 37(2):209– 226, February 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Park:2008:CAP

[Par08] Hyungju Park. Computer algebra in pure and applied mathematics: a foreword.

Journal of Symbolic Computation, 43(6–7):395–396, June/ July 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Park:2022:SMD

[Par22] Euisung Park. On surfaces of minimal degree in P_{5}. Journal of Symbolic Computation, 109 (??):116–123, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000547.

Pasztor:1986:NSA

[Pas86] Ana Pasztor. Non-standard algorithmic and dynamic logic. Journal of Symbolic Computation, 2(1):59– 81, March 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Paul:1985:EMF

[Pau85] Etienne Paul. Equational methods in first order predicate calculus. Journal of Symbolic Computation, 1(1): 7–29, March 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Paulson:1986:CRO

[Pau86] Lawrence C. Paulson. Constructing recursion operators in intuitionistic type theory. Journal of Symbolic Computation, 2(4):325–355, December 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pauer:1992:LIG

[Pau92a] Franz Pauer. On lucky ideals for Gröbner basis computations. Journal of Symbolic Computation, 14(5):471– 482, November 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Paul:1992:GRC

[Pau92b] Etienne Paul. A general refutational completeness result for an inference procedure based on associative — commutative unification. Journal of Symbolic Computation, 14 (6):577–618, December 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Paule:1995:GFF

[Pau95] Peter Paule. Greatest factorial factorization and symbolic summation. Journal of Symbolic Computation, 20(3):235– 268, September 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pauli:2001:FPL

[Pau01]  Sebastian Pauli.  Factor-

ing polynomials over local fields. Journal of Symbolic Computation, 32(5): 533–547, November 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0493;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0493/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0493/ref.

Pauer:2007:GBC

[Pau07] Franz Pauer. Gröbner bases with coefficients in rings. Journal of Symbolic Computation, 42(11–12):1003–1011, November/December 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Picart:2007:SSA

[PB07] Philippe Saux Picart and Cyril Brunie. Symmetric subresultants and applications. Journal of Symbolic Computation, 42(9):884–919, September 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Paternoster:1998:CIS

[PC98] B. Paternoster and M. Cafaro. Computation of the interval of stability of Runge–Kutta– Nyström methods. Journal of Symbolic Computation, 25(3): 383–394, March 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pison-Casares:2008:LS

[PCVT08] Pilar Pisón-Casares and Alberto Vigneron-Tenorio. On Lawrence semigroups. Journal of Symbolic Computation, 43 (11):804–810, November 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Perez-Diaz:2007:CSP

[PD07]  Sonia Pérez-Díaz.  Com-

putation of the singularities of parametric plane curves. Journal of Symbolic Computation, 42(8):835–857, August 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pickering:2024:EAI

[PdRAEC24] Lynn Pickering, Tereso del Río Almajano, Matthew England, and Kelly Cohen. Explainable AI insights for symbolic computation: a case study on selecting the variable ordering for cylindrical algebraic decomposition. Journal of Symbolic Computation, 123(??):??, July/August 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000901.

Perez-Diaz:2003:CAP

[PDS03] Sonia Pérez-Díaz and J. Rafael Sendra. Computing all parametric solutions for blending parametric surfaces. Journal of Symbolic Computation, 36 (6):925–964, December 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Perez-Diaz:2008:URB

[PDS08] Sonia Pérez-Díaz and J. Rafael Sendra. A univariate resultantbased implicitization algorithm for surfaces. Journal of Symbolic Computation, 43(2): 118–139, February 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pfalzgraf:2011:F

[PEGS11] Jochen Pfalzgraf, Hartmut Ehrig, Ulrike Golas, and Thomas Soboll. Foreword. Journal of Symbolic Computation, 46(3):219–221, March 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001641.

Peltier:1997:IMB

[Pel97] Nicolas Peltier. Increasing model building capabilities by constraint solving on terms with integer exponents. Journal of Symbolic Computation, 24(1):59–102 (or 59–101??), July 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Peltier:2003:CCR

[Pel03a] Nicolas Peltier. A calculus combining resolution and enumeration for building finite models. Journal of Symbolic Computation, 36(1– 2):49–77, July/August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Peltier:2003:MBO

[Pel03b] Nicolas Peltier. Model building with ordered resolution: extracting models from saturated clause sets. Journal of Symbolic Computation, 36(1– 2):5–48, July/August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Perdry:2004:SNR

[Per04]  Hervé Perdry.  Strongly

Noetherian rings and constructive ideal theory. Journal of Symbolic Computation, 37 (4):511–535, April 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Prebet:2024:CRU

[PES24] Rémi Prébet, Mohab Safey El Din, and Éric Schost. Computing roadmaps in unbounded smooth real algebraic sets I: Connectivity results. Journal of Symbolic Computation, 120(??):Article 102234, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000482.

Petho:1987:RTI

[Pet87] Attila Pethö. On the resolution of Thue inequalities. Journal of Symbolic Computation, 4(1):103–109, August 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Petkovsek:1992:HSL

[Pet92] Marko Petkovšek. Hypergeometric solutions of linear recurrences with polynomial coefficients. Journal of Symbolic Computation, 14(2–3): 243–264, August/September 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Petermann:2000:CCT

[Pet00] Uwe Petermann. Connection calculus theorem proving with multiple built-in theories. Journal of Symbolic Computation, 29(2):373–392, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0365;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0365/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0365/ref.

Peternell:2010:RTP

[Pet10]  Martin Peternell.  Ratio-

nal two-parameter families of spheres and rational offset surfaces. Journal of Symbolic Computation, 45(1):1– 18, January 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Plaisted:1986:SPC

[PG86] David A. Plaisted and Steven Greenbaum. A structure preserving clause form translation. Journal of Symbolic Computation, 2(3):293– 304, September 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Paige:1987:MTS

[PH87] Robert Paige and Fritz Henglein. Mechanical translation of set theoretic problem specifications into efficient RAM code — a case study. Journal of Symbolic Computation, 4(2):207–232, October 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Poole:2011:SCC

[PH11] Douglas Poole and Willy Hereman. Symbolic computation of conservation laws for nonlinear partial differential equations in multiple space dimensions. Journal of Symbolic Computation, 46(12):1355–1377, December 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711100126X.

Picart:1998:SCA

[Pic98] Ph. Saux Picart. The Schur– Cohn algorithm revisited. Journal of Symbolic Computation, 26(4):387–408, October 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pichler:2000:SAA

[Pic00] Reinhard Pichler. Speeding up algorithms on atomic representations of Herbrand models via new redundancy criteria. Journal of Symbolic Computation, 29(2): 213–257, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0361;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0361/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0361/ref.

Pichler:2003:CEP

[Pic03] Reinhard Pichler. On the complexity of equational problems in CNF. Journal of Symbolic Computation, 36 (1–2):235–269, July/August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pieper:2024:TNS

[Pie24]  Andreas Pieper.  Theta

nullvalues of supersingular Abelian varieties. Journal of Symbolic Computation, 123 (??):??, July/August 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123001104.

Pilnikova:2007:TCS

[Píl07] Jana Pílniková. Trivializing a central simple algebra of degree 4 over the rational numbers. Journal of Symbolic Computation, 42(6):579–586, June 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Piquette:1991:MSE

[Piq91] Jean C. Piquette. A method for symbolic evaluation of indefinite integrals containing special functions or their products. Journal of Symbolic Computation, 11(3):231– 250 (or 231–249??), March 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pisabarro:2004:CLI

[Pis04] María-Jesús Pisabarro. Computing lattice ideals of unions of monomial curves. Journal of Symbolic Computation, 38 (2):1025–1042, August 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Plesken:1987:TSQ

[Ple87] W. Plesken. Towards a soluble quotient algorithm. Journal of Symbolic Computation, 4(1): 111–122, August 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Paulin-Mohring:1993:SMP

[PMW93] Christine Paulin-Mohring and Benjamin Werner. Synthesis of ML programs in the system Coq. Journal of Symbolic Computation, 15(5– 6):607–640, May/June 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Policarpo:2013:HAE

[PNM13] H. Policarpo, M. M. Neves, and N. M. M. Maia. On a hybrid analytical-experimental technique to assess the storage modulus of resilient materials using symbolic computation. Journal of Symbolic Computation, 61–62(??):31–52, ???? 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001260.

Pohst:1987:MLR

[Poh87a] M. Pohst. A modification of the LLL-reduction algorithm. Journal of Symbolic Computation, 4(1):123–127, August 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pohst:1987:I

[Poh87b] Michael Pohst. Introduction.

Journal of Symbolic Computation, 4(1):1–2, 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pohst:1997:VCA

[Poh97] Michael E. Pohst. On validated computing in algebraic number fields. Journal of Symbolic Computation, 24(6):657–666 (or 657– 665??), December 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pohst:2005:FPG

[Poh05] Michael E. Pohst. Factoring polynomials over global fields I. Journal of Symbolic Computation, 39(6):617–630, June 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Poli:1995:DCN

[Pol95]  Alain Poli.  A determin-

istic construction of normal bases with complexity O(n^{3} + n log n log(log n) log q). Journal of Symbolic Computation, 19(4):305–320 (or 305– 319??), April 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pommaret:2011:MIS

[Pom11] J. F. Pommaret. Macaulay inverse systems revisited. Journal of Symbolic Computation, 46(9):1049–1069, September 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000782.

Ponder:1991:PMP

[Pon91] Carl G. Ponder. Parallel multiplication and powering of polynomials. Journal of Symbolic Computation, 11(4): 307–320, April 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Popescu:2015:ACH

[Pop15] Adrian Popescu. An algorithm to compute the Hilbert depth. Journal of Symbolic Computation, 66(??):1– 7, January/February 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000339.

Posur:2018:CEV

[Pos18] Sebastian Posur. Constructing equivariant vector bundles via the BGG correspondence. Journal of Symbolic Computation, 91(??):57–73, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300786.

Pozar:2015:SCA

[Poz15] Rok Pozar. Some computational aspects of solvable regular covers of graphs. Journal of Symbolic Computation, 70(??):1–13, September/October 2015. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000911.

Pozar:2019:CSE

[Poz19] Rok Pozar. Computing stable epimorphisms onto finite groups. Journal of Symbolic Computation, 92 (??):22–30, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117301244.

Pozar:2024:FCC

[Poz24] Rok Pozar. Fast computation of the centralizer of a permutation group in the symmetric group. Journal of Symbolic Computation, 123 (??):??, July/August 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123001013.

Pais:1991:UFP

[PP91a] John Pais and Gerald E. Peterson. Using forcing to prove completeness of resolution and paramodulation. Journal of Symbolic Computation, 11(1– 2):3–19 (or 3–20??), January/February 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Plaisted:1991:TRS

[PP91b] David A. Plaisted and Richard C. Potter. Term rewriting: Some experimental results. Journal of Symbolic Computation, 11(1–2):149–180, January/February 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Peternell:1997:CRP

[PP97] Martin Peternell and Helmut Pottmann. Computing rational parametrizations of canal surfaces. Journal of Symbolic Computation, 23(2– 3):255–266, February/March 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Pilehrood:2011:ABB

[PP11] Kh. Hessami Pilehrood and T. Hessami Pilehrood. A q-analogue of the Bailey– Borwein–Bradley identity. Journal of Symbolic Computation, 46(6):699–711, June 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000186.

Pfister:2017:CGN

[PP17] Gerhard Pfister and Dorin Popescu. Constructive General Neron Desingularization for one dimensional local rings. Journal of Symbolic Computation, 80 (part 3) (??):570–580, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300712.

Peter:2013:RSL

[PPR13] Thomas Peter, Gerlind Plonka, and Daniela Rosca. Representation of sparse Legendre expansions. Journal of Symbolic Computation, 50 (??):159–169, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001101.

Petrovic:2020:GBP

[PPR20] Zoran Z. Petrović, Branislav I.

Prvulović, and Marko Radovanović. Gröbner bases for (partial) flag manifolds. Journal of Symbolic Computation, 101(??):90–108, November/December 2020. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300628.

Poteaux:2012:GRP

[PR12] Adrien Poteaux and Marc Rybowicz. Good reduction of Puiseux series and applications. Journal of Symbolic Computation, 47 (1):32–63, January 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001209.

Perrucci:2022:NGF

[PR22] Daniel Perrucci and Marie-Françoise Roy. A new general formula for the Cauchy index on an interval with subresultants. Journal of Symbolic Computation, 109(??): 465–481, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300717.

Prank:2013:TES

[Pra13] Rein Prank. A tool for evaluating solution economy of algebraic transformations. Journal of Symbolic Computation, 61–62(??):100–115, ???? 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001338.

Previtali:2006:ICM

[Pre06] Andrea Previtali. Irreducible constituents of monomial representations. Journal of Symbolic Computation, 41(12): 1345–1359, December 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pritchard:1996:IMP

[Pri96]  F. Leon Pritchard.  The

ideal membership problem in non-commutative polynomial rings. Journal of Symbolic Computation, 22(1):27–48, July 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Prosper:2000:SPS

[Pro00] Vincent Prosper. SFA, a package on symmetric functions considered as operators over the ring of polynomials for the computer algebra system Maple. Journal of Symbolic Computation, 29(1):83–94, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0291;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0291/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0291/ref.

Poor:2018:AOA

[PRR18] Jamal Hossein Poor, Clemens G.

Raab, and Georg Regensburger. Algorithmic operator algebras via normal forms in tensor rings. Journal of Symbolic Computation, 85(??): 247–274, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300779.

Pittaluga:1989:COA

[PS89] Marilena Pittaluga and Elisabetta Strickland. A computer oriented algorithm for the determination of the dimension and character of a modular irreducible sl(n, K)- module. Journal of Symbolic Computation, 7(2):155– 162 (or 155–161??), February 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Puel:1993:CPM

[PS93] Laurence Puel and Ascánder Suárez. Compiling pattern matching by term decomposition. Journal of Symbolic Computation, 15(1):1– 26, January 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Paule:1995:SIS

[PS95a] P. Paule and V. Strehl. Special issue: Symbolic computation in combinatorics Δ_{1}: Foreword of the Guest Editors. Journal of Symbolic Computation, 20(5–6):483– 486, November/December 1995.

Paule:1995:MVZ

[PS95b] Peter Paule and Markus A.

Schorn. A mathematica version of Zeilberger’s algorithm for proving binomial coefficient identities. Journal of Symbolic Computation, 20(5–6):673–698, November/ December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Pirastu:1995:PCI

[PS95c] Roberto Pirastu and Kurt Siegl. Parallel computation and indefinite summation: a ∥MAPLE∥ application for the rational case. Journal of Symbolic Computation, 20(5–6):603–616, November/December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Pirastu:1995:RSG

[PS95d] Roberto Pirastu and Volker Strehl. Rational summation and Gosper–Petkovšek representation. Journal of Symbolic Computation, 20(5–6):617– 636 (or 617–635??), November/December 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Policriti:1995:EP

[PS95e] Alberto Policriti and Jacob T. Schwartz. T -theorem proving. I. Journal of Symbolic Computation, 20(3):315– 342, September 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Plesken:1997:AFP

[PS97a] W. Plesken and B. Souvignier. Analysing finitely presented groups by constructing representations. Journal of Symbolic Computation, 24(3–4):335–350 (or 335– 349??), September/October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Plesken:1997:CIL

[PS97b] W. Plesken and B. Souvignier. Computing isometries of lattices. Journal of Symbolic Computation, 24(3–4): 327–334, September/October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Pau:2000:QET

[PS00] Petru Pau and Josef Schicho. Quantifier elimination for trigonometric polynomials by cylindrical trigonometric decomposition. Journal of Symbolic Computation, 29(6):971–983, June 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0352;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0352/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0352/ref.

Pope:2009:NMS

[PS09] Scott R. Pope and Agnes Szanto. Nearest multivariate system with given root multiplicities. Journal of Symbolic Computation, 44(6):606–625, June 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Poteaux:2013:CCZ

[PS13] Adrien Poteaux and Éric Schost. On the complexity of computing with zerodimensional triangular sets. Journal of Symbolic Computation, 50(??):110–138, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001083.

Panina:2018:MPC

[PS18a] Gaiane Panina and Dirk Siersma. Motion planning and control of a planar polygonal linkage. Journal of Symbolic Computation, 88(??):5– 20, September/October 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300129.

Pernet:2018:TSE

[PS18b] Clément Pernet and Arne Storjohann. Time and space efficient generators for quasiseparable matrices. Journal of Symbolic Computation, 85(??):224–246, March/April 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300767.

Prabhakar:2020:IBA

[PS20] Swaroop N. Prabhakar and Vikram Sharma. Improved bounds on absolute positiveness of multivariate polynomials. Journal of Symbolic Computation, 101(??):170–188, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300914.

Pfister:2021:PDS

[PS21] Gerhard Pfister and Andreas Steenpaß. On the primary decomposition of some determinantal hyperedge ideal. Journal of Symbolic Computation, 103(??):14–21, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301348.

Portakal:2023:NCI

[PSA23] Irem Portakal and Javier Sendra-Arranz. Nash conditional independence curve. Journal of Symbolic Computation, 122(??):Article 102255, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712300069X.

Peltier:2012:FOT

[PSS12] Nicolas Peltier and Viorica Sofronie-Stokkermans. Firstorder theorem proving: Foreword. Journal of Symbolic Computation, 47(9): 1009–1010, September 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002239.

Plaumann:2011:QCT

[PSV11] Daniel Plaumann, Bernd Sturmfels, and Cynthia Vinzant. Quartic curves and their bitangents. Journal of Symbolic Computation, 46(6):712–733, June 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000198.

Potocnik:2013:CVT

[PSV13] Primoz Potocnik, Pablo Spiga, and Gabriel Verret. Cubic vertex-transitive graphs on up to 1280 vertices. Journal of Symbolic Computation, 50(??):465–477, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001563.

Pasini:1991:SCO

[PSZ91] P. Pasini, F. Semeria, and C. Zannoni. Symbolic computation of orientational correlation function moments. Journal of Symbolic Computation, 12(2):221–231 (or 221– 232??), August 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Petho:1998:TPQ

[PT98] Attila Pethö and Robert F. Tichy. On two-parametric quartic families of Diophantine problems. Journal of Symbolic Computation, 26 (2):151–171, 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pillwein:2014:LFC

[PT14] Veronika Pillwein and Stefan Takacs. A local Fourier convergence analysis of a multigrid method using symbolic computation. Journal of Symbolic Computation, 63(??):1–20, May 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001752.

Pan:2016:NOR

[PT16] Victor Y. Pan and Elias P.

Tsigaridas. Nearly optimal refinement of real roots of a univariate polynomial. Journal of Symbolic Computation, 74(??):181–204, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000590.

Puel:1989:UUS

[Pue89] Laurence Puel. Using unavoidable set of trees to generalize Kruskal’s theorem. Journal of Symbolic Computation, 8(4):335–382, October 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Puschel:2002:DMR

[Püs02] Markus Püschel. Decomposing monomial representations of solvable groups. Journal of Symbolic Computation, 34(6): 561–596, December 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Poulakis:2000:PSG

[PV00] Dimitrios Poulakis and Evaggelos Voskos. On the practical solution of genus zero Diophantine equations. Journal of Symbolic Computation, 30(5):573–582, October 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0420;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0420/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0420/ref.

Poulakis:2002:SGZ

[PV02] Dimitrios Poulakis and Evaggelos Voskos. Solving genus zero Diophantine equations with at most two infinite valuations. Journal of Symbolic Computation, 33(4):479–491, April 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Phillips:2005:LGA

[PV05] J. D. Phillips and Petr Vojtěchovský. Linear groupoids and the associated wreath products. Journal of Symbolic Computation, 40(3): 1106–1125, September 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Plaumann:2013:DRH

[PV13] Daniel Plaumann and Cynthia Vinzant. Determinantal representations of hyperbolic plane curves: an elementary approach. Journal of Symbolic Computation, 57(??):48–60, October 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000679.

Pavelle:1985:M

[PW85] Richard Pavelle and Paul S.

Wang. MACSYMA from F to G. Journal of Symbolic Computation, 1(1):69–100, March 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Parker:1990:CCM

[PW90] R. A. Parker and R. A. Wilson. The computer construction of matrix representations of finite groups over finite fields. Journal of Symbolic Computation, 9(5– 6):583–590, May/June 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Priestley:1994:MBA

[PW94] H. A. Priestley and M. P.

Ward. A multipurpose backtracking algorithm. Journal of Symbolic Computation, 18 (1):1–40, July 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pistone:2006:CV

[PW06] Giovanni Pistone and Henry P. Wynn. Cumulant varieties. Journal of Symbolic Computation, 41(2):210–221, February 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Puchinger:2018:FOL

[PWZ18] Sven Puchinger and Antonia Wachter-Zeh. Fast operations on linearized polynomials and their applications in coding theory. Journal of Symbolic Computation, 89(??):194–215, November/December 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117301190.

Pedersen:1994:TRC

[PY94] John Pedersen and Margaret Yoder. Term rewriting for the conjugacy problem and the Braid groups. Journal of Symbolic Computation, 18(6): 563–572, December 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Palacian:2005:AIT

[PY05] Jesús F. Palacián and Patricia Yanguas. Asymptotic invariant tori of perturbed two-body problems. Journal of Symbolic Computation, 40(4–5):1256–1268, October/November 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Paule:1992:SIJ

[PZ92] P. Paule and D. Zeilberger.

Special issue of JSC on “Symbolic Computation in Combinatorics” — foreward of the Guest Editors. Journal of Symbolic Computation, 14(2–3):123–126, August/September 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Pauer:1996:GBR

[PZ96] Franz Pauer and Sandro Zampieri. Gröbner bases with respect to generalized term orders and their application to the modelling problem. Journal of Symbolic Computation, 21(2):155–168, February 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Prest:2012:NLP

[PZ12] Thomas Prest and Paul Zimmermann. Non-linear polynomial selection for the number field sieve. Journal of Symbolic Computation, 47(4):401–409, April 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001398.

Qin:2013:LIS

[QHL^{+}13] Shengchao Qin, Guanhua He, Chenguang Luo, Wei-Ngan Chin, and Xin Chen. Loop invariant synthesis in a combined abstract domain. Journal of Symbolic Computation, 50(??):386–408, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711200140X.

Qi:2005:ERS

[Qi05] Liqun Qi. Eigenvalues of a real supersymmetric tensor. Journal of Symbolic Computation, 40(6):1302–1324, December 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Qi:2006:RES

[Qi06] Liqun Qi. Rank and eigenvalues of a supersymmetric tensor, the multivariate homogeneous polynomial and the algebraic hypersurface it defines. Journal of Symbolic Computation, 41(12):1309– 1327, December 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Qi:2023:TBA

[Qi23] Jiayue Qi. A tree-based algorithm for the integration of monomials in the Chow ring of the moduli space of stable marked curves of genus zero.

Journal of Symbolic Computation, 122(??):Article 102253, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000676.

Quadrat:2007:CBF

[QR07] Alban Quadrat and Daniel Robertz. Computation of bases of free modules over the Weyl algebras. Journal of Symbolic Computation, 42(11–12):1113–1141, November/December 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Quaresma:2019:TGP

[QSGB19] Pedro Quaresma, Vanda Santos, Pierluigi Graziani, and Nuno Baeta. Taxonomies of geometric problems. Journal of Symbolic Computation, 97(??):31–55, ???? 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118301305.

Qureshi:2017:CIO

[Qur17] Muhammad Imran Qureshi.

Computing isolated orbifolds in weighted flag varieties. Journal of Symbolic Computation, 79 (part 2)(??): 457–474, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116000341.

Qian:1996:MHO

[QW96] Zhenyu Qian and Kang Wang. Modular higher-order equational preunification. Journal of Symbolic Computation, 22(4):401–424, October 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Qin:2022:HSE

[QZYC22] Xiaolin Qin, Lige Zhang, Lu Yang, and Sheng Cao. Heuristics to sift extraneous factors in Dixon resultants. Journal of Symbolic Computation, 112(??): 105–121, September/October 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000037.

Raab:2012:UGB

[Raa12]  Clemens G. Raab.  Us-

ing Gröbner bases for finding the logarithmic part of the integral of transcendental functions. Journal of Symbolic Computation, 47(10): 1290–1296, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002483.

Radu:2015:AAR

[Rad15] Cristian-Silviu Radu. An algorithmic approach to Ramanujan– Kolberg identities. Journal of Symbolic Computation, 68 (part 1)(??):225–253, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000868.

Rajaee:2006:NAG

[Raj06] Saeed Rajaee. Non-associative Gröbner bases. Journal of Symbolic Computation, 41(8): 887–904, August 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ranise:2012:VSA

[Ran12] Silvio Ranise. On the verification of security-aware Eservices. Journal of Symbolic Computation, 47(9): 1066–1088, September 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002264.

Rapallo:2006:MBS

[Rap06] Fabio Rapallo. Markov bases and structural zeros. Journal of Symbolic Computation, 41(2):164–172, February 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ratschan:2002:QCU

[Rat02] Stefan Ratschan. Quantified constraints under perturbation. Journal of Symbolic Computation, 33(4): 493–505, April 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Raum:2011:HIM

[Rau11] Martin Raum. How to implement a modular form. Journal of Symbolic Computation, 46(12):1336–1354, December 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001258.

Rodriguez-Carbonell:2007:GAP

[RCK07] E. Rodríguez-Carbonell and D. Kapur. Generating all polynomial invariants in simple loops. Journal of Symbolic Computation, 42(4): 443–476, April 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Roque:1991:CAS

[Rd91] Waldir L. Roque and Renato P. dos Santos. Computer algebra in spacetime embedding. Journal of Symbolic Computation, 12(3):381– 389, September 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Roune:2013:CAE

[RdC13] Bjarke Hammersholt Roune and Eduardo Sáenz de Cabezón. Complexity and algorithms for Euler characteristic of simplicial complexes. Journal of Symbolic Computation, 50(??):170–196, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001198.

Roney-Dougal:2003:APP

[RDU03] Colva M. Roney-Dougal and William R. Unger. The affine primitive permutation groups of degree less than 1000. Journal of Symbolic Computation, 35(4):421–439, April 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Reeves:1998:PIB

[Ree98] Alyson A. Reeves. A parallel implementation of Buchberger’s algorithm over Z_{p} for p ≤ 31991. Journal of Symbolic Computation, 26(2):229– 244, 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Reinhart:1999:SKC

[Rei99] Georg M. Reinhart. The Schmidt-Kolchin conjecture. Journal of Symbolic Computation, 28(4–5):611–630, October/November 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0317/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0317/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0317/production/ref;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0318/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0318/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0318/production/ref.

Reinert:2006:GBF

[Rei06] Birgit Reinert. Gröbner bases in function rings — a guide for introducing reduction relations to algebraic structures. Journal of Symbolic Computation, 41(11):1264– 1294, November 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Renegar:1992:CCGa

[Ren92a] James Renegar. On the computational complexity and geometry of the first-order theory of the reals, Part I: Introduction. Preliminaries. The geometry of semi-algebraic sets. The decision problem for the existential theory of the reals. Journal of Symbolic Computation, 13(3):255–299 (or 255–300??), March 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Renegar:1992:CCGb

[Ren92b] James Renegar. On the computational complexity and geometry of the first-order theory of the reals, Part II: The general decision problem. Preliminaries for quantifier elimination. Journal of Symbolic Computation, 13(3):301– 328 (or 301–327??), March 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Renegar:1992:CCGc

[Ren92c] James Renegar. On the computational complexity and geometry of the first-order theory of the reals. Part III: Quantifier elimination. Journal of Symbolic Computation, 13(3):329–352, March 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rennert:2004:PMM

[Ren04] Nicolas Rennert. A parallel multi-modular algorithm for computing Lagrange resolvents. Journal of Symbolic Computation, 37(5):547–556, May 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Renaudineau:2017:TCF

[Ren17] Arthur Renaudineau. A tropical construction of a family of real reducible curves. Journal of Symbolic Computation, 80 (part 2)(??):251–272, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300037.

Rashid:2018:FAC

[RH18] Adnan Rashid and Osman Hasan. Formal analysis of continuous-time systems using Fourier transform. Journal of Symbolic Computation, 90(??):65–88, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300336.

Richardson:1991:WMK

[Ric91]  Daniel Richardson.  Wu’s

method and the Khovanskiĭ finiteness theorem. Journal of Symbolic Computation, 12(2): 127–142 (or 127–141??), August 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Richardson:1992:CTB

[Ric92a] Daniel Richardson. Computing the topology of a bounded non-algebraic curve in the plane. Journal of Symbolic Computation, 14(6):619– 644 (or 619–643??), December 1992. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Richardson:1992:BBP

[Ric92b] J. Richardson. The Blockhandler and the Bitfield package. Journal of Symbolic Computation, 14(1):93–102 (or 93– 101??), July 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Richardson:1997:HRZ

[Ric97]  Daniel Richardson.  How

to recognize zero. Journal of Symbolic Computation, 24(6):627–646 (or 627– 645??), December 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Richardson:1999:WWS

[Ric99] Dan Richardson. Weak Wu stratification in R^{n}. Journal of Symbolic Computation, 28 (1–2):213–223, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0273;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0273/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0273/ref.

Rieger:1993:CVG

[Rie93] J. H. Rieger. Computing view graphs of algebraic surfaces.

Journal of Symbolic Computation, 16(3):259–272, September 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Riese:2003:QPP

[Rie03] Axel Riese. qMultiSum — a package for proving qhypergeometric multiple summation identities. Journal of Symbolic Computation, 35(3): 349–376, March 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rincon:2013:CTL

[Rin13] Felipe Rincón. Computing tropical linear spaces. Journal of Symbolic Computation, 51(??):86–98, April 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001174.

Rioboo:2003:TFR

[Rio03] Renaud Rioboo. Towards faster real algebraic numbers. Journal of Symbolic Computation, 36(3–4): 513–533, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Risler:1988:SAC

[Ris88] Jean-Jacques J. Risler. Some aspects of complexity in real algebraic geometry. Journal of Symbolic Computation, 5 (1–2):109–120 (or 109–119??), February/April 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Robbiano:1986:TGS

[Rob86] Lorenzo Robbiano. On the theory of graded structures. Journal of Symbolic Computation, 2(2):139–170, June 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Robertson:1988:TTW

[Rob88] Edmund F. Robertson. Tietze transformations with weighted substring search. Journal of Symbolic Computation, 6(1):59–64, August 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Robidoux:1997:CAI

[Rob97] N. Robidoux. Computer algebra and interpolation: a lesson plan. Journal of Symbolic Computation, 23(5– 6):551–576, May/June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Roblot:2004:PFA

[Rob04] Xavier-François Roblot. Polynomial factorization algorithms over number fields. Journal of Symbolic Computation, 38(5):1429–1443, November 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Robertz:2009:NNG

[Rob09] Daniel Robertz. Noether normalization guided by monomial cone decompositions. Journal of Symbolic Computation, 44(10):1359–1373, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Roche:2011:CES

[Roc11] Daniel S. Roche. Chunky and equal-spaced polynomial multiplication. Journal of Symbolic Computation, 46(7):791–806, July 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001434.

Rochera:2022:AEC

[Roc22] David Rochera. Algebraic equations for constant width curves and Zindler curves. Journal of Symbolic Computation, 113(??):139–147, November/December 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000190.

Rodriguez:2015:CEI

[Rod15] Jose Israel Rodriguez. Combinatorial excess intersection. Journal of Symbolic Computation, 68 (part 2)(??):

297–307, May/June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001023.

Rojas:1999:SDS

[Roj99]  J. Maurice Rojas.  Solv-

ing degenerate sparse polynomial systems faster. Journal of Symbolic Computation, 28 (1–2):155–186, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0271;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1998.0271/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0271/ref.

Rolletschek:1986:NDE

[Rol86] Heinrich Rolletschek. On the number of divisions of the Euclidean algorithm applied to Gaussian integers. Journal of Symbolic Computation, 2(3):261–292 (or 261– 291??), September 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rolletschek:1990:SDC

[Rol90] Heinrich Rolletschek. Shortest division chains in imaginary quadratic number fields. Journal of Symbolic Computation, 9(3):321–354, March 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ronyai:1990:CSF

[Rón90] Lajos Rónyai. Computing the structure of finite algebras. Journal of Symbolic Computation, 9(3):355–374 (or 355– 373??), March 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Roozemond:2013:CSM

[Roo13] Dan Roozemond. Computing split maximal toral subalgebras of Lie algebras over fields of small characteristic. Journal of Symbolic Computation, 50(??):335–349, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711200137X.

Roque:2013:SNA

[Roq13] C. M. C. Roque. Symbolic and numerical analysis of plates in bending using Matlab. Journal of Symbolic Computation, 61–62(??):3–11, ???? 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001247.

Rosenmann:1993:ACG

[Ros93]  Amnon Rosenmann.  An

algorithm for constructing Gröbner and free Schreier bases in free group algebras. Journal of Symbolic Computation, 16(6):523–550 (or 523– 549??), December 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rosenkranz:2005:NSM

[Ros05] Markus Rosenkranz. A new symbolic method for solving linear two-point boundary value problems on the level of operators. Journal of Symbolic Computation, 39(2):171– 199, February 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rosowski:2023:FCM

[Ros23]  Andreas Rosowski.  Fast

commutative matrix algorithms. Journal of Symbolic Computation, 114(??): 302–321, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000499.

Roune:2008:STD

[Rou08] Bjarke Hammersholt Roune.

Solving thousand-digit Frobenius problems using Gröbner bases. Journal of Symbolic Computation, 43(1):1–7, January 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Roune:2009:SAI

[Rou09] Bjarke Hammersholt Roune. The Slice Algorithm for irreducible decomposition of monomial ideals. Journal of Symbolic Computation, 44(4): 358–381, April 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Royle:1987:TGD

[Roy87] Gordon F. Royle. The transitive groups of degree twelve. Journal of Symbolic Computation, 4(2):255–268, October 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Royle:1989:CVT

[RP89] Gordon F. Royle and Cheryl E. Präger. Constructing the vertex-transitive graphs of order 24. Journal of Symbolic Computation, 8(4):309– 326, October 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ramesh:1990:PTP

[RR90a] R. Ramesh and I. V. Ramakrishnan. Parallel tree pattern matching. Journal of Symbolic Computation, 9(4): 485–502 (or 485–501??), April 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Risler:1990:TP

[RR90b] Jean-Jacques J. Risler and Felice Ronga. Testing polynomials. Journal of Symbolic Computation, 10(1):1–5, July 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Roman:2005:GBS

[RR05] Manuel García Román and Socorro García Román. Gröbner bases and syzygies on bimodules over PBW algebras. Journal of Symbolic Computation, 40(3):1039–1052, September 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rosenkranz:2008:SFB

[RR08] Markus Rosenkranz and Georg Regensburger. Solving and factoring boundary problems for linear ordinary differential equations in differential algebras. Journal of Symbolic Computation, 43(8): 515–544, August 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Romero:2012:CHG

[RR12] Ana Romero and Julio Rubio. Computing the homology of groups: the geometric way. Journal of Symbolic Computation, 47(7):752–770, July 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001994.

Romero:2006:CSS

[RRS06] A. Romero, J. Rubio, and F. Sergeraert. Computing spectral sequences. Journal of Symbolic Computation, 41(10):1059–1079, October 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Romero:2019:IEH

[RRS19] Ana Romero, Julio Rubio, and Francis Sergeraert. An implementation of effective homotopy of fibrations. Journal of Symbolic Computation, 94(??): 149–172, September/October 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300920.

Roy:1990:CCR

[RS90] Marie-Françoise F. Roy and Aviva Szpirglas. Complexity of computation on real algebraic numbers. Journal of Symbolic Computation, 10(1): 39–52 (or 39–51??), July 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Reeves:1993:NPR

[RS93] Alyson Reeves and Bernd Sturmfels. A note on polynomial reduction. Journal of Symbolic Computation, 16(3):273–278 (or 273– 277??), September 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Recio:1997:RRR

[RS97] Tomas Recio and J. Rafael Sendra. Real reparametrizations of real curves. Journal of Symbolic Computation, 23(2– 3):241–254, February/March 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Rees:2000:AAF

[RS00] Sarah Rees and Leonard H.

Soicher. An algorithmic approach to fundamental groups and covers of combinatorial cell complexes. Journal of Symbolic Computation, 29(1):59–77, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0292;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0292/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0292/ref.

Rueda:2010:LCD

[RS10] Sonia L. Rueda and J. Rafael Sendra. Linear complete differential resultants and the implicitization of linear DPPEs. Journal of Symbolic Computation, 45(3):324– 341, March 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See corrigendum [RS11b].

Roy:2011:SDS

[RS11a] Marie-Françoise Roy and Aviva Szpirglas. Sylvester double sums and subresultants. Journal of Symbolic Computation, 46(4): 385–395, April 2011. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001872.

Rueda:2011:CLC

[RS11b] Sonia L. Rueda and J. Rafael Sendra. Corrigendum to “Linear complete differential resultants and the implicitization of linear DPPEs” [J. Symbolic Comput. 45 (3) March (2010) 324–341]. Journal of Symbolic Computation, 46(1):91–92, January 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711000132X.

See [RS10].

Roy:2013:NFE

[RS13] Marie-Françoise Roy and Sidi Mohamed Sedjelmaci. New fast Euclidean algorithms. Journal of Symbolic Computation, 50(??): 208–226, March 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001216.

Rauh:2016:LMB

[RS16] Johannes Rauh and Seth Sullivant. Lifting Markov bases and higher codimension toric fiber products. Journal of Symbolic Computation, 74 (??):276–307, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000723.

Rosenkranz:2019:IDS

[RS19] Markus Rosenkranz and Nitin Serwa. An integro-differential structure for Dirac distributions. Journal of Symbolic Computation, 92(??): 156–189, May/June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300105.

Roy:2020:SDS

[RS20] Marie-Françoise Roy and Aviva Szpirglas. Sylvester double sums, subresultants and symmetric multivariate Hermite interpolation. Journal of Symbolic Computation, 96(??):85–107, January/February 2020. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300240.

Radu:2021:MVP

[RS21a] Cristian-Silviu Radu and Nicolas Allen Smoot. A method of verifying partition congruences by symbolic computation. Journal of Symbolic Computation, 104(??):105–133, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300237.

Rosenkilde:2021:ASH

[RS21b] Johan Rosenkilde and Arne Storjohann. Algorithms for simultaneous Hermite– Padé approximations. Journal of Symbolic Computation, 102(??):279–303, January/February 2021. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119301300.

Rybalchenko:2010:CSI

[RSS10] Andrey Rybalchenko and Viorica Sofronie-Stokkermans. Constraint solving for interpolation. Journal of Symbolic Computation, 45(11):1212– 1233, November 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rueda:2013:APA

[RSS13] Sonia L. Rueda, Juana Sendra, and J. Rafael Sendra. An algorithm to parametrize approximately space curves. Journal of Symbolic Computation, 56(??):80–106, September 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000618.

Rudnicki:2001:CAM

[RST01] Piotr Rudnicki, Christoph Schwarzweller, and Andrzej Trybulec. Commutative algebra in the Mizar system. Journal of Symbolic Computation, 32(1–2): 143–169, July/August 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0456;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0456/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0456/ref.

Recio:2016:TVS

[RSTV16] Tomás Recio, J. Rafael Sendra, Luis-Felipe Tabera, and Carlos Villarino. On tubular vs. swung surfaces. Journal of Symbolic Computation, 72(??):55–64, January/February 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711400131X.

Rubio:2009:DRS

[RSV09] R. Rubio, J. M. Serradilla, and M. P. Vélez. Detecting real singularities of a space curve from a real rational parametrization. Journal of Symbolic Computation, 44(5): 490–498, May 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rizzi:1985:USC

[RT85] Nicola Rizzi and Amabile Tatone. Using symbolic computation in buckling analysis. Journal of Symbolic Computation, 1(3):317–322 (or 317– 321??), September 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ronveaux:1989:DES

[RT89] A. Ronveaux and G. Thiry. Differential equations of some orthogonal families in RE-DUCE. Journal of Symbolic Computation, 8(5):537– 541, November 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rylands:1998:MGO

[RT98] L. J. Rylands and D. E. Taylor. Matrix generators for the orthogonal groups. Journal of Symbolic Computation, 25(3): 351–360, March 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rodriguez:2017:PAC

[RT17] Jose Israel Rodriguez and Xiaoxian Tang. A probabilistic algorithm for computing data-discriminants of likelihood equations. Journal of Symbolic Computation, 83(??):342–364, November/December 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301389.

Ruano:2009:SGT

[Rua09] Diego Ruano. On the structure of generalized toric codes. Journal of Symbolic Computation, 44(5):499–506, May 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rueda:2011:PDR

[Rue11] Sonia L. Rueda. A perturbed differential resultant based implicitization algorithm for linear DPPEs. Journal of Symbolic Computation, 46(9):977–996, September 2011. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111000617.

Rupprecht:2004:SNA

[Rup04]  David Rupprecht.  Semi-

numerical absolute factorization of polynomials with integer coefficients. Journal of Symbolic Computation, 37(5): 557–574, May 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rusinowitch:1987:PSO

[Rus87] Michael Rusinowitch. Path of subterms ordering and recursive decomposition ordering revisited. Journal of Symbolic Computation, 3(1– 2):117–132 (or 117–131??), February/April 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Rewriting techniques and applications (Dijon, 1985).

Rusinowitch:1991:TPR

[Rus91] Michael Rusinowitch. Theoremproving with resolution and superposition. Journal of Symbolic Computation, 11(1– 2):21–50 (or 21–49??), January/February 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rutman:1992:GBP

[Rut92] Elizabeth W. Rutman. Gröbner bases and primary decomposition of modules. Journal of Symbolic Computation, 14(5):483–504 (or 483– 503??), November 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rutman:1993:PDM

[Rut93] Elizabeth W. Rutman. Primary decomposition of modules: Two variables over a field. Journal of Symbolic Computation, 15(3):267–276 (or 267–275??), March 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Riazanov:2003:LRS

[RV03] Alexandre Riazanov and Andrei Voronkov. Limited resource strategy in resolution theorem proving. Journal of Symbolic Computation, 36 (1–2):101–115, July/August 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rety:2005:TAR

[RV05] Pierre Réty and Julie Vuotto.

Tree automata for rewrite strategies. Journal of Symbolic Computation, 40(1): 749–794, July 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ratner:1990:PRR

[RW90] Daniel Ratner and Manfred Warmuth. The (n^{2} − 1)- puzzle and related relocation problems. Journal of Symbolic Computation, 10(2):111– 137 (or 111–136??), August 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Raymond:1994:GCL

[RW94] D. Raymond and D. Wood. Grail: A C++ library for automata and expressions. Journal of Symbolic Computation, 17(4):341–350, April 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ryba:1990:CCM

[Ryb90]  A. J. E. Ryba.  Com-

puter condensation of modular representations. Journal of Symbolic Computation, 9(5– 6):591–600, May/June 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Ryba:2001:CST

[Ryb01] A. J. E. Ryba. Condensation of symmetrized tensor powers. Journal of Symbolic Computation, 32 (3):273–289, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.

com/links/doi/10.1006/jsco.
2001.0459;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0459/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0459/ref.

Rybowicz:2003:NNF

[Ryb03] Marc Rybowicz. On the normalization of numbers and functions defined by radicals. Journal of Symbolic Computation, 35(6):651–672, June 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Reid:2009:SPS

[RZ09] Greg Reid and Lihong Zhi.

Solving polynomial systems via symbolic-numeric reduction to geometric involutive form. Journal of Symbolic Computation, 44(3):280– 291, March 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ronveaux:1999:DPR

[RZAG99] A. Ronveaux, A. Zarzo, I. Area, and E. Godoy. Decomposition of polynomials with respect to the cyclic group of order m. Journal of Symbolic Computation, 28(6):755–765, December 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.

                                        1999.0337/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0337/production/pdf;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0337/production/ref.

Sakai:1989:CTB

[SA89] Kô Sakai and Akira Aiba. CAL: a theoretical background of constraint logic programming and its applications. Journal of Symbolic Computation, 8(6):589– 604 (or 589–603??), December 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sadek:2016:PCF

[Sad16] Mohammad Sadek. Periodic continued fractions and elliptic curves over quadratic fields. Journal of Symbolic Computation, 76(??): 200–218, September/October 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116000043.

Sadykov:2017:PDH

[Sad17]  Timur Sadykov.  Polyno-

mial dynamics of human blood genotypes frequencies. Journal of Symbolic Computation, 79 (part 2)(??): 342–355, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000213.

Sage:1988:ATQ

[Sag88] Martin L. Sage. An algebraic treatment of quantum vibrations using REDUCE. Journal of Symbolic Computation, 5(3):377–384, June 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See erratum [Sag89].

Sage:1989:EAT

[Sag89] Martin L. Sage. Erratum:

“An algebraic treatment of quantum vibrations using RE-DUCE” [J. Symbolic Comput. 5 (1988), no. 3, 377–384]. Journal of Symbolic Computation, 7(1):101, 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See [Sag88].

Sagraloff:2014:CDM

[Sag14] Michael Sagraloff. On the complexity of the Descartes method when using approximate arithmetic. Journal of Symbolic Computation, 65 (??):79–110, November 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000157.

Sakata:1988:FMS

[Sak88] Shojiro Sakata. Finding a minimal set of linear recurring relations capable of generating a given finite twodimensional array. Journal of Symbolic Computation, 5 (3):321–338 (or 321–337??), June 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Smolka:1989:IHS

[SAK89] Gert Smolka and Hassan Aït-Kaci. Inheritance hierarchies: Semantics and unification. Journal of Symbolic Computation, 7(3–4): 343–370, March/April 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sakamoto:2021:LRL

[Sak21] Ryuichi Sakamoto. Lexicographic and reverse lexicographic quadratic Gröbner bases of cut ideals. Journal of Symbolic Computation, 103 (??):201–212, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711930152X.

Salvy:1994:FCS

[Sal94] Bruno Salvy. Fast computation of some asymptotic functional inverses. Journal of Symbolic Computation, 17(3): 227–236, March 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Salem:2008:ESA

[Sal08] Fatima K. Abu Salem. An efficient sparse adaptation of the polytope method over F_{p} and a record-high binary bivariate factorisation. Journal of Symbolic Computation, 43 (5):311–341, May 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Santas:1995:TSC

[San95] Philip S. Santas. A type system for computer algebra. Journal of Symbolic Computation, 19(1/2/3):79– 110 (or 79–109??), January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Sanchez:1996:MMS

[San96]  Nestor E. Sanchez.  The

method of multiple scales: Asymptotic solutions and normal forms for nonlinear oscillatory problems. Journal of Symbolic Computation, 21(2): 245–252, February 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

SauxPicart:1993:SCS

[Sau93] M. Ph. Saux Picart. Schur– Cohn sub-transforms of a polynomial. Journal of Symbolic Computation, 16(1):1–8 (or 1–7??), July 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sausse:2001:NAP

[Sau01] Alain Sausse. A new approach to primary decomposition. Journal of Symbolic Computation, 31(1–2): 243–257, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.1017;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.1017/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.1017/ref.

Sauer:2018:PMS

[Sau18] Tomas Sauer. Prony’s method in several variables: Symbolic solutions by universal interpolation. Journal of Symbolic Computation, 84(??):95– 112, January/February 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300342.

Savage:1990:SCF

[Sav90] Stuart B. Savage. Symbolic computation of the flow of granular avalanches. Journal of Symbolic Computation, 9 (4):515–530, April 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sit:1999:FGE

[SB99] William Sit and Manuel Bronstein. Foreword of the Guest Editors. Journal of Symbolic Computation, 28(4–5): 435–440, October/November 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0298/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0298/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0298/production/ref.

Sterling:1989:SSE

[SBB^{+}89] Leon Sterling, Alan Bundy, Lawrence Byrd, Richard O’Keefe, and Bernard Silver. Solving symbolic equations with PRESS. Journal of Symbolic Computation, 7 (1):71–84, January 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schwarz:1985:ADP

[Sch85]  Fritz Schwarz.  An algo-

rithm for determining polynomial first integrals of autonomous systems of ordinary differential equations. Journal of Symbolic Computation, 1(2):229–234 (or 229–233??), June 1985. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Schneider:1990:CER

[Sch90a] Gerhard J. A. Schneider. Computing with endomorphism rings of modular representations. Journal of Symbolic Computation, 9(5– 6):607–636, May/June 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Schneider:1990:DCT

[Sch90b] Gerhard J. A. Schneider. Dixon’s character table algorithm revisited. Journal of Symbolic Computation, 9(5–6):601–606 (or 600– 606??), May/June 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Schend:1991:MDS

[Sch91] B. Schend. A methodology for detecting shared variable dependencies in logic programs. Journal of Symbolic Computation, 12(3):275–298, September 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schicho:1992:CPP

[Sch92] Josef Schicho. On the choice of pencils in the parametrization of curves. Journal of Symbolic Computation, 14(6):

557–576, December 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schorn:1993:AAR

[Sch93] Peter Schorn. An axiomatic approach to robust geometric programs. Journal of Symbolic Computation, 16(2): 155–166 (or 155–165??), August 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schorn:1994:ESS

[Sch94] P. Schorn. Evolution of a software system: Interaction, interfaces and applications in the XYZ GeoBench. Journal of Symbolic Computation, 17 (4):311–320, April 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schreiner:1996:PFP

[Sch96] Wolfgang Schreiner. A parafunctional programming interface for a parallel computer algebra package. Journal of Symbolic Computation, 21 (4/5/6):593–614, April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Schicho:1998:RPS

[Sch98a]  Josef Schicho.  Rational

parametrization of surfaces.

Journal of Symbolic Computation, 26(1):1–29, July 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schiemann:1998:CHF

[Sch98b] A. Schiemann. Classification of Hermitian forms with the neighbour method. Journal of Symbolic Computation, 26(4): 487–508, October 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schwartz:1999:PDE

[Sch99] Alan L. Schwartz. Partial differential equations and bivariate orthogonal polynomials. Journal of Symbolic Computation, 28(6):827–845, December 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0342/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0342/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0342/production/ref.

Schicho:2000:PPR

[Sch00]  Josef Schicho.  Proper

parametrization of real tubular surfaces. Journal of Symbolic Computation, 30(5): 583–593, October 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0393;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0393/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0393/ref.

Schulze:2001:AGM

[Sch01]  Mathias Schulze.  Algo-

rithms for the Gauss-Manin connection. Journal of Symbolic Computation, 32 (5):549–564, November 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0482;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0482/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0482/ref.

Schicho:2003:SSP

[Sch03a] Josef Schicho. Simplification of surface parametrizations — a lattice polygon approach. Journal of Symbolic Computation, 36(3–4): 535–554, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schost:2003:CRT

[Sch03b] Éric Schost. Complexity results for triangular sets. Journal of Symbolic Computation, 36(3–4):555–594, September/ October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schulze:2004:NFA

[Sch04] Mathias Schulze. A normal form algorithm for the Brieskorn lattice. Journal of Symbolic Computation, 38 (4):1207–1225, October 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schulze:2005:GBT

[Sch05] Mathias Schulze. Good bases for tame polynomials. Journal of Symbolic Computation, 39(1):103–126, January 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schonhage:2006:PRS

[Sch06] Arnold Schönhage. Polynomial root separation examples. Journal of Symbolic Computation, 41(10):1080– 1090, October 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schauenburg:2007:GBT

[Sch07]  Peter Schauenburg.  A

Gröbner-based treatment of elimination theory for affine varieties. Journal of Symbolic Computation, 42(9):859– 870, September 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schneider:2008:RDF

[Sch08] Carsten Schneider. A refined difference field theory for symbolic summation. Journal of Symbolic Computation, 43 (9):611–644, September 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Scheiblechner:2010:GST

[Sch10]  Peter Scheiblechner.  On

a generalization of Stickelberger’s Theorem. Journal of Symbolic Computation, 45(12):1459–1470, December 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schneider:2016:DRT

[Sch16] Carsten Schneider. A difference ring theory for symbolic summation. Journal of Symbolic Computation, 72(??):82– 127, January/February 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000103.

Scheicher:2017:GBT

[Sch17a] Martin Scheicher. Gröbner bases and their application to the Cauchy problem on finitely generated affine monoids. Journal of Symbolic Computation, 80 (part 2) (??):416–450, May/June 2017.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300335.

Schneider:2017:STI

[Sch17b] Carsten Schneider. Summation theory II: Characterizations of RΠΣ^{∗}-extensions and algorithmic aspects. Journal of Symbolic Computation, 80 (part 3)(??):616–664, May/ June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300748.

Schrempf:2019:FNC

[Sch19]  Konrad Schrempf.  On

the factorization of noncommutative polynomials (in free associative algebras). Journal of Symbolic Computation, 94(??):126–148, September/October 2019. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300919.

Subramani:2005:OQE

[SD05] K. Subramani and D. Desovski. Out of order quantifier elimination for Standard Quantified Linear Programs. Journal of Symbolic Computation, 40(6): 1383–1396, December 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sedoglavic:2002:PAT

[Sed02]  Alexandre Sedoglavic.  A

probabilistic algorithm to test local algebraic observability in polynomial time. Journal of Symbolic Computation, 33(5): 735–755, May 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Seiler:2002:TLR

[Sei02] Werner M. Seiler. Taylor and Lyubeznik resolutions via Gröbner bases. Journal of Symbolic Computation, 34(6): 597–608, December 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Seigal:2020:RSR

[Sei20] Anna Seigal. Ranks and symmetric ranks of cubic surfaces. Journal of Symbolic Computation, 101(??):304–317, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711930104X.

Sekigawa:2009:RFR

[Sek09] Hiroshi Sekigawa. On real factors of real interval polynomials. Journal of Symbolic Computation, 44(7):908–922, July 2009. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Sekigawa:2011:CNP

[Sek11] Hiroshi Sekigawa. Computing the nearest polynomial with a zero in a given domain by using piecewise rational functions. Journal of Symbolic Computation, 46(12): 1318–1335, December 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001246.

Sendra:2002:NPA

[Sen02] J. Rafael Sendra. Normal parametrizations of algebraic plane curves. Journal of Symbolic Computation, 33(6): 863–885, June 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sakkalis:1990:SPA

[SF90] Takis Sakkalis and Rida Farouki. Singular points of algebraic curves. Journal of Symbolic Computation, 9(4): 405–422 (or 405–421??), April 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Snyder:1989:HOU

[SG89] Wayne Snyder and Jean Gallier. Higher order unification revisited: Complete sets of transformations. Journal of Symbolic Computation, 8(1–2):101–140, July/August 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sederberg:1997:IRC

[SGD97] Tom Sederberg, Ron Goldman, and Hang Du. Implicitizing rational curves by the method of moving algebraic curves. Journal of Symbolic Computation, 23(2– 3):153–176 (or 153–175??), February/March 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Sinanan:2017:APF

[SH17a] S. K. Sinanan and D. F. Holt.

Algorithms for polycyclic-by-finite groups. Journal of Symbolic Computation, 79 (part 2)(??):269–284, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000171.

Sun:2017:BNL

[SH17b] Xianbo Sun and Wentao Huang. Bounding the number of limit cycles for a polynomial Liénard system by using regular chains. Journal of Symbolic Computation, 79 (part 2)(??):197–210, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116000134.

Shackell:1990:GEE

[Sha90a] John Shackell. Growth estimates for Exp-log functions. Journal of Symbolic Computation, 10(6):611–632, December 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shallit:1990:WCT

[Sha90b] Jeffrey Shallit. On the worst case of three algorithms for computing the Jacobi symbol. Journal of Symbolic Computation, 10(6):593–610, December 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shaska:2001:CGD

[Sha01]  T. Shaska.  Curves of

genus 2 with (N, N) decomposable Jacobians. Journal of Symbolic Computation, 31(5):603–617, May 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0439;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0439/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0439/ref.

Shafer:2012:SCC

[Sha12] Douglas S. Shafer. Symbolic computation and the cyclicity problem for singularities. Journal of Symbolic Computation, 47(10):1140–1153, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002306.

Shaska:2013:CAG

[Sha13] T. Shaska. Computational algebraic geometry. Journal of Symbolic Computation, 57(??):1–2, October 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000643.

Shen:1992:SCG

[She92] Ba Zhong Shen. Solving a congruence on a graded algebra by a subresultant sequence and its application. Journal of Symbolic Computation, 14(5): 505–522, November 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sherman:1997:TDG

[She97a] G. J. Sherman. Trying to do group theory with undergraduates and computers. Journal of Symbolic Computation, 23(5–6):577–588, May/ June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shevchenko:1997:NDS

[She97b] Ivan I. Shevchenko. Numeric deduction in symbolic computation. application to normalizing transformations. Journal of Symbolic Computation, 24(1):103–112 (or 103–111??), July 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shi:2004:CFD

[Shi04] Hongbo Shi. Computation of the finitistic dimension of monomial algebras. Journal of Symbolic Computation, 37 (4):537–546, April 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shoup:1994:FCI

[Sho94] Victor Shoup. Fast construction of irreducible polynomials over finite fields. Journal of Symbolic Computation, 17 (5):371–391, May 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shoup:1995:NPF

[Sho95] Victor Shoup. A new polynomial factorization algorithm and its implementation. Journal of Symbolic Computation, 20(4):363–398 (or 363– 397??), October 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shparlinski:2014:PVL

[Shp14] Igor E. Shparlinski. Products with variables from lowdimensional affine spaces and shifted power identity testing in finite fields. Journal of Symbolic Computation, 64(??):35–41, August 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001715.

Shtokhamer:1988:LCA

[Sht88] R. Shtokhamer. Lifting canonical algorithms from a ring R to the ring R[x]. Journal of Symbolic Computation, 6 (2–3):169–182 (or 169–181??), October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Sidebottom:1993:ICI

[Sid93] G. Sidebottom. Implementing CLP(IB) using the connection theorem proving method and a clause management system. Journal of Symbolic Computation, 15(1):27–48, January 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Siekmann:1989:UT

[Sie89] Jörg H. Siekmann. Unification theory. Journal of Symbolic Computation, 7(3–4): 207–274, March/April 1989.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sills:2004:RMP

[Sil04] Andrew V. Sills. RRtools — a Maple package for aiding the discovery and proof of finite Rogers–Ramanujan type identities. Journal of Symbolic Computation, 37(4): 415–448, April 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sims:1987:VN

[Sim87] Charles C. Sims. Verifying nilpotence. Journal of Symbolic Computation, 3(3): 231–247, June 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sims:1990:COS

[Sim90a] Charles C. Sims. Computing the order of a solvable permutation group. Journal of Symbolic Computation, 9(5–6):699–706 (or 699– 705??), May/June 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Sims:1990:IBC

[Sim90b] Charles C. Sims. Implementing the Baumslag-Cannonito-Miller polycyclic quotient algorithm. Journal of Symbolic Computation, 9(5–6):

707–724 (or 707–723??), May/ June 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Sims:1991:KBP

[Sim91] Charles C. Sims. The Knuth– Bendix procedure for strings as a substitute for coset enumeration. Journal of Symbolic Computation, 12(4–5): 439–442, October/November 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part 2.

Singer:1990:FSD

[Sin90] Michael F. Singer. Formal solutions of differential equations. Journal of Symbolic Computation, 10(1):59–94, July 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Singer:1991:LSL

[Sin91] Michael F. Singer. Liouvillian solution of linear differential equations with Liouvillian coefficients. Journal of Symbolic Computation, 11(3):251– 274 (or 251–273??), March 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sato:2011:BGB

[SIS^{+}11] Yosuke Sato, Shutaro Inoue, Akira Suzuki, Katsusuke Nabeshima, and Ko Sakai.

Boolean Gröbner bases. Journal of Symbolic Computation, 46(5):622–632, May 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711000180X.

Sit:1992:ASP

[Sit92] William Y. Sit. An algorithm for solving parametric linear systems. Journal of Symbolic Computation, 13(4): 353–394, April 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sit:1997:MNC

[Sit97] W. Y. Sit. Mathematica notebooks for a conventional differential equations course. Journal of Symbolic Computation, 23(5–6):589–623, May/ June 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schrammel:2012:AAA

[SJ12] Peter Schrammel and Bertrand Jeannet. Applying abstract acceleration to (co- )reachability analysis of reactive programs. Journal of Symbolic Computation, 47(12):1512–1532, December 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002549.

San-Juan:2001:ASM

[SJA01] Félix San-Juan and Alberto Abad. Algebraic and symbolic manipulation of Poisson series. Journal of Symbolic Computation, 32 (5):565–572, November 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0396;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0396/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0396/ref.

Saraswat:1996:TDC

[SJG96] V. Saraswat, R. Jagadeesan, and V. Gupta. Timed default concurrent constraint programming. Journal of Symbolic Computation, 22 (5–6):475–520, November/ December 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shi:2013:UBR

[SJG13] Xiaoran Shi, Xiaohong Jia, and Ron Goldman. Using a bihomogeneous resultant to find the singularities of rational space curves. Journal of Symbolic Computation, 53(??):1–25, June 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711200171X.

Szilagyi:2006:LPC

[SJS06] Ibolya Szilágyi, Bert Jüttler, and Josef Schicho. Local parametrization of cubic surfaces. Journal of Symbolic Computation, 41(1):30– 48, January 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sattler-Klein:1991:ECS

[SK91] Andrea Sattler-Klein. Elimination of composite superpositions may cause abortion. Journal of Symbolic Computation, 11(3):205–212 (or 205– 211??), March 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sprenger:2012:ADP

[SK12] Torsten Sprenger and Wolfram Koepf. Algorithmic determination of q-power series for q-holonomic functions. Journal of Symbolic Computation, 47(5):519–535, May 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001969.

Sendra:1992:EPG

[SL92] J. Rafael Sendra and Juan Llovet. An extended polynomial GCD algorithm using Hankel matrices. Journal of Symbolic Computation, 13 (1):25–40 (or 25–39??), January 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Slattery:1986:CCD

[Sla86] Michael C. Slattery. Computing character degrees in p-groups. Journal of Symbolic Computation, 2(1):51– 58, March 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Slattery:2001:CDC

[Sla01] Michael C. Slattery. Computing double cosets in soluble groups. Journal of Symbolic Computation, 31(1–2): 179–192, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1005;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.1005/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1005/ref.

Slattery:2007:GGS

[Sla07] Michael C. Slattery. Generation of groups of square-free order. Journal of Symbolic Computation, 42(6):668–677, June 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schulz:2011:CFT

[SLK11] Christoph Schulz, Michael Löwe, and Harald König. A categorical framework for the transformation of objectoriented systems: Models and data. Journal of Symbolic Computation, 46 (3):316–337, March 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001690.

She:2013:DPL

[SLX^{+}13] Zhikun She, Haoyang Li, Bai Xue, Zhiming Zheng, and Bican Xia. Discovering polynomial Lyapunov functions for continuous dynamical systems. Journal of Symbolic Computation, 58 (??):41–63, November 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113000941.

Sagraloff:2016:CRR

[SM16] Michael Sagraloff and Kurt Mehlhorn. Computing real roots of real polynomials. Journal of Symbolic Computation, 73(??): 46–86, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000292.

Schreck:2018:UJG

[SM18] Pascal Schreck and Pascal Mathis. Using jointly geometry and algebra to determine RC-constructibility. Journal of Symbolic Computation, 90(??):124–148, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711830035X.

Smart:1996:SDF

[Sma96] Nigel Smart. Solving discriminant form equations via unit equations. Journal of Symbolic Computation, 21(3): 367–374, March 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schreiner:2003:DMP

[SMB03] Wolfgang Schreiner, Christian Mittermaier, and Karoly Bosa. Distributed Maple: parallel computer algebra in networked environments. Journal of Symbolic Computation, 35 (3):305–347, March 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Semjonov:1987:SPK

[SME87] A. L. Semjonov, L. S.

Mel′nikov, and V. A. Evstigneev. Solving the problems of kinetics of complex reactions by symbolic algebraic manipulation methods. Journal of Symbolic Computation, 3(3): 303–307, June 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Smith:1993:CSM

[Smi93] Douglas R. Smith. Constructing specification morphisms. Journal of Symbolic Computation, 15(5–6):571–606, May/ June 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Smith:2000:CGE

[Smi00] Gregory G. Smith. Computing global extension modules. Journal of Symbolic Computation, 29(4–5):729–746, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0399;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0399/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0399/ref.

Smith:2002:GBO

[Smi02] David Andrew Smith. Gröbner bases in orders of algebraic number fields. Journal of Symbolic Computation, 33(2):209–220, February 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.

2001.0501;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0501/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0501/ref.

Smith:2005:OFG

[Smi05] Lawrence H. Smith. On ordering free groups. Journal of Symbolic Computation, 40(6): 1285–1290, December 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). See corrigendum [CS09].

Selakovic:2019:NDD

[SMJ19] Milica Selaković, Vesna Marinković, and Predrag Janicić. New dynamics in dynamic geometry: Dragging constructed points. Journal of Symbolic Computation, 97(??):3–15, ???? 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118301287.

Sailer:1991:PRG

[SML91] V. Sailer and G. Morales-Luna. On the polynomial representation of generalized Liouville operators. Journal of Symbolic Computation, 12(3):373–380 (or 373– 379??), September 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Smolka:1998:SIO

[Smo98] G. Smolka. Special issue on order-sorted rewriting foreword by the Guest Editor. Journal of Symbolic Computation, 25(4):395, April 1998.

Smoot:2021:CIR

[Smo21] Nicolas Allen Smoot. On the computation of identities relating partition numbers in arithmetic progressions with eta quotients: an implementation of Radu’s algorithm. Journal of Symbolic Computation, 104(??):276–311, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771712030033X.

Snellman:1998:GBN

[Sne98] Jan Snellman. Gröbner bases and normal forms in a subring of the power series ring on countably many variables. Journal of Symbolic Computation, 25(3):315–328, ???? 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Snyder:1993:FAG

[Sny93] Wayne Snyder. A fast algorithm for generating reduced ground rewriting systems from a set of ground equations. Journal of Symbolic Computation, 15(4): 415–450, April 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Scholten:1989:GMA

[SO89] Hans W. Scholten and Mark H.

Overmars. General methods for adding range restrictions to decomposable searching problems. Journal of Symbolic Computation, 7(1):1–10, January 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Socher:1991:RBR

[Soc91] Rolf Socher. On the relation between resolution based and completion based theorem proving. Journal of Symbolic Computation, 11(1–2): 129–148 (or 129–147??), January/February 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sodan:1996:SAM

[Sod96] A. Sodan. A semi-automatic multiple-strategy approach to mapping tree-structured symbolic processing programs. Journal of Symbolic Computation, 21(4/5/6):615–634, April, May & June 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sofroniou:1994:SDR

[Sof94]  M. Sofroniou.  Symbolic

derivation of Runge–Kutta methods. Journal of Symbolic Computation, 18(3):265– 296, September 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sofroniou:1996:OSL

[Sof96] M. Sofroniou. Order stars and linear stability theory. Journal of Symbolic Computation, 21(1):101–131, January 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Soprunov:2013:TCI

[Sop13] Ivan Soprunov. Toric complete intersection codes. Journal of Symbolic Computation, 50(??):374–385, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001393.

Sorea:2022:MLN

[Sor22] Miruna-Stefana Sorea. Measuring the local non-convexity of real algebraic curves. Journal of Symbolic Computation, 109(??):482–509, March/April 2022. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300729.

Sato:2010:EIS

[SP10] Masahiko Sato and Randy Pollack. External and internal syntax of the λ calculus. Journal of Symbolic Computation, 45(5):598–616, May 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shen:2014:CRR

[SPD14] Li-Yong Shen and Sonia Pérez-Díaz. Characterization of rational ruled surfaces. Journal of Symbolic Computation, 63(??):21–45, May 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001661.

Sitharam:2010:OPS

[SPZ10] Meera Sitharam, Jörg Peters, and Yong Zhou. Optimized parametrization of systems of incidences between rigid bodies. Journal of Symbolic Computation, 45(4):481–498, April 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Steinberg:1986:UMW

[SR86] Stanly Steinberg and Patrick J.

Roache. Using Macsyma to write FORTRAN subroutines. Journal of Symbolic Computation, 2(2):213–216, June 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sangwin:2007:LSC

[SR07] C. J. Sangwin and P. Ramsden. Linear syntax for communicating elementary mathematics. Journal of Symbolic Computation, 42(9):920– 934, September 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shannon:1988:UGB

[SS88] David Shannon and Moss Sweedler. Using Gröbner bases to determine algebra membership, split surjective algebra homomorphisms determine birational equivalence. Journal of Symbolic Computation, 6(2–3): 267–274 (or 267–273??), October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Schmidt-Schauss:1989:UCA

[SS89a] Manfred Schmidt-Schauss. Unification in a combination of arbitrary disjoint equational theories. Journal of Symbolic Computation, 8(1–2):51–100 (or 51– 99??), July/August 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schmidt-Schauss:1989:UPE

[SS89b] Manfred Schmidt-Schauss. Unification in permutative equational theories is undecidable. Journal of Symbolic Computation, 8(4):415– 422 (or 415–421??), October 1989. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Schwartz:1990:TDD

[SS90] Jacob T. Schwartz and Micha Sharir. On the two-dimensional Davenport– Schinzel problem. Journal of Symbolic Computation, 10 (3–4):371–393, September/ October 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sasaki:1992:TNA

[SS92] Tateaki Sasaki and Masayuki Suzuki. Three new algorithms for multivariate polynomial GCD. Journal of Symbolic Computation, 13(4): 395–412 (or 395–411??), April 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shallit:1994:ALS

[SS94] Jeffrey Shallit and Jonathan Sorenson. Analysis of a leftshift binary GCD algorithm. Journal of Symbolic Computation, 17(6):473–486, June 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shackell:1995:AFA

[SS95] John Shackell and Bruno Salvy. Asymptotic forms and algebraic differential equations. Journal of Symbolic Computation, 20(2):169–178 (or 169–177??), August 1995.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Salinier:1996:ESF

[SS96a] Bruno Salinier and Robert Strandh. Efficient simulation of forward-branching systems with constructor systems. Journal of Symbolic Computation, 22(4):381–400 (or 381–399??), October 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schmidt-Schauss:1996:DUT

[SS96b] Manfred Schmidt-Schauß. Decidability of unification in the theory of one-sided distributivity and a multiplicative unit. Journal of Symbolic Computation, 22(3):315– 344, September 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Salvy:1998:SAF

[SS98a] Bruno Salvy and John Shackell. Symbolic asymptotics: Functions of two variables, implicit functions. Journal of Symbolic Computation, 25(3):329–350 (or 329– 349??), March 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shirayanagi:1998:RAA

[SS98b] K. Shirayanagi and M. Sweedler. Remarks on automatic algorithm stabilization. Journal of Symbolic Computation, 26(6): 761–766, December 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Salvy:1999:SAM

[SS99] Bruno Salvy and John Shackell. Symbolic asymptotics: Multiseries of inverse functions. Journal of Symbolic Computation, 27(6):543–563, June 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0281/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0281/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0281/production/ref.

Sofronie-Stokkermans:2003:RBD

[SS03a] Viorica Sofronie-Stokkermans.

Resolution-based decision procedures for the universal theory of some classes of distributive lattices with operators. Journal of Symbolic Computation, 36(6):891–924, December 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Suzuki:2003:AAC

[SS03b] Akira Suzuki and Yosuke Sato. An alternative approach to comprehensive Gröbner bases. Journal of Symbolic Computation, 36(3–4): 649–667, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schicho:2005:NSS

[SS05] Josef Schicho and Ibolya Szilágyi. Numerical stability of surface implicitization. Journal of Symbolic Computation, 40(6): 1291–1301, December 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Slavkovic:2006:SCF

[SS06] Aleksandra B. Slavkovic and Seth Sullivant. The space of compatible full conditionals is a unimodular toric variety. Journal of Symbolic Computation, 41(2):196–209, February 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

SanSegundo:2009:PDF

[SS09] F. San Segundo and J. R. Sendra. Partial degree formulae for plane offset curves. Journal of Symbolic Computation, 44(6):635–654, June 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sendra:2011:RPA

[SS11] J. Rafael Sendra and David Sevilla. Radical parametrizations of algebraic curves by adjoint curves. Journal of Symbolic Computation, 46(9): 1030–1038, September 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000769.

Scott:2016:CIK

[SS16] Leonard L. Scott and Timothy Sprowl. Computing individual Kazhdan–Lusztig basis elements. Journal of Symbolic Computation, 73(??): 244–249, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000486.

Sitharam:2018:FSI

[SS18] Meera Sitharam and Audrey St. John. Foreword to special issue. Journal of Symbolic Computation, 88(??):1– 4, September/October 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300117.

Szemberg:2024:SPF

[SS24] Tomasz Szemberg and Justyna Szpond. Sextactic points on the Fermat cubic curve and arrangements of conics. Journal of Symbolic Computation, 120(??):Article 102228, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000354.

Schmidt-Schauss:2002:SCE

[SSS02] Manfred Schmidt-Schauß and Klaus U. Schulz. Solvability of context equations with two context variables is decidable. Journal of Symbolic Computation, 33(1):77–122, January 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0438;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0438/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0438/ref.

Schmidt-Schauss:2005:DBH

[SSS05] Manfred Schmidt-Schauß and Klaus U. Schulz. Decidability of bounded higher-order unification. Journal of Symbolic Computation, 40(2):905– 954, August 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Salvy:2011:OPF

[SSS^{+}11] Bruno Salvy, Bob Sedgewick, Michele Soria, Wojciech Szpankowski, and Brigitte Vallee. Obituary. Philippe Flajolet. Journal of Symbolic Computation, 46(9): 1085–1086, September 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000873.

Saha:2023:BSP

[SSS23] Joydip Saha, Indranath Sengupta, and Pranjal Srivastava. Betti sequence of the projective closure of affine monomial curves. Journal of Symbolic Computation, 119(??):101– 111, November/December 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000172.

Schmidt-Schauss:2018:NUA

[SSSK18]  Manfred  Schmidt-Schauß,

David Sabel, and Yunus D. K. Kutz. Nominal unification with atom-variables. Journal of Symbolic Computation, 90(??):42–64, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300324.

Saha:2018:IF

[SST18] Joydip Saha, Indranath Sengupta, and Gaurab Tripathi. Ideals of the form I_{1}(XY). Journal of Symbolic Computation, 91(??):17–29, ???? 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300762.

Shahidi:2023:DKV

[SSV23] Zahra Shahidi, Luca Sodomaco, and Emanuele Ventura. Degrees of Kalman varieties of tensors. Journal of Symbolic Computation, 114 (??):74–98, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000360.

Sato:1989:FOC

[ST89a] Taisuke Sato and Hisao Tamaki. First order compiler: a deterministic logic program synthesis algorithm. Journal of Symbolic Computation, 8(6):605–628 (or 605– 627??), December 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Suppes:1989:ICT

[ST89b] Patrick Suppes and Shuzo Takahashi. An interactive calculus theorem-prover for continuity properties. Journal of Symbolic Computation, 7 (6):573–590, June 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Strzebonski:2019:URR

[ST19] Adam Strzebonski and Elias Tsigaridas. Univariate real root isolation in an extension field and applications. Journal of Symbolic Computation, 92(??):31–51, May/ June 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117301256.

Shibuta:2020:ACH

[ST20] Takafumi Shibuta and Shinichi Tajima. An algorithm for computing the Hilbert– Samuel multiplicities and reductions of zero-dimensional ideals of Cohen–Macaulay local rings. Journal of Symbolic Computation, 96(??): 108–121, January/February 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300252.

Seynnaeve:2024:UEM

[ST24a] Tim Seynnaeve and Nafie Tairi. Universal equations for maximal isotropic Grassmannians. Journal of Symbolic Computation, 121(??): Article 102260, March/April 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000743.

Sodomaco:2024:SST

[ST24b] Luca Sodomaco and Ettore Teixeira Turatti. The span of singular tuples of a tensor beyond the boundary format. Journal of Symbolic Computation, 120(??):Article 102230, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000378.

Statman:1989:WPS

[Sta89] Rick Statman. The word problem for Smullyan’s lark combinator is decidable. Journal of Symbolic Computation, 7(2):103–112, February 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sawada:1994:PCG

[STA94] Hiroyuki Sawada, Satoshi Terasaki, and Akira Aiba. Parallel computation of Gröbner bases on distributed memory machines. Journal of Symbolic Computation, 18(3):207– 222, September 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Stagliano:2016:ESQ

[Sta16] Giovanni Staglianò. Examples of special quadratic birational transformations into complete intersections of quadrics. Journal of Symbolic Computation, 74(??):635–649, May/ June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115001029.

Stainsby:2018:TBI

[Sta18] Hayden D. Stainsby. Triangular bases of integral closures. Journal of Symbolic Computation, 87(??): 140–175, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300780.

Staffolani:2023:SA

[Sta23] Reynaldo Staffolani. Schur apolarity. Journal of Symbolic Computation, 114(??):37– 73, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000372.

Soeken:2016:AFS

[STDD16] Mathias Soeken, Laura Tague, Gerhard W. Dueck, and Rolf Drechsler. Ancilla-free synthesis of large reversible functions using binary decision diagrams. Journal of Symbolic Computation, 73 (??):1–26, March/April 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000188.

Stembridge:1995:MPS

[Ste95] John R. Stembridge. A Maple package for symmetric functions. Journal of Symbolic Computation, 20(5–6):755– 768, November/December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Steel:1997:NAC

[Ste97] Allan Steel. A new algorithm for the computation of canonical forms of matrices over fields. Journal of Symbolic Computation, 24(3–4): 409–432, September/October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Steel:2005:CIP

[Ste05]  Allan Steel.  Conquering

inseparability: Primary decomposition and multivariate factorization over algebraic function fields of positive characteristic. Journal of Symbolic Computation, 40(3): 1053–1075, September 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Steel:2010:CAC

[Ste10] Allan K. Steel. Computing with algebraically closed fields. Journal of Symbolic Computation, 45(3):342– 372, March 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Steidel:2013:GBS

[Ste13] Stefan Steidel. Gröbner bases of symmetric ideals. Journal of Symbolic Computation, 54(??):72–86, July 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711300014X.

Stifter:1987:GRR

[Sti87]  Sabine Stifter.  A gener-

alization of reduction rings. Journal of Symbolic Computation, 4(3):351–364, December 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Stillman:2003:CAG

[Sti03] Michael Stillman. Computing in algebraic geometry and commutative algebra using Macaulay 2. Journal of Symbolic Computation, 36(3–4): 595–611, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Stokkermans:1999:CCP

[Sto99] K. Stokkermans. A categorical critical-pair completion algorithm. Journal of Symbolic Computation, 27(5):435–478, May 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Storjohann:2003:HOL

[Sto03] Arne Storjohann. High-order lifting and integrality certification. Journal of Symbolic Computation, 36(3–4): 613–648, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Stoutemyer:2011:TCG

[Sto11] David R. Stoutemyer. Ten commandments for good default expression simplification. Journal of Symbolic Computation, 46(7):859–887, July 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001471.

Stout:2017:AIZ

[Sto17] Andrew R. Stout. On the auto Igusa-zeta function of an algebraic curve. Journal of Symbolic Computation, 79 (part 1)(??):156–185, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300906.

Stoimenow:2020:HCG

[Sto20] A. Stoimenow. Hyperbolicity of the canonical genus two knots. Journal of Symbolic Computation, 101(??):242– 269, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300951.

Strzebonski:1997:CFC

[Str97] Adam Wojciech Strzeboński.

Computing in the field of complex algebraic numbers. Journal of Symbolic Computation, 24(6):647–656, December 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Strzebonski:2000:SSS

[Str00]  Adam Strzeboński.  Solv-

ing systems of strict polynomial inequalities. Journal of Symbolic Computation, 29(3):471–480, March 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0327;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0327/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0327/ref.

Stratulat:2001:GFB

[Str01]  Sorin Stratulat.  A gen-

eral framework to build contextual cover set induction provers. Journal of Symbolic Computation, 32 (4):403–445, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0469;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0469/

                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0469/ref.

Strzebonski:2006:CAD

[Str06] Adam W. Strzeboński. Cylindrical Algebraic Decomposition using validated numerics. Journal of Symbolic Computation, 41(9): 1021–1038, September 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Strzebonski:2011:CDS

[Str11] Adam Strzeboński. Cylindrical decomposition for systems transcendental in the first variable. Journal of Symbolic Computation, 46(11): 1284–1290, November 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001210.

Strzebonski:2012:RRI

[Str12]  Adam Strzeboński.  Real

root isolation for exp–log– arctan functions. Journal of Symbolic Computation, 47(3):282–314, March 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111001908.

Strzebonski:2016:CAD

[Str16] Adam Strzeboński. Cylindrical algebraic decomposition using local projections. Journal of Symbolic Computation, 76(??):36– 64, September/October 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115001170.

Strobel:2019:NSA

[Str19]  Michael Strobel.  Non-

standard analysis in dynamic geometry. Journal of Symbolic Computation, 97(??):69–108, ???? 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118301329.

Sturmfels:1991:CAG

[Stu91] Bernd Sturmfels. Computational algebraic geometry of projective configurations. Journal of Symbolic Computation, 11(5–6):595–618, May/ June 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Invarianttheoretic algorithms in geometry (Minneapolis, MN, 1987).

Sturm:2000:LPV

[Stu00] Thomas Sturm. Linear problems in valued fields. Journal of Symbolic Computation, 30(2):207–219, August 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.

                                        1999.0303;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0303/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0303/ref.

Sturmfels:2017:HFP

[Stu17] Bernd Sturmfels. The Hurwitz form of a projective variety. Journal of Symbolic Computation, 79 (part 1)(??): 186–196, March/April 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116300918.

Sturmfels:2024:TGE

[STVvR24] Bernd Sturmfels, Simon Telen, François-Xavier Vialard, and Max von Renesse. Toric geometry of entropic regularization. Journal of Symbolic Computation, 120(??):Article 102221, January/February 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000287.

Sitharam:2018:CRI

[STW18] Meera Sitharam, Mohamad Tarifi, and Menghan Wang. Combinatorial rigidity of incidence systems and application to dictionary learning. Journal of Symbolic Computation, 88(??):21– 46, September/October 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717118300130.

Singer:1993:GGS

[SU93a] Michael F. Singer and Felix Ulmer. Galois groups of second and third order linear differential equations. Journal of Symbolic Computation, 16(1):9–36, July 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Singer:1993:LAS

[SU93b] Michael F. Singer and Felix Ulmer. Liouvillian and algebraic solutions of second and third order linear differential equations. Journal of Symbolic Computation, 16(1):37– 74 (or 37–73??), July 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sutherland:2012:ECM

[Sut12]  Nicole Sutherland.  Effi-

cient computation of maximal orders in radical (including Kummer) extensions. Journal of Symbolic Computation, 47(5):552–567, May 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002562.

Sutherland:2013:ECM

[Sut13] Nicole Sutherland. Efficient computation of maximal orders in Artin–Schreier extensions. Journal of Symbolic Computation, 53(??):26–39, June 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001873.

Sutherland:2015:CGG

[Sut15] Nicole Sutherland. Computing Galois groups of polynomials (especially over function fields of prime characteristic). Journal of Symbolic Computation, 71(??): 73–97, November/December 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001114.

Sutherland:2016:ECM

[Sut16]  Nicole Sutherland.  Effi-

cient computation of maximal orders in Artin–Schreier– Witt extensions. Journal of Symbolic Computation, 77(??):189–216, November/December 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000122.

Sarlet:1992:RPS

[SV92] W. Sarlet and J. Vanden Bonne. REDUCE-procedures for the study of adjoint symmetries of second-order differential equations. Journal of Symbolic Computation, 13 (6):683–693, June 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Shkaravska:2014:UPS

[SvE14] O. Shkaravska and M. van Eekelen. Univariate polynomial solutions of algebraic difference equations. Journal of Symbolic Computation, 60(??):15–28, January 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001296.

Shkaravska:2021:PSA

[SvE21] Olha Shkaravska and Marko van Eekelen. Polynomial solutions of algebraic difference equations and homogeneous symmetric polynomials. Journal of Symbolic Computation, 103(??):22–45, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711930135X.

Szutkoski:2019:CCA

[SvH19] Jonas Szutkoski and Mark van Hoeij. The complexity of computing all subfields of an algebraic number field. Journal of Symbolic Computation, 93(??):161–182, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717118300658.

Sendra:1991:SPC

[SW91a] J. Rafael Sendra and Franz Winkler. Symbolic parametrization of curves. Journal of Symbolic Computation, 12 (6):607–632 (or 607–631??), December 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sturmfels:1991:SFP

[SW91b] Bernd Sturmfels and Walter Whiteley. On the synthetic factorization of projectively invariant polynomials. Journal of Symbolic Computation, 11(5–6):439–454 (or 439– 453??), May/June 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Invariant-theoretic algorithms in geometry (Minneapolis, MN, 1987).

Strehl:1995:FSS

[SW95] Volker Strehl and Herbert S. Wilf. Five surprisingly simple complexities. Journal of Symbolic Computation, 20(5– 6):725–730 (or 725–729??), November/December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Sendra:1997:PAC

[SW97a] J. Rafael Sendra and Franz Winkler. Parametrization of algebraic curves over optimal field extensions. Journal of Symbolic Computation, 23 (2–3):191–208 (or 191–207??), February/March 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

Suleiman:1997:MCC

[SW97b] Ibrahim A. I. Suleiman and Robert A. Wilson. The 2modular characters of Conway’s third group Co_{3}. Journal of Symbolic Computation, 24(3–4):493–506, September/ October 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational algebra and number theory (London, 1993).

Shank:2002:CMI

[SW02] R. James Shank and David L.

Wehlau. Computing modular invariants of p-groups. Journal of Symbolic Computation, 34(5):307–327, November 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Schonberg:2011:VCW

[SWF11] Christian Schönberg, Franz Weitl, and Burkhard Freitag. Verifying the consistency of Web-based technical documentations. Journal of Symbolic Computation, 46(2):183–206, February 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001379.

Shimoyama:1996:LPD

[SY96] Takeshi Shimoyama and Kazuhiro Yokoyama. Localization and primary decomposition of polynomial ideals. Journal of Symbolic Computation, 22(3):247–278 (or 247– 277??), September 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Szanto:2008:SDS

[Sza08] Agnes Szanto. Solving overdetermined systems by the subresultant method (with an appendix by Marc Chardin). Journal of Symbolic Computation, 43(1):46–74, January 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Szilagyi:2017:CJK

[Szi17] Zsolt Szilágyi. Computation of Jeffrey–Kirwan residues using Gröbner bases. Journal of Symbolic Computation, 79 (part 2)(??):327– 341, March/April 2017. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000201.

Szpond:2022:UHM

[Szp22] Justyna Szpond. Unexpected hypersurfaces with multiple fat points. Journal of Symbolic Computation, 109(??): 510–519, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300730.

Tiden:1987:UPO

[TA87] Erik Tidén and Stefan Arnborg. Unification problems with one-sided distributivity. Journal of Symbolic Computation, 3(1–2):183–202, February/April 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Rewriting techniques and applications (Dijon, 1985).

Tabera:2011:OAR

[Tab11] Luis Felipe Tabera. Optimal affine reparametrization of rational curves. Journal of Symbolic Computation, 46(8):967–976, August 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111000502.

Tabera:2013:CHM

[Tab13] Luis Felipe Tabera. Computing hypercircles by moving hyperplanes. Journal of Symbolic Computation, 50 (??):450–464, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717112001551.

Takahashi:1989:PRC

[Tak89] Masako Takahashi. Parallel reduction in λ-calculus. Journal of Symbolic Computation, 7(2):113–124 (or 113– 123??), February 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Takayama:1991:ERF

[Tak91] Yukihide Takayama. Extraction of redundancy-free programs from constructive natural deduction proofs. Journal of Symbolic Computation, 12(1):29–70 (or 29–69??), July 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Takayama:1992:AZR

[Tak92] Nobuki Takayama. An approach to the zero recognition problem by Buchberger algorithm. Journal of Symbolic Computation, 14(2–3): 265–282, August/September 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Takayama:1993:CCP

[Tak93] Yukihide Takayama. QPC_{2}:

a constructive calculus with parameterized specifications. Journal of Symbolic Computation, 15(5–6):641–672, May/ June 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Takayama:1995:AFR

[Tak95] Nobuki Takayama. An algorithm for finding recurrence relations of binomial sums and its complexity. Journal of Symbolic Computation, 20(5–6):637–652 (or 637– 651??), November/December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Taylor:2002:IGB

[Tay02]  Amelia Taylor.  The in-

verse Gröbner basis problem in codimension two. Journal of Symbolic Computation, 33(2):221–238, February 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0511;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0511/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0511/ref.

Torrente:2017:PRZ

[TBS17] Maria-Laura Torrente, Mauro C. Beltrametti, and Andrew J. Sommese. Perturbation results on the zero-locus of a polynomial. Journal of Symbolic Computation, 80 (part 2) (??):307–328, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711630013X.

Tonelli-Cueto:2023:CNC

[TCT23] Josué Tonelli-Cueto and Elias Tsigaridas. Condition numbers for the cube. I: Univariate polynomials and hypersurfaces. Journal of Symbolic Computation, 115(??): 142–173, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000797.

Tefera:2002:MMP

[Tef02] Akalu Tefera. MultInt, a MAPLE package for multiple integration by the WZ method. Journal of Symbolic Computation, 34(5):329– 353, November 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Teske:1999:PHM

[Tes99] Edlyn Teske. The Pohlig-Hellman method generalized for group structure computation. Journal of Symbolic Computation, 27(6):521–534, June 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0279/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0279/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0279/production/ref.

Thatte:1993:FAT

[Tha93] Satish R. Thatte. Finite acyclic theories are unitary. Journal of Symbolic Computation, 15(2):235–249, February 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Theobald:2006:FPC

[The06] Thorsten Theobald. On the frontiers of polynomial computations in tropical geometry. Journal of Symbolic Computation, 41(12):1360– 1375, December 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Thome:2002:SCV

[Tho02]  Emmanuel Thomé.  Sub-

quadratic computation of vector generating polynomials and improvement of the block Wiedemann algorithm. Journal of Symbolic Computation, 33(5):757–775, May 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tummarakota:1996:SFE

[TL96] Srinivas Tummarakota and Junghsen Lieh. Symbolic finite element modeling of structural systems. Journal of Symbolic Computation, 22 (1):105–119, July 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tombal:1985:MCD

[TM85] Ph. Tombal and A. Moussiaux. MACSYMA computation of the Dirac–Bergmann algorithm for Hamiltonian systems with constraints. Journal of Symbolic Computation, 1(4):419–421, December 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tombal:1989:APG

[TM89] Ph. Tombal and A. Moussiaux. Algebraic programming of geometrical calculus and Clifford algebra. Journal of Symbolic Computation, 7(1):85–92 (or 85–91??), January 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tajima:2009:AIA

[TN09] Shinichi Tajima and Yayoi Nakamura. Annihilating ideals for an algebraic local cohomology class. Journal of Symbolic Computation, 44(5): 435–448, May 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tohaneanu:2010:BPM

[Toh10]  Ştefan O. Tohaňeanu. On the

De Boer–Pellikaan method for computing minimum distance. Journal of Symbolic Computation, 45(10):965–974, October 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Topuzoglu:2014:CRP

[Top14] Alev Topuzoglu. The Carlitz rank of permutations of finite fields: a survey. Journal of Symbolic Computation, 64(??):53–66, August 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001739.

Torgersen:1993:PSR

[Tor93]  T. Torgersen.  Parallel

scheduling of recursively defined arrays: Revisited. Journal of Symbolic Computation, 16(2):189–226, August 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Traugott:1989:DSS

[Tra89] J. Traugott. Deductive synthesis of sorting programs.

Journal of Symbolic Computation, 7(6):533–572, June 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Traverso:1996:HFB

[Tra96] Carlo Traverso. Hilbert functions and the Buchberger algorithm. Journal of Symbolic Computation, 22(4):355– 376, October 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tran:1998:SNM

[Tra98] Q.-N. Tran. A symbolicnumerical method for finding a real solution of an arbitrary system of nonlinear algebraic equations. Journal of Symbolic Computation, 26(6):739– 760, December 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tran:2000:FAG

[Tra00] Quoc-Nam Tran. A fast algorithm for gröbner basis conversion and its applications. Journal of Symbolic Computation, 30(4):451–467, October 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0416;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0416/
pdf;   http://www.idealibrary.

com/links/doi/10.1006/jsco.
1999.0416/ref.

Traves:2006:DOO

[Tra06]  William N. Traves.  Dif-

ferential operators on orbifolds. Journal of Symbolic Computation, 41(12):1295– 1308, December 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tran:2007:NCT

[Tra07a] Quoc-Nam Tran. A new class of term orders for elimination. Journal of Symbolic Computation, 42(5):533–548, May 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tran:2007:SIA

[Tra07b] Quoc-Nam Tran. Special issue on Applications of Computer Algebra. Journal of Symbolic Computation, 42(5):495–496, May 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Treinen:1992:NMU

[Tre92] Ralf Treinen. A new method for undecidability proofs of first order theories. Journal of Symbolic Computation, 14(5):437–458 (or 437– 457??), November 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Trindle:1986:AMR

[Tri86] C. Trindle. Application of the MuMath(R) symbol manipulation system to chemically significant permutation groups. Journal of Symbolic Computation, 2(2):207–212, June 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tran:2010:CCT

[TRRK10] Duc-Khanh Tran, Christophe Ringeissen, Silvio Ranise, and Hélène Kirchner. Combination of convex theories: Modularity, deduction completeness, and explanation. Journal of Symbolic Computation, 45(2):261–286, February 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Temiz:2024:CCT

[TS24] Fatih Temiz and Irfan Siap. On cyclic codes over Z_{q}[u]/⟨u^{2}⟩ and their enumeration. Journal of Symbolic Computation, 120(??):Article 102227, January/February 2024. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000342.

Tsai:2000:WCL

[Tsa00] Harrison Tsai. Weyl closure of a linear differential operator. Journal of Symbolic Computation, 29(4–5):747–775, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0400;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0400/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0400/ref.

Tsai:2016:ENT

[Tsa16]  Ya-Lun Tsai.  Estimat-

ing the number of tetrahedra determined by volume, circumradius and four face areas using Groebner basis. Journal of Symbolic Computation, 77(??):162–174, November/December 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116000109.

Tsai:2023:ECY

[Tsa23] Ya-Lun Tsai. Exploring circulations yielding no vortex equilibria. Journal of Symbolic Computation, 114(??): 231–245, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000396.

Tsuji:2009:IEG

[Tsu09] Kuniaki Tsuji. An improved EZ-GCD algorithm for multivariate polynomials. Journal of Symbolic Computation, 44(1):99–110, January 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tengely:2018:PP

[TU18] Sz. Tengely and M. Ulas. On a problem of Pethö. Journal of Symbolic Computation, 89(??):216–226, November/December 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117301207.

Tung:2002:ASP

[Tun02] Shih Ping Tung. Approximate solutions of polynomial equations. Journal of Symbolic Computation, 33(2):239–254, February 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0505;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0505/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0505/ref.

Tung:2009:ANS

[Tun09]  Shih Ping Tung.  Algo-

rithms for near solutions to polynomial equations. Journal of Symbolic Computation, 44(10):1410–1424, October 2009. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Torstensson:2005:URS

[TUÖ05] Anna Torstensson, Victor Ufnarovski, and Hans Öfverbeck. Using resultants for SAGBI basis verification in the univariate polynomial ring. Journal of Symbolic Computation, 40(3):1087–1105, September 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Torrente:2018:CBT

[TV18] Maria-Laura Torrente and Matteo Varbaro. Computing the Betti table of a monomial ideal: a reduction algorithm. Journal of Symbolic Computation, 87(??): 87–98, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117300615.

tenEikelder:1991:NFC

[tW91] H. M. M. ten Eikelder and J. C. F. Wilmont. Normal forms for a class of formulas. Journal of Symbolic Computation, 12(2):143–160, August 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tsai:2001:CHB

[TW01] Harrison Tsai and Uli Walther.

Computing homomorphisms between holonomic D-modules.

Journal of Symbolic Computation, 32(6):597–617, December 1, 2001. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0485;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0485/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0485/ref.

Thanatipanonda:2021:MCE

[TZ21] Thotsaporn “Aek” Thanatipanonda and Doron Zeilberger. A multi-computational exploration of some games of pure chance. Journal of Symbolic Computation, 104 (??):38–68, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300183.

U:2005:CD

[U.05] Orlando Villamayor U. On constructive desingularization. Journal of Symbolic Computation, 39(3–4):465– 491, March/April 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Uteshev:1998:SMP

[UC98] Alexei Yu. Uteshev and Timofei M. Cherkasov. The search for the maximum of a polynomial. Journal of Symbolic Computation, 25(5):587–618, May 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ucha:2004:CBS

[UCJ04] J. M. Ucha and F. J. Castro-Jiménez. On the computation of Bernstein–Sato ideals. Journal of Symbolic Computation, 37(5):629–639, May 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ulmer:1994:ILD

[Ulm94] Felix Ulmer. Irreducible linear differential equations of prime order. Journal of Symbolic Computation, 18(4):385– 401, October 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ulmer:2003:LST

[Ulm03] Felix Ulmer. Liouvillian solutions of third order differential equations. Journal of Symbolic Computation, 36(6): 855–889, December 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Unger:2006:CCT

[Ung06] W. R. Unger. Computing the character table of a finite group. Journal of Symbolic Computation, 41(8):847– 862, August 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Unger:2019:ACS

[Ung19] W. R. Unger. An algorithm for computing Schur indices of characters. Journal of Symbolic Computation, 93(??):148–160, July/August 2019. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711830066X.

Uray:2024:ANF

[Ura24] M. J. Uray. Algebraic number fields and the LLL algorithm. Journal of Symbolic Computation, 121(??): Article 102261, March/April 2024. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000755.

Ulmer:1996:NKA

[UW96] Felix Ulmer and Jacques-Arthur Weil. Note on Kovacic’s algorithm. Journal of Symbolic Computation, 22(2): 179–200, August 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Uteshev:2015:MPQ

[UY15] Alexei Yu. Uteshev and Marina V. Yashina. Metric problems for quadrics in multidimensional space. Journal of Symbolic Computation, 68 (part 1)(??):287–315, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000893.

Umeno:1989:SCA

[UYSA89] Takaji Umeno, Syuichi Yamashita, Osami Saito, and Kenichi Abe. Symbolic computation application for the design of linear multivariable control systems. Journal of Symbolic Computation, 8(6): 581–588, December 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vaccon:2017:MFA

[Vac17] Tristan Vaccon. Matrix-F5 algorithms over finite-precision complete discrete valuation fields. Journal of Symbolic Computation, 80 (part 2) (??):329–350, May/June 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116300293.

Vaccon:2018:MFA

[Vac18]  Tristan Vaccon.  Matrix–

F5 algorithms and tropical Gröbner bases computation. Journal of Symbolic Computation, 89(??):227–254, November/December 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117301219.

Valibouze:2011:GBA

[Val11] Annick Valibouze. Gröbner basis of the alternating galoisian ideal. Journal of Symbolic Computation, 46(4): 396–405, April 2011. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001896.

vandenEssen:1993:ACI

[van93] Arno van den Essen. An algorithm to compute the invariant ring of a G_{a}-action on an affine variety. Journal of Symbolic Computation, 16(6):551– 556 (or 551–555??), December 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanHoeij:1994:ACI

[van94] Mark van Hoeij. An algorithm for computing an integral basis in an algebraic function field. Journal of Symbolic Computation, 18(4): 353–364 (or 353–363??), October 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanHoeij:1997:FDO

[van97a] Mark van Hoeij. Factorization of differential operators with rational functions coefficients. Journal of Symbolic Computation, 24(5):537–562 (or 537– 561??), November 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanHoeij:1997:FSF

[van97b] Mark van Hoeij. Formal solutions and factorization of differential operators with power series coefficients. Journal of Symbolic Computation, 24 (1):1–30, July 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanHoeij:1997:RPA

[van97c] Mark van Hoeij. Rational parametrizations of algebraic curves using a canonical divisor. Journal of Symbolic Computation, 23(2–3):209–228 (or 209–227??), February/March 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parametric algebraic curves and applications (Albuquerque, NM, 1995).

VanDerKallen:2000:CHM

[Van00] Wilberd Van Der Kallen.

Complexity of the Havas, Majewski, Matthews LLL Hermite normal form algorithm. Journal of Symbolic Computation, 30(3): 329–337, September 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.

2000.0374;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0374/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0374/ref.

VanDerHoeven:2002:FLM

[Van02] Joris Van Der Hoeven. FFTlike multiplication of linear differential operators. Journal of Symbolic Computation, 33(1):123–127, January 1, 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0496;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0496/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0496/ref.

Vasconcelos:2000:DEC

[Vas00] Wolmer V. Vasconcelos. Divisorial extensions and the computation of integral closures. Journal of Symbolic Computation, 30(5):595–604, October 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0405;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0405/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0405/ref.

Vatter:2006:FLG

[Vat06] Vincent Vatter. Finitely labeled generating trees and restricted permutations. Journal of Symbolic Computation, 41(5):559–572, May 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vatter:2012:FRI

[Vat12] Vincent Vatter. Finding regular insertion encodings for permutation classes. Journal of Symbolic Computation, 47(3):259–265, March 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001787.

Velev:2003:EUB

[VB03] Miroslav N. Velev and Randal E. Bryant. Effective use of Boolean satisfiability procedures in the formal verification of superscalar and VLIW microprocessors. Journal of Symbolic Computation, 35(2): 73–106, February 2003. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderHoeven:2001:FEH

[vdH01] Joris van der Hoeven. Fast evaluation of holonomic functions near and in regular singularities. Journal of Symbolic Computation, 31(6):717–743, June 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0474;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2000.0474/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2000.0474/ref.

vanderHoeven:2002:RDT

[vdH02] Joris van der Hoeven. Relax, but don’t be too lazy. Journal of Symbolic Computation, 34(6):479–542, December 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderHoeven:2005:EAF

[vdH05] Joris van der Hoeven. Effective analytic functions. Journal of Symbolic Computation, 39(3–4):433–449, March/April 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderHoeven:2006:CWC

[vdH06] Joris van der Hoeven. Counterexamples to witness conjectures. Journal of Symbolic Computation, 41(9):959– 963, September 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderHoeven:2007:ANS

[vdH07a] Joris van der Hoeven. Around the numeric–symbolic computation of differential Galois groups. Journal of Symbolic Computation, 42(1–2): 236–264, January/February 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderHoeven:2007:EAS

[vdH07b] Joris van der Hoeven. Efficient accelero-summation of holonomic functions. Journal of Symbolic Computation, 42 (4):389–428, April 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderHoeven:2007:GPS

[vdH07c] Joris van der Hoeven. Generalized power series solutions to linear partial differential equations. Journal of Symbolic Computation, 42(8):771– 791, August 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderHoeven:2007:NAR

[vdH07d] Joris van der Hoeven. New algorithms for relaxed multiplication. Journal of Symbolic Computation, 42(8):792– 802, August 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderHoeven:2009:AE

[vdH09] Joris van der Hoeven. On asymptotic extrapolation.

Journal of Symbolic Computation, 44(8):1000–1016, August 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderHoeven:2010:NMF

[vdH10] Joris van der Hoeven. Newton’s method and FFT trading. Journal of Symbolic Computation, 45(8):857–878, August 2010. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderHoeven:2011:MET

[vdH11] Joris van der Hoeven. Metaexpansion of transseries. Journal of Symbolic Computation, 46(4):339–359, April 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001495.

vanderHoeven:2013:GSD

[vdH13] Joris van der Hoeven. Guessing singular dependencies. Journal of Symbolic Computation, 59(??):54–80, December 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000977.

vanderHoeven:2015:TSM

[vdH15] Joris van der Hoeven. Towards semantic mathematical editing. Journal of Symbolic Computation, 71(??): 1–46, November/December 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114001084.

vanderHoeven:2013:BCS

[vdHL13] Joris van der Hoeven and Grégoire Lecerf. On the bitcomplexity of sparse polynomial and series multiplication. Journal of Symbolic Computation, 50(??):227–254, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001228.

vanderHoeven:2006:CBZ

[vdHS06] Joris van der Hoeven and John Shackell. Complexity bounds for zero-test algorithms. Journal of Symbolic Computation, 41(9): 1004–1020, September 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderPut:1999:GTD

[vdP99] Marius van der Put. Galois theory of differential equations, algebraic groups and Lie algebras. Journal of Symbolic Computation, 28(4–5): 441–472, October/November 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0310/production;

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0310/production/pdf;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0310/production/ref.

vanderPut:2005:GTA

[vdP05] Marius van der Put. Galois theory and algorithms for linear differential equations. Journal of Symbolic Computation, 39(3–4): 451–463, March/April 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanderPut:2015:SOO

[vdPT15] Marius van der Put and Jaap Top. Stratified order one differential equations in positive characteristic. Journal of Symbolic Computation, 68 (part 2)(??):308–315, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114001059.

Veigneau:1997:SPS

[Vei97] Sébastien Veigneau. SP, a package for Schubert polynomials realized with the computer algebra system MAPLE. Journal of Symbolic Computation, 23(4):413–426 (or 413–425??), April 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vela:2000:ESG

[Vel00] Montserrat Vela. Explicit solutions of Galois embedding problems by means of generalized Clifford algebras. Journal of Symbolic Computation, 30(6):811–842, December 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0384;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0384/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0384/ref.

Verschelde:2000:TNM

[Ver00] Jan Verschelde. Toric Newton method for polynomial homotopies. Journal of Symbolic Computation, 29(4–5):777–793, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0296;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0296/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0296/ref.

vonzurGathen:1990:CNB

[vG90] Joachim von zur Gathen and Mark Giesbrecht. Constructing normal bases in finite fields. Journal of Symbolic Computation, 10(6):547– 570, December 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vira:1990:ASC

[VGT90] N. Vira, T. Gill, and E. Tunstel. Application of symbolic computation in robot pose error modeling. Journal of Symbolic Computation, 10(5):509– 524 (or 509–523??), November 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vo:2018:DER

[VGW18] N. Thieu Vo, Georg Grasegger, and Franz Winkler. Deciding the existence of rational general solutions for firstorder algebraic ODEs. Journal of Symbolic Computation, 87(??):127–139, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300639.

VonZurGathen:1998:FMP

[VH98] J. Von Zur Gathen and S. Hartlieb. Factoring modular polynomials. Journal of Symbolic Computation, 26(5): 583–606, November 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanHoeij:2013:GS

[vHKN13] Mark van Hoeij, Jürgen Klüners, and Andrew Novocin.

Generating subfields. Journal of Symbolic Computation, 52(??):17–34, May 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001277.

Vidunas:1999:DEO

[Vid99] Raimundas Vidunas. Differential equations of order two with one singular point. Journal of Symbolic Computation, 28(4–5): 495–520, October/November 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0312/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0312/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0312/production/ref.

Villard:1995:GSC

[Vil95] Gilles Villard. Generalized subresultants for computing the Smith normal form of polynomial matrices. Journal of Symbolic Computation, 20 (3):269–286, September 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Villard:2002:P

[Vil02] Gilles Villard. Preface. Journal of Symbolic Computation, 33(5):519, May 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Villard:2011:KDF

[Vil11]  Gilles Villard.  Kaltofen’s

division-free determinant algorithm differentiated for matrix adjoint computation. Journal of Symbolic Computation, 46(7):773–790, July 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717110001422.

Vill:2023:GST

[Vil23] Julian Vill. Gram spectrahedra of ternary quartics. Journal of Symbolic Computation, 116(??):263–283, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122001018.

Viry:1993:FMP

[Vir93] Guy Viry. Factorization of multivariate polynomials with coefficients in F_{p}. Journal of Symbolic Computation, 15(4): 371–392 (or 371–391??), April 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Viry:1999:EC

[Vir99]  Patrick Viry.  Elimina-

tion of conditions. Journal of Symbolic Computation, 28(3):381–401, September 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0288/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0288/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0288/production/ref.

Visser:2005:SSR

[Vis05] Eelco Visser. A survey of strategies in rule-based program transformation systems. Journal of Symbolic Computation, 40(1):831–873, July 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Verrill:2007:CTV

[VJ07] Helena Verrill and David Joyner. Computing with toric varieties. Journal of Symbolic Computation, 42(5):511–532, May 2007. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vaskouski:2016:SDC

[VK16] Maksim Vaskouski and Nikita Kondratyonok. Shortest division chains in unique factorization domains. Journal of Symbolic Computation, 77(??):175–188, November/December 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717116000110.

Vaskouski:2021:KVT

[VK21] Maksim Vaskouski and Nikita Kondratyonok. The Kronecker– Vahlen theorem fails in real quadratic norm-Euclidean fields. Journal of Symbolic Computation, 104(??): 134–141, May/June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717120300249.

vanDeursen:1993:OT

[vKT93] A. van Deursen, P. Klint, and F. Tip. Origin tracking. Journal of Symbolic Computation, 15(5–6):523–546 (or 523– 545??), May/June 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vaughan-Lee:1990:CL

[VL90] M. R. Vaughan-Lee. Collection from the left. Journal of Symbolic Computation, 9(5– 6):725–733, May/June 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational group theory, Part I.

Vaughan-Lee:1993:ACG

[VL93] Michael Vaughan-Lee. An algorithm for computing graded algebras. Journal of Symbolic Computation, 16(4):345– 354, October 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vrsek:2010:CAC

[VL10] Jan Vršek and Miroslav Lávička. On convolutions of algebraic curves. Journal of Symbolic Computation, 45(6): 657–676, June 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vrsek:2016:RIG

[VL16] Jan Vrsek and Miroslav Lávicka. Recognizing implicitly given rational canal surfaces. Journal of Symbolic Computation, 74(??): 367–377, May/June 2016. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717115000760.

Vivek:2014:CSC

[VM14] Srinivas Vivek and C. E. Veni Madhavan. Cubic Sieve Congruence of the Discrete Logarithm Problem, and fractional part sequences. Journal of Symbolic Computation, 64(??):22–34, August 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717113001703.

vonzurGathen:1987:FAC

[von87] Joachim von zur Gathen. Feasible arithmetic computations: Valiant’s hypothesis. Journal of Symbolic Computation, 4(2):137–172, October 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vonzurGathen:1990:FDPa

[von90a] Joachim von zur Gathen. Functional decomposition of polynomials: The tame case. Journal of Symbolic Computation, 9(3):281–300 (or 281– 299??), March 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vonzurGathen:1990:FDPb

[von90b] Joachim von zur Gathen. Functional decomposition of polynomials: The wild case. Journal of Symbolic Computation, 10(5):437–452, November 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vonzurGathen:1990:IFF

[von90c] Joachim von zur Gathen. Inversion in finite fields using logarithmic depth. Journal of Symbolic Computation, 9(2):175–184 (or 175– 183??), February 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

VonMohrenschildt:1998:NFF

[Von98] M. Von Mohrenschildt. Normal form for function rings of piecewise functions. Journal of Symbolic Computation, 26 (5):607–620, November 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vorontsov:1989:EAS

[Vor89] S. A. Vorontsov. An example of the application of symbolic computation in solid state physics. Journal of Symbolic Computation, 7(2):179– 182 (or 179–181??), February 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vorobjov:1992:CDC

[Vor92] Nikolaj N. Vorobjov, Jr. The complexity of deciding consistency of systems of polynomials in exponent inequalities. Journal of Symbolic Computation, 13(2):139–173 (or 139– 174??), February 1992. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vorobjov:1999:CCL

[Vor99] Nicolai Vorobjov. Complexity of computing the local dimension of a semialgebraic set. Journal of Symbolic Computation, 27(6):565–579, June 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0282/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0282/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0282/production/ref.

VanHoeij:1999:LSL

[VRUW99] Mark Van Hoeij, Jean-François Ragot, Felix Ulmer, and Jacques-Arthur Weil. Liouvillian solutions of linear differential equations of order three and higher. Journal of Symbolic Computation, 28(4–5): 589–609, October/November 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0316/production;
http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0316/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0316/production/ref.

Vakhidov:1997:NAA

[VV97] Akmal A. Vakhidov and Nickolay N. Vasiliev. A new approach for analytical computation of Hamiltonian of a satellite perturbed motion. Journal of Symbolic Computation, 24(6):705–710, December 1997. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Vavra:2018:PUG

[VV18] T. Vávra and F. Veneziano.

Pisot unit generators in number fields. Journal of Symbolic Computation, 89(??): 94–108, November/December 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717117301128.

Vaccon:2021:ATF

[VVY21] Tristan Vaccon, Thibaut Verron, and Kazuhiro Yokoyama. On affine tropical F5 algorithms. Journal of Symbolic Computation, 102(??): 132–152, January/February 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301208.

vonzurGathen:1995:HBD

[vW95] Joachim von zur Gathen and Jürgen Weiss. Homogeneous bivariate decompositions. Journal of Symbolic Computation, 19(5): 409–434, April 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Verdoolaege:2008:CRG

[VW08] Sven Verdoolaege and Kevin Woods. Counting with rational generating functions.

Journal of Symbolic Computation, 43(2):75–91, February 2008. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vonzurGathen:2013:LBD

[vzG13] Joachim von zur Gathen. Lower bounds for decomposable univariate wild polynomials. Journal of Symbolic Computation, 50(??): 409–430, March 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112001411.

vonzurGathen:2021:CIS

[vzGGZ21] Joachim von zur Gathen, Mark Giesbrecht, and Konstantin Ziegler. Counting invariant subspaces and decompositions of additive polynomials. Journal of Symbolic Computation, 105 (??):214–233, ???? 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771712030050X.

vonzurGathen:2022:SVD

[vzGM22] Joachim von zur Gathen and Guillermo Matera. Shifted varieties and discrete neighborhoods around varieties. Journal of Symbolic Computation, 109(??):31–49, March/April 2022. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000420.

vonzurGathen:2010:APG

[vzGMS10] Joachim von zur Gathen, Maurice Mignotte, and Igor E. Shparlinski. Approximate polynomial GCD: Small degree and small height perturbations. Journal of Symbolic Computation, 45(8):879– 886, August 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vonzurGathen:2001:FPF

[vzGP01] Joachim von zur Gathen and Daniel Panario. Factoring polynomials over finite fields: a survey. Journal of Symbolic Computation, 31 (1–2):3–17, January/February 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1002;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.1002/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.1002/ref.

Walther:2000:ACR

[Wal00] Uli Walther. Algorithmic computation of de Rham cohomology of complements of complex affine varieties. Journal of Symbolic Computation, 29(4–5):795–839, May 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0328;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.1999.0328/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        1999.0328/ref.

Waldmann:2002:CAMa

[Wal02a] Uwe Waldmann. Cancellative Abelian monoids and related structures in refutational theorem proving (Part I). Journal of Symbolic Computation, 33 (6):777–829, June 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Waldmann:2002:CAMb

[Wal02b] Uwe Waldmann. Cancellative Abelian monoids and related structures in refutational theorem proving (Part II). Journal of Symbolic Computation, 33(6):831–861, June 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Walther:2003:CSP

[Wal03] Uli Walther. Cohomology, stratifications and parametric Gröbner bases in characteristic zero. Journal of Symbolic Computation, 35(5):527–542, May 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Walther:2005:ASR

[Wal05] Uli Walther. Algorithmic stratification of RHom_{D}(M, N) for regular algebraic Dmodules on C^{n}. Journal of Symbolic Computation, 39 (3–4):493–499, March/April 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:1986:FSS

[Wan86] Paul S. Wang. FINGER:

a symbolic system for automatic generation of numerical programs in finite element analysis. Journal of Symbolic Computation, 2(3):305– 316, September 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:1991:MMC

[Wan91] Dong Ming Wang. Mechanical manipulation for a class of differential systems. Journal of Symbolic Computation, 12(2): 233–254, August 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:1993:EMP

[Wan93] Dong Ming Wang. An elimination method for polynomial systems. Journal of Symbolic Computation, 16(2): 83–114, August 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:1994:DII

[Wan94a] Dong Ming Wang. Differentiation and integration of indefinite summations with respect to indexed variables — some rules and applications. Journal of Symbolic Computation, 18(3):249–264 (or 249– 263??), September 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:1994:HO

[Wan94b] Xiang Dong Wang. The Hecke operators on S_{k}(Γ_{1}(N)). Journal of Symbolic Computation, 18(3):187–198, September 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:1996:PPO

[Wan96] Paul S. Wang. Parallel polynomial operations on SMPs: an overview. Journal of Symbolic Computation, 21(4/5/6): 397–410, April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Wang:1998:DPS

[Wan98] Dongming Wang. Decomposing polynomial systems into simple systems. Journal of Symbolic Computation, 25(3): 295–314, March 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:1999:PSC

[Wan99] Dongming Wang. Polynomial systems from certain differential equations. Journal of Symbolic Computation, 28 (1–2):303–315, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0278;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1998.0278/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0278/ref.

Wang:2000:CTS

[Wan00]  Dongming Wang.  Com-

puting triangular systems and regular systems. Journal of Symbolic Computation, 30(2):221–236, August 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0355;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0355/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0355/ref;   http://
www.idealibrary.com/links/
doi/10.1006/jsco.2000.0408;
 http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0408/pdf.

Wang:2004:SMI

[Wan04] Dongming Wang. A simple method for implicitizing rational curves and surfaces. Journal of Symbolic Computation, 38(1):899–914, July 2004. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wan:2006:ASI

[Wan06] Zhendong Wan. An algorithm to solve integer linear systems exactly using numerical methods. Journal of Symbolic Computation, 41(6):621–632, June 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:2018:DIQ

[Wan18] Jie Wang. Difference indices of quasi-prime difference algebraic systems. Journal of Symbolic Computation, 87 (??):1–13, July/August 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300536.

Wang:2023:TCD

[Wan23] Charles Wang. Towards cluster duality for Lagrangian and orthogonal Grassmannians. Journal of Symbolic Computation, 114(??):102– 121, January/February 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717122000384.

Wolf:1999:CAA

[WBM99] T. Wolf, A. Brand, and M. Mohammadzadeh. Computer algebra algorithms and routines for the computation of conservation laws and fixing of gauge in differential expressions. Journal of Symbolic Computation, 27(2):221– 238, February 1999. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:2012:IPS

[WC12] Xuhui Wang and Falai Chen.

Implicitization, parameterization and singularity computation of Steiner surfaces using moving surfaces. Journal of Symbolic Computation, 47(6):733–750, June 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002227.

Weber:1995:CCA

[Web95]  Andreas Weber.  On co-

herence in computer algebra. Journal of Symbolic Computation, 19(1/2/3):25– 38, January, February, March 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Weber:1996:PIA

[Web96] Kenneth Weber. Parallel implementation of the accelerated integer GCD algorithm. Journal of Symbolic Computation, 21(4/5/6):457–466, April, May & June 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Weispfenning:1988:CAL

[Wei88] Volker Weispfenning. The complexity of almost linear Diophantine problems. Journal of Symbolic Computation, 5(1–2):3–28 (or 3–27??), February/April 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Weispfenning:1990:CAL

[Wei90] V. Weispfenning. The complexity of almost linear Diophantine problems. Journal of Symbolic Computation, 10(5):395–404 (or 395– 403??), November 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Weispfenning:1992:CGB

[Wei92] Volker Weispfenning. Comprehensive Gröbner bases. Journal of Symbolic Computation, 14(1):1–30 (or 1–29??), July 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Weiermann:1994:CBS

[Wei94] Andreas Weiermann. Complexity bounds for some finite forms of Kruskal’s theorem. Journal of Symbolic Computation, 18(5):463–488, November 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Weispfenning:1997:SOQ

[Wei97] Volker Weispfenning. Simulation and optimization by quantifier elimination. Journal of Symbolic Computation, 24(2):189–208, August 1997. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Applications of quantifier elimination (Albuquerque, NM, 1995).

Weilert:2000:AGC

[Wei00] André Weilert. (1 + i)-ary GCD computation in Z[i] as an analogue to the binary GCD algorithm. Journal of Symbolic Computation, 30(5):605–617, October 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0422;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0422/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0422/ref.

Weispfenning:2003:CCG

[Wei03] Volker Weispfenning. Canonical comprehensive Gröbner bases. Journal of Symbolic Computation, 36(3–4): 669–683, September/October 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Weispfenning:2006:CGB

[Wei06] Volker Weispfenning. Comprehensive Gröbner bases and regular rings. Journal of Symbolic Computation, 41(3–4): 285–296, March/April 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Weimann:2013:FBP

[Wei13] Martin Weimann. Factoring bivariate polynomials using adjoints. Journal of Symbolic Computation, 58 (??):77–98, November 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113000837.

Weng:2006:CGT

[Wen06] Annegret Weng. Computing generators of the tame kernel of a global function field. Journal of Symbolic Computation, 41(9):964–979, September 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Werner:1998:SAO

[Wer98]  Andreas Werner.  A se-

mantic approach to ordersorted rewriting. Journal of Symbolic Computation, 25(4): 527–569, April 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wernhard:2012:PSD

[Wer12] Christoph Wernhard. Projection and scope-determined circumscription. Journal of Symbolic Computation, 47(9): 1089–1108, September 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002276.

Wirth:1994:CBA

[WG94] Claus-Peter Wirth and Bernhard Gramlich. A constructorbased approach to positive/ negative-conditional equational specifications. Journal of Symbolic Computation, 17 (1):51–90, January 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Conditional term rewriting systems (Pont-à-Mousson, 1992).

Wang:2018:STS

[WG18] Haohao Wang and Ron Goldman. Syzygies for translational surfaces. Journal of Symbolic Computation, 89(??):73–93, November/December 2018. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117301116.

White:1991:MCF

[Whi91a] Neil L. White. Multilinear Cayley factorization. Journal of Symbolic Computation, 11(5–6):421–438, May/ June 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Invarianttheoretic algorithms in geometry (Minneapolis, MN, 1987).

Whiteley:1991:ICA

[Whi91b]  Walter Whiteley.  Invari-

ant computations for analytic projective geometry. Journal of Symbolic Computation, 11(5–6):549–578, May/ June 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Invarianttheoretic algorithms in geometry (Minneapolis, MN, 1987).

Wibmer:2007:GBF

[Wib07] Michael Wibmer. Gröbner bases for families of affine or projective schemes. Journal of Symbolic Computation, 42(8): 803–834, August 2007. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Widiger:2001:SWP

[Wid01]  Alfred Widiger.  Solving

the word problem for two classes of non-associative rings by rewriting. Journal of Symbolic Computation, 32 (3):291–301, September 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0404;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.0404/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.0404/ref.

Wilson:1990:MCC

[Wil90] Robert A. Wilson. The 2- and 3-modular characters of J_{3}, its covering group and automorphism group. Journal of Symbolic Computation, 10(6):647– 657 (or 647–656??), December 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wilson:1993:BTC

[Wil93] Robert A. Wilson. The Brauer tree for J_{3} in characteristic 17. Journal of Symbolic Computation, 15(3):325–330, March 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wilf:1995:CAD

[Wil95]  Herbert S. Wilf.  The

computer-aided discovery of a theorem about Young tableaux. Journal of Symbolic Computation, 20(5–6):731– 736 (or 731–735??), November/December 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Winkler:1988:AAC

[Win88] Franz Winkler. A p-adic approach to the computation of Gröbner bases. Journal of Symbolic Computation, 6(2–3):287–304, October/December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Computational aspects of commutative algebra.

Windsteiger:2006:APZ

[Win06] Wolfgang Windsteiger. An automated prover for Zermelo– Fraenkel set theory in Theorema. Journal of Symbolic Computation, 41(3–4): 435–470, March/April 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Winterhof:2014:GCM

[Win14] Arne Winterhof. Generalizations of complete mappings of finite fields and some applications. Journal of Symbolic Computation, 64(??):42–52, August 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717113001727.

Wirth:2009:SCC

[Wir09] Claus-Peter Wirth. Shallow confluence of conditional term rewriting systems. Journal of Symbolic Computation, 44(1): 60–98, January 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wirth:2012:NPS

[Wir12] Claus-Peter Wirth. lim +, δ^{+}, and non-permutability of β-steps. Journal of Symbolic Computation, 47(9): 1109–1135, September 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111002288.

Wolf:1991:GCP

[WK91] Kurt Bernardo Wolf and Guillermo Krötzsch. Groupclassified polynomials of phase space in higher-order aberration expansions. Journal of Symbolic Computation, 12(6):673–694 (or 673– 693??), December 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wouodjie:2020:SHT

[WK20] Merlin Mouafo Wouodjié and Wolfram Koepf. On the solutions of holonomic thirdorder linear irreducible differential equations in terms of hypergeometric functions.

Journal of Symbolic Computation, 101(??):202–241, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711930094X.

Weber:1994:APP

[WKA94] G. Weber, L. Knipping, and H. Alt. An application of point pattern matching in astronautics. Journal of Symbolic Computation, 17(4): 321–340, April 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Watowich:1986:SAO

[WKB86] S. J. Watowich, J. L. Krause, and R. S. Berry. Stability analysis of an optimally controlled light-driven engine. Journal of Symbolic Computation, 2(1):103–108, March 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wada:2006:GBH

[WO06] Takashi Wada and Hidefumi Ohsugi. Gröbner bases of Hilbert ideals of alternating groups. Journal of Symbolic Computation, 41(8):905– 908, August 2006. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wolf:2000:SIE

[Wol00a] Thomas Wolf. The symbolic integration of exact PDEs. Journal of Symbolic Computation, 30(5):619–629, October 1, 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.1069;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2000.1069/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2000.1069/ref.

Wolfram:2000:FGS

[Wol00b]  D. A. Wolfram.  A for-

mula for the general solution of a constant-coefficient difference equation. Journal of Symbolic Computation, 29(1):79–82, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0350;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1999.0350/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0350/ref.

Wolf:2002:SRP

[Wol02] Thomas Wolf. Size reduction and partial decoupling of systems of equations. Journal of Symbolic Computation, 33(3):367–383, March 1, 2002. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0509;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jsco.2001.0509/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jsco.
                                        2001.0509/ref.

Wolf:2003:ISL

[Wol03] Thomas Wolf. The integration of systems of linear PDEs using conservation laws of syzygies. Journal of Symbolic Computation, 35(5):499–526, May 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Worfolk:1994:ZEV

[Wor94] Patrick A. Worfolk. Zeros of equivariant vector fields: Algorithms for an invariant approach. Journal of Symbolic Computation, 17(6):487–512 (or 487–511??), June 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wrathall:1988:WPF

[Wra88] C. Wrathall. The word problem for free partially commutative groups. Journal of Symbolic Computation, 6(1): 99–104, August 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wu:2009:IRB

[WRI09] Wenyuan Wu, Greg Reid, and Silvana Ilie. Implicit Riquier Bases for PDAE and their semi-discretizations. Journal of Symbolic Computation, 44 (7):923–941, July 2009. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Watt:1998:SIS

[WS98] S. M. Watt and H. J. Stetter. Special issue on symbolic numeric algebra for polynomials: Foreword of the Guest Editors. Journal of Symbolic Computation, 26(6):649– 652, December 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wietecha:2009:PMA

[WS09] T. Wietecha and K. Sokalski.

Plus–minus algorithm — a method for derivation of the Bäcklund transformations. Journal of Symbolic Computation, 44(10):1511–1528, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:2023:LBR

[WS23] Kexin Wang and Anna Seigal. Lower bounds on the rank and symmetric rank of real tensors. Journal of Symbolic Computation, 118(??): 69–92, September/October 2023. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000044.

Wursthorn:1993:IMG

[Wur93] Martin Wursthorn. Isomorphisms of modular group algebras: An algorithm and its application to groups of order 2^{6}. Journal of Symbolic Computation, 15(2):211–228 (or 211– 227??), February 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Welzl:1994:SRB

[WW94] Emo Welzl and Barbara Wolfers. Surface reconstruction between simple polygons via angle criteria. Journal of Symbolic Computation, 17 (4):351–369, April 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Algorithms: implementation, libraries and use (Schloss Dagstuhl, 1993).

Wang:2023:EGA

[WWWX23] Dingkang Wang, Hesong Wang, Jingjing Wei, and Fanghui Xiao. An extended GCRD algorithm for parametric univariate polynomial matrices and application to parametric Smith form. Journal of Symbolic Computation, 115 (??):248–265, March/April 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000645.

Wilkening:2011:LCS

[WY11] Jon Wilkening and Jia Yu.

A local construction of the Smith normal form of a matrix polynomial. Journal of Symbolic Computation, 46(1):1–22, January 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717110001136.

Woo:2020:TSP

[WY20] Alexander Woo and Alexander Yong. Tropicalization, symmetric polynomials, and complexity. Journal of Symbolic Computation, 99(??): 242–249, July/August 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119300550.

Wang:1993:ACI

[WyW93] Rongdong Wang and Peng yung Woo. Automatic computation of z- & inverse z-transformations by Maple. Journal of Symbolic Computation, 15(3):349–363, March 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wu:2012:DSS

[WZ12] Xiaoli Wu and Lihong Zhi.

Determining singular solutions of polynomial systems via symbolic-numeric reduction to geometric involutive forms. Journal of Symbolic Computation, 47 (3):227–238, March 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717111001672.

Wang:2023:RRA

[WZ23] Rong-Hua Wang and Michael X. X. Zhong. q-rational reduction and q-analogues of series for π. Journal of Symbolic Computation, 116 (??):58–71, May/June 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717122000876.

Xu:2013:STA

[XL13] Ming Xu and Zhi-Bin Li. Symbolic termination analysis of solvable loops. Journal of Symbolic Computation, 50(??):28–49, March 2013. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717112000958.

Xu:2015:QEC

[XLY15] Ming Xu, Zhi-Bin Li, and Lu Yang. Quantifier elimination for a class of exponential polynomial formulas. Journal of Symbolic Computation, 68 (part 1)(??):146–168, May/ June 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717114000832.

Xu:2020:EFE

[XW20] Ce Xu and Weiping Wang.

Explicit formulas of Euler sums via multiple zeta values. Journal of Symbolic Computation, 101(??):109–127, November/December 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074771711930063X.

Xu:2023:SNR

[XWL23] Juan Xu, Dongming Wang, and Dong Lu. Squarefree normal representation of zeros of zero-dimensional polynomial systems. Journal of Symbolic Computation, 122(??): Article 102273, ???? 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000871.

Xin:2023:MDL

[XXZZ23] Guoce Xin, Xinyu Xu, Chen Zhang, and Yueming Zhong. On magic distinct labellings of simple graphs. Journal of Symbolic Computation, 119(??):22–37, November/December 2023. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000135.

Xia:2002:AIR

[XY02] Bican Xia and Lu Yang. An algorithm for isolating the real solutions of semialgebraic systems. Journal of Symbolic Computation, 34(5): 461–477, November 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Xia:2010:TLP

[XZ10] Bican Xia and Zhihai Zhang. Termination of linear programs with nonlinear constraints. Journal of Symbolic Computation, 45(11):1234– 1249, November 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yahl:2023:CGG

[Yah23]  Thomas Yahl.  Comput-

ing Galois groups of Fano problems. Journal of Symbolic Computation, 119(??): 81–89, November/December 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717123000123.

Yamamoto:1994:SBZ

[Yam94] Osami Yamamoto. On some bounds for zeros of normbounded polynomials. Journal of Symbolic Computation, 18(5):403–428 (or 403– 427??), November 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yan:1998:GDS

[Yan98] Thomas Yan. The geobucket data structure for polynomials. Journal of Symbolic Computation, 25(3):285–294 (or 285–293??), March 1998. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yang:1999:RAD

[Yan99] Lu Yang. Recent advances on determining the number of real roots of parametric polynomials. Journal of Symbolic Computation, 28 (1–2):225–242, July/August 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0274;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.1998.0274/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1998.0274/ref.

Yap:1990:STG

[Yap90] Chee-Keng K. Yap. Symbolic treatment of geometric degeneracies. Journal of Symbolic Computation, 10(3–4): 349–370, September/October 1990. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-

855X (electronic).

Yap:1991:NLB

[Yap91] Chee-K. K. Yap. A new lower bound construction for the word problem for commutative Thue systems. Journal of Symbolic Computation, 12(1):1–28 (or 1–27??), July 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ye:2017:SDP

[Ye17] Liangjie Ye. A symbolic decision procedure for relations arising among Taylor coefficients of classical Jacobi theta functions. Journal of Symbolic Computation, 82(??): 134–163, September/October 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117300135.

Ye:2018:EFB

[Ye18] Liangjie Ye. Elliptic function based algorithms to prove Jacobi theta function relations. Journal of Symbolic Computation, 89(??):171–193, November/December 2018. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717117301189.

Yelick:1987:UCC

[Yel87] Katherine A. Yelick. Unification in combinations of collapse-free regular theories. Journal of Symbolic Computation, 3(1–2):153–181, February/April 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Rewriting techniques and applications (Dijon, 1985).

Yesil:2021:GKZ

[Yes21] Mehmet Yesil. A generalization of the Katzman–Zhang algorithm. Journal of Symbolic Computation, 103(??): 213–233, March/April 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717119301531.

Yamasaki:1994:TLP

[YI94] Susumu Yamasaki and Kentaro Iida. Transformation of logic programs to FP programs based on dataflows. Journal of Symbolic Computation, 18(2):157–182, August 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yokoyama:1989:CPE

[YNT89] Kazuhiro Yokoyama, Masayuki Noro, and Taku Takeshima. Computing primitive elements of extension fields. Journal of Symbolic Computation, 8(6):

553–580, December 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yokoyama:1992:SSA

[YNT92] Kazuhiro Yokoyama, Masayuki Noro, and Taku Takeshima. Solutions of systems of algebraic equations and linear maps on residue class rings. Journal of Symbolic Computation, 14(4):399–417, October 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yokoyama:1994:MMA

[YNT94] Kazuhiro Yokoyama, Masayuki Noro, and Taku Takeshima. Multi-modular approach to polynomial-time factorization of bivariate integral polynomials. Journal of Symbolic Computation, 17(6):545–563, June 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yokoyama:2017:SIC

[Yok17] Kazuhiro Yokoyama. Special issue on the conference ISSAC 2015: Symbolic computation and computer algebra. Journal of Symbolic Computation, 83(??):1– 3, November/December 2017. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717116301225.

You:1989:EON

[You89] Jia-Huai You. Enumerating outer narrowing derivations for constructor-based term rewriting systems. Journal of Symbolic Computation, 7(3–4):319–342 (or 319– 341??), March/April 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yamartino:1991:ACA

[YP91] Robert J. Yamartino and Richard Pavelle. An application of computer algebra to a problem in stratified fluid flow. Journal of Symbolic Computation, 12(6):669– 672, December 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Young:1987:GGU

[YW87] D. A. Young and P. S. Wang.

GI/S: a graphical user interface for symbolic computation systems. Journal of Symbolic Computation, 4(3):365– 380, December 1987. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yi:1995:ATT

[YX95] Bo Yi and Jia Fu Xu. Analogical type theory. Journal of Symbolic Computation, 19(1/2/3):3–24, January, February, March 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Ying:1999:CMD

[YXL99] J. Q. Ying, L. Xu, and Z. Lin. A computational method for determining strong stabilizability of n-D systems. Journal of Symbolic Computation, 27(5):479–500, May 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yu:2003:MPT

[YY03] Pei Yu and Yuan Yuan. A matching pursuit technique for computing the simplest normal forms of vector fields. Journal of Symbolic Computation, 35(5):591–615, May 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Yu:2012:DPH

[YYZ12] Zhiheng Yu, Lu Yang, and Weinian Zhang. Discussion on polynomials having polynomial iterative roots. Journal of Symbolic Computation, 47(10):1154–1162, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111002318.

Yuan:2021:NBE

[YZ21] Chun-Ming Yuan and Zhi-Yong Zhang. New bounds and an efficient algorithm for sparse difference resultants. Journal of Symbolic Computation, 107(??):279–298, November/December 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717121000262.

Zantema:1994:TTR

[Zan94] H. Zantema. Termination of term rewriting: interpretation and type elimination. Journal of Symbolic Computation, 17(1):23–50, January 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Conditional term rewriting systems (Pont-à-Mousson, 1992).

Zantema:1995:TTT

[Zan95] Hans Zantema. Total termination of term rewriting is undecidable. Journal of Symbolic Computation, 20(1): 43–60, July 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhang:1996:PDP

[ZBH96] Hantao Zhang, Maria Paola Bonacina, and Jieh Hsiang. PSATO: a distributed propositional prover and its application to quasigroup problems. Journal of Symbolic Computation, 21(4/5/6):543– 560, April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Zhao:2002:MPD

[ZD02] Xuan Zhao and Haibao Duan. A Mathematica program for the degrees of certain Schubert varieties. Journal of Symbolic Computation, 33(4): 507–517, April 2002. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zeilberger:1991:MCT

[Zei91]  Doron Zeilberger.  The

method of creative telescoping. Journal of Symbolic Computation, 11(3):195– 204, March 1991. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zeilberger:1995:TRH

[Zei95]  Doron Zeilberger.  Three

recitations on holonomic systems and hypergeometric series. Journal of Symbolic Computation, 20(5–6):699– 724, November/December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Zeng:2006:DTR

[Zen06] Guangxing Zeng. Determination of the tangents for a real plane algebraic curve. Journal of Symbolic Computation, 41(8):863–886, August 2006. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhang:2009:DOD

[ZG09] Mingbo Zhang and Xiao-Shan Gao. Decomposition of ordinary difference polynomials. Journal of Symbolic Computation, 44(10):1394–1409, October 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhang:2023:CFC

[ZGG23] Tianjie Zhang, Xing Gao, and Li Guo. Construction of free commutative Reynolds algebras by Gröbner–Shirshov bases. Journal of Symbolic Computation, 119(??): 64–80, November/December 2023. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717123000196.

Zhang:1990:APR

[Zha90] Hantao Zhang. Automated proof of ring commutativity problems by algebraic methods. Journal of Symbolic Computation, 9(4):423–428 (or 423–427??), April 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhang:1992:SPF

[Zha92] Louxin Zhang. Some properties of finite special stringrewriting systems. Journal of Symbolic Computation, 14(4): 359–370 (or 359–369??), October 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zharkov:1993:CCI

[Zha93] A. Yu. Zharkov. Computer classification of the integrable coupled Kdv-like systems with unit main matrix. Journal of Symbolic Computation, 15(1): 85–90, January 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhang:1994:NMB

[Zha94] Hantao Zhang. A new method for the Boolean ring based theorem proving. Journal of Symbolic Computation, 17(2):189–212 (or 189– 211??), February 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zharkov:1995:CFS

[Zha95] Alexey Zharkov. Coefficient fields of solutions in Kovacic’s algorithm. Journal of Symbolic Computation, 19(5): 403–408, April 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhang:1996:SCC

[Zha96] Jun Zhang. Symbolic computation on complex polynomial solution of differential equations. Journal of Symbolic Computation, 22(3):345– 354, September 1996. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhang:2003:NEA

[Zha03] Bao-Yin Zhang. A new elementary algorithm for proving q-hypergeometric identities. Journal of Symbolic Computation, 35(3):293–303, March 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zheng:2021:FAC

[Zhe21] Tao Zheng. A fast algorithm for computing multiplicative relations between the roots of a generic polynomial. Journal of Symbolic Computation, 104(??):381–401, May/ June 2021. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0747717120300821.

Ziegler:2016:TDC

[Zie16] Konstantin Ziegler. Tame decompositions and collisions. Journal of Symbolic Computation, 75(??): 244–268, July/August 2016.

CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717115001169.

Zippel:1985:SEI

[Zip85]  Richard Zippel.  Simplifi-

cation of expressions involving radicals. Journal of Symbolic Computation, 1(2): 189–210, June 1985. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zippel:1990:IPT

[Zip90] Richard Zippel. Interpolating polynomials from their values. Journal of Symbolic Computation, 9(3):375–403, March 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhou:2012:EAO

[ZL12] Wei Zhou and George Labahn. Efficient algorithms for order basis computation. Journal of Symbolic Computation, 47(7):793–819, July 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074771711100201X.

Zheng:2001:DAC

[ZS01] Jianmin Zheng and Thomas W. Sederberg. A direct approach to computing the μbasis of planar rational curves.

Journal of Symbolic Computation, 31(5):619–629, May 1, 2001. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0437;   http://www.
idealibrary.com/links/doi/
10.1006/jsco.2001.0437/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
2001.0437/ref.

Zhang:1993:EAP

[ZSY93] C. N. Zhang, B. Shirazi, and D. Y. Y. Yun. An efficient algorithm and parallel implementations for binary and residue number systems. Journal of Symbolic Computation, 15(4):451–462, April 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhou:2008:CDD

[ZW08] Meng Zhou and Franz Winkler. Computing differencedifferential dimension polynomials by relative Gröbner bases in difference-differential modules. Journal of Symbolic Computation, 43(10): 726–745, October 2008. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Zhao:2011:SFC

[ZWH11] Ting Zhao, Dongming Wang, and Hoon Hong. Solution formulas for cubic equations without or with constraints. Journal of Symbolic Computation, 46(8):904–918, August 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717111000344.

Zankl:2015:BPP

[ZWM15] Harald Zankl, Sarah Winkler, and Aart Middeldorp. Beyond polynomials and Peano arithmetic — automation of elementary and ordinal interpretations. Journal of Symbolic Computation, 69(??): 129–158, July/August 2015. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717114001011.

Zeng:2020:DLB

[ZX20] Xiaoning Zeng and Shuijing Xiao. Determining the limits of bivariate rational functions by Sturm’s theorem. Journal of Symbolic Computation, 96 (??):1–21, January/February 2020. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0747717119300215.